A Complete Bibliography of Publications in Nordisk Tidskrift for Informationsbehandling, BIT, and BIT

Numerical Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

07 October 2025 Version 3.64

Title word cross-reference

#24873 [1089].

(3) [2450]. (A − λB) [1414]. 0/1 [1242]. 1 [3037, 2711]. 1 − 2 [1097]. 1.0 [3043]. 10^{12} [838]. 11 [1311]. 2 [2144, 1953, 2291, 2162, 3047, 886, 2551, 957, 2007, 2552, 1832, 949, 3024, 3219, 2194]. 2, 3 [1490]. 2^{n} − 1 [320]. 2^{p} − 1 [100]. 2m − 4 [1105]. 3 [2119, 1953, 2531, 1351, 2551, 1292, 1793, 949, 1356, 2711, 2227, 570]. 3, 000, 000, 000 [575, 637]. 33 [1437]. 4 [2465, 1490, 1737]. 5 [1670]. 6 [3306]. 682 [1270, 1338]. 7 [1670]. 8 [2937]. ^{4} [591]. A [3105, 328, 469, 655, 896, 524, 873, 455, 779, 946, 2944, 297, 1752, 670, 2582, 1409, 1987, 915, 808, 761, 916, 2071, 2198, 1449, 780, 959, 1105, 1021, 497, 2589]. A(α) [896, 2594, 333]. A^{∗} [2013]. A^{∗}Ax = b [2369]. A_{0} [1640, 566, 947, 1580, 1460]. A_{n} = a^{2^{n}} + 1 [334]. α [824, 1580]. AN [1622]. A(ϑ) [3439]. AX − XB = C [2195, 2006]. AXD − BXC = E [1101]. B [2187, 1575, 1267, 3734, 1409, 1489, 1991, 1191, 979, 1819, 1597, 1823, 1773]. β [3639, 824]. BN [1181]. BS [1773]. BSI [1446]. C^{0} [2906]. C^{1} [3202]. C^{2} [3108, 2422, 3000, 2036]. C^{r} [3643]. χ^{2} [30, 31]. Cl_{n}(θ), (n ≥ 2) [2929]. cos [228]. D [1390, 1754, 314]. δ^{2} [1142]. ΔH [2585]. E [1475]. e^{−z} [775]. ℓ_{1} [3201]. ℓ_{p} [3308]. ℓ_{q} [3308]. F [3112]. G [1476, 2755, 3208, 3306, 3107, 946, 762]. g(|f(x)|) [2171]. G^{1} [2813]. G^{2} [2221, 2530]. γ [1061]. GF(2) [1490]. GF(2^{n}) [1463]. GF(2^{p}) [1177]. H [3180, 3409, 1353, 2612, 3223, 3178]. H(curl, Ω) [3701]. H^{1} [3487, 3510, 3169]. h^{2} [1540]. HC^{2} [3200]. I [1898, 959]. I_{MN} [2279]. ILU [2400]. iR [1391]. k [581, 1921, 1955, 3178, 2029, 1875]. K < 10 [100]. L [2354, 2495, 1909, 2165, 906, 3471, 3724, 3412]. L[2] [2744]. L[p] [2744]. L^{1} [2677, 2153]. L^{∞} [3156]. L^{∞}(L^{2}) [3374]. L^{p} [3594]. L_{1} [1831, 1932]. L_{2} [3223]. L_{∞} [2908, 2694, 2273]. L_{p} [584, 619, 1302]. LALR(1) [660]. LALR(k) [1173]. λ [1940, 1210, 1197]. LDL^{H} [1903]. LDU [3132]. LL(1) [789]. LR [3600]. LR(k) [1173, 1245]. LU [1903, 2784, 1701, 2374]. M [3392, 1785, 2308, 2202, 2540, 2612, 2127, 2400, 1458, 3197, 2226]. M/G/1 [2296]. F_{2} [3135]. R^{3} [3307, 3367]. L_{∈} [3135]. O(N) [3176]. O(N^{∈}) [3176]. MK [2091]. N [1501, 1980, 3176, 2535, 1064, 320, 1362]. N = h · 2^{n} − 1 [2092]. N = k · 2^{m}3^{n} + 1 [1976]. N = k · 3^{n} + 1 [1976]. n^{1} [1463]. n × n [1010]. O(h^{6}) [1663]. O(log n) [1294]. O(n^{2}) [2027]. O(Nlog N) [1501]. P [1262, 1398, 1371, 1162, 1252, 1016, 1368, 1626, 2390, 932, 1063, 1753, 2131, 2312, 707, 3639, 3339]. p(n, m) [806]. p < 15000 [100]. P [α, β] [1262]. P_{1} [2689, 164]. π [1335, 1511]. Q [3279, 2437, 719, 2873, 1874]. q = 2Kp + 1 [100]. qd [2096]. QR [3120, 1314, 1913, 642, 218, 1901, 2156, 1907]. R [2895, 1530, 2542, 3023, 1227, 3139]. R^{3} [1150]. R^{d} [2530]. ℜ [3590]. S [2036, 610]. sin [228]. sin(θ) [3498]. SR [3129]. SU(n) [410]. T [3597]. τ [371]. θ [3747, 2347, 2388]. φ [3735]. φ [3442]. W [2550, 2711, 2290, 2681, 2522]. W (m, m − 1)2(0, 1)W_{2}^{(m,m−1)}(0, 1) [3091]. X [1329]. XA + AY = F [844]. XY = A [1302]. Y [1329]. y′′ = f(x, y) [628, 1505]. y′′ = f(z, y) [454]. y′′ = f(x, y) [1195, 1081]. Z [3132, 1574]. z′ = f(t, x) [448]. ζ(3) [2459].

-Acceptability [2582, 1640]. -adic [932]. -algorithm [2096, 1476]. -approximations [584]. -Arnoldi-type [3279]. -bases [3409]. -BEM [3178]. -Bernstein [2873]. -Bézier [3223]. -Colour [570]. -conditioned [2522]. -consistency [1773]. -continuity [2036]. -contractivity [1460]. -convergence [1575, 1489, 3594, 1819, 1597]. -convex [2036]. -cube [1737]. -curve [2354, 2165]. -curves [2495]. -cyclic [1753, 2131]. -Decomposition [642, 1701]. -dimension [1227]. -dimensional [1921, 1356, 1875]. -elliptic [3180]. -error [3487]. -estimates [2677]. -estimation [2308, 2202]. -estimator [2540]. -factors [1785]. -Formula [1142]. -free [1210]. -functions [3442, 1458]. -Galerkin [3169]. -Grammars [610]. -Hermitian [3338]. -Hessenberg [3197]. -Interchange [1242]. -languages [906]. -Leja [3590]. -linear [3023]. -Matrices [2585, 2390, 3132, 1353, 2612, 2127]. -matrix [1197, 2400]. -method [218]. -methods [2347, 2711, 619, 2290, 2681]. -node [2937]. -Norm [2744, 3223, 3112]. -obstacle [3176]. -order [3735, 2895, 1530, 2437, 2542]. -parametric [1063]. -point [957]. -problem [1414]. -Projection [2744]. -quadratic [2906]. -Riccati [3597]. -samplable [1909]. -separation [1874]. -series [3734]. -simplex [1574]. -solution [3339, 1302]. -splines [3471]. -splitting [3392]. -stability [2312, 333, 3439, 896, 1267, 1622, 2755, 946, 1446, 947, 1409, 1390, 1987, 1773, 915, 808, 761, 916, 3724, 979, 1181, 780, 959, 1021]. -Stable [469, 524, 455, 566, 297, 670, 2594, 497, 3105, 328, 655, 2187, 1262, 873, 1398, 1371, 1162, 779, 1252, 2944, 1752, 1016, 2071, 2198, 1475, 3412, 1898, 1991, 762, 1449, 1580, 1368, 1626, 1105]. -Step [581, 719, 2531, 1351]. -StepSCAN [1980].

-subspace [3747]. -suitability [1754]. -symmetric [3338]. -symplectic [3208, 3306]. -symplecticity [3107]. -terms [1940]. -th [707]. -times [2226]. -Transformation [2550]. -trees [1953, 1391, 1823, 1191]. -type [1580]. -weak [3169]. -weighted [2091].

/Ferromagnetic [2642].

1 [27, 58, 81, 28]. 10 [394]. 11 [465]. 12 [562, 543]. 13 [637]. 16 [890]. 17 [1089]. 18 [1002]. 19 [3485]. 1961 [20]. 1962 [56]. 1988 [1703]. 1993 [2064]. 1D [3538].

2 [78, 3130]. 20 [1320]. 2000-2005 [2714]. 2008 [2843]. 21 [1202]. 22 [1308, 1338]. 24 [1398]. 27 [1629, 1630, 1651]. 2D [3738]. 2nd [2867].

3 [3092]. 3003 [181]. 31 [1964]. 35 [2316].

4 [188, 120, 498, 499, 135, 267, 3104]. 40-45 [2714]. 4000 [245, 256]. 44 [2655]. 4th [273].

500 [61]. 54-3 [3143]. 58 [1089].

6 [517]. 60 [283, 6, 13, 215]. 60th [1703]. 64 [40]. 66 [274]. 68 [232, 294, 273, 396].

75th [2064].

80b [1002]. 82e [1320]. 84b [1338]. 88k [1651].

91m [1964].

a-posteriori [2845]. A-priori [2845]. A-Stability [2389]. A-stable [3544]. A.D.I. [886]. AAA [3632]. Abel [3551, 702, 2219, 2248, 2322, 607, 3507]. Abelian [2996]. ABS [1701]. ABS-method [1701]. abscissae [2516]. absolute [1195, 1166]. Absorbing [2720]. abstract [1486, 881, 1572]. abstraction [2210]. Abstractions [1278]. Accelerated [988, 3151, 3250, 1017, 3381, 3184, 2796, 2240, 3248, 1107, 1214, 3560, 1545]. Acceleration [312, 2453, 693, 144, 194, 155, 254, 1349, 3474, 1408, 1977, 1931, 1834]. Acceptability [2582, 1640]. Access [398, 3753, 1259, 690, 1998, 1705, 1391, 2026, 2000]. Accessibility [358]. Accessing [1657]. accretive [2226]. accumulated [773]. Accuracy [27, 3754, 2321, 12, 28, 2653, 643, 3340, 215, 608, 966, 1528, 2119, 912, 3663, 2420, 3357, 2107, 1249]. Accurate [3132, 2562, 3687, 983, 875, 3408, 3150, 706, 3520, 2156, 3489, 2187, 3184, 3496, 3526, 3596, 2066, 3177, 2970, 1488, 2460, 2122, 3506, 3320]. accurately [1846]. Achieving [2847]. Ackermann [703, 417, 765]. Acknowledgements [349, 408, 466, 542, 614, 681, 752, 821, 892, 963, 1046, 1134, 1206, 1274, 1341, 1498, 1569, 1634, 1723, 1800, 1870, 1934, 1994, 2050, 2100, 2142, 2200, 2262, 2317, 2363, 2414, 2491, 2544]. Acoustic [2703, 3184, 3176, 3181]. acquired [992]. Action [332, 3121, 3383]. activities [1235]. Activity [41, 1555]. acyclic [1361]. ADA [1688]. Adams [3016, 2393, 3199, 2511, 3006, 3618]. Adams-type [2393, 3006]. Adaptation [1759, 1482]. Adapted [2326, 3070, 3325]. adapting [3673, 1988]. Adaptive [1782, 3133, 3046, 2320, 2566, 2370, 3180, 1484, 3348, 2750, 2765, 2939, 3533, 2517, 3723, 2791, 3614, 3493, 3295, 1369, 3580, 1751, 3462, 3500, 3708, 1096, 3620, 3644, 2384, 1303, 2244, 2021, 2539, 3643, 2973, 1318, 2502, 1250, 3610, 2854]. Adaptive-order [3133]. adaptivity [2500]. ADB [10]. Addendum [1320]. Addition [147]. Additive [2689, 3268, 2671, 3654, 2945, 3474, 2944, 2607, 2896, 1987, 3033, 3020, 3089]. Address [1434]. Addresses [1093, 1113, 1136, 1156, 1170, 1189, 1208, 1224, 1240]. addressing [2038, 1672]. ADI [3196, 3515]. Adiabatic [2644, 2676]. adic [932]. adjacent [2164]. Adjoint [3503, 530, 2436, 990, 3389, 2888]. Adjoint-based [3503]. Administration [486, 745, 134]. Administrative [152, 145, 42]. Admissible [3682, 1678]. admitting [1815]. ADP [134, 180, 145]. advanced [1031]. Advances [3241]. advancing [2113]. advection [3614, 3619, 3063, 2814, 3366]. advection-diffusion-reaction [3614]. af [15, 25]. Affine [3447, 2934, 1218, 3339]. affine-linear [3339]. after [309]. aggregation [3696]. Agreements [3484]. ahead [2111, 2164]. aid [39]. aided [2996]. Air [440, 1424]. Air-Pollution [440]. Airy [2498]. Ait [3670, 2983]. Aitken [1349, 1142, 1119, 155]. aLFA [3463]. Algebra [2713, 446, 154, 2614, 1094, 3241, 2756, 2998]. Algebraic [377, 1459, 693, 2441, 670, 329, 342, 2552, 2698, 1898, 2700, 2711, 677, 240, 3493, 2119, 2291, 2416, 1436, 2602, 3597, 1622, 2519, 2307, 2162, 2796, 1283, 940, 1385, 1439, 947, 2757, 1725, 3254, 1699, 1732, 3619, 2084, 2007, 3037, 2542, 1918, 2348, 3311, 2787, 2515, 3434, 2227, 3260, 2338, 847, 1835, 3328, 2259, 3540]. Algebraically [2884, 945, 1575, 930, 1087]. ALGOL [124, 267, 38, 135, 340, 9, 36, 396, 283, 153, 359, 6, 13, 170, 128, 87, 91, 129, 119, 161, 70, 351, 71, 62, 88, 141, 188, 260, 78, 389, 63, 177, 38, 58, 517, 89, 241, 112, 120, 207, 336, 130, 498, 499, 95, 121, 314, 218, 200, 250, 156, 215, 149]. ALGOL-GENIUS [128]. Algorithm [340, 410, 1049, 2016, 488, 705, 599, 1950, 536, 182, 235, 329, 342, 354, 1618, 745, 642, 550, 321, 492, 552, 2750, 725, 357, 229, 370, 1059, 1767, 496, 514, 416, 557, 1067, 3535, 217, 174, 534, 310, 610, 403, 1225, 2353, 2600, 1452, 1764, 1476, 2145, 1921, 2943, 1704, 3225, 1602, 1574, 2146, 1010, 3553, 1512, 1314, 1697, 2095, 78, 3197, 816, 1077, 1119, 2175, 1429, 2893, 2235, 2860, 3741, 1050, 1501, 2605, 1350, 1669, 2796, 2024, 899, 3185, 1643, 2096, 1402, 3030, 2236, 3529, 825, 885, 2533, 2652, 3212, 3732, 1645, 2283, 2474, 1468, 1013, 1838, 3176]. algorithm [2222, 1361, 1706, 1407, 2774, 2130, 1378, 1726, 3282, 2481, 1096, 1956, 1658, 1745, 1822, 2760, 2011, 1292, 3752, 1832, 1276, 1518, 1896, 1700, 1202, 1146, 1128, 1088, 3744, 2027, 2000, 3636, 1285, 1502, 1990, 1902, 3126, 879, 2678, 3011, 1471, 3390, 1984, 1865, 1548, 2009, 3647, 1985, 2217, 1866, 2168, 2680, 1807, 2063, 2961, 3170, 2526, 1028, 1709, 1841, 3141, 1261, 1365, 1381, 1768, 2502, 2503, 3676, 2588, 1504, 814, 1022, 2820, 2832, 2031, 1366, 3671, 1876, 3659]. algorithmic [928]. Algorithms [318, 1617, 1731, 840, 1844, 1679, 1158, 585, 2406, 307, 380, 336, 2551, 1537, 1619, 213, 429, 2207, 1747, 1456, 495, 1017, 533, 516, 1159, 93, 1326, 3717, 1343, 2159, 1425, 1971, 3116, 2579, 2456, 1555, 1980, 1560, 1981, 2177, 2992, 3300, 2906, 1137, 1346, 2310, 2756, 902, 2537, 1227, 1831, 2865, 1265, 2136, 1659, 3471, 2595, 2234, 2499, 1958, 1123, 1315, 1863, 2867, 997, 3546, 2061, 935, 2028, 2917, 3747, 1954, 1787, 2337, 1472, 1503, 829, 1707, 3129, 2928, 2210, 1758, 2090, 883, 1250, 1974, 3157, 1209, 1383, 2030, 3674]. alignment [2832]. all-nearest-neighbors [1921]. Allen [3419, 3458, 3738, 3365]. allocated [2012]. allocating [899]. Allocation [13, 429, 335, 1922, 1516, 1453, 788, 807, 1235, 1334, 1175]. allocations [1603]. Almost [581, 589, 3132, 3093, 1368]. alpha [3423]. alpha-stable [3423]. Alternate [1799]. Alternate-Block-Factorization [1799]. Alternating [400, 805, 967, 3242, 1803, 1349, 2044, 874, 990, 1126, 3572, 2192, 793, 3318, 1021]. alternation [2109]. Alternative [1584, 2269]. Altran [774]. Alwac [106]. always [2018]. American [3287]. AMLI [2602]. Amodio [2452]. among [1612]. Amortized [1980, 1518]. amount [992]. Ampère [3195]. Analogue [30, 31]. analyses [2163]. Analysis [453, 3368, 2353, 929, 2365, 1463, 504, 738, 1657, 2320, 2686, 2843, 2746, 599, 1432, 3084, 3244, 558, 3254, 3619, 671, 2268, 2568, 2334, 359, 745, 3648, 3702, 1806, 1052, 3653, 643, 1857, 103, 1191, 3465, 2999, 2435, 1747, 1456, 3273, 726, 3286, 2328, 2961, 3711, 534, 628, 2670, 2832, 187, 3685, 3683, 1313, 1445, 2342, 3418, 2159, 1764, 3242, 2988, 1476, 1921, 2494, 2044, 1892, 2053, 3436, 3383, 786, 1376, 3527, 1505, 1980, 2039, 2024, 3198, 3213, 2430, 2991, 3427, 1327, 2404, 2394, 1433, 3228, 3462, 3450, 39, 1164, 2481, 2006, 2957, 2133, 3374, 3164, 2191, 3752, 3463, 2675]. analysis [3725, 3713, 3246, 3517, 2205, 3215, 3333, 3113, 1315, 2801, 863, 3326, 2060, 3311, 3128, 1353, 3188, 3352, 1097, 3643, 3698, 3432, 2680, 2245, 3098, 2048, 2228, 1028, 3535, 2196, 3328, 2447, 2918, 2856, 980, 1875, 1464, 3402, 3521, 796, 2807, 2049, 3731, 2986, 3205, 3536, 3720, 3470, 2845, 2503]. Analytic [366, 425, 526, 427, 2356, 1268, 2003, 1750, 2907, 1492, 1441, 1895, 2300, 2061, 1563, 792, 3286, 3472]. Analytical [1677, 193, 774, 1356]. Analyzing [3008, 1555]. and/or [3521]. Andersen [78]. Angled [99]. angles [3439, 3461]. angular [2541]. Anisotropic [2854, 3096]. Announcement [2303, 2543, 2261]. Anti [2642, 3351, 2581, 3460]. Anti- [2642]. Anti-/Ferromagnetic [2642]. anti-Gauss [2581]. Anti-Gaussian [3351]. anti-symmetric [3460]. Antiaircraft [5]. antiperiodic [2345]. any [1526, 3329, 1648]. AOR [1214, 2341]. appearing [2060]. apples [1689]. Applicability [295]. applicable [2333]. Application [3441, 2281, 1657, 235, 606, 2489, 646, 1424, 2231, 21, 2844, 1476, 1401, 1086, 2256, 1101, 2105, 2522, 2783, 2886, 3331, 3581, 3398, 1766, 2463, 3678, 1852, 1254, 1301, 3722, 3193, 3454, 1020, 2213, 2277, 2068]. Applications [2713, 1879, 2686, 1956, 2340, 2666, 40, 145, 1330, 1200, 2343, 3116, 2344, 3588, 2954, 3436, 1804, 1635, 2797, 2189, 2440, 2644, 2955, 1795, 2537, 2311, 3517, 1527, 3167, 3090, 2272, 3440, 2842, 2127, 2614]. Applied [2549, 3380, 2942, 3183, 2162, 3213, 1821, 1789, 3308, 2610, 3052, 3453]. Approach [2471, 285, 670, 493, 677, 2739, 2558, 3673, 2342, 3564, 2519, 3106, 3280, 2944, 3736, 1012, 2510, 3134, 2477, 1725, 1872, 2511, 2996, 3570, 2948, 2300, 3740, 2358, 3022, 3288, 1875, 1559]. Approaches [101, 3685, 3542, 2243]. Approximant [629]. approximants [1529, 1506, 791]. Approximate [1785, 2561, 702, 739, 2105, 1433, 193, 2747, 1537, 725, 2280, 2238, 2473, 1184, 3187, 2347, 3665, 3112, 2136, 2386, 3647, 3076, 3649]. Approximating [1952, 2709, 2223, 2172, 3318]. Approximation [48, 1927, 2705, 3479, 2455, 377, 638, 2458, 600, 641, 2749, 2424, 238, 383, 747, 647, 3102, 302, 44, 2671, 855, 3223, 3729, 3445, 804, 3131, 3714, 3081, 3477, 968, 3553, 3586, 2379, 856, 1880, 955, 3580, 3539, 867, 3071, 2632, 2945, 3706, 1219, 900, 1072, 1013, 1547, 1407, 2846, 3048, 1831, 1525, 2132, 3085, 2775, 3672, 3180, 3721, 2634, 3516, 2908, 3743, 3152, 3662, 845, 3645, 2008, 3475, 3124, 3546, 3336, 3481, 3558, 1216, 2253, 2818, 2917, 3465, 2361, 792, 2804, 3705, 1544, 3272, 1667, 1867, 3549, 2209, 2123, 3695, 3274, 2928, 1042, 1372, 3689, 907]. approximation [3013, 3723, 2589, 3536, 2109]. Approximations [2810, 639, 656, 720, 11, 33, 81, 353, 444, 569, 3029, 1399, 2140, 2218, 2293, 1640, 1696, 3477, 3082, 3382, 1946, 3383, 2003, 1079, 584, 1894, 3264, 3589, 1565, 3487, 2957, 1040, 3627, 3628, 2674, 3442, 2136, 3020, 3089, 60, 3430, 3455, 3333, 1417, 2654, 846, 975, 1884, 976, 3433, 775, 2841, 2448, 2767, 3315, 3468, 869, 3470]. approximieren [60]. Arbitrary [203, 3544, 288, 3225, 3513, 3299, 3443, 3281, 1872, 331, 1247, 1564, 1461, 2155]. arborescences [1744]. Arc [2706, 646, 1887, 1770, 1128]. architecture [1585]. architectures [2039]. archive [2969]. area [1667]. Argument [177, 206, 2103]. Arioli [3050]. arising [2439, 3668, 1176, 2511, 1080, 1714]. Arithmetic [896, 505, 127, 329, 342, 354, 604, 494, 506, 553, 384, 2350, 465, 415, 215, 386, 310, 70, 838, 3225, 3687, 787, 932, 3358, 3408, 1488, 1145, 3066, 2075, 851, 1270, 1338]. arithmetics [1357]. arity [3329]. Arnoldi [3133, 3279, 2177, 2774, 3094, 3338, 3349, 3140, 3431, 3218, 2184, 2640]. Arnoldi-type [3133, 2774]. Arranging [1689]. Array [1539, 1683, 1572, 1689, 1841, 1193]. Arrays [1746, 1558, 1686, 1073, 788, 921]. arrival [2277]. Arrivals [2564]. Arsenal [70]. Artificial [2782, 2824]. Aspects [710, 1160, 3106, 2521, 1060, 3040]. Asselin [3164]. assemble [3263]. Assessment [1537]. asset [3287, 3401]. assign [2000]. Assignment [654, 381, 1639, 534, 1010, 1984, 1809, 1022]. assimilation [2974]. Associated [2697, 240, 3551, 3586, 1015, 2826, 1753, 3442, 1353, 1973, 2829]. assumption [1962]. assumptions [1120]. Asymmetric [481]. Asymptotic [2578, 3488, 2467, 2319, 1142, 2546, 2745, 2320, 2189, 3514, 2070, 290, 2899, 3390, 2304, 2858, 3186, 2088, 1417, 3643, 1761]. asymptotic-preserving [3186]. Asymptotical [2710]. Asymptotically [3645, 2965, 1420, 2099]. asymptotics [1608]. Asynchronous [568]. atmospheric [2902]. Attainable [846, 2387]. attractivity [2290]. Attribute [1840, 2040, 1469, 1364]. attributed [1842, 1312]. attributes [1705]. Aubin [2635]. augmentation [3246].

Augmented

[2419, 3508, 2052, 2297, 3005, 3347, 3250]. August [56]. Author [2714, 2364, 2377, 2415, 2545]. Authors [1093, 1113, 1136, 1156, 1170, 1189, 1208, 1224, 1240]. autocorrelation [3017]. autocorrelation-based [3017]. Autocorrelations [107]. Automata [163]. Automated [3463, 1427, 2855]. Automatic [741, 1750, 3605, 673, 2445, 3571, 129, 920, 43, 2171, 134, 1369, 2874, 3748, 3500]. Automatically [1746]. Automorphism [488]. autonomous [1265, 1299]. autoregulation [2971]. auxiliary [1802, 3570]. av [92]. Average [1570, 1391, 1960, 1377, 2224, 2447, 1768]. Average-case [1570]. averaged [3473, 1735, 3272]. averages [1830]. averaging [2816, 3497, 3671]. AVL [1959]. avoid [3555]. avoidance [1602]. Avoiding [2615, 978]. award [2303]. Axel [3194]. axiomatic [1519, 1809, 764]. axiomatization [1243]. axis [1143, 916].

B [2216, 2934, 2991, 2800, 2087, 3544, 3600, 1598, 3026, 3075, 2213]. B-convergent [3544]. B-series [2991, 2800]. B-spline [2934]. B-splines [2087, 3600, 1598, 3026, 3075, 2213]. B. [2719]. back [3695, 2128]. Backtracking [1084, 1435, 1908]. Backup [574]. Backus [212]. Backward [3383, 3755, 2331, 719, 471, 2404, 1353, 2255, 1166, 2526, 3205, 3651, 3439, 2342, 3294, 3418, 2306, 3602, 3567, 2395, 2369, 2006, 2251, 2483, 3215, 772, 3650, 3338, 3594, 2461, 2438, 3680, 3548, 1963, 830, 2124, 2212, 3402, 3076]. Backward-Difference [471]. Bagley [2520, 3711]. Bairstow [244, 2166]. balanced [3371, 3186, 1875, 2771]. balancing [3605]. Banach [2504]. Band [2604, 3556, 2998, 3559, 478, 1539, 270, 2476]. Band-times-circulant [3559]. Band-Toeplitz [2604, 3556]. banded [3392, 3384, 2912, 2901]. banker [1602]. barrier [1410, 1411]. Barycentric [3726, 3035, 2917]. Base [859, 3639]. Based [2016, 2696, 524, 296, 2566, 552, 137, 2326, 2571, 1059, 1767, 2669, 2559, 2472, 3685, 3729, 3717, 3493, 2823, 3616, 2159, 1094, 3159, 3381, 2439, 3371, 2392, 3602, 3426, 1309, 3700, 3437, 2605, 2104, 3108, 3163, 3307, 3367, 2240, 1478, 1296, 1297, 2673, 2431, 3617, 3017, 2149, 3062, 2480, 3085, 2645, 3005, 3149, 2298, 2311, 3087, 3503, 3112, 2136, 2997, 3284, 2198, 2252, 1173, 2909, 3138, 3486, 3728, 2950, 2586, 997, 3311, 3340, 3351, 3011, 3140, 3259, 1217, 3000, 3074, 3549, 3507, 3237, 2961, 3288, 2805, 3141, 1365, 3509, 3622, 3205, 2068, 1914]. Bases [574, 3446, 2497, 3409]. Basic [1679, 2573, 345, 1419]. Basis [1792, 477, 1973, 2469, 2235, 3070, 3209, 900, 1468, 3617, 3287, 764]. Batch [2564]. batched [1806, 1637]. Battles [192]. BCH [2450]. BCH-Formula [2450]. BDF [2577, 3684, 2119, 2942, 2531, 2894, 3461, 1351, 3316, 3524]. BDF-formula [1351]. BDF2 [3294]. BDF2-Maruyama [3294]. BDFs [1460]. BDM [3645]. Be [2800, 1679, 3064, 102, 2327]. Bearbejdning [15]. Bech [717]. Behavior [480, 1191, 3425, 1344, 2925, 1130]. behaviour [1959, 2242, 3385, 1835, 811, 2638, 1319]. Below [575, 637, 634]. Beltrami [3023]. BEM [3178]. Benchmarking [1747]. benchmarks [3638]. Benjamin [3596, 2675, 2301]. Bernoulli [1698, 2440]. Bernstein [2189, 2798, 2873, 2618, 2217]. Bessel [3664, 2498, 3725, 3713, 1317, 3077]. Best [644, 2424, 976, 3223, 2632, 584, 2077, 2917, 1503]. best-bound [1503]. better [1355]. Bettis [1664]. Between [2401, 621, 3201, 1059, 3212, 1789, 107, 2072, 1543, 287, 3129, 1421, 1925, 371]. Bézier [3223, 2798]. Bezout [2338]. BFGS [3550]. Bi [411, 2111]. BI-CGSTAB [2111]. Bi-Unique [411]. Biased [1219]. bicubic [2044, 2036]. bidiagonal [2343, 2423]. bidiagonalization [3080, 3525, 1697, 2095, 2354, 2915]. Bifurcations [2325, 2773]. Biharmonic [567, 990, 3138, 2623]. bijective [3342]. bilinear [1080]. Bin [1586]. Binary [1553, 1158, 1328, 77, 554, 1357, 1535, 884, 824, 1676, 1516, 1346, 875, 1873, 1075, 1588, 1779, 1902, 793, 1175, 1808, 1261]. binary-exponent [793]. binomial [2860]. biochemical [3166]. biomolecular [3445]. Biot [3360, 3653]. Birkhoff [2135, 675]. birth [3506]. birth-and-death [3506]. birthday [1703, 2064]. bisection [3125]. bisections [3044]. bisectors [2232]. bistability [3021]. BIT [27, 188, 78, 637, 1398, 1320, 1629, 517, 81, 28, 120, 1089, 562, 2316, 394, 1202, 498, 499, 1308, 135, 890, 1630, 543, 1651, 1002, 465, 1964, 1338, 1836, 1874, 866, 433, 1496, 1655, 1920, 2364, 2377, 2415, 2545, 2655, 3206, 3276, 2413, 2378, 3171, 3015, 3067, 3079, 3092, 3104, 3130, 3158]. BIT50 [2963]. Bivariate [3277, 136, 2804, 2556, 3729, 3224, 2407, 2514, 2909, 3074]. black [1972, 1789, 3094, 2623, 2973]. black-box [3094, 2623]. Blasted [214]. blended [1194]. Blending [1263, 3062]. Block [1799, 418, 3704, 744, 1846, 2733, 1793, 2595, 3724, 2337, 2328, 1200, 461, 497, 628, 2400, 3511, 564, 1796, 3404, 3080, 2439, 1314, 1027, 3057, 1561, 3384, 2163, 2105, 3568, 1753, 3309, 2134, 3246, 3519, 3581, 1259, 2776, 1857, 1418, 3390, 2274, 1449, 2680, 2099, 2425, 3671, 2625]. block-implicit [1449]. block-Jacobi [3390]. block-matrix [1796]. Block-projections [2337, 2274]. block-tridiagonal [2400]. Blocked [2865]. Blocks [1074, 2733, 3404, 2233, 2337]. blockwise [3081]. Blow [2850]. Bo [517]. body [3269, 2090, 3454, 3270]. Boltzmann [2831]. Bona [3596]. Books [64, 79, 90, 96, 114, 123, 132, 159, 167, 190, 202, 209, 231, 243, 272, 281, 292, 338, 361, 392, 420, 435, 450, 540, 560, 578, 594, 650, 679, 697, 732, 782, 802, 819, 853, 871, 889, 909, 925, 961, 985, 1007, 1024, 1044, 1069, 1092, 1111, 1132, 1154, 1169, 1204, 1238, 1257, 1272, 1307, 1324, 1395, 1359, 1374, 1451, 1466, 1481, 1495, 1534, 1551, 1582, 1601, 1615, 1632, 1654, 1675, 1722, 1801, 1740, 1979, 2023, 2079, 2230, 373, 501, 325, 519, 613, 767, 484, 633]. Books [316, 252, 714, 834, 348, 464, 46, 666, 54, 750, 406, 72, 104, 179, 220, 262, 304, 363, 942, 1187, 1222, 1533, 1567, 1837, 1871, 1995, 2051, 2080, 2158]. Boolean [517, 617, 207, 748, 1827]. bordered [1846]. Boris [3385]. Bose [3509]. bottom [2464]. bottom-up [2464]. Bound [321, 2341, 2246, 1904, 1268, 3613, 3742, 1700, 2020, 971, 1503, 1628, 1392, 1925]. boundaries [3496, 3286]. Boundary [275, 489, 224, 567, 414, 664, 608, 387, 45, 2892, 2032, 1399, 1661, 800, 755, 3159, 3207, 2468, 3444, 3058, 3555, 1054, 1230, 858, 839, 969, 1125, 1320, 1540, 1642, 3145, 3173, 3528, 3030, 2946, 1893, 3047, 3175, 1057, 2995, 3482, 3450, 1440, 2285, 2720, 3049, 957, 1663, 3088, 2136, 2886, 3216, 1413, 1417, 2463, 2979, 3424, 1717, 1060, 2841, 3289, 1082, 1249, 1318, 3737, 2727, 3689, 849, 2806, 3014, 3720, 3511]. Boundary-Value [45]. Bounded [3685, 1425, 3297, 2202, 3178, 2909, 2181, 792, 3415]. Boundedness [2546, 3515, 3007]. Bounding [1164, 3181, 1860, 948, 3170]. Bounds [684, 3262, 2344, 824, 3755, 67, 1604, 1530, 2340, 2397, 2409, 827, 445, 549, 1147, 688, 2360, 385, 2598, 259, 479, 3029, 1959, 2140, 1812, 1903, 3477, 1892, 3419, 2631, 2861, 2922, 3372, 2331, 3071, 787, 1385, 1894, 3264, 1847, 2382, 3612, 2634, 3442, 1562, 2251, 917, 2008, 2206, 1332, 1367, 2012, 2461, 2438, 1578, 499, 1246, 1308, 762, 3705, 2981, 3235, 2170, 2257, 3722, 1198, 1255, 1901, 1993, 2212, 3156, 1295, 1462, 3676, 3076, 3205]. Boussinesq [3534]. box [1972, 3094, 3000, 2623, 3191]. brain [140].

Branch [1765, 321, 1700, 1503]. branch-and-bound [1503]. Branch-Bound [321]. branching [1700]. breakdown [3309]. breakdown-free [3309]. Breakdowns [2225]. Brent [1124]. Brickell [1706]. Bridging [2650]. broadcasting [1997]. brother [949, 1097]. Brouwer [2847]. Brownian [1502, 3651]. Broyden [2990, 1905]. Brugnano [3192]. Bruno [550]. BS [3464, 2911]. BSOR [3581]. BSQI [3348]. Bucket [1192]. buckets [1376]. buddy [807]. buffering [1981]. BUG [3700]. building [1292]. built [3002, 1160]. built-in [1160]. bundle [3553]. Burgers’ [3348, 3347, 2699, 3153]. Butcher [2138, 3607, 1962]. BV [3017]. BVPs [3341]. Byte [433].

C [3072]. Cahn

[3419, 3458, 3654, 3738, 3365, 3524]. Cake [300]. Calculate [376, 1847]. Calculating [550, 416, 3134]. Calculation [30, 31, 62, 63, 298, 584, 474, 537, 394, 355, 3225, 3621, 2929]. calculations [2183]. calculi [1258]. calculus [1942]. Calderón [3704]. calling [765]. Calls [709, 760]. Camassa [3530]. can [2309, 3064]. Cannot [2800]. Canonical [1953, 528, 562, 1001, 1947, 2358]. Caputo [3450, 3216]. Carathéodory [3013]. Card [76]. Cardano [3153]. cardinality [3633]. Carl [2064]. Carl-Erik [2064]. Carlo [3082, 192, 3525, 3068, 1077, 3185, 2250, 2284, 5, 3584, 3217, 3607, 3643, 730, 3234, 3354, 216, 3649]. Carousel [40]. carrier [2605]. Cartesian [1922, 1516]. Cartography [225, 322]. cascades [1357]. Case [622, 2619, 16, 216, 589, 1688, 1344, 1570, 1072, 1882, 1276, 3430, 3455, 2224, 3423, 1587, 2371, 2194, 2447, 795]. cases [1128, 1715]. Casteljau [2217]. Casulli [3192]. Cauchy [84, 3587, 2293, 2281, 618, 1576, 2282, 472, 1029, 1184, 3094, 1774, 1058, 1895, 3648, 401, 1756, 2117, 882].

Cauchy-Problems [84]. Cauchy-type [1058]. Cavity [299]. Cayley [2508, 2089]. CCC [3441]. celestial [148]. cell [1211]. Cellular [365]. centered [1621, 1641, 1949, 1716]. Central [3406, 1413, 84, 216]. centre [55]. Centroids [2571]. Certain [439, 246, 482, 320, 605, 569, 478, 45, 93, 206, 3404, 989, 905, 2994, 1771, 1530, 1488, 86, 1128, 1109, 775]. Certification [188, 1617, 498, 499]. CFG [1210]. CFL [3616]. CG [1798, 2815, 2611, 1787]. CG-type [1798]. CGSTAB [2111]. chain [2947, 3415]. chains [2296]. Chance [514]. Chance-Constrained [514]. changed [3480, 3516]. Channel [196, 1984]. channel-assignment [1984]. Channels [1220]. chaos [3627]. chaotic [2141]. Character [1484, 336]. characterisation [1109]. Characteristic [376, 504, 692, 2995, 2302]. Characteristics [66, 2722]. Characterization [3021, 1105, 2239, 789, 1898, 1549, 2074, 2612]. characters [1542]. charged [3385, 3625]. charged-particle [3385, 3625]. charges [3014]. Charts [735]. Chawla [1852]. Cheap [3299, 2586]. chebfun [2917]. chebop [2874]. Chebyshev [81, 700, 855, 804, 2977, 3635, 638, 1219, 3061, 1541, 1671, 33, 1184, 1231, 1407, 3358, 3408, 549, 1149, 3662, 1977, 3114, 2926, 923, 3401, 289, 302]. Chebyshevian [3464, 1104]. check [3576, 1298, 1311]. Checking [2160, 741, 7, 161, 139]. Chemical [2517, 2724, 2723, 2825, 3324, 2849]. chemistry [2179, 1405]. chemotaxis [2724]. Chernoff [3232]. Chevalley [1542]. Chief [3317]. Choice [2732, 2688, 2632, 2482, 1061]. Cholesky [2163, 2114, 2398, 1558, 1901]. cholesteric [3508]. Choosing [1227]. Chr [78]. Christoffel [2219, 2248, 2322]. Cichelli [1725]. CIR [3423, 3649]. Circle [565, 490, 1510, 3125]. Circles [2329, 2098]. circuit [53]. circuits [1694]. Circulant [1928, 2366, 1989, 2487, 3554, 2151, 3559]. circular [1887, 1621]. circular-arc [1887]. Circularly [759]. Cited [2413]. Class [328, 469, 655, 487, 524, 454, 566, 720, 299, 2340, 657, 2570, 605, 659, 2751, 725, 692, 402, 2670, 2964, 3144, 1280, 2418, 1769, 1078, 1859, 2592, 857, 1629, 1592, 1079, 3659, 1816, 2367, 2534, 3694, 1541, 2536, 1408, 914, 2131, 1813, 2007, 3471, 3702, 950, 3231, 2911, 1620, 2462, 2679, 1548, 1316, 1333, 774, 1549, 810, 1042, 1372, 3220, 3595, 3563]. classes [1771, 1126, 3139, 1612]. classical [3538, 2955, 1196, 2355, 1492, 2252, 2768]. Classification [50]. clause [848]. Clausen [3621, 631, 2929]. Clausus [476]. Clenshaw [3332, 3398, 343, 2481, 445, 3257, 1317, 1255, 3012, 3203, 2526, 592, 2740]. closed [3681]. Closest [1625, 1211]. Closure [357, 906, 2780]. clustered [2182]. Clustering [1767, 1227]. Clusters [2356, 332, 384]. CMRH [3039]. coarse [3731]. coarsening [3044]. COBOL [26, 17, 34, 181, 26, 17, 34]. COBOL-grammatiken [26]. COBOL-kompilering [34]. Code [170, 119, 2605]. Codes [7, 2439, 1277, 972, 1036]. Coefficient [249, 270, 2837, 3584, 847, 3523]. Coefficients [581, 193, 306, 2700, 3223, 1927, 1014, 2785, 3644, 3430, 3455, 2357, 2924, 2978, 3038, 3698, 3651, 3455]. cographs [2031]. Collaborators [2913]. Collection [669, 223]. collector [813]. Collocation [2743, 2661, 933, 2702, 627, 778, 371, 2965, 2044, 2859, 2470, 2953, 3341, 2054, 2247, 2532, 1749, 1629, 1592, 1056, 1610, 1752, 2970, 1663, 3216, 2198, 1081, 3285, 1814, 3322, 3679, 2951, 2980, 3352, 2229, 1109, 1650, 3052, 2387, 3711, 2806, 2791, 2260, 1914]. collocation-based [2198, 1914]. Color [3583, 3370, 3369]. Coloring [2275]. Colour [570]. column [2457, 3040, 2588]. columns [2337]. combination [2934]. Combinations [69, 1500, 1536, 1727]. combinatorial [1425, 1680]. combined [2858, 2605]. Comments [867, 323]. commit [1877]. Common [237, 334, 1347, 883]. Commutator [3565]. Commutator-free [3565]. commuting [2214]. Compact [3443, 1030, 710, 3662, 2027, 2635, 3100, 3220, 3523, 3484]. compactifying [813]. compactly [3002, 2074]. companion [3117, 3541]. Company [85]. Comparing [2102, 686, 3017, 1817]. Comparison [3121, 2663, 3568, 3187, 287, 1686, 3525, 829, 1974, 921]. comparisons [3696, 1707]. Compatibility [743]. Compilation [740, 531, 625, 34]. Compile [901]. Compiled [153]. Compiler [124, 106, 58, 296, 181, 345, 87, 91]. Compilers [125, 170, 798, 161]. Complement [1677]. complementarity [2940]. complementary [62, 63]. Complete [1909, 1295, 1559]. Completeness [1258, 1447]. completion [3137]. Complex [177, 169, 2701, 59, 2063, 70, 1621, 2629, 3687, 786, 2905, 3528, 3732, 2886, 1896, 2722, 1202, 1146, 1563, 1386, 1216, 1579, 218, 998, 3563]. complexes [3448, 3449]. Complexity [2553, 1969, 1942, 1688, 1937, 2476, 1407, 1744, 2041, 1518, 1502, 1348, 1856, 2852, 1486]. Component [1892, 1657, 2984, 1745]. Component-wise [1892]. Components [85, 92]. Componentwise [2327, 1993, 2880]. Composite [582, 3362, 3173, 1990]. Composition [585, 2115, 3734, 2162]. compound [922, 1917]. compressible [3290]. Compression [1484, 2038, 2810, 1345]. compression-oriented [2038]. computability [848]. Computable [3442, 3165, 3419]. Computation [88, 3405, 1403, 2, 970, 89, 2498, 2701, 136, 1542, 2887, 52, 689, 530, 2461, 3038, 3639, 626, 1434, 711, 206, 55, 3117, 2173, 3118, 3045, 1354, 2237, 2330, 981, 930, 3209, 1404, 1844, 3175, 2222, 932, 1163, 2883, 1441, 2761, 3503, 3409, 2996, 2459, 2410, 3532, 1215, 879, 2725, 2927, 1579, 1317, 3273]. Computational [575, 637, 3251, 2328, 2614, 417, 3133, 1160, 2895, 2969, 3410, 1620, 2386, 634, 3622]. Computations [2719, 154, 706, 763, 832, 2691, 2823, 2466, 3116, 1688, 1766, 2501]. compute [2241, 2191]. computed [2797, 2653]. Computer [37, 340, 106, 118, 378, 32, 2378, 457, 39, 76, 228, 1114, 446, 1220, 225, 322, 240, 570, 572, 1428, 2756, 245, 2996, 1684, 1383]. computer-aided [2996]. Computer-Related [32]. Computers [27, 65, 12, 23, 28, 77, 103, 1971, 117, 1418, 2245, 140]. Computing [439, 410, 580, 684, 1879, 2174, 1553, 1677, 2266, 1593, 2443, 2731, 2457, 2826, 1347, 2300, 3361, 2700, 2787, 426, 527, 2900, 2208, 217, 2985, 433, 2214, 1225, 3241, 988, 2175, 3057, 3121, 3383, 2066, 1125, 1062, 1320, 1937, 1561, 3727, 3732, 1014, 3176, 2476, 3748, 2774, 948, 1177, 2817, 1991, 3647, 3339, 3040, 2854]. comrade [3117]. Concave [2383, 919, 1000, 1235]. Concept [173]. Concepts [1679, 1483]. concerning [1267, 1037]. Concluding [585]. concrete [881]. Concurrency [1617, 1519, 1687]. Concurrent [1657, 1538]. condensates [3509]. Condition [2585, 688, 2741, 3294, 2880, 2629, 1446, 3421, 2369, 3265, 1073, 2180, 3693, 1861, 3629, 3702, 1415, 979, 1819, 2090, 3467]. Conditionally [11]. conditioned [840, 1404, 2534, 2522, 2968, 3556, 2077, 2349, 2099, 2425, 3103]. Conditionedness [2327]. Conditioning [489, 3472]. Conditions [275, 2567, 102, 1399, 1661, 3355, 1965, 3208, 1947, 3198, 3735, 1439, 2946, 2240, 3047, 3175, 2285, 2720, 3088, 2205, 1897, 952, 1000, 3424, 2170, 3289, 3737, 2091]. Conduction [183, 138, 196, 3101]. Cones [557, 816]. Conference [294, 2719, 2614]. confidence [2466]. configuration [3270, 3269]. configurations [3176, 2041]. conflicts [1173]. confluence [1804]. conformal [21, 3652]. conforming [3435]. Congress [56, 57, 232]. Congruences [2590]. congruential [1828, 1281, 2330, 1711, 2622]. Conjecture [718, 630, 184, 278, 139, 432, 2216, 3660, 1129, 1919]. conjugacy [1126]. Conjugate [2467, 532, 2712, 2690, 2417, 1781, 3669, 2104, 1561, 3107, 1784, 2178, 1795, 3309, 1790, 1665, 1794, 1418, 2626]. conjugate-symplecticity [3107]. connected [1745, 2014]. connecting [1321, 1337]. Connection [479, 1214, 3038]. Connections [1789, 2617, 2655]. Connectivity [416]. conquer [2476, 1822, 1709]. consecutive [1370, 1678]. Consequence [631]. Conservation [3069, 2685, 2664, 2549, 2824, 3256, 3579, 3462, 2916, 2722, 3234, 3354, 2194]. Conservative [2723, 3473, 3004, 2442, 3501, 3570, 3375, 2411, 3169, 3434, 3400]. conservative/dissipative [3570, 3434]. conserved [3291]. conserving [2152]. considered [1685]. Consistency [1140, 1609, 1773, 1185, 1321, 1337]. Consistent [2536, 3047, 2887, 1545]. consistently [1785]. Consolidation [673, 3360, 3653]. Constant [581, 306, 2710, 1225, 1984]. constant-time [1225, 1984]. Constants [605, 2375, 2797]. constellations [1678]. Constrained [3754, 3755, 2321, 2209, 514, 2093, 2871, 2579, 3741, 1972, 3106, 2331, 3635, 1264, 2510, 2069, 2130, 2150, 3581, 3270, 3269, 3452, 2015, 2351, 3310]. Constraining [3090]. Constraint [1765, 590, 919]. constraints [3242, 3081, 2432, 3180, 1332, 2777, 3126, 2804, 2037, 3601, 2258]. Constructing [2550, 1779, 1435, 2944, 1822, 2316, 1748, 2129]. Construction [3295, 117, 2591, 2366, 3135, 212, 171, 2421, 477, 3091, 3638, 2740, 2187, 2367, 3062, 1823, 3505, 2427, 2155]. Constructive [285, 3106, 2358]. Container [214]. Containing [621, 2337]. Contents [375, 395, 422, 437, 452, 468, 485, 503, 522, 545, 563, 579, 596, 616, 636, 652, 668, 683, 699, 716, 734, 754, 769, 784, 803, 823, 837, 854, 894, 927, 944, 965, 987, 1009, 1026, 1048, 1071, 2575, 2627, 2656, 145, 2659, 2682, 2692, 2704, 2716, 2729, 2770, 2793, 2809, 2822, 2781, 2834, 2857, 2870, 2903, 2879, 2742, 2891, 2919, 2931, 2975, 2987, 3028, 3042]. context [1363, 1485]. context-free [1363, 1485]. Continuation [2174, 1764, 3185, 1122, 2273, 3472]. Continued [629, 229, 1266, 1596, 1651, 1269]. continuity [2036]. Continuous [2744, 3019, 2167, 2126, 2530, 1904, 2221, 2482, 2268, 1488, 2085, 3285, 3464, 3154, 3098, 3099, 3170]. Contour [2686]. Contraction [738]. Contractive [953, 973, 1510]. Contractivity [2319, 1915, 2762, 1817, 1460, 1421, 1613]. Contribution [71, 62, 88, 63, 89, 241, 112, 207, 336, 130, 121, 135, 200, 250, 156, 149, 188, 3057, 498, 499]. Control [1617, 721, 256, 587, 551, 53, 1229, 248, 61, 1278, 3004, 3393, 1929, 1662, 1122, 2268, 3149, 1507, 3180, 2026, 2948, 3581, 2000, 1509, 830, 3533, 980, 1423]. Controlled [1190]. controlling [3493, 982]. Convection [2746, 3368, 1796, 3703, 3003, 3226, 3427, 2995, 3325, 3688, 2978, 2979, 3259, 3453, 3220]. Convection-Diffusion [2746, 3368, 1796, 3003, 2995, 3325, 3688, 3259]. convection-diffusion-reaction [3226]. convection-dominated [3703]. convection-subdiffusion [3220]. Convergence [254, 2110, 3016, 312, 3719, 2176, 425, 2914, 505, 693, 1576, 144, 2521, 826, 638, 2894, 2706, 3325, 2783, 1524, 2323, 3450, 194, 306, 2046, 3179, 2957, 3721, 1883, 2674, 2007, 2437, 3033, 2324, 1178, 1647, 2313, 2815, 2959, 3113, 3679, 506, 553, 2750, 2193, 3023, 2352, 2680, 258, 288, 2047, 2227, 2557, 3203, 2425, 609, 387, 3731, 3716, 3622, 3470, 3540, 2433, 1106, 3717, 3294, 2305, 1608, 2186, 805, 3242, 2439, 2118, 3382, 3056, 284, 3554, 3458, 3341, 1575, 2453, 3654, 1904, 2002, 1730, 1629, 1592, 2096, 757, 913, 1029, 3699, 2896, 1231, 902, 797, 1408]. convergence [1753, 2131, 3111, 1058, 3008, 3665, 3088, 2897, 3644, 3151, 2483, 1975, 2485, 3020, 3089, 1489, 2542, 3285, 3650, 3475, 3594, 3326, 3494, 1031, 1624, 2678, 2889, 3390, 3698, 1513, 2117, 3433, 2108, 1963, 3400, 1819, 1180, 2154, 2527, 1004, 1597, 3289, 3156, 3227, 2638, 882, 2504, 3315, 3467, 3221, 1218]. Convergent [3604, 629, 725, 1335, 3297, 3707, 948, 3544, 2192, 1931, 3424, 3275, 1182]. converges [2607]. Conversational [661, 922]. Conversion [3328, 859]. Converting [99]. Convex [3018, 2383, 919, 1000, 1247, 1564, 1546, 1900, 557, 1225, 1343, 816, 1585, 3490, 1838, 1830, 1956, 2634, 3032, 1292, 1832, 2410, 1362, 1235, 1329, 2073, 2028, 3560, 1666, 383, 2036, 2257]. Convex-Concave [2383, 919, 1000]. convexity [2652]. Convolution [313, 2648, 2988, 3160, 2795, 2799, 2385, 3678, 1549]. convolution-type [2385]. Cooling [214]. Coordinate [146, 3212, 1141]. Copenhagen [1684]. Copies [1840]. corner [2851]. Coroutine [461]. Correct [368, 2965, 1832]. Corrected [2086, 976]. Correction [1661, 3370, 3576, 3456, 3537, 3725, 3745, 3455, 2445, 3611, 3577, 3457, 3354, 43, 3673, 1399, 2578, 3602, 2161, 3444, 2381, 841, 1089, 2478, 3619, 3544, 1254, 1301, 3204, 1020, 3222, 3691, 3591, 3379]. Corrections [224, 1618, 2688]. Correctness [285, 508, 1538, 764]. Corrector [509, 3199, 1035, 1424]. correctors [331, 1424]. Correlation [2403]. corresponds [1701]. cosets [1854]. cosine [1696, 3069, 1079]. Cosmo [440]. Cost [2592, 2836, 1970, 2895, 1391]. cost-effective [1970]. costs [1960]. COSY [3362]. COSY-GLMs [3362]. Cotes [425, 2566, 1917]. coth [1611]. coth-kernel [1611]. Coulomb [290]. Count [866]. Counter [2664]. Counter-Examples [2664]. counterexamples [2059]. counterparts [2934]. Counting [1808, 1571, 1433]. Counts [654, 586, 2457]. coupled [1798, 3690, 2416, 2976, 3475, 3260, 3467]. coupling [2825, 2916, 2464, 3533]. Course [1679, 345]. covariance [3547]. covers [974]. COVID [3485]. COVID-19 [3485]. Cowell [1760]. CP [3583]. CPU [1220]. Crack [2237]. Craig [2139]. Crank [3487, 3164, 3188, 2597]. criss [3071]. criss-cross [3071]. criteria [2417, 1922, 3112, 1004]. Criterion [642, 137, 333, 358, 2456, 991, 1665, 3598, 2312, 3064]. Critical [956, 1002]. cross [3071, 1404, 3546]. cross-validation [1404]. crossing [2041, 2014]. crossing-free [2041]. Crouch [2427]. Crouzeix [3736, 2867]. Crucial [1537]. Crude [216]. cryptographic [2000]. cryptosystem [1037]. cryptosystems [956, 1002]. crystal [3174, 3175]. crystals [3251, 3508]. Cubature [1018, 1670, 1710, 2977, 2882, 2462, 2805, 3633]. cube [977, 1737, 2014, 1827]. cubes [1148]. Cubic [298, 474, 641, 3168, 676, 937, 968, 1511, 3106, 2221, 1525, 2813, 790, 2108, 1867, 2169, 3202, 1546, 1668, 1900]. cubical [3449]. Cullen [1003]. CUR [3543, 3728]. CUR-type [3543]. cure [3628]. curl [3667, 3180, 2721, 3435]. curl-conforming [3435]. curl-preserving [2721]. Current [2309, 3666]. Curtis [3332, 343, 445, 3398, 3257, 1317, 1255, 3012, 3203, 2740]. curvature [2257, 3141]. curvature-based [3141]. curvatures [2332]. Curve [49, 99, 742, 2706, 211, 3062, 146, 137, 311, 3002, 2175, 2354, 2165, 3149, 3009, 2213]. Curve-fitting [137]. curved [3513, 2933].

Curves [2694, 2565, 2530, 2495, 2798, 1644, 1525, 3334, 3312, 1867, 2169, 2338, 1886, 1964, 3681]. Customers [2564]. cut [3709]. Cuthill [1050]. Cutting [109]. cycle [2503, 3716]. cycles [2014, 776]. Cyclic [2606, 857, 1690, 1753, 2131, 2128]. cycling [811]. cyclotomic [763, 832]. cylinder [3732]. Cylindrical [138, 214].

D [1793, 3047, 1292, 1832, 3024, 3219, 2194]. D. [2216]. D.E. [297]. DAE [2836, 2536, 2538, 3388, 3630]. DAEs [2144, 2965, 2102, 3668, 2551, 3640, 3237, 3260, 2789, 2110, 2887, 2435]. Dafermos [3598]. Dahlquist [1410, 1411, 2768]. damped [1664, 3192, 2624]. damping [3010]. Dangerous [674]. Danish [15, 25, 53]. Darcy [2600, 3520]. DASK [15, 15]. Data [2705, 152, 125, 3087, 50, 1174, 128, 42, 1787, 2571, 1330, 677, 574, 2329, 335, 403, 3717, 134, 3341, 2281, 1603, 3393, 3566, 2310, 1426, 1999, 2512, 2915, 2936, 3632, 3648, 2910, 3231, 1824, 3608, 1605, 1825, 1985, 383, 2540, 2828, 1638, 2974, 2210, 3170, 865, 2841, 1028, 2197, 3618, 3575, 3320]. database [1431, 1587, 1910]. databases [1968]. Datalogy [1684]. Datamaskinen [140]. Davidson [3446, 3476, 2185]. Davis [917]. Days [1116]. Deadbeat [1423]. Deadlock [1602]. death [3506]. deblurring [3147]. Debugging [277]. Decay [48, 3431]. Decay-Type [48]. decentralized [1877]. decide [816, 3064]. Deciding [1183]. Decision [283, 1295]. decoders [25]. Decompilation [625]. decomposability [1425, 1729]. Decomposed [704]. Decomposition [617, 389, 642, 2749, 3652, 2376, 1200, 479, 928, 2232, 2145, 3132, 1812, 3583, 3703, 3527, 2631, 2922, 2473, 3300, 2240, 2220, 2906, 2163, 3707, 817, 2537, 1414, 1854, 3543, 3745, 3739, 2020, 3486, 3728, 1587, 2009, 1985, 1701, 3433, 2842, 2156, 2314, 1933, 2277]. decompositions [1903, 2441, 2267, 815, 1993, 1788, 3674]. deconvolution [2946, 2673, 3087]. decreasing [1475]. Deep [3661]. deepening [2013]. defect [3159, 1929, 841, 1254, 1301, 1020, 3691]. defect-correction [3691]. Deferred [2478, 2688, 3602, 2381, 3544, 3204, 3591, 3379]. deficiency [3498]. Deficient [589, 2523, 2267, 2274]. Defined [2687, 1844, 336, 1666]. Definite [112, 120, 390, 2328, 3701, 2398, 2486, 3249, 972, 3747, 1197, 2639, 3403, 2626]. Definiteness [2752]. Definition [212, 247, 60, 3236, 3190, 1005]. definitions [1940, 1941]. deflating [3123, 2826]. Deflation [441, 3446, 3402]. degeneracy [1305]. Degenerate [472, 3003, 2634, 2674, 1300, 3688]. degree [3223, 1792, 1670, 3281, 1108, 1247, 1564, 2254, 977, 1065, 1150]. Delaunay [1377, 1362, 2015]. Delay [2319, 3323, 2334, 2325, 2555, 2702, 2558, 2433, 2279, 2081, 2201, 2953, 2521, 3254, 3148, 2133, 2513, 2584, 2071, 2198, 3573, 3675, 3052, 2387, 3027, 3115, 2426, 1988, 2270]. Delay-dependent [3323]. delay-differential [2133]. Delays [2743, 2953, 2430]. Dennis [1234]. denoising [2796, 2587, 3141]. Dense [2366, 2747, 2509, 2431, 2233, 788, 2398]. Dependence [722]. dependencies [1968, 3339]. Dependent [2693, 237, 2615, 2604, 3436, 3172, 2773, 3323, 3361, 3188, 3453, 811, 3321, 3536]. Depth [703, 1448, 1522]. depth-first [1448]. derangement [1726]. Derangements [1049]. Derivation [307, 529, 240, 416]. derivations [1448]. Derivative [1508, 3029, 1282, 2522, 3450, 2056, 3311, 3289]. Derivatives [380, 430, 3071, 1712, 2132, 2481, 3628, 3415, 3678, 880, 2037, 3191, 3041, 3337, 2589]. derive [2880]. derived [968, 1676, 994, 2960]. Deriving [3457, 1455]. Descent [302, 3144, 3212, 3490, 2611]. described [156]. describing [2210]. description [3024]. descriptions [1694]. Design [49, 1554, 475, 154, 661, 1376, 1483, 1051, 1692, 3064, 1766, 87, 91, 1682, 938, 1695, 2213]. designed [2460]. destabilizes [2010]. destruction [3375]. detailed [1433]. detect [1872]. Detecting [2992, 3747, 2652, 2773, 2041]. detection [1311, 2089, 2314]. deteriorated [3252]. determinant [3519]. Determinantal [2324]. determinants [3225, 2773]. Determination [900, 2839, 370, 983, 791, 3321]. determined [3550]. Determining [488, 266, 446, 725, 2090]. Deterministic [257, 346, 929, 3095]. deunifications [1856]. Development [512, 1118, 1939, 1051, 1260]. Deviates [136]. Deviation [423]. Devices [212]. DGELS [2474]. diagnosis [39]. Diagnostics [1173]. Diagonal [2168, 2353, 3704, 3086, 2316, 1214, 2423, 2129]. diagonalisable [2295]. Diagonalizable [2395, 3372]. diagonalization [2680]. Diagonally [2033, 2647, 3105, 2043, 3132, 2454, 1087, 3150, 2878, 3603]. Diagram [585]. diagrammatic [2596]. Diagrams [340]. Diamond [1157]. dichotomy [2992]. Dictionary [1484, 3288]. dielectrics [2923]. diffeomorphism [3661]. Difference [740, 719, 471, 11, 341, 388, 379, 2709, 306, 621, 475, 659, 387, 2703, 3439, 1399, 2306, 1262, 284, 858, 839, 969, 3297, 3060, 3443, 1055, 1056, 1038, 2442, 1893, 3686, 3450, 957, 2107, 1178, 3165, 2898, 60, 1413, 1333, 3415, 3432, 3424, 3400, 1422, 3289, 1249, 1019, 3220, 3378, 3523, 3622]. Differences [663, 611, 2947, 84, 3051, 1103]. Differencing [2686, 3486]. Different [102, 3031, 2190, 2243, 3688, 1852, 3425]. differentiable [2989, 3139]. Differential [294, 116, 1799, 701, 2319, 2743, 582, 487, 2643, 597, 2745, 740, 2685, 203, 404, 275, 536, 2458, 657, 2334, 2552, 2325, 2698, 2688, 2555, 674, 2700, 185, 2711, 2735, 2736, 3638, 592, 2702, 2558, 2671, 2433, 3105, 3043, 3224, 2119, 2291, 2416, 3131, 2279, 3159, 2218, 1927, 3068, 1262, 1314, 2081, 2507, 1027, 3602, 904, 2393, 2201, 2470, 2953, 897, 1078, 2147, 2246, 2616, 2033, 2429, 2651, 989, 3526, 3069, 2307, 2672, 2162, 3437, 3623, 969, 1162, 1336, 3473, 3697, 2054, 3448, 2810, 1252, 2862, 912, 3474, 3499, 3579, 2944, 3529, 3210, 3480, 3211, 2521, 2874]. differential [2381, 3567, 2510, 1176, 1101, 3699, 886, 1057, 841, 2773, 2757, 3748, 902, 3395, 2935, 1662, 3254, 1232, 1699, 1732, 2408, 2479, 2837, 2775, 3672, 3063, 2759, 2084, 3615, 3620, 2799, 2133, 3323, 2513, 1144, 2286, 2335, 2525, 1016, 2007, 2885, 2897, 3033, 3516, 84, 1178, 3570, 2876, 3034, 2584, 2205, 3095, 2071, 2198, 2948, 60, 2088, 3037, 3333, 1300, 3573, 2676, 1918, 2348, 3679, 1246, 1308, 3311, 2021, 2463, 3128, 2951, 2787, 953, 973, 2515, 3155, 2853, 1333, 3548, 3718, 774, 1973, 2819, 3098, 3434, 2788, 1719, 2227, 3052, 3274, 2765, 1422, 1835, 2387, 3328, 3303, 1249, 1318]. differential [1035, 2528, 2259, 2449, 1201, 849, 3053, 3618, 3027, 3723, 2791, 3142, 3595, 1393, 3320, 3291, 3511, 3651, 3540, 1988, 2270, 1053]. Differential-Algebraic [2552, 2698, 2700, 2119, 2291, 2416, 2307, 2162, 3254, 2007, 3037, 1918, 2787, 2227, 2259, 3540]. differential-difference [1262]. differential-geometric [2510]. Differentialgleichungen [60, 84]. Differentiation [724, 727, 3602, 3566, 2395, 2483, 3548, 1109, 830, 1166, 812]. Differenzen [84, 60]. difficulties [881]. Diffusion [2746, 3368, 3614, 3445, 3304, 1796, 3392, 3552, 3703, 3161, 3146, 3003, 3226, 3427, 2995, 3325, 3255, 3374, 3677, 3430, 3455, 3688, 3376, 3531, 2801, 3584, 2979, 3188, 3259, 3705, 3399, 3302, 3585, 3366, 3157, 3321, 3536]. diffusions [3627]. diffusive [3072]. digit [1311]. Digital [378, 228, 1660, 103, 3748, 2969]. digraph [825]. digraphs [1954]. Dijkstra [1128]. Dijkstra-like [1128]. dimension [3244, 1227, 3611, 3609, 2999, 3718]. dimension-wise [3244]. Dimensional [3344, 367, 2746, 169, 2554, 2571, 3683, 3131, 967, 2218, 3160, 1921, 1576, 3176, 3186, 759, 3672, 3592, 2357, 1356, 2371, 3258, 2853, 3432, 3286, 3535, 2529, 2125, 2155, 1875, 3491]. dimensionality [2832]. Dimensionering [25]. Dimensioning [25]. Dimensions [224, 2668, 2966, 3443, 886, 2535, 1872, 3018, 2957, 3333, 3489, 3681]. DIMSIMs [2053]. Dinsmore [771]. diode [25]. diodeafkodninger [25]. diodes [1]. dipole [3286]. dipoles [2883]. Dirac [2823, 3186, 3637]. Direct [3437, 3336, 3184, 3211, 1296, 3177, 3374, 1058, 3740, 3711, 3491]. Directed [295, 350, 439, 653, 1361, 1985, 776]. directed-undirected [1985]. Direction [400, 805, 967, 3045, 874, 990, 3000, 1021, 2626, 2277]. direction-of-arrival [2277]. Directional [1880, 3429, 2652, 2432]. Directly [710]. Dirichlet [2580, 3444, 3529, 2533, 3492, 3022, 3424, 3286, 831, 3737]. DIRK [3221]. discipline [1229]. discontinuities [3673]. Discontinuous [702, 3368, 3616, 3697, 3003, 3256, 3456, 3451, 3210, 3211, 3427, 2993, 3542, 3663, 2837, 3667, 3374, 3230, 3592, 3574, 2978, 3705, 3399, 3453, 2529, 3522, 3321, 3710]. Discrepancy [2554, 1828, 3135]. Discrete [525, 2687, 2732, 3096, 790, 2621, 855, 3223, 3525, 2943, 3370, 3369, 1463, 2824, 2495, 3070, 3448, 2441, 3122, 3646, 3582, 1844, 1478, 3265, 3619, 2480, 1861, 3063, 991, 2972, 3387, 3087, 2514, 2609, 2958, 2674, 3438, 3089, 3348, 3728, 3267, 3073, 1235, 3548, 1544, 1715, 3434, 3170, 2984, 3157]. discretisation [2894, 2993, 1246, 1308, 777]. discretised [3312]. Discretization [3640, 2689, 3381, 2201, 2531, 2976, 1055, 1487, 3324, 3253, 1507, 2525, 2814, 1123, 2348, 3268, 3520, 2802, 3622, 2270]. Discretizations [2323, 2615, 2110, 1400, 2365, 3512, 2391, 3072, 3694, 3530, 2864, 2512, 3148, 1080, 2721, 3375, 1416, 2228, 2855]. discretized [3392, 2118, 1184]. discussion [1751]. Disk [1516, 1759, 1482, 1603, 1657, 1922, 1980, 3726, 1705, 958, 1637, 1824]. disks [1376, 1957]. dispersive [3596]. Display [367, 742]. dissimilarity [3608]. dissipation [3327]. dissipative [3668, 3334, 3570, 3169, 3434]. dissipative/conservative [3169]. Dissipativity [2204, 2116]. Distance [2694, 2660, 1067, 1588, 3125, 2912]. Distant [474]. distinct [1733]. Distributed [1910, 1521, 1639, 1698, 1075]. Distribution [376, 504, 525, 1074, 1174, 1085, 1707]. distributional [3448]. distributions [3586, 1909, 2911, 3649]. Distributive [3751, 1226, 1607]. Divergence [2661, 3709, 16, 3078]. Divergence-free [3709, 3078]. divide [2476, 1822, 1709]. divide-and-conquer [2476, 1822, 1709]. Divided [663, 3051, 611, 1038, 2947, 1103]. Division [24, 885]. divisor [883]. Divisors [68, 51, 100, 86, 108]. DNT [3531]. do [1710, 2420]. dogleg [2351]. Domain [2569, 2232, 3527, 3280, 3297, 3071, 2473, 3072, 2240, 2220, 3707, 3177, 3745, 3739, 3486, 2253, 1985]. domain-specific [3072]. domains [2726, 1753, 2909, 3633, 3453]. dominance [1214]. Dominant [446, 3132, 3150, 2878, 2001]. dominated [3703, 3453]. dot [3025]. Double [172, 147, 664, 515, 1311, 3109, 1563, 1607, 2199]. Double-Precision [172, 147]. doubled [1357]. doubled-mantissa [1357]. Doubling [3722]. Doubly [2566, 2502]. Douglas [3304]. Down [346, 330, 1606, 2464]. Downdating [2684, 2145, 2163, 1865, 2156]. downsampling [3313]. drift [3105, 2811, 2837, 3230, 3592, 3705, 3425]. drift-implicit [3105]. drifts [3516]. drive [25]. Driven [721, 2671, 1431, 2775, 3651, 3298]. drivtrin [25]. dropping [3112]. drum [929, 899]. drum-like [899]. Dual [3162, 2750, 3551, 2880, 1872, 2948, 3677, 3076, 2505]. duality [1430]. due [2219, 2248, 2322]. dummy [3311]. Durand [1882, 1565, 1412]. Dvorcuk [2078]. Dynamic [3753, 918, 1191, 726, 1911, 709, 1250, 2416, 2081, 2118, 2453, 1998, 1908, 2040, 1956, 807, 1228, 1594, 3390, 1559]. Dynamical [2697, 3615, 2384, 3689, 3553, 3580, 3539, 3620, 3743, 3124, 3481, 2974, 3695, 2956]. dynamically [788]. dynamics [3599, 2644, 2811, 3385, 2268, 2355, 3721, 3625, 3251]. dynamization [1638].

E.A.D.I. [874, 1063]. each [919]. Early [1116]. ECM [40]. Economical [2862, 957]. edge [3002, 3049, 3110, 3180, 1843]. edges [1967]. Editor [3317]. Editor-in-Chief [3317]. Editorial [2599, 2628, 2641, 2657, 2715, 3241, 2869, 2560, 2576, 2769, 2808, 2821, 2833, 2941, 2952, 3001, 3055, 2613, 2614]. EDP [10]. Education [42, 180]. Effect [1969, 448, 3060, 1784, 2915, 980, 3366]. Effective [1976, 3600, 1970, 1315, 1004]. Efficiency [655, 195, 555, 1686, 2304, 2097, 935, 765]. Efficient [1326, 1523, 1887, 1425, 2642, 3045, 3586, 3197, 1398, 1371, 2958, 319, 3713, 1088, 3376, 3728, 3505, 3267, 3532, 1954, 3313, 2374, 1758, 3303, 1626, 1720, 3077, 3562, 1053, 1961, 1226, 2965, 3132, 2860, 2592, 899, 3263, 931, 905, 841, 1999, 3176, 3621, 1122, 1832, 3216, 887, 3561, 3378, 2929, 3610, 2820, 2314, 1877, 3115, 3524, 3725]. Efficiently [2015, 3521]. eigen [2178]. eigen-like [2178]. eigendecomposition [2676]. eigenpairs [3372, 3476]. eigenproblems [3174, 2134, 2185]. eigensolvers [3384]. Eigenvalue [3056, 2547, 2434, 3127, 2063, 2669, 627, 168, 1399, 2578, 3118, 2494, 2629, 3436, 3279, 3420, 2177, 2430, 3421, 2508, 3061, 2394, 2864, 2405, 3667, 3111, 3397, 3733, 3629, 3702, 3645, 3097, 3338, 2878, 971, 1991, 2925, 809, 890, 2725, 2009, 3233, 3680, 2183, 2184, 2501, 2124, 1423, 2640, 3041, 3222, 3337, 2049, 3610, 3691, 3469, 3205, 3550, 3470]. Eigenvalues [2264, 525, 530, 121, 135, 218, 385, 708, 1528, 1661, 2823, 3117, 2173, 1927, 2379, 1125, 1062, 1320, 2861, 930, 2295, 2476, 2826, 2076, 2316, 2421, 2089, 1386, 2182, 2129]. Eigenvector [2598, 3436, 3337]. eigenvector-dependent [3436]. eigenvectors [2295, 2476, 2653, 121, 135, 3041]. Eigenwertprobleme [168]. eight [1212]. eighth [1477, 858]. eikonal [3604]. Einstein [3509]. Elastic [103]. elasticity [1790, 3136, 3610]. elastodynamics [3228]. election [15]. electromagnetic [3172]. electromagnetism [3133]. Electronic [118, 367, 23, 2183]. electronics [92]. Electrostatic [2747]. electrostatics [2848, 3014]. elektronikkomponenter [92]. Element [2693, 2744, 2707, 2734, 477, 2750, 3683, 1661, 3587, 3195, 2218, 2428, 3044, 1792, 2726, 3173, 3528, 3213, 3004, 3263, 3307, 3367, 3736, 1172, 2906, 3658, 2113, 3709, 3325, 2864, 2896, 3449, 2149, 3396, 3693, 3657, 2785, 3672, 3049, 3179, 2957, 3180, 3088, 1790, 1794, 3020, 3089, 3167, 3677, 2243, 3138, 1300, 2814, 3645, 3360, 3653, 2866, 2937, 3326, 2359, 2867, 3268, 2021, 1179, 2029, 3698, 2245, 3098, 2961, 2841, 2918, 2624, 2831, 3222, 3510, 3666, 3470]. element-based [2961]. element-singly [2021]. Elementary [33, 81, 1059, 2122, 776]. Elements [264, 351, 2689, 1468, 2995, 3110, 3136, 2359, 3342, 2845]. Elimination [760, 460, 249, 2094, 2496, 2368, 1846, 2451, 1471, 149]. Ellipses [2329, 2098]. ellipsoidal [2540]. Elliptic [2, 2663, 475, 2689, 967, 2428, 3382, 284, 2054, 990, 3542, 3017, 1063, 2785, 3180, 933, 2634, 2851, 3745, 3739, 3113, 3475, 3231, 2867, 3268, 3584, 2253, 3723]. elliptic-parabolic [3475]. Embedded [1419, 2126, 3575, 1670, 3072, 3574, 1682]. Embeddings [1827, 2635, 1810]. Empirical [44, 3728]. employing [1313, 1445]. empty [816, 1649]. Enclosed [138]. enclosing [1961]. Enclosure [2562, 3413]. enclosures [3056]. encoding [921]. Encroaching [1708]. End [1074, 1523, 1857]. endpoints [3066]. Energy [2685, 2811, 2664, 3385, 3334, 3616, 3738, 2967, 3047, 3570, 3572, 3154, 3169, 2152, 3521, 2985, 3524]. energy-based [3616]. energy-preserving [2967, 3154, 3521]. Engquist [2719]. Enhanced [3572, 3538, 2028]. enhancement [3663]. Enhances [1913]. Enhancing [2152]. enlarging [2231]. enrichment [3736]. Enterprise [173]. entries [2344, 2316, 2129]. Entropy [3290, 3468, 3228, 3598, 3656, 3415]. Entrywise [2127]. Enumeration [704, 3607, 1305, 1444, 1504]. envelope [2234]. envelopes [2410]. Environment [709, 461, 1965]. Environments [1118, 1278, 1259, 936]. equal [2466]. Equality [3754, 2321, 3242, 2579, 3126, 936, 3601, 2258, 2351]. equals [3223]. Equation [580, 702, 443, 313, 2746, 2663, 2549, 341, 388, 2401, 2325, 2699, 2569, 2668, 268, 607, 426, 527, 2557, 2671, 2703, 2559, 3445, 21, 3616, 2600, 3538, 3587, 2293, 1927, 3160, 3419, 3586, 3554, 3458, 3426, 904, 1078, 3496, 989, 3058, 3654, 3596, 3146, 3738, 3280, 3046, 3297, 3060, 3514, 1749, 3004, 3529, 3443, 2520, 2442, 1893, 1176, 2994, 2723, 3530, 2535, 3324, 2896, 902, 3186, 3178, 3589, 3501, 2935, 2849, 3487, 844, 3347, 3637, 2107, 2675, 3648, 3517, 2886, 60, 3348, 3138, 1300, 3376, 1416, 3153, 3009, 3312, 3259, 3217, 3023, 2999, 2489, 3424, 3098, 3400, 2623, 3533]. equation [1850, 3365, 3534, 3328, 2855, 2301, 2624, 2125, 2155, 1019, 3689, 849, 3101, 3378, 3711, 3575, 2831, 3014, 3321, 1393, 3524, 3536, 1302]. equational [2042]. Equations [700, 71, 116, 1799, 701, 2319, 2743, 582, 487, 470, 2643, 597, 2661, 2745, 352, 418, 740, 2685, 693, 203, 227, 279, 404, 639, 656, 399, 275, 509, 620, 744, 640, 536, 2458, 1913, 657, 329, 342, 354, 2334, 319, 2552, 475, 491, 2688, 689, 2555, 674, 459, 2710, 604, 2700, 185, 390, 428, 2711, 2735, 240, 198, 2736, 249, 289, 270, 323, 592, 2702, 2558, 164, 2433, 1798, 3105, 3368, 2577, 3684, 3043, 1961, 3293, 1313, 1445, 966, 2119, 1869, 3131, 3304, 2279, 2942, 3159, 3295, 3195, 1400, 1436, 3392, 2218, 3068, 3182, 3477, 3597, 2391, 1262]. equations [3371, 988, 1697, 2081, 2507, 1027, 3602, 284, 2393, 2161, 2201, 2470, 2953, 3161, 897, 2246, 2616, 2033, 2429, 2651, 2103, 2281, 3526, 3069, 2307, 2795, 2672, 2905, 2162, 3437, 2726, 3623, 969, 1162, 1336, 3473, 3697, 2054, 2914, 2810, 3003, 3059, 1459, 1252, 2862, 1283, 1629, 1592, 3163, 3307, 3367, 912, 3474, 3499, 3579, 1385, 2944, 3706, 3210, 3480, 2955, 1055, 1439, 2521, 1487, 3613, 2874, 3427, 3658, 2381, 3567, 546, 2608, 3707, 1101, 3699, 817, 886, 990, 3694, 3730, 1057, 841, 3663, 2773, 2757, 1184, 1231, 3748, 902, 1524, 3395, 1662, 3254, 1232, 1699, 1732]. equations [2408, 2479, 2837, 2070, 2511, 1164, 3148, 3283, 3537, 3502, 2775, 3672, 2333, 2759, 2875, 2956, 3615, 3620, 1507, 2346, 2799, 1080, 2133, 2513, 3255, 2058, 1058, 1284, 1144, 2286, 2335, 2525, 2646, 2916, 1016, 1442, 2634, 2674, 2721, 2885, 2897, 3033, 1989, 3516, 84, 1790, 2299, 1178, 861, 3570, 2876, 3034, 3151, 2584, 2898, 2205, 3095, 2071, 2198, 3020, 3089, 2948, 60, 1647, 3096, 2088, 1032, 1102, 1149, 3037, 3483, 3333, 3572, 3592, 2866, 2959, 3486, 3573, 3688, 3285, 3650, 2357, 2137, 2151, 2385, 3531, 1266, 863, 1352, 1714, 1213, 2676, 2500, 2411, 1594, 1918, 2348, 3322, 3679, 1124, 2060]. equations [2061, 1595, 1246, 1308, 3311, 2289, 3022, 2924, 3169, 3330, 2021, 2463, 2802, 3128, 3481, 2951, 2787, 2889, 1443, 953, 973, 975, 3188, 2980, 3259, 2852, 3155, 2226, 1151, 3698, 1333, 1233, 3548, 3718, 774, 1818, 2819, 1493, 1254, 1301, 3434, 2788, 1719, 3052, 3274, 2765, 1422, 847, 2211, 2962, 2637, 3290, 2196, 2387, 3303, 812, 1249, 1318, 1035, 2528, 3193, 2259, 1201, 1462, 1020, 2984, 3053, 3220, 3275, 3618, 3027, 3723, 3468, 1167, 2791, 3157, 3077, 2276, 2505, 2820, 2589, 3142, 3595, 149, 3666, 3320, 2260, 3291, 3511, 3651, 3710, 2426, 1988, 2270, 1053, 1106, 294]. Equations [3638]. Equi [474]. Equi-Distant [474]. equidistant [3221]. equilateral [788]. Equilibration [772]. equilibria [2630, 2249]. Equilibrium [2290, 3255]. Equipment [85]. equispaced [3632]. Equivalence [3129, 1622, 1183, 1923, 1409, 1773, 1210, 1841]. Equivalent [689, 946, 2635]. Eratosthenes [838]. Erdos [432]. EREW [1878]. Erik [2064]. Errata [19, 98, 520, 1202, 1651]. Erratum [133, 282, 1290, 1396, 637, 1398, 1629, 3367, 2655, 327, 81, 562, 394, 1308, 3270, 135, 890, 1630, 543, 465, 1964, 1338]. Error [453, 1811, 3159, 1400, 2467, 3495, 3623, 177, 2320, 3755, 3071, 1385, 3567, 2394, 3264, 1163, 307, 529, 1699, 1732, 1930, 2481, 2568, 2409, 1562, 2675, 445, 549, 1537, 2698, 3331, 723, 2008, 1123, 1315, 530, 1660, 1849, 428, 747, 1628, 162, 588, 3025, 2960, 810, 831, 2341, 2228, 248, 3365, 2712, 259, 571, 386, 1462, 628, 3720, 2375, 3683, 1523, 2342, 3493, 3551, 3418, 1869, 2159, 3344, 1608, 2942, 2965, 2988, 1085, 1945, 1892, 2278, 1160, 2454, 3383, 2630, 3527, 1268, 3213, 2331, 3393, 787, 2955, 982, 1487, 1038, 3427, 1577, 2163, 1894, 1751, 2404, 1881]. error [2369, 1547, 3742, 3462, 2241, 2477, 2970, 758, 3178, 2482, 2070, 2512, 3487, 1847, 2006, 2268, 3374, 2106, 2525, 2916, 2634, 3442, 1265, 3725, 3713, 3165, 917, 3215, 2949, 3677, 1862, 1303, 3645, 1311, 2866, 1814, 773, 2801, 3097, 3338, 2586, 1246, 1308, 2677, 1735, 2978, 1353, 762, 1509, 2926, 3257, 3705, 3548, 3432, 3219, 3235, 3098, 1066, 2789, 2890, 3722, 1198, 1255, 3012, 1392, 3239, 1130, 1835, 2447, 777, 3402, 3522, 3676, 2588, 3205, 2260, 2109, 2845]. Error-free [1163]. Errors [685, 366, 67, 3749, 2445, 689, 706, 43, 3262, 3477, 1055, 842, 3680, 980, 1932]. essentially [2886]. establish [1210]. Estimate [2467, 428, 747, 1751, 1547, 2978, 3219, 2960, 3365, 2109]. Estimates [2518, 2744, 2707, 530, 3231, 588, 571, 1523, 2578, 3551, 1869, 3344, 1400, 2428, 3588, 1160, 3459, 3427, 3567, 3123, 1881, 3264, 2477, 2970, 1930, 2512, 3487, 2916, 3165, 2949, 3677, 3331, 1303, 3645, 2866, 3097, 1897, 2677, 1849, 3431, 3548, 2635, 2789, 2890, 810, 831, 2841, 2624, 3522, 2588, 3353, 2068]. Estimating [366, 982, 1345, 1637, 3239, 3296]. Estimation [2354, 1730, 2706, 2251, 2698, 308, 706, 476, 162, 2712, 431, 3159, 1945, 2454, 3495, 3623, 2308, 2772, 2509, 2202, 1731, 875, 2484, 1975, 1814, 1123, 2586, 2272, 3313, 3433, 3025, 1066]. Estimations [707]. estimator [2965, 1645, 2540]. estimators [3519, 3584]. Euclidean [3223, 3116, 1276, 1896, 1141]. Euler [3294, 2144, 3082, 2362, 3480, 3670, 3567, 2837, 3148, 2850, 1299, 3166, 2325, 1647, 3650, 3594, 2802, 3188, 2199, 1963, 1757, 3290, 3468, 3651]. Euler-integral [1299]. Eulerian [2442]. Evalquote [507]. Evaluate [465, 415]. evaluating [2481, 2950, 1756]. Evaluation [2683, 1828, 904, 618, 65, 3346, 3358, 427, 712, 146, 493, 401, 606, 938, 1034, 515, 156, 3685, 3380, 2893, 1576, 2282, 2083, 1611, 1851, 2935, 3408, 1774, 1895, 1469, 1563, 1000, 3607, 3506, 1217, 3363, 3414, 880, 903, 3012, 882, 1018, 1312]. Evans [2683]. EVD [3390]. even [2920, 3345, 1108]. Event [721, 3211]. Event-Driven [721]. events [1243]. Evidence [736]. evolution [2942, 3295, 1400, 2293, 2894, 2993, 3283, 3537, 3502, 3020, 3089, 2411, 3169, 3379]. evolutionary [3623, 2955, 2889]. Exact [3384, 3461, 2131, 3718, 1749, 2955, 932, 3395, 3503, 3645, 3153, 3329, 1866]. Exactness [538, 3606]. examined [102]. Examples [2664, 2485, 368]. exchange [1981]. excluding [1952]. exclusion [2962]. Executed [710]. Execution [2013]. exist [2907]. existence [1459, 1283, 1439, 1184, 2056]. Exits [1538]. expanded [3611, 3609]. Expansion [2440, 548, 290, 1969, 3363, 2392, 1335, 2189, 2070, 773, 2677, 3313, 1152]. Expansions [1142, 658, 2964, 2106, 2899, 3271, 831]. Expected [1982, 1959, 1362, 1470]. Experience [58, 1483]. Experiment [473, 512]. experimental [1315]. Experiments [786, 2280, 1057, 2022]. explanation [3628]. Explicit [1926, 3490, 307, 341, 2709, 474, 2647, 3410, 2649, 3350, 2751, 569, 3024, 2126, 3054, 3078, 3417, 2577, 3604, 3614, 3690, 2291, 1511, 1477, 3654, 1904, 1195, 1355, 1505, 3213, 2944, 2757, 1039, 958, 861, 3359, 3284, 2289, 1864, 1906, 3585, 2779, 1019]. Exploiting [2732, 3279]. exponent [793]. Exponential [3093, 2686, 2355, 3006, 2409, 2751, 1295, 3240, 1640, 1946, 3121, 3383, 3479, 1194, 3641, 3059, 3735, 2955, 2992, 3699, 3149, 827, 3165, 3166, 845, 3486, 2950, 2776, 2411, 3505, 3398, 3532, 2654, 846, 1884, 2803, 2465, 3318, 1820, 3534, 3676, 869, 3054, 3575, 3595, 3353]. exponential-type [2950, 3575]. Exponentially [675, 1475, 3154]. exponentially-fitted [3154]. exponentials [3565, 3442]. Exponents [511, 746]. Expressing [621]. expression [156]. Expressions [517, 193, 307, 2397, 493, 465, 415, 310, 207, 1735]. Extended [1331, 3114, 155, 3414, 1194, 3212, 2395, 3423, 1166, 1019, 3521]. Extendible [3753, 1192, 1211]. Extending [2186].

Extension [551, 592, 2004, 1816, 3465, 1021]. Extensions [2383, 2274, 1119, 2167, 3239]. exterior [2853]. External [709, 2190]. extra [3399]. extracting [3476]. Extraction [2666, 247, 2542]. Extraneous [2573]. Extrapolated [400, 2698, 2488, 805, 967, 874, 1386, 1301]. Extrapolation [1142, 444, 413, 571, 1086, 1730, 2189, 2055, 993, 1196, 2960, 1834, 2154, 1152]. extrapolations [1771]. extrapolative [1816]. Extreme [704, 2153, 1444]. Ezer [3626].

F. [2452]. Fàa [550]. Facit [58, 40].

Facit-Algol [58]. factor

[3008, 2206, 763, 832, 1490]. factored [1416]. Factorials [22]. Factoring [3086, 2827]. Factorised [2448]. Factorization [1799, 2662, 22, 1913, 730, 1558, 2738, 2342, 1796, 2418, 1077, 2294, 1296, 1297, 2457, 2114, 2784, 914, 2149, 2883, 2347, 2484, 2398, 972, 1786, 2679, 1438, 2374, 2982, 2901, 1907, 2625]. Factorizations [1539, 1785, 2052, 1797, 2595, 2360, 1901]. factorizing [1088]. Factors [2403, 320, 334, 1785, 2237]. Fadeev [2157]. fail [2309]. failure [3144, 1877]. Fair [2257, 2169]. Falsi} [426, 580]. families [906, 1035]. Family [2324, 2573, 458, 2749, 1978, 897, 2970, 1227, 957, 808, 2542, 1949, 3412, 3090, 1151, 1420, 3053, 2845]. Farey [1277]. Farm [85]. Fast [3117, 3642, 2034, 2795, 2366, 1635, 3209, 2946, 886, 3753, 76, 2058, 2459, 1958, 3581, 2385, 1213, 2423, 2668, 2927, 2350, 2740, 2353, 3588, 2893, 3058, 2066, 2921, 3474, 1829, 1706, 2298, 2311, 3471, 1416, 2462, 879, 1471, 3647, 2999, 2168, 3585, 813, 1083, 2807, 3491, 2532]. fast-oscillating [2462]. Faster [3221, 2607, 3490, 2474, 1905]. fault [1877]. fault-tolerant [1877]. FCFS [1965]. FCFS-scheduling [1965]. FCV [1227]. FDE [3588]. FE [1793]. feasibility [1830]. feature [3721]. Features [267, 2496, 2431, 2790]. Fejér [1899, 3013, 3102, 2740]. Fekete [3715]. FEM [2801, 2855]. Fermi [88, 89]. Ferromagnetic [2642]. FETI [2243]. few [2173, 930]. Feynman [2580, 3217]. FFT [2004, 2447, 3622]. FFT-based [3622]. Fibonacci [176, 555]. fibre [3553, 3174]. fidelity [3727, 3571]. Field [2174, 3479, 3473, 3360]. fields [3224, 1896, 3662, 2649, 763, 832, 3078]. Fifth [1442, 236, 458, 2112, 1151]. Fifth-Order [458, 1151]. fifty [1095]. File [398, 1074, 673, 8, 14, 1840, 1456, 726, 1767, 1376, 1483, 294]. FILE-68 [294]. Files [175, 3753, 1759, 247, 1482, 1998, 1922, 1516, 1030, 1345, 1705, 1851, 2026, 1228, 811, 868]. fill [1970]. Filon [2858, 1577, 3332, 3398, 1849]. Filon/asymptotic [2858]. Filter [300, 2256, 3164]. filtered [3650]. filtering [2534]. filters [3642, 3357, 3258]. finally [1625]. financial [3082]. find [2232]. finder [2868]. finders [843]. Finding [2094, 1955, 2324, 603, 2166, 552, 1759, 1141, 557, 1891, 2505, 1889, 1482, 1769, 1669, 867, 1468, 1882, 1565, 1744, 1412, 2078, 2027, 1587, 1990, 1902, 3169, 1130, 2001, 2962, 2820, 2277]. Fine [3588]. Finite [2693, 2695, 2744, 969, 740, 2633, 1893, 2707, 2709, 712, 3088, 2410, 2734, 2937, 2434, 3326, 1863, 2689, 477, 2750, 2399, 163, 387, 2703, 3683, 1399, 1661, 2218, 2428, 3044, 1792, 2176, 2726, 858, 839, 3297, 3213, 3060, 3004, 3263, 3736, 3443, 1056, 2442, 2906, 3658, 3686, 2113, 3709, 2995, 3325, 2864, 2896, 3449, 3450, 932, 3396, 3693, 3657, 2785, 3672, 2957, 957, 3255, 1441, 2721, 1790, 1794, 1178, 3165, 2898, 3020, 3089, 3167, 3677, 3138, 1300, 2814, 3645, 3360, 3653, 2866, 2959, 3268, 2021, 1179, 1165, 1233, 3432, 3424, 2245, 3098, 3400, 2961, 1372, 2841, 3365, 2918]. finite [1249, 2624, 2448, 3220, 3523, 2831, 3222, 3622, 3666, 3470, 3342].

Finite-Difference [740, 387, 3622]. finite-element [2624]. Finite-Part [2633, 2937, 1165]. finitely [3066]. First [656, 640, 2647, 459, 1445, 966, 3586, 3058, 3495, 3541, 2994, 1771, 914, 3620, 1058, 1410, 1411, 3032, 861, 3096, 1448, 1595, 1386, 763, 832, 2226, 3639, 812, 1201, 1167, 1522, 3379]. first-order [3620, 3032, 3379]. first-passage [3586]. Fisher [3030]. Fit [211, 238, 2673]. Fitted [675, 2920, 2778, 2838, 3229, 3154, 1885]. Fitting [2694, 742, 2329, 3045, 1194, 3393, 3593, 2098, 2432, 2407, 2910, 3740, 137, 2197, 3054]. FitzHugh [3514]. Five [2597, 3062]. fix [1543]. fix-point [1543]. Fixed [2573, 24, 1085, 898, 758, 1392]. fixed-point [758, 1392]. Fixed-Points [2573]. Flexible [3422, 568, 3668, 3249, 3218]. flipped [3411]. Floating [27, 685, 12, 28, 127, 643, 706, 147, 386, 3642, 3687, 3358, 3408, 3331, 3258, 3025, 3235]. Floating-Point [685, 127, 643, 706, 386, 3642, 3687, 3358, 3408, 3331, 3258, 3025, 3235]. Floquet [2057]. Flow [365, 735, 66, 585, 299, 341, 3496, 1795, 3709, 2936, 3267, 1179, 980, 2902]. Flow-Mode [365]. Flowchart [397]. flows [3334, 3466]. fluid [1795, 3744, 3251]. flux [3619, 3688]. focusing [2332]. Fokker [3586, 2723]. folding [1376]. folkbokföring [10]. Following [1251, 2886]. forces [3298]. Ford [1429]. Ford-Johnson [1429]. forest [2954]. Forests [1330]. Form [438, 611, 270, 310, 2353, 591, 3197, 1789, 241, 3568, 2865, 1587, 996, 2423, 974, 1217, 2779, 2171, 212]. formal [2494]. formalisms [1681]. Formalization [1260, 2062]. Format [197, 110, 3131, 3543]. formatted [1030]. Forming [384]. Forms [320, 569, 1621, 1641, 3448, 1976, 3063, 1949, 2358, 3169, 2168, 1716]. Formula [1142, 2362, 165, 2450, 343, 550, 430, 417, 2307, 3406, 1038, 1351, 1284, 2459, 3504, 3217, 2465, 1288, 1628, 3548, 2199, 2829, 2960, 1255, 3012, 3203, 2805, 1248]. Formulae [694, 2665, 3439, 3602, 3576, 1670, 1710, 826, 1577, 3384, 1298, 1912, 2498, 1184, 2882, 2462, 2516, 2926, 3257, 3351, 1316, 1849, 1978, 850, 810, 812, 1368, 1626]. Formulas [424, 470, 583, 618, 739, 2322, 670, 585, 601, 445, 602, 344, 240, 1945, 2904, 3626, 1350, 3726, 1576, 2219, 2248, 2083, 2097, 1671, 2395, 3617, 876, 1089, 1121, 759, 1411, 3746, 2483, 1917, 2073, 3301, 1580, 1034, 1104, 1166, 1199, 1018]. formulation [3482, 3096, 3348, 3666]. Forsythe [3660]. FORTRAN [340, 507, 287]. found [1625]. four [2895, 1885, 1460]. four-step [1885]. Fourier [2713, 2532, 2792, 2892, 2044, 504, 3161, 2066, 2004, 298, 2746, 1478, 3463, 3518, 2196, 2075, 1083]. Fourth [319, 2411, 2703, 1927, 1125, 1162, 1320, 1642, 1252, 1663, 808, 1587, 1626, 849, 3378]. fourth-order [3378]. FP [2943]. Fractal [3106, 2141, 3447]. Fraction [629, 229, 1266]. Fractional [2972, 3440, 3585, 3368, 3392, 3552, 3382, 3161, 3641, 3437, 3146, 3697, 3003, 3210, 3482, 3450, 2132, 3374, 3628, 3478, 3573, 3376, 3531, 3679, 919, 952, 1000, 3465, 1548, 3399, 1322, 3289, 3535, 3220, 3315, 3378, 3618, 3142, 3321, 3562, 3524, 3536, 3651]. fractional-derivative [3289]. fractional-in-space [3161]. Fractional-order [3440, 3585]. fractions [955, 1277, 1596, 1651, 1269]. frame [3706]. Frames [2326, 2812]. framework [2977, 2113, 3266, 2298, 2311, 3169, 2974]. Frank [3508]. Fredholm [470, 3163, 3730, 2070, 491, 1213, 2980, 1301]. Free [2424, 2556, 3565, 2772, 705, 1296, 1297, 1346, 3709, 1163, 1508, 2041, 3309, 1363, 1485, 1210, 2209, 2123, 3078]. Frequencies [398]. Frequency [654, 2685, 586, 995, 2668, 3177, 3361, 2999]. frequency-domain [3177]. Fresnel [62, 63].

Freudenthal [340]. Friedrichs [3448]. Frisch [416]. Frobenius [2180, 1728]. Fröberg [2064]. Front [2936, 2113, 3234, 3354]. front-tracking [3234, 3354]. frontal [2067]. frozen [2924]. Full [2835, 623, 2574, 3040]. Fully [3646, 2406, 3063, 3548, 1737, 3614, 3357, 3267, 3053, 3563, 3089]. Function [2683, 49, 663, 376, 504, 703, 639, 656, 286, 2701, 600, 672, 308, 157, 358, 431, 417, 3380, 88, 2880, 2844, 1640, 1946, 2724, 3405, 1792, 3070, 806, 2977, 1404, 89, 1253, 1593, 3617, 336, 760, 934, 1733, 3167, 3688, 2336, 1578, 1563, 846, 1884, 1317, 1866, 2302, 3287, 3012, 3203, 156, 765, 869, 3106]. Functional [701, 2743, 657, 1686, 2296, 1178]. functional-differential [1178]. Functionally [3229, 2778, 2838]. functionally-fitted [2778]. Functionals [2687, 1844, 3570, 3073, 792, 3452, 3272, 3549]. Functions [48, 617, 366, 425, 526, 720, 33, 81, 2582, 2708, 548, 658, 427, 712, 290, 2573, 2356, 1969, 477, 314, 676, 725, 447, 2446, 692, 2328, 626, 44, 403, 206, 1435, 3729, 2344, 1770, 2989, 2893, 2932, 3459, 3447, 3437, 1268, 2440, 2282, 3071, 3664, 3732, 1750, 546, 1948, 3686, 1478, 1541, 3264, 2443, 2498, 842, 3621, 2114, 3048, 3094, 3062, 2132, 3386, 1492, 1441, 1895, 1040, 3442, 2085, 3725, 3713, 1527, 3036, 2499, 3139, 2950, 2122, 2899, 3494, 2462, 1612, 3051, 3361, 2926, 2073, 3090, 937, 1033, 792, 1815, 1579, 3301, 3643, 3431, 3680, 3414, 3415]. functions [3440, 2153, 864, 2529, 3454, 2197, 3102, 2171, 3275, 2929, 2807, 1458]. functors [974]. Fundamental [2696, 2739, 199, 1212, 1798, 3002]. Further [876, 1089, 2154, 3402, 3749].

G [2884]. G-matrix [2884]. G. [547].

Gaffney [1738]. Galerkin

[3368, 3616, 2218, 3382, 3697, 2810, 3003, 3173, 3210, 2906, 3427, 2993, 3542, 3663, 3667, 3374, 2949, 2699, 3662, 3572, 3475, 3574, 3169, 2226, 3399, 3453, 3435, 1462, 3522, 2671, 3321, 2986, 3470, 3710]. Games [199]. gamma [3405, 983, 1253]. Gander [2749]. gap [3416, 1270, 1338]. Garbage [669, 223, 813]. Garey [1102]. GARK [3569]. Garwick [1780, 814]. Gas [196, 2936]. gauge [2914, 3710]. Gauss [3045, 1369, 2265, 2581, 618, 1268, 1350, 1576, 2219, 2248, 2322, 3356, 2665, 1184, 1231, 2748, 3428, 3036, 492, 2677, 2926, 3301, 3272, 1017, 831, 1198, 2805, 882, 3469]. Gaussian [3576, 663, 2496, 3664, 2368, 1298, 2498, 1526, 2106, 3428, 3036, 3662, 2451, 460, 3351, 2352, 3239, 1778, 149]. GCV [3335, 3140]. Gear [982]. Gegenbauer [3736, 3271, 1198, 3012, 3203]. General [3355, 2429, 2591, 2730, 2565, 2190, 128, 213, 623, 2751, 370, 1802, 1399, 1078, 1590, 1591, 2033, 2651, 3278, 2039, 3072, 2131, 2884, 2513, 1713, 1414, 1794, 3246, 1416, 3169, 1693, 3362, 1787, 2302, 3302, 3440, 2047, 2227, 1229, 3731]. Generalisation [244, 1706]. Generalised [629, 856, 3670, 1269]. Generalization [288, 1161, 2798, 3332, 1917, 3498, 1124, 2154]. Generalized [523, 3183, 1946, 2665, 211, 2550, 993, 394, 355, 154, 1311, 2618, 675, 1214, 3301, 3272, 1545, 1461, 1474, 2792, 3717, 2794, 939, 3587, 3118, 3426, 3243, 2631, 2872, 1431, 3421, 1264, 1404, 1327, 1310, 2443, 2405, 1196, 3308, 2011, 1040, 1414, 3246, 1527, 3429, 3518, 2020, 2287, 951, 1285, 1107, 3249, 3336, 2925, 2678, 2217, 2185, 3534, 1573, 2277, 2713]. Generated [741, 3044, 2085, 1210, 2153]. Generating [1049, 1803, 1500, 705, 1571, 2756, 1690, 1873, 382, 2328, 2632, 2514, 1485, 1841, 3454, 3275]. Generation [2403, 228, 136, 170, 356, 69, 1281, 1536, 2113, 1726, 1854, 1727, 1366]. Generator [543, 513]. Generators [402, 2330, 1711, 2622]. generic [3127]. genetic [3425]. GENIUS [128]. Geodesics [2472]. Geometric [2642, 3119, 3631, 2708, 2479, 2813, 3010, 2718, 2507, 2921, 1430, 2510, 2367, 2087, 3640, 2776, 2649, 3469]. Geometrically [3464]. geometries [3116, 3366]. German [168, 84, 60, 148]. Gershgorin [21]. Gibbs [2964]. GIER [94, 124, 76, 94, 24, 87, 91, 267]. Ginzburg [2886, 3378]. Given [535, 556, 1904, 1783]. Givens [2610]. GKB [2943]. GKB-FP [2943]. gleichungen [60]. GLMs [3362]. Global [116, 902, 2970, 2698, 2499, 1066, 1198, 1255, 3294, 2456, 1860, 1012, 982, 3730, 3617, 2525, 2586]. Globally [725, 1182, 948]. GLT [3564]. GMRES [2305, 2604, 2495, 2188, 2473, 2120, 2863, 3422, 2242, 2397, 2815, 2619, 3050, 3039, 3193, 2930]. Goldbach [718, 1129, 184, 278, 139]. Golden [176]. Golub [3525, 2915]. good [95]. GOTO [764, 95]. Goursat [462]. governed [2894]. GPS [2605]. GPU [3011]. GPU-based [3011]. Graded [2743]. Gradient [2690, 2417, 1781, 2391, 2104, 1561, 2178, 3490, 3582, 1795, 2956, 3309, 1790, 1665, 1794, 3334, 3138, 3636, 1418, 3560, 3452, 3434]. gradient-like [2178]. Gradients [2467, 3733, 532, 2712, 1784, 3187, 2866, 3470]. Grading [129, 3600]. Gram [453, 2684, 222, 2114, 2242, 2130, 3129, 2588, 2156]. Grammar [974, 26]. Grammars [625, 610, 789, 1469, 1485, 974, 1245, 1842, 1312, 1005]. grammatiken [26]. granular [2650]. Graph [488, 2035, 645, 495, 217, 1639, 1383, 2232, 2146, 1635, 2104, 1822, 2027, 776]. Graphic [742]. Graphics [570]. Graphs [439, 646, 1887, 1425, 2094, 3552, 1517, 1361, 1967, 3745, 3739, 1001, 1380, 1810, 1843, 2780, 1986, 2030]. Grassmann [2928]. greatest [883]. Greedy [1636, 3717, 2235, 1958, 2001, 3723]. Green [2805]. Gregory [399]. Greibach [974]. Grid [3554, 1456, 3183, 1483, 1636, 1659]. grids [2239, 3658, 1351, 2849, 2243, 2828, 3391, 3238, 3221, 3510]. Grossman [2427]. ground [3509]. grounded [2035]. Group [340, 2548, 359, 2553, 3565, 2592, 2404, 1126, 1542, 2436]. Grouping [673]. Groups [2590, 3661, 1854, 3577, 3578, 2271, 2372, 2853]. Growth [2568, 2026, 3656, 763, 832, 2228, 2985]. Grünwald [3573]. GSVD [1772]. Guaranteed [787, 1750]. guided [3175]. guides [3175]. Gunilla [3317]. gyroscopic [3612].

Haar [2250, 2284, 2352]. Hadamard [1183]. Half [2291, 58, 672, 3344, 1585, 1912]. Half-explicit [2291]. half-infinite [1912]. Half-Period [672]. half-plane [1585]. Hall [3060]. Hamblin [310]. Hamiltonian [3418, 2921, 3198, 3668, 3394, 3387, 3327, 3154, 3010, 3350, 3362, 1718, 3364, 3521]. Hamming [1447]. Handling [110, 1073]. Hankel [3300, 2361]. Hanoi [3457, 1327, 1310, 1455]. hard [1965]. Harmonic [2666, 676, 387, 1350, 3658]. Hartley [2085]. Hash [1969, 1435]. Hashing [729, 1012, 3753, 558, 918, 1052, 1190, 1473, 1840, 1192, 1382, 1826, 662, 1998, 1725, 2040]. Having [2328, 1733, 1991, 2302, 3302]. head [1637]. headed [1824]. Heap [396, 1172, 2025, 2029]. heaps [1344, 1604, 1748, 1982]. Heapsort [1570]. Heat [341, 388, 183, 138, 196, 2557, 3477, 1893, 2535, 2896, 3627, 2802, 1179, 2489, 3533, 3101]. Heating [2693, 2957, 3088, 3326]. Height [1875, 2094, 2029]. Height-analysis [1875]. height-balanced [1875]. heights [1982]. Heiki [1920]. Hellgren [517]. Helmholtz [3058, 3178, 3648, 2569, 2668, 2999]. hemivariational [3657]. Henriksen [1518]. Hensel [1277]. Herlestam [1463, 1002]. Hermite [2044, 3277, 2893, 3106, 3162, 3736, 2005, 1038, 3084, 2443, 3428, 3518, 3036, 996, 2777, 675, 2422, 2817, 2911, 624, 569, 1899, 588, 3712, 3630]. Hermitian [1782, 2603, 1928, 3338, 3249, 2925, 3390, 3747, 2009, 3431, 2124, 2639, 3403]. Hessenberg [3197, 441, 3568, 2865, 3414]. Hessenberg-triangular [2865]. Hessian [3503]. heterogeneous [3004, 3646, 3410, 1910]. Heuristic [2141, 1767, 448, 2456, 2024, 1072, 2001]. hidden [1471, 3536]. hidden-memory [3536]. Hierarchical [3081, 3131, 3277, 3070]. hierarchically [1790, 3608]. High [2794, 580, 582, 2507, 945, 1230, 471, 2944, 297, 3589, 2907, 3283, 83, 2172, 2554, 2289, 2021, 2668, 531, 1840, 608, 966, 3131, 2466, 3207, 2215, 2392, 3184, 2920, 1792, 1054, 2176, 969, 2836, 1970, 3299, 2946, 1829, 797, 1080, 1793, 3095, 3544, 3505, 2999, 1694, 811, 1368, 3454, 2125, 3523, 3630]. high-confidence [2466]. High-Dimensional [2554, 3131]. high-index [2836, 3630]. High-Level [531, 1694]. High-Order [608, 3283, 2021, 3184, 1829, 3523]. High-Speed [83]. Higher [2379, 2697, 2898, 2151, 3578, 1097, 2229, 3142, 3512, 1282, 2995, 1712, 1063, 3558, 3674, 1053, 3577, 2449]. Higher-Order [2697, 2151, 2995, 3674]. Highest [1087]. Highly [566, 3686, 2667, 2122, 2858, 3579, 2534, 3332, 2525, 2646, 2761, 2876, 3034, 2676, 2818, 3521, 3676, 2807, 3077]. highly-oscillatory [2525]. Highly-Stable [566]. Hilbert [2455, 366, 2083, 1593, 2443, 2204, 1284, 3036, 792, 3549]. Hille [2170]. Hilliard [3654, 3524]. Himmelsmechanik [148]. Hiptmair [3701]. histogram [1667]. histopolation [3319]. history [39, 2766]. Hjärnan [140]. HJM [3082]. Hoare [1258]. Hoare-calculi [1258]. Hochbaum [1876]. Hodgkin [2709]. Hölder [2471, 2499]. Holm [3530]. Holman [2503]. holomorphic [2140]. Holt [1384]. homoclinic [3253]. Homogeneous [3574, 2096]. Hopf [2151]. Horn [848]. Horner [3687, 3214, 3235]. hour [1965]. Householder [2303, 260, 2680, 3129]. hp} [3697, 3667]. hp}-version [3697, 3667]. HSS [3305, 3243, 3113, 3249]. HSS-multigrid [3113]. Huber [2308, 1645, 1731, 2540]. Hugger [3261]. hull [1343, 1585, 1838, 1956, 1292, 1832, 1362, 1329, 2028, 1666]. hulls [1225]. hump [3318]. Hunter [2216, 3297, 1029, 3501, 3424, 3400]. Huxley [2709, 3153]. Hybrid [3445, 2971, 3543, 1226, 2849, 3388, 1287, 3232, 1316, 3561]. hydraulics [1176]. hydrodynamic [3710]. Hydrodynamics [2549]. hyperbolic [2824, 3492, 1989, 1527, 1300, 2357, 2500, 3011, 3098, 2196, 3101, 2854]. hypercube [1981]. Hypercubes [1743, 2014]. hypergeometric [1040]. hypergreedy [2011]. hyperinterpolation [3606, 2882]. hyperplane [3373]. hyperpower [1847]. Hyperquicksort [1747]. hypersingular [2994]. Hypothesis [30, 31].

I/O [1220]. ICC [55, 126]. IDeC [1608]. idempotent [3090]. identification [3547, 3655]. Identifiers [411]. Identities [2213, 2396]. Identity [439, 741, 1738, 1852]. Identity-Kernels [439]. IDEs [2939]. if [816]. IFIP [56, 232, 126]. II [265, 2237, 2248, 2284, 2241, 2311, 342, 354, 3034, 746, 3089, 14, 506, 91, 2787, 1247, 973, 1777, 832, 2679, 1900, 2890, 2449, 2842]. III [124, 2322, 1564, 2063]. Il’in [3219]. Ill [2380, 2687, 2732, 2621, 2327, 3525, 2871, 2943, 3370, 3369, 1697, 2095, 2495, 3122, 840, 1404, 1844, 2534, 2165, 2480, 1861, 2972, 3556, 2609, 2997, 3322, 3073, 2611, 2349, 2099, 2425, 3472]. ill-conditioned [840, 1404, 2534, 3556, 2349, 2099, 2425]. Ill-Conditionedness [2327]. ill-conditioning [3472]. Ill-Posed [2380, 2687, 2732, 2621, 3525, 2871, 2943, 3370, 3369, 1697, 2095, 2495, 3122, 1844, 2165, 2480, 1861, 2972, 2609, 2997, 3322, 3073, 2611]. Ill-posedness [2327]. ILU [1970]. Image [2748, 3370, 3369, 3583, 3441, 2673, 3147, 2593, 3308, 3571, 3090, 3288, 3141]. images [3585]. imaginary [1696, 1143, 916]. Imaging [2923, 2871]. immersed [3683]. immiscible [2936]. impact [3485, 2466]. Implementation [340, 442, 598, 1483, 2667, 6, 3032, 691, 1104, 1100, 785, 1050, 905, 2242, 2285, 2335, 1088, 887, 3505, 1985, 1229, 3535, 1318, 3013, 1022, 2679]. Implementing [396]. Implications [2688]. Implicit [2693, 3120, 2616, 737, 598, 1296, 1297, 353, 2406, 2432, 400, 3148, 2552, 2647, 538, 2500, 2373, 2558, 497, 371, 1907, 3105, 2577, 2043, 564, 805, 967, 1945, 2215, 2095, 3719, 1100, 785, 873, 1161, 2033, 2454, 3145, 3738, 779, 1459, 2148, 1283, 2256, 3480, 1101, 874, 1999, 1712, 3569, 3357, 1087, 1232, 2847, 2511, 2046, 915, 1442, 1775, 3665, 1733, 1489, 1754, 1817, 3724, 1714, 3412, 2021, 1612, 1735, 2463, 2802, 1605, 3051, 1543, 2764, 1449, 3432, 3655, 1978, 1719, 2681, 1461, 1034, 1066, 2211, 1597, 1835, 2197, 3053, 1061, 1021, 3316, 3591, 3720, 3603, 1053]. implicit-explicit [2577]. implicitly [3080]. implies [3107, 860, 959]. importance [2982]. Impressions [1116]. improve [2630, 2420]. Improved [967, 1959, 2684, 968, 2392, 1894, 603, 1242, 3235, 3676, 1209, 2304, 1077, 1282, 2234, 1807, 3722, 3534, 1022]. Improvement [2980, 664, 771, 1123, 1261]. improvements [1429]. Improving [2119, 1344, 720, 643, 1249, 805, 3328]. In-Place [508, 1159]. In-trees [1810]. Incidence [645, 1001]. Including [162, 2194]. Inclusion [1621, 430, 898, 3280, 2596]. inclusions [1847, 3498, 3692]. Incomplete [2660, 1796, 2, 1786, 2418, 1797, 2149, 1805, 2134, 2398, 2679, 2828, 2982, 1788]. incompletely [3231]. incompressible [3706, 2608, 2850]. Inconsistent [289, 3407, 2274]. Incorporating [2541]. increase [3357, 1210]. Increasing [511, 746, 3433]. increment [1102].

Incremental [2509, 1235, 3674, 1312, 1279]. Increments [99]. Indefinite [2738, 35, 1782, 2579, 3126, 2487, 2842, 2351, 2625, 3563]. indent [798]. indentation [920]. independence [2780]. Independent [3363, 187, 1608, 1952, 3554, 1669, 3600, 1630, 1589, 1130, 1986]. independently [3178]. Index [2714, 409, 467, 544, 615, 682, 753, 822, 893, 964, 1047, 1135, 1207, 1275, 1342, 1499, 1724, 1742, 1935, 2101, 2143, 2263, 2318, 2364, 2377, 2415, 2492, 2545, 2551, 2552, 2698, 2711, 2119, 2144, 2291, 2965, 1309, 2307, 2162, 2836, 1966, 2007, 3037, 2887, 2227, 2939, 3630, 3540]. index- [2162]. index-1 [2965]. index-2 [2307]. indexed [935]. Indexes [2348]. Indexing [1382]. Indices [1746, 550]. induction [3060, 1941, 2898, 2042, 2924]. Inductive [1691, 199, 1940]. inequalities [3448]. inequality [3657, 1332, 2830, 3522]. Inexact [2547, 2863, 2405, 3537, 3502, 3483, 3431, 2501, 3420, 2783, 3248, 3338]. inf [1937]. infants [2475]. inference [1691]. Inferring [1968]. Infinite [312, 556, 216, 1400, 1816, 1912, 3179, 1813, 3517, 3333, 3349, 2853, 1866, 1834, 1042, 1182]. infinitesimal [2902]. Infix [310]. Inflexion [266]. Influence [146]. Informatics [1679]. Information [273, 1554, 2732, 8, 14, 101, 154, 1140, 111, 992, 1805, 1051, 1299, 2908, 2541]. Inherent [689, 3640]. Inheritance [3300, 3265, 1454]. inhomogeneous [3482]. Initial [1565, 2568, 723, 2032, 2954, 989, 3555, 1398, 1371, 787, 2478, 1883, 3007, 1016, 1713, 1775, 2136, 1178, 2483, 1475, 2887, 2167, 1215, 828, 3352, 937, 1033, 1885, 1613, 2228, 3523, 778, 3320, 1848]. initial-boundary [2136]. initial-value [2954]. initial-value-problems [1033]. initialization [2536]. Ink [570]. Inner [3191, 2417, 3742, 876, 1089, 1121, 2486, 3065]. Input [23]. insert [1704]. Inserting [1172]. Instability [471, 460, 176, 1388, 1437]. instruction [2024]. instructions [1137]. Instruments [37]. Insulated [183]. Integer [2016, 644, 1700, 1578, 1596, 1651]. Integers [411, 511, 746, 1976, 1277, 1778]. Integral [470, 702, 2661, 352, 418, 177, 443, 227, 279, 313, 639, 656, 2663, 399, 509, 620, 744, 2731, 640, 120, 412, 491, 2688, 459, 268, 611, 428, 607, 3293, 1313, 1445, 21, 966, 3587, 2966, 2161, 2201, 3496, 2103, 2281, 3058, 2795, 1749, 1629, 1592, 3163, 3307, 3367, 1487, 546, 2994, 3694, 3730, 983, 1184, 1231, 2070, 2799, 1058, 1284, 861, 1032, 1102, 1149, 3285, 2385, 863, 1213, 3322, 2060, 1595, 1216, 2980, 2852, 2226, 3507, 1254, 1301, 3181, 812, 3193, 1462, 1167, 3077, 3014, 3720, 2260, 1106, 1299]. Integrals [618, 2633, 2, 298, 112, 548, 658, 246, 427, 712, 2667, 136, 538, 401, 35, 62, 3144, 1354, 2281, 63, 1576, 2282, 3664, 1816, 1029, 3332, 1611, 3373, 1813, 1526, 1774, 2761, 2996, 3725, 3713, 2459, 2937, 2899, 3398, 3532, 1563, 3652, 2818, 2927, 1165, 1317, 978, 1756, 1321, 1834, 903, 1042, 1372, 882, 1018, 3489, 3681]. Integrand [380, 3299]. integrands [1523, 1899, 831]. integrate [3555]. integrated [3437, 1705]. integrating [2630, 2055]. Integration [2683, 116, 582, 565, 2686, 275, 670, 2708, 380, 529, 601, 672, 2554, 459, 492, 624, 447, 162, 1493, 608, 3690, 2782, 3295, 1369, 1770, 1262, 1027, 2454, 2519, 2860, 2672, 857, 3460, 3213, 1331, 3108, 3635, 3668, 1750, 757, 913, 2250, 2284, 3748, 3500, 3373, 878, 759, 2479, 3640, 2909, 3504, 1990, 2839, 1595, 2649, 3010, 828, 2515, 3362, 1316, 977, 1065, 1150, 3643, 978, 1513, 1899, 1885, 3074, 2681, 1992, 864, 2805, 1737, 849, 2502, 1250, 1916, 2807, 2986, 1053]. integration/interpolation [1262]. integrations [2921]. Integrator [2642, 3614, 3624, 3580, 3539, 3700, 3124, 3625, 3534]. Integrators [2620, 3105, 3488, 3182, 2392, 2507, 3526, 2592, 2967, 3059, 3198, 3735, 3613, 3394, 2367, 3699, 2355, 3615, 3620, 3387, 3516, 3743, 2676, 2888, 1962, 3364, 1886, 1964, 3454, 3521, 3240, 3575, 3595, 3320]. Integro [597, 404, 536, 2218, 2470, 3697, 2810, 3529, 1176, 2799, 3679, 2951, 3548, 3098, 3052, 3511]. Integro-Differential [597, 404, 536, 2218, 2470, 3697, 2810, 3529, 1176, 2799, 3679, 2951, 3548, 3098, 3052, 3511]. intensity [2237]. interacting [3592]. Interaction [2747, 3744]. interactions [3514, 3176]. interactive [1939, 1076]. Interchange [1242]. interest [3670, 1405, 2983]. interface [3683, 3256, 3103, 3316]. interior [816, 3667, 3633]. Intermediate [224, 708]. intermittent [2974]. internal [2191]. International [134, 55, 57]. interpolant [1231, 2829, 2960]. interpolants [3071, 3108, 2909, 3168, 3152, 2817, 2911, 3074, 3026, 3202, 3052]. Interpolating [49, 2565, 474, 3163, 1247, 1564, 2539, 2117, 3102]. Interpolation [1309, 3106, 2933, 638, 2706, 211, 675, 2422, 724, 676, 588, 311, 2530, 3717, 939, 3551, 1971, 1262, 3277, 3715, 3083, 3345, 3406, 3070, 3513, 3447, 3726, 3590, 2835, 3307, 3367, 3566, 2005, 2345, 1038, 3061, 1829, 1845, 2221, 2310, 878, 1196, 336, 3632, 1284, 2135, 3471, 2813, 1806, 3728, 1285, 790, 996, 1356, 2254, 2352, 1899, 1217, 1833, 3000, 1736, 1546, 1668, 1900, 2036, 2257, 3712, 3091, 3722, 3238, 1318, 1236, 1627, 3634, 1988]. Interpolation-based [1309]. interpolational [2462]. interpolators [1185]. Interpolatory [1121, 3002, 968, 1529, 2455, 2097, 3200, 2516, 1513, 1756, 1449]. interpretation [1944]. Interpreting [507]. Interpretive [382]. Intersection [557, 1957, 816, 1585, 1117]. intersections [1872, 1925]. Interval [738, 505, 1860, 380, 604, 494, 506, 553, 1443, 2350, 3066, 3314, 626, 2276, 2691, 2304, 2094, 898, 2456, 1057, 948, 1164, 2596, 2137, 3330, 1387, 3413, 2207, 1233, 1761, 1818, 3692, 1546, 1900, 3339, 1464, 1868]. Interval-Arithmetic [494]. interval-Newton-method [2304]. Intervals [1195, 1400, 1912, 1441, 1668, 3100]. Introduction [2575, 2627, 2656, 2659, 2682, 2692, 2704, 2716, 2729, 2770, 2793, 2809, 2822, 2781, 2834, 2857, 2870, 2891, 2903, 2879, 2919, 2931, 2975, 2987, 3028, 3042, 2742, 17]. introduktion [17]. intrusive [3462]. intuitionistic [1942]. INV [1561]. invariance [2644, 1855]. Invariant [1644, 2748, 3312, 2182, 385, 1886, 1964, 3123, 2084, 2515, 1218]. Invariantization [2812]. Invariants [410, 3284, 2102, 2082, 3069, 3473, 3579]. Inverse [2561, 2547, 2747, 183, 138, 532, 1059, 2469, 2280, 2990, 955, 2238, 3420, 2309, 2508, 2179, 983, 2783, 2405, 1847, 1507, 3008, 3112, 2899, 1991, 3399, 2981, 3040, 1423, 2049, 3469, 3550]. Inverses [3, 2731, 573, 2105, 2396, 3086, 1353]. Inversion [344, 478, 692, 2636, 2034, 2004, 1478, 794, 2780]. invert [2089]. inverted [1851, 812, 2490]. Investigation [429, 574, 1148, 2141]. involutive [2449]. Involving [246, 2034, 969, 3664, 886, 902, 1526, 1774, 3516, 2889, 849]. IRK [2551]. Irreducible [2730, 446]. irrelevant [1685]. isarithmic [117]. isarithms [117]. isn’t [3752]. isolation [934]. isomorphic [1902]. isomorphism [2146, 825]. isospectral [3466, 3603]. isothermal [3468]. isotonicity [1621]. isotropic [3139]. Issue [2575, 2627, 2656, 2659, 2682, 2692, 2704, 2716, 2729, 2770, 2793, 2809, 2822, 2781, 2834, 2857, 2870, 2903, 2879, 2742, 2891, 2919, 2931, 2975, 2987, 3028, 3042]. issues [3109]. Itô [3105, 3577, 3578, 2735, 3303]. Items [411]. Iterated [841, 2991, 3245, 2226, 1254, 1301, 1020]. Iteration [693, 720, 2573, 695, 532, 1059, 2416, 2602, 3305, 3392, 2603, 2081, 2118, 1511, 2176, 2309, 2521, 3356, 2179, 2296, 2783, 3244, 2405, 3008, 3740, 2137, 2244, 1594, 1977, 2288, 3682, 1443, 1755, 1777, 1543, 2372, 1579, 1233, 2639, 3509, 3403]. Iterations [475, 2417, 2604, 3426, 2147, 2453, 2895, 3665, 1392]. Iterative [2173, 260, 255, 265, 898, 2564, 2069, 400, 1913, 860, 2786, 1145, 491, 664, 3685, 3599, 2601, 1280, 1436, 1796, 1791, 2871, 3196, 2161, 3083, 3145, 2148, 3122, 2283, 1310, 2673, 2783, 2241, 2477, 2420, 2482, 1530, 2915, 2875, 1508, 2191, 3696, 2172, 1975, 2225, 2385, 1123, 3656, 2763, 2611, 2013, 2272, 1455, 3457, 1357, 1648, 2527, 2940, 3141, 3041]. iterative-deepening [2013]. Iteratively [1974]. Ivar [717]. IVP [2836]. IVPs [3412]. Iyengar [2830].

Jackson [1527]. Jacobi [3446, 618, 1948, 1593, 2516, 3390, 3301, 3476, 258, 288, 1059, 2185, 2049]. Jacobian [2787]. Jacobians [1880, 3665]. Jan [1780]. Jens [3261]. Jet [570]. job [1051]. Johannesson [1463]. Johnson [1429]. join [1968]. joins [1805]. Jordan [3286]. Joule [2693, 2957, 3088, 3326]. Journal [2378]. Journals [64, 79, 90, 96, 114, 123, 132, 159, 167, 190, 202, 209, 231, 243, 272, 281, 292, 338, 361, 392, 420, 435, 450, 540, 594, 650, 679, 697, 732, 782, 802, 819, 853, 871, 889, 909, 925, 961, 985, 1007, 1024, 1044, 1069, 1092, 1111, 1132, 1154, 1169, 1204, 1238, 1257, 1272, 1324, 1359, 1374, 1466, 1481, 1495, 1533, 1534, 1551, 1582, 1601, 1615, 1632, 1654, 1675, 1722, 1740, 1979, 2023, 373, 501, 325, 519, 767, 484, 633, 316, 252, 714, 834, 348, 464, 46, 666]. Journals [54, 750, 406, 72, 104, 179, 220, 262, 304, 363, 942, 1187, 1222, 1567, 1837, 1871, 1995, 2051, 2080, 2158]. Julia [3638]. jump [3445, 3423, 3705]. jump-diffusion [3445, 3705]. jump-extended [3423]. Jumping [3451, 3456]. June [2064].

Kac [2580, 3217]. Kaczmarz [3212, 3389, 3545, 3497, 2938, 3671]. Kahan [3525, 2915]. KdV [3575]. Kernel [443, 3363, 3717, 1611, 1284, 3428, 3549, 3723]. kernel-based [3717]. Kernels [439, 620, 2470, 3697, 3456, 3451, 2799, 3627, 3036, 1213, 3077]. Kerner [1882, 1565, 1412]. Key [1554, 956, 2026, 1002, 1037]. Key-Lock-Pair [1554, 2026]. Keys [558, 2000]. KH [610]. Kind [639, 656, 509, 640, 459, 268, 1445, 966, 3058, 3173, 1749, 1487, 2994, 3730, 2070, 1058, 2135, 861, 1032, 1102, 1595, 2980, 2226, 1254, 1301, 812, 1462, 1167, 1106]. Kinematic [2856]. Kinetics [2517, 2825]. King [1151]. Knot [2424, 2556, 2087, 2911]. Knots [474, 2336, 1461, 2209, 2123]. Known [376, 643, 3231]. Kogbetliantz [3057]. kompilering [34]. Kontorovich [2731]. Korteweg [3273]. Kosloff [3626]. Kreiss [3317, 2059, 1415, 1919]. Kronrod [1815]. Krylov [2159, 2305, 3552, 3459, 3209, 3668, 3593, 3394, 2697, 3264, 2773, 3265, 3048, 3537, 3502, 3308, 2382, 3442, 3519, 2776, 3494, 2193, 2725, 2063, 2290, 3227, 3240, 3076]. Krylov- [2290]. Kuhn [952]. Kummer [630]. Kutta [2391, 2265, 1947, 3499, 3229, 2463, 890, 3054, 1914, 3417, 2043, 2615, 3355, 2291, 2315, 2292, 2988, 3160, 1945, 2187, 2215, 3719, 945, 897, 1100, 1575, 2246, 2616, 737, 785, 873, 1401, 1477, 2112, 1859, 2402, 2920, 2795, 1904, 3093, 2162, 3145, 455, 598, 1459, 1491, 2148, 2862, 1283, 2945, 1385, 1439, 2521, 2548, 2567, 353, 1712, 2406, 2970, 2323, 2550, 2444, 1440, 1039, 1087, 1699, 2056, 2076, 2847, 1987, 2046, 2057, 2084, 2204, 2778, 2838, 2223, 2334, 915, 2286, 2191, 1775, 2007, 2552, 3665, 958, 2762, 1733, 2584, 3284, 3095, 3375, 2071, 2198, 2252, 2312, 1915, 1754]. Kutta [3037, 2647, 538, 1714, 3505, 1918, 2586, 3154, 2021, 1612, 1735, 1898, 1991, 2138, 2271, 809, 2386, 2803, 1864, 1906, 2326, 2735, 2819, 1978, 1388, 1718, 240, 1304, 1719, 1819, 1461, 1850, 2211, 1835, 3303, 2528, 2412, 2126, 371, 3595, 3603, 3540, 2472, 1988, 2433, 1106].

Labelling [110]. labels [1808]. Lack [2116]. Lacunary [2345]. lag [1505]. lagging [2103]. Lagrange [2144, 3444, 3686, 2877, 2352, 1217, 3440]. Lagrangian [2718, 165, 2442, 3005, 3063, 146, 3364, 3508]. Lagrangian-based [3005]. Laguerre [2443, 877, 1287]. laminates [2410]. Lanczos [3080, 2174, 2354, 2177, 2394, 2480, 2486, 369, 2245, 2490, 371, 2690]. Land [358]. Land-Use [358]. Landau [2886, 3378]. Langevin [3215]. Language [264, 274, 128, 551, 495, 1094, 1431, 1171, 3072, 922, 1364, 1683, 1229, 1695, 1457]. language-driven [1431]. Languages [212, 1454, 330, 382, 237, 531, 710, 186, 346, 3263, 906, 1363, 1485]. LAPACK [2474]. Laplace [2904, 2004, 183, 3645, 138, 3217, 344, 2196]. Laplacian [525, 1416, 2035, 3535]. Large [2563, 2687, 4, 2572, 2747, 384, 206, 2173, 3714, 2943, 1463, 1697, 3588, 930, 2017, 1795, 3123, 3176, 2233, 2936, 3397, 2134, 3483, 3625, 2461, 3073, 3747, 2183, 3601, 2184, 3320]. Large-Scale [2687, 2943, 1795, 3073]. large-stepsize [3625]. Largest [958, 930, 1649]. last [3639]. Lattice [2570, 2914, 1402, 3135, 1776, 2358, 2839, 3238]. lattices [2977, 1990, 1916]. Lavrentiev [2763, 2226]. law [2162, 2847]. laws [2824, 3462, 2916, 2722, 3234, 3354, 2194]. Lawson [3555, 3499, 3579]. layer [3325, 3363, 3181]. layer-adapted [3325]. layered [3593]. Layers [2851, 2995, 3096, 3361, 3353]. leaf [1875]. leap [3232]. Leapfrog [2010, 3164]. Learning [3688, 992, 3288]. Least [453, 684, 260, 255, 222, 265, 3754, 3755, 2321, 3356, 2098, 2407, 2666, 642, 687, 2571, 747, 3476, 2373, 707, 1017, 2327, 2123, 2669, 2739, 1067, 2197, 2329, 589, 2741, 2559, 3685, 2880, 1476, 2754, 1892, 2579, 3626, 2881, 1972, 2331, 1402, 3635, 3659, 900, 840, 1264, 1404, 2510, 2332, 1547, 2180, 2069, 2130, 2233, 2150, 2299, 2251, 3201, 2287, 1332, 1367, 2461, 3126, 3336, 2678, 2361, 2804, 2828, 2274, 829, 1544, 3601, 1028, 2258, 2212, 907, 1974, 3469, 3550]. least- [3201]. least-square [907]. Least-squares [2098, 3626, 2881, 2233, 3201, 2274]. Lebedev [2731]. Lebesgue [605]. lecture [2719]. Lee [1618]. Left [350, 1588, 974, 2374]. Left-Recursive [350]. Leg [2594, 1883, 2779, 3027]. Legendre [1268, 3271, 831]. Lehmer [883]. Leja [3080, 3590, 1833]. Length [411, 2706, 386, 2305, 1377, 1345, 1619, 2650, 1036]. lengths [1128]. Less [535, 3714, 3384, 1449, 1460]. Letnikov [3573]. Letter [528, 562]. Level [223, 341, 671, 531, 1860, 1970, 1096, 2915, 1790, 1139, 950, 2998, 3204, 1694, 3222, 3731]. Levenberg [2593]. LeVeque [1919]. Lévy [3165, 3584]. Lévy-type [3584]. Lexicographical [356]. lexicographically [1803]. Lichtenstein [21]. Lie [446, 340, 3565, 2592, 2548, 2404, 3148, 2553, 3577, 3578, 2138, 2271, 3128, 2372, 2436]. Lie-Group [2548, 2404, 2436]. Like [2547, 2748, 723, 800, 755, 2034, 2002, 899, 1643, 2178, 2968, 2419, 1454, 1128, 3250, 936, 1180, 2636]. likelihood [994]. Limit [535, 216, 3538, 1086, 3637]. Limitations [2165]. limited [1328, 3310]. limiting [2420]. Limits [68]. Lindelof [2322, 2219, 2248, 1977, 1755, 1777]. Line [290, 186, 1656, 2455, 2992, 1241, 81, 277, 1471, 1839, 53, 3476, 3655, 3489, 3681]. line-implicit-solvers [3655]. line-sweep [1471]. Linear [2713, 453, 700, 3224, 71, 804, 2561, 260, 255, 222, 265, 487, 1590, 1591, 2591, 2730, 2563, 740, 2967, 693, 704, 75, 144, 203, 227, 2202, 2687, 399, 2697, 670, 489, 2323, 2572, 1913, 722, 329, 342, 354, 2334, 130, 642, 3035, 1672, 2445, 689, 2555, 1190, 333, 1294, 2751, 390, 2540, 2828, 1473, 2373, 3434, 301, 2339, 2574, 2614, 249, 289, 270, 323, 592, 590, 2558, 226, 93, 2671, 1393, 3685, 855, 3604, 2615, 966, 254, 2279, 1782, 1656, 2942, 1436, 2880, 3118, 3305, 3525, 2160, 2466, 3370, 3369, 3669, 2468, 1828, 988, 1697, 1892]. linear [2052, 2507, 3588, 1281, 2111, 2330, 2470, 1098, 2033, 2429, 2651, 3278, 2402, 2881, 1194, 1050, 839, 2308, 2148, 1402, 940, 912, 3474, 1143, 3480, 1711, 2955, 3031, 2863, 3122, 900, 584, 1645, 1797, 2534, 886, 905, 1013, 2369, 1406, 2477, 3186, 1846, 2180, 2069, 2130, 3254, 3407, 2758, 3319, 1930, 2512, 2233, 2884, 991, 2972, 957, 933, 1745, 2513, 3150, 2058, 2760, 1883, 2335, 2190, 916, 2086, 1790, 2251, 1983, 2876, 3034, 1145, 2997, 3020, 3089, 3430, 3455, 1862, 995, 2121, 2224, 2815, 3581, 1266, 1352, 2676, 1332, 2411, 2348, 3545, 1863, 1902, 2461, 3675, 1215, 1362, 2061]. linear [1031, 1108, 3327, 3249, 879, 3330, 3577, 3578, 1210, 2938, 2678, 2889, 1353, 828, 3250, 1470, 3362, 1786, 2979, 792, 2255, 3023, 1701, 2273, 2804, 2226, 3452, 2302, 1580, 1761, 1818, 1510, 3549, 2047, 2227, 3601, 3189, 3339, 1104, 3170, 1648, 2527, 847, 2258, 1841, 1322, 2212, 2940, 3192, 3227, 2779, 2790, 1295, 3689, 849, 907, 999, 1319, 2639, 3723, 2780, 1974, 2276, 3403, 2031, 149, 2930, 3205, 2625, 3563, 3511, 2426]. Linearization [735, 2888, 3597, 2672, 2944, 3541, 3629, 1315, 2897, 3033]. Linearization-preserving [2888]. Linearized [3298, 3056, 2637]. Linearly [2693, 3707, 2764, 3316, 3569, 1931, 3600, 1597, 1835]. linearly-implicit [3569]. Lines [49, 2285]. linieskriver [53]. link [1139, 950]. link-level [1139, 950]. Linked [1158]. linking [1457]. Linpack [1865]. Liouville [3450, 1399, 2578, 2990]. Lipschitz [3651]. liquid [3251, 3508]. LISP [936, 507]. LISP-like [936]. List [100, 200, 1937]. lists [1960, 1708]. literature [2278]. Litt [38]. little [38].

LLGAUSS [78]. LMF [2439]. LMF-based [2439]. loading [995]. Lobatto [116, 2581, 2665, 2552, 3301, 850, 3540]. Lobatto-type [3540]. Local [3044, 1945, 3210, 1881, 1814, 1509, 2789, 2890, 3453, 3368, 3493, 2754, 2672, 3514, 2944, 2955, 3427, 2431, 3325, 1073, 2268, 3463, 1862, 3399, 3691, 3321, 2897, 3033]. localized [3703, 3686]. Locally [3298, 3159, 2442, 2114, 2359, 3600]. Locating [2356, 1492]. Location [707, 3211]. Lock [1554, 2026]. locus [1060]. LOD [3146, 1020]. Löf [1379, 1943]. log [3519]. log-determinant [3519]. logarithm [1463, 1715]. Logarithmic [2362, 443, 2538, 2401, 412, 2300, 2766]. Logarithmic-Trigonometric [412]. logarithms [1177]. Logic [276, 301, 2013]. Logical [4, 245]. Long [3198, 3059, 3362, 2790, 3521]. Long-term [3198, 3521]. long-time [3362, 2790]. longest [1347]. Look [2111, 110, 576, 2164, 3039]. Look-ahead [2111, 2164]. Look-up [110, 576]. looking [2374]. Loop [705, 1346]. Loop-free [705, 1346]. Loops [748]. Lossless [1805]. Lösung [84]. Lösungsmethode [168]. Lotka [3182, 3557]. Low [2695, 442, 3641, 3613, 3137, 2553, 3547, 2254, 2361, 3633, 2466, 3553, 3624, 3460, 3580, 3539, 3700, 2476, 3589, 3708, 3657, 3615, 3620, 3483, 3743, 3636, 3124, 3608, 3481, 3695, 2928, 2841, 3689, 3575, 3624]. low- [2466]. low-complexity [2476]. low-order [3657]. Low-rank [3641, 3137, 3547, 3553, 3624, 3460, 3580, 3539, 3700, 3615, 3620, 3743, 3124, 3608, 3695, 3689, 3624]. low-rank-adaptive [3708]. low-regularity [3575]. Lower [3477, 2571, 3705, 1904, 971, 1295]. lowest [3072]. lowest-order [3072]. LPMODSIMPLEX [188, 498, 499]. LR [2222]. LSQR [2609, 2139, 2208]. LSS [2438]. LU [2342, 2662, 2294, 2457].

Lyapunov [3426, 2724, 2932, 2708, 3589, 2401, 3483, 3009, 3377].

M [3568]. M-matrices [3568]. M.I.T. [1037]. Macdonald [2701]. Mach [2695]. Machine [170, 119, 237, 572, 3748]. Machine-Code [119]. Machines [397, 1157, 1486, 1572]. Maclaurin [2362, 2199]. Macro [653, 3474, 2825]. macroscopic [2816]. made [1355]. Magnetic [40, 3060, 2898, 2924, 3666]. magnetic-current [3666]. magneto [3710]. magneto-hydrodynamic [3710]. Magnus [2683, 2392, 2335, 2488]. Mahony [3596]. maintain [2084]. Maintenance [205, 1309, 2969, 1051]. Majorization [3308]. Majorization-minimization [3308]. Making [215, 1229]. male [1855]. Management [669, 726, 134, 1910]. Manifold [2553, 2472]. Manifolds [2326, 3553, 2408, 2479, 3662, 2515, 3712]. Manipulation [433]. Manipulative [417]. Mannila [1920]. Manteuffel [3282]. mantissa [1357]. Manufacturing [85]. Many [1220, 1710, 3066, 3454, 2155, 2503]. many-body [3454]. many-dimensional [2155]. Map [322]. Maple [3461]. Mapping [1277, 1435, 21, 2141, 2239, 3652]. Mappings [738, 2300, 3342]. maps [117]. marching [2136, 2061]. Marking [533]. Markov [736, 2296, 585]. Marquardt [2593]. Marriage [381, 534, 1293, 1472, 1381, 1768, 1729]. Martensen [2737]. Martin [1379, 1943]. Martin-Löf [1379, 1943]. Maruyama [3294, 3480, 3670, 3567, 2837]. masks [2934]. mass [3004]. mass-conservative [3004]. master [3445, 3324, 2849]. Match [756, 1473, 1840, 1115, 1228]. matched [3096, 3361]. Matching [1639, 2011, 1636, 1855, 2243, 1924, 1858, 2031]. Matematiken [140]. Material [138, 970]. Mathematical [276]. Mathematicians [57]. Mathematics [2364, 2719, 140, 2377, 2415, 2545].

MATLAB [2313]. Matrices [2590, 441, 2366, 2403, 169, 2241, 2747, 2733, 2749, 2585, 688, 2619, 478, 258, 288, 708, 2328, 2390, 3117, 3404, 2160, 3714, 3132, 3207, 2494, 2264, 3119, 3631, 3183, 3479, 2034, 3460, 3372, 1183, 3263, 3668, 3421, 2508, 2509, 2295, 3384, 1948, 2105, 3123, 2431, 1771, 2266, 2476, 2968, 3568, 2536, 3515, 2149, 1753, 2131, 3147, 2223, 3086, 2827, 3246, 772, 2421, 2398, 3546, 2060, 2423, 2878, 3411, 2089, 3127, 2035, 972, 121, 135, 1353, 2912, 3390, 1001, 2349, 3431, 2063, 2153, 2928, 2636, 3040, 2612, 2127, 1907]. Matrix [438, 295, 350, 2231, 684, 729, 3714, 2366, 2772, 2906, 2401, 2524, 2583, 154, 1415, 494, 623, 3259, 390, 645, 646, 309, 370, 1539, 2328, 2376, 1200, 249, 270, 662, 3380, 1802, 1796, 2173, 3081, 3116, 2344, 3553, 3120, 3121, 3383, 3479, 3405, 3296, 2835, 930, 2177, 2045, 3541, 3384, 931, 3264, 2114, 3708, 3094, 1847, 2884, 3615, 3620, 844, 2316, 3442, 827, 3517, 2059, 3702, 3498, 3572, 2776, 1348, 3125, 3494, 2654, 3577, 3578, 3747, 2361, 2465, 3680, 3414, 3272, 218, 1625, 2981, 3318, 2464, 1197, 1648, 865, 1919, 847, 1993, 2901, 2049, 1933, 2129, 2214]. matrix [2625, 1302, 2249, 2400]. Matrix-equation-based [3259]. Matrix-free [2772]. Matrix-less [3714, 3384]. Matrix-Valued [2328]. matroid [1468]. Max [1334, 1604, 1748, 1365]. Max-min [1334]. maxima [2153]. maximal [2305, 2107]. maximization [3318]. maximizes [1468]. Maximum [3439, 2428, 423, 2709, 2801, 1630, 1589, 431, 2624, 1952, 2824, 2246, 2307, 1669, 3613, 3515, 994, 3546, 3219, 1858, 2985, 1986, 2031]. Maximum-norm [2801, 2624]. Maximumnorm [379]. Maxwell [3344, 2914, 3174, 3658, 3663, 2984]. Mazzia [2452]. McKean [3592, 3558]. McKee [1050]. MDS [3740]. Mean [3105, 3294, 2745, 2389, 3333, 3467, 3135, 3141, 3315]. Mean-Square [2745, 2389, 3105, 3294, 3333, 3467, 3315]. Means [194, 385, 3119, 3631, 2002, 3641, 2222, 2166]. Measure [1537, 2775, 992, 3608, 3038, 3692, 1708]. Measured [2329]. Measurement [654, 586, 3399]. measurements [1426]. measures [2860, 2097, 2072]. Measuring [1114, 3125]. mechanical [2093]. mechanics [148]. Mechanism [1554]. mechanisms [1259]. media [3004, 3177, 3410, 2650, 3491]. median [1831]. medical [39]. Mellin [2904, 794]. memoriam [1780]. Memory [486, 669, 40, 2039, 788, 2374, 3115, 3536]. memory-efficient [3115]. Merge [195, 2784, 1261, 1747]. Merging [1159, 1326, 1138, 1605]. Meromorphic [2181]. Mersenne [482, 68, 51, 86, 108]. Mersenne-Type [482]. Mesh [1743, 2734, 3371, 2113, 2777, 1108, 1337, 2210, 2854]. mesh-based [3371]. Meshes [2743, 3044, 1997, 3072, 1056, 3427, 1845, 3325, 3449, 3167, 2978, 1984, 2028, 3698, 2961, 3523, 3562, 3353]. meshfree [3723]. meshless [3371, 2806, 2986]. meso [2825]. metadata [2969]. Method [453, 580, 523, 2518, 312, 244, 376, 2174, 2546, 418, 442, 2366, 2547, 2320, 3754, 2321, 480, 399, 3753, 275, 67, 112, 120, 246, 266, 195, 687, 603, 7, 2325, 2553, 695, 2699, 2734, 109, 552, 1759, 2689, 137, 494, 477, 2668, 268, 185, 555, 447, 730, 426, 527, 2435, 258, 288, 532, 2557, 2574, 2739, 2341, 592, 608, 2702, 627, 302, 728, 2690, 1435, 2559, 168, 3368, 3684, 2304, 3293, 1802, 1528, 3616, 1811, 1482, 3538, 3587, 2858, 3144, 3195, 3404, 3080, 3525, 2871, 3068, 2306, 3207, 2718, 2468, 2990, 3196, 1027, 2989, 1281, 2379, 1157]. method [2161, 2470, 3057, 3306, 3496, 2672, 1904, 3083, 1054, 2066, 981, 2921, 858, 839, 1125, 1062, 1320, 1540, 1642, 1922, 2039, 2308, 3146, 3697, 3420, 3003, 867, 3173, 1561, 1860, 1629, 1592, 3004, 3107, 2017, 2256, 3480, 3670, 982, 3668, 2508, 2442, 2220, 2906, 3658, 3646, 1829, 3567, 2608, 2993, 990, 1432, 1789, 2105, 2522, 1705, 2394, 2179, 2825, 3730, 1882, 1014, 2476, 2896, 1231, 1565, 2149, 2285, 3396, 3693, 3357, 3407, 1753, 1039, 3385, 2883, 1286, 2070, 3667, 2165, 877, 948, 1623, 3085, 3537, 3502, 2785, 2333, 1813, 3086, 3255, 3374, 2269, 3164, 3347, 1492, 1441, 1663]. method [2335, 2646, 2958, 1442, 1664, 3032, 3478, 3309, 2134, 84, 130, 1665, 1794, 2299, 861, 3644, 862, 1412, 2078, 2157, 3216, 3518, 1173, 2949, 3167, 3677, 1032, 3037, 1905, 2243, 3138, 2910, 1300, 3230, 1287, 2287, 2959, 3247, 3573, 3248, 3581, 2357, 1760, 3728, 863, 1107, 2950, 3389, 2166, 3679, 3153, 3125, 2359, 997, 1124, 1214, 3682, 2817, 2486, 3349, 3249, 1418, 1386, 3154, 2587, 2763, 2951, 1179, 1387, 1364, 3250, 3140, 2371, 2980, 3560, 1033, 3155, 1701, 1317, 2273, 2226, 2999, 3399, 1885, 3476, 1682, 1544, 2108, 2245, 218, 3318, 3098, 3286, 1060, 1545, 1254, 1301, 2022]. method [3237, 2622, 3287, 2184, 3339, 2010, 864, 1648, 3401, 1850, 3204, 3453, 3638, 2258, 3534, 2425, 3192, 3435, 2196, 1211, 2918, 2301, 1019, 3013, 1462, 1020, 3466, 2490, 2640, 3103, 3561, 2075, 3220, 3523, 3618, 2807, 3041, 3337, 3509, 3610, 3691, 2505, 3403, 3469, 3321, 1218, 2351, 3666, 3550, 3720, 2260, 3651, 3379, 3630, 3710, 2068, 2771, 2956]. Methoden [148]. Methodology [1114]. Methods [2693, 581, 328, 469, 655, 2684, 2743, 582, 487, 524, 2745, 2380, 737, 2591, 2730, 454, 2564, 471, 566, 75, 203, 297, 1677, 2548, 686, 1101, 2664, 2567, 509, 744, 640, 2419, 2406, 224, 2550, 2708, 2707, 444, 400, 1913, 2551, 657, 2666, 2334, 2524, 2733, 2667, 2800, 2594, 319, 2552, 2324, 458, 2698, 2647, 491, 2749, 2688, 2434, 674, 643, 459, 2710, 675, 333, 624, 2751, 176, 554, 2326, 2711, 2399, 707, 2735, 1539, 1992, 249, 2126, 497, 609, 226, 628, 371, 2472, 2433, 1798, 3685, 1926, 2892, 3599, 3417, 2577, 2043, 1961, 2231, 1313, 1445]. methods [2032, 966, 564, 3355, 800, 2119, 2144, 2159, 2601, 2291, 3304, 2315, 2292, 755, 3016, 3295, 1436, 1796, 1781, 1791, 2602, 2173, 2418, 3242, 3305, 3392, 2102, 3565, 1945, 2187, 2215, 3370, 3369, 3552, 3669, 1769, 2044, 2859, 1262, 988, 3133, 3183, 3602, 2237, 2111, 2393, 2953, 3161, 1584, 1880, 2265, 945, 897, 1100, 1098, 1078, 1099, 1120, 1575, 2246, 2616, 785, 873, 1161, 1477, 1590, 1591, 2033, 2112, 2429, 2454, 2519, 2651, 2954, 3208, 3278, 1859, 1947, 2402, 2920, 2932, 3687, 3279, 3069, 3555, 2795, 898, 1792, 3093, 2002, 857, 1194, 1230, 1398, 1371, 2905, 2176]. methods [1282, 2162, 3527, 969, 1195, 1162, 1336, 1505, 3145, 3473, 779, 2054, 2247, 3046, 3528, 2104, 3213, 1491, 2148, 1252, 2862, 1749, 2067, 1283, 912, 3209, 2945, 2991, 3299, 1143, 1385, 2944, 3210, 2955, 1439, 2521, 1487, 1056, 3031, 1516, 3427, 2381, 3122, 2178, 3582, 1644, 619, 1893, 3707, 2250, 2284, 3686, 2332, 874, 886, 1296, 1297, 1795, 905, 2404, 2297, 1057, 3709, 3325, 1610, 1752, 3542, 1771, 3492, 2773, 3265, 1406, 2757, 1712, 3199, 2241, 2970, 3515, 2444, 1408, 914, 2478, 1122, 1087, 1232, 1699, 1732, 2056, 2076, 2408, 2758, 2847, 2482, 2511, 3657, 2480, 2907, 2849]. methods [1987, 2046, 1847, 3063, 2884, 2204, 2778, 2838, 3229, 3006, 3179, 2346, 3557, 957, 933, 2133, 3323, 2513, 3308, 1883, 2347, 2382, 1058, 1144, 2286, 2525, 1508, 2190, 808, 2191, 3637, 1016, 1713, 1775, 3696, 2007, 3665, 761, 958, 916, 1562, 2762, 2786, 1790, 2172, 3725, 3713, 1178, 1733, 2876, 3034, 1975, 2997, 3359, 2584, 3284, 3095, 2071, 2198, 2252, 2312, 3544, 1915, 1081, 3702, 1475, 3021, 1754, 2088, 1102, 3483, 2542, 1862, 887, 3360, 3572, 3653, 1817, 2225, 3113, 3486, 3519, 3724, 1814, 2385, 3656, 1266, 1714, 3338, 2950, 2776, 2411, 3574, 3505, 1918, 2348, 3322, 2167, 2427, 1931]. methods [1031, 1246, 1308, 2115, 2289, 3412, 2021, 1612, 1735, 2463, 1898, 1991, 3577, 3578, 2138, 2271, 3340, 2889, 762, 828, 953, 973, 2193, 3350, 809, 890, 1419, 1509, 3362, 1427, 2979, 3352, 2488, 2803, 1864, 1906, 2679, 2272, 3233, 2274, 1151, 3698, 2302, 3302, 2229, 3414, 2152, 2764, 1449, 1438, 2819, 3065, 1493, 1510, 1963, 3434, 1304, 1719, 2290, 2681, 3099, 1650, 1819, 1180, 1461, 2047, 2227, 2464, 3052, 2275, 1834, 2501, 830, 2185, 3391, 2527, 3377, 148, 1597, 1835, 2940, 2387, 3328, 3303, 3227, 1249, 1318, 2638, 2779, 1720, 2624, 1035, 2528, 2449, 1201, 1319]. methods [2504, 2639, 3053, 3467, 3522, 3711, 1105, 1061, 1021, 1182, 3027, 2902, 3723, 1167, 921, 2791, 3054, 3077, 3222, 3731, 3716, 3078, 3142, 3316, 2626, 2436, 3591, 2986, 3291, 3511, 1868, 3603, 3540, 3298, 1988, 1458, 1053, 1106, 1914, 2418]. Metric [1687, 3745, 3739, 3740, 1596, 1651]. MHD [3056, 3666]. micro [3474, 3410]. micro-heterogeneous [3410]. micro-macro [3474]. micromechanics [2650, 3622]. Microprocessor [568]. Microprogramming [709]. microscale [2844]. microstructures [2410]. mid [1337, 1365, 1236]. mid-points [1236]. Midknots [722]. midpoint [3700, 1489]. mildly [1926]. Miller [496]. Milne [174]. Milstein [3455, 3644, 3430, 3432, 3053]. Mimetic [3415]. min [1604, 1334, 1748, 1365]. min-max [1604, 1748]. min-mid-max [1365]. Mindlin [2845]. Miniaturization [92]. Miniatyrisering [92]. minima [2153]. minima/maxima [2153]. Minimal [2471, 691, 1826, 2188, 1501, 3198, 1345, 1276, 1367, 1141, 3560, 1709, 3466, 1891]. minimal-variable [3466]. minimality [1853]. Minimax [955, 2017, 2236, 1894, 2086, 2008, 1868]. Minimization [3118, 2687, 3073, 3242, 3308, 3112, 3572, 2499, 2816, 2831]. minimizers [3459]. Minimizing [1923, 3091]. Minimum [2297, 3403, 2094, 3565, 2760, 2786, 1983, 1619, 1825, 815]. minimum-residual [2786]. minors [3225]. mismatched [3389]. missing [2928]. Mixed [387, 2880, 1171, 3657, 3019, 3672, 3088, 3167, 3645, 1179, 1985, 3666]. Möbius [1478, 204]. mock [3635]. mock-Chebyshev [3635]. Mode [365]. Model [440, 192, 3593, 2746, 2709, 671, 5, 1174, 386, 2517, 2724, 3030, 3670, 2645, 2936, 3557, 3360, 3653, 2725, 938, 3585, 2983, 1295, 3508, 3115, 1572]. Modeling [2093, 2024, 2210, 2475]. modelling [3174]. Models [2642, 1140, 3082, 1645, 3438, 3165, 2814, 2541, 1974]. moderately [2789, 2890]. Modern [1438, 2766]. modes [3175]. Modifiability [1114]. Modification [182, 343, 496, 2149, 1492, 1664, 994, 1865, 1876]. Modifications [2304, 2686, 877, 3038]. Modified [2577, 3304, 663, 1142, 487, 618, 3473, 2633, 235, 307, 458, 606, 1317, 426, 2736, 3275, 2892, 1811, 1602, 1770, 2581, 990, 2242, 2130, 1526, 3665, 130, 3375, 1102, 3348, 3247, 3322, 1543, 3129, 3400, 3737, 2588, 3054, 2156]. Modifying [2277, 2314]. modular [1706]. Modules [1639, 1457]. moduli [2330, 1711]. modulo [1922, 1064, 1311, 314]. molecular [2644, 2811, 2355, 3721, 3014]. Moments [606, 2241]. momentum [3560]. monadic [1452]. Monge [3195]. Monitors [1015, 1073, 901]. Mono [2463]. Mono-implicit [2463]. monolithic [3533]. monomial [3446]. Monotone [3196, 693, 3255, 2527, 1004, 2894, 2993, 2837, 2482, 2875, 2956, 3644, 2584, 2801, 3594, 2036, 1180]. monotonic [3459]. Monotonicity [2705, 2539, 3294, 3694, 3467]. monotonicity-preserving [3694]. Monte [192, 3525, 3068, 3082, 1077, 3185, 2250, 2284, 5, 3584, 3217, 3607, 3643, 730, 3234, 3354, 216, 3649]. Moore [3040]. Mortar [2689, 2533, 2243, 2359, 2867, 2831]. Most [2413, 1967]. motion [1502, 3270, 3269, 3651]. motivated [3553]. Motzkin [3407]. Moulton [2511]. Movement [726, 1637, 1825, 2090]. moving [2812, 3201, 3336]. MR [1320, 1089, 1651, 1002, 1964, 1338, 3227]. MR2057361 [2655]. MRA [3085]. Mühlbach [1119]. Multi [1657, 2685, 671, 1840, 2399, 2681, 2529, 2260, 2218, 3704, 3727, 3567, 3266, 2056, 2040, 1096, 3547, 3287, 3401, 3204, 1843, 3222]. multi-asset [3287, 3401]. Multi-Attribute [1840, 2040]. multi-derivative [2056]. Multi-dimensional [2529, 2218]. Multi-disk [1657]. multi-edge [1843]. multi-fidelity [3727]. Multi-Frequency [2685]. Multi-implicit [2681]. Multi-Level [671, 1096, 3204, 3222]. Multi-parameter [2260, 3266]. multi-parametric [3547]. multi-screens [3704]. Multi-Symplectic [2399]. multi-valued [3567]. multibody [2856]. Multics [709]. Multics-type [709]. Multiderivative [670, 1440, 1098, 1099, 761, 1201, 3591]. multidiagonal [2105]. Multidimensional [1191, 883, 3146, 2945, 3627, 3054]. Multidisk [1376]. multidomain [3381]. Multifilters [2506]. Multigrid [2237, 2247, 3031, 2746, 2733, 1594, 2670, 3183, 2796, 3146, 2004, 2239, 2179, 3113, 3574, 2973, 2623, 2425, 2196, 3509, 3610, 3716]. multihop [1744]. Multilayer [2569]. multilayered [3491]. Multilevel [3068, 2524, 3584, 3232, 3217, 3234, 3691, 3649, 1791, 2602, 3554, 2247, 3185, 1821, 3049, 2997, 2244, 3354]. multiphase [3228]. Multiple [3380, 480, 601, 1454, 414, 1840, 707, 1464, 1997, 981, 3473, 3646, 1882, 3492, 1508, 862, 2137, 1417, 2925, 809, 890, 1579, 3239]. multiple-shooting [1417]. Multiplication [1064, 1085, 1706, 758, 2596]. Multiplicative [402, 2046, 2802, 3585, 3053]. Multiplicities [446]. multiplicity [1130]. Multiplier [365, 3242].

Multiplier-Summation [365]. multipliers [1281, 2330, 3444, 1711, 2877, 2622]. Multipoint [319, 1508, 1442]. Multipole [2668, 2353, 2921, 2168]. multiprocessor [1013, 2012]. multiprocessors [2374]. Multiprogrammering [74]. Multiprogramming [74]. Multirate [1406, 2444, 3569, 2902, 1963, 2788, 3533]. multiresolution [2910]. Multiscale [3322, 3599, 3646, 1417, 3320]. multisplitting [2601]. multistage [2454, 1411]. Multistep [582, 487, 2730, 471, 75, 670, 2594, 2698, 675, 333, 226, 966, 2110, 2144, 3016, 3069, 2307, 1194, 1282, 912, 1143, 1406, 2758, 3006, 1883, 761, 916, 1411, 1562, 3021, 1862, 1814, 1031, 828, 1580, 1510, 2022, 1104, 3391, 2638, 2779, 1319]. Multistep-Multiderivative [670]. multivalue [1098, 1099, 1120, 1610, 1752]. Multivariate [2705, 2562, 3345, 1506, 3393, 2989, 1529, 1643, 2096, 3214, 3566, 2250, 2284, 2947, 2761, 3152, 3139, 3051, 880]. multiwavelet [3587]. multiwavelets [3507]. multiway [1890]. Munich [56]. Munkre [1022]. mutation [3030]. mutually [2337].

n [591]. Nagumo [3514]. Naimark [2325]. national [1453]. Natural [564, 1661, 1284, 2071]. nature [1426]. Naur [1703]. Navier [3706, 2608, 3657, 2959, 3022, 1850, 2211, 2637]. Near [378, 480, 2327, 1770, 1894, 797, 2008, 2109]. near-minimax [1894, 2008]. Near-Ring [378]. nearest [1921, 3498, 1596, 1651, 2900]. nearly [3652, 3489, 3681]. Necessary [3289, 1446, 979, 1819]. necessity [1229]. Need [42, 2327]. Needed [465, 415]. Negative [377, 2564, 308, 1128]. neighborhood [1822]. neighbors [1921]. nematic [3251, 3508]. Nested [475, 695, 3526, 1519, 1891]. Network [41, 3727, 1744, 3721, 3438, 1734]. Networking [1428]. Networks [233, 3445, 3661, 1555, 1620, 1244, 1474].

NEUCC [273]. Neuere [148]. Neumann [2533, 3492, 2646, 3745, 3739, 2867, 2841]. neural [3661, 3727, 3721]. Neutral [2319, 2558, 2133, 3323, 1178, 2198, 2088, 3027]. Neville [1119, 993, 1152]. Newer [148]. Newton [2304, 800, 2601, 755, 2518, 3045, 3406, 2547, 981, 425, 2308, 3420, 3590, 1829, 2608, 2332, 2566, 2522, 3356, 2748, 3665, 458, 1917, 996, 552, 1543, 2371, 2372, 1833, 1017, 198, 2940, 3192, 2702, 2504, 3509, 3469, 1218, 2559]. Newton-Like [2547, 800, 755]. Newton-type [2504]. Nial [1683]. Nicolson [3164, 3487, 3188, 2597]. Niels [717]. Nielsen [394, 355]. Nielson [3513]. Nitsche [2635, 3103]. no [71, 62, 88, 188, 78, 63, 2655, 89, 241, 112, 207, 336, 130, 498, 499, 121, 135, 200, 250, 156, 149]. Noble [1032]. Nodal [3100]. node [2937, 1875]. Nodes [1553, 2688, 3428, 3359, 2060, 3239]. Noise [2735, 2671, 3382, 3654, 2945, 3474, 2944, 2896, 1987, 2046, 2915, 3033, 3020, 3089, 3037, 2802, 3585, 2890, 3053, 3315]. noisy [3648, 2938]. Non [3002, 1098, 377, 398, 2664, 2334, 257, 308, 2339, 2558, 93, 3729, 3604, 929, 1927, 3116, 2990, 3586, 816, 1590, 3446, 839, 3198, 2934, 2282, 3071, 3480, 2431, 2995, 3462, 2056, 2512, 957, 3255, 1883, 2634, 3516, 3034, 3151, 906, 2949, 2801, 3249, 2802, 974, 3559, 3431, 3065, 2639, 3618, 3221, 3691, 3403, 3651]. non-closure [906]. non-convex [2634]. Non-Deterministic [257, 929]. non-equidistant [3221]. non-equilibrium [3255]. non-Euclidean [3116]. non-existence [2056]. non-Hermitian [3249, 3431, 2639, 3403]. non-intrusive [3462]. non-left-recursive [974]. Non-Linear [2334, 2339, 2558, 93, 1098, 3604, 1590, 839, 3480, 957, 1883, 3034]. non-Lipschitz [3651]. non-monomial [3446]. non-monotone [2801]. Non-Negative [377, 308]. non-random [3516]. non-regular [3586]. non-resonance [3198]. non-selfadjoint [2949, 3691]. non-smooth [3729, 1927, 2431, 2512, 3151, 3618]. non-standard [2282, 2995, 3065]. non-stationary [2934]. non-symmetric [2990, 3559]. Non-Symplectic [2664]. Non-Uniform [398, 3002, 3071, 2802]. nonaffine [3658]. Nonautonomous [3253, 2323]. Nonconforming [2689, 2243, 2359, 3222]. nonconvex [2726, 2816, 3636]. nonconvexity [1818]. nondifferentiable [1869]. nonexistence [1477]. nonexpansive [3090]. Nonhomogeneous [1269]. nonisomorphism [1380]. Nonlinear [597, 1078, 1099, 352, 1055, 509, 744, 1039, 2538, 2513, 687, 604, 2620, 414, 2245, 1017, 1319, 2559, 3729, 2577, 1961, 2601, 2186, 3295, 3207, 2824, 1591, 3436, 3555, 2795, 3437, 858, 3059, 2430, 1403, 2220, 3061, 1731, 2894, 2993, 817, 2332, 2522, 1524, 3395, 3693, 2070, 3148, 2593, 3487, 2333, 2875, 2956, 3637, 2876, 1700, 1102, 1124, 3169, 828, 2979, 1548, 3233, 3302, 829, 3273, 774, 1719, 2290, 3339, 1422, 2962, 2983, 1082, 1720, 2640, 2504, 3378, 3618, 780, 2831, 2276, 2505, 2820, 3316, 2832, 1302, 1458]. nonlinearly [1664]. nonlocal [1893]. Nonnegative [2328]. nonnegatively [2871]. Nonnormality [2759, 2176]. nonoverlapping [3707]. Nonpolynomial [1717]. nonpositive [3086]. nonrelativistic [3637]. nonresidues [1370]. nonselfadjoint [1797, 3113]. Nonseparable [2663]. nonsingular [2127]. nonsingularity [2160]. Nonsmooth [2962, 3341, 1487, 3562]. Nonstandard [3449]. nonstationary [3245]. Nonsteady [299]. nonstep [331]. nonsymmetric [3597, 2177, 2134, 2089, 2063, 3227, 2400, 3252]. nonuniform [3262, 1056, 1351, 2911, 3391, 3523]. NordDATA [273]. NordDATA-68 [273]. Nordic [1496, 1655, 2378, 1428]. Nordsam [20]. Nordsieck [185]. Norm [2744, 2401, 309, 3223, 3493, 2428, 3459, 2509, 2266, 2241, 3515, 2180, 3112, 2801, 2926, 3257, 2361, 1544, 3219, 2981, 3091, 2766, 2624, 1932]. Normal [585, 136, 212, 2619, 258, 647, 164, 241, 1587, 1216, 974, 1625]. normality [2912]. Normalized [3026, 2893, 3281]. Norms [2471, 2538]. Norwegian [38, 94]. notation [1766]. Note [350, 654, 469, 1084, 694, 1158, 120, 558, 388, 2572, 510, 957, 712, 228, 2437, 761, 1975, 695, 1220, 460, 172, 2452, 1103, 2341, 226, 346, 576, 728, 45, 2771, 939, 2315, 1476, 2118, 884, 1194, 981, 779, 2388, 1816, 1404, 931, 817, 905, 547, 1547, 1407, 797, 1063, 1390, 2837, 3487, 1127, 807, 2316, 1040, 3665, 2077, 1075, 2157, 917, 1469, 2722, 887, 1129, 1890, 331, 2192, 1448, 1649, 2427, 1735, 3270, 3269, 923, 2029, 1151, 1472, 2092, 850, 1613, 1748, 2124, 3156, 868, 2125, 907, 1236, 869, 2129, 1522, 2930]. Notes [18, 20, 47, 73, 80, 97, 105, 124, 134, 115, 143, 150, 160, 191, 210, 221, 232, 253, 263, 294, 273, 293, 305, 317, 326, 339, 362, 364, 374, 393, 407, 421, 436, 451, 502, 521, 541, 561, 595, 635, 651, 667, 680, 698, 715, 733, 751, 768, 783, 801, 820, 836, 872, 891, 910, 926, 943, 962, 986, 1008, 1025, 1045, 1070, 1091, 1112, 1133, 1155, 1188, 1205, 1223, 1239, 1273, 1291, 1325, 1340, 1360, 1375, 1397, 1467, 1480, 1497, 1515, 1532, 1552, 1568, 1583, 1600, 1616, 1633, 1655, 1653]. Notes [1674, 1741, 1763, 2340, 330, 1486]. Notice [29]. notion [1181]. Noumerov [1062, 1336]. Noumerov-type [1336]. NP [1447, 2252, 1295]. NP-complete [1295]. NP-completeness [1447]. NP-stability [2252]. NU [3464]. Null [2600, 3754, 2321, 1968, 3647]. Number [535, 2570, 465, 415, 402, 1828, 884, 1281, 2330, 2920, 1904, 1711, 1923, 2369, 1586, 875, 1126, 2180, 859, 1391, 1967, 1896, 331, 1776, 1807, 1716, 2622, 2338, 1490, 1248, 1925]. Numbers [2695, 575, 637, 736, 107, 68, 51, 86, 108, 621, 688, 334, 2741, 3616, 2880, 3421, 1163, 3629, 3702, 3331, 634, 2090, 1003, 2014]. Numerical [2713, 294, 2364, 2377, 2415, 2545, 2683, 49, 2280, 582, 718, 1148, 470, 736, 597, 352, 418, 454, 565, 618, 443, 2685, 3108, 472, 3226, 227, 279, 313, 404, 462, 3566, 2520, 2120, 297, 2, 298, 1677, 546, 639, 686, 2994, 2825, 204, 620, 2701, 536, 2242, 299, 3501, 3373, 567, 194, 380, 1286, 2799, 2568, 672, 712, 1895, 537, 266, 183, 3409, 3478, 3438, 7, 394, 355, 2909, 2554, 3573, 2710, 2515, 138, 624, 268, 724, 2818, 185, 3607, 344, 401, 447, 607, 3233, 16, 3065, 370, 3074, 2022, 692, 2557, 3040, 847, 448, 1820]. Numerical [727, 61, 2983, 516, 794, 608, 2558, 3027, 628, 2807, 187, 2794, 3144, 967, 3241, 3477, 2718, 3597, 3586, 1354, 1027, 2278, 2065, 2393, 3058, 2402, 2672, 857, 1054, 1230, 2921, 2797, 3280, 3198, 2862, 2644, 1749, 2282, 1402, 3307, 3367, 1437, 2004, 1431, 3030, 2083, 3211, 3732, 1176, 2250, 2284, 1101, 990, 2296, 2475, 3530, 2756, 1611, 3462, 3395, 2935, 1662, 3254, 1930, 2775, 1530, 1507, 844, 2133, 3323, 3007, 1284, 1144, 3637, 860, 3725, 3713, 861, 3334, 2812, 3598, 3544, 1081, 3702, 772, 3230, 3592, 2816, 3423, 1315, 1213, 3153, 2060, 1563, 2924, 3128, 828]. numerical [2979, 1033, 1316, 3023, 3155, 3302, 978, 1217, 1885, 2764, 1109, 3189, 2195, 903, 1042, 1372, 3012, 1613, 3638, 1004, 148, 3535, 1082, 2301, 156, 2449, 980, 849, 3315, 3521, 3468, 3425, 3142, 2986, 2426, 2068, 1848, 2843]. Numerically [3419, 441, 2403, 2580, 1832, 2927]. numerische [148]. Numerov [1528, 1355]. Numerus [476]. NURBS [3045, 3633]. NURBS-shaped [3633]. Nyström [3293, 2265, 1947, 1195, 1505, 1231, 2070, 3229, 2167, 2463, 3352, 3054, 1914].

O [1220]. O.D.E [686]. O.D.E. [850]. O.D.E.s [857]. Obituary [3261]. Object [153]. objects [2141, 3173]. oblique [3742]. Obreschkoff [3630]. observation [1923]. observations [2754, 2466, 1412, 3688, 3310]. obstacle [3172, 3542, 3176, 3396]. obtain [1904]. Obtained [1660, 3340]. occurring [919, 1848]. October [1703]. Odd [672, 3345, 320]. ODE [2439, 2976, 2067, 3708, 1286, 1265, 1299, 2384, 1543, 3607]. ODE-solvers [1543]. ode23 [2313]. ODES [1034, 2102, 1403, 3668, 3544, 3640, 2586, 2290, 2789, 1992, 1250, 2019, 2472]. ODEs/DAEs [3668]. ODR [3045]. Off [442, 67, 453, 685, 758]. oil [1388]. om [38]. On-Line [186, 1656, 277]. One [3059, 3211, 657, 528, 562, 2594, 624, 247, 301, 1201, 497, 866, 3683, 3344, 2859, 1512, 3513, 2933, 1749, 3107, 1644, 2773, 3186, 1846, 1883, 1469, 2088, 3592, 2410, 3376, 3531, 3127, 3413, 2999, 3377, 1597, 431, 2779, 2125, 999, 1573, 3027, 1053]. One-Bit [866]. one-dimensional [3683, 3186, 3592, 2125]. one-half [3344]. One-Leg [2594, 1883, 2779, 3027]. one-parameter [2088]. one-parametric [999]. one-pass [1469]. One-sided [3211, 3376, 3531]. One-stage [3059]. One-Step [657, 497, 1201, 1749, 3107, 1644, 3377, 1053]. one-way [1512, 1573]. online [3696]. only [1640, 1696, 1488, 809, 846, 890, 2302]. Ono [2675, 2301]. onto [2744, 2239, 1277]. Open [2606, 557, 1672]. Operand [161]. Operating [587]. operational [3207]. operations [1365]. Operator [11, 224, 2409, 3685, 3056, 3635, 2955, 2894, 3048, 3589, 1352, 3507, 3287]. operator-based [3685]. Operators [530, 659, 3224, 2428, 3183, 3441, 3641, 3513, 3213, 2933, 1676, 2522, 2846, 2875, 2956, 3138, 2121, 2224, 2288, 2193, 2181, 2168, 3415, 3000, 1973, 2504]. optical [1705]. optics [2718]. Optimal [2043, 669, 1641, 2859, 1281, 2330, 2366, 526, 565, 739, 1981, 1711, 2323, 1783, 2040, 3374, 3397, 3136, 602, 586, 1174, 2734, 3139, 1115, 2438, 1824, 2462, 2978, 2073, 2028, 554, 3698, 572, 574, 1639, 1036, 3510, 1878, 2375, 966, 2989, 2235, 3661, 1669, 824, 3742, 3500, 3487, 3180, 1823, 3032, 1265, 2251, 2484, 2948, 1855, 3581, 2000, 1890, 1587, 1990, 2461, 3608, 792, 3219, 1109, 2622, 1841, 2212, 3315, 1022, 2214]. Optimality [1922, 1299, 2585, 952, 1000, 3705]. Optimally [2968, 3359, 3588]. optimisation [3490]. Optimization [2603, 82, 842, 205, 247, 3599, 1280, 2456, 1860, 2645, 3308, 2086, 3137, 951, 3310, 2371, 3560, 2351, 3470]. optimize [772]. Optimized [2164, 3295, 2507, 2065, 2720]. optimizing [1586, 2653]. Optimum [175, 1453, 1771, 228, 109, 1010, 1648]. option [3165, 3401]. options [3287]. orbits [2270]. Order [3029, 580, 582, 3208, 1947, 1540, 471, 203, 297, 3582, 2567, 2697, 67, 319, 2324, 458, 2585, 356, 2711, 102, 1192, 608, 954, 2703, 3105, 2892, 928, 2043, 3043, 2615, 3690, 2794, 3355, 3159, 1280, 1927, 3512, 2306, 3207, 2215, 2392, 2507, 3133, 284, 3184, 945, 1120, 1575, 2246, 1477, 2112, 3306, 2920, 3069, 3555, 2307, 1904, 1054, 1230, 3700, 1282, 858, 839, 969, 1125, 1162, 1320, 1336, 1642, 3146, 3738, 1252, 912, 3735, 2945, 3299, 2944, 3443, 3072, 1487, 2946, 3593, 1829, 874, 990, 2895, 2995, 1771, 797, 3589, 914, 1063, 1087, 3657, 2907, 3283, 788]. order [2333, 1530, 3615, 3620, 957, 1080, 1526, 1663, 1410, 808, 2191, 3628, 1016, 1442, 1713, 1775, 2437, 3032, 1989, 3136, 2172, 861, 1975, 2898, 1793, 2205, 3095, 3375, 3471, 3544, 3096, 1475, 3743, 2542, 3152, 2460, 2410, 3113, 2151, 2676, 2411, 3505, 1990, 2867, 2167, 2289, 1386, 2021, 3577, 3578, 3558, 846, 1419, 2725, 1097, 3155, 2465, 1151, 2229, 1885, 1449, 2819, 3440, 1460, 3585, 830, 2387, 3303, 1368, 1626, 1720, 3454, 2125, 2155, 2449, 1201, 849, 3378, 3523, 1105, 3142, 3674, 3115, 3320, 3524, 3536, 3379, 1053]. Ordered [318, 1557, 1759, 1482, 1785, 1999, 1312]. Ordering [264, 351, 1888, 2104, 1784, 3390, 2464]. orderings [3406, 1545]. Orders [2701, 2126, 968, 1730, 2190, 1867]. Ordinary [487, 2745, 203, 275, 2458, 307, 2688, 185, 3043, 3159, 1027, 3602, 904, 2393, 2246, 2033, 2651, 3526, 3210, 2381, 1101, 1057, 841, 1662, 1232, 2479, 2286, 2525, 2885, 2876, 3034, 3095, 2948, 1246, 1308, 2463, 2788, 2765, 1249, 1318, 1035, 2449, 1201, 3618, 2791]. ordinates [3199]. Organization [677]. orientation [74, 3258]. Oriented [382, 2038, 2026]. orientering [74]. original [1210]. Orthogonal [2660, 1879, 2964, 3380, 2418, 2044, 2052, 3703, 2034, 1253, 1405, 2018, 2514, 2595, 3127, 3038, 2679, 2337, 3722, 2048]. Orthogonality [3452, 2835]. Orthogonalization [222, 369, 2574, 2134, 3065]. orthonormal [3209, 1420, 2074]. orthonormality [2203]. Oscillating [548, 2785, 2462]. oscillation [975, 1720]. oscillations [2597]. oscillator [3189, 2790]. oscillators [1664, 3298]. Oscillatory [2685, 712, 2667, 447, 1926, 2858, 3144, 3526, 3198, 3579, 1816, 3332, 1813, 2525, 2646, 2761, 2876, 3034, 2676, 2899, 3398, 3532, 2818, 2927, 1834, 864, 3181, 3193, 3521, 3676, 2807, 3077, 3320]. Osculatory [694, 826, 1288, 1628]. Oseen [3005, 3508]. Ostrovsky [3297, 3424]. Other [411, 884, 2111, 2881]. Outer [359, 1830, 1805]. outerplanar [1810, 1474]. Outline [2719]. Output [197, 3099]. Overdetermined [700, 2273]. Overflow [1190, 1382, 1998]. Overlapping [2535, 3486, 3527, 3427]. Overrelaxation [59, 1771, 2287, 1107, 1214, 1545].

p [81, 562, 394, 135, 543, 1785, 3741].

P-spline [3741]. P. [2452]. Paardekooper [2108]. package [3638]. packed [1586]. packets [2207]. packing [1586]. Padé [791, 3729, 3016, 1529, 1946, 867, 1506, 2632, 905, 2164, 1040]. Padé-based [3729].

Padé-type [1529, 2632]. Page [1966, 2774]. paged [899]. Paige [3659]. Pair [1554, 416, 1670, 2026]. pairs [3295, 2365, 3111, 2269, 3747, 2214]. pairwise [3696]. Palindromic [3597]. Pantograph [2710, 3395]. Paper [83, 2452, 547]. Papers [2413, 2768]. Parabolic [2726, 472, 2323, 306, 659, 2671, 2577, 3684, 2615, 2942, 2218, 2365, 3068, 2306, 3196, 2531, 2905, 2810, 3529, 3613, 2442, 3732, 3646, 3707, 886, 905, 3492, 1930, 2512, 3148, 2775, 1507, 1144, 2674, 2460, 2866, 3486, 3724, 3650, 3475, 1594, 2061, 2289, 2021, 2978, 3548, 2841, 1020, 2589, 3316, 2426]. paradox [1938]. Parallel [2601, 1879, 2330, 1555, 1536, 2067, 1929, 1746, 2569, 1747, 3655, 2350, 2670, 2030, 1225, 1452, 1971, 3225, 1688, 3624, 1500, 1635, 2976, 2521, 1829, 1432, 1361, 1378, 1726, 1967, 1202, 1146, 3743, 2451, 1607, 3390, 1954, 1985, 1472, 1503, 2245, 2681, 1992, 1365, 1381, 1768, 1383, 1522, 1878, 3379]. parallel-in-time [3624]. parallelizing [2001]. Parameter [2732, 2979, 3302, 206, 2109, 3596, 2773, 3266, 3605, 2482, 1286, 2484, 2088, 3653, 3247, 3322, 3249, 2587, 3233, 3433, 2918, 1035, 2260]. parameter-robust [3653]. Parameter-uniform [2979, 3302]. parameterized [3421, 3248]. Parameterizing [590]. parameters [3002, 779, 3226, 3149, 2115, 3481, 2274, 3313]. Parametric [2694, 2567, 641, 3342, 2536, 1063, 1525, 3547, 2840, 3024, 3413, 1867, 2169, 3339, 999, 2541]. parametrization [3050]. parareal [3676, 3624]. Parent [1553]. Parsers [543, 513, 660]. Parsing [295, 1279]. Part [2633, 531, 2876, 3034, 2542, 2937, 2042, 1165, 2842, 3295, 1688, 505, 2477, 2537, 506, 87, 91, 953, 973, 1755, 1777, 2789, 2890]. Part-Compilation [531]. Partial [264, 351, 1799, 756, 740, 1940, 1228, 1538, 1473, 1840, 2711, 2780, 2671, 1953, 3068, 1574, 989, 3069, 3623, 2054, 2772, 3071, 1055, 886, 2368, 3335, 2773, 2784, 2775, 3672, 84, 3570, 60, 1300, 1115, 2348, 1943, 1333, 3102, 3723, 3041]. Partial-Match [1840, 1228, 1115]. Partially [318, 1999]. Particle [2549, 3599, 3385, 3592, 3625]. Particular [538, 2981]. partielle [84, 60]. Partition [644, 314, 1603, 806, 1609, 1658, 1578, 1839, 1911, 3287]. Partitioned [2976, 2473, 2548, 2550, 3744, 3355, 1120, 2444, 1835, 3540]. Partitioned-GMRES [2473]. Partitioning [350, 488, 3751, 1843, 1226, 1314, 1517, 1983, 1348, 1607, 2765, 2939, 1559]. Partitions [1011, 884, 875, 3168, 1619, 3074, 3202]. Parts [85, 52, 3267, 3156, 2639]. Pascal [865, 868, 1076]. pass [1469]. passage [3586]. Passive [2567]. Patankar [3375]. Patch [2244]. Patch-adaptive [2244]. Path [2886, 1887, 1128, 1244, 1041, 2351]. Paths [217, 1954, 1925, 1251]. Pathwise [2458, 2885, 3274]. pattern [1783]. Patterns [1660, 1787]. PCG [1789, 2105]. PDE [2645, 2311, 3581, 3310, 3723]. PDE-constrained [3310, 3581]. PDE-greedy [3723]. PDE-model [2645]. PDES [2399, 2604, 3382, 3584]. Peano [2797]. peer [3485, 3359, 3412, 2681, 3099, 3240]. Pegasus [603, 527]. penalty [3667, 2273, 1866]. pencil [931]. Pencils [2376, 3120, 3123, 3702, 2249]. penetrable [3172]. penetration [2844]. Penrose [3040]. people [10]. percent [1095]. Perfect [1969, 1826, 1435, 1725, 2011]. perfectly [3096, 3361]. Performance [1603, 1851, 2012, 1840, 3685, 1343, 2039, 1432, 1768]. Period [672]. Periodic [641, 427, 672, 2702, 2270, 1661, 2966, 1398, 1371, 2922, 3296, 1403, 2655, 2617, 2826, 2935, 1930, 1016, 1691, 2923, 1931, 2852, 3491, 2655]. periodic-linearly [1931]. periodic-to-periodic [2655, 2617]. periodicity [1195]. permanents [1680].

Permutation [508, 1690, 1986, 2030]. permutational [1996]. Permutations [705, 356, 69, 1803, 1536, 1854, 1366]. permutative [1485]. Perron [971]. Perspectives [2843]. Perturbation [1812, 1903, 2494, 3436, 2922, 2497, 2163, 2396, 2006, 3612, 1332, 2925, 385, 2339, 2048, 1901, 479, 573, 2794, 3262, 1892, 2631, 2861, 3372, 2045, 3541, 2295, 827, 2251, 3517, 2020, 2206, 2461, 2438, 3127, 1380, 2047, 1993, 2212, 3076, 2127, 3205]. perturbations [3685, 2629, 2336, 2878, 3127, 2124]. perturbed [2391, 3226, 3017, 2970, 2851, 3677, 2801, 3327, 3010, 2978, 2979, 3302, 2229, 1820, 2918, 2638]. Peter [1703]. Peterson [1672]. Phase [1505, 1368, 3538, 2605, 1885]. phase-fitted [1885]. Phase-lag [1505]. phenomenon [2964]. photonic [3174, 3175]. PI [1662]. Picard [3526, 3265, 1861, 2688, 1977, 1755, 1777, 2702]. Pictorial [1330]. pieces [1586]. Piecewise [2661, 2044, 3496, 3084, 2268, 3285, 3464, 2338, 3170, 3192, 3689, 3711, 2791]. Piper [3026]. Pivoting [2662, 1444, 2496, 2368, 2784, 3086, 2451, 2255, 2588]. Place [508, 1159, 3639]. placement [1824]. places [1453]. Plana [2219, 2248, 2322]. Planar [2565, 3280, 3334, 2813, 1954]. planarity [1209]. Planck [3586, 2723]. plane [1501, 1585, 3732, 1525, 2041, 1867, 1810]. Planning [85, 41, 358, 970]. plate [3396, 2845]. plateau [3085]. Plotter [570]. plus [2423, 2998]. Poincaré= [3448]. Poincaré=-Friedrichs [3448]. Point [27, 2518, 685, 144, 2617, 275, 12, 28, 127, 24, 643, 706, 147, 225, 414, 2272, 386, 1523, 254, 800, 755, 3305, 1085, 3642, 2603, 3687, 3243, 1054, 1230, 858, 839, 969, 1125, 1320, 1540, 1642, 2236, 2533, 1057, 3244, 758, 3450, 1440, 3282, 3149, 957, 3358, 3408, 3255, 1663, 3246, 3216, 3331, 3247, 3248, 3611, 3609, 1543, 2979, 3258, 3025, 3235, 1717, 1392, 3289, 1211, 1082, 1444, 3252, 2655].

Point-In-Polygon [225].

Point-to-periodic [2617, 2655]. Pointers [1746]. Points [654, 704, 490, 266, 2573, 586, 724, 898, 1501, 2761, 331, 1362, 2927, 1666, 1833, 1109, 1337, 2259, 1236, 1251, 1889]. Pointwise [3551, 2707, 2949, 3353]. Poisson [3381, 2871, 3182, 3554, 3280, 3046, 2967, 3443, 3019, 2775, 1080, 3518, 1416, 3689, 2831]. Poisson-type [1080]. Pol [3312]. Polar [2749, 1812, 2631, 2020, 2206]. pole [3494, 975, 1884]. poles [1640, 1696, 1770, 1526, 1774, 1733, 845, 1991, 846, 2302]. Policy [175]. Pollution [440, 1424]. Pólya [547]. Polygon [1618, 3750, 225, 1853]. polygonal [2726, 3698]. polygons [2937, 1619, 1329, 2805, 1891]. Polyhedral [557, 816, 1872]. Polylogarithms [629, 537, 394, 355, 2459]. Polynomial [3223, 2661, 739, 2836, 638, 720, 2706, 2222, 600, 722, 146, 1826, 1782, 2094, 1971, 3715, 3383, 2188, 3446, 3345, 955, 2632, 3566, 3541, 1609, 1014, 3135, 2481, 2298, 2311, 1202, 1146, 1147, 1088, 3285, 1949, 2166, 3125, 2649, 1108, 2917, 1579, 976, 1420, 1544, 1716, 1549, 2185, 847, 3402, 3711, 2791, 3078]. Polynomials [2562, 524, 377, 1253, 2340, 250, 2375, 3380, 856, 2034, 3345, 2002, 981, 2189, 2440, 2005, 2345, 2873, 3346, 3084, 1882, 1405, 2018, 3358, 3408, 2514, 1527, 1149, 2542, 3504, 2618, 3125, 3038, 3351, 923, 2217, 1217, 880, 1668, 1490, 1848]. Polyphase [195]. polytope [787, 3373]. polytropic [3468]. poroelasticity [3690, 3360, 3520]. porous [3004]. port [3668]. port-Hamiltonian [3668]. Portable [2446, 1193]. Posed [2380, 2687, 2732, 2621, 3525, 2871, 2943, 3370, 3369, 1697, 2095, 2495, 3122, 1844, 2165, 2480, 1861, 2972, 2609, 2997, 3322, 3073, 2611]. posedness [3256, 3672, 3592, 3475, 2327, 3737]. positioning [2605]. Positions [490]. Positive [2340, 2752, 1736, 2835, 1710, 3701, 902, 2085, 3375, 2398, 3249, 972, 1815, 1668, 2036, 3633, 1199, 1248, 2639, 3403, 2626].

Positivity [3670, 3557, 2516, 2803, 1668, 2779, 3182, 3423]. Positivity-preserving [3670, 3557]. possessing [831]. Possibilities [153]. Possible [646, 215, 1087]. Post [180, 528, 562]. Post-secondary [180]. posteriori [3493, 3344, 3419, 3527, 3213, 3462, 2916, 3677, 3645, 2866, 3097, 3322, 3548, 3098, 2845]. posteriori-based [3493]. posteriori-bounds [3419]. Postfix [310]. potential [2894, 2889]. potentials [3151, 3363, 3181]. Powell [3281, 3168, 3074, 3202]. Power [1084, 204, 3515, 236, 707, 198, 1996, 1430, 1176, 1408, 2908, 1931, 2181, 1380]. Powers [511, 746]. pp [27, 188, 637, 1398, 1320, 1629, 517, 28, 120, 1089, 2316, 1202, 498, 499, 1308, 890, 1630, 1651, 1002, 465, 1964, 1338]. Pracniques [3749]. Practical [3295, 2591, 235, 369, 571, 1829, 2097, 1665, 1757, 1181]. practice [3057, 2412]. PRAM [1878]. Precedence [1639, 1005]. Precise [1753]. Precisely [2594]. Precision [739, 172, 147, 664, 3225, 2176, 1488, 1145, 3639, 2448]. precision- [3639]. Preconditioned [1798, 2416, 1781, 2695, 1795, 2712, 1436, 1791, 2439, 2604, 2104, 1784, 2783, 2609, 1790, 2151, 2385, 1418, 2245, 2930]. preconditioner [3493, 3243, 3145, 1970, 3701, 2239, 2150, 3246, 3376, 3531, 3268, 2973, 2623, 3252, 3563]. Preconditioners [2696, 2366, 2673, 2747, 2555, 2349, 2280, 2065, 1928, 2238, 2532, 2534, 2431, 1783, 3019, 2298, 2311, 3556, 1989, 2085, 2077, 2288, 3310, 2487, 2998, 3559, 2982, 3508, 2400, 2842]. Preconditioning [2561, 2147, 2608, 2524, 2253, 2099, 1782, 3404, 2344, 3554, 1792, 3704, 2772, 990, 2149, 3005, 3112, 1794, 3653, 3611, 3609, 3411, 2637]. preconditionings [1561, 1793, 2448]. predator [3557]. predator-prey [3557]. predecessor [2900]. predicates [3642]. prediction [2644, 1402, 2764]. predictive [3433]. Predictor [509, 3199, 1035, 1424]. Predictor-Corrector [509, 3199, 1035, 1424]. Preface [1951, 2658, 3241, 3206, 3276, 3292, 3317, 3171, 3343, 1936, 2728, 2781, 2879, 3015, 3067, 3079, 3092, 3104, 3143, 3130, 3158, 2717, 2493, 2753, 2742, 1781]. Prefilters [2506]. Preparation [118]. preparations [901]. prescribed [2421]. prescriptive [1695]. Presence [2362, 3673, 2115, 1130]. Presentation [83]. preservation [3641]. Preserve [2708]. Preserving [2705, 2565, 2800, 2776, 2422, 1192, 3488, 2932, 3473, 2967, 3514, 3030, 3670, 3047, 3694, 3324, 3186, 2203, 3557, 2721, 3284, 3200, 3423, 2888, 3154, 2539, 1667, 2790, 3521, 3078, 3595, 3666]. presortedness [1708]. Pressure [3706, 3737]. Pressure-robust [3706]. Previous [621]. prewhitening [2267]. prey [3557]. Pricing [3401, 3165, 3287]. primality [1976, 1127, 2092]. primary [3167]. Prime [575, 637, 286, 100, 621, 815, 334, 2330, 1711, 763, 832, 634, 1490, 1270, 1338]. Primes [535, 3, 482, 621, 384, 851, 838, 591, 1678]. Primitive [320, 1937]. primitives [1193]. Principal [1657, 401, 1576, 2282, 1029, 1774, 1756, 2117]. Principle [2709, 2824, 1821, 994]. Principles [345, 3118, 3613]. printer [53]. Printout [175]. priori [1523, 3228, 3219, 3522, 2845]. Priority [745, 1996, 1557, 935]. Prize [1496, 1655, 1920, 1836]. Probabilistic [1921, 1127, 1293, 3022]. probabilities [3506]. probability [1873, 1216]. probing [995]. Problem [2693, 2547, 3754, 2321, 704, 75, 462, 285, 567, 1646, 2666, 183, 642, 321, 382, 138, 664, 2669, 589, 387, 45, 1384, 939, 3587, 1452, 3118, 2293, 1921, 2306, 2494, 1010, 1148, 2579, 3715, 2235, 2954, 989, 3436, 1540, 1642, 3256, 787, 3529, 2508, 3659, 1893, 1844, 1327, 1310, 1447, 2394, 2673, 2995, 1680, 1908, 3450, 3396, 3693, 878, 3282, 3667, 3019, 1930, 3085, 3386, 2057, 2957, 1745, 3374, 1138, 1293, 3088, 3478, 1414, 1832, 2251, 1178, 3216, 1924, 2801, 1332, 1367, 3326, 2438, 1649, 3073, 3126, 3022, 952, 1991, 1179, 3258, 3465, 3399, 1472, 829, 1715, 1728, 1117]. problem [2048, 1820, 1322, 2124, 2212, 2940, 3435, 1858, 2918, 3737, 1212, 1381, 3454, 849, 1423, 1729, 1932, 1474, 1022, 3337, 2049, 3610, 1559, 1986, 2031, 3205]. Problem-Oriented [382]. Problems [453, 113, 122, 131, 142, 151, 158, 166, 178, 189, 201, 208, 219, 230, 242, 251, 261, 271, 280, 291, 303, 315, 324, 337, 347, 360, 372, 391, 405, 419, 434, 449, 463, 483, 500, 518, 539, 559, 577, 593, 612, 632, 649, 665, 678, 696, 713, 731, 749, 766, 781, 799, 818, 833, 852, 870, 888, 908, 924, 941, 960, 984, 1006, 1023, 1043, 1068, 1090, 1110, 1131, 1153, 1168, 1186, 1203, 1221, 1237, 1256, 1271, 1289, 1306, 1323, 1339, 1358, 1373, 1394, 1450, 1465, 1479, 1494, 1514, 1531, 1550]. Problems [1566, 1581, 1599, 1614, 1631, 1652, 1673, 1702, 1721, 1739, 1762, 701, 222, 2380, 2563, 3755, 472, 65, 2687, 276, 4, 489, 2323, 299, 2572, 2732, 2666, 2568, 673, 84, 687, 2434, 2689, 414, 2621, 2373, 646, 1017, 2327, 358, 608, 627, 2741, 2559, 168, 3685, 2892, 3683, 3599, 3614, 2615, 2032, 2794, 1399, 1661, 800, 1887, 1425, 2858, 1888, 755, 2186, 3159, 967, 1796, 2844, 3242, 3525, 2365, 2871, 1476, 2966, 2943, 3207, 2718, 3370, 3369, 2603, 3552, 3669, 2468, 2990, 3196, 3703, 2580, 2281, 2495, 2531, 3279, 3555, 3243, 3093, 2881, 1054, 1230, 1398, 1371, 858]. problems [839, 969, 1125, 1320, 1585, 2308, 3420, 2331, 1860, 2177, 2430, 1402, 3226, 2017, 2236, 3701, 2955, 3072, 2946, 2533, 3122, 3646, 2178, 3061, 840, 1264, 1404, 2510, 2894, 2993, 905, 3356, 2523, 1057, 3325, 3482, 3542, 3244, 3492, 3228, 3177, 3017, 2405, 2970, 1909, 2180, 2130, 1440, 2478, 2407, 1063, 1699, 1732, 3619, 2165, 2480, 2512, 1861, 2233, 2645, 2785, 3005, 2972, 1507, 3179, 957, 3180, 933, 1744, 3111, 3087, 3397, 3733, 1883, 3007, 1663, 2609, 1016, 1713, 1775, 1488, 1051, 2150, 2136, 2851, 2483, 2886, 2997, 1700, 3745, 3739, 1413, 3629, 1475, 2949, 3677, 2460, 3645]. problems [2287, 2816, 3113, 3247, 3724, 3248, 3581, 3097, 3338, 3267, 3611, 3609, 3231, 2461, 2867, 3268, 2167, 1417, 2978, 2611, 2925, 828, 2253, 2979, 3352, 1548, 1666, 937, 1033, 1455, 3457, 2828, 3233, 2274, 3302, 2168, 1885, 2063, 2183, 1717, 1852, 3286, 2047, 1197, 3601, 2789, 2890, 2184, 3401, 2841, 1613, 2228, 3453, 2258, 3289, 1211, 1082, 2638, 1720, 1295, 2640, 3103, 3523, 2806, 778, 3041, 3222, 3691, 3316, 3469, 3562, 3510, 3252, 3353, 3491, 3550, 3379, 1424, 1868, 1876, 1848]. Procedure [1799, 211, 162, 590, 2240, 999, 2314, 2854, 1988]. Procedures [517, 207, 6, 215, 70, 1160, 770, 1296, 1297, 788, 765]. Process [673, 369, 2455, 3748, 1524, 992, 3215, 3423, 2386, 3506, 2154, 3649]. Processes [455, 598, 693, 538, 3621, 1440, 1530, 915, 1004, 980]. Processing [152, 94, 8, 14, 128, 15, 42, 200, 1639, 134, 10, 1572, 273]. Processor [295, 653, 365]. processors [1015, 1294, 1841, 1193]. Procrustes [1646, 2048]. producing [2141]. Product [424, 470, 583, 3635, 757, 601, 359, 459, 1595, 1899, 2805, 2417, 2494, 1335, 2111, 1263, 806, 1922, 1576, 1331, 2873, 1516, 913, 1468, 3742, 876, 1089, 1121, 2407, 2827, 2486, 1513, 1756, 3415, 3065, 3025, 1907, 1933]. Product-integration [757]. product-product [1933]. product-sum [806]. Product-Type [583, 1263]. Production [296, 85, 52, 970, 3375]. production-destruction [3375]. Products [456, 3725, 3713, 3235, 3191]. Program [439, 440, 741, 285, 153, 233, 277, 586, 587, 1969, 512, 951, 1260, 1455, 3457, 920, 998, 795]. programmes [1866]. Programming [274, 71, 36, 62, 88, 188, 63, 704, 144, 517, 1679, 276, 89, 241, 4, 112, 120, 207, 336, 473, 257, 130, 1174, 128, 622, 498, 499, 95, 129, 121, 135, 237, 332, 1116, 314, 495, 247, 301, 514, 1067, 200, 250, 156, 590, 149, 895, 9, 141, 254, 3242, 260, 78, 389, 786, 177, 770, 359, 2086, 1700, 919, 952, 1000, 218, 1683, 1322, 999, 1193, 1182, 1076, 2276, 1559]. Programs [654, 740, 1114, 368, 225, 1118, 710, 572, 1688, 2012, 1379, 2013, 764]. progression [851]. progressions [838]. Projected [3245, 1918, 2351]. Projection [2744, 2932, 687, 3291, 1921, 988, 3213, 3325, 2408, 2675, 3340, 3722, 3453, 3671]. projection-based [3340]. Projections [1879, 1830, 2274, 2337, 2208]. projector [3124]. projector-splitting [3124]. projectors [3163]. Proof [508, 213, 2062, 199, 310, 1267, 1939, 1073, 1234, 1410, 1210, 1738, 976, 780]. proofs [2042]. Propagation [723, 2278, 2630, 3410, 2727]. proper [978]. properly [2887]. Properties [617, 583, 719, 2549, 2567, 641, 2288, 238, 1330, 775, 516, 371, 2102, 2330, 2872, 3300, 876, 1089, 3007, 2312, 2596, 1817, 2776, 3411, 2035, 3431, 3415, 1650, 2153, 3203, 1368, 2589]. Property [737, 674, 2071, 2198]. proportional [2387]. Proposed [743, 1709]. proposition [1944]. propositional [1942]. Propositions [1379]. Prosessdataenhet [94]. Protection [1554]. protocol [2831, 1877]. protocols [1139, 950]. Proving [368, 1427, 764]. Proximal [3636, 2236]. Pseudo [2292, 107, 402, 573, 2315, 1828, 2330, 1859, 1711, 2645, 3329, 3350, 3401]. Pseudo-Inverses [573]. Pseudo-Random [107, 402, 1828, 1711]. pseudo-spectral [3401]. pseudo-splines [3329]. Pseudo-symplectic [2292, 2315, 3350]. pseudo-timestepping [2645]. pseudoinverse [988, 1264, 2283]. pseudoinverses [2146, 2396, 2091]. pseudorandom [1281]. Pseudospectra [2174, 2430, 2249]. Pseudospectral [3599, 2468, 3534]. pseudospectrum [2175]. PSS [3252]. Pták [1869, 3050]. public [134, 956, 1002, 1037]. public-key [956, 1002, 1037]. Publish [3484]. Publisher [2261, 2543]. Pulawy [256]. Pulp [66]. pure [2719]. purely [1686]. Pyramids [865].

Q [3227]. Q-MR [3227]. Q-OR [3227]. QD [1643]. QD-like [1643]. QR [2992, 2457, 2484, 2360, 2738, 2376]. quad [3667, 3435]. quad-curl [3667, 3435]. Quadratic [2054, 505, 1845, 456, 3319, 320, 1174, 506, 553, 1666, 258, 288, 2669, 3242, 2859, 2082, 1370, 3279, 2002, 3459, 2247, 2532, 3071, 2906, 3111, 2286, 1896, 3629, 2909, 3125, 3390, 3560, 3074, 1197, 1236, 1627, 3041, 3337, 3510]. quadratic-like [2002]. Quadratic/linear [3319]. quadratically [1972, 2510]. Quadrature [116, 663, 424, 470, 583, 2904, 618, 203, 694, 2566, 1912, 2701, 640, 3621, 194, 2707, 343, 380, 2667, 445, 549, 602, 2876, 3034, 723, 2688, 2648, 344, 3363, 1199, 259, 216, 2740, 1523, 3606, 2858, 2988, 1369, 1770, 3576, 1263, 2860, 3626, 1268, 2797, 826, 1577, 1029, 1751, 1881, 2097, 1541, 1671, 2370, 3332, 1298, 2498, 1184, 3617, 2241, 876, 1089, 1121, 1813, 1526, 2106, 1492, 1441, 3428, 3035, 1917, 2252, 1303, 2950, 3652, 2516, 2926, 3257, 3351, 2073, 1165, 1815, 2352, 1849, 1288, 1628, 3272, 3678, 2829, 2960, 810, 831, 1198, 1255, 3012, 3203]. quadrature [1248, 2529, 882, 2171, 1167, 3634, 3489, 3681]. quadrature-type [1441]. Quadratures [366, 425, 526, 3500, 3036, 2151, 2677, 3239, 2767]. quadrilateral [2937]. quadtree [1966]. Quadtrees [1330]. Qualitative [50]. Quality [66, 3692]. quantification [3495, 3561]. Quantifying [3472]. Quantitative [992]. quantities [1073, 3366, 3291]. Quantum [3075, 3488, 2179, 2355]. quantum-classical [2355]. Quartic [86, 3108, 3000, 2960]. Quasi [2743, 2471, 3307, 2250, 2284, 2332, 172, 147, 3000, 2559, 2966, 3631, 3277, 3071, 3108, 3163, 2652, 3085, 3517, 2909, 3168, 3152, 2911, 2117, 2829, 2960, 3074, 3026, 3202, 3522, 3491, 3367]. quasi-convexity [2652]. Quasi-Graded [2743]. quasi-interpolant [2960]. quasi-interpolants [3071, 3108, 2909, 3168, 3152, 2911, 3074, 3026, 3202]. quasi-interpolating [3163, 2117]. Quasi-interpolation [3307, 3000, 3277, 3367]. Quasi-Minimal [2471]. Quasi-Monte [2250, 2284]. Quasi-Newton [2332, 2559]. quasi-periodic [2966, 3491]. quasi-plateau [3085]. quasi-Toeplitz [3631, 3517]. quasi-variational [3522]. quasilinear [3707, 2357, 2259]. quasilinearization [3437]. quasiseparable [3421]. quaternion [2610, 3728]. quaternionic [3346]. queens [1212]. queries [2040, 1957]. query [1094, 1364]. Question [2606]. Queue [745, 1996, 935]. Queues [269, 1557]. Queuing [2564]. Quick [1747]. quicksort [1512, 1432, 1573, 1878]. quintic [1663]. Quotient [3528, 1517, 2267, 2653, 2501]. Quotient-space [3528]. quotient-type [2501]. quotients [3397, 3733].

R [3464, 779]. R-K [779]. Rabin [1127]. Radau [2665, 2926, 3301, 850, 2022]. Radial [3287, 3070, 3437, 3617]. radiation [3255]. radii [3327]. radio [1744]. radius [3134, 3675, 1848]. radix [1778]. Raising [3152]. Ramanujan [2459, 556]. Random [423, 3529, 2458, 2747, 107, 136, 171, 402, 1828, 2330, 2672, 1711, 3462, 1690, 2775, 3721, 2885, 2897, 3516, 3662, 3481, 1097, 1041, 3234, 3354, 2622, 3425, 3298]. Randomized [3519, 3545, 3497, 2938, 3671, 3525, 3719, 3212, 3702, 2908, 3389, 3647]. Randomly [741, 760]. Range [322, 2040, 1019]. Ranges [228, 2068]. Rank [2267, 3620, 2570, 3413, 589, 3553, 3120, 3624, 3641, 3460, 3580, 3539, 3700, 2835, 2523, 2774, 3589, 3708, 3615, 3696, 3137, 3483, 3743, 3498, 2410, 3636, 3124, 2839, 3127, 3608, 3481, 3547, 2361, 2274, 3433, 3695, 2928, 3040, 3689, 2314, 3624]. Rank-adaptive [3620, 3580]. Rank-deficient [2267, 2523, 2274]. rank-one [2410]. rank-structured [3120]. Ranking [1322, 1548]. Rapidly [548, 1335, 2785, 2799]. RAT [1039]. Rate [2896, 306, 2897, 2439, 284, 3033, 3598, 3594, 3722, 2983]. Rates [2750, 805, 3458, 3654, 3721, 2674, 2485, 3475, 1834, 3156, 3315, 3467]. Rational [48, 3552, 33, 81, 353, 2582, 2846, 845, 3494, 2725, 2063, 626, 44, 3016, 1640, 1696, 2468, 3133, 955, 2003, 2256, 2798, 1079, 3593, 1219, 2266, 1163, 3094, 1039, 3319, 3537, 3502, 3035, 1285, 2654, 846, 975, 2254, 3680, 3414, 1598, 1667, 1736, 2169, 2036, 775, 2767, 3013, 869]. rationality [3114]. Raviart [3736, 2867]. Rayleigh [2653, 3010, 2501]. Razumikhin [3395]. Razumikhin-type [3395]. RBF [3209, 3686]. RC [245, 256]. RD [2654]. RD-rational [2654]. Re [369]. Re-Orthogonalization [369]. reaction [3614, 3445, 3304, 3161, 3226, 3374, 2958, 3166, 3677, 2801, 3302, 3523, 3157]. reaction-diffusion [3304, 3161, 2801]. reaction-subdiffusion [3523]. reactions [2723]. reactive [1692]. Read [3484]. Readability [1114]. Reader [76, 83]. Real [256, 1884, 725, 215, 1640, 1965, 2629, 3687, 2455, 2440, 2992, 1478, 1014, 3134, 2076, 948, 1441, 1733, 3331, 845, 2166, 3327, 1991, 121, 135, 809, 846, 890, 975, 2009, 2465, 2302, 2063, 1685, 2626]. real-pole [975]. Real-Time [256, 1965, 1685]. Received [560, 578, 1307, 1395, 1451, 613, 1801]. reciprocal [1804, 1253, 1088]. reciprocals [1560]. Recognition [350, 212]. reconfigurable [1984]. Reconstructing [1556]. Reconstruction [3271, 1584, 3090, 3452, 1758, 3288]. Record [1767]. Records [1746, 110]. recovered [2866]. Recovering [3541]. Recovery [1660, 2989, 3138, 3139, 3636, 2462, 1388]. rectangle [2440, 1649, 1117]. rectangles [1889]. Rectangular [109, 196, 1997, 3071, 3136, 2777, 2139, 1925, 3510]. Recurrence [496, 928, 3668, 2018, 3358, 1149, 879, 3506, 796]. recurrences [2164]. Recursion [703, 1937, 1941, 1693, 1598]. recursions [2279, 1215, 1269]. Recursive [350, 911, 1743, 551, 1473, 2390, 1888, 770, 2474, 760, 1806, 974]. red [1789]. red/black [1789]. Reduce [448, 1279]. Reduced [3228, 3564, 1789, 2895, 3302]. Reducibility [2762]. Reducible [1491, 1167, 1313, 1445]. Reducing [1617, 3749, 2597]. Reduction [438, 623, 2853, 2517, 3223, 2615, 1887, 3197, 3555, 3031, 3593, 3628, 2865, 2234, 2423, 2725, 3643, 3718, 3218, 2128, 2832, 3115]. redundancy [1444]. Reference [175, 866, 3706]. Refined [2666, 3111]. Refinement [260, 255, 265, 1743, 1913, 2038, 1846, 2069, 860, 1145, 3600, 1357]. Refining [153, 3123]. Reflected [1304, 1735]. Reflections [1059]. Regarding [575, 637, 634]. regime [2955]. region [2881, 3740, 2877]. Regional [358]. Regions [2546, 1556, 2568, 1282, 2131, 759, 1416, 1624, 3286, 1166, 2228]. Register [205]. Registers [465, 415]. Registration [2748, 10, 2593]. Regression [2694, 2660, 2466, 2308, 619, 1863]. Regressions [707]. Regula [580, 426]. Regular [481, 212, 3388, 3586, 2921, 2255]. regularisations [2881]. Regularity [3174, 3512, 3613, 3329, 2841, 3315, 3575]. Regularization [2563, 2572, 1772, 2484, 2621, 3237, 3673, 840, 2673, 3335, 3266, 3422, 3605, 2482, 2480, 1623, 2972, 1507, 3245, 3087, 2786, 3032, 2485, 3429, 3388, 3571, 2763, 2611, 3140, 2868, 3218, 3313, 3433, 3260]. regularization-correction [3673]. Regularized [3305, 2669, 3459, 2396, 2675, 3746]. regularizing [2915]. regularly [1075]. regulation [2475]. rehashed [1672]. Reissner [2845]. rejection [2456]. Related [2322, 32, 2378, 136, 195, 556, 3117, 2219, 2248, 1680, 2498, 2059, 3109, 3411, 3257, 1906, 1919]. Relation [2401, 1059, 3212, 1999, 2018, 1149, 3506, 3039, 1598, 1421]. relational [1968]. Relations [722, 928, 1676, 1609, 1108, 879, 1321, 1337, 1841, 796]. Relationship [2072]. Relationships [1612, 371]. Relative [3372, 2295, 490, 2206, 2878, 2630, 2045, 3228, 1822, 2127]. Relativistic [3344]. Relaxation [2606, 2147, 2017, 2535, 3244, 3492, 2720, 2814, 2274, 1963, 2194, 2196, 3157, 1458]. Relaxed [1797, 2296, 1830]. Relay [233]. Reliable [2343, 2454, 2860]. Remainder [165, 2677]. Remark [165, 517, 2316, 78]. Remarkable [86]. Remarks [223, 548, 1196, 3749, 1755, 1777, 1943, 648, 2339, 610, 1079, 956, 3746, 2078, 1002, 1037]. remedy [3144]. Remez [2917]. removal [2964, 3585]. Renormalization [2643]. Reorganization [1767]. repacking [814]. reparametrization [3661]. Repeated [1052, 571, 1730, 879, 1815]. replacement [771]. Replacing [1746]. Reply [28]. Representation [1478, 2567, 127, 77, 320, 146, 2571, 1330, 2966, 2580, 2955, 3281, 3270, 3269, 1244, 3507, 865, 1778]. Representations [340, 359, 511, 746, 446, 1686, 1953, 1588, 1348]. reproducing [3036, 3549]. Requirements [85, 52, 711, 3565, 970]. resampling [3441]. Research [378, 42]. Residuacity [236]. Residual [3426, 2471, 3727, 2420, 2732, 2397, 2750, 3416, 2297, 2268, 2760, 3008, 2786, 3009, 3431, 3403, 3205]. Residual-based [3426, 3205]. residuals [2948]. Residue [329, 342, 354, 859, 2300]. Residues [456, 1370]. resilient [3685, 3735]. resolution [2260]. Resolvent [3048, 2428, 1415, 1897, 2170]. resolvents [2181]. Resolving [3656]. resonance [3198]. resonances [3361]. Resource [711, 1235, 1334]. Resources [690]. respect [2860, 1923, 1253, 1967]. response [3118, 3205]. Restarted [3279, 2574, 3080]. restarting [3349]. restoration [3370, 3369, 3583, 3308, 3571]. restricted [855, 2607, 1295, 1036]. Restriction [2590]. result [1906, 2302, 1761, 869]. Results [575, 637, 718, 2045, 630, 2323, 1180, 3417, 1570, 2148, 2295, 3617, 906, 2313, 15, 1864, 2679, 634, 2047, 2154, 1597, 3315, 1218, 1106]. retarded [1262]. Retracing [3009]. retractions [3712, 3695]. Retrieval [756, 8, 14, 111, 1473, 1840, 788, 1115, 1228]. reuse [3565]. revealed [2360]. revealing [2915, 1734]. reversal [2961]. reverse [1050]. reversible [3362]. review [3485, 3189]. Reviews [835]. Revisited [2648, 1759, 278, 1860, 1506, 3214, 1095, 2370, 1562, 2762, 2865, 1258, 1117, 2830]. reweighted [1974]. rewriting [1807]. rewriting-number [1807]. Reynolds [174]. Rham [3162, 3449]. Rham-type [3162]. Riccati [3597, 2935, 2184]. rich [3585]. Richards [3004]. Richardson [3083, 1730, 67, 1524, 2154, 571]. Riemann [3450, 157]. Riemannian [3420, 3137, 3662, 3550]. Riesel [648]. Riesz [3146]. Right [99, 2374]. Right-Angled [99]. rightmost [2089]. Rigid [2326, 3270, 3269, 2090]. Rigorous [3242, 2446]. Ring [378, 3346, 1521]. risk [3433]. Ritz [3111, 2269]. RK [2229, 1061]. RKMK [2592, 2367]. RKN [2229, 3521]. RKS [2256]. Robert [3164]. Robin [2240, 3175]. Robin-to-Robin [3175]. Robust [2382, 3112, 3360, 3310, 2160, 3642, 2466, 3580, 3539, 3700, 2308, 3706, 619, 3743, 3653, 1863, 3013]. robustness [2368]. Rock [214]. Rodas5P [3638]. Role [1537, 2417]. Rollbacks [1617]. Romberg [120, 112, 182, 235, 307, 529, 878, 246, 723, 492, 259]. Romberg-Like [723]. rook [2496]. roommates [1858]. Root [580, 2016, 480, 2340, 934, 2324, 603, 552, 1759, 791, 426, 527, 1482, 1769, 867, 1296, 1297, 2222, 1565, 1412, 2078, 1130]. Root-Finding [2324, 552, 1759, 1412, 2078, 1130]. root-free [1296, 1297]. Rooted [2205]. rootfinding [2002]. roots [2375, 1560, 1882, 948, 1508, 2172, 862, 2542, 2137, 2166]. roots-finding [1882]. ROS3P [2460]. Rosenbrock [1160, 1732, 2460, 2711, 3638, 1061]. rotated [3581]. rotating [3706]. Rotation [2748, 1328]. Rotational [3147]. rotations [1391]. rough [2132, 3718, 3575]. Round [453, 442, 685, 565, 67, 758]. Round-Off [442, 67, 453, 685, 758]. Rounding [2445, 723, 689, 706, 773, 3109, 2900, 980]. Roundoff [2159, 1085, 386, 2588, 3236, 2447]. routine [2474, 2313]. Routines [2566]. row [2017, 2457, 2784, 1880, 2205]. ROW-type [2205]. rows [2164, 2337]. Ruhe [3194]. Rule [640, 413, 1770, 3700, 1095, 3664, 1029, 2947, 2482, 1526, 1700, 1489, 994, 2358, 1165, 1756, 1066, 882, 777, 3562]. Rules [350, 565, 2322, 296, 2566, 2570, 2740, 93, 1313, 1445, 1523, 2871, 1263, 2581, 1331, 2219, 2248, 3635, 757, 913, 1073, 3135, 1776, 3398, 1595, 977, 1065, 1150, 1513, 3415, 3272, 1737, 2502, 1444]. run [901]. Runge [2265, 2552, 2584, 3375, 890, 3417, 2043, 2615, 3355, 2291, 2315, 2292, 2988, 3160, 1945, 2187, 2215, 2391, 3719, 945, 897, 1100, 1575, 2246, 2616, 737, 785, 873, 1401, 1477, 2112, 1859, 1947, 2402, 2920, 2795, 1904, 3093, 2162, 3145, 455, 598, 1459, 1491, 2148, 2862, 1283, 2945, 3499, 1385, 1439, 2521, 2548, 2567, 353, 1712, 2406, 2970, 2323, 2550, 2444, 1440, 1039, 1087, 1699, 2056, 2076, 2847, 1987, 2046, 2057, 2084, 2204, 2778, 2838, 3229, 2223, 2334, 915, 2286, 2191, 1775, 2007, 3665, 958, 2762, 1733, 3284, 3095, 2071, 2198, 2252, 2312, 1915, 1754, 3037, 2647, 538]. Runge [1714, 3505, 1918, 2586, 3154, 2021, 1612, 1735, 2463, 1898, 1991, 2138, 2271, 809, 2386, 2803, 1864, 1906, 2326, 2735, 2819, 1978, 1388, 1718, 240, 1304, 1719, 1819, 1461, 1850, 2211, 1835, 3303, 2528, 2412, 2126, 371, 3054, 3595, 3603, 3540, 2472, 1988, 2433, 1106, 1914]. Runge-Kutta [2584, 2391]. runs [1370]. Runtime [1118]. rush [1965]. rush-hour [1965].

s [1286, 1785]. S/P [1785]. Saab [61]. Sabin [3281, 3168, 3074, 3202]. Sacker [2325, 3377]. saddle [3305, 2603, 3243, 2533, 3244, 3282, 3246, 3247, 3248, 3611, 3609, 3252]. saddle-point [3305, 2603, 3248]. Safe [895]. Sahalia [3670, 2983]. Sahalia-type [3670, 2983]. salesman [1447]. samplable [1909]. Sample [1765]. samples [3634]. samplesort [911, 691]. Sampling [3549, 3262, 2379, 3746]. sandwich [975]. Satisfaction [1765]. satisfying [3294]. saturated [2485]. Saxton [3501, 3400]. Scalar [3570, 2735, 3179, 2916, 1265, 1299, 3037, 3234, 3354, 3366]. Scale [2687, 2943, 1795, 2864, 3483, 3475, 3073]. scaled [3456, 3451, 2968, 3428, 2878]. scales [3646, 2825, 2650]. Scaling [2373, 940, 3361, 3440]. SCAN [1980]. Scanline [1659]. Scattered [2705, 2281, 3566, 2310, 2910]. scattering [2353, 2966, 3184, 3172, 3173, 3528, 3704, 3177, 3361, 2168, 3491]. Schatz [2707]. Scheduling [41, 52, 1242, 572, 929, 1965, 1980, 2024, 1013, 901, 2012, 1876]. Scheme [2565, 341, 388, 2709, 2748, 13, 549, 3604, 1463, 3654, 2038, 3297, 3214, 3047, 3186, 3450, 2837, 2593, 2875, 2191, 2026, 3165, 2205, 3166, 3348, 2816, 3423, 3650, 3475, 2978, 2802, 2226, 3432, 3424, 3219, 3235, 1852, 3400, 3533, 1130, 3365, 3378, 1152, 3222, 3524]. Schemes [2695, 719, 2686, 353, 379, 306, 571, 2703, 3043, 3673, 2794, 3294, 2942, 3295, 805, 967, 1280, 1369, 3371, 3719, 284, 3458, 3341, 3596, 3437, 3738, 3162, 2148, 3060, 3514, 3474, 3499, 3579, 2944, 3443, 1030, 3346, 3462, 3569, 993, 1063, 3619, 3283, 2084, 2634, 2107, 2885, 2897, 3033, 3334, 2812, 3598, 3284, 2898, 3089, 1413, 3592, 3594, 3267, 2586, 2924, 3169, 3270, 3269, 3188, 1333, 1217, 1388, 1718, 3189, 2211, 3289, 3156, 2125, 2155, 3101, 1444, 3275, 3425, 3622, 1424]. Schmidt [453, 2684, 222, 2242, 2130, 3129, 2588, 2156]. Schnabel [1234]. Schoenberg [3441]. Scholes [2973]. School [273, 118, 1679]. Schrödinger [3538, 3059, 2955, 3487, 3151, 2889]. Schur [524, 2922]. Schwarz [2044, 3527, 2607, 2535, 2720, 2689, 3268, 2940, 3157]. Science [187, 1684]. Sciences [32]. Scientific [2340, 3241]. Scientists [1496, 1655]. scrambled [3135]. screens [3528, 3704]. SDAEs [2890]. SDC [3221]. SDE [3294]. SDEs [3299, 3644, 3230, 3594, 3577, 3578, 3558, 3705, 3467, 2945, 2890]. SDFEM [2995, 3353]. SDIRK [1419, 1509]. SDIRK-methods [1419, 1509]. Search [68, 176, 555, 554, 645, 1434, 1639, 1361, 1328, 1806, 2013, 1960, 3476, 2622, 3190, 776, 1036, 796, 1522]. Searchability [1605]. Searching [1759, 1482, 1635, 1999, 1637]. Seasonal [239]. Secant [480, 728, 2559, 1811, 862, 994, 997, 1387]. Second [3146, 3738, 639, 509, 67, 567, 3743, 2585, 184, 278, 268, 2819, 3029, 3105, 2892, 928, 3690, 3159, 1280, 2306, 284, 3069, 3700, 1282, 1162, 1336, 3173, 1252, 1749, 912, 1487, 874, 990, 2522, 3730, 1771, 2070, 3615, 3620, 1016, 1713, 1775, 1989, 3136, 3096, 1475, 1032, 1102, 3113, 2676, 2167, 2980, 3155, 3415, 1885, 1254, 1301, 3303, 1462, 3320, 3524, 1106]. second-kind [3173, 3730]. Second-Order [2585, 3146, 3743, 3069, 3700, 3615, 3620, 3113, 2676, 3155, 3303, 3524]. secondary [180, 1175]. Section [176]. sections [3373]. segment [932]. segmental [3715]. segmented [838]. Segments [1556, 1241, 1839]. select [1704]. selected [760]. Selection [1997, 724, 3717, 1792, 771, 3605, 1378, 3571, 3322, 3494, 3311, 2587, 1250]. selective [3244]. Self [365, 454, 530, 990, 1088, 2888, 1620]. Self-Adjoint [530, 990, 2888]. self-reciprocal [1088]. Self-Starting [454]. Self-Steering [365]. self-timed [1620]. selfadjoint [2949, 3691, 2436]. Sell [3377]. semantics [1688, 1519, 1695, 1687]. Semi [3690, 3480, 2061, 2211, 2637, 3157, 2626, 3604, 3538, 2718, 2402, 3738, 2955, 1296, 1232, 3619, 1956, 3063, 3637, 2674, 861, 3348, 1817, 2423, 2486, 1866, 3091, 2984, 1182]. Semi-analytic [2061]. semi-classical [3538, 2955]. Semi-conjugate [2626]. semi-definite [2486]. semi-direct [1296]. Semi-discrete [3157, 3619, 2674, 3348, 2984]. semi-dynamic [1956]. Semi-explicit [3690, 3604, 861]. Semi-implicit [3480, 2211, 3738, 1232, 1817]. semi-infinite [1866, 1182]. semi-Lagrangian [2718, 3063]. semi-linear [2402, 2061]. semi-nonrelativistic [3637]. semi-norm [3091]. semi-separable [2423]. Semi-Toeplitz [2637]. semiconductor [1798]. semidefinite [3546]. semidirect [1297]. Semidiscrete [2671, 3475]. semigroups [2140, 2003, 2907]. semilinear [3196, 3093, 3613, 3283, 3486]. seminar [134]. Semismooth [2940, 3192]. Sensitivity [2662, 2294, 2536, 2336]. Sentence [543, 513]. Separable [687, 2423, 2820]. Separate [1190, 1851]. separation [3296, 1874]. separators [1998]. September [56, 2316]. Sequence [171, 796, 1766, 1270, 1338]. Sequences [433, 2305, 1656, 1828, 3296, 3590, 1691, 1578]. Sequencing [461]. sequential [2086, 997, 811, 814]. serial [2245]. Series [312, 204, 193, 2700, 384, 556, 648, 515, 3262, 3551, 1349, 3734, 2991, 2250, 2284, 1408, 2335, 2646, 2800, 1967, 2192, 1931, 2787, 3607, 2199, 3170, 831]. series-parallel [1967]. Set [2562, 481, 689, 215, 647, 1435, 1094, 1669, 1478, 1908, 2826, 1347, 3231, 1329, 1630, 1589, 3413, 1818, 1559, 1986]. Sets [318, 381, 1624, 804, 1952, 1860, 2266, 1999, 2027, 2840, 3024, 2001, 1889, 1957]. Setting [2548, 2908]. seven [3000]. seven-direction [3000]. Seventh [977, 1065, 1150]. Several [490, 2153, 1974, 1424]. Shallow [2549, 3371, 2346, 2721]. shallow-water [3371]. Shannon [3017, 3746]. Shape [2565, 3200, 2422, 3661, 3187, 3149, 2072, 2777, 2371, 1667, 3470]. shaped [3633]. shared [1851, 2374]. Sharing [671]. Sharp [1578, 2841, 3315]. Sharper [1303]. Sharpness [2467]. Sheets [109]. shells [2854]. Shellsort [599, 1072, 1728]. Shift [2089, 2095, 3249, 1279]. Shift-invert [2089]. shift-parameter [3249]. shift-reduce [1279]. Shifted [2561, 2574, 3562]. shiftedCholeskyQR3 [3742]. shifts [3080, 2063]. Ship [49]. Ship-Lines [49]. Shishkin [2978, 3353]. Shishkin-type [2978]. Shmoys [1876]. Shooting [609, 1417, 1464]. Short [1738, 1267, 3668, 797, 1234, 780]. shortest [1128, 1954]. should [798]. shrinkage [2587, 2489]. side [3513, 2933]. sided [3211, 3376, 3531, 1246, 1308]. sideways [2489]. Siemens [181]. sieve [838]. sign [2255]. sign-regular [2255]. Signature [1705, 3237]. Signatures [1660]. Significant [1429]. signs [923]. SIMD [1695, 1809]. Similarity [309]. Simple [440, 507, 1765, 3504, 3750, 3258, 197, 2435, 1159, 1961, 1094, 1441, 1508, 1292, 1542, 3506, 1598, 3239, 1005]. simpler [2474]. Simplex [248, 1574, 130, 3504, 2505]. Simplex-Technique [248]. simplices [3373]. simplicial [3044]. simplification [1063]. Simplified [705, 3243]. Simplifying [1120, 517, 207, 1962]. Simpson [777, 3618]. Simpson-type [3618]. Simula [440, 486, 1454, 287]. Simula-like [1454]. Simulation [192, 457, 233, 5, 2475, 2971, 2958, 3648, 3166, 2816, 3153, 3286, 1388, 2983, 2902, 3298]. simulations [2811, 3155]. Simultaneous [2518, 2645, 249, 3197, 2002, 930, 2497, 2542, 1124, 1579]. Simultaneously [720, 2166]. Sinc [2823, 3352, 1770, 2980, 3352]. Sinc-based [2823]. Sinc-collocation [3352, 2980]. sine [3534]. Single [1998, 1713, 3299, 1488, 919, 1974, 1877]. Singly [2647, 1945, 1100, 1161, 3412, 2021, 1461, 1061]. singly-implicit [1945, 1100, 1161, 1461, 1061]. Singular [684, 389, 618, 620, 1231, 2666, 642, 3629, 479, 93, 2794, 2965, 1354, 2237, 2470, 3341, 2281, 1540, 1642, 2861, 3697, 3420, 1629, 1592, 2045, 1487, 2388, 2497, 2863, 1264, 1611, 1184, 1058, 1284, 1713, 1414, 2886, 3702, 1149, 2421, 3247, 863, 3679, 3398, 3125, 2060, 2878, 3652, 2951, 3340, 1786, 2360, 2852, 1899, 3313, 3433, 1973, 2981, 1717, 1852, 2047, 903, 2527, 2529, 3193, 882, 1018, 2259, 1462, 3674, 1933, 2277, 2930, 3511, 3489, 3681]. singularities [1523, 2237, 2055, 2851]. Singularity [2362, 246, 3681, 978, 3523, 3489]. singularly [2391, 3226, 3017, 2970, 2851, 3677, 2801, 2978, 2979, 3302, 1820, 2918, 2638]. site [1877]. situ [1825]. six [1449]. sixth [839, 957, 1720]. Sixty [3190]. Size [2733, 660, 647, 3031, 1122, 1075, 3517, 1210, 2010, 1886, 1964]. skatteuppbörd [10]. Skew [2639, 2603, 1928, 3338, 2465, 3403, 1928]. skew-circulant [1928]. Skew-Hermitian [2639, 2603, 3403, 1928]. skew-symmetric [2465]. skinny [3479]. skip [1960]. SLD [2013]. SLD-tree [2013]. slip [3693]. slow [3474]. slow-fast [3474]. slowly [2192]. Small [2733, 3627, 661, 3416, 3280, 3226, 842, 3135, 1286, 760, 2115, 2890, 2001]. smallest [930, 1955]. Smith [241]. Smooth [49, 2661, 211, 311, 3729, 1927, 3496, 2431, 2512, 3151, 2789, 2890, 3391, 3618]. Smooth-Curve [211]. Smoothed [2549]. Smoothers [2670]. Smoothing [66, 413, 403, 2589, 3416, 2420, 3471, 2817, 2037]. Snapshots [213]. SOFT [197]. software [904, 3121, 2309, 1682]. Soil [239]. solitary [3153]. solitary-wave [3153]. soluble [1715]. Solution [453, 700, 71, 294, 3370, 3369, 2052, 470, 597, 702, 352, 418, 454, 443, 472, 227, 279, 313, 404, 462, 297, 656, 399, 743, 620, 536, 299, 567, 341, 388, 657, 1080, 2552, 2569, 689, 604, 268, 185, 390, 607, 1233, 2183, 2139, 448, 249, 289, 270, 323, 164, 628, 45, 149, 168, 1384, 2032, 2794, 1869, 967, 2880, 3525, 2343, 3597, 1262, 3588, 2393, 3058, 2402, 1054, 1230, 3145, 2976, 2862, 1749, 3163, 3307, 3367, 3226, 1385, 2473, 3211, 2520, 2874, 546, 1844, 1176, 931, 1101, 1310, 886, 990, 2296, 2994, 2723, 841, 3177, 1662, 3254, 1930, 1507]. solution [2799, 844, 1058, 1284, 1488, 84, 861, 1081, 1149, 3483, 3740, 1266, 1416, 1352, 1213, 2500, 3231, 3073, 2273, 2840, 3024, 3413, 3302, 1818, 1034, 2195, 3339, 1613, 847, 1820, 812, 1082, 1035, 2449, 1932, 3509, 2276, 1302, 2426, 1848]. Solutions [260, 255, 265, 2696, 2661, 639, 2687, 2734, 414, 2739, 2702, 2965, 988, 1892, 2237, 1459, 1283, 1403, 1439, 1487, 2523, 3663, 1184, 3501, 3395, 2407, 1164, 2084, 2850, 3150, 1647, 1475, 3388, 3285, 3153, 1215, 1000, 3607, 1455, 3457, 774, 2962, 3290, 1212, 2505, 2820, 3562]. solvability [1714]. solve [2333, 3330, 2226, 1472, 1381]. Solved [183, 138]. solver [3381, 3184, 3446, 3070, 2460, 3009, 2938, 3023, 2973, 3491]. Solvers [2663, 2669, 3599, 1972, 2532, 2477, 2935, 2420, 2384, 1543, 3655, 3731, 1393]. solves [2783, 1846]. Solving [222, 2580, 1972, 693, 3754, 2321, 203, 2510, 2233, 1913, 329, 342, 354, 319, 687, 491, 2555, 1367, 1124, 3022, 2700, 2787, 1701, 1197, 3601, 1868, 1798, 3043, 2600, 3392, 1921, 3207, 1697, 897, 2307, 2836, 2177, 3004, 3659, 3730, 3542, 3017, 902, 3178, 2511, 2875, 2956, 2760, 3397, 1442, 1790, 1265, 1299, 1700, 3247, 863, 3126, 1417, 1386, 3412, 2951, 1033, 2999, 1151, 2489, 1493, 3189, 3585, 1648, 3227, 1249, 3220, 3521, 1250, 2831, 3142, 2930, 2625, 1424]. Some [351, 1961, 1343, 2754, 575, 637, 2496, 2631, 2148, 472, 1056, 1079, 1057, 193, 2969, 380, 658, 877, 2747, 528, 562, 641, 2106, 319, 673, 51, 108, 3746, 2078, 101, 643, 368, 3109, 3546, 2060, 1931, 1108, 1864, 556, 1715, 634, 881, 1321, 1337, 648, 1060, 323, 516, 1193, 1037, 1159, 610, 371, 1686, 2032, 1888, 1400, 1349, 284, 1119, 1980, 1405, 902, 2233, 956, 1744, 2459, 1975, 2677, 1612, 2979, 1579, 2272, 606, 829, 1002, 1973, 2170, 1993, 1295]. SOR [2419, 2131, 3250]. SOR-like [2419, 3250]. Sort [691, 2025, 1137]. Sorting [3751, 234, 1470, 1607, 1747, 1226, 1656, 1704, 1157, 1429, 1821, 1825, 1294, 1486, 1707, 1365]. Source [2848, 951, 3018, 2674, 3321]. Space [3754, 2321, 1074, 1639, 3321, 3368, 2600, 3161, 1401, 3528, 3209, 886, 1276, 1983, 887, 3376, 3531, 2500, 1902, 1857, 3064, 1141, 3647, 3098, 3091, 3378, 3014, 3524, 3536]. Space-dependent [3321]. space-time [2500, 3098, 3536]. spaced [3634]. Spaces [2744, 2661, 366, 3224, 2365, 2954, 3459, 2872, 3209, 2204, 3049, 3036, 3285, 3114, 3128, 792, 3549, 2504]. spacings [2087]. spanning [2600, 1501, 1955, 1030, 1345, 1967, 1276, 1141, 1709, 1504]. Spare [85]. Sparse [2792, 2849, 3110, 646, 708, 249, 662, 1476, 1697, 2065, 2238, 930, 2509, 817, 1296, 1297, 3123, 2457, 2784, 2233, 3147, 3112, 2086, 1348, 2461, 972, 1701, 3271, 3507, 2139, 2374, 2464, 1197]. Sparsity [3160, 1783, 1787]. spatial [3392, 3382, 3646, 2957, 3689]. SPDE [3432]. SPDEs [3315]. Special [701, 75, 3007, 897, 1137, 1541, 2459, 1527, 2348, 1423, 1152]. species [2724]. specific [3072]. Specification [1364]. specifications [1379, 1944]. Specified [2570, 2826, 2316, 3498, 2129]. Spectra [1879]. Spectral [3207, 2381, 3411, 2328, 2339, 3587, 3381, 3714, 3588, 3161, 1792, 2266, 3386, 2675, 2981, 2153, 3401, 3204, 3377, 3379, 2068]. spectrally [3177]. spectroscopic [1028]. Spectrum [3246]. Speed [83, 811]. Speeding [1137]. speedup [1503]. sphere [3726, 2910, 1065]. spheres [3686, 3336]. spherical [1713]. splaying [1606]. Spline [49, 3195, 639, 656, 211, 600, 238, 676, 2037, 403, 2530, 3224, 564, 2044, 2859, 2989, 3277, 2470, 3741, 2054, 2247, 2532, 2934, 2872, 3071, 3108, 3163, 1845, 546, 2221, 1610, 1752, 3319, 1663, 2909, 3168, 2937, 790, 2336, 3064, 2911, 2254, 937, 1033, 2207, 3000, 2829, 2960, 3074, 1321, 1337, 3202, 1668, 1236, 1627]. spline-collocation [1610, 1752]. Spline-Fit [211]. Splines [2633, 298, 474, 641, 722, 2424, 2556, 2737, 3002, 968, 1609, 3281, 2407, 1525, 3478, 2087, 3471, 2813, 1108, 3329, 1247, 1564, 2539, 3600, 2117, 1973, 1549, 1598, 1717, 1667, 1736, 1867, 3026, 1546, 1900, 2036, 2209, 2123, 3091, 3100, 3075, 3191, 1820, 1082, 2213]. Split [224, 1890, 3155, 3026]. split-step [3155]. Splitting [2905, 3101, 3304, 3295, 3392, 2603, 3669, 3553, 3458, 3596, 3145, 2955, 3244, 2907, 3148, 3283, 2346, 3347, 2958, 2674, 887, 2357, 3124, 2035, 2889, 3287, 2984, 2639, 3078, 3403]. Splittings [2663, 2583, 2409, 3568, 3049]. Springer [3484]. Square [2016, 2745, 2389, 2583, 623, 3105, 3294, 3404, 1560, 900, 3135, 2172, 3333, 1648, 847, 907, 3315, 3467]. Squares [453, 260, 255, 222, 265, 3754, 3755, 2321, 3356, 2666, 642, 687, 2571, 747, 2373, 1017, 2327, 2669, 2739, 2329, 589, 2741, 2559, 3685, 2880, 1476, 2754, 1892, 1011, 2579, 3626, 2881, 1972, 2331, 1402, 3635, 3300, 3659, 840, 1264, 1404, 2510, 2332, 1547, 2098, 2180, 2069, 2130, 2407, 2233, 2150, 2299, 2251, 3201, 2287, 1332, 1367, 2461, 3126, 3336, 2678, 2804, 2828, 2274, 829, 3476, 1544, 2123, 3601, 1028, 2258, 2212, 2197, 1974, 3469, 3550, 1889]. SSOR [523, 1753, 1793, 3247]. SSTF [1980]. Stabbing [1241]. stabilisation [3325]. Stabilised [3693, 3490, 3396]. Stabilitätsdefinition [60]. Stability [1313, 1445, 2032, 2279, 1476, 701, 2319, 2695, 2953, 2546, 2745, 2662, 737, 3754, 2321, 719, 75, 1143, 3427, 639, 3199, 640, 1913, 2389, 2538, 2568, 2133, 2334, 2347, 3164, 916, 1081, 3650, 2710, 2620, 2451, 3327, 2668, 659, 333, 3188, 1333, 3432, 1963, 3400, 1719, 1650, 1850, 628, 371, 3591, 2625, 2426, 2068, 1914, 1848, 2375, 3417, 3439, 2942, 1696, 2145, 2102, 896, 3719, 1098, 1078, 1099, 1267, 1622, 1590, 1591, 2755, 2103, 1282, 1195, 1355, 3256, 1459, 2148, 1749, 2003, 1402, 912, 946, 1446, 1056, 2120, 3134, 947, 1610, 3461, 3617, 3515, 1351, 3395, 1408]. stability [1039, 1232, 1409, 1390, 2512, 1987, 2046, 2057, 991, 2838, 3323, 2513, 1773, 915, 808, 860, 761, 958, 1733, 2071, 2198, 2252, 2312, 60, 3096, 2088, 3333, 772, 2959, 3573, 3724, 1315, 1897, 3675, 3520, 1624, 2924, 1612, 1864, 1906, 2255, 2302, 1217, 3415, 3065, 2170, 979, 2154, 1166, 2526, 3391, 1421, 2228, 3377, 1019, 1319, 1181, 780, 954, 959, 2768, 1021, 3027, 2806, 2049, 2128, 1393, 3634, 2249, 1106]. Stabilization [3708, 2690, 2093, 3648, 3453]. Stabilized [2144, 3657, 3614, 3658, 1797, 2949, 2959, 3710]. Stable [3043, 469, 582, 524, 441, 455, 3060, 566, 2403, 297, 11, 670, 379, 2132, 3049, 2594, 1983, 3398, 1215, 381, 1825, 2558, 497, 226, 534, 3105, 328, 655, 2187, 1262, 945, 1575, 873, 2651, 1398, 1371, 1282, 1162, 1336, 3738, 779, 1252, 2944, 1752, 2018, 1087, 2884, 2346, 1293, 3612, 1016, 2107, 3359, 2071, 2198, 3544, 1475, 1855, 3423, 2423, 3412, 1898, 1991, 762, 2927, 1472, 3302, 1449, 1580, 1858, 1368, 1626, 1381, 1768, 1729, 1105, 3468, 3524]. stack [1075]. stackless [1512, 1573]. Stage [2647, 2601, 2215, 3719, 1477, 3059, 1524, 2191, 3154, 2528]. stages [2291, 2920, 1904, 2190]. stagnation [2225]. Standard [2446, 2282, 2995, 2908, 3065]. stars [1410, 954]. Starting [454, 2406, 2551, 268]. State [1639, 1647, 3509]. State-Space [1639]. stated [2887]. Statement [510, 951]. Statements [721, 301, 922, 95]. static [2453]. statically [2012]. stationary [2934, 2283, 2761, 2959, 1213, 2927, 3666]. Statistical [274, 706, 429, 386, 574]. Staudt [631]. steady [1647]. Steepest [2611, 3144]. Steering [365]. Steffensen [1180]. Steffensen-like-methods [1180]. Steiner [1474]. Steklov [3691]. Step [581, 524, 2745, 719, 657, 624, 497, 3684, 1945, 2859, 3602, 2531, 1162, 1252, 1749, 3107, 2521, 1644, 2894, 3694, 3748, 1351, 1122, 1144, 1713, 3744, 3650, 3155, 1885, 3548, 1460, 2681, 3099, 2010, 3533, 3377, 1886, 1964, 1597, 2779, 2412, 1201, 2902, 1974, 1053]. step-by-step [3748]. step-parallel [2521]. steplength [830]. stepping [2203, 3644, 3270, 3269, 2764, 2788]. steps [2306, 1896, 25, 2589]. StepSCAN [1980]. Stepsize [2710, 2293, 1662, 1760, 3625, 2638]. Stiefel [1664, 2553, 2472]. Stieltjes [3264, 3094, 3494, 3351]. Stiff [582, 487, 686, 674, 624, 3105, 1926, 2186, 896, 1314, 1027, 1160, 897, 2429, 2402, 857, 841, 1699, 1732, 1286, 1883, 2483, 3166, 1413, 3412, 953, 973, 1316, 2788, 1719, 2290, 2194, 1992, 1835, 3053, 1393, 3630, 1053]. stiff-stability [896]. Stiffly [582, 1282, 2187, 2970]. Stiffness [3735, 2019, 3190]. Stochastic [2546, 2745, 2380, 2389, 3387, 3696, 3095, 2735, 2557, 2736, 2671, 3105, 3043, 3355, 3418, 3068, 3477, 3419, 3382, 3458, 2393, 2246, 2616, 3654, 3473, 3046, 2862, 3474, 3499, 3579, 2944, 3480, 3567, 3699, 2723, 2825, 2896, 2477, 3395, 2837, 2775, 2759, 3557, 2958, 3033, 3516, 2205, 3166, 3020, 3089, 3021, 3037, 3333, 3572, 3350, 2073, 3155, 2819, 3189, 3274, 3303, 2528, 2790, 3053, 3275, 3467, 3595, 3291, 3651, 3298]. Stockholm [57]. Stokes [3496, 3706, 2608, 3709, 3693, 3657, 2959, 3097, 3022, 1850, 2211, 2637, 2855, 3737]. Stopping [2871, 2417, 1004]. Storage [486, 223, 125, 13, 40, 1190, 691, 429, 335, 3565, 899, 1030, 1705, 807, 1175, 811, 814]. store [1802]. Storing [1839, 1041]. Störmer [2010]. Störmer/Leapfrog/Verlet [2010]. Strakos [3050]. Strang [2555]. Strang-Type [2555]. Strategies [726, 2012, 3259, 2099]. Strategy [1765, 2451, 2255, 2788, 2374, 776]. stratified [2502]. Stratonovich [2819]. Straus [432]. stress [2237, 3136]. Strict [259, 431]. strictly [3560, 3090]. String [457, 336, 1138]. Strings [1484, 77, 1347, 1874]. Strong [3654, 2582, 2512, 3516, 3644, 3043, 3477, 2246, 3514, 3033, 3577, 3578, 2639]. Strongly [469, 3490, 1813, 3032, 2983, 2624]. Structural [3311, 2435, 3260, 2042, 2245, 3328]. Structural-algebraic [3260]. Structure [3324, 370, 335, 1888, 2942, 3553, 3082, 2082, 3641, 3122, 1999, 843, 245, 1096, 2150, 2000, 1985, 2183, 2982, 3595, 3666]. Structure-preserving [3324, 3595, 3666]. Structured [654, 3116, 3421, 2369, 473, 3517, 622, 3680, 461, 2741, 2559, 3714, 3119, 3120, 3420, 2239, 3338, 1259, 3127, 3608, 2912, 2361, 1383, 3563]. Structures [1538, 103, 2091, 2519, 3253, 1605, 1787, 3680, 1638, 3434]. Structuring [398, 1051]. Student [129]. students [1453]. studied [1699, 1732]. Studies [204]. Study [3474, 5, 300, 417, 1688, 2065, 3530, 2596, 2224, 795, 765]. studying [2430]. Sturm [1399, 2578, 2990]. style [95]. Styrekredse [53]. subclass [1245]. subclasses [1942]. subdiffusion [3478, 3220, 3523, 3562]. subdivision [3673, 3002, 2934, 3162, 3084, 3200, 2777]. subdomain [2473]. subdomains [3492]. subgraphs [1952]. subharmonicity [2266]. Subject [275, 842, 1332, 1367, 3399, 2170]. sublogarithmic [1838]. subproblems [2877]. Subroutine [709]. Subroutinized [234]. subscribers [29]. Subscripts [729, 662]. subsequence [1347]. subset [1944]. Subspace [2609, 2159, 3668, 2388, 3264, 3265, 3048, 3308, 2269, 3519, 2776, 3682, 3747]. Subspaces [2697, 385, 2497, 3123, 2826, 3682, 2182, 3076]. substitution [2128]. Substructuring [3701]. subtrees [1902]. Successive [600, 59, 2176, 1771, 2287, 1254, 1301, 2448]. successor [2900]. Sufficiency [952]. Suggestion [152]. suitability [1754]. suitable [1705]. Sum [3, 806, 3300, 815, 2199]. sum-of-squares [3300]. summability [2075]. summands [1734]. Summation [365, 442, 685, 2322, 193, 2199, 433, 2219, 2248, 3331, 3267, 2192, 3025, 3156]. summation-by-parts [3267, 3156]. Summer [273]. Summing [1734]. Sums [1879, 456, 511, 746, 643, 198, 3641, 1698, 1948, 2193, 793, 3102]. Super [3703, 3281, 3430, 3455]. super-linear [3430, 3455]. Super-localized [3703]. Supercloseness [3167]. supercomputers [2233]. Superconvergence [2743, 3663, 2814, 3097, 2918, 3679]. Superconvergent [3293, 3202, 3052, 2859, 3452]. superoptimal [2534]. superscalar [2024]. Superstable [1775]. support [3446, 3018]. supported [3002, 2114, 2074]. supports [2848]. surface [3083, 2432, 2407, 1018]. Surfaces [2694, 3184, 3307, 3367, 2817]. survey [1425, 1436, 1791, 2602]. SURVO [274]. SVD [3525, 3057, 2992, 3335, 1623, 2267, 3340, 3011]. Sverige [10]. swap [3681]. swapping [3489]. Sweden [10, 180]. Swedish [1, 74, 10, 152, 92, 140, 26, 17, 34]. sweep [1471]. sweeps [3221]. Switching [1543]. symbol [38]. symbolbruken [38]. Symbolic [301, 786]. Symbols [3329, 2835]. Symmetric [438, 3669, 441, 3278, 2581, 2537, 2408, 2758, 1646, 359, 3128, 369, 390, 478, 708, 2842, 2703, 2173, 2990, 2920, 3069, 3460, 930, 3122, 3384, 2097, 2523, 2476, 3086, 2316, 3136, 1854, 2398, 3338, 1107, 3268, 3546, 972, 3329, 121, 135, 3362, 3559, 2465, 1449, 1197, 1737, 1778, 3521, 2806, 2129, 2625, 3563]. Symmetric-conjugate [3669]. Symmetric-triangular [2537, 2842]. symmetrical [759]. symmetries [3640]. symmetrizable [2264]. symmetry [3279, 1713, 2853, 3680]. Symplectic [2664, 2550, 2076, 2620, 3481, 2399, 2315, 2292, 2082, 3208, 3306, 2920, 2394, 2056, 3557, 2595, 2888, 3127, 3350, 2229, 1962, 1978, 3129, 3454, 3466, 3275, 3521, 3054, 3603]. symplecticity [3107]. symplecticness [2116]. Symposium [273]. Synchronous [163]. Synnoetics [32]. Syntax [295, 264, 350, 653, 881]. Syntax-Directed [295, 350]. System [486, 669, 296, 66, 23, 256, 145, 154, 1220, 119, 267, 390, 61, 173, 249, 661, 110, 1798, 928, 3604, 3344, 3488, 3526, 786, 1171, 1939, 2874, 1789, 1453, 2756, 3637, 3518, 2815, 3475, 2917, 938, 3273, 1818, 2194, 1648, 3192, 1076, 2787]. systematic [3564, 2622]. Systems [700, 3418, 1799, 2319, 2561, 1554, 2564, 693, 297, 686, 275, 2697, 2419, 2956, 528, 562, 5, 2698, 101, 587, 2555, 1660, 624, 2574, 173, 2472, 3685, 1926, 2093, 2823, 2601, 2291, 2416, 1782, 1436, 3305, 2343, 2439, 988, 1314, 1697, 2095, 2052, 3588, 1027, 2111, 1160, 2147, 2402, 2795, 857, 1928, 2921, 2967, 3198, 2976, 2430, 940, 3668, 3211, 3031, 1571, 2863, 3394, 1797, 931, 817, 2534, 1789, 2297, 841, 1426, 2369, 1940, 2536, 3462, 3253, 1846, 859, 3407, 1941, 1164, 1987, 2046, 2084, 991, 3556, 1080, 807, 2133, 3323, 3150, 2760, 3612, 2958, 1051, 2007, 2085, 2299]. systems [2876, 3034, 1145, 1793, 3166, 3375, 2071, 3388, 1692, 2384, 3592, 3581, 1266, 1714, 3545, 3675, 1824, 1417, 3327, 3249, 1386, 3154, 3010, 3330, 3128, 2938, 1443, 1353, 2487, 2515, 3350, 3250, 3362, 1786, 2998, 3559, 3352, 1316, 2255, 2352, 1701, 2273, 2853, 1910, 1233, 1682, 2974, 1761, 1718, 2139, 2227, 3260, 2099, 3339, 1992, 2527, 1835, 2425, 3328, 3227, 1457, 2449, 2856, 1685, 2639, 3521, 3676, 2780, 2842, 2128, 2505, 2820, 3403, 2626, 149, 2930, 3115, 2625, 3563, 3603, 3540]. Systolic [1539, 1558, 1585, 1745, 1727, 1766]. Szilard [1363].

Table [1969, 576, 806, 2164, 760]. Tables [118, 283, 236, 2122]. Tabulation [35]. Tal [3626]. Talbot [2767]. tame [3616]. tamed [3699]. Tandem [269]. tangram [1389]. Tank [192]. Tape [234, 83, 40]. Task [1639]. tau [3232]. tau-leap [3232]. Taught [1679]. tax [10]. taxonomy [1535]. Taylor [2991, 165, 2885, 2699, 773, 2700, 2787, 3364]. Teaching [152]. Technique [382, 248, 2630, 770, 3177, 3395, 3611, 3609, 1210, 1356]. Techniques [368, 387, 2880, 82, 2810, 2104, 3213, 1012, 3211, 1816, 2608, 2113, 2296, 2673, 3228, 2420, 3388, 3652, 2765, 2939, 3181]. Temperature [239]. tempered [3535]. temporal [3646]. Temporary [690]. ten [1477]. ten-stage [1477]. tension [2872]. Tensor [2873, 3583, 3641, 3659, 2407, 3110, 3137, 3636, 3545, 3349, 3288]. tensor-based [3288]. tensors [3460, 3300, 3543]. term [3082, 3198, 1577, 3358, 2677, 3506, 2194, 3012, 3521, 2426]. terminal [1967]. Terminating [1693]. termination [1665, 1521]. Terminology [126]. Terms [307, 1478, 1940, 2674, 3571]. Test [65, 2390, 947, 2883, 1127, 1520, 1185, 2962]. tested [1860]. Testing [30, 31, 490, 2121, 2224, 1660, 936, 543, 513, 1999, 1173, 1380]. tests [1343, 896, 1976, 2092]. tetrahedral [2961]. tetrahedron [2072]. Text [77, 2571, 43, 1705]. texture [3585]. texture-rich [3585]. th [3139, 2029, 707]. Their [2403, 2686, 2667, 796, 2231, 1584, 1676, 757, 2731, 2191, 212, 3725, 3713, 394, 355, 1856, 3169, 1612, 646, 880, 3440, 247, 3239, 3191, 3275, 371]. Theo [1939]. Theorem [631, 199, 3734, 1234, 1895, 1672, 2059, 3498, 1427, 1919, 1415, 216]. theorems [1056, 2388, 3568, 1527, 954, 2537]. Theoretical [2097, 300, 3352, 417, 2986, 1405]. theories [2418]. Theory [378, 1950, 101, 573, 2353, 2186, 2145, 2579, 3057, 2914, 1938, 3300, 2497, 3582, 878, 1196, 2006, 3127, 2138, 1379, 1943, 1596, 1651, 1944, 1683, 2766, 3377, 2504, 2768, 2127, 2298, 2412]. thermo [2475]. thermo-regulation [2475]. Thermostatted [2642]. thermoviscoelastic [3326]. Theta [2546, 3467]. thin [2854]. Thinking [1520]. Thinning [2310]. third [2945, 2333, 861, 3375, 2460]. third-order [2333, 2460]. Thirty [2755]. Threaded [1553]. Three [1389, 367, 2864, 224, 341, 997, 2668, 429, 967, 3160, 2102, 3176, 3018, 2957, 3358, 3360, 2882, 1419, 2371, 3506, 1907, 3489]. Three-Dimensional [367, 967, 3160, 2371]. three-field [3360]. Three-Level [341]. Three-scale [2864]. three-term [3358, 3506]. Three-triangle-tangram [1389]. threshold [3495, 3623]. tight [1925]. tighter [2464]. Tightly [1190, 2182]. Tikhonov [2563, 3335, 3605, 2396, 2572, 2972, 3245, 2485, 3140, 2868, 3218, 3433]. Time [2693, 118, 2531, 3460, 3172, 2686, 2993, 3357, 256, 671, 2203, 2136, 1685, 1826, 3368, 1225, 2615, 2782, 2094, 1656, 2160, 3512, 2604, 1965, 3586, 3624, 1050, 3495, 3623, 1669, 3213, 2976, 2430, 3480, 3668, 3658, 2894, 1013, 901, 1930, 1823, 3374, 3164, 3627, 3516, 1983, 3644, 2814, 3486, 3573, 3744, 3650, 2676, 2500, 1902, 3675, 1362, 2061, 2802, 3270, 3269, 1470, 3362, 3188, 1984, 3399, 2764, 3098, 2788, 2681, 3170, 3533, 3453, 811, 2790, 3220, 3523, 2589, 3321, 2031, 3115, 3320, 3536]. time-changed [3480, 3516]. time-delay [3675, 3115]. Time-Dependent [2693, 2615, 2604, 3188, 3453]. time-fractional [3374, 3399, 3321, 3536]. time-harmonic [3658]. Time-marching [2136, 2061]. time-optimal [1823]. time-periodic [1930]. Time-Sharing [671]. time-space [3368]. time-step [3533]. Time-stepping [2203, 3644]. time-varying [2676]. timed [1620]. times [2905, 3059, 3559, 2226]. timestepping [2645, 2384]. timesteps [2503]. Toeplitz [2604, 3631, 3183, 1928, 2508, 3031, 3384, 3556, 2733, 2085, 2150, 2299, 3517, 2815, 2288, 3411, 2487, 2912, 2998, 3559, 2349, 2099, 2153, 2328, 2425, 2637]. tolerant [1877]. tomographic [3288]. tool [1315]. tools [2969]. Top [330, 346, 134, 1606, 2464]. Top-Down [346, 330, 1606, 2464]. Topological [516]. Tore [1002]. Torvik [2520, 3711]. Total [3356, 2669, 2796, 1981, 3422, 3032, 2299, 3429, 2361, 764]. total-variation [2796]. totally [3086]. Towers [1327, 1310, 1455, 3457]. trace [3519]. tracing [2175]. tracking [2936, 3234, 3354]. tradition [1684]. transactions [1910]. transcendental [2137]. transcription [2971]. Transfer [1118, 3183]. Transform [183, 138, 344, 2455, 1478, 1593, 2443, 2253, 2075, 1083, 2713, 2532, 2083]. Transformation [260, 411, 2550, 2734, 309, 2053, 2610, 1757, 795]. Transformations [169, 146, 3398, 3532, 1455, 3457]. Transforms [2731, 2792, 2904, 2066, 2004, 2946, 2298, 2311, 2058, 2089, 794]. Transient [196, 2985]. Transition [290, 3506]. transitions [1923]. Transitive [357]. Translating [301]. Translation [350, 2748, 625, 310, 2971, 2168]. Translations [1245]. Translator [153, 197, 1431]. Translators [382]. transmission [2240]. transparent [3175]. transport [3514, 3608]. transposed [1304]. transposition [1802]. transversal [3253]. Transverse [457]. Trapezoid [688]. Trapezoidal [413, 1489, 1066, 3562]. Traversal [1158]. traversals [1535, 1584, 1779]. traversing [1346]. treatment [3538, 1809, 2727]. Tree [617, 691, 1535, 2954, 1501, 1804, 1669, 1517, 2784, 1276, 2205, 2000, 1588, 1779, 1902, 2013, 1758, 1709, 1295, 1383]. tree-structured [1383]. Trees [1553, 2643, 1158, 2600, 1953, 2094, 1959, 1584, 824, 1938, 1955, 1030, 1345, 1346, 1328, 1391, 1823, 1967, 1873, 1075, 949, 1890, 1141, 1097, 1191, 1839, 1630, 1589, 1911, 1175, 1808, 1810, 1875, 1036, 796, 1504]. Trefethen [1919]. triangle [1389, 3513, 2933, 1670]. Triangular [684, 688, 3197, 1571, 2537, 2223, 2865, 2961, 2210, 2639, 2842, 3353]. Triangularization [169]. triangulation [1853, 3373, 1958, 1362]. triangulations [1377, 3071, 3281, 2015]. Tridiagonal [438, 2662, 441, 2160, 858, 839, 3150, 2058, 2815, 2423, 2128, 2400, 2625]. tridiagonals [2182]. Tries [1192]. Trigonometric [444, 412, 3238, 1971, 3488, 2005, 2345, 2846, 3062, 2135, 1527, 996, 1103, 2213, 3622]. trigonometrically [2920]. triple [3247, 3128]. triple-parameter [3247]. triplets [1933]. Triply [1158]. trivariate [3108, 3386, 3000, 3026]. trivial [1764, 3752]. true [1667]. Truncated [3736, 2523, 3670, 1623, 1772, 923, 1788, 3674]. Truncation [3749, 642, 3262, 1862]. Trust [2881, 3740, 2877]. Trust-region [2881, 3740, 2877]. tubal [3636]. Tucker [3131, 3460, 3543, 952]. Tuning [814]. tunnel [1]. Tunneldioder [1]. Turán [2677]. Turing [1520]. turning [2979, 1251]. Tutorial [507, 1677]. TV [3605]. twice [2989]. Twin [3]. Twisted [2901, 1656].

Two

[524, 2745, 1162, 2861, 223, 1252, 2083, 2746, 275, 169, 3542, 2583, 1144, 1790, 1412, 621, 906, 2059, 2234, 2243, 1246, 1308, 269, 414, 1747, 3099, 1035, 2528, 2412, 3684, 800, 2601, 755, 3404, 2966, 1945, 2280, 3719, 2724, 2002, 1054, 1230, 1603, 858, 839, 969, 1125, 1320, 1540, 1642, 3296, 1576, 3226, 2894, 1057, 1524, 3450, 1440, 759, 3149, 2827, 1058, 1663, 1967, 3725, 3713, 3216, 887, 3581, 3475, 1824, 1563, 935, 2193, 2998, 3258, 3233, 1472, 3548, 3432, 1717, 3286, 2681, 3289, 3535, 1925, 2918, 1082, 2779, 921, 3731, 3491, 3489, 3681]. two-by-two [3404, 3581]. two-chemical [2724]. Two-Dimensional [2746, 169, 1576, 759, 3258, 3432, 3286, 3535]. two-disk [1603]. two-headed [1824]. Two-Level [223, 1790, 2998, 3731]. Two-parameter [1035, 3233, 2918]. Two-Point [275, 414, 800, 755, 1054, 1230, 858, 839, 969, 1125, 1320, 1540, 1642, 1057, 3450, 1440, 1663, 3216, 1717, 3289, 1082]. two-scale [3475]. Two-sided [1246, 1308]. Two-stage [2528, 2601, 3719, 1524]. Two-Step [524, 2745, 1162, 1252, 1144, 3099, 2412, 3684, 1945, 2894, 3548, 2681, 2779]. two-terminal [1967]. Type [37, 580, 48, 2684, 2319, 424, 470, 583, 618, 601, 482, 2555, 624, 2472, 1798, 3016, 2718, 3133, 1529, 2393, 1263, 3279, 955, 1928, 1336, 3162, 1938, 1576, 2632, 3670, 2296, 2369, 1940, 1610, 1752, 1771, 1184, 2774, 3395, 993, 3005, 2956, 3006, 1080, 3323, 1058, 1441, 2135, 1178, 917, 2205, 3543, 3745, 3739, 2385, 2950, 3584, 2978, 3558, 1379, 1943, 1580, 1944, 2501, 2185, 3012, 3203, 865, 2983, 709, 2727, 882, 2504, 3618, 1152, 3575, 3222, 3540, 1106]. type-systems [1940]. Types [1171, 161, 3688, 1944].

Ubiquitous [3363]. ufp [3639]. ulp [3639]. ultimately [1691]. ULV [2145, 2267, 2314]. Unbiased [3433]. unbounded [3570, 2889]. uncertain [3366]. uncertainties [3330]. Uncertainty [1352, 358, 3495, 2540, 3561]. Uncle [1842]. Uncle-attributed [1842]. unconditional [912]. Unconditionally [2651, 1336, 2346, 3375, 226]. unconventional [3539]. under-determined [3550]. Underestimating [2691]. underlying [3107, 3169]. Undirected [488, 1517, 1985]. Unequal [381]. unevenly [3634]. unfeasible [1924]. unfitted [3103]. unification [1452]. Unifications [1856]. Unified [493, 3683, 1196, 2504]. Uniform [856, 398, 659, 3551, 3002, 2039, 1629, 1592, 3071, 1609, 1845, 1873, 1108, 2802, 2979, 3302, 2117, 1544, 3676, 3562]. Uniformly [3526, 3596, 1698, 379, 3320, 3103]. unifying [2977]. union [1908]. unions [3330, 3314]. uniparametric [2970]. Unique [411, 320, 645]. Uniqueness [736, 1439, 1906, 2091]. Unit [565, 40, 1804, 94, 3125, 3236, 3639]. Unitary [169, 3669, 2206, 3127]. united [3413]. unity [3287]. Univariate [2892, 2829, 2257]. Universal [296]. Universities [152, 1428]. University [32, 1453]. unknown [1774]. Unlimited [215]. Unrestricted [1560]. Unstable [275, 1215]. unsteady [3267]. unstructured [2239, 3167]. update [3270, 3269, 1960, 2275]. updated [1254]. updates [1574]. Updating [389, 1558, 2015]. upon [3002, 1478, 2136, 1365]. Upper [2570, 3477, 1268, 3442, 499, 1503, 2257, 1925]. upscaling [3438]. upwind [2794, 2978]. Urysohn [3293]. Usage [1220, 1966]. Use [742, 558, 444, 369, 358, 433, 1369, 38, 932, 2191, 791, 1757, 1082]. Used [49, 65, 532, 3064]. User [186]. users [2026]. Using [2232, 2880, 470, 2174, 2471, 378, 223, 380, 1847, 329, 342, 354, 7, 3598, 446, 154, 3330, 119, 1473, 403, 3080, 3225, 2146, 3583, 3045, 2052, 1688, 3554, 2580, 3444, 3345, 1972, 3590, 2810, 867, 2772, 787, 2239, 2510, 2055, 2179, 3084, 3542, 2266, 2756, 3373, 2432, 2407, 1164, 2233, 760, 1488, 934, 2459, 2812, 2484, 1149, 3688, 2899, 1417, 3336, 3361, 1548, 1899, 3363, 2229, 3548, 3678, 2199, 774, 2245, 2139, 3239, 1695, 812, 1464, 2502, 1932, 796, 2276, 2505, 3321, 3649]. USORT [1226]. Utilization [572]. Uzawa [3248, 3710].

V [1780]. Valgresultater [15]. validation [1404]. Value [684, 389, 489, 567, 2666, 2568, 412, 414, 664, 608, 479, 387, 45, 2892, 2032, 800, 755, 3159, 3207, 2468, 3444, 2954, 989, 3555, 1054, 1230, 1398, 1371, 3495, 3623, 858, 839, 969, 1125, 1320, 1540, 1642, 3145, 3420, 1576, 2282, 787, 1893, 1029, 1057, 3482, 2266, 3450, 1440, 2478, 957, 1883, 3007, 1774, 1663, 1016, 1713, 1775, 1414, 2136, 1178, 2483, 2886, 3216, 1413, 1475, 2167, 2878, 1215, 1417, 2463, 828, 3352, 937, 1033, 1756, 1885, 3313, 3433, 1717, 1613, 2228, 3289, 1082, 1249, 1318, 849, 1181, 2806, 778, 3674, 1933, 2277, 3511, 1848]. Valued [2328, 748, 3567, 2610, 2818]. Values [2174, 401, 3479, 2861, 2045, 1264, 842, 760, 2191, 934, 2887, 2421, 2360, 2117, 2981, 1321, 1337, 2928]. Vandermonde [2034, 2968, 2636, 1200]. Vandermonde-like [2034, 2968, 2636]. Vanishing [2743, 2953, 2799]. Varför [1]. Variable [411, 3602, 193, 687, 1760, 2010, 431, 110, 3684, 3242, 2293, 2306, 2201, 2531, 2894, 3694, 3177, 3570, 1145, 3359, 3744, 3650, 2357, 919, 3698, 1460, 1886, 1964, 2638, 3466, 3523, 2589, 3536]. variable-precision [1145]. Variable-step [3602, 3694, 3650, 1460]. variables [1698, 2202, 1073, 902, 887, 1367, 2882, 1563, 1932]. variably [3456, 3451]. Variance [216, 1671, 3643]. variant [3094, 2490, 3252]. Variants [3416, 3248]. Variation [239, 2796, 3422, 3032, 3429, 3415]. Variational [3512, 3482, 137, 3603, 3530, 3387, 3648, 3571, 3364, 3141, 3522]. variations [2388]. Varying [386, 788, 3430, 3455, 2676]. Vector [557, 3224, 1688, 1157, 3473, 3263, 2649, 1418, 2818, 156, 1193, 3078]. vector-valued [2818]. vectorize [1787]. vectorizing [1083]. vectors [3111, 2610, 3340]. velocity [2093]. Velte [2441]. Verfahren [84]. Verification [1139, 950, 1620, 2752, 998]. Verified [2797, 3498, 2981]. Verlet [2320, 2010]. Verrier [2157]. version [3404, 2724, 3697, 3667, 1293, 1387, 1022]. versus [2177, 2062]. Vertex [416, 2055]. Vertex-Pair [416]. vertices [1548, 1322]. very [1335, 2854]. Via [2568, 3207, 504, 2354, 3593, 3708, 1699, 1732, 2268, 3544, 3498, 2653, 2924, 3218, 3313, 3181, 2228, 1820, 3454, 3275]. Vibrations [457]. video [3583]. View [144, 254]. Virtual [486, 3698]. viscoelastic [3698]. viscosity [2824]. viscous [2722]. Visibility [1618, 3750]. Visible [1556]. Visual [175]. Visualizing [2840]. vital [1967]. Vlasov [3344, 3663, 3518, 3592, 3558, 3689]. VLSI [1658, 1694]. Volterra [1313, 1445, 966, 3182, 597, 2161, 2201, 2470, 2661, 2103, 352, 418, 2795, 1749, 227, 279, 1487, 546, 639, 656, 3694, 3730, 509, 620, 744, 640, 657, 3557, 2799, 861, 1032, 1102, 1149, 3285, 459, 3679, 1595, 268, 428, 2226, 1254, 3052, 812, 1167, 3077, 2260, 3511, 1106]. Volume [2714, 2364, 2415, 2492, 2545, 2695, 2800, 2434, 2399, 3004, 2721, 2959, 3268, 3546, 2913, 3365, 3078]. volume-finite [3004]. Volume-Preserving [2800, 3078]. Volumes [2377]. Vries [3273]. vs [3525]. VSDKs [3456, 3451]. VSV [2523]. VW [3716]. VW-cycle [3716].

walk [3529, 1041]. Walks [423]. Walls [196]. Wanner [3638]. Waring [1148]. warping [2961]. Washing [300]. Wasted [1074, 1857]. Water [2549, 3371, 2346, 2721]. Wave [11, 290, 2703, 3616, 3160, 1176, 3175, 3176, 2107, 3096, 3153, 3410, 3481, 3698, 3533, 2624, 2727, 2125, 2155, 1019]. wave-guides [3175]. Wave-Operator [11]. wave-power [1176]. waveform [2147, 2535, 3492, 2720, 1963, 2196, 3157]. wavefront [2234]. waveguide [3179]. waveguides [3593]. Wavelet [2238, 2431, 2997, 2699, 2810, 3046, 2587, 2207, 2489]. Wavelet-based [2431, 2997]. wavelet-Galerkin [2810]. Wavelet-Taylor [2699]. wavelets [3017, 2074]. way [1512, 3263, 1573]. ways [2597]. Weak [3418, 3382, 3458, 2025, 3699, 3215, 3020, 3089, 2557, 3105, 2945, 2896, 3629, 3169, 3558, 3155, 3435, 3303, 3523]. Weak-heap [2025]. weakly [3690, 2470, 3697, 1487, 3136, 3679, 2852, 1717, 3275, 3511]. weather [2644, 2764]. Wedin [3498]. Wegematic [106]. Weight [340, 658, 1815, 1669, 2977, 2282, 1948, 1541, 1593, 2443, 2926, 1630, 1589, 3301, 3012, 3203]. Weighted [2744, 2104, 1831, 2401, 2554, 2750, 2739, 3159, 2455, 3736, 1264, 1830, 1908, 3135, 2180, 2069, 2130, 2396, 807, 1823, 2011, 2150, 2948, 777, 2091, 1986]. weighting [2466, 2258]. Weights [446, 2690, 3160, 1710, 1468, 2516, 1199, 1248]. Well [3371, 3256, 3672, 3592, 3737, 3186, 3475, 3103]. Well-balanced [3371, 3186]. well-conditioned [3103]. Well-posedness [3256, 3672, 3592, 3737, 3475]. WENO [3348]. where [847]. whether [3064]. which [1468, 102]. white [3382]. Whittaker [3551]. whole [3014]. whose [1764, 842, 3752, 390]. Widely [2413]. Wiener [2151]. Wilf [1202, 1146]. Williams [3164]. Wilson [1088]. Wire [183, 138]. Wiring [321]. Wisdom [2503]. wise [1892, 3244]. Within [215, 2971, 119, 3638]. Without [664, 804, 2506, 2662, 1981, 585, 1874, 3520, 1734, 1190, 2128]. WKB [3538]. WKB-method [3538]. Work [126, 145]. workarounds [2961]. Working [647]. Worsey [3026]. Worst [2619, 2447, 1344, 1072, 1276, 1587]. Worst-Case [2619, 1344, 1587]. Wrapping [448]. Wright [3030]. WV [3716]. WV-cycle [3716].

X [2303]. Xu [3701].

Year [58, 32]. years [2755, 1437, 3190]. Yee [3047]. yielding [86]. Yosida [2170]. Young [1496, 1655]. Yugoslav [273].

zag [788]. Zakai [3432]. Zakharov [3488]. Zakian [2279]. zentralen [84]. Zero [2562, 127, 247, 301, 797, 843, 1351, 3359, 2868, 3391]. zero-finder [2868]. Zero-One [247, 301]. zero-stability [1351, 3391]. Zeros [720, 2356, 725, 250, 1961, 981, 3732, 1014, 1492, 1441, 2300, 1202, 1146, 1443, 3351, 1579, 2504]. Zeta [286, 157]. Zig [788]. Zig-zag [788]. zipper [3447]. ZP [3307, 3367]. ZPelement [3307, 3367]. zur [84].

References

Bragnum:1961:VTS

[1] Sigurd Bragnum. Varför tunneldioder? (Swedish) [Why tunnel diodes?]. Nordisk tidskrift for informationsbehandling, 1(1):2– 7, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=1&spage=2.

Ehrling:1961:NCI

[2] G. Ehrling. On the numerical computation of incomplete elliptic integrals. Nordisk tidskrift for informationsbehandling, 1(1):8– 14, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=1&spage=8.

Froberg:1961:SIP

[3] Carl-Erik Fröberg. On the sum of inverses of primes and of twin primes. Nordisk tidskrift for informationsbehandling, 1(1):15–20, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=1&
issue=1&spage=15.

Garwick:1961:PLL

[4] Jan V. Garwick. The programming of large logical problems. Nordisk tidskrift for informationsbehandling, 1(1): 21–26, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=1&spage=21.

Jansson:1961:SAS

[5] Birger Jansson. Study of antiaircraft systems by simulation with a Monte Carlo model. Nordisk tidskrift for informationsbehandling, 1(1):27– 37, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=1&spage=27.

Jensen:1961:IAP

[6] J. Jensen and P. Naur. An implementation of ALGOL 60 procedures. Nordisk tidskrift for informationsbehandling, 1(1):38–47, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=1&
issue=1&spage=38.

Kivivuori:1961:MCN

[7] Risto Kivivuori. A method for checking numerical codes using the 1401. Nordisk tidskrift for informationsbehandling, 1(1):48–53, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=1&
issue=1&spage=48.

Langefors:1961:IRFa

[8] Börje Langefors. Information retrieval in file processing I. Nordisk tidskrift for informationsbehandling, 1(1):54– 63, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=1&spage=54.

Anonymous:1961:APa

[9] Anonymous. ALGOL programming. Nordisk tidskrift for informationsbehandling, 1(1):64, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=1&
issue=1&spage=64.

Dopping:1961:AFO

[10] Olle Dopping. ADB för folkbokföring och skatteuppbörd i Sverige. (Swedish) [EDP for people registration and tax processing in Sweden]. Nordisk tidskrift for informationsbehandling, 1(2): 65–68, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=2&spage=65.

Friberg:1961:CSD

[11] Jöran Friberg. Conditionally stable difference approximations for the wave-operator. Nordisk tidskrift for informationsbehandling, 1(2):69– 86, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=2&spage=69.

Garwick:1961:AFP

[12] Jan V. Garwick. The accuracy of floating point computers. Nordisk tidskrift for informationsbehandling, 1(2): 87–88, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=2&spage=87.  See

also comment and reply [27, 28].

Jensen:1961:SAS

[13] J. Jensen, P. Mondrup, and P. Naur. A storage allocation scheme for AL-GOL 60. Nordisk tidskrift for informationsbehandling, 1(2):89–102, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=2&spage=89.

Langefors:1961:IRFb

[14] Börje Langefors. Information retrieval in file processing II. Nordisk tidskrift for informationsbehandling, 1(2):103– 112, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=2&spage=103.

Melbye:1961:BAV

[15] Aage Melbye. Bearbejdning af valgresultater på DASK. (Danish) [Processing of election results on DASK]. Nordisk tidskrift for informationsbehandling, 1(2):113–129, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=2&spage=113.

Riesel:1961:CND

[16] Hans Riesel. A case of numerical divergence. Nordisk tidskrift for informationsbehandling, 1(2):130– 131, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=2&spage=130.

Volby:1961:CIS

[17] Kurt Volby. COBOL: en introduktion. (Swedish) [COBOL: an introduction]. Nordisk tidskrift for informationsbehandling, 1(2):132–137, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=2&spage=132.

Anonymous:1961:N

[18] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 1(2):138– 140, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=2&spage=138.

Anonymous:1961:E

[19] Anonymous. Errata. Nordisk tidskrift for informationsbehandling, 1(2): 140, June 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=2&spage=140.

Anonymous:1961:NN

[20] Anonymous. Notes: Nordsam 1961. Nordisk tidskrift for informationsbehandling, 1(3):138–140, 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Andersen:1961:ALG

[21] Chr. Andersen. The application of the Lichtenstein–Gershgorin integral equation in conformal mapping. Nordisk tidskrift for informationsbehandling, 1(3):141–166, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=3&spage=141.

Comet:1961:FF

[22] Stig Comét. Factorization of factorials. Nordisk tidskrift for informationsbehandling, 1(3):167–176, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=3&spage=167.

Garwick:1961:ISE

[23] Jan V. Garwick. An input system for electronic computers. Nordisk tidskrift for informationsbehandling, 1(3): 177–199, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=3&spage=177.

Krarup:1961:FPD

[24] Torben Krarup. The fixed point division in GIER. Nordisk tidskrift for informationsbehandling, 1(3):200–201, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=3&spage=200.

Petersen:1961:DAD

[25] Bent Scharøe Petersen. Dimensionering af drivtrin for diodeafkodninger. (Danish) [Dimensioning of drive steps for diode decoders]. Nordisk tidskrift for informationsbehandling, 1(3):202–205, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=3&spage=202.

Volby:1961:CGS

[26] K. Volby. COBOL-grammatiken. (Swedish) [COBOL grammar]. Nordisk tidskrift for informationsbehandling, 1 (3):206–219, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=3&spage=206.

Amble:1961:AFP

[27] O. Amble and Jan V. Garwick. On the accuracy of floating point computers [BIT 1(2), 1961, pp. 87–88]. Nordisk tidskrift for informationsbehandling, 1(3):220–222, September 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=1&issue=3&spage=220. See

[12, 28].

Garwick:1961:RAF

[28] J. V. Garwick. Reply to “On the Accuracy of Floating Point Computers” [BIT 1(3), 1961, pp. 220–221].

Nordisk tidskrift for informationsbehandling, 1(3):222, 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [12, 27].

Anonymous:1961:NS

[29] Anonymous. Notice to subscribers. Nordisk tidskrift for informationsbehandling, 1(4):223, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.
com/openurl.asp?genre=article&issn=
0006-3835&volume=1&issue=4&spage=
223.

Alesen:1961:ACT

[30] Erik Magnus Alesen. Analogue calculation of χ^{2} for the testing of hypothesis. Nordisk tidskrift for informationsbehandling, 1(4):224–226, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=4&spage=224.

Alfsen:1961:ACT

[31] E. M. Alfsen. Analogue calculation of χ^{2} for the testing of hypothesis. Nordisk tidskrift for informationsbehandling, 1(4):224–226, 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Fein:1961:CRS

[32] Louis Fein. The computer-related sciences (synnoetics) at a university in the year 1975. Nordisk tidskrift for informationsbehandling, 1(4):227–255, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=4&spage=227.

Froberg:1961:RCA

[33] Carl-Erik Fröberg. Rational Chebyshev approximations of elementary functions. Nordisk tidskrift for informationsbehandling, 1(4):256–262, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=4&spage=256. See

erratum [81].

Volby:1961:CKS

[34] Kurt Volby. COBOL-kompilering. (Swedish) [COBOL compilation]. Nordisk tidskrift for informationsbehandling, 1(4):263–285, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=4&spage=263.

Wynn:1961:TII

[35] P. Wynn. On the tabulation of indefinite integrals. Nordisk tidskrift for informationsbehandling, 1(4):286–289, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=4&spage=286.

Anonymous:1961:APb

[36] Anonymous. ALGOL programming. Nordisk tidskrift for informationsbehandling, 1(4):290, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=1&issue=4&spage=290.

Aaltonen:1962:CTI

[37] Aarre Aaltonen. Computer type instruments. Nordisk tidskrift for informationsbehandling, 2(1):1–6, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=2&
issue=1&spage=1.

Dahl:1962:LOS

[38] Ole-Johan Dahl. Litt om symbolbruken i Algol. (Norwegian) [A little about symbol use in Algol]. Nordisk tidskrift for informationsbehandling, 2(1): 7–8, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=1&spage=7.

Greko:1962:CAM

[39] Bertil Greko. Computer analysis of medical history as an aid to diagnosis. Nordisk tidskrift for informationsbehandling, 2(1):9– 15, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=1&spage=9.

Karlqvist:1962:AMT

[40] Olle Karlqvist. Applications to the magnetic tape storage unit, Facit ECM 64 (the carousel memory). Nordisk tidskrift for informationsbehandling, 2(1): 16–20, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=1&spage=16.

Langefors:1962:ANP

[41] Börje Langefors. Activity network for planning and scheduling. Nordisk tidskrift for informationsbehandling, 2(1): 21–34, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=1&spage=21.

Melbye:1962:NER

[42] Aage Melbye and Poul Sveistrup. The need for education and research in administrative data processing. Nordisk tidskrift for informationsbehandling, 2(1):35–44, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=2&
                                        issue=1&spage=35.

Thorelli:1962:ACE

[43] Lars Erik Thorelli. Automatic correction of errors in text. Nordisk tidskrift for informationsbehandling, 2(1): 45–52, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=1&spage=45.

Wittmeyer:1962:RAE

[44] Linde Wittmeyer. Rational approximation of empirical functions. Nordisk tidskrift for informationsbehandling, 2(1): 53–60, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=1&spage=53.

Wynn:1962:NSC

[45] P. Wynn. Note on the solution of a certain boundary-value problem. Nordisk tidskrift for informationsbehandling, 2(1):61–64, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=1&spage=61.

Lokki:1962:NBJ

[46] Olli Lokki. New books and journals. Nordisk tidskrift for informationsbehandling, 2(1):65–67, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=2&
issue=1&spage=65.

Anonymous:1962:Na

[47] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 2(1):66– 67, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=1&spage=66.

Appel:1962:RAD

[48] Klaus Appel. Rational approximation of decay-type functions. Nordisk tidskrift for informationsbehandling, 2(2): 69–75, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=2&spage=69.

Asker:1962:SCS

[49] Bengt Asker. The spline curve, a smooth interpolating function used in numerical design of shiplines. Nordisk tidskrift for informationsbehandling, 2(2):76–82, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=2&
                                        issue=2&spage=76.

Hyvarinen:1962:CQD

[50] Lassi Hyvärinen. Classification of qualitative data. Nordisk tidskrift for informationsbehandling, 2(2):83– 89, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=2&issue=2&spage=83.

Karst:1962:SND

[51] Edgar Karst. Some new divisors of Mersenne numbers. Nordisk tidskrift for informationsbehandling, 2(2): 90, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=2&spage=90.

Langefors:1962:CPR

[52] Börje Langefors. Computation of parts requirements for production scheduling. Nordisk tidskrift for informationsbehandling, 2(2):91–111, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=2&spage=91.

Petersen:1962:SLD

[53] Bent Scharøe Petersen. Styrekredse for linieskriver. (Danish) [Control circuit for a line printer]. Nordisk tidskrift for informationsbehandling, 2(2):112– 117, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=2&spage=112.

Melbye:1962:NBJ

[54] Aage Melbye. New books and journals. Nordisk tidskrift for informationsbehandling, 2(2):118–119, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=2&spage=118.

Anonymous:1962:ICC

[55] Anonymous. International computation centre (ICC). Nordisk tidskrift for informationsbehandling, 2(2):120– 121, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=2&spage=120.

Anonymous:1962:ICMa

[56] Anonymous. IFIP Congress 1962, Munich, August 27 to September 1, 1962. Nordisk tidskrift for informationsbehandling, 2(2):122–132, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=2&spage=122.

Anonymous:1962:ICMb

[57] Anonymous. International Congress of Mathematicians, Stockholm. Nordisk tidskrift for informationsbehandling, 2(2):132–136, June 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=2&spage=132.

Dahlstrand:1962:HYE

[58] Ingemar Dahlstrand. A half year’s experience with the Facit-Algol 1 compiler. Nordisk tidskrift for informationsbehandling, 2(3):137–142, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=3&spage=137.

Kredell:1962:CSO

[59] Bengt Kredell. On complex successive overrelaxation. Nordisk tidskrift for informationsbehandling, 2(3):143–152, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=3&spage=143.

Kreiss:1962:SDG

[60] Heinz-Otto Kreiss. Über die Stabilitätsdefinition für Differenzen gleichungen die partielle Differentialgleichungen approximieren. (German) [On the stability definition for difference equations of partial differential equation approximations]. Nordisk tidskrift for informationsbehandling, 2(3):153–181, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=3&spage=153.

Sundstrom:1962:SNC

[61] Lars-Olof Sundström. Saab 500 — A numerical control system. Nordisk tidskrift for informationsbehandling, 2(3):182–191, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=3&spage=182.

Anonymous:1962:APC

[62] Anonymous. Algol programming: Contribution no. 3: Calculation of complementary Fresnel integrals. Nordisk tidskrift for informationsbehandling, 2(3):192–194, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=3&spage=192.

Christiansen:1962:APC

[63] Søren Christiansen. Algol programming: Contribution no. 3: Calculation of complementary Fresnel integrals. Nordisk tidskrift for informationsbehandling, 2(3):192–194, 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1962:NBJb

[64] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 2(3):195–196, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=3&spage=195.

Dopping:1962:TPU

[65] O. Dopping. Test problems used for evaluation of computers. Nordisk tidskrift for informationsbehandling, 2(4):

197–202, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=4&spage=197.

Eriksson:1962:SPQ

[66] L. Eriksson, O. Lokki, and N. Ryti. On smoothing of pulp quality characteristics in a flow system. Nordisk tidskrift for informationsbehandling, 2 (4):203–211, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=4&spage=203.

Golub:1962:BRE

[67] Gene H. Golub. Bounds for the roundoff errors in the Richardson second order method. Nordisk tidskrift for informationsbehandling, 2(4):212–223, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=4&spage=212.

Karst:1962:SLD

[68] Edgar Karst. Search limits on divisors of Mersenne numbers. Nordisk tidskrift for informationsbehandling, 2 (4):224–227, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=4&spage=224.

Shen:1962:GPC

[69] Mok-Kong Shen. On the generation of permutations and combinations. Nordisk tidskrift for informationsbehandling, 2(4):228–231, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=4&spage=228.

Wynn:1962:AAP

[70] P. Wynn. An arsenal of AL-GOL procedures for complex arithmetic. Nordisk tidskrift for informationsbehandling, 2(4):232–255, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=4&spage=232.

Andersen:1962:APC

[71] Christian Andersen. Algol programming: Contribution no. 4: Solution of linear equations. Nordisk tidskrift for informationsbehandling, 2(4):256–257, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=2&issue=4&spage=256. See

remark [78].

Sigurd:1962:NBJ

[72] Bengt Sigurd. New books and journals. Nordisk tidskrift for informationsbehandling, 2(4):258–259, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=4&spage=258.

Anonymous:1962:Nb

[73] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 2(4): 260–262, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=2&issue=4&spage=260.

Bubenko:1963:MOS

[74] Janis A. Bubenko. Multiprogrammering — en orientering. (Swedish) [Multiprogramming: an orientation]. Nordisk tidskrift for informationsbehandling, 3(1):1–26, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=3&
issue=1&spage=1.

Dahlquist:1963:SSP

[75] Germund G. Dahlquist. A special stability problem for linear multistep methods. Nordisk tidskrift for informationsbehandling, 3(1):27– 43, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=1&spage=27.

Hansen:1963:FCR

[76] Leif Prøhl Hansen and Bent Scharøe Petersen. A fast card reader for the GIER computer. Nordisk tidskrift for informationsbehandling, 3(1):44– 51, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=1&spage=44.

Karlgren:1963:RTS

[77] Hans Karlgren. Representation of text strings in binary computers. Nordisk tidskrift for informationsbehandling, 3(1):52–59, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=3&
                                        issue=1&spage=52.

Bohman:1963:APR

[78] Jan Bohman. Algol programming: Remark on algorithm no. 4, LLGAUSS (Chr. Andersen, BIT 2, 4 (1962)). Nordisk tidskrift for informationsbehandling, 3(1):60, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=1&spage=60.  See

[71].

Anonymous:1963:NBJa

[79] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 3(1):61–66, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=1&spage=61.

Anonymous:1963:Na

[80] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 3(1):67– 68, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=1&spage=67.

Froberg:1963:ERC

[81] C.-E. Fröberg. Erratum: “Rational Chebyshev Approximations of Elementary Functions” [BIT 1(4), 1961, p. 261, line 12]. Nordisk tidskrift for informationsbehandling, 3(1):68, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=1&spage=68.  See

[33].

Carter:1963:OT

[82] I. P. V. Carter. Optimization techniques. Nordisk tidskrift for informationsbehandling, 3(2):69–92, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=2&spage=69.

Hult:1963:PNH

[83] Ture Hult. Presentation of a new highspeed paper tape reader. Nordisk tidskrift for informationsbehandling, 3(2): 93–96, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=2&spage=93.

Johansson:1963:VZD

[84] Olov Johansson and Heinz-Otto Kreiss. Über das Verfahren der zentralen Differenzen zur Lösung des Cauchy-Problems für partielle Differentialgleichungen. (German) [On the method of central differences for the solution of Cauchy problems for partial differential equations]. Nordisk tidskrift for informationsbehandling, 3(2):97– 107, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=2&spage=97.

Johansson:1963:RPP

[85] Carl-Axel Johansson. Requirements planning of production components and spare parts at a farm equipment manufacturing company. Nordisk tidskrift for informationsbehandling, 3(2): 108–121, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=2&spage=108.

Karst:1963:RQY

[86] Edgar Karst. A remarkable quartic yielding certain divisors of Mersenne numbers. Nordisk tidskrift for informationsbehandling, 3(2):122–123, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=2&spage=122.

Naur:1963:DGAa

[87] Peter Naur. The design of the GIER ALGOL compiler. Part I. Nordisk tidskrift for informationsbehandling, 3(2): 124–140, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=2&spage=124.

Anonymous:1963:APC

[88] Anonymous. Algol programming: Contribution no. 5: Computation of the Fermi function. Nordisk tidskrift for informationsbehandling, 3(2):141– 142, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=2&spage=141.

Froberg:1963:APC

[89] Carl-Erik Fröberg. Algol programming: Contribution no. 5: Computation of the Fermi function. Nordisk tidskrift for informationsbehandling, 3 (2):141–142, 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1963:NBJb

[90] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 3(2):143–144, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=2&spage=143.

Naur:1963:DGAb

[91] Peter Naur. The design of the GIER ALGOL compiler. Part II. Nordisk tidskrift for informationsbehandling, 3(3):145–166, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=3&spage=145.

Jiewertz:1963:MAE

[92] Bengt Jiewertz. Miniatyrisering av elektronikkomponenter. (Swedish) [Miniaturization of electronics components]. Nordisk tidskrift for informationsbehandling, 3(3):167–174, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=3&spage=167.

Wynn:1963:SRC

[93] P. Wynn. Singular rules for certain non-linear algorithms. Nordisk tidskrift for informationsbehandling, 3(3): 175–195, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=3&spage=175.

Hestvik:1963:PGN

[94] Ole R. Hestvik and Hans J. Levold. Prosessdataenhet for GIER. (Norwegian) [Processing unit for GIER]. Nordisk tidskrift for informationsbehandling, 3(3):196–203, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=3&spage=196.

Naur:1963:APG

[95] Peter Naur. Algol programming: Goto statements and good Algol style. Nordisk tidskrift for informationsbehandling, 3(3):204–208, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=3&spage=204.

Anonymous:1963:NBJc

[96] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 3(3):209–211, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=3&spage=209.

Anonymous:1963:Nb

[97] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 3(3): 211–212, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=3&spage=211.

Anonymous:1963:Eb

[98] Anonymous. Errata. Nordisk tidskrift for informationsbehandling, 3(3):212, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=3&spage=212.

Danielsson:1963:CCR

[99] Per-Erik Danielsson. Converting a curve to right-angled increments. Nordisk tidskrift for informationsbehandling, 3(4):213–221, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=4&spage=213.

Karst:1963:LAP

[100] Edgar Karst. List of all prime divisors q = 2kp + 1 of 2^{p} − 1, K < 10, p < 15000. Nordisk tidskrift for informationsbehandling, 3(4):222–228, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=4&spage=222.

Langefors:1963:SAT

[101] Börje Langefors. Some approaches to the theory of information systems. Nordisk tidskrift for informationsbehandling, 3(4):229–254, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=4&spage=229.

Riesel:1963:WOD

[102] Hans Riesel. In which order are different conditions to be examined? Nordisk tidskrift for informationsbehandling, 3(4):255–256, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=4&spage=255.

Olsen:1963:AES

[103] T. Vahl Olsen. Analysis of elastic structures on digital computers. Nordisk tidskrift for informationsbehandling, 3(4):257–272, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=3&issue=4&spage=257.

Anonymous:1963:NBJd

[104] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 3(4):273–274, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=4&spage=273.

Anonymous:1963:Nc

[105] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 3(4):275, December 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=3&issue=4&spage=275.

Appel:1964:CAW

[106] Klaus Appel. A compiler for the Alwac–Wegematic computer. Nordisk tidskrift for informationsbehandling, 4 (1):1–5, March 1964. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=1&spage=1.

Jansson:1964:ABP

[107] Birger Jansson. Autocorrelations between pseudo-random numbers. Nordisk tidskrift for informationsbehandling, 4(1):6–27, March 1964. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0006-3835&volume=4&issue=1&spage=
                                        6.

Karst:1964:SND

[108] Edgar Karst. Some new divisors of Mersenne numbers. Nordisk tidskrift for informationsbehandling, 4(1):28– 29, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=1&spage=28.

Lindecrantz:1964:MOC

[109] N. Lindecrantz. Method for optimum cutting of rectangular sheets. Nordisk tidskrift for informationsbehandling, 4(1):30–35, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=1&spage=30.

Tollet:1964:LLS

[110] Gustav Tollet. A labelling and look-up system for handling variable format records. Nordisk tidskrift for informationsbehandling, 4(1):36– 45, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=1&spage=36.

Ostlund:1964:IR

[111] Lars Östlund. Information retrieval. Nordisk tidskrift for informationsbehandling, 4(1):46–53, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=4&
issue=1&spage=46.

Gram:1964:APC

[112] Christian Gram. Algol programming: Contribution no. 8: Definite integrals by Romberg’s method. Nordisk tidskrift for informationsbehandling, 4(1): 54–60, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=1&spage=54.

Anonymous:1964:Pa

[113] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 4(1): 61–64, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=1&spage=61.

Anonymous:1964:NBJa

[114] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 4(1):65–68, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=1&spage=65.

Anonymous:1964:Na

[115] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 4(1): 68, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=1&spage=68.

Axelsson:1964:GID

[116] Owe Axelsson. Global integration of differential equations through Lobatto quadrature. Nordisk tidskrift for informationsbehandling, 4(2):69– 86, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=2&spage=69.

Bengtsson:1964:CII

[117] Bengt-Erik Bengtsson and Stig Nordbeck. Construction of isarithms and isarithmic maps by computers. Nordisk tidskrift for informationsbehandling, 4(2):87–105, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=2&spage=87.

Berghuis:1964:PST

[118] J. Berghuis, A. J. van der Heiden, and R. Bakker. The preparation of school time tables by electronic computer. Nordisk tidskrift for informationsbehandling, 4(2):106– 114, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=2&spage=106.

Naur:1964:UMC

[119] Peter Naur. Using machine-code within an ALGOL system. Nordisk tidskrift for informationsbehandling, 4(2):115– 117, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=2&spage=115.

Gram:1964:APN

[120] Christian Gram. Algol programming: Note on “Definite Integral by Romberg’s Method” [BIT 4(1), 1964, pp. 54–60]. Nordisk tidskrift for informationsbehandling, 4(2):118– 119, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=2&spage=118.

Naur:1964:APC

[121] Peter Naur. Algol programming: Contribution no. 9: Eigenvalues and eigenvectors of real symmetric matrices. Nordisk tidskrift for informationsbehandling, 4(2):120–130, 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See erratum [135].

Anonymous:1964:Pb

[122] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 4(2): 131–133, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=2&spage=131.

Anonymous:1964:NBJb

[123] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 4(2):134, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=4&
issue=2&spage=134.

Anonymous:1964:NGA

[124] Anonymous. Notes: The GIER AL-GOL III compiler. Nordisk tidskrift for informationsbehandling, 4(2):135– 136, June 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=2&spage=135.

Garwick:1964:DSC

[125] Jan v. Garwick. Data storage in compilers. Nordisk tidskrift for informationsbehandling, 4(3):137–140, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=137.

Gould:1964:TWI

[126] I. H. Gould and G. C. Tootill. The terminology work of IFIP and ICC. Nordisk tidskrift for informationsbehandling, 4(3):141–155, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=141.

Gram:1964:RZF

[127] Christian Gram. On the representation of zero in floating-point arithmetic. Nordisk tidskrift for informationsbehandling, 4(3):156–161, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=156.

Langefors:1964:AGP

[128] Börje Langefors. ALGOL-GENIUS, a programming language for general data processing. Nordisk tidskrift for informationsbehandling, 4(3):162–176, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=162.

Naur:1964:AGS

[129] Peter Naur. Automatic grading of students’ ALGOL programming. Nordisk tidskrift for informationsbehandling, 4(3):177–188, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=177.

Josefsen:1964:APC

[130] Harold Josefsen. Algol programming: Contribution no. 10: Linear programming by the modified simplex method. Nordisk tidskrift for informationsbehandling, 4(3):189–196, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=189. See

certification [188, 498, 499].

Anonymous:1964:Pc

[131] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 4(3): 197–201, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=197.

Anonymous:1964:NBJc

[132] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 4(3):202–203, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=202.

Anonymous:1964:E

[133] Anonymous. Erratum. Nordisk tidskrift for informationsbehandling, 4(3):204, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=3&spage=204.

Anonymous:1964:NIS

[134] Anonymous. Notes: International seminar on automatic data processing (ADP) for top management in public administration. Nordisk tidskrift for informationsbehandling, 4(3):204, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=3&spage=204.

Naur:1964:EAP

[135] Peter Naur. Erratum: “Algol Programming: Contribution no. 9: Eigenvalues and eigenvectors of real symmetric matrices” [BIT 4(2), 1964, p. 127]. Nordisk tidskrift for informationsbehandling, 4(3):204, 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [121].

Jansson:1964:GRB

[136] Birger Jansson. Generation of random bivariate normal deviates and computation of related integrals. Nordisk tidskrift for informationsbehandling, 4(4):205–212, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=4&spage=205.

Mehlum:1964:CFM

[137] Even Mehlum. A curve-fitting method based on a variational criterion. Nordisk tidskrift for informationsbehandling, 4(4):213–223, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=4&spage=213.

Norden:1964:PHC

[138] Harry v. Nordén and Ilari Seppä. A problem on heat conduction in a wire enclosed by a cylindrical material solved by numerical inverse Laplace transform. Nordisk tidskrift for informationsbehandling, 4(4):224–242, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=4&spage=224.

Shen:1964:CGC

[139] Mok-Kong Shen. On checking the Goldbach conjecture. Nordisk tidskrift for informationsbehandling, 4(4): 243–245, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=4&spage=243.

Thorelli:1964:HDO

[140] Lars-Erik Thorelli. Hjärnan, datamaskinen och matematiken. (Swedish) [The brain, computers, and mathematics]. Nordisk tidskrift for informationsbehandling, 4(4):246–249, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=4&issue=4&spage=246.

Anonymous:1964:APd

[141] Anonymous. Algol programming. Nordisk tidskrift for informationsbehandling, 4(4):250–260, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=4&spage=250.

Anonymous:1964:Pd

[142] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 4(4): 261–265, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=4&spage=261.

Anonymous:1964:Nd

[143] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 4(4): 266–267, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=4&issue=4&spage=266.

Dahlquist:1965:CAP

[144] Germund Dahlquist, Sven-Åke Gustafson, and Károly Siklósi. Convergence acceleration from the point of view of linear programming. Nordisk tidskrift for informationsbehandling, 5(1): 1–16, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=1&spage=1.

Kostamo:1965:CAS

[145] Eero Kostamo. The contents of ADP system work in administrative ADP applications. Nordisk tidskrift for informationsbehandling, 5(1):17– 24, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=1&spage=17.

Laurikainen:1965:ICT

[146] J. J. Laurikainen. On the influence of coordinate transformations on the Lagrangian polynomial and its evaluation in curve representation. Nordisk tidskrift for informationsbehandling, 5(1): 25–36, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=1&spage=25.

Moller:1965:QDP

[147] Ole Møller. Quasi double-precision in floating point addition. Nordisk tidskrift for informationsbehandling, 5(1): 37–50, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=1&spage=37.  See

also [172].

Stiefel:1965:NNM

[148] E. Stiefel. Neuere numerische Methoden in der Himmelsmechanik. (German) [Newer numerical methods in celestial mechanics]. Nordisk tidskrift for informationsbehandling, 5(1):51– 60, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=1&spage=51.

Zachariassen:1965:APC

[149] J. Zachariassen. Algol programming: Contribution no. 12: Solution of systems of linear equations by Gaussian elimination. Nordisk tidskrift for informationsbehandling, 5(1):61– 67, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=1&spage=61.

Anonymous:1965:Na

[150] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 5(1): 67, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=1&spage=67.

Anonymous:1965:Pa

[151] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 5(1): 68–72, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=1&spage=68.

Dopping:1965:STA

[152] O. Dopping. Suggestion for teaching of administrative data processing at Swedish universities. Nordisk tidskrift for informationsbehandling, 5(2): 73–84, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=2&spage=73.

Duncan:1965:PRO

[153] F. G. Duncan. Possibilities for refining an object program compiled with an ALGOL translator. Nordisk tidskrift for informationsbehandling, 5(2): 85–95, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=2&spage=85.

Langefors:1965:ISD

[154] Börje Langefors. Information system design computations using generalized matrix algebra. Nordisk tidskrift for informationsbehandling, 5(2): 96–121, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=2&spage=96.

Overholt:1965:EAA

[155] K. J. Overholt. Extended Aitken acceleration. Nordisk tidskrift for informationsbehandling, 5(2):122–132, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=2&spage=122.

Thorelli:1965:APC

[156] Lars Erik Thorelli. Algol programming: Contribution no. 13: Evaluation of a function expression described by a numerical vector. Nordisk tidskrift for informationsbehandling, 5(2):133– 141, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=2&spage=133.

Markman:1965:RZF

[157] B. Markman. The Riemann zeta function. Nordisk tidskrift for informationsbehandling, 5(2):138–141, 1965. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1965:Pb

[158] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 5(2): 142–146, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=2&spage=142.

Anonymous:1965:NBJa

[159] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 5(2):147, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=5&
                                        issue=2&spage=147.

Anonymous:1965:Nb

[160] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 5(2):148– 150, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=2&spage=148.

Naur:1965:COT

[161] Peter Naur. Checking of operand types in ALGOL compilers. Nordisk tidskrift for informationsbehandling, 5(3): 151–163, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=3&spage=151.

Reimer:1965:IPI

[162] Manfred Reimer. An integration procedure including error estimation. Nordisk tidskrift for informationsbehandling, 5(3):164–174, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=3&spage=164.

Thorelli:1965:FSA

[163] Lars-Erik Thorelli. Finite synchronous automata. Nordisk tidskrift for informationsbehandling, 5(3):175–202, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=3&spage=175.

Villumsen:1965:SNE

[164] P. V. Villumsen. On the solution of normal equations p_{1}. Nordisk tidskrift for informationsbehandling, 5(3):203–210, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=3&spage=203.

Ditlevsen:1965:RLR

[165] O. Ditlevsen. A remark on the Lagrangian remainder in Taylor’s formula. Nordisk tidskrift for informationsbehandling, 5(3):211–213, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=3&spage=211.

Anonymous:1965:Pc

[166] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 5(3): 214–218, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=3&spage=214.

Anonymous:1965:NBJb

[167] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 5(3):219, September 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=3&spage=219.

Aasen:1965:LEG

[168] Jan Ole Aasen and Werner Romberg. Eine Lösungsmethode für Eigenwertprobleme. (German) [A solution method for eigenvalue problems]. Nordisk tidskrift for informationsbehandling, 5(4):221–229, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=221.

Froberg:1965:TCM

[169] Carl-Erik Fröberg. On triangularization of complex matrices by two-dimensional unitary transformations. Nordisk tidskrift for informationsbehandling, 5(4):230–234, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=230.

Jensen:1965:GMC

[170] Jørn Jensen. Generation of machine code in ALGOL compilers. Nordisk tidskrift for informationsbehandling, 5(4):235–245, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=235.

Knuth:1965:CRS

[171] Donald E. Knuth. Construction of a random sequence. Nordisk tidskrift for informationsbehandling, 5(4):246–250, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=4&spage=246.

Moller:1965:NQD

[172] Ole Møller. Note on quasi doubleprecision. Nordisk tidskrift for informationsbehandling, 5(4):251–255, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=4&spage=251. See

[147].

Sveistrup:1965:SCE

[173] Poul Sveistrup. The systems concept and the enterprise as a system. Nordisk tidskrift for informationsbehandling, 5(4):256–275, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=4&spage=256.

Timlake:1965:AMR

[174] W. P. Timlake. On an algorithm of Milne and Reynolds. Nordisk tidskrift for informationsbehandling, 5(4):276–281, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=5&issue=4&spage=276.

Dopping:1965:OPP

[175] Olle Dopping. Optimum printout policy for visual reference files.

Nordisk tidskrift for informationsbehandling, 5(4):282–286, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=282.

Overholt:1965:IFG

[176] K. J. Overholt. An instability in the Fibonacci and Golden Section search methods. Nordisk tidskrift for informationsbehandling, 5(4):284–286, 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Christiansen:1965:APE

[177] S. Christiansen. Algol programming: Error integral with complex argument. Nordisk tidskrift for informationsbehandling, 5(4):287–293, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=287.

Anonymous:1965:Pd

[178] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 5(4): 294–296, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=294.

Anonymous:1965:NBJc

[179] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 5(4):297–299, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=297.

Anonymous:1965:PSA

[180] Anonymous. Post-secondary ADP education in Sweden. Nordisk tidskrift for informationsbehandling, 5(4):300–307, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=5&issue=4&spage=300.

Hansen:1966:CCS

[181] Per Brinch Hansen and Roger House. The COBOL compiler for the Siemens 3003. Nordisk tidskrift for informationsbehandling, 6(1):1–23, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=6&
issue=1&spage=1.

Haavie:1966:MRA

[182] T. Håvie. On a modification of Romberg’s algorithm. Nordisk tidskrift for informationsbehandling, 6(1):24– 30, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=24.

Jarvelainen:1966:PHC

[183] Martti Järveläinen and Harry V. Nordén. A problem on heat conduction in an insulated wire solved by numerical inverse Laplace transform. Nordisk tidskrift for informationsbehandling, 6(1):31–47, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=31.

Mayoh:1966:SGC

[184] Brian H. Mayoh. On the second Goldbach conjecture. Nordisk tidskrift for informationsbehandling, 6(1):48– 50, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=48.

Osborne:1966:NMN

[185] M. R. Osborne. On Nordsieck’s method for the numerical solution of ordinary differential equations. Nordisk tidskrift for informationsbehandling, 6(1):51– 57, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=51.

Weizenbaum:1966:LUL

[186] Joseph Weizenbaum. On-line user languages. Nordisk tidskrift for informationsbehandling, 6(1):58–65, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=58.

vanWijngaarden:1966:NAI

[187] A. van Wijngaarden. Numerical analysis as an independent science. Nordisk tidskrift for informationsbehandling, 6(1):66–81, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=6&
issue=1&spage=66.

Asker:1966:APC

[188] Bengt Asker. Algol programming: Certification of contribution no. 10: LPMODSIMPLEX [BIT 4(3), 1964, pp. 189–196]. Nordisk tidskrift for informationsbehandling, 6(1):82, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=82.  See

[130, 498, 499].

Anonymous:1966:Pa

[189] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 6(1): 83–84, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=83.

Anonymous:1966:NBJa

[190] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 6(1):85–86, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=85.

Anonymous:1966:Na

[191] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 6(1):87– 88, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=1&spage=87.

Andersson:1966:MCM

[192] S.-E. Andersson. A Monte Carlo model for simulation of tank battles. Nordisk tidskrift for informationsbehandling, 6(2):89–100, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=2&spage=89.

Garwick:1966:SSS

[193] Jan V. Garwick. The summation of some series with variable coefficients by approximate analytical expressions. Nordisk tidskrift for informationsbehandling, 6(2):101– 116, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=2&spage=101.

Gustafson:1966:CAM

[194] Sven-Åke Gustafson. Convergence acceleration by means of numerical quadrature. Nordisk tidskrift for informationsbehandling, 6(2):117– 128, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=2&spage=117.

Johnsen:1966:EPM

[195] Th. Lunde Johnsen. Efficiency of the polyphase merge and a related method. Nordisk tidskrift for informationsbehandling, 6(2):129–143, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=2&spage=129.

Norden:1966:THC

[196] H. V. Nordén and I. Seppä. On transient heat conduction in the walls of a rectangular gas channel. Nordisk tidskrift for informationsbehandling, 6(2): 144–154, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=2&spage=144.

Palme:1966:SSO

[197] J. Palme. SOFT — simple output format translator. Nordisk tidskrift for informationsbehandling, 6(2):155– 157, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=2&spage=155.

Selmer:1966:NEP

[198] E. S. Selmer. On Newton’s equations for the power sums. Nordisk tidskrift for informationsbehandling, 6(2):158– 160, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=2&spage=158.

Strom:1966:IPF

[199] Torsten Ström. An inductive proof of the fundamental theorem of games. Nordisk tidskrift for informationsbehandling, 6(2):161–163, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=2&spage=161.

Svejgaard:1966:APC

[200] Bj. Svejgaard. Algol programming: Contribution no. 17: List processing. Nordisk tidskrift for informationsbehandling, 6(2):164–175, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=2&spage=164.

Anonymous:1966:Pb

[201] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 6(2): 176–179, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=2&spage=176.

Anonymous:1966:NBJb

[202] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 6(2):180, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=6&
                                        issue=2&spage=180.

Day:1966:QMA

[203] J. T. Day. Quadrature methods of arbitrary order for solving linear ordinary differential equations. Nordisk tidskrift for informationsbehandling, 6(3):181–190, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=3&spage=181.

Froberg:1966:NSM

[204] Carl-Erik Fröberg. Numerical studies of the Möbius power series. Nordisk tidskrift for informationsbehandling, 6(3):191–211, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=3&spage=191.

Seppala:1966:OMR

[205] Yrjö Seppälä. On optimization of the maintenance of a register. Nordisk tidskrift for informationsbehandling, 6(3):212–227, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=3&spage=212.

Wynn:1966:CCF

[206] P. Wynn. On the computation of certain functions of large argument and parameter. Nordisk tidskrift for informationsbehandling, 6(3):228–259, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=3&spage=228.

Hellgren:1966:APC

[207] Bo Hellgren. Algol programming: Contribution no. 18: Procedures for simplifying Boolean expressions. Nordisk tidskrift for informationsbehandling, 6(3):260–269, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=3&spage=260.

Anonymous:1966:Pc

[208] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 6(3): 270–272, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=3&spage=270.

Anonymous:1966:NBJc

[209] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 6(3):273, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=3&spage=273.

Anonymous:1966:Nb

[210] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 6(3): 274–276, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=3&spage=274.

Glass:1966:SCI

[211] J. M. Glass. Smooth-curve interpolation: a generalized splinefit procedure. Nordisk tidskrift for informationsbehandling, 6(4):277– 293, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=4&spage=277.

Johansen:1966:CRD

[212] Peter Johansen. Construction of recognition devices for regular languages from their Backus Normal Form definition. Nordisk tidskrift for informationsbehandling, 6(4):294– 309, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=4&spage=294.

Naur:1966:PAG

[213] Peter Naur. Proof of algorithms by general snapshots. Nordisk tidskrift for informationsbehandling, 6(4):310– 316, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=4&spage=310.

Norden:1966:CCC

[214] Harry V. Nordén and Ilari Seppä. Cooling of a cylindrical container blasted into the rock. Nordisk tidskrift for informationsbehandling, 6(4):317– 331, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=4&spage=317.

Tienari:1966:SPM

[215] M. Tienari and V. Suokonautio. A set of procedures making real arithmetic of unlimited accuracy possible within Algol 60. Nordisk tidskrift for informationsbehandling, 6(4):332– 338, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=4&spage=332.

Torn:1966:CMC

[216] A. Törn. Crude Monte Carlo quadrature in infinite variance case and the Central Limit Theorem. Nordisk tidskrift for informationsbehandling, 6(4): 339–346, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=4&spage=339.

Thorelli:1966:ACA

[217] Lars-Erik E. Thorelli. An algorithm for computing all paths in a graph. Nordisk tidskrift for informationsbehandling, 6(4):347–349, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=4&spage=347.

Ruhe:1966:APE

[218] Axel Ruhe. Algol programming: Eigenvalues of a complex matrix by the QR-method. Nordisk tidskrift for informationsbehandling, 6(4):350– 358, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=6&issue=4&spage=350.

Anonymous:1966:Pd

[219] Anonymous. Problems. Nordisk tidskrift for informationsbehandling, 6(4): 359–361, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=4&spage=359.

Anonymous:1966:NBJd

[220] Anonymous. New books and journals. Nordisk tidskrift for informationsbehandling, 6(4):362–365, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=4&spage=362.

Anonymous:1966:Nc

[221] Anonymous. Notes. Nordisk tidskrift for informationsbehandling, 6(4): 366, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=6&issue=4&spage=366.

Bjorck:1967:SLL

[222] Åke Björck. Solving linear least squares problems by Gram–Schmidt orthogonalization. BIT, 7(1):1–21, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=1.

Cohen:1967:RGC

[223] Jacques Cohen and Laurent Trilling. Remarks on ‘garbage collection’ using a two-level storage. BIT, 7(1):22– 30, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=22.

Gourlay:1967:IBC

[224] A. R. Gourlay and A. R. Mitchell. Intermediate boundary corrections for split operator methods in three dimensions. BIT, 7(1):31–38, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=7&
issue=1&spage=31.

Nordbeck:1967:CCP

[225] Stig Nordbeck and Bengt Rystedt. Computer cartography pointin-polygon programs. BIT, 7(1):39– 64, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=39.

Widlund:1967:NUS

[226] O. B. Widlund. A note on unconditionally stable linear multistep methods. BIT, 7(1):65–70, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=65.

Day:1967:NSL

[227] J. T. Day. On the numerical solution of linear Volterra integral equations. BIT, 7(1):71–72, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=71.

James:1967:NOR

[228] W. James and P. Jarratt. A note on optimum ranges for the generation of sin and cos on a digital computer. BIT, 7(1):73–75, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=73.

Riesel:1967:CFA

[229] Hans Riesel. A continued fraction algorithm. BIT, 7(1):76–80, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=1&spage=76.

Anonymous:1967:Pa

[230] Anonymous. Problems. BIT, 7(1):81– 83, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=7&issue=1&spage=81.

Anonymous:1967:NBJa

[231] Anonymous. New books and journals. BIT, 7(1):84–86, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=7&
                                        issue=1&spage=84.

Anonymous:1967:NIC

[232] Anonymous. Notes: IFIP Congress 68. BIT, 7(1):87–90, March 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=7&issue=1&spage=87.

Gunnarsson:1967:PSR

[233] L. Gunnarsson. Program for simulation of relay networks. BIT, 7(2): 91–95, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=7&issue=2&spage=91.

Hansen:1967:STS

[234] H. B. Hansen and Alan Wessel. Subroutinized tape sorting. BIT, 7(2):96– 102, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=7&issue=2&spage=96.

Haavie:1967:PAM

[235] T. Håvie. On the practical application of the modified Romberg algorithm. BIT, 7(2):103–113, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=103.

Karst:1967:TFP

[236] Edgar Karst. Tables on fifth power residuacity. BIT, 7(2):114– 122, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=114.

Naur:1967:MDP

[237] Peter Naur. Machine dependent programming in common languages. BIT, 7(2):123–131, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=123.

Nord:1967:APS

[238] Stig Nord. Approximation properties of the spline fit. BIT, 7(2):132– 144, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=132.

Norden:1967:STV

[239] Harry v. Nordén and Ilari Seppä. Seasonal temperature variation of soil. BIT, 7(2):145–155, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=7&
issue=2&spage=145.

Sarafyan:1967:CDA

[240] Diran Sarafyan and Robert Brown. Computer derivation of algebraic equations associated with Runge– Kutta formulas. BIT, 7(2):156– 162, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=156.

Froberg:1967:APC

[241] Carl-Erik Fröberg and Arne Sundström. Algol programming: Contribution no. 20: Smith’s normal form. BIT, 7(2):163–169, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=163.

Anonymous:1967:Pb

[242] Anonymous. Problems. BIT, 7(2):170– 173, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=2&spage=170.

Anonymous:1967:NBJb

[243] Anonymous. New books and journals. BIT, 7(2):174, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=7&
issue=2&spage=174.

Birtwistle:1967:GBM

[244] G. M. Birtwistle and D. J. Evans. On the generalisation of Bairstow’s method. BIT, 7(3):175–190, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=175.

Hansen:1967:LSR

[245] Per Brinch Hansen. The logical structure of the RC 4000 computer. BIT, 7(3):191–199, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=191.

Hunter:1967:RMC

[246] D. B. Hunter. Romberg’s method for certain integrals involving a singularity. BIT, 7(3):200–205, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=200.

Seppala:1967:DEF

[247] Yrjö Seppälä. Definition of extraction files and their optimization by zeroone programming. BIT, 7(3):206–215, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=206.

Storoy:1967:ECS

[248] Sverre Storøy. Error control in the simplex-technique. BIT, 7(3):216–225, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=216.

Tewarson:1967:SSS

[249] R. P. Tewarson. Solution of a system of simultaneous linear equations with a sparse coefficient matrix by elimination methods. BIT, 7(3):226–239, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=226.

Svejgaard:1967:APC

[250] Bj. Svejgaard. Algol programming: Contribution no. 21: Zeros of polynomials. BIT, 7(3):240–246, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=240.

Anonymous:1967:Pc

[251] Anonymous. Problems. BIT, 7(3): 247–249, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=247.

Hansen:1967:NBJ

[252] H. B. Hansen and Aage Melbye. New books and journals. BIT, 7(3):250–254, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=250.

Anonymous:1967:Nb

[253] Anonymous. Notes. BIT, 7(3):255–256, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=255.

Anonymous:1967:CAP

[254] Anonymous. Convergence acceleration from the point of view of linear programming. BIT, 7(3):256, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=3&spage=256.

Bjorck:1967:IRLa

[255] Åke Björck. Iterative refinement of linear least squares solutions I. BIT, 7(4):257–278, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=257.

Hansen:1967:RRT

[256] Per Brinch Hansen. The RC 4000 real-time control system at Pulawy. BIT, 7(4):279–288, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=279.

Johansen:1967:NDP

[257] Peter Johansen. Non-deterministic programming. BIT, 7(4):289–304, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=289.

Ruhe:1967:QCJ

[258] Axel Ruhe. On the quadratic convergence of the Jacobi method for normal matrices. BIT, 7(4):305–313, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=305.

Strom:1967:SEB

[259] Torsten Ström. Strict error bounds in Romberg quadrature. BIT, 7(4):

314–321, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=314.

Bjorck:1967:API

[260] Åke Björck and Gene H. Golub. Algol programming: Iterative refinement of linear least squares solutions by Householder transformation. BIT, 7(4):322–337, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=322.

Anonymous:1967:Pe

[261] Anonymous. Problems. BIT, 7(4): 338–340, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=338.

Anonymous:1967:NBJd

[262] Anonymous. New books and journals. BIT, 7(4):341–344, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=7&issue=4&spage=341.

Anonymous:1967:Nc

[263] Anonymous. Notes. BIT, 7(4):345, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=7&issue=4&spage=345.

Abramson:1968:POS

[264] H. D. Abramson. Partial ordering of the syntax elements of a language. BIT, 8(1):1–7, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=8&
                                        issue=1&spage=1.

Bjorck:1968:IRL

[265] Åke Björck. Iterative refinement of linear least squares solutions II. BIT, 8(1): 8–30, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=1&spage=8.

Jarratt:1968:NMD

[266] P. Jarratt. A numerical method for determining points of inflexion. BIT, 8(1):31–35, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=8&
                                        issue=1&spage=31.

Naur:1968:FGA

[267] Peter Naur. Features of the GIER ALGOL 4 system. BIT, 8(1):36– 42, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=1&spage=36.

ONeill:1968:SMN

[268] John C. O’Neill and George D. Byrne. A starting method for the numerical solution of Volterra’s integral equation of the second kind. BIT, 8(1): 43–47, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=1&spage=43.

Paananen:1968:TQT

[269] Eero Paananen. Two queues in tandem. BIT, 8(1):48–52, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=8&
issue=1&spage=48.

Tewarson:1968:SLE

[270] R. P. Tewarson. Solution of linear equations with coefficient matrix in band form. BIT, 8(1):53– 58, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=1&spage=53.

Anonymous:1968:Pa

[271] Anonymous. Problems. BIT, 8(1):59– 61, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=1&spage=59.

Anonymous:1968:NBJa

[272] Anonymous. New books and journals. BIT, 8(1):62–66, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=8&
issue=1&spage=62.

Anonymous:1968:NNN

[273] Anonymous. Notes: NordDATA-68; NEUCC Summer School 1968; 4th Yugoslav Symposium on Information Processing. BIT, 8(1):67– 68, March 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=1&spage=67.

Alanko:1968:SPL

[274] T. Alanko, S. Mustonen, and M. Tienari. A statistical programming language SURVO 66. BIT, 8(2):69– 85, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=2&spage=69.

Falkenberg:1968:MIU

[275] J. C. Falkenberg. A method for integration of unstable systems of ordinary differential equations subject to twopoint boundary conditions. BIT, 8(2):

86–103, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=2&spage=86.

Foxley:1968:PPM

[276] Eric Foxley. Programming problems in mathematical logic. BIT, 8(2):104– 121, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=2&spage=104.

Jonsson:1968:LPD

[277] Sven Ingvar Jönsson. On-line program debugging. BIT, 8(2):122– 127, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=2&spage=122.

Mayoh:1968:SGC

[278] B. H. Mayoh. The second Goldbach conjecture revisited. BIT, 8(2):128– 133, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=2&spage=128.

Day:1968:NSV

[279] J. T. Day. On the numerical solution of Volterra integral equations. BIT, 8(2): 134–137, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=2&spage=134.

Anonymous:1968:Pb

[280] Anonymous. Problems. BIT, 8(2):138– 141, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=2&spage=138.

Anonymous:1968:NBJb

[281] Anonymous. New books and journals. BIT, 8(2):142–146, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=8&
                                        issue=2&spage=142.

Anonymous:1968:E

[282] Anonymous. Erratum. BIT, 8(2): 145, June 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=2&spage=145.

Bjork:1968:DTA

[283] Harry Björk. Decision tables in ALGOL 60. BIT, 8(3):147–153, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=3&spage=147.

Bramble:1968:RCS

[284] J. H. Bramble, R. B. Kellogg, and V. Thomée. On the rate of convergence of some difference schemes for second order elliptic equations. BIT, 8(3):154–173, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=3&spage=154.

Dijkstra:1968:CAP

[285] E. W. Dijkstra. A constructive approach to the problem of program correctness. BIT, 8(3):174–186, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=3&spage=174.

Froberg:1968:PZF

[286] Carl-Erik Fröberg. On the prime zeta function. BIT, 8(3):187–202, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=3&spage=187.

Palme:1968:CBS

[287] Jacob Palme. A comparison between Simula and Fortran. BIT, 8(3): 203–209, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=3&spage=203.

Ruhe:1968:QCG

[288] Axel Ruhe. On the quadratic convergence of a generalization of the Jacobi method to arbitrary matrices. BIT, 8(3):210–231, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=3&spage=210.

Tewarson:1968:CSI

[289] R. P. Tewarson. On the Chebyshev solution of inconsistent linear equations. BIT, 8(3):232–242, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=3&spage=232.

Isacson:1968:AEC

[290] Tore Isacson. Asymptotic expansion of Coulomb wave functions on the transition line. BIT, 8(3):243–245, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=3&spage=243.

Anonymous:1968:Pc

[291] Anonymous. Problems. BIT, 8(3): 246–249, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=3&spage=246.

Anonymous:1968:NBJc

[292] Anonymous. New books and journals. BIT, 8(3):250–251, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=3&spage=250.

Anonymous:1968:Nb

[293] Anonymous. Notes. BIT, 8(3):251–252, September 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=3&spage=251.

Anonymous:1968:NFC

[294] Anonymous. Notes: FILE-68: Conference on the Numerical Solution of Differential Equations. BIT, 8(3):252, 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Abramson:1968:AMS

[295] H. D. Abramson. The applicability matrix of a syntax-directed parsing processor. BIT, 8(4):253–261, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=253.

Domolki:1968:UCS

[296] Bálint Dömölki. A universal compiler system based on production rules. BIT, 8(4):262–275, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=262.

Ehle:1968:HOS

[297] Byron L. Ehle. High order A-stable methods for the numerical solution of systems of D.E.’s. BIT, 8(4):276–278, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=276.

Einarsson:1968:NCF

[298] Bo Einarsson. Numerical calculation of Fourier integrals with cubic splines. BIT, 8(4):279–286, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=279.

Greenspan:1968:NSC

[299] Donald Greenspan. Numerical solution of a class of nonsteady cavity flow problems. BIT, 8(4):287–294, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=287.

Jarvelainen:1968:TSF

[300] Martti Järveläinen and Harry V. Nordén. A theoretical study of filter cake washing. BIT, 8(4):295–309, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=295.

Seppala:1968:TSS

[301] Yrjö Seppälä. Translating statements of symbolic logic into zero-one linear programming. BIT, 8(4):310–327, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=310.

Wittmeyer-Koch:1968:MDC

[302] L. Wittmeyer-Koch. A method of descent for Chebyshev approximation. BIT, 8(4):328–342, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=328.

Anonymous:1968:Pd

[303] Anonymous. Problems. BIT, 8(4): 343–346, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=8&issue=4&spage=343.

Anonymous:1968:NBJd

[304] Anonymous. New books and journals. BIT, 8(4):347–348, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=4&spage=347.

Anonymous:1968:Nc

[305] Anonymous. Notes. BIT, 8(4):349–350, December 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=8&issue=4&spage=349.

Hedstrom:1969:RCP

[306] G. W. Hedström. The rate of convergence of parabolic difference schemes with constant coefficients. BIT, 9(1): 1–17, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=1&spage=1.

Haavie:1969:DEE

[307] T. Håvie. Derivation of explicit expressions for the error terms in the ordinary and the modified Romberg algorithms. BIT, 9(1):18– 29, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=1&spage=18.

Larkin:1969:ENN

[308] F. M. Larkin. Estimation of a nonnegative function. BIT, 9(1):30– 52, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=1&spage=30.

Ruhe:1969:NMA

[309] Axel Ruhe. The norm of a matrix after a similarity transformation. BIT, 9(1): 53–58, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=1&spage=53.

Wood:1969:PHA

[310] D. Wood. A proof of Hamblin’s algorithm for translation of arithmetic expressions from infix to postfix form. BIT, 9(1):59–68, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=1&spage=59.

Woodford:1969:SCI

[311] C. H. Woodford. Smooth curve interpolation. BIT, 9(1):69–77, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=1&spage=69.

Beckman:1969:MAC

[312] Anders Beckman, Bengt Fornberg, and Arne Tengvald. A method for acceleration of the convergence of infinite series. BIT, 9(1):78– 80, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=1&spage=78.

Day:1969:NSC

[313] J. T. Day. Numerical solution of the convolution integral equation. BIT, 9(1):81–82, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=9&
                                        issue=1&spage=81.

Overholt:1969:APP

[314] K. J. Overholt. Algol programming: Partition functions (modulo d). BIT, 9(1):83–86, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=9&
                                        issue=1&spage=83.

Anonymous:1969:Pa

[315] Anonymous. Problems. BIT, 9(1):87– 90, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=1&spage=87.

Froberg:1969:NBJ

[316] Carl-Erik Fröberg and Peter Naur. New books and journals. BIT, 9(1): 91–93, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=1&spage=91.

Anonymous:1969:Na

[317] Anonymous. Notes. BIT, 9(1):94– 96, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=1&spage=94.

Baer:1969:APO

[318] R. M. Baer and O. Østerby. Algorithms over partially ordered sets. BIT, 9(2): 97–118, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=2&spage=97.

Jarratt:1969:SEF

[319] P. Jarratt. Some efficient fourth order multipoint methods for solving equations. BIT, 9(2):119–124, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=2&spage=119.

Karst:1969:URP

[320] Edgar Karst. Unique representation of primitive factors of 2^{n} − 1. n odd, in certain quadratic forms. BIT, 9(2): 125–132, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=2&spage=125.

Krarup:1969:BBA

[321] Jakob Krarup. A branch-bound algorithm for a wiring problem. BIT, 9(2): 133–156, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=2&spage=133.

Nordbeck:1969:CCR

[322] Stig Nordbeck and Bengt Rystedt. Computer cartography — range map. BIT, 9(2):157–166, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=2&spage=157.

Tewarson:1969:SCS

[323] R. P. Tewarson and B. Ramnath. Some comments on the solution of linear equations. BIT, 9(2):167– 173, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=2&spage=167.

Anonymous:1969:Pb

[324] Anonymous. Problems. BIT, 9(2):174– 177, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=2&spage=174.

Dahl:1969:NBJ

[325] Ole Johan Dahl and Helge Bratt. New books and journals. BIT, 9(2):178– 182, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=2&spage=178.

Anonymous:1969:Nb

[326] Anonymous. Notes. BIT, 9(2):183– 184, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=2&spage=183.

Einarsson:1969:E

[327] B. Einarsson. Erratum. BIT, 9(2):183– 184, June 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=2&spage=183.

Axelsson:1969:CSM

[328] Owe Axelsson. A class of Astable methods. BIT, 9(3):185–199, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=185.

Howell:1969:ASLa

[329] Jo Ann Howell and Robert T. Gregory. An algorithm for solving linear algebraic equations using residue arithmetic I. BIT, 9(3):200–224, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=200.

Kurki-Suonio:1969:NTL

[330] R. Kurki-Suonio. Notes on topdown languages. BIT, 9(3):225–238, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=225.

Lyche:1969:NCA

[331] Tom Lyche. A note on correctors with an arbitrary number of nonstep points. BIT, 9(3):239–249, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=239.

Naur:1969:PAC

[332] Peter Naur. Programming by action clusters. BIT, 9(3):250–258, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=250.

Norsett:1969:CSL

[333] Syvert P. Nørsett. A criterion for A(α)-stability of linear multistep methods. BIT, 9(3):259–263, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=259.

Riesel:1969:CPF

[334] Hans Riesel. Common prime factors of the numbers A_{n} = a^{2^{n}} + 1. BIT, 9(3):264–269, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=264.

Wodon:1969:DSS

[335] P. L. Wodon. Data structure and storage allocation. BIT, 9(3):270–282, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=270.

Hellgren:1969:APC

[336] Bo Hellgren. Algol programming: Contribution no. 24: Algorithms for the interpolation of a function defined by a character string. BIT, 9(3):283–293, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=3&spage=283. See

remark [517].

Anonymous:1969:Pc

[337] Anonymous. Problems. BIT, 9(3): 294–296, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=3&spage=294.

Anonymous:1969:NBJc

[338] Anonymous. New books and journals. BIT, 9(3):297–298, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=3&spage=297.

Anonymous:1969:Nc

[339] Anonymous. Notes. BIT, 9(3):298–300, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=3&spage=298.

Agrawala:1969:WDL

[340] Vishnu K. Agrawala and Johan G. Belinfante. Weight diagrams for Lie group representations: A computer implementation of Freudenthal’s algorithm in ALGOL and FOR-TRAN. BIT, 9(4):301–314, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=301.

Hadjidimos:1969:NET

[341] Apostolos Hadjidimos. A new explicit three-level difference scheme for the solution of the heat flow equation. BIT, 9(4):315–323, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=4&spage=315.

Howell:1969:ASLb

[342] Jo Ann Howell and Robert T. Gregory. An algorithm for solving linear algebraic equations using residue arithmetic II. BIT, 9(4):324–337, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=4&spage=324.

Haavie:1969:MCC

[343] T. Håvie. On a modification of the Clenshaw–Curtis quadrature formula. BIT, 9(4):338–350, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=4&spage=338.

Piessens:1969:NQF

[344] Robert Piessens. New quadrature formulas for the numerical inversion of the Laplace transform. BIT, 9(4):351–361, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=351.

Wirth:1969:BCC

[345] Niklaus Wirth. A basic course on compiler principles. BIT, 9(4):362–386, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=362.

Wood:1969:NTD

[346] D. Wood. A note on top-down deterministic languages. BIT, 9(4):387–399, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=387.

Anonymous:1969:Pd

[347] Anonymous. Problems. BIT, 9(4): 400–403, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=400.

Hornstrom:1969:NBJ

[348] Lars Hörnström and Karl Erik Karlsson. New books and journals. BIT, 9(4):404–407, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=404.

Anonymous:1969:A

[349] Anonymous. Acknowledgements. BIT, 9(4):407, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=9&issue=4&spage=407.

Abramson:1970:NLR

[350] H. D. Abramson. A note on leftrecursive rules and the partitioning of a recognition matrix for syntaxdirected translation. BIT, 10(1):1– 5, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=1&spage=1.

Abramson:1970:POS

[351] H. D. Abramson. A partial ordering of some of the elements of Algol. BIT, 10(1):6–9, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=10&
issue=1&spage=6.

Campbell:1970:NSN

[352] G. M. Campbell and J. T. Day. The numerical solution of nonlinear Volterra integral equations. BIT, 10(1):10– 19, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=1&spage=10.

Gear:1970:RAI

[353] C. W. Gear. Rational approximations by implicit Runge–Kutta schemes. BIT, 10(1):20–22, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=10&
                                        issue=1&spage=20.

Howell:1970:SLE

[354] Jo Ann Howell and Robert T. Gregory. Solving linear equations using residue arithmetic — algorithm II. BIT, 10(1): 23–37, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=1&spage=23.

Kolbig:1970:NGP

[355] K. S. Kölbig, J. A. Mignaco, and E. Remiddi. On Nielsen’s generalized polylogarithms and their numerical calculation. BIT, 10(1):38– 73, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=1&spage=38. See

erratum [394].

Mossige:1970:GPL

[356] Svein Mossige. Generation of permutations in lexicographical order. BIT, 10(1):74–75, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=10&
issue=1&spage=74.

Purdom:1970:TCA

[357] Paul Purdom. A transitive closure algorithm. BIT, 10(1):76–94, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=1&spage=76.

Seppala:1970:CFA

[358] Yrjö Seppälä. Criterion function for accessibility, uncertainty problems in regional land-use planning. BIT, 10(1):95–105, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=10&
issue=1&spage=95.

Hunter:1970:APA

[359] D. B. Hunter. ALGOL programming: Analysis of the outer product of symmetric group representations. BIT, 10(1):106–114, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=1&spage=106.

Anonymous:1970:Pa

[360] Anonymous. Problems. BIT, 10 (1):115–117, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=1&spage=115.

Anonymous:1970:NBJa

[361] Anonymous. New books and journals. BIT, 10(1):117, 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1970:Na

[362] Anonymous. Notes. BIT, 10(1): 117–123, 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1970:NBJb

[363] Anonymous. New books and journals. BIT, 10(1):118, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=10&
                                        issue=1&spage=118.

Anonymous:1970:Nb

[364] Anonymous. Notes. BIT, 10(1):118– 123, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=1&spage=118.

Bjorner:1970:FMS

[365] Dines Bjørner. A flow-mode, selfsteering, cellular multiplier-summation processor. BIT, 10(2):125–144, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=10&issue=2&spage=125.

Chawla:1970:HSE

[366] M. M. Chawla. Hilbert spaces for estimating errors of quadratures for analytic functions. BIT, 10(2):145– 155, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=2&spage=145.

Danielsson:1970:TDE

[367] P. E. Danielsson. Three-dimensional electronic display. BIT, 10(2):156– 167, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=2&spage=156.

London:1970:PPC

[368] Ralph L. London. Proving programs correct: Some techniques and examples. BIT, 10(2):168– 182, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=2&spage=168.

Paige:1970:PUS

[369] C. C. Paige. Practical use of the symmetric Lánczos process with reorthogonalization. BIT, 10(2):183– 195, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=2&spage=183.

Ruhe:1970:AND

[370] Axel Ruhe. An algorithm for numerical determination of the structure of a general matrix. BIT, 10(2):196– 216, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=2&spage=196.

Wright:1970:SRB

[371] K. Wright. Some relationships between implicit Runge–Kutta, collocation and Lánczos τ methods, and their stability properties. BIT, 10(2):217– 227, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=2&spage=217.

Anonymous:1970:Pb

[372] Anonymous. Problems. BIT, 10(2): 228–231, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=2&spage=228.

Apelkrans:1970:NBJ

[373] Matts Apelkrans. New books and journals. BIT, 10(2):232–234, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=10&issue=2&spage=232.

Anonymous:1970:Nc

[374] Anonymous. Notes. BIT, 10(2): 235, June 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=2&spage=235.

Anonymous:1970:Cb

[375] Anonymous. Contents. BIT, 10 (2):236, 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bohman:1970:MCD

[376] Harald Bohman. A method to calculate the distribution function when the characteristic function is known. BIT, 10(3):237–242, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=237.

Chawla:1970:ANN

[377] M. M. Chawla. Approximation by non-negative algebraic polynomials. BIT, 10(3):243–248, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=243.

Clay:1970:RNR

[378] James R. Clay. Research in nearring theory using a digital computer. BIT, 10(3):249–265, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=249.

Hakberg:1970:UMS

[379] Bengt Hakberg. Uniformly maximumnorm stable difference schemes. BIT, 10(3):266–276, September 1970. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=10&issue=3&spage=
266.

Haavie:1970:SAN

[380] T. Håvie. Some algorithms for numerical quadrature using the derivatives of the integrand in the integration interval. BIT, 10(3):277–294, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=277.

McVitie:1970:SMA

[381] D. G. McVitie and L. B. Wilson. Stable marriage assignment for unequal sets. BIT, 10(3):295–309, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=295.

Murphree:1970:TGI

[382] E. L. Murphree, Jr. and S. J. Fenves. A technique for generating interpretive translators for problem-oriented languages. BIT, 10(3):310–323, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=310.

Phillips:1970:ACD

[383] G. M. Phillips and P. J. Taylor. Approximation of convex data. BIT, 10(3):324–332, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=324.

Riesel:1970:PFA

[384] Hans Riesel. Primes forming arithmetic series and clusters of large primes. BIT, 10(3):333–342, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=333.

Ruhe:1970:PBM

[385] Axel Ruhe. Perturbation bounds for means of eigenvalues and invariant subspaces. BIT, 10(3):343–354, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=3&spage=343.

Tienari:1970:SMR

[386] Martti Tienari. A statistical model of roundoff error for varying length floating-point arithmetic. BIT, 10(3): 355–365, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=3&spage=355.

Whiteman:1970:CFD

[387] J. R. Whiteman and J. C. Webb. Convergence of finite-difference techniques for a harmonic mixed boundary value problem. BIT, 10(3):366–374, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=3&spage=366.

Hadjidimos:1970:NDS

[388] Apostolos Hadjidimos. A note on a difference scheme for the solution of the heat equation. BIT, 10(3):375, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=3&spage=375.

Businger:1970:APU

[389] P. Businger. Algol programming: Updating a singular value decomposition. BIT, 10(3):376–385, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=376.

Parlett:1970:SSL

[390] B. Parlett and J. K. Reid. On the solution of a system of linear equations whose matrix is symmetric but not definite. BIT, 10(3):386–397, 1970. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1970:Pc

[391] Anonymous. Problems. BIT, 10(3): 398–399, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=398.

Anonymous:1970:NBJd

[392] Anonymous. New books and journals. BIT, 10(3):400–401, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=400.

Anonymous:1970:Nd

[393] Anonymous. Notes. BIT, 10(3): 402–403, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=3&spage=402.

Kolbig:1970:ENG

[394] K. S. Kölbig, J. A. Mignaco, and E. Remiddi. Erratum: “On Nielsen’s Generalized Polylogarithms and their Numerical Calculation” [BIT 10(1), 1970, p. 59]. BIT, 10(3):403, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=3&spage=403. See

[355].

Anonymous:1970:Cc

[395] Anonymous. Contents. BIT, 10 (3):404, 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Baecker:1970:IAH

[396] H. D. Baecker. Implementing the AL-GOL 68 heap. BIT, 10(4):405–414, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=4&spage=405.

Bjorner:1970:FM

[397] Dines Bjørner. Flowchart machines. BIT, 10(4):415–442, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=4&spage=415.

Coffman:1970:FSN

[398] E. G. Coffman and J. Bruno. On file structuring for non-uniform access frequencies. BIT, 10(4):443–456, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=4&spage=443.

Espinosa-Maldonado:1970:SLI

[399] Ruben Espinosa-Maldonado and George D. Byrne. Solution of linear integral equations by Gregory’s method. BIT, 10 (4):457–464, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=4&spage=457.

Hadjidimos:1970:EAD

[400] Apostolos Hadjidimos. Extrapolated alternating direction implicit iterative methods. BIT, 10(4):465–475, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=4&spage=465.

Piessens:1970:NEC

[401] Robert Piessens. Numerical evaluation of Cauchy principal values of integrals. BIT, 10(4):476–480, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=4&spage=476.

Stoneham:1970:NCM

[402] R. G. Stoneham. On a new class of multiplicative pseudo-random number generators. BIT, 10(4):481–500, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=4&spage=481.

Woodford:1970:ADS

[403] C. H. Woodford. An algorithm for data smoothing using spline functions. BIT, 10(4):501–510, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=4&spage=501.

Day:1970:NSI

[404] J. T. Day. On the numerical solution of integro-differential equations. BIT, 10(4):511–514, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=4&spage=511.

Anonymous:1970:Pd

[405] Anonymous. Problems. BIT, 10(4): 515–517, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=10&issue=4&spage=515.

Persson:1970:NBJ

[406] Inge Persson and P. E. Danielsson. New books and journals. BIT, 10(4): 518–520, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=4&spage=518.

Anonymous:1970:Ne

[407] Anonymous. Notes. BIT, 10(4): 521–523, December 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=10&issue=4&spage=521.

Anonymous:1970:A

[408] Anonymous. Acknowledgements. BIT, 10(4):524, 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1970:I

[409] Anonymous. Index. BIT, 10(4): 525–526, 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Agrawala:1971:ACI

[410] Vishnu K. Agrawala and Johan G. Belinfante. An algorithm for computing su(n) invariants. BIT, 11(1):1– 15, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=1.

Bjork:1971:BUT

[411] Harry Björk. A bi-unique transformation into integers of identifiers and other variable length items. BIT, 11(1): 16–20, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=16.

Kolbig:1971:VLT

[412] K. S. Kölbig. On the value of a logarithmic-trigonometric integral. BIT, 11(1):21–28, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=11&
issue=1&spage=21.

Lindberg:1971:SET

[413] Bengt Lindberg. On smoothing and extrapolation for the trapezoidal rule. BIT, 11(1):29–52, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=29.

Parter:1971:NTP

[414] Seymour V. Parter and Víctor Pereyra. Nonlinear two-point boundary value problems with multiple solutions. BIT, 11(1):53–83, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=53.

Schneider:1971:NRN

[415] Victor Schneider. On the number of registers needed to evaluate arithmetic expressions. BIT, 11(1):84– 93, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=84. See

erratum [465].

Steiglitz:1971:NDF

[416] Kenneth Steiglitz and John Bruno. A new derivation of Frisch’s algorithm for calculating vertex-pair connectivity. BIT, 11(1):94–106, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=94.

Sundblad:1971:AFT

[417] Y. Sundblad. The Ackermann function. A theoretical, computational, and formula manipulative study. BIT, 11(1):107–119, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=1&spage=107.

Campbell:1971:BBM

[418] G. M. Campbell and J. T. Day. A block by block method for the numerical solution of Volterra integral equations. BIT, 11(1):120– 124, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=1&spage=120.

Anonymous:1971:Pa

[419] Anonymous. Problems. BIT, 11 (1):125–126, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=1&spage=125.

Anonymous:1971:NBJa

[420] Anonymous. New books and journals. BIT, 11(1):127–128, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=11&
                                        issue=1&spage=127.

Anonymous:1971:Na

[421] Anonymous. Notes. BIT, 11(1):129– 131, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=1&spage=129.

Anonymous:1971:Ca

[422] Anonymous. Contents. BIT, 11 (1):132, 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bohman:1971:MDR

[423] H. Bohman. On the maximum deviation in random walks. BIT, 11(2): 133–138, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=133.

Boland:1971:PTQ

[424] W. Robert Boland and C. S. Duris. Product type quadrature formulas. BIT, 11(2):139– 158, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=139.

Chawla:1971:CNC

[425] M. M. Chawla. Convergence of Newton–Cotes quadratures for analytic functions. BIT, 11(2):159– 167, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=159.

Powell:1971:MRF

[426] M. Powell and P. Jarratt. A modified regula falsi method for computing the root of an equation. BIT, 11(2):168– 174, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=168.

Hunter:1971:EIP

[427] D. B. Hunter. The evaluation of integrals of periodic analytic functions. BIT, 11(2):175–180, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=2&spage=175.

Phillips:1971:EEV

[428] G. M. Phillips. An error estimate for Volterra integral equations. BIT, 11(2): 181–186, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=2&spage=181.

Purdom:1971:SIT

[429] P. W. Purdom, S. M. Stigler, and Tat-Ong Cheam. Statistical investigation of three storage allocation algorithms. BIT, 11(2):187– 195, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=2&spage=187.

Strom:1971:IFD

[430] Torsten Ström. An inclusion formula for derivatives. BIT, 11(2):196– 198, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=2&spage=196.

Strom:1971:SEM

[431] Torsten Ström. Strict estimation of the maximum of a function of one variable. BIT, 11(2):199–211, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=199.

Terzi:1971:CES

[432] P. G. Terzi. On a conjecture by Erdős–Straus. BIT, 11(2):212– 216, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=212.

Wunderlich:1971:UBB

[433] M. C. Wunderlich. The use of bit and byte manipulation in computing summation sequences. BIT, 11(2):217– 224, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=217.

Anonymous:1971:Pb

[434] Anonymous. Problems. BIT, 11(2): 225–227, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=2&spage=225.

Anonymous:1971:NBJb

[435] Anonymous. New books and journals. BIT, 11(2):228–230, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=11&
                                        issue=2&spage=228.

Anonymous:1971:Nb

[436] Anonymous. Notes. BIT, 11(2): 231, June 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=2&spage=231.

Anonymous:1971:Cb

[437] Anonymous. Contents. BIT, 11 (2):232, 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Aasen:1971:RSM

[438] Jan Ole Aasen. On the reduction of a symmetric matrix to tridiagonal form. BIT, 11(3):233–242, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=3&spage=233.

Abramson:1971:PCI

[439] H. D. Abramson. A program for computing identity-kernels from certain directed graphs. BIT, 11(3):243–253, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=243.

Amble:1971:CSP

[440] E. Amble. A Cosmo–Simula program for a simple air-pollution model. BIT, 11(3):254–261, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=254.

Businger:1971:NSD

[441] P. A. Businger. Numerically stable deflation of Hessenberg and symmetric tridiagonal matrices. BIT, 11(3): 262–270, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=262.

Caprani:1971:ILR

[442] Ole Caprani. Implementation of a low round-off summation method. BIT, 11 (3):271–275, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=271.

Christiansen:1971:NSI

[443] Søren Christiansen. Numerical solution of an integral equation with a logarithmic kernel. BIT, 11(3):276–287, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=3&spage=276.

Haavie:1971:UTA

[444] T. Håvie. On the use of trigonometric approximations in extrapolation methods. BIT, 11(3):288–298, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=3&spage=288.

Kambo:1971:EBC

[445] N. S. Kambo. Error bounds for the Clenshaw–Curtis quadrature formulas. BIT, 11(3):299–309, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=3&spage=299.

Krusemeyer:1971:DMD

[446] Mark I. Krusemeyer. Determining multiplicities of dominant weights in irreducible lie algebra representations using a computer. BIT, 11(3):310–316, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=3&spage=310.

Piessens:1971:NMI

[447] R. Piessens and F. Poleunis. A numerical method for the integration of oscillatory functions. BIT, 11(3):317–327, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=317.

Stewart:1971:HRW

[448] N. F. Stewart. A heuristic to reduce the wrapping effect in the numerical solution of z′ = f(t, x). BIT, 11(3): 328–337, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=328.

Anonymous:1971:Pc

[449] Anonymous. Problems. BIT, 11(3): 338–339, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=338.

Anonymous:1971:NBJc

[450] Anonymous. New books and journals. BIT, 11(3):340–341, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=340.

Anonymous:1971:Nc

[451] Anonymous. Notes. BIT, 11(3): 341–343, September 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=3&spage=341.

Anonymous:1971:Cc

[452] Anonymous. Contents. BIT, 11 (3):344, 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Abdelmalek:1971:REA

[453] Nabih N. Abdelmalek. Round-off error analysis for Gram–Schmidt method and solution of linear least squares problems. BIT, 11(4):345–367, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=345.

Chakravarti:1971:CSS

[454] P. C. Chakravarti and P. B. Worland. A class of self-starting methods for the numerical solution of y′′ = f(z, y). BIT, 11(4):368–383, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=368.

Chipman:1971:SRK

[455] F. H. Chipman. A-stable Runge– Kutta processes. BIT, 11(4):384–388, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=384.

Froberg:1971:SPQ

[456] Carl-Erik E. Fröberg. On sums and products of quadratic residues. BIT, 11 (4):389–398, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=389.

Greenspan:1971:CST

[457] Donald Greenspan. Computer simulation of transverse string vibrations. BIT, 11(4):399–408, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=399.

King:1971:FOF

[458] Richard F. King. A fifth-order family of modified Newton methods. BIT, 11(4):409–412, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=409.

Linz:1971:PIM

[459] Peter Linz. Product integration methods for Volterra integral equations of the first kind. BIT, 11(4):413–421, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=413.

Miller:1971:NIG

[460] Webb Miller. A note on the instability of Gaussian elimination. BIT, 11 (4):422–424, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=4&spage=422.

Wang:1971:CSB

[461] A. Wang and O. J. Dahl. Coroutine sequencing in a block structured environment. BIT, 11(4):425–449, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=4&spage=425.

Day:1971:NSG

[462] J. T. Day. On the numerical solution of the Goursat problem. BIT, 11(4):450–454, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=4&spage=450.

Anonymous:1971:Pd

[463] Anonymous. Problems. BIT, 11(4): 455–457, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=11&issue=4&spage=455.

Hugoson:1971:NBJ

[464] Mats-Åke Hugoson. New books and journals. BIT, 11(4):458–459, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=458.

Schneider:1971:ENR

[465] V. Schneider. Erratum: “On the Number of Registers Needed to Evaluate Arithmetic Expressions” [BIT 11(1), 1971, pp. 87, 92]. BIT, 11(4):460, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=460. See

[415].

Anonymous:1971:A

[466] Anonymous. Acknowledgements. BIT, 11(4):461, December 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=11&issue=4&spage=461.

Anonymous:1971:I

[467] Anonymous. Index. BIT, 11(4): 462–463, 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1971:Cd

[468] Anonymous. Contents. BIT, 11 (4):464, 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Axelsson:1972:NCS

[469] Owe Axelsson. A note on a class of strongly A-stable methods. BIT, 12(1): 1–4, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=1.

Boland:1972:NSF

[470] W. Robert Boland. The numerical solution of Fredholm integral equations using product type quadrature formulas. BIT, 12(1):5–16, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=5.

Cryer:1972:IHO

[471] Colin W. Cryer. On the instability of high order backward-difference multistep methods. BIT, 12(1):17– 25, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=17.

Dahl:1972:NSS

[472] Olav Dahl. Numerical solution of some degenerate parabolic Cauchy problems. BIT, 12(1):26–37, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=12&
issue=1&spage=26.

Henderson:1972:ESP

[473] P. Henderson and R. Snowdon. An experiment in structured programming. BIT, 12(1):38–53, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=12&
issue=1&spage=38.

Hoskins:1972:ECI

[474] W. D. Hoskins and P. J. Ponzo. Explicit calculation of interpolating cubic splines on equi-distant knots. BIT, 12(1):54–62, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=12&
issue=1&spage=54.

Kronsjo:1972:DNI

[475] Lydia Kronsjö and Germund Dahlquist. On the design of nested iterations for elliptic difference equations. BIT, 12(1):63–71, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=12&
issue=1&spage=63.

Lukka:1972:ENC

[476] Anita Lukka. On the estimation of Numerus Clausus. BIT, 12(1):72– 80, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=1&spage=72.

Mitchell:1972:CBF

[477] A. R. Mitchell and G. M. Phillips. Construction of basis functions in the finite element method. BIT, 12(1):81– 89, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=1&spage=81.

Rehnqvist:1972:ICS

[478] Lars Rehnqvist. Inversion of certain symmetric band matrices. BIT, 12(1): 90–98, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=1&spage=90.

Wedin:1972:PBC

[479] P.-Å. Wedin. Perturbation bounds in connection with singular value decomposition. BIT, 12(1):99–111, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=1&spage=99.

Espelid:1972:BSM

[480] T. Espelid. On the behavior of the secant method near a multiple root. BIT, 12(1):112–115, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=112.

Hosken:1972:ARS

[481] W. H. Hosken. An asymmetric regular set. BIT, 12(1):115–117, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=115.

Jonsson:1972:CPM

[482] Ingemar Jönsson. On certain primes of Mersenne-type. BIT, 12(1):117– 118, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=117.

Anonymous:1972:Pa

[483] Anonymous. Problems. BIT, 12 (1):119–121, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=1&spage=119.

Eriksson:1972:NBJ

[484] Göran Eriksson. New books and journals. BIT, 12(1):121–123, March 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=

                                        article&issn=0006-3835&volume=12&
                                        issue=1&spage=121. Includes review

of Introduction to FORTRAN, by S. Kallin.

Anonymous:1972:Ca

[485] Anonymous. Contents. BIT, 12 (1):124, 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Arnborg:1972:SAV

[486] Stefan Arnborg. Storage administration in a virtual memory Simula system. BIT, 12(2):125–141, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=125.

Bjurel:1972:MLM

[487] Gunnar Bjurel. Modified linear multistep methods for a class of stiff ordinary differential equations. BIT, 12(2): 142–160, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=142.

Corneil:1972:ADA

[488] D. G. Corneil. An algorithm for determining the automorphism partitioning of an undirected graph. BIT, 12(2): 161–171, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=161.

George:1972:CLB

[489] John H. George and Robert W. Gunderson. Conditioning of linear boundary value problems. BIT, 12(2):172– 181, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=2&spage=172.

Hilden:1972:TRP

[490] J. Hilden. Testing the relative positions of several points on a circle. BIT, 12(2): 182–187, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=2&spage=182.

Laidlaw:1972:IMS

[491] B. H. Laidlaw and G. M. Phillips. Iterative methods for solving Fredholm integral equations. BIT, 12(2):188– 193, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=2&spage=188.

Lyness:1972:AGR

[492] J. N. Lyness. An algorithm for Gauss– Romberg integration. BIT, 12(2):194– 203, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=2&spage=194.

Mavaddat:1972:UAE

[493] F. Mavaddat. A unified approach to evaluation of expressions. BIT, 12(2): 204–212, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=204.

Miller:1972:IAM

[494] Webb Miller. On an interval-arithmetic matrix method. BIT, 12(2):213– 219, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=213.

Rheinboldt:1972:PLG

[495] Werner C. Rheinboldt, Victor R. Basilli, and Charles K. Mesztenyi. On a programming language for graph algorithms. BIT, 12(2):220– 241, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=220.

Scraton:1972:MMR

[496] R. E. Scraton. A modification of Miller’s recurrence algorithm. BIT, 12(2):242–251, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=242.

Watts:1972:SBI

[497] H. A. Watts and L. F. Shampine. Astable block implicit one-step methods. BIT, 12(2):252–266, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=12&
issue=2&spage=252.

McKay:1972:APCa

[498] A. C. McKay. Algol programming: Certification of contribution no. 10: LPMODSIMPLEX [BIT 4(4), 1964, pp. 189–196]. BIT, 12(2):267– 268, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=2&spage=267. See

[130, 188, 499].

McKay:1972:APCb

[499] A. C. McKay. Algol programming: Certification of contribution no. 10: LPMODSIMPLEX [BIT 4(4), 1964, pp. 189–196] with upper bounds. BIT, 12(2):268–271, 1972. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [130, 188, 498].

Anonymous:1972:Pb

[500] Anonymous. Problems. BIT, 12(2): 272–273, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=2&spage=272.

Bergh:1972:NBJ

[501] Peter Bergh. New books and journals. BIT, 12(2):274–276, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=12&
                                        issue=2&spage=274.  Includes re-

view of Introduktion till Operativsystem (Swedish) [Introduction to Operating Systems], by J. Bubenko and T. Ohlin.

Anonymous:1972:Na

[502] Anonymous. Notes. BIT, 12(2): 277, June 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=2&spage=277.

Anonymous:1972:Cb

[503] Anonymous. Contents. BIT, 12 (2):278, 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bohman:1972:CFD

[504] Harald Bohman. From characteristic function to distribution function via Fourier analysis. BIT, 12(3):279–283, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=3&spage=279.

Chuba:1972:QCI

[505] William Chuba and Webb Miller. Quadratic convergence in interval arithmetic, Part I. BIT, 12(3):284–290, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=284.

Miller:1972:QCI

[506] Webb Miller. Quadratic convergence in interval arithmetic, Part II. BIT, 12(3):291–298, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=291.

Cohen:1972:ESF

[507] Jacques Cohen and Carl Zuckerman. Evalquote in simple Fortran: a tutorial on interpreting Lisp. BIT, 12(3):299–317, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=299.

Duijvestijn:1972:CPP

[508] A. J. W. Duijvestijn. Correctness proof of an in-place permutation. BIT, 12(3):318–324, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=318.

Garey:1972:PCM

[509] L. Garey. Predictor-corrector methods for nonlinear Volterra integral equations of the second kind. BIT, 12(3): 325–333, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=3&spage=325.

Hoare:1972:NS

[510] C. A. R. Hoare. A note on the For statement. BIT, 12(3):334–341, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=3&spage=334.

Klove:1972:RIS

[511] Torleiv Kløve. Representations of integers as sums of powers with increasing exponents. BIT, 12(3):342–346, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=3&spage=342.

Naur:1972:EPD

[512] Peter Naur. An experiment on program development. BIT, 12(3):347–365, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=3&spage=347.

Purdom:1972:SGT

[513] Paul Purdom. A sentence generator for testing parsers. BIT, 12(3):366–375, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=366. See

erratum [543].

Seppala:1972:CCP

[514] Yrjö Seppälä. A chance-constrained programming algorithm. BIT, 12(3): 376–399, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=376.

Streit:1972:EDS

[515] Roland F. Streit. The evaluation of double series. BIT, 12(3):400–408, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=400.

Tienari:1972:STP

[516] Martti Tienari. On some topological properties of numerical algorithms. BIT, 12(3):409–433, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=409.

Emery:1972:APR

[517] J. Emery and J. Vine. Algol programming: Remark on “Procedures for Simplifying Boolean Expressions” [BIT 6(3), 1966, pp. 260–269, by Bo Hellgren]. BIT, 12(3):434, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=434. See

[336].

Anonymous:1972:Pc

[518] Anonymous. Problems. BIT, 12(3): 435–437, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=435.

Dahlstrand:1972:NBJ

[519] I. Dahlstrand. New books and journals. BIT, 12(3):438–439, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=438. In-

cludes review of Programmeringsspråk (Swedish) [Programming languages], by Jacob Palme.

Anonymous:1972:E

[520] Anonymous. Errata. BIT, 12(3):440, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=440.

Anonymous:1972:Nc

[521] Anonymous. Notes. BIT, 12(3):440, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=3&spage=440.

Anonymous:1972:Cc

[522] Anonymous. Contents. BIT, 12 (3):442, 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Axelsson:1972:GSM

[523] O. Axelsson. A generalized SSOR method. BIT, 12(4):443–467, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=443.

Brunner:1972:CST

[524] H. Brunner. A class of A-stable twostep methods based on Schur polynomials. BIT, 12(4):468–474, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=468.

Burden:1972:DED

[525] R. L. Burden and G. W. Hedstrom. The distribution of the eigenvalues of the discrete Laplacian. BIT, 12(4):475–488, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=475.

Chawla:1972:OQA

[526] M. M. Chawla and B. L. Raina. Optimal quadratures for analytic functions. BIT, 12(4):489–502, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=489.

Powell:1972:PMC

[527] M. Powell and P. Jarratt. The ‘Pegasus’ method for computing the root of an equation. BIT, 12(4):503–508, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=503.

Hosken:1972:SPC

[528] W. H. Hosken. Some post canonical systems in one letter. BIT, 12 (4):509–515, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=509. See

erratum [562].

Haavie:1972:EDR

[529] Tore Håvie. Error derivation in Romberg integration. BIT, 12(4): 516–527, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=516.

Linz:1972:EEC

[530] Peter Linz. Error estimates for the computation of eigenvalues of selfadjoint operators. BIT, 12(4):528–533, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=528.

Palme:1972:PCH

[531] Jacob Palme. Part-compilation in highlevel languages. BIT, 12(4):534–542, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=534.

Ruhe:1972:MCG

[532] Axel Ruhe and Torbjörn Wiberg. The method of conjugate gradients used in inverse iteration. BIT, 12(4):543–554, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=543.

Thorelli:1972:MA

[533] Lars-Erik Thorelli. Marking algorithms. BIT, 12(4):555–568, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=555.

Wilson:1972:ASM

[534] L. B. Wilson. An analysis of the stable marriage assignment algorithm. BIT, 12(4):569–575, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=569.

Bohman:1972:NPL

[535] J. Bohman. On the number of primes less than a given limit. BIT, 12(4): 576–577, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=576.

Goldfine:1972:ANS

[536] A. Goldfine. An algorithm for the numerical solution of integro-differential equations. BIT, 12(4):578–580, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=578.

Jacobs:1972:NCP

[537] D. Jacobs and F. Lambert. On the numerical calculation of polylogarithms. BIT, 12(4):581–585, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=581.

Lawson:1972:EIR

[538] J. Douglas Lawson. On the exactness of implicit Runge–Kutta processes for particular integrals. BIT, 12(4): 586–588, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=586.

Anonymous:1972:Pd

[539] Anonymous. Problems. BIT, 12(4): 589–590, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=589.

Anonymous:1972:NBJd

[540] Anonymous. New books and journals. BIT, 12(4):591, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=12&issue=4&spage=591.

Anonymous:1972:Nd

[541] Anonymous. Notes. BIT, 12(4): 592–594, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=592.

Anonymous:1972:A

[542] Anonymous. Acknowledgements. BIT, 12(4):595, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=595.

Purdom:1972:ESG

[543] P. Purdom. Erratum: “A Sentence Generator for Testing Parsers” [BIT 12(3), 1972, p. 372]. BIT, 12(4):595, December 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=12&issue=4&spage=595. See

[513].

Anonymous:1972:I

[544] Anonymous. Index. BIT, 12(4): 596–597, 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1972:Cd

[545] Anonymous. Contents. BIT, 12 (4):598, 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

ElTom:1973:NSV

[546] M. E. A. El Tom. Numerical solution of Volterra integral equations by spline functions. BIT, 13(1):1– 7, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=1.

Farmer:1973:NPG

[547] M. R. Farmer and G. Loizou. A note on a paper by G. Pólya. BIT, 13(1): 8–15, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=8.

Haavie:1973:REI

[548] Tore Håvie. Remarks on an expansion for integrals of rapidly oscillating functions. BIT, 13(1):16– 29, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=16.

Kambo:1973:EBC

[549] N. S. Kambo. Error bounds for a Chebyshev quadrature scheme. BIT, 13(1):30–37, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=30.

Klimko:1973:ACI

[550] Eugene M. Klimko. An algorithm for calculating indices in Fàa di Bruno’s formula. BIT, 13(1):38– 49, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=38.

Lindstrom:1973:CER

[551] Gary Lindstrom. Control extension in a recursive language. BIT, 13(1):50– 70, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=50.

Madsen:1973:RFA

[552] Kaj Madsen. A root-finding algorithm based on Newton’s method. BIT, 13(1): 71–75, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=71.

Miller:1973:MQC

[553] Webb Miller. More on quadratic convergence in interval arithmetic. BIT, 13(1):76–83, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=76.

Overholt:1973:OBS

[554] K. J. Overholt. Optimal binary search methods. BIT, 13(1):84–91, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=84.

Overholt:1973:EFS

[555] K. J. Overholt. Efficiency of the Fibonacci search method. BIT, 13(1): 92–96, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=92.

Riesel:1973:SSR

[556] Hans Riesel. Some series related to infinite series given by Ramanujan. BIT, 13(1):97–113, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=97. See

remarks [648].

Storoy:1973:AFV

[557] S. Storøy. An algorithm for finding a vector in the intersection of open convex polyhedral cones. BIT, 13(1):114– 119, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=114.

Gurski:1973:NAK

[558] Aaron Gurski. A note on analysis of keys for use in hashing. BIT, 13(1):120– 122, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=120.

Anonymous:1973:Pa

[559] Anonymous. Problems. BIT, 13 (1):123–124, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=123.

Anonymous:1973:BRa

[560] Anonymous. Books received. BIT, 13(1):125, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=125.

Anonymous:1973:Na

[561] Anonymous. Notes. BIT, 13(1):126– 129, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=1&spage=126.

Hosken:1973:ESP

[562] W. H. Hosken. Erratum: “Some Post Canonical Systems in One Letter” [BIT 12(4), 1972, p. 511]. BIT, 13(1): 129, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=1&spage=129. See

[528].

Anonymous:1973:Ca

[563] Anonymous. Contents. BIT, 13 (1):130, 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Andria:1973:NSB

[564] George D. Andria, George D. Byrne, and David R. Hill. Natural spline block implicit methods. BIT, 13(2): 131–144, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=2&spage=131.

Chawla:1973:ORN

[565] M. M. Chawla and Veena Kaul. Optimal rules for numerical integration round the unit circle. BIT, 13(2):145– 152, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=2&spage=145.

Cryer:1973:NCH

[566] Colin W. Cryer. A new class of highly-stable methods: A_{0}-stable methods. BIT, 13(2):153–159, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=

                                        article&issn=0006-3835&volume=13&
                                        issue=2&spage=153.

Gupta:1973:NSS

[567] Murli M. Gupta. Numerical solution of a second biharmonic boundary value problem. BIT, 13(2):160– 164, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=160.

Lawson:1973:FAM

[568] Harold W. Lawson, Jr. and Bengt Malm. A flexible asynchronous microprocessor. BIT, 13(2):166– 176, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=165.

Phillips:1973:EFC

[569] G. M. Phillips. Explicit forms for certain Hermite approximations. BIT, 13(2):177–180, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=177.

Smeds:1973:CIJ

[570] Boris Smeds. A 3-colour ink jet plotter for computer graphics. BIT, 13(2): 181–195, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=13&issue=2&spage=181.

Strom:1973:PEE

[571] Torsten Ström. Practical error estimates for repeated Richardson extrapolation schemes. BIT, 13(2):196– 205, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=2&spage=196.

Thesen:1973:SCP

[572] Arne Thesen. Scheduling of computer programs for optimal machine utilization. BIT, 13(2):206– 216, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=2&spage=206.

Wedin:1973:PTP

[573] Per-Åke Wedin. Perturbation theory for pseudo-inverses. BIT, 13(2):217– 232, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=2&spage=217.

Vold:1973:OBD

[574] H. Vold and B. H. Sjøgren. Optimal backup of data bases: a statistical investigation. BIT, 13(2):233– 241, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=233.

Bohman:1973:SCR

[575] J. Bohman. Some computational results regarding the prime numbers below 3, 000, 000, 000. BIT, 13(2):242– 244, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=242. See

erratum [637].

Wood:1973:NTL

[576] Derick Wood. A note on table look-up. BIT, 13(2):245–246, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=245.

Anonymous:1973:Pb

[577] Anonymous. Problems. BIT, 13(2): 247–249, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=247.

Anonymous:1973:BRb

[578] Anonymous. Books received. BIT, 13 (2):250, June 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=2&spage=250.

Anonymous:1973:Cb

[579] Anonymous. Contents. BIT, 13 (2):251, 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anderson:1973:NHO

[580] Ned Anderson and Åke Björck. A new high order method of Regula Falsi type for computing a root of an equation. BIT, 13(3):253–264, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=253.

Andreassen:1973:SMA

[581] Dagfinn Andreassen. On k-step methods with almost constant coefficients. BIT, 13(3):265–271, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=265.

Bickart:1973:HOS

[582] Theodore A. Bickart and Zdenek Picel. High order stiffly stable composite multistep methods for numerical integration of stiff differential equations. BIT, 13(3):272–286, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=272.

Boland:1973:PPT

[583] W. Robert Boland. Properties of product-type quadrature formulas. BIT, 13(3):287–291, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=287.

Ekblom:1973:CLB

[584] Håkan Ekblom. Calculation of linear best L_{p}-approximations. BIT, 13(3): 292–300, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=292.

Germano:1973:FDC

[585] G. Germano and A. Maggiolo-Schettini. A flow diagram composition of Markov’s normal algorithms without concluding formulas. BIT, 13(3):301–312, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=301.

Knuth:1973:OMP

[586] Donald E. Knuth and Francis R. Stevenson. Optimal measurement points for program frequency counts. BIT, 13(3):313–322, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=
0006-3835&volume=13&issue=3&spage=
313.

Lauesen:1973:PCO

[587] Søren Lauesen. Program control of operating systems. BIT, 13(3):323–337, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=323.

Riess:1973:EEH

[588] R. D. Riess. Error estimates of Hermite interpolation. BIT, 13(3):338–343, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=338.

Wedin:1973:ARD

[589] Per-Åke Wedin. On the almost rank deficient case of the least squares problem. BIT, 13(3):344–354, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=344.

Valiaho:1973:PPC

[590] H. Väliaho. A procedure for parameterizing a constraint in linear programming. BIT, 13(3):355–369, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=3&spage=355.

Bohman:1973:NPF

[591] J. Bohman. New primes of the form n ^{4} +1. BIT, 13(3):370–372, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=3&spage=370.

Tuan:1973:ECM

[592] P. D. Tuan. An extension of Clenshaw’s method for linear differential equations. BIT, 13(3):372–374, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=3&spage=372.

Anonymous:1973:Pc

[593] Anonymous. Problems. BIT, 13(3): 375–378, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=3&spage=375.

Anonymous:1973:NBJ

[594] Anonymous. New books and journals. BIT, 13(3):378, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=3&spage=378.

Anonymous:1973:Nb

[595] Anonymous. Notes. BIT, 13(3):379, September 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=3&spage=379.

Anonymous:1973:Cc

[596] Anonymous. Contents. BIT, 13 (3):380, 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Brunner:1973:NSN

[597] H. Brunner. On the numerical solution of nonlinear Volterra integrodifferential equations. BIT, 13(4): 381–390, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=381.

Chipman:1973:IRK

[598] F. H. Chipman. The implementation of Runge–Kutta implicit processes. BIT, 13(4):391–393, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=391.

Espelid:1973:ASA

[599] Terje O. Espelid. Analysis of a shellsort algorithm. BIT, 13(4):394–400, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=394.

Hoskins:1973:SPS

[600] W. D. Hoskins and D. S. Meek. Successive polynomial spline function approximation. BIT, 13(4):401–407, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=401.

Hunkins:1973:PTM

[601] D. R. Hunkins. Product type multiple integration formulas. BIT, 13 (4):408–414, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=408.

Kambo:1973:OQF

[602] N. S. Kambo. Optimal quadrature formulas. BIT, 13(4):415–422, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=415.

King:1973:IPM

[603] Richard F. King. An improved Pegasus method for root finding. BIT, 13 (4):423–427, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=423.

Madsen:1973:SNE

[604] Kaj Madsen. On the solution of nonlinear equations in interval arithmetic. BIT, 13(4):428–433, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=428.

McCabe:1973:CCL

[605] J. H. McCabe and G. M. Phillips. On a certain class of Lebesgue constants. BIT, 13(4):434–442, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=434.

Piessens:1973:EAS

[606] Robert Piessens and Maria Branders. The evaluation and application of some modified moments. BIT, 13(4):443–450, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=443.

Piessens:1973:NSA

[607] R. Piessens and P. Verbaeten. Numerical solution of the Abel integral equation. BIT, 13(4):451–457, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=4&spage=451.

Usmani:1973:MHO

[608] Riaz A. Usmani. A method of high-order accuracy for the numerical integration of boundary value problems. BIT, 13(4):458–469, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=4&spage=458.

Weiss:1973:CSM

[609] Richard Weiss. The convergence of shooting methods. BIT, 13(4):470–475, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=4&spage=470.

Wood:1973:SRK

[610] D. Wood. Some remarks on the KH algorithm for s-grammars. BIT, 13 (4):476–489, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=13&issue=4&spage=476.

Overholt:1973:IFD

[611] K. J. Overholt. Integral form of divided differences. BIT, 13(4):490–492, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=490.

Anonymous:1973:Pd

[612] Anonymous. Problems. BIT, 13(4): 493–495, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=493.

Ekenberg:1973:BR

[613] Bertil Ekenberg. Books received. BIT, 13(4):496–497, December 1973. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=13&issue=4&spage=

496. Includes review of Introduktion till grafisk informationsbehandling [English: Introduction to graphical information processing], by L. A. Bergström and M. Jern.

Anonymous:1973:A

[614] Anonymous. Acknowledgements. BIT, 13(4):497, December 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=13&issue=4&spage=497.

Anonymous:1973:I

[615] Anonymous. Index. BIT, 13(4): 498–499, 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1973:Cd

[616] Anonymous. Contents. BIT, 13 (4):500, 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bochmann:1974:PBF

[617] G. V. Bochmann and W. W. Armstrong. Properties of Boolean functions with a tree decomposition. BIT, 14(1): 1–13, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=1.

Chawla:1974:MGJ

[618] M. M. Chawla and T. R. Ramakrishnan. Modified Gauss–Jacobi quadrature formulas for the numerical evaluation of Cauchy type singular integrals. BIT, 14(1):14–21, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=14.

Ekblom:1974:MRR

[619] Håkan Ekblom. L_{p}-methods for robust regression. BIT, 14(1):22–32, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=22.

Garey:1974:NSV

[620] L. Garey. The numerical solution of Volterra integral equations with singular kernels. BIT, 14(1):33– 39, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=1&spage=33.

Klove:1974:EPD

[621] Torleiv Kløve. Expressing a prime as difference between two numbers containing all the previous primes. BIT, 14(1):40–44, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=14&
issue=1&spage=40.

Ledgard:1974:CSP

[622] Henry F. Ledgard. The case for structured programming. BIT, 14(1):45– 57, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=1&spage=45.

Novak:1974:FRG

[623] Marketa Novak. Full reduction of a general square matrix. BIT, 14(1):58– 62, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=1&spage=58.

Norsett:1974:OSM

[624] Syvert P. Nørsett. One step methods of Hermite type for numerical integration of stiff systems. BIT, 14(1):63– 77, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=63.

Schneider:1974:TGC

[625] Victor Schneider and Gary Winiger. Translation grammars for compilation and decompilation. BIT, 14(1):78– 86, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=78.

Skelboe:1974:CRI

[626] Stig Skelboe. Computation of rational interval functions. BIT, 14(1):87– 95, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=87.

Winther:1974:CME

[627] R. Winther. A collocation method for eigenvalue problems. BIT, 14(1):96– 105, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=96.

Worland:1974:SEA

[628] P. B. Worland. A stability and error analysis of block methods for the numerical solution of y′′ = f(x, y). BIT, 14(1):106–111, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=14&
issue=1&spage=106.

Barlow:1974:CCF

[629] R. H. Barlow. Convergent continued fraction approximant’s to generalised polylogarithms. BIT, 14(1):112– 116, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=1&spage=112.

Froberg:1974:NRK

[630] Carl-Erik E. Fröberg. New results on the Kummer conjecture. BIT, 14 (1):117–119, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=1&spage=117.

Riesel:1974:CSC

[631] Hans Riesel. A consequence of the Von Staudt–Clausen theorem. BIT, 14(1):120–121, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=1&spage=120.

Anonymous:1974:Pa

[632] Anonymous. Problems. BIT, 14 (1):122–124, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=122.

Eriksson:1974:NBJ

[633] Göran Eriksson. New books and journals. BIT, 14(1):125–127, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=14&
                                        issue=1&spage=125.  Includes re-

view of SIMULA BEGIN, by G. M. Birtwistle, O.-J. Dahl, B. Myhrhaug and K. Nygaard.

Rohman:1974:SCR

[634] Jan Rohman. Some computational results regarding the prime numbers below 2,000,000,000. BIT, 14(1): 127, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=127.

Anonymous:1974:Na

[635] Anonymous. Notes. BIT, 14(1):128– 131, March 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=1&spage=128.

Anonymous:1974:Ca

[636] Anonymous. Contents. BIT, 14 (1):132, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bohman:1974:ESC

[637] J. Bohman. Erratum: “Some Computational Results Regarding the Prime Numbers Below 3, 000, 000, 000” [BIT 13(2), 1973, pp. 242–244]. BIT, 14(1):132, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [575].

Dunham:1974:CIP

[638] Charles B. Dunham. Convergence of interpolation in polynomial Chebyshev approximation. BIT, 14(2):133– 135, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=2&spage=133.

ElTom:1974:NSS

[639] M. E. A. El Tom. On the numerical stability of spline function approximations to solutions of Volterra integral equations of the second kind. BIT, 14(2): 136–143, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=2&spage=136.

Gladwin:1974:SQR

[640] C. J. Gladwin and R. Jeltsch. Stability of quadrature rule methods for first kind Volterra integral equations. BIT, 14(2):144– 151, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=2&spage=144.

Hoskins:1974:SAP

[641] W. D. Hoskins and P. J. Ponzo. Some approximation properties of periodic parametric cubic splines. BIT, 14(2): 152–155, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=2&spage=152.

Karasalo:1974:CTD

[642] Ilkka Karasalo. A criterion for truncation of the QR-decomposition algorithm for the singular linear least squares problem. BIT, 14(2):156– 166, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=2&spage=156.

Linnainmaa:1974:ASK

[643] Seppo Linnainmaa. Analysis of some known methods of improving the accuracy of floating-point sums. BIT, 14(2): 167–202, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=2&spage=167.

Magidin:1974:BPI

[644] Mario Magidin. The ‘best’ partition of an integer. BIT, 14(2):203– 208, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=2&spage=203.

Proskurowski:1974:SUI

[645] Andrzej Proskurowski. Search for a unique incidence matrix of a graph. BIT, 14(2):209–226, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=2&spage=209.

Rheinboldt:1974:AGT

[646] Werner C. Rheinboldt and Charles K. Mesztenyi. Arc graphs and their possible application to sparse matrix problems. BIT, 14(2):227– 239, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=2&spage=227.

Vantilborgh:1974:WSS

[647] H. Vantilborgh. On the working set size and its normal approximation. BIT, 14(2):240–251, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=2&spage=240.

Salzer:1974:SRR

[648] Herbert E. Salzer. Some remarks on Riesel’s series. BIT, 14(2):252– 253, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&

                                        volume=14&issue=2&spage=252. See

[556].

Anonymous:1974:Pb

[649] Anonymous. Problems. BIT, 14(2): 254–256, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=2&spage=254.

Anonymous:1974:NBJb

[650] Anonymous. New books and journals. BIT, 14(2):256, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=14&
                                        issue=2&spage=256.

Anonymous:1974:Nb

[651] Anonymous. Notes. BIT, 14(2):257– 259, June 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=2&spage=257.

Anonymous:1974:Cb

[652] Anonymous. Contents. BIT, 14 (2):260, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Abramson:1974:SDM

[653] H. D. Abramson. A syntax directed macro processor. BIT, 14(3):261–272, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=261.

Arnborg:1974:NAM

[654] Stefan Arnborg. A note on the assignment of measurement points for frequency counts in structured programs. BIT, 14(3):273–278, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=273.

Axelsson:1974:ECS

[655] O. Axelsson. On the efficiency of a class of A-stable methods. BIT, 14 (3):279–287, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=279.

ElTom:1974:SFA

[656] M. E. A. El Tom. On spline function approximations to the solution of Volterra integral equations of the first kind. BIT, 14(3):288–297, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=288.

Hill:1974:NCO

[657] David R. Hill. A new class of one-step methods for the solution of Volterra functional differential equations. BIT, 14(3):298–305, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=298.

Haavie:1974:SEI

[658] Tore Håvie. Some expansions for integrals with weight functions. BIT, 14(3):306–313, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=306.

Nordmark:1974:USC

[659] Staffan Nordmark. Uniform stability of a class of parabolic difference operators. BIT, 14(3):314–325, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=314.

Purdom:1974:SP

[660] Paul Purdom. The size of LALR(1) parsers. BIT, 14(3):326–337, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=326.

Thorelli:1974:DSC

[661] Lars-Erik Thorelli. Design of a small conversational system. BIT, 14(3):

338–346, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=338.

deVillers:1974:HSS

[662] E. v. d. S. de Villers and L. B. Wilson. Hashing the subscripts of a sparse matrix. BIT, 14(3):347–358, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=347.

Billauer:1974:GQD

[663] A. Billauer. On Gaussian quadrature by divided differences of a modified function. BIT, 14(3):359–361, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=359.

Pohl:1974:IID

[664] Peter Pohl. Iterative improvement without double precision in a boundary value problem. BIT, 14(3):361–365, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=3&spage=361.

Anonymous:1974:Pc

[665] Anonymous. Problems. BIT, 14(3): 366–368, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=366.

Maartensson:1974:NBJ

[666] C. Mårtensson. New books and journals. BIT, 14(3):369–371, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=369. In-

cludes review of Avancerad COBOL [English: Advanced COBOL], by O. Björner.

Anonymous:1974:Nc

[667] Anonymous. Notes. BIT, 14(3): 372–373, September 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=14&issue=3&spage=372.

Anonymous:1974:Cc

[668] Anonymous. Contents. BIT, 14 (3):374, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Arnborg:1974:OMM

[669] Stefan Arnborg. Optimal memory management in a system with garbage collection. BIT, 14(4):375–381, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=375.

Genin:1974:AAS

[670] Y. Genin. An algebraic approach to A-stable linear multistepmultiderivative integration formulas. BIT, 14(4):382–406, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=382.

Heacox:1974:AML

[671] Harry C. Heacox, Jr. and Paul W. Purdom, Jr. Analysis of a multi-level timesharing model. BIT, 14(4):407–412, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=407.

Hunter:1974:NIO

[672] D. B. Hunter and Susan J. Parsons. On the numerical integration of an odd periodic function over a half-period. BIT, 14(4):413–415, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=413.

Jarvi:1974:SPA

[673] Timo Järvi and Markku Nurminen. Some problems in automatic process grouping and file consolidation. BIT, 14(4):416–429, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=416.

Lindberg:1974:DPM

[674] Bengt Lindberg. On a dangerous property of methods for stiff differential equations. BIT, 14(4):430–436, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=430.

Makela:1974:EFM

[675] Matti Mäkelä, Olavi Nevanlinna, and Aarne H. Sipilä. Exponentially fitted multistep methods by generalized Hermite–Birkhoff interpolation. BIT, 14(4):437–451, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=437.

Papamichael:1974:CSI

[676] N. Papamichael and J. R. Whiteman. Cubic spline interpolation of harmonic functions. BIT, 14(4):452–459, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=452.

Rus:1974:AAD

[677] Teodor Rus and Ulrich Sinn. An algebraic approach to data organization. BIT, 14(4):460–481, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=460.

Anonymous:1974:Pd

[678] Anonymous. Problems. BIT, 14(4): 482–484, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=482.

Anonymous:1974:NBJd

[679] Anonymous. New books and journals. BIT, 14(4):485, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=485.

Anonymous:1974:Nd

[680] Anonymous. Notes. BIT, 14(4):486, December 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=14&issue=4&spage=486.

Anonymous:1974:A

[681] Anonymous. Acknowledgements. BIT, 14(4):487, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1974:I

[682] Anonymous. Index. BIT, 14(4): 488–489, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1974:Cd

[683] Anonymous. Contents. BIT, 14 (4):490, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anderson:1975:CBL

[684] Ned Anderson and Ilkka Karasalo. On computing bounds for the least singular value of a triangular matrix. BIT, 15(1):1–4, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=1&spage=1.

Caprani:1975:REF

[685] Ole Caprani. Round-off errors in floating-point summation. BIT, 15(1): 5–9, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=1&spage=5.

Enright:1975:CNM

[686] W. H. Enright, T. E. Hull, and B. Lindberg. Comparing numerical methods for stiff systems of O.D.E’s. BIT, 15(1): 10–48, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=10.

Kaufman:1975:VPM

[687] Linda Kaufman. A variable projection method for solving separable nonlinear least squares problems. BIT, 15(1): 49–57, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=49.

Lemeire:1975:BCN

[688] Frans Lemeire. Bounds for condition numbers of triangular and trapezoid matrices. BIT, 15(1):58–64, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=58.

Lemeire:1975:CEI

[689] Frans Lemeire. Computation of equivalent inherent rounding errors in the solution of a set of linear equations. BIT, 15(1):65–71, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=65.

Maus:1975:ATR

[690] Arne Maus. On access to temporary resources. BIT, 15(1):72– 84, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=1&spage=72.

Peters:1975:ISM

[691] J. G. Peters and P. S. Kritzinger. Implementation of Samplesort: a minimal storage tree sort. BIT, 15(1):85– 93, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=1&spage=85.

Schorr:1975:NIC

[692] B. Schorr. Numerical inversion of a class of characteristic functions. BIT, 15(1):94–102, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=15&
                                        issue=1&spage=94.

Corradi:1975:ACM

[693] C. Corradi. On the acceleration of the convergence of monotone iteration processes for solving systems of linear algebraic equations. BIT, 15(1):103– 104, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=1&spage=103.

Donaldson:1975:NOQ

[694] J. D. Donaldson. A note on osculatory quadrature formulae. BIT, 15(1):105– 106, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=105.

Kronsjo:1975:NNI

[695] Lydia Kronsjö. A note on the ‘nested iterations’ method. BIT, 15(1):107– 110, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=107.

Anonymous:1975:Pa

[696] Anonymous. Problems. BIT, 15 (1):111–113, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=1&spage=111.

Anonymous:1975:NBJa

[697] Anonymous. New books and journals. BIT, 15(1):114, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=15&
issue=1&spage=114.

Anonymous:1975:Na

[698] Anonymous. Notes. BIT, 15(1):114– 115, March 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=1&spage=114.

Anonymous:1975:Ca

[699] Anonymous. Contents. BIT, 15 (1):116, 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Abdelmalek:1975:CSO

[700] Nabih N. Abdelmalek. Chebyshev solution of overdetermined systems of linear equations. BIT, 15(2):117– 129, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=117.

Barwell:1975:SSP

[701] V. K. Barwell. Special stability problems for functional differential equations. BIT, 15(2):130– 135, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=130.

Brunner:1975:ASA

[702] H. Brunner. On the approximate solution of the Abel integral equation with discontinuous solution. BIT, 15(2): 136–143, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=136.

Cornelius:1975:DRA

[703] B. J. Cornelius and G. H. Kirby. Depth of recursion and the Ackermann function. BIT, 15(2):144– 150, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=144.

Dahl:1975:DEE

[704] Grete Dahl and Sverre Storøy. Decomposed enumeration of extreme points in the linear programming problem. BIT, 15(2):151– 157, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=151.

Dershowitz:1975:SLF

[705] Nachum Dershowitz. A simplified loop-free algorithm for generating permutations. BIT, 15(2):158– 164, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=158.

Linnainmaa:1975:TAS

[706] Seppo Linnainmaa. Towards accurate statistical estimation of rounding errors in floating-point computations. BIT, 15(2):165–173, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=165.

Rey:1975:LPM

[707] W. Rey. On least p-th power methods in multiple regressions and location estimations. BIT, 15(2):174– 184, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=174.

Sameh:1975:IES

[708] A. Sameh, J. Lermit, and K. Noh. On the intermediate eigenvalues of symmetric sparse matrices. BIT, 15(2): 185–191, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=185.

Tafvelin:1975:DME

[709] Sven Tafvelin. Dynamic microprogramming and external subroutine calls in a Multics-type environment. BIT, 15(2): 192–202, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=192.

Tafvelin:1975:ACP

[710] Sven Tafvelin and Åke Wikström. Aspects of compact programs and directly executed languages. BIT, 15(2):203– 214, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=203.

Tarnlund:1975:CRR

[711] S.-Å. Tärnlund. Computation of resource requirements. BIT, 15(2):215– 220, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=215.

Hunter:1975:NNE

[712] D. B. Hunter and Susan J. Parsons. A note on the numerical evaluation of finite integrals of oscillatory functions. BIT, 15(2):221– 223, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=221.

Anonymous:1975:Pb

[713] Anonymous. Problems. BIT, 15(2): 224–226, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=2&spage=224.

Henrickson:1975:NBJ

[714] Sten Henrickson. New books and journals. BIT, 15(2):227–231, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=227. In-

cludes review of Concise Survey of Computer Methods, by Peter Naur.

Anonymous:1975:Nb

[715] Anonymous. Notes. BIT, 15(2):232– 234, June 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=2&spage=232.

Anonymous:1975:Cb

[716] Anonymous. Contents. BIT, 15 (2):235, 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Froberg:1975:NIB

[717] Carl-Erik Fröberg. Niels Ivar Bech: 22.8.1920 – 25.7.1975. BIT, 15(3): 237–238, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=237.

Bohman:1975:NRG

[718] Jan Bohman and Carl-Erik Fröberg. Numerical results on the Goldbach conjecture. BIT, 15(3):239–243, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=239.

Creedon:1975:SPS

[719] David M. Creedon and John J. H. Miller. The stability properties of qstep backward difference schemes. BIT, 15(3):244–249, September 1975. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=15&issue=3&spage=
244.

Farmer:1975:CIF

[720] M. R. Farmer and G. Loizou. A class of iteration functions for improving, simultaneously, approximations to the zeros of a polynomial. BIT, 15(3):250–258, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=3&spage=250.

Halaas:1975:EDC

[721] Arne Halaas. Event-driven control statements. BIT, 15(3):259–271, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=3&spage=259.

Hoskins:1975:LDR

[722] W. D. Hoskins and D. S. Meek. Linear dependence relations for polynomial splines at midknots. BIT, 15(3): 272–276, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=272.

Laurie:1975:PIR

[723] D. P. Laurie. Propagation of initial rounding error in Romberg-like quadrature. BIT, 15(3):277–282, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=277.

Oliver:1975:SIP

[724] J. Oliver and A. Ruffhead. The selection of interpolation points in numerical differentiation. BIT, 15(3):283–295, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=283.

Patrick:1975:GCA

[725] Merrell L. Patrick and Donald G. Saari. A globally convergent algorithm for determining approximate real zeros of a class of functions. BIT, 15(3):296–303, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=296.

Rose:1975:AFM

[726] Lawrence L. Rose and Malcolm H. Gotterer. An analysis of file movement under dynamic file management strategies. BIT, 15(3):304–313, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=3&spage=304.

Strom:1975:ND

[727] T. Ström and J. N. Lyness. On numerical differentiation. BIT, 15(3):314–322, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=3&spage=314.

Woodhouse:1975:NSM

[728] D. Woodhouse. A note on the secant method. BIT, 15(3):323–327, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=3&spage=323.

Atkinson:1975:HMS

[729] L. V. Atkinson. Hashing matrix subscripts. BIT, 15(3):328–330, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=3&spage=328.

Pollard:1975:MCM

[730] J. M. Pollard. A Monte Carlo method for factorization. BIT, 15(3):331–334, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=331.

Anonymous:1975:Pc

[731] Anonymous. Problems. BIT, 15(3): 335–337, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=335.

Anonymous:1975:NBJc

[732] Anonymous. New books and journals. BIT, 15(3):337, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=337.

Anonymous:1975:Nc

[733] Anonymous. Notes. BIT, 15(3): 338–339, September 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=3&spage=338.

Anonymous:1975:Cc

[734] Anonymous. Contents. BIT, 15 (3):340, 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Boffey:1975:LFC

[735] T. B. Boffey. The linearization of flow charts. BIT, 15(4):341–350, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=341.

Borosh:1975:MNE

[736] I. Borosh. More numerical evidence on the uniqueness of Markov numbers. BIT, 15(4):351–357, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=351.

Butcher:1975:SPI

[737] J. C. Butcher. A stability property of implicit Runge–Kutta methods. BIT, 15(4):358–361, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=358.

Caprani:1975:CMI

[738] Ole Caprani and Kaj Madsen. Contraction mappings in interval analysis. BIT, 15(4):362–366, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=362.

Chawla:1975:OAF

[739] M. M. Chawla. Optimal approximate formulas with polynomial precision. BIT, 15(4):367–372, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=367.

Cohen:1975:CLP

[740] Jacques Cohen and Peter Grossman. Compilation of linear partial differential equations into finite-difference programs. BIT, 15(4):373–380, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=373.

Dahlquist:1975:RGP

[741] Germund Dahlquist, Ingemar Ingemarsson, and Hans Riesel. A randomly generated program for automatic identity checking. BIT, 15(4):381–384, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=381.

Ekenberg:1975:CFU

[742] Bertil Ekenberg. Curve fitting by use of graphic display. BIT, 15(4):385–393, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=385.

Fries:1975:CPS

[743] Göran Fries. On compatibility — A proposed solution. BIT, 15(4):394–400, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=394.

Garey:1975:BMN

[744] L. Garey. Block methods for nonlinear Volterra integral equations. BIT, 15(4):401–408, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=401.

Jonassen:1975:AAP

[745] Arne Jonassen and Ole-Johan Dahl. Analysis of an algorithm for priority queue administration. BIT, 15(4): 409–422, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=409.

Klove:1975:RIS

[746] Torleiv Kløve. Representations of integers as sums of powers with increasing exponents — II. BIT, 15(4):423–425, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=423.

Phillips:1975:EEL

[747] G. M. Phillips and B. N. Sahney. An error estimate for least squares approximation. BIT, 15(4):426–430, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=426.

Wise:1975:BVL

[748] David S. Wise, Daniel P. Friedman, Stuart C. Shapiro, and Mitchell Wand. Boolean valued loops. BIT, 15(4): 431–451, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=431.

Anonymous:1975:Pd

[749] Anonymous. Problems. BIT, 15(4): 452–454, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=452.

Naur:1975:NBJ

[750] P. Naur. New books and journals. BIT, 15(4):455–457, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=15&issue=4&spage=455. In-

cludes review of Operating Systems Principles, by P. Brinch Hansen.

Anonymous:1975:Nd

[751] Anonymous. Notes. BIT, 15(4): 458–459, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=458.

Anonymous:1975:A

[752] Anonymous. Acknowledgements. BIT, 15(4):460, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=15&issue=4&spage=460.

Anonymous:1975:I

[753] Anonymous. Index. BIT, 15(4): 461–462, 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1975:Cd

[754] Anonymous. Contents. BIT, 15 (4):463, 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Aulin:1976:NLM

[755] Allan Aulin. Newton-like methods for two-point boundary value problems. BIT, 16(1):1–12, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=16&
issue=1&spage=1.

Burkhard:1976:PMR

[756] Walter A. Burkhard. Partial match retrieval. BIT, 16(1):13–31, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=13.

Elliott:1976:PIR

[757] David Elliott and D. F. Paget. Product-integration rules and their convergence. BIT, 16(1):32–40, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=32.

Goodman:1976:REF

[758] Richard Goodman. On roundoff error in fixed-point multiplication. BIT, 16(1):41–51, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=16&
issue=1&spage=41.

Haegemans:1976:CSI

[759] Ann Haegemans. Circularly symmetrical integration formulas for two-dimensional circularly symmetrical regions. BIT, 16(1):52–59, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=52.

Hilden:1976:ERC

[760] J. Hilden. Elimination of recursive calls using a small table of ‘randomly’ selected function values. BIT, 16(1):60– 73, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=60.

Jeltsch:1976:NSM

[761] Rolf Jeltsch. Note on A-stability of multistep multiderivative methods. BIT, 16(1):74–78, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=16&
issue=1&spage=74.

Nevanlinna:1976:EBS

[762] Olavi Nevanlinna. On error bounds for G-stable methods. BIT, 16(1):79– 84, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=79.

Pajunen:1976:CGF

[763] Seppo Pajunen. Computations on the growth of the first factor for prime cyclotomic fields. BIT, 16(1):85– 87, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=85.

Wang:1976:ABP

[764] Arne Wang. An axiomatic basis for proving total correctness of GOTO programs. BIT, 16(1):88– 102, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=88.

Wichmann:1976:AFS

[765] B. A. Wichmann. Ackermann’s function: a study in the efficiency of calling procedures. BIT, 16(1):103– 110, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=103.

Anonymous:1976:Pa

[766] Anonymous. Problems. BIT, 16 (1):111–113, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=111.

Ekenberg:1976:NBJ

[767] Bertil Ekenberg. New books and journals. BIT, 16(1):114–115, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=16&
issue=1&spage=114. Includes review

of Grafisk databehandling [English: Graphical data processing], by K. Bø and S. Hov Moen.

Anonymous:1976:Na

[768] Anonymous. Notes. BIT, 16(1):116– 122, March 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=1&spage=116.

Anonymous:1976:Ca

[769] Anonymous. Contents. BIT, 16 (1):123, 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cornelius:1976:PTR

[770] B. J. Cornelius and G. H. Kirby. A programming technique for recursive procedures. BIT, 16(2):125– 132, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=125.

Espelid:1976:RSD

[771] Terje O. Espelid. On replacement selection and Dinsmore’s improvement. BIT, 16(2):133–142, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=133.

Lemeire:1976:EMO

[772] Frans Lemeire. Equilibration of matrices to optimize backward numerical stability. BIT, 16(2):143– 145, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=143.

Linnainmaa:1976:TEA

[773] Seppo Linnainmaa. Taylor expansion of the accumulated rounding error. BIT, 16(2):146–160, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=146.

Rink:1976:ASC

[774] R. A. Rink and B. P. Guru. Analytical solutions for a class of nonlinear differential equations using Altran. BIT, 16(2):161–171, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=161.

Siemieniuch:1976:PCR

[775] J. L. Siemieniuch. Properties of certain rational approximations to e^{−z}. BIT, 16(2):172–191, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=16&
issue=2&spage=172.

Szwarcfiter:1976:SSE

[776] Jayme L. Szwarcfiter and Peter E. Lauer. A search strategy for the elementary cycles of a directed graph. BIT, 16(2):192–204, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=192.

Wang:1976:DEW

[777] Jesse Y. Wang. On the discretisation error of the weighted Simpson rule. BIT, 16(2):205– 214, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=205.

Wouk:1976:CIV

[778] A. Wouk. Collocation for initial value problems. BIT, 16(2):215– 222, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=215.

Chipman:1976:NIS

[779] F. H. Chipman. A note on implicit A-stable R-K methods with parameters. BIT, 16(2):223–225, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=2&spage=223.

Wanner:1976:SPN

[780] Gerhard Wanner. A short proof on nonlinear A-stability. BIT, 16(2):226– 227, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=2&spage=226.

Anonymous:1976:Pb

[781] Anonymous. Problems. BIT, 16(2): 228–230, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=2&spage=228.

Anonymous:1976:NBJb

[782] Anonymous. New books and journals. BIT, 16(2):231, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=16&
                                        issue=2&spage=231.

Anonymous:1976:Nb

[783] Anonymous. Notes. BIT, 16(2):232– 234, June 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=2&spage=232.

Anonymous:1976:Cb

[784] Anonymous. Contents. BIT, 16 (2):235, 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Butcher:1976:IIR

[785] J. C. Butcher. On the implementation of implicit Runge–Kutta methods. BIT, 16(3):237–240, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=237.

Campbell:1976:ESP

[786] J. A. Campbell, J. G. Kent, and R. J. Moore. Experiments with a symbolic programming system for complex analysis. BIT, 16(3):241–256, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=241.

Davey:1976:GEB

[787] D. P. Davey and N. F. Stewart. Guaranteed error bounds for the initial value problem using polytope arithmetic. BIT, 16(3):257–268, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=257.

Hansson:1976:ZZP

[788] Åke Hansson and Sten-Åke Tärnlund. Zig-zag procedures for memory allocation and retrieval of dense equilateral arrays of dynamically varying order. BIT, 16(3):269–274, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=269.

Johnson:1976:CG

[789] Donald B. Johnson and Ravi Sethi. A characterization of LL(1) grammars. BIT, 16(3):275–280, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=275.

Lyche:1976:DCS

[790] Tom Lyche. Discrete cubic spline interpolation. BIT, 16(3):281–290, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=281.

Nourein:1976:RDU

[791] Abdel-Wahab M. Nourein. Root determination by use of padé approximants. BIT, 16(3):291–297, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=291. See

comments [867].

Paulik:1976:OAB

[792] Augustin Paulik. On the optimal approximation of bounded linear functionals in Hilbert spaces of analytic functions. BIT, 16(3):298–307, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=298.

Pohl:1976:BEA

[793] Peter Pohl. On binary-exponent alternating sums. BIT, 16(3):308–312, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=308.

Tsamasphyros:1976:NIM

[794] G. Tsamasphyros and P. S. Theocaris. Numerical inversion of Mellin transforms. BIT, 16(3):313–321, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=313.

Wang:1976:CSP

[795] Arne Wang. A case study in program transformation. BIT, 16(3):322–331, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=322.

Wilson:1976:SST

[796] L. B. Wilson. Sequence search trees: Their analysis using recurrence relations. BIT, 16(3):332–337, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=332.

Gonnet:1976:SNC

[797] Gaston H. Gonnet. A short note on convergence near a high order zero. BIT, 16(3):338–339, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=338.

Knobe:1976:CSI

[798] Bruce Knobe and Gideon Yuval. Compilers should indent. BIT, 16(3): 340–343, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=340.

Anonymous:1976:Pc

[799] Anonymous. Problems. BIT, 16(3): 344–346, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=3&spage=344.

Anonymous:1976:NLM

[800] Anonymous. Newton-like methods for two-point boundary value problems. BIT, 16(3):346, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=346.

Anonymous:1976:Nc

[801] Anonymous. Notes. BIT, 16(3): 346–347, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=346.

Anonymous:1976:NBJc

[802] Anonymous. New books and journals. BIT, 16(3):347, September 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=3&spage=347.

Anonymous:1976:Cc

[803] Anonymous. Contents. BIT, 16 (3):348, 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Andreassen:1976:LCA

[804] D. O. Andreassen and G. A. Watson. Linear Chebyshev approximation without Chebyshev sets. BIT, 16 (4):349–362, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=349.

Avdelas:1976:ICR

[805] G. Avdelas and A. Hadjidimos. On improving the convergence rates of extrapolated alternating direction implicit schemes. BIT, 16(4):363–373, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=363.

Chawla:1976:TPS

[806] L. M. Chawla and John E. Maxfield. On a table of the product-sum partition function p(n, m). BIT, 16(4):374–377, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=374.

Hsu:1976:NWB

[807] Harry T. Hsu. A note on weighted buddy systems for dynamic storage allocation. BIT, 16(4):378–382, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=378.

Jackson:1976:SFF

[808] L. W. Jackson. The A-stability of a family of fourth order methods. BIT, 16(4):383–387, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=383.

Norsett:1976:RKM

[809] Syvert P. Nørsett. Runge–Kutta methods with a multiple real eigenvalue only. BIT, 16(4):388–393, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=388. See

erratum [890].

Smith:1976:EEC

[810] H. V. Smith. Error estimates for a class of quadrature formulae. BIT, 16(4):394–399, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=394.

Tafvelin:1976:TDB

[811] Sven Tafvelin. The time dependent behaviour of high speed sequential files on cycling storage. BIT, 16(4):400–415, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=400.

Taylor:1976:SVI

[812] P. J. Taylor. The solution of Volterra integral equations of the first kind using inverted differentiation formulae. BIT, 16(4):416–425, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=4&spage=416.

Thorelli:1976:FCG

[813] Lars-Erik Thorelli. A fast compactifying garbage collector. BIT, 16(4): 426–441, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=4&spage=426.

Wise:1976:TGA

[814] David S. Wise and Dan C. Watson. Tuning Garwick’s algorithm for repacking sequential storage. BIT, 16(4): 442–450, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=4&spage=442.

Osterby:1976:PDM

[815] O. Østerby. Prime decompositions with minimum sum. BIT, 16(4):451–458, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=4&spage=451.

Boyer:1976:ADI

[816] M. Boyer and L. Paquette. An algorithm to decide if the intersection of convex polyhedral cones has a non empty interior. BIT, 16(4):459–461, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=459.

Eriksson:1976:NDS

[817] Jan Eriksson. A note on the decomposition of systems of sparse nonlinear equations. BIT, 16(4):462–465, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=462.

Anonymous:1976:Pd

[818] Anonymous. Problems. BIT, 16(4): 466–468, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=466.

Anonymous:1976:NBJd

[819] Anonymous. New books and journals. BIT, 16(4):469, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=16&issue=4&spage=469.

Anonymous:1976:Nd

[820] Anonymous. Notes. BIT, 16(4): 470–473, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=4&spage=470.

Anonymous:1976:A

[821] Anonymous. Acknowledgements. BIT, 16(4):473, December 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=16&issue=4&spage=473.

Anonymous:1976:I

[822] Anonymous. Index. BIT, 16(4): 474–475, 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1976:Cd

[823] Anonymous. Contents. BIT, 16 (4):476, 1976. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Choy:1977:BOB

[824] David M. Choy and C. K. Wong. Bounds for optimal α-β binary trees. BIT, 17(1):1–15, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=17&
                                        issue=1&spage=1.

Deo:1977:NAD

[825] Narsingh Deo, J. M. Davis, and R. E. Lord. A new algorithm for digraph isomorphism. BIT, 17(1):16– 30, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=17&issue=1&spage=16.

Donaldson:1977:COQ

[826] John D. Donaldson. Convergence of osculatory quadrature formulae. BIT, 17(1):31–38, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=17&
issue=1&spage=31.

Kaagstrom:1977:BPB

[827] Bo Kågström. Bounds and perturbation bounds for the matrix exponential. BIT, 17(1):39–57, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=1&spage=39.

Nevanlinna:1977:NIN

[828] Olavi Nevanlinna. On the numerical integration of nonlinear initial value problems by linear multistep methods. BIT, 17(1):58–71, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=1&spage=58.

Ramsin:1977:CSA

[829] Håkan Ramsin and Per-Åke Wedin. A comparison of some algorithms for the nonlinear least squares problem. BIT, 17(1):72–90, March 1977. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=17&
                                        issue=1&spage=72.

Skelboe:1977:COS

[830] Stig Skelboe. The control of order and steplength for backward differentiation methods. BIT, 17(1):91– 107, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=1&spage=91.

Smith:1977:EEG

[831] H. V. Smith. Error estimates for Gauss–Legendre quadrature of integrands possessing Dirichlet series expansions. BIT, 17(1):108–112, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=1&spage=108.

Pajunen:1977:CGF

[832] Seppo Pajunen. Computations of the growth of the first factor for prime cyclotomic fields II. BIT, 17(1):113– 114, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=1&spage=113.

Anonymous:1977:Pa

[833] Anonymous. Problems. BIT, 17 (1):115–116, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=1&spage=115.

Henriksson:1977:NBJ

[834] S. Henriksson. New books and journals. BIT, 17(1):116–117, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=17&
issue=1&spage=116. Includes review

of Niels Ivar Bech — En epoke i EDB-Udvikling i Danmark (Danish) [Niels Ivar Bech — A time of EDP Development in Denmark], by P. Sveistrup, P. Naur, H. B. Hansen and Chr. Gram.

Henriksson:1977:R

[835] Sten Henriksson. Reviews. BIT, 17 (1):117–118, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=1&spage=117.

Anonymous:1977:Na

[836] Anonymous. Notes. BIT, 17(1): 119, March 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=1&spage=119.

Anonymous:1977:Ca

[837] Anonymous. Contents. BIT, 17 (1):120, 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bays:1977:SSE

[838] Carter Bays and Richard H. Hudson. The segmented sieve of Eratosthenes and primes in arithmetic progressions to 10^{12}. BIT, 17(2):121– 127, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=121.

Chawla:1977:SOT

[839] M. M. Chawla. A sixth order tridiagonal finite difference method for non-linear two-point boundary value problems. BIT, 17(2):128– 133, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=128.

Elden:1977:ARI

[840] Lars Eldén. Algorithms for the regularization of ill-conditioned least squares problems. BIT, 17(2):134– 145, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=134.

Frank:1977:IDC

[841] Reinhard Frank and Christoph W. Ueberhuber. Iterated defect correction for the efficient solution of stiff systems of ordinary differential equations. BIT, 17(2):146– 159, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=146.

Glad:1977:OFW

[842] Torkel Glad and Allen A. Goldstein. Optimization of functions whose values are subject to small errors. BIT, 17(2): 160–169, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=160.

Gonnet:1977:SZF

[843] Gaston H. Gonnet. On the structure of zero finders. BIT, 17(2):170– 183, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=170.

Hoskins:1977:NSM

[844] W. D. Hoskins, D. S. Meek, and D. J. Walton. The numerical solution of the matrix equation XA + AY = F. BIT, 17(2):184–190, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=184.

Lau:1977:REA

[845] Terence C.-Y Y. Lau. Rational exponential approximation with real poles. BIT, 17(2):191–199, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=191.

Norsett:1977:AOR

[846] Syvert P. Nørsett and Arne Wolfbrandt. Attainable order of rational approximations to the exponential function with only real poles. BIT, 17(2): 200–208, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=200. See

note [869].

Steihaug:1977:NSL

[847] Trond Steihaug. Numerical solution of linear algebraic equations where the coefficient matrix is a polynomial of a square matrix. BIT, 17(2):209– 214, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=209.

Tarnlund:1977:HCC

[848] Sten-Åke Tärnlund. Horn clause computability. BIT, 17(2):215–226, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=215.

Usmani:1977:NIB

[849] R. A. Usmani. On the numerical integration of a boundary value problem involving a fourth order linear differential equation. BIT, 17(2):227– 234, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=227.

Scherer:1977:NRL

[850] R. Scherer. A note on Radau and Lobatto formulae for O.D.E.’s. BIT, 17(2):235–238, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=235.

Weintraub:1977:PAP

[851] Sol Weintraub. Primes in arithmetic progression. BIT, 17(2):239– 243, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=2&spage=239.

Anonymous:1977:Pb

[852] Anonymous. Problems. BIT, 17(2): 244–246, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=2&spage=244.

Anonymous:1977:NBJb

[853] Anonymous. New books and journals. BIT, 17(2):247, June 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=17&
                                        issue=2&spage=247.

Anonymous:1977:Cb

[854] Anonymous. Contents. BIT, 17 (2):248, 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Abdelmalek:1977:DLR

[855] Nabih N. Abdelmalek. The discrete linear restricted Chebyshev approximation. BIT, 17(3):249–261, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=249.

Brannigan:1977:UAG

[856] M. Brannigan. Uniform approximation by generalised polynomials. BIT, 17(3):262–269, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=262.

Cash:1977:CCM

[857] J. R. Cash. On a class of cyclic methods for the numerical integration of stiff systems of O.D.E.s. BIT, 17(3):270–280, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=270.

Chawla:1977:EOT

[858] M. M. Chawla. An eighth order tridiagonal finite difference method for nonlinear two-point boundary value problems. BIT, 17(3):281–285, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=281.

Gregory:1977:BCR

[859] Robert Todd Gregory and David W. Matula. Base conversion in residue number systems. BIT, 17(3):286–302, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=286.

Jankowski:1977:IRI

[860] M. Jankowski and H. Woźniakowski. Iterative refinement implies numerical stability. BIT, 17(3):303–311, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=303.

Keech:1977:TOS

[861] Malcolm S. Keech. A third order, semi-explicit method in the numerical solution of first kind Volterra integral equations. BIT, 17(3):312–320, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=312.

King:1977:SMM

[862] Richard F. King. A secant method for multiple roots. BIT, 17(3):321–328, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=321.

Linz:1977:AMS

[863] Peter Linz. An analysis of a method for solving singular integral equations. BIT, 17(3):329–337, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=329.

Smith:1977:MIO

[864] H. V. Smith. A method for the integration of oscillatory functions. BIT, 17(3):338–343, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=338.

Solntseff:1977:PDT

[865] N. Solntseff and D. Wood. Pyramids: a data type for matrix representation in Pascal. BIT, 17(3):344–350, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=344.

Wise:1977:OBR

[866] David S. Wise and Daniel P. Friedman. The one-bit reference count. BIT, 17(3):351–359, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=351.

Claessens:1977:CRF

[867] G. Claessens, G. Loizou, and L. Wuytack. Comments on a root finding method using Padé approximation. BIT, 17(3):360–361, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=360. See

[791].

Tennent:1977:NFP

[868] R. D. Tennent. A note on files in Pascal. BIT, 17(3):362–366, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=362.

Wolfbrandt:1977:NRR

[869] Arne Wolfbrandt. A note on a recent result of rational approximations to the exponential function. BIT, 17(3):367–368, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=367. See

[846].

Anonymous:1977:Pc

[870] Anonymous. Problems. BIT, 17(3): 369–370, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=369.

Anonymous:1977:NBJc

[871] Anonymous. New books and journals. BIT, 17(3):371, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=3&spage=371.

Anonymous:1977:Nb

[872] Anonymous. Notes. BIT, 17(3): 372–373, September 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=3&spage=372.

Butcher:1977:SIR

[873] J. C. Butcher. On A-stable implicit Runge–Kutta methods. BIT, 17(4): 375–378, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=375.

Evans:1977:EAD

[874] D. J. Evans and G. Avdelas. On extrapolated alternating direction implicit (E.A.D.I.) methods of second order. BIT, 17(4):379–385, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=379.

Froberg:1977:AEN

[875] Carl-Erik Fröberg. Accurate estimation of the number of binary partitions. BIT, 17(4):386–391, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=386.

Gribble:1977:FPI

[876] J. D. Gribble. Further properties of inner product quadrature formulas. BIT, 17(4):392–408, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=392. See

correction [1089].

Hansen:1977:SML

[877] E. Hansen, M. Patrick, and J. Rusnak. Some modifications of Laguerre’s method. BIT, 17(4):409–417, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=409.

Haavie:1977:RIP

[878] Tore Håvie. Romberg integration as a problem in interpolation theory. BIT, 17(4):418–429, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=418.

Miklosko:1977:FAR

[879] Jozef Mikloško. A fast algorithm for repeated computation of linear recurrence relations. BIT, 17(4):430–436, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=430.

Rheinboldt:1977:EMP

[880] W. C. Rheinboldt, C. K. Mesztenyi, and J. M. Fitzgerald. On the evaluation of multivariate polynomials and their derivatives. BIT, 17(4):437–450, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=437.

Sajaniemi:1977:SDA

[881] Jorma Sajaniemi. Some difficulties with abstract and concrete syntax. BIT, 17(4):451–457, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=451.

Tsamasphyros:1977:CGQ

[882] G. J. Tsamasphyros and P. S. Theocaris. On the convergence of a Gauss quadrature rule for evaluation of Cauchy type singular integrals. BIT, 17(4):458–464, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=458.

Waterman:1977:MGC

[883] M. S. Waterman. Multidimensional greatest common divisor and Lehmer algorithms. BIT, 17(4):465–478, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=465.

Bohman:1977:NNP

[884] J. Bohman. A note on the number of partitions other than binary. BIT, 17(4):479–480, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=479.

Dhawan:1977:NDA

[885] A. K. Dhawan and V. K. Srivastava. On a new division algorithm. BIT, 17(4):481–485, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=481.

Evans:1977:FDM

[886] D. J. Evans. Fast A.D.I. methods for the solution of linear parabolic partial differential equations involving 2 space dimensions. BIT, 17(4):486–491, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=486.

Lawson:1977:NEI

[887] J. D. Lawson and J. Ll. Morris. A note on the efficient implementation of splitting methods in two space variables. BIT, 17(4):492–493, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=492.

Anonymous:1977:Pd

[888] Anonymous. Problems. BIT, 17(4): 494–496, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=494.

Anonymous:1977:NBJd

[889] Anonymous. New books and journals. BIT, 17(4):497, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=497.

Norsett:1977:ERK

[890] Syvert P. Nørsett. Erratum: “Runge– Kutta methods with a multiple real eigenvalue only” [BIT 16(4), 1976, pp. 388–393]. BIT, 17(4):497, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=17&issue=4&spage=497. See

[809].

Anonymous:1977:Nd

[891] Anonymous. Notes. BIT, 17(4): 498–499, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=17&issue=4&spage=498.

Anonymous:1977:A

[892] Anonymous. Acknowledgements. BIT, 17(4):499, 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1977:I

[893] Anonymous. Index. BIT, 17(4): 500–501, 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1977:Cd

[894] Anonymous. Contents. BIT, 17(4): 502–504, 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anderson:1978:SP

[895] T. Anderson and R. W. Witty. Safe programming. BIT, 18(1):1– 8, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=1.

Bickart:1978:ATS

[896] T. A. Bickart and E. I. Jury. Arithmetic tests for A-stability, A(α)-stability, and stiff-stability. BIT, 18(1):9– 21, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=9.

Burrage:1978:SFR

[897] Kevin Burrage. A special family of Runge–Kutta methods for solving stiff differential equations. BIT, 18(1):22– 41, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=1&spage=22.

Caprani:1978:IMI

[898] Ole Caprani and Kaj Madsen. Iterative methods for interval inclusion of fixed points. BIT, 18(1):42– 51, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=1&spage=42.

Coffman:1978:EAA

[899] E. G. Coffman, Jr., Donald B. Johnson, and Joseph Y.-T. Leung. An efficient algorithm for allocating paged drum-like storage. BIT, 18(1):52– 66, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=1&spage=52.

Einbu:1978:DBL

[900] John M. Einbu. Determination of a basis for linear least square approximation. BIT, 18(1):67–71, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=67.

Gjessing:1978:CTP

[901] Stein Gjessing. Compile time preparations for run time scheduling in monitors. BIT, 18(1):73– 83, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=72.

Goh:1978:GCS

[902] B. S. Goh. Global convergence of some differential equation algorithms for solving equations involving positive variables. BIT, 18(1):84–90, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=84.

Sloan:1978:NES

[903] I. H. Sloan. On the numerical evaluation of singular integrals. BIT, 18(1):91–102, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=18&
                                        issue=1&spage=91.

Brown:1978:EOD

[904] R. L. Brown. Evaluation of ordinary differential equation software. BIT, 18(1):103–105, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=18&
issue=1&spage=103.

Fairweather:1978:NEI

[905] G. Fairweather. A note on the efficient implementation of certain Padé methods for linear parabolic problems. BIT, 18(1):106–109, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=1&spage=106.

Klove:1978:TNC

[906] Torleiv Kløve. Two non-closure results on families of L-languages. BIT, 18(1):109–111, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=1&spage=109.

Usmani:1978:NLS

[907] Riaz A. Usmani and F. S. Chebib. A note on the least-square linear approximation. BIT, 18(1):112–115, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=1&spage=112.

Anonymous:1978:Pa

[908] Anonymous. Problems. BIT, 18 (1):115–117, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=115.

Anonymous:1978:NBJa

[909] Anonymous. New books and journals. BIT, 18(1):117, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=18&
                                        issue=1&spage=117.

Anonymous:1978:Na

[910] Anonymous. Notes. BIT, 18(1):118– 123, March 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=1&spage=118.

Apers:1978:RS

[911] Peter M. G. Apers. Recursive samplesort. BIT, 18(2):125–132, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=2&spage=125.

Dahlquist:1978:AUS

[912] Germund Dahlquist. On accuracy and unconditional stability of linear multistep methods for second order differential equations. BIT, 18(2):133– 136, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=133.

Elliott:1978:CPI

[913] David Elliott and D. F. Paget. The convergence of product integration rules. BIT, 18(2):137–141, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=137.

Gustafsson:1978:CFO

[914] Ivar Gustafsson. A class of first order factorization methods. BIT, 18(2): 142–156, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=142.

Iserles:1978:SIR

[915] Arieh Iserles. On the A-stability of implicit Runge–Kutta processes. BIT, 18(2):157–169, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=157.

Jeltsch:1978:SIA

[916] R. Jeltsch. Stability on the imaginary axis and A-stability of linear multistep methods. BIT, 18(2):170– 174, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=2&spage=170.

Knauff:1978:NDT

[917] W. Knauff and A. Paulik. A note on Davis type error bounds. BIT, 18(2): 175–183, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=2&spage=175.

Larson:1978:DH

[918] Per-Åke Larson. Dynamic hashing. BIT, 18(2):184–201, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=18&
                                        issue=2&spage=184. See [1456].

Mjelde:1978:CCF

[919] K. M. Mjelde. Convex-concave fractional programming with each variable occurring in a single constraint. BIT, 18(2):202–210, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=18&
                                        issue=2&spage=202.

Purdom:1978:API

[920] P. W. Purdom. Automatic program indentation. BIT, 18(2):211– 218, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=211.

Wood:1978:CTM

[921] D. Wood. A comparison of two methods of encoding arrays. BIT, 18(2): 219–229, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=219.

Georges:1978:CSC

[922] J. Georges and F. Heymans. On compound statements in a conversational language. BIT, 18(2):230– 233, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=230.

Oliver:1978:NST

[923] J. Oliver. A note on the signs of truncated Chebyshev polynomials. BIT, 18(2):233–235, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=18&
issue=2&spage=233.

Anonymous:1978:Pb

[924] Anonymous. Problems. BIT, 18(2): 236–237, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=2&spage=236.

Anonymous:1978:NBJb

[925] Anonymous. New books and journals. BIT, 18(2):237, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=18&
                                        issue=2&spage=237.

Anonymous:1978:Nb

[926] Anonymous. Notes. BIT, 18(2): 238, June 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=2&spage=238.

Anonymous:1978:Cb

[927] Anonymous. Contents. BIT, 18 (2):239, 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Aggarwal:1978:ADS

[928] Vijay B. Aggarwal and James W. Burgmeier. An algorithmic decomposition system for second order recurrence relations. BIT, 18(3):241–258, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=241.

Arnborg:1978:AND

[929] Stefan Arnborg. Analysis of nondeterministic drum scheduling. BIT, 18 (3):259–264, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=259.

Cullum:1978:SCF

[930] Jane Cullum. The simultaneous computation of a few of the algebraically largest and smallest eigenvalues of a large, sparse, symmetric matrix. BIT, 18(3):265–275, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=265.

Enright:1978:NES

[931] W. H. Enright and Steven M. Serbin. A note on the efficient solution of matrix pencil systems. BIT, 18(3):276–281, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=276.

Gregory:1978:UFS

[932] Robert Todd Gregory. The use of finite segment p-adic arithmetic for exact computation. BIT, 18(3):282–300, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=282.

Houstis:1978:CML

[933] E. N. Houstis. Collocation methods for linear elliptic problems. BIT, 18 (3):301–310, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=301.

Jones:1978:RIU

[934] Bush Jones, W. G. Waller, and Arnold Feldman. Root isolation using function values. BIT, 18(3):311–319, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=311.

Nevalainen:1978:ETI

[935] Olli Nevalainen and Jukka Teuhola. The efficiency of two indexed priority queue algorithms. BIT, 18(3):320–333, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=320.

Pacini:1978:TEL

[936] Giuliano Pacini and Maria Simi. Testing equality in LISP-like environments. BIT, 18(3):334–341, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=334.

Patricio:1978:CSF

[937] Fernanda Patrício. Cubic spline functions and initial value problems. BIT, 18(3):342–347, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=342.

Revay:1978:ESD

[938] P. Révay. Evaluation of a system design model. BIT, 18(3):348–361, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=348.

Anderson:1978:NGI

[939] Ned Anderson. A note on a generalized interpolation problem. BIT, 18 (3):362–363, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=362.

Dahl:1978:SLA

[940] Grete Dahl. On scaling in linear algebraic systems. BIT, 18(3):363–365, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=3&spage=363.

Anonymous:1978:Pc

[941] Anonymous. Problems. BIT, 18(3): 366–368, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=366.

Anonymous:1978:NBJc

[942] Anonymous. New books and journals. BIT, 18(3):368, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=368.

Anonymous:1978:Nc

[943] Anonymous. Notes. BIT, 18(3): 369–371, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=3&spage=369.

Anonymous:1978:Cc

[944] Anonymous. Contents. BIT, 18 (3):372, 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Burrage:1978:HOA

[945] Kevin Burrage. High order algebraically stable Runge–Kutta methods. BIT, 18(4):373–383, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=373.

Dahlquist:1978:SES

[946] Germund Dahlquist. G-stability is equivalent to A-stability. BIT, 18(4):

384–401, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=384.

Friedli:1978:ATS

[947] A. Friedli and R. Jeltsch. An algebraic test for A_{0}-stability. BIT, 18 (4):402–414, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=402.

Hansen:1978:GCI

[948] Eldon Hansen. A globally convergent interval method for computing and bounding real roots. BIT, 18(4): 415–424, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=415.

Kriegel:1978:BT

[949] H. P. Kriegel, V. K. Vaishnavi, and D. Wood. 2-3 brother trees. BIT, 18 (4):425–435, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=425.

Krogdahl:1978:VCL

[950] Stein Krogdahl. Verification of a class of link-level protocols. BIT, 18(4): 436–448, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=436.

Lichtman:1978:SPO

[951] Zavdi L. Lichtman. Source program optimization with a generalized for statement. BIT, 18(4):449–453, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=449.

Mjelde:1978:SKT

[952] K. M. Mjelde. Sufficiency of Kuhn–Tucker optimality conditions for a fractional programming problem. BIT, 18(4):454–456, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=454.

Nevanlinna:1978:CMS

[953] Olavi Nevanlinna and Werner Liniger. Contractive methods for stiff differential equations, Part I. BIT, 18(4):457–474, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=457.

Wanner:1978:OSS

[954] G. Wanner, E. Hairer, and S. P. Nørsett. Order stars and stability theorems. BIT, 18(4):475–489, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=475.

Carta:1978:MAR

[955] David G. Carta. Minimax approximation by rational fractions of the inverse polynomial type. BIT, 18(4):490–492, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=490.

Herlestam:1978:CRS

[956] Tore Herlestam. Critical remarks on some public-key cryptosystems. BIT, 18(4):493–496, December 1978. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=18&issue=4&spage=

493. See remarks [1002].

Hopkins:1978:NFE

[957] T. R. Hopkins. Note on a family of economical sixth order finite difference methods for nonlinear 2-point boundary value problems. BIT, 18(4):497–499, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=497.

Jeltsch:1978:LDS

[958] Rolf Jeltsch and Olavi Nevanlinna. Largest disk of stability of explicit Runge–Kutta methods. BIT, 18(4): 500–502, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=500.

Wanner:1978:WSI

[959] G. Wanner, E. Hairer, and S. P. Nørsett. When I-stability implies A-stability. BIT, 18(4):503, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=503.

Anonymous:1978:Pd

[960] Anonymous. Problems. BIT, 18(4): 504–505, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=504.

Anonymous:1978:NBJd

[961] Anonymous. New books and journals. BIT, 18(4):506, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=18&issue=4&spage=506.

Anonymous:1978:Nd

[962] Anonymous. Notes. BIT, 18(4): 506–509, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=506.

Anonymous:1978:A

[963] Anonymous. Acknowledgements. BIT, 18(4):509, December 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=18&issue=4&spage=509.

Anonymous:1978:I

[964] Anonymous. Index. BIT, 18(4): 510–511, 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1978:Cd

[965] Anonymous. Contents. BIT, 18 (4):512, 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Andrade:1979:OHA

[966] Celia Andrade and S. McKee. On optimal high accuracy linear multistep methods for first kind Volterra integral equations. BIT, 19(1):1– 11, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=1.

Avdelas:1979:IEA

[967] G. Avdelas, A. Hadjidimos, and N. P. Krimnianiotis. Improved extrapolated alternating direction implicit schemes for the numerical solution of three-dimensional elliptic problems. BIT, 19(1):12–18, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=12.

Behforooz:1979:IOA

[968] G. H. Behforooz and N. Papamichael. Improved orders of approximation derived from interpolatory cubic splines. BIT, 19(1):19–26, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=1&spage=19.

Chawla:1979:FDM

[969] M. M. Chawla and C. P. Katti. Finite difference methods for twopoint boundary value problems involving high order differential equations. BIT, 19(1):27–33, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=1&spage=27.

Francis:1979:CMR

[970] N. D. Francis. Computation of material requirements for production planning. BIT, 19(1):34–38, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=34.

Merikoski:1979:LBP

[971] Jorma Kaarlo Merikoski. On a lower bound for the Perron eigenvalue. BIT, 19(1):39–42, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=1&spage=39.

Munksgaard:1979:NFC

[972] Niels Munksgaard. New factorization codes for sparse, symmetric and positive definite matrices. BIT, 19(1):43– 52, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=43.

Nevanlinna:1979:CMS

[973] Olavi Nevanlinna and Werner Liniger. Contractive methods for stiff differential equations Part II. BIT, 19(1):53– 72, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=53.

Nijholt:1979:GFC

[974] Anton Nijholt. Grammar functors and covers: From non-left-recursive to Greibach normal form grammars. BIT, 19(1):73–78, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=1&spage=73.

Norsett:1979:RPS

[975] S. P. Nørsett and G. Wanner. The real-pole sandwich for rational approximations and oscillation equations. BIT, 19(1):79–94, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=79.

Phillips:1979:BPA

[976] G. M. Phillips. Best polynomial approximations: a corrected proof. BIT, 19(1):95–97, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=1&spage=95.

Phillips:1979:SDI

[977] G. M. Phillips. Seventh degree integration rules for the cube. BIT, 19(1):98– 103, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=98.

Rabinowitz:1979:ASN

[978] Philip Rabinowitz. On avoiding the singularity in the numerical integration of proper integrals. BIT, 19(1):104– 110, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=104.

Scherer:1979:NCS

[979] Rudolf Scherer. A necessary condition for B-stability. BIT, 19(1):111– 115, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=111.

Ukkonen:1979:AER

[980] Esko Ukkonen. An analysis of the effect of rounding errors on the flow of control in numerical processes. BIT, 19(1):116– 133, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=116.

Channabasappa:1979:NCM

[981] M. N. Channabasappa. A note on the computation of multiple zeros of polynomials by Newton’s method. BIT, 19(1):134–135, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=1&spage=134.

Dew:1979:ECG

[982] P. M. Dew and M. R. West. Estimating and controlling the global error in Gear’s method. BIT, 19(1):135– 137, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=1&spage=135.

Fransen:1979:ADI

[983] Arne Fransén. Accurate determination of the inverse gamma integral. BIT, 19(1):137–138, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=19&
                                        issue=1&spage=137.

Anonymous:1979:Pa

[984] Anonymous. Problems. BIT, 19 (1):139–140, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=1&spage=139.

Anonymous:1979:NBJa

[985] Anonymous. New books and journals. BIT, 19(1):140, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=19&
                                        issue=1&spage=140.

Anonymous:1979:Na

[986] Anonymous. Notes. BIT, 19(1):141– 143, March 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=1&spage=141.

Anonymous:1979:Ca

[987] Anonymous. Contents. BIT, 19 (1):144, 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjorck:1979:APM

[988] Åke Björck and Tommy Elfving. Accelerated projection methods for computing pseudoinverse solutions of systems of linear equations. BIT, 19(2): 145–163, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=145.

Cahlon:1979:IVP

[989] Baruch Cahlon. On the initial value problem for a certain partial differential equation. BIT, 19(2):164– 171, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=164.

Evans:1979:MAD

[990] D. J. Evans and N. M. Missirlis. The modified alternating direction preconditioning method for the numerical solution of the elliptic selfadjoint second order and biharmonic equations. BIT, 19(2):172– 185, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=172.

Hillion:1979:NSC

[991] Pierre Hillion. A new stability criterion for linear discrete systems. BIT, 19(2): 186–195, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=186.

Hisdal:1979:QMA

[992] Ellen Hisdal. Quantitative measure of the amount of information acquired in a learning process. BIT, 19(2):196– 203, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=196.

Haavie:1979:GNT

[993] T. Håvie. Generalized Neville type extrapolation schemes. BIT, 19(2):204– 213, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=204.

Larkin:1979:MSR

[994] F. M. Larkin. A modification of the secant rule derived from a maximum likelihood principle. BIT, 19(2):214– 222, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=214.

Larson:1979:FLL

[995] Per-Åke Larson. Frequency loading and linear probing. BIT, 19(2):223– 228, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=223.

Lyche:1979:NFT

[996] Tom Lyche. A Newton form for trigonometric Hermite interpolation. BIT, 19(2):229–235, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=2&spage=229.

Martinez:1979:TNA

[997] José Mario Martínez. Three new algorithms based on the sequential secant method. BIT, 19(2):236– 243, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=236.

Sanden:1979:VPC

[998] Bo Sandén. Verification of a program complex. BIT, 19(2):244– 255, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=244.

Valiaho:1979:POP

[999] Hannu Väliaho. A procedure for one-parametric linear programming. BIT, 19(2):256–269, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=2&spage=256.

Mjelde:1979:CCF

[1000] K. M. Mjelde. Convex-concave fractional programming — evaluation of solutions and optimality conditions. BIT, 19(2):270, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=270.

Overton:1979:CIM

[1001] Michael L. Overton and Andrzej Proskurowski. Canonical incidence matrices of graphs. BIT, 19(2):271– 273, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=271.

Rivest:1979:CRC

[1002] Ronald L. Rivest. Critical remarks on: “Critical remarks on some publickey cryptosystems” [BIT 18(4), 1978, pp. 493–496; MR 80b:94033 ] by Tore Herlestam. BIT, 19(2):274– 275, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=274. See

[956].

Steiner:1979:CN

[1003] Ray P. Steiner. On Cullen numbers. BIT, 19(2):276–277, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=19&
issue=2&spage=276.

Stepleman:1979:MCE

[1004] R. S. Stepleman. Monotone convergence and effective stopping criteria for numerical processes. BIT, 19(2):278– 281, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=2&spage=278.

Yehudai:1979:NDS

[1005] Amiram Yehudai. A new definition for simple precedence grammars. BIT, 19(2):282–284, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=19&
                                        issue=2&spage=282.

Anonymous:1979:Pb

[1006] Anonymous. Problems. BIT, 19(2): 285–286, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=2&spage=285.

Anonymous:1979:NBJb

[1007] Anonymous. New books and journals. BIT, 19(2):286, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=19&
                                        issue=2&spage=286.

Anonymous:1979:Nb

[1008] Anonymous. Notes. BIT, 19(2): 287, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=2&spage=287.

Anonymous:1979:Cb

[1009] Anonymous. Contents. BIT, 19 (2):288, 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bhat:1979:AOA

[1010] Kabekode V. S. Bhat and Bharat Kinariwala. An algorithm for the n × n optimum assignment problem. BIT, 19(3):289–296, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=289.

Bohman:1979:PS

[1011] Jan Bohman, Carl-Erik Fröberg, and Hans Riesel. Partitions in squares. BIT, 19(3):297–301, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=297.

Devillers:1979:HTG

[1012] R. Devillers and G. Louchard.

Hashing techniques. A global approach. BIT, 19(3):302–311, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=302.

Finn:1979:LTA

[1013] Greg Finn and Ellis Horowitz.

A linear time approximation algorithm for multiprocessor scheduling. BIT, 19(3):312–320, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=312.

Freeman:1979:MCA

[1014] T. L. Freeman. A method for computing all the zeros of a polynomial with real coefficients. BIT, 19(3):321–333, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=321.

Gjessing:1979:MAP

[1015] Stein Gjessing. Monitors with associated processors. BIT, 19(3):334–346, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=334.

Jain:1979:SMP

[1016] M. K. Jain, R. K. Jain, and U. Anantha Krishnaiah. P -stable methods for periodic initial value problems of second order differential equations. BIT, 19(3):347–355, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=347.

Ruhe:1979:AGN

[1017] Axel Ruhe. Accelerated Gauss– Newton algorithms for nonlinear least squares problems. BIT, 19(3):356–367, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=356.

Tsamasphyros:1979:CFE

[1018] G. Tsamasphyros and P. S. Theocaris. Cubature formulas for the evaluation of surface singular integrals. BIT, 19(3):368–377, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=368.

Twizell:1979:EDM

[1019] E. H. Twizell. An explicit difference method for the wave equation with extended stability range. BIT, 19(3): 378–383, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=378.

Verwer:1979:AID

[1020] J. G. Verwer. The application of iterated defect correction to the LOD method for parabolic equations. BIT, 19(3):384–394, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=384.

Warming:1979:ESA

[1021] R. F. Warming and Richard M.

Beam. An extension of A-stability to alternating direction implicit methods. BIT, 19(3):395–417, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=395.

Wong:1979:NIA

[1022] Jin Kue Wong. A new implementation of an algorithm for the optimal assignment problem: An improved version of Munkre’s algorithm. BIT, 19(3):418–424, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=418.

Anonymous:1979:Pc

[1023] Anonymous. Problems. BIT, 19(3): 425–426, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=425.

Anonymous:1979:NBJc

[1024] Anonymous. New books and journals. BIT, 19(3):427, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=3&spage=427.

Anonymous:1979:Nc

[1025] Anonymous. Notes. BIT, 19(3):427, September 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=3&spage=427.

Anonymous:1979:Cc

[1026] Anonymous. Contents. BIT, 19 (3):428, 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bond:1979:BMN

[1027] J. E. Bond and J. R. Cash. A block method for the numerical integration of stiff systems of ordinary differential equations. BIT, 19(4):429–447, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=429.

StClair:1979:ALS

[1028] D. C. St. Clair and A. K. Rigler.

An algorithm for least squares analysis of spectroscopic data. BIT, 19 (4):448–456, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=448.

Elliott:1979:CHQ

[1029] David Elliott. On the convergence of Hunter’s quadrature rule for Cauchy principal value integrals. BIT, 19(4):457–462, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=457.

Ernvall:1979:CSS

[1030] J. Ernvall and O. Nevalainen. Compact storage schemes for formatted files by spanning trees. BIT, 19(4):463–475, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=463.

McKee:1979:CAL

[1031] S. McKee and N. Pitcher. On the convergence of advanced linear multistep methods. BIT, 19(4):476–481, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=476.

Kumar:1979:MNS

[1032] Sheo Kumar. On a method of Noble for second kind Volterra integral equations. BIT, 19(4):482–488, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=482.

Patricio:1979:NMS

[1033] F. Patrício. A numerical method for solving initial-value-problems with spline functions. BIT, 19(4):489–494, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=489.

Shampine:1979:EIF

[1034] L. F. Shampine. Evaluation of implicit formulas for the solution of ODES. BIT, 19(4):495–502, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=495.

Thomsen:1979:TPF

[1035] Per Grove Thomsen and Zahari Zlatev. Two-parameter families of predictor-corrector methods for the solution of ordinary differential equations. BIT, 19(4):503–517, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=503.

Wikstrom:1979:OST

[1036] Å. Wikström. Optimal search trees and length restricted codes. BIT, 19 (4):518–524, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=518.

Williams:1979:SRC

[1037] H. C. Williams and B. Schmid. Some remarks concerning the M.I.T. publickey cryptosystem. BIT, 19(4):525–538, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=525.

Dokken:1979:DDF

[1038] T. Dokken and T. Lyche. A divided difference formula for the error in Hermite interpolation. BIT, 19(4):539–540, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=539.

Hairer:1979:NSR

[1039] E. Hairer. Nonlinear stability of RAT, an explicit rational Runge– Kutta method. BIT, 19(4):540–542, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=540.

Iserles:1979:NPA

[1040] Arieh Iserles. A note on Padé approximations and generalized hypergeometric functions. BIT, 19(4):543–545, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=543.

Pollard:1979:SPR

[1041] J. M. Pollard. On not storing the path of a random walk. BIT, 19(4):545–548, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=545.

Smith:1979:NAC

[1042] H. V. Smith. The numerical approximation of a class of infinite integrals. BIT, 19(4):549–551, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=549.

Anonymous:1979:Pd

[1043] Anonymous. Problems. BIT, 19(4):

552–553, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=552.

Anonymous:1979:NBJd

[1044] Anonymous. New books and journals. BIT, 19(4):554, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=19&issue=4&spage=554.

Anonymous:1979:Nd

[1045] Anonymous. Notes. BIT, 19(4):

555–556, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=555.

Anonymous:1979:A

[1046] Anonymous. Acknowledgements. BIT, 19(4):556, December 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=19&issue=4&spage=556.

Anonymous:1979:I

[1047] Anonymous. Index. BIT, 19(4): 557–559, 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1979:Cd

[1048] Anonymous. Contents. BIT, 19 (4):560, 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Akl:1980:NAG

[1049] Selim G. Akl. A new algorithm for generating derangements. BIT, 20(1): 2–7, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=2.

Chan:1980:LTI

[1050] W. M. Chan and Alan George. A linear time implementation of the reverse Cuthill–McKee algorithm. BIT, 20(1): 8–14, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=8.

Jarvinen:1980:SPJ

[1051] Pertti Järvinen. On structuring problems of job design met in the development and maintenance of information systems. BIT, 20(1):15– 24, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=15.

Larson:1980:ARH

[1052] Per-Åke Larson. Analysis of repeated hashing. BIT, 20(1):25–32, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=25.

vanBokhoven:1980:EHO

[1053] W. M. G. van Bokhoven. Efficient higher order implicit one-step methods for integration of stiff differential equations. BIT, 20(1):34– 43, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=34.

Cash:1980:HOM

[1054] J. R. Cash and D. R. Moore. A high order method for the numerical solution of two-point boundary value problems. BIT, 20(1):44–52, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=44.

Dey:1980:NDE

[1055] Suhrit K. Dey. Nonlinear discretization errors in partial difference equations. BIT, 20(1):53–57, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=53.

Doedel:1980:SST

[1056] Eusebius J. Doedel. Some stability theorems for finite difference collocation methods on nonuniform meshes. BIT, 20(1):58–66, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=58.

Fox:1980:SEI

[1057] L. Fox and M. R. Valenca. Some experiments with interval methods for twopoint boundary value problems in ordinary differential equations. BIT, 20(1): 67–82, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=67.

Ioakimidis:1980:CTD

[1058] N. I. Ioakimidis and P. S. Theocaris. On convergence of two direct methods for solution of Cauchytype singular integral equations of the first kind. BIT, 20(1):83–87, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=83.

Ruhe:1980:RBJ

[1059] Axel Ruhe. The relation between the Jacobi algorithm and inverse iteration and a Jacobi algorithm based on elementary reflections. BIT, 20(1):88– 96, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=88.

Sanz-Serna:1980:SAB

[1060] J. M. Sanz-Serna. Some aspects of the boundary locus method. BIT, 20(1):97–101, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=97.

Wanner:1980:CSI

[1061] G. Wanner. On the choice of γ for singly-implicit RK or Rosenbrock methods. BIT, 20(1):102– 106, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=102.

Chawla:1980:NMC

[1062] M. M. Chawla and C. P. Katti. On Noumerov’s method for computing eigenvalues. BIT, 20(1):107– 109, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=107.

Hadjidimos:1980:NSP

[1063] A. Hadjidimos. A note on the simplification of p-parametric E.A.D.I. schemes for elliptic problems of higher order. BIT, 20(1):110–114, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=110.

Head:1980:MM

[1064] A. K. Head. Multiplication modulo n. BIT, 20(1):115–116, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=20&
                                        issue=1&spage=115.

Phillips:1980:SDI

[1065] G. M. Phillips. Seventh degree integration rules for the sphere. BIT, 20(1):117–119, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=117.

Shniad:1980:GEE

[1066] H. Shniad. Global error estimation for the implicit trapezoidal rule. BIT, 20(1):120–121, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=120.

Storoy:1980:ALD

[1067] Sverre Storøy. An algorithm for least distance programming. BIT, 20 (1):122–123, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=122.

Anonymous:1980:Pa

[1068] Anonymous. Problems. BIT, 20 (1):124–125, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=1&spage=124.

Anonymous:1980:NBJa

[1069] Anonymous. New books and journals. BIT, 20(1):126, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=20&
issue=1&spage=126.

Anonymous:1980:Na

[1070] Anonymous. Notes. BIT, 20(1): 127, March 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=1&spage=127.

Anonymous:1980:C

[1071] Anonymous. Contents. BIT, 20 (1):128, 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Erkio:1980:HAW

[1072] Hannu Erkiö. A heuristic approximation of the worst case of Shellsort. BIT, 20(2):130–136, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=2&spage=130.

Gjessing:1980:MAC

[1073] Stein Gjessing. Monitors with arrays of condition variables and proof rules handling local quantities. BIT, 20(2): 137–144, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=2&spage=137.

Hakola:1980:DWS

[1074] Juha Hakola and Ari Heiskanen. On the distribution of wasted space at the end of file blocks. BIT, 20(2):145– 156, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=145.

Kemp:1980:NSS

[1075] R. Kemp. A note on the stack size of regularly distributed binary trees. BIT, 20(2):157–162, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=157.

Wilander:1980:IPS

[1076] Jerker Wilander. An interactive programming system for Pascal. BIT, 20(2):163–174, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=20&
issue=2&spage=163.

Brent:1980:IMC

[1077] Richard P. Brent. An improved Monte Carlo factorization algorithm. BIT, 20(2):176–184, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=176.

Burrage:1980:NSG

[1078] Kevin Burrage and J. C. Butcher. Nonlinear stability of a general class of differential equation methods. BIT, 20(2): 185–203, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=185.

Dougalis:1980:SRC

[1079] Vassilios A. Dougalis and Steven M. Serbin. Some remarks on a class of rational approximations to the cosine. BIT, 20(2):204–211, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=204.

Hoskins:1980:SBS

[1080] W. D. Hoskins, G. M. Pathan, and D. J. Walton. Solution of bilinear systems arising from high order discretizations of Poisson-type equations. BIT, 20(2):212–214, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=212.

Kramarz:1980:SCM

[1081] L. Kramarz. Stability of collocation methods for the numerical solution of y′′ = f(x, y). BIT, 20(2):215– 222, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=215.

Tewarson:1980:USN

[1082] R. P. Tewarson. On the use of splines for the numerical solution of nonlinear two-point boundary value problems. BIT, 20(2):223– 232, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=223.

Wang:1980:VFF

[1083] H. H. Wang. On vectorizing the fast Fourier transform. BIT, 20(2):233– 243, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=233.

Degano:1980:NPB

[1084] P. Degano, G. Pacini, and F. Turini.

A note on the power of backtracking. BIT, 20(2):244–246, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=20&
issue=2&spage=244.

Bareiss:1980:RED

[1085] E. H. Bareiss and J. L. Barlow.

Roundoff error distribution in fixed point multiplication. BIT, 20(2):247– 250, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=247.

Caldwell:1980:AEL

[1086] James Caldwell. An application of extrapolation to the limit. BIT, 20(2): 251–253, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=2&spage=251.

Hairer:1980:HPO

[1087] E. Hairer. Highest possible order of algebraically stable diagonally implicit Runge–Kutta methods. BIT, 20(2): 254–256, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=2&spage=254.

Laurie:1980:EIW

[1088] D. P. Laurie. Efficient implementation of Wilson’s algorithm for factorizing a self-reciprocal polynomial. BIT, 20(2): 257–259, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=2&spage=257.

Gribble:1980:CFP

[1089] J. Gribble. A correction to: “Further properties of inner product quadrature formulas” [BIT 17(4), 1977, pp. 392– 408; MR 58 #24873]. BIT, 20(2): 260, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&

volume=20&issue=2&spage=260. See

[876].

Anonymous:1980:Pb

[1090] Anonymous. Problems. BIT, 20(2):

261–262, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=261.

Anonymous:1980:Nb

[1091] Anonymous. Notes. BIT, 20(2):

263, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=2&spage=263.

Anonymous:1980:NBJb

[1092] Anonymous. New books and journals. BIT, 20(2):263, June 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=20&
issue=2&spage=263.

Anonymous:1980:AAb

[1093] Anonymous. Addresses of the authors. BIT, 20(2):264, 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Arnborg:1980:SQL

[1094] Stefan Arnborg. A simple query language based on set algebra. BIT, 20 (3):266–278, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=265.

Davies:1980:FPR

[1095] D. Julian M. Davies. The fifty percent rule revisited. BIT, 20(3):279–288, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=279.

Hemker:1980:SAM

[1096] P. W. Hemker. On the structure of an adaptive multi-level algorithm. BIT, 20(3):289–301, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=289.

Ottmann:1980:HOA

[1097] Th. Ottmann and W. Stucky. Higher order analysis of random 1 − 2 brother trees. BIT, 20(3):302–314, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=302.

Burrage:1980:NLSb

[1098] K. Burrage. Non-linear stability of multivalue multiderivative methods. BIT, 20(3):315–325, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=315.

Burrage:1980:NSM

[1099] K. Burrage. Nonlinear stability of multivalue multiderivative methods. BIT, 20(3):316–325, 1980. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Burrage:1980:ISI

[1100] K. Burrage, J. C. Butcher, and F. H. Chipman. An implementation of singly-implicit Runge–Kutta methods. BIT, 20(3):326–340, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=326.

Epton:1980:MSA

[1101] Michael A. Epton. Methods for the solution of AXD − BXC = E and its application in the numerical solution of implicit ordinary differential equations. BIT, 20(3):341–345, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=341.

Kumar:1980:MIM

[1102] Sheo Kumar. On modified increment methods of Garey for nonlinear second kind Volterra integral equations. BIT, 20(3):346–350, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=346.

Salzer:1980:NTD

[1103] Herbert E. Salzer. Note on trigonometric divided differences. BIT, 20(3): 351–355, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=351.

Skelboe:1980:ICL

[1104] Stig Skelboe. Implementation of Chebyshevian linear multistep formulas. BIT, 20(3):356–366, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=356.

Wanner:1980:CAS

[1105] G. Wanner. Characterization of all A-stable methods of order 2m − 4. BIT, 20(3):367–374, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=367.

vanderHouwen:1980:CSR

[1106] P. J. van der Houwen. Convergence and stability results in Runge–Kutta type methods for Volterra integral equations of the second kind. BIT, 20(3):375–377, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=375.

Loizou:1980:GAS

[1107] G. Loizou. The generalized accelerated symmetric overrelaxation method. BIT, 20(3):378–381, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=378.

Meek:1980:SNL

[1108] D. Meek. Some new linear relations for even degree polynomial splines on a uniform mesh. BIT, 20(3):382–384, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=382.

Ruffhead:1980:CCO

[1109] A. Ruffhead and J. Oliver. A characterisation of certain optimal collocation points for numerical differentiation. BIT, 20(3):385–388, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=3&spage=385.

Anonymous:1980:Pc

[1110] Anonymous. Problems. BIT, 20(3): 389–390, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=389.

Anonymous:1980:NBJc

[1111] Anonymous. New books and journals. BIT, 20(3):391, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=391.

Anonymous:1980:Nc

[1112] Anonymous. Notes. BIT, 20(3):391, September 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=3&spage=391.

Anonymous:1980:AAc

[1113] Anonymous. Addresses of the authors. BIT, 20(3):392, 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jorgensen:1980:MMR

[1114] Anker Helms Jørgensen. A methodology for measuring the readability and modifiability of computer programs. BIT, 20(4):394–405, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=393.

Lloyd:1980:OPM

[1115] J. W. Lloyd. Optimal partial-match retrieval. BIT, 20(4):406–413, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=406.

Naur:1980:IED

[1116] Peter Naur. Impressions of the early days of programming. BIT, 20(4): 414–425, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=414.

Six:1980:RIP

[1117] H.-W. Six and D. Wood. The rectangle intersection problem revisited. BIT, 20(4):426–433, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=426.

Stromberg:1980:TPD

[1118] Dan Strömberg and Peter Fritzson. Transfer of programs from development to runtime environments. BIT, 20(4):434–442, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=434.

Brezinski:1980:MNA

[1119] C. Brezinski. The Mühlbach–Neville– Aitken algorithm and some extensions. BIT, 20(4):443–451, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=443.

Burrage:1980:SAO

[1120] Kevin Burrage and Pamela Moss. Simplifying assumptions for the order of partitioned multivalue methods. BIT, 20(4):452–465, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=452.

Gribble:1980:IIP

[1121] J. D. Gribble. Interpolatory inner product quadrature formulas. BIT, 20(4): 466–474, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=466.

Hackl:1980:ESS

[1122] J. Hackl, Hj. Wacker, and W. Zulehner. An efficient step size control for continuation methods. BIT, 20(4):475–485, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=475.

Lindberg:1980:EEI

[1123] Bengt Lindberg. Error estimation and iterative improvement for discretization algorithms. BIT, 20(4):486–500, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=486.

Martinez:1980:SNS

[1124] José Mario Martínez. Solving nonlinear simultaneous equations with a generalization of Brent’s method. BIT, 20(4):501–510, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=501.

Chawla:1980:NFO

[1125] M. M. Chawla and C. P. Katti. A new fourth order method for computing eigenvalues of two-point boundary value problems. BIT, 20(4):511–514, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=511. See

addendum [1320].

Girse:1980:NCC

[1126] Robert D. Girse. The number of conjugacy classes of the alternating group. BIT, 20(4):515–517, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=515.

Herlestam:1980:NRP

[1127] Tore Herlestam. A note on Rabin’s probabilistic primality test. BIT, 20 (4):518–521, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=518.

Kundu:1980:DLS

[1128] Sukhamay Kundu. A Dijkstra-like shortest path algorithm for certain cases of negative arc lengths. BIT, 20 (4):522–524, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=522.

Light:1980:NGC

[1129] W. A. Light, J. Forrest, N. Hammond, and S. Roe. A note on Goldbach’s conjecture. BIT, 20(4):525, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=525.

Stewart:1980:BMI

[1130] G. W. Stewart. The behavior of a multiplicity independent rootfinding scheme in the presence of error. BIT, 20(4):526–528, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=526.

Anonymous:1980:Pd

[1131] Anonymous. Problems. BIT, 20(4):

529–530, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=529.

Anonymous:1980:NBJd

[1132] Anonymous. New books and journals. BIT, 20(4):530–531, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=530.

Anonymous:1980:Nd

[1133] Anonymous. Notes. BIT, 20(4):

531–532, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=20&issue=4&spage=531.

Anonymous:1980:A

[1134] Anonymous. Acknowledgements. BIT, 20(4):532, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=20&issue=4&spage=532.

Anonymous:1980:I

[1135] Anonymous. Index. BIT, 20(4): 533–535, 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1980:AAd

[1136] Anonymous. Addresses of the authors. BIT, 20(4):536, 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Erkio:1981:SSA

[1137] Hannu Erkiö. Speeding sort algorithms by special instructions. BIT, 21(1):2– 19, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=2.

Itoga:1981:SMP

[1138] Stephen Y. Itoga. The string merging problem. BIT, 21(1):20–30, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=20.

Knuth:1981:VLL

[1139] Donald E. Knuth. Verification of link-level protocols. BIT, 21(1):31– 36, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=31.

Lundberg:1981:CIM

[1140] Bengt Lundberg. On consistency of information models. BIT, 21(1):37– 45, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=37.

Nevalainen:1981:FMS

[1141] O. Nevalainen, J. Ernvall, and J. Katajainen. Finding minimal spanning trees in a Euclidean coordinate space. BIT, 21(1):46–54, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=46.

Bjorstad:1981:EAE

[1142] Petter E. Bjørstad, Germund Dahlquist, and Eric H. Grosse. Extrapolation of asymptotic expansions by a modified Aitken δ^{2}-formula. BIT, 21(1): 56–65, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=56.

Dekker:1981:SLM

[1143] K. Dekker. Stability of linear multistep methods on the imaginary axis. BIT, 21(1):66–79, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=66.

Iserles:1981:TSN

[1144] A. Iserles. Two-step numerical methods for parabolic differential equations. BIT, 21(1):80–96, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=80.

Kielbasinski:1981:IRL

[1145] Andrzej Kiełbasiński. Iterative refinement for linear systems in variableprecision arithmetic. BIT, 21(1):97– 103, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=97.

Krishnamurthy:1981:PWA

[1146] E. V. Krishnamurthy and H. Venkateswaran. A parallel Wilf algorithm for complex zeros of a polynomial. BIT, 21(1):104– 111, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=104. See

erratum [1202].

Lane:1981:BP

[1147] Jeffrey M. Lane and R. F. Riesenfeld. Bounds on a polynomial. BIT, 21(1):112–117, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=112.

Bohman:1981:NIW

[1148] Jan Bohman and Carl-Erik Fröberg.

Numerical investigation of Waring’s problem for cubes. BIT, 21(1):118– 122, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=118.

Kumar:1981:RRS

[1149] Sheo Kumar. A recurrence relation for solution of singular Volterra integral equations using Chebyshev polynomials. BIT, 21(1):123–125, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=123.

Phillips:1981:SDI

[1150] G. M. Phillips. Seventh degree integration rules for R^{3}. BIT, 21(1):126– 128, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=126.

Popovski:1981:NKF

[1151] D. B. Popovski. A note on King’s fifth-order family of methods for solving equations. BIT, 21(1):129–130, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=129.

Wu:1981:NTE

[1152] Wen Da Wu. Neville type extrapolation scheme for a special expansion. BIT, 21(1):131–135, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=131.

Anonymous:1981:Pa

[1153] Anonymous. Problems. BIT, 21 (1):136–137, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=1&spage=136.

Anonymous:1981:NBJa

[1154] Anonymous. New books and journals. BIT, 21(1):138, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=21&
                                        issue=1&spage=138.

Anonymous:1981:Na

[1155] Anonymous. Notes. BIT, 21(1):

139, March 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=1&spage=139.

Anonymous:1981:AAa

[1156] Anonymous. Addresses of the authors. BIT, 21(1):140, 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Brock:1981:DSM

[1157] Harvey K. Brock, Barbara J. Brooks, and Francis Sullivan. Diamond: a sorting method for vector machines. BIT, 21(2):142–152, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=21&
issue=2&spage=142.

Fenner:1981:NTA

[1158] T. I. Fenner and G. Loizou. A note on traversal algorithms for triply linked binary trees. BIT, 21(2):153– 156, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=153.

Wong:1981:SSP

[1159] Jin Kue Wong. Some simple in-place merging algorithms. BIT, 21(2):157– 166, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=157.

Bui:1981:CAR

[1160] T. D. Bui and S. W. H. Poon. On the computational aspects of Rosenbrock procedures with built-in error estimates for stiff systems. BIT, 21(2): 168–174, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=168.

Butcher:1981:GSI

[1161] J. C. Butcher. A generalization of singly-implicit methods. BIT, 21(2): 175–189, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=175.

Chawla:1981:TSF

[1162] M. M. Chawla. Two-step fourth order P -stable methods for second order differential equations. BIT, 21(2):190– 193, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=190.

Gregory:1981:EFC

[1163] R. T. Gregory. Error-free computation with rational numbers. BIT, 21(2):194– 202, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=194.

Hansen:1981:BSS

[1164] Eldon Hansen and Saumyendra Sengupta. Bounding solutions of systems of equations using interval analysis. BIT, 21(2):203–211, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=203.

Paget:1981:QRF

[1165] D. F. Paget. A quadrature rule for finite-part integrals. BIT, 21(2):212– 220, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=212.

Skelboe:1981:BDF

[1166] Stig Skelboe and Brian Christensen.

Backward differentiation formulas with extended regions of absolute stability. BIT, 21(2):221–231, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=2&spage=221.

Wolkenfelt:1981:RQM

[1167] P. H. M. Wolkenfelt. Reducible quadrature methods for Volterra integral equations of the first kind. BIT, 21(2): 232–241, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=2&spage=232.

Anonymous:1981:Pb

[1168] Anonymous. Problems. BIT, 21(2): 242–243, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=2&spage=242.

Anonymous:1981:NBJb

[1169] Anonymous. New books and journals. BIT, 21(2):243, June 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=21&
                                        issue=2&spage=243.

Anonymous:1981:AAb

[1170] Anonymous. Addresses of the authors. BIT, 21(2):244, 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Darondeau:1981:TML

[1171] Ph. Darondeau, P. Le Guernic, and M. Raynal. Types in a mixed language system. BIT, 21(3):246–254, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=3&spage=245.

Doberkat:1981:INE

[1172] Ernst-Erich Doberkat. Inserting a new element into a heap. BIT, 21(3): 255–269, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=255.

Kristensen:1981:DCB

[1173] Bent Bruun Kristensen and Ole Lehrmann Madsen. Diagnostics on LALR(k) conflicts based on a method for LR(k) testing. BIT, 21(3):270–293, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=270.

Lam:1981:QPM

[1174] C. M. Lam and K. T. Fung. An quadratic programming model for optimal data distribution. BIT, 21(3): 294–304, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=294.

Sprugnoli:1981:ABT

[1175] Renzo Sprugnoli. On the allocation of binary trees to secondary storage. BIT, 21(3):305–316, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=305.

Elliott:1981:NSI

[1176] Charles M. Elliott and Sean Mc-Kee. On the numerical solution of an integro-differential equation arising from wave-power hydraulics. BIT, 21(3):317–325, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=3&spage=317.

Herlestam:1981:CL

[1177] Tore Herlestam and Rolf Johannesson. On computing logarithms over GF(2^{p}). BIT, 21(3):326–334, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=3&spage=326.

Kazakova:1981:CFD

[1178] N. G. Kazakova and D. D. Baĭnov. Convergence of finite difference methods for the initial value problem for functional-differential equations of neutral type. BIT, 21(3):335–341, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=3&spage=335.

Neittaanmaki:1981:MFE

[1179] P. Neittaanmäki and J. Saranen. A mixed finite element method for the heat flow problem. BIT, 21(3):342–346, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=342.

Schneider:1981:RAM

[1180] Norbert Schneider. Results about monotone convergence of Steffensenlike-methods. BIT, 21(3):347–354, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=347.

Verwer:1981:PVN

[1181] J. G. Verwer. On the practical value of the notion of BNstability. BIT, 21(3):355–361, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=355.

Watson:1981:GCM

[1182] G. A. Watson. Globally convergent methods for semi-infinite programming. BIT, 21(3):362–373, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=362.

Colbourn:1981:DHE

[1183] Charles J. Colbourn and Marlene J.

Colbourn. Deciding Hadamard equivalence of Hadamard matrices. BIT, 21(3):374–376, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=374.

Gerasoulis:1981:EAS

[1184] Apostolos Gerasoulis. On the existence of approximate solutions for singular integral equations of Cauchy type discretized by Gauss–Chebyshev quadrature formulae. BIT, 21(3):377–380, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=377.

Newbery:1981:CTI

[1185] A. C. R. Newbery. A consistency test for interpolators. BIT, 21(3):381–384, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=381.

Anonymous:1981:Pc

[1186] Anonymous. Problems. BIT, 21(3): 385–386, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=385.

Anonymous:1981:NBJc

[1187] Anonymous. New books and journals. BIT, 21(3):386, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=386.

Anonymous:1981:Nb

[1188] Anonymous. Notes. BIT, 21(3):387, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=387.

Anonymous:1981:AAc

[1189] Anonymous. Addresses of the authors. BIT, 21(3):388, September 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=3&spage=388.

Mullin:1981:TCL

[1190] James K. Mullin. Tightly controlled linear hashing without separate overflow storage. BIT, 21(4):390–400, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=390.

Ouksel:1981:MTA

[1191] Mohamed Ouksel and Peter Scheuermann. Multidimensional B-trees:

Analysis of dynamic behavior. BIT, 21 (4):401–418, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=401.

Tamminen:1981:OPE

[1192] Markku Tamminen. Order preserving extendible hashing and bucket tries. BIT, 21(4):419–435, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=419.

Wallis:1981:SPP

[1193] Peter J. L. Wallis. Some primitives for the portable programming of array and vector processors. BIT, 21(4):436–448, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=436.

Cash:1981:NEF

[1194] J. R. Cash. A note on the exponential fitting of blended, extended linear multistep methods. BIT, 21(4):450–454, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=450.

Chawla:1981:IPA

[1195] M. M. Chawla and S. R. Sharma. Intervals of periodicity and absolute stability of explicit Nyström methods for y′′ = f(x, y). BIT, 21(4):455–464, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=455.

Haavie:1981:RUT

[1196] Tore Håvie. Remarks on a unified theory for classical and generalized interpolation and extrapolation. BIT, 21(4):465–474, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=465.

Scott:1981:SSS

[1197] D. S. Scott. Solving sparse symmetric definite quadratic λ-matrix problems. BIT, 21(4):475–480, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=475.

Smith:1981:GEB

[1198] H. V. Smith. Global error bounds for Gauss–Gegenbauer quadrature. BIT, 21(4):481–490, December 1981. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/

                                        openurl.asp?genre=article&issn=
                                        0006-3835&volume=21&issue=4&spage=
                                        481.

Sottas:1981:QFP

[1199] G. Sottas. Quadrature formulas with positive weights. BIT, 21(4):491–504, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=4&spage=491.

Tang:1981:BDV

[1200] W. P. Tang and G. H. Golub. The block decomposition of a Vandermonde matrix and its applications. BIT, 21(4):505–517, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=4&spage=505.

Twizell:1981:OSM

[1201] E. H. Twizell and A. Q. M. Khaliq. One-step multiderivative methods for first order ordinary differential equations. BIT, 21(4):518–527, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=4&spage=518.

Krishnamurthy:1981:EPW

[1202] E. V. Krishnamurthy and H. Venkateswaran. Errata: “A parallel Wilf algorithm for complex zeros of a polynomial” [BIT 21(1), 1981, pp.

104–111]. BIT, 21(4):528, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=528. See

[1146].

Anonymous:1981:Pd

[1203] Anonymous. Problems. BIT, 21(4):

529–530, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=529.

Anonymous:1981:NBJd

[1204] Anonymous. New books and journals. BIT, 21(4):530, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=530.

Anonymous:1981:Nd

[1205] Anonymous. Notes. BIT, 21(4):531, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=21&issue=4&spage=531.

Anonymous:1981:A

[1206] Anonymous. Acknowledgements. BIT, 21(4):532, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=21&issue=4&spage=532.

Anonymous:1981:I

[1207] Anonymous. Index. BIT, 21(4): 533–535, 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1981:AAd

[1208] Anonymous. Addresses of the authors. BIT, 21(4):536, 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Yeh:1982:IPA

[1209] Da Shing Yeh. Improved planarity algorithms. BIT, 22(1):2– 16, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=1&spage=2.

Navlakha:1982:NPT

[1210] Jainendra Navlakha. A new proof technique to establish equivalence of the original and the generated λ-free CFG with linear increase in size. BIT, 22(1): 17–26, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=1&spage=17.

Tamminen:1982:ECM

[1211] Markku Tamminen. The extendible cell method for closest point problems. BIT, 22(1):27–41, March 1982. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=22&
issue=1&spage=27.

Topor:1982:FSE

[1212] Rodney W. Topor. Fundamental solutions of the eight queens problem. BIT, 22(1):42–52, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=22&
issue=1&spage=42.

Ljung:1982:FNS

[1213] Stefan Ljung and Lennart Ljung.

Fast numerical solution of Fredholm integral equations with stationary kernels. BIT, 22(1):54–72, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=1&spage=54.

Martins:1982:GDD

[1214] M. Madalena Martins. Generalized diagonal dominance in connection with the accelerated overrelaxation (AOR) method. BIT, 22(1):73– 78, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=1&spage=73.

Mattheij:1982:SCS

[1215] R. M. M. Mattheij. Stable computation of solutions of unstable linear initial value recursions. BIT, 22(1):79– 93, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=1&spage=79.

Nelson:1982:ACN

[1216] Peter R. Nelson. An approximation for the complex normal probability integral. BIT, 22(1):94–100, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=1&spage=94.

Rack:1982:NSE

[1217] Heinz-Joachim Rack and Manfred Reimer. The numerical stability of evaluation schemes for polynomials based on the Lagrange interpolation form. BIT, 22(1):101–107, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=1&spage=101.

Ypma:1982:AIC

[1218] T. J. Ypma. Affine invariant convergence results for Newton’s method. BIT, 22(1):108–118, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=22&
issue=1&spage=108.

Dunham:1982:BRC

[1219] Charles B. Dunham. Biased rational Chebyshev approximation. BIT, 22(1):119–122, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=1&spage=119.

Lazos:1982:NCU

[1220] C. Lazos. A note on the CPU usage of a computer system with many I/O channels. BIT, 22(1):123– 124, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=1&spage=123.

Anonymous:1982:Pa

[1221] Anonymous. Problems. BIT, 22 (1):125–126, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=1&spage=125.

Anonymous:1982:NBJa

[1222] Anonymous. New books and journals. BIT, 22(1):126, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=22&
issue=1&spage=126.

Anonymous:1982:Na

[1223] Anonymous. Notes. BIT, 22(1): 127, March 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=1&spage=127.

Anonymous:1982:AAa

[1224] Anonymous. Addresses of the authors. BIT, 22(1):128, 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Akl:1982:CTP

[1225] Selim G. Akl. A constant-time parallel algorithm for computing convex hulls. BIT, 22(2):130–134, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=2&spage=129.

Allison:1982:UEH

[1226] D. C. S. Allison and M. T. Noga. USORT: an efficient hybrid of distributive partitioning sorting. BIT, 22(2): 135–139, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=2&spage=135.

Gunderson:1982:CDF

[1227] Robert W. Gunderson. Choosing the r-dimension for the FCV family of clustering algorithms. BIT, 22(2):140– 149, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=140.

Lloyd:1982:PMR

[1228] John W. Lloyd and K. Ramamohanarao. Partial-match retrieval for dynamic files. BIT, 22(2):150– 168, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=150.

Soffa:1982:CDN

[1229] Mary Lou Soffa. Control discipline necessity: making the language as general as the implementation. BIT, 22(2): 169–182, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=169.

Cash:1982:HOM

[1230] J. R. Cash and A. Singhal. High order methods for the numerical solution of two-point boundary value problems. BIT, 22(2):183– 199, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=183.

Gerasoulis:1982:SIE

[1231] Apostolos Gerasoulis. Singular integral equations — the convergence of the Nyström interpolant of the Gauss– Chebyshev method. BIT, 22(2):200– 210, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=200.

Hairer:1982:SSI

[1232] E. Hairer, G. Bader, and Ch. Lubich. On the stability of semiimplicit methods for ordinary differential equations. BIT, 22(2):211– 232, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=211.

Rall:1982:SFS

[1233] L. B. Rall. Solution of finite systems of equations by interval iteration. BIT, 22(2):233–251, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=233.

Griewank:1982:SPD

[1234] Andreas Griewank. A short proof of the Dennis–Schnabel theorem. BIT, 22(2): 252–256, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=252.

Mjelde:1982:IAD

[1235] K. M. Mjelde. Incremental allocation of a discrete and convex resource to concave activities. BIT, 22(2):257– 260, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=257.

Usmani:1982:NQS

[1236] Riaz A. Usmani and Manabu Sakai.

A note on quadratic spline interpolation at mid-points. BIT, 22(2):261– 267, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=261.

Anonymous:1982:Pb

[1237] Anonymous. Problems. BIT, 22(2):

268–269, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=2&spage=268.

Anonymous:1982:NBJb

[1238] Anonymous. New books and journals. BIT, 22(2):270, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=22&
issue=2&spage=270.

Anonymous:1982:Nb

[1239] Anonymous. Notes. BIT, 22(2): 271, June 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=2&spage=271.

Anonymous:1982:AAb

[1240] Anonymous. Addresses of the authors. BIT, 22(2):272, 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Edelsbrunner:1982:SLS

[1241] H. Edelsbrunner, H. A. Maurer, F. P. Preparata, A. L. Rosenberg, E. Welzl, and D. Wood. Stabbing line segments. BIT, 22(3):274–281, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=274.

Langston:1982:IIS

[1242] Michael A. Langston. Improved 0/1interchange scheduling. BIT, 22(3): 282–290, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=282.

Lundberg:1982:AE

[1243] Bengt Lundberg. An axiomatization of events. BIT, 22(3):291–299, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=291.

Perko:1982:PRN

[1244] A. Perko. On the path representation of networks. BIT, 22(3):300–302, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=300.

Soisalon-Soininen:1982:TSG

[1245] Eljas Soisalon-Soininen. Translations on a subclass of LR(k) grammars. BIT, 22(3):303–312, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=303.

McKee:1982:TSE

[1246] S. McKee and N. Pitcher. Twosided error bounds for discretisation methods in ordinary differential equations. BIT, 22(3):314–330, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=314. See

erratum [1308].

Neuman:1982:CIS

[1247] Edward Neuman. Convex interpolating splines of arbitrary degree.

II. BIT, 22(3):331–338, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=331.

Sottas:1982:NPW

[1248] G. Sottas and G. Wanner. The number of positive weights of a quadrature formula. BIT, 22(3):339–352, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=339.

Tewarson:1982:IAF

[1249] R. P. Tewarson and S. Gupta. Improving the accuracy of finite difference methods for solving boundary value ordinary differential equations. BIT, 22(3):353–360, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=353.

Wolfbrandt:1982:DAS

[1250] Arne Wolfbrandt. Dynamic adaptive selection of integration algorithms when solving ODEs. BIT, 22(3): 361–367, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=361.

Ypma:1982:FPT

[1251] T. J. Ypma. Following paths through turning points. BIT, 22(3):368–383, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=368.

Costabile:1982:TSF

[1252] F. Costabile and C. Costabile. Twostep fourth order P -stable methods for second order differential equations. BIT, 22(3):384–386, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=384.

Gautschi:1982:POR

[1253] Walter Gautschi. Polynomials orthogonal with respect to the reciprocal gamma function. BIT, 22(3):387–389, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=3&spage=387.

Sathiaraj:1982:MSU

[1254] D. Daniel Sathiaraj and R. Sankar.

The method of successive updated iterated defect correction and its application to second kind Volterra integral equations. BIT, 22(3):390–394, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=390.

Smith:1982:GEB

[1255] H. V. Smith. Global error bounds for the Clenshaw–Curtis quadrature formula. BIT, 22(3):395–398, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=395.

Anonymous:1982:Pc

[1256] Anonymous. Problems. BIT, 22(3): 399–400, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=399.

Anonymous:1982:NBJc

[1257] Anonymous. New books and journals. BIT, 22(3):400, September 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=400.

Kirkerud:1982:CHC

[1258] Bjørn Kirkerud. Completeness of Hoare-calculi revisited. BIT, 22(4): 401–418, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=401.

Lofgren:1982:AMB

[1259] Mats Löfgren and Boris Magnusson.

Access mechanisms in block structured environments. BIT, 22(4):419–436, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=419.

Naur:1982:FPD

[1260] Peter Naur. Formalization in program development. BIT, 22(4):437–453, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=437.

Thanh:1982:IBM

[1261] Mai Thanh and T. D. Bui. An improvement of the binary merge algorithm. BIT, 22(4):454–462, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=454.

Bickart:1982:SSI

[1262] Theodore A. Bickart. P -stable and P [α, β]-stable integration/interpolation methods in the solution of retarded differential-difference equations. BIT, 22(4):463–476, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=463.

Bunse:1982:BPT

[1263] Wolfgang Bunse. Blending producttype quadrature rules. BIT, 22(4): 477–486, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=477.

Elden:1982:WPG

[1264] Lars Eldén. A weighted pseudoinverse, generalized singular values, and constrained least squares problems. BIT, 22(4):487–502, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=487.

Kacewicz:1982:OEA

[1265] Bolesław Kacewicz. On the optimal error of algorithms for solving a scalar autonomous ODE. BIT, 22(4):503–518, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=503.

Lindskog:1982:CFM

[1266] Gunhild Lindskog. The continued fraction methods for the solution of systems of linear equations. BIT, 22(4):519–527, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=519.

Butcher:1982:SPC

[1267] J. C. Butcher. A short proof concerning B-stability. BIT, 22(4):528–529, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=528.

Chen:1982:UBG

[1268] T. H. Charles Chen. An upper bound for the Gauss–Legendre quadrature error for analytic functions. BIT, 22(4):530–532, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=530.

VanderCruyssen:1982:NRG

[1269] P. Van der Cruyssen. Nonhomogeneous recursions and generalised continued fractions. BIT, 22(4):533–537, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=22&issue=4&spage=533.

Weintraub:1982:PGP

[1270] S. Weintraub. A prime gap of 682 and a prime arithmetic sequence. BIT, 22(4):538, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=538. See

erratum [1338].

Anonymous:1982:Pd

[1271] Anonymous. Problems. BIT, 22(4):539, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=539.

Anonymous:1982:NBJd

[1272] Anonymous. New books and journals. BIT, 22(4):540–541, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=540.

Anonymous:1982:Nc

[1273] Anonymous. Notes. BIT, 22(4): 540–541, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=540.

Anonymous:1982:A

[1274] Anonymous. Acknowledgements. BIT, 22(4):542, December 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=22&issue=4&spage=542.

Anonymous:1982:I

[1275] Anonymous. Index 1982. BIT, 22 (4):543–544, 1982. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Katajainen:1983:WCM

[1276] Jyrki Katajainen. On the worst case of a minimal spanning tree algorithm for Euclidean space. BIT, 23(1):1– 8, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=1.

Kornerup:1983:MIH

[1277] Peter Kornerup and R. T. Gregory.

Mapping integers and Hensel codes onto Farey fractions. BIT, 23(1):9– 20, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=9.

Turini:1983:ACE

[1278] Franco Turini. Abstractions of control environments. BIT, 23(1):21– 35, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=21.

Yeh:1983:ISR

[1279] Da Shing Yeh. On incremental shift-reduce parsing. BIT, 23(1):36– 48, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=36.

Avdelas:1983:OCS

[1280] G. Avdelas, S. Galanis, and A. Hadjidimos. On the optimization of a class of second order iterative schemes. BIT, 23(1):50–64, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=50.

Borosh:1983:OMP

[1281] Itshak Borosh and Harald Niederreiter. Optimal multipliers for pseudorandom number generation by the linear congruential method. BIT, 23(1): 65–74, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=65.

Chakravarti:1983:SSS

[1282] P. C. Chakravarti and M. S. Kamel. Stiffly stable second derivative multistep methods with higher order and improved stability regions. BIT, 23(1): 75–83, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=75.

Crouzeix:1983:ESA

[1283] M. Crouzeix, W. H. Hundsdorfer, and M. N. Spijker. On the existence of solutions to the algebraic equations in implicit Runge– Kutta methods. BIT, 23(1):84–91, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=84.

Ioakimidis:1983:NIF

[1284] N. I. Ioakimidis. A natural interpolation formula for the numerical solution of singular integral equations with Hilbert kernel. BIT, 23(1):92– 104, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=92.

Loi:1983:AGR

[1285] S. L. Loi and A. W. McInnes. An algorithm for generalized rational interpolation. BIT, 23(1):105–117, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=105.

Han:1983:NSP

[1286] Tian Min Han. Numerical small parameter method for stiff ODE:s. BIT, 23(1):118–131, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=1&spage=118.

Leuze:1983:HLM

[1287] Michael R. Leuze. A hybrid Laguerre method. BIT, 23(1):132– 138, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=1&spage=132.

Rabinowitz:1983:OQF

[1288] Philip Rabinowitz. On an osculatory quadrature formula. BIT, 23(1):139– 140, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=1&spage=139.

Anonymous:1983:Pa

[1289] Anonymous. Problems. BIT, 23 (1):141–142, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=1&spage=141.

Anonymous:1983:E

[1290] Anonymous. Erratum. BIT, 23(1): 142, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=1&spage=142.

Anonymous:1983:Na

[1291] Anonymous. Notes. BIT, 23(1):143– 144, March 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=1&spage=143.

Johansen:1983:SAB

[1292] Gunner Helweg Johansen and Chr.

Gram. A simple algorithm for building the 3D convex hull. BIT, 23(2): 145–160, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=145.

Itoga:1983:PVS

[1293] Stephen Y. Itoga. A probabilistic version of the stable marriage problem. BIT, 23(2):161– 169, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=161.

Orenstein:1983:LSP

[1294] J. A. Orenstein, T. H. Merrett, and L. Devroye. Linear sorting with O(log n) processors. BIT, 23(2):170– 180, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=170.

Ukkonen:1983:ELB

[1295] Esko Ukkonen. Exponential lower bounds for some NP-complete problems in a restricted linear decision tree model. BIT, 23(2):181– 192, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=181.

Evans:1983:ISD

[1296] David J. Evans and Elias A. Lipitakis. Implicit semi-direct methods based on root-free sparse factorization procedures. BIT, 23(2):193– 208, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=193.

Evans:1983:ISM

[1297] David J. Evans and Elias A. Lipitakis. Implicit semidirect methods based on root-free sparse factorization procedures. BIT, 23(2):194–208, 1983. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gautschi:1983:HHC

[1298] Walter Gautschi. How and how not to check Gaussian quadrature formulae. BIT, 23(2):209– 216, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=209. See

correction [3576].

Kacewicz:1983:OEI

[1299] Bolesław Kacewicz. Optimality of Euler-integral information for solving a scalar autonomous ODE. BIT, 23(2): 217–230, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=217.

Layton:1983:FEM

[1300] William J. Layton. The finite element method for a degenerate hyperbolic partial differential equation. BIT, 23(2):231–238, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=231.

Sathiaraj:1983:MSE

[1301] D. Daniel Sathiaraj and R. Sankar.

The method of successive extrapolated iterated defect correction and its application to second kind Fredholm’s integral equations. BIT, 23(2):239– 247, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=239.

Zietak:1983:SNM

[1302] K. Zietak. The l_{p}-solution of the nonlinear matrix equation XY = A. BIT, 23(2):248–257, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=23&
                                        issue=2&spage=248.

Laurie:1983:SEE

[1303] D. P. Laurie. Sharper error estimates in adaptive quadrature. BIT, 23(2): 258–261, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=2&spage=258.

Scherer:1983:RTR

[1304] Rudolf Scherer and Helmut Türke. Reflected and transposed Runge– Kutta methods. BIT, 23(2):262– 266, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=2&spage=262.

Wallace:1983:DE

[1305] Stein W. Wallace. On degeneracy in enumeration. BIT, 23(2):267– 270, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=2&spage=267.

Anonymous:1983:Pb

[1306] Anonymous. Problems. BIT, 23(2): 271–272, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=271.

Anonymous:1983:BR

[1307] Anonymous. Books received. BIT, 23 (2):272, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=272.

McKee:1983:ETS

[1308] S. McKee and N. Pitcher. Erratum: “Two-sided error bounds for discretisation methods in ordinary differential equations” [BIT 22(3), 1982, pp. 314–330]. BIT, 23(2): 272, June 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=2&spage=272. See

[1246].

Burkhard:1983:IBI

[1309] Walter A. Burkhard. Interpolationbased index maintenance. BIT, 23(3): 274–294, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=274. Dis-

cusses construction of order-preserving hash functions for use in range queries.

Er:1983:ISG

[1310] M. C. Er. An iterative solution to the generalized Towers of Hanoi problem. BIT, 23(3):295–302, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=295.

LegindLarsen:1983:GDM

[1311] H. Legind Larsen. Generalized double modulo 11 check digit error detection. BIT, 23(3):303–307, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=303.

Yeh:1983:IEO

[1312] Da Shing Yeh. On incremental evaluation of ordered attributed grammars. BIT, 23(3):308–320, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=308.

Amini:1983:SAM

[1313] S. Amini. Stability analysis of methods employing reducible rules for Volterra integral equations. BIT, 23(3):322–328, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=322.

Bjorck:1983:BAP

[1314] Åke Björck. A block QR algorithm for partitioning stiff differential systems. BIT, 23(3):329–345, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=329.

Linnainmaa:1983:ELE

[1315] Seppo Linnainmaa. Error linearization as an effective tool for experimental analysis of the numerical stability of algorithms. BIT, 23(3):346–359, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=346.

Patricio:1983:CHF

[1316] F. Patrício. A class of hybrid formulae for the numerical integration of stiff systems. BIT, 23(3):360–369, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=360.

Piessens:1983:MCC

[1317] Robert Piessens and Maria Branders.

Modified Clenshaw–Curtis method for the computation of Bessel function integrals. BIT, 23(3):370–381, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=370.

Tewarson:1983:AII

[1318] R. P. Tewarson and N. S. Huslak. An adaptive implementation of interpolation methods for boundary value ordinary differential equations. BIT, 23(3):382–387, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=382.

Vanselow:1983:NSB

[1319] Reiner Vanselow. Nonlinear stability behaviour of linear multistep methods. BIT, 23(3):388–396, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=388.

Chawla:1983:ANF

[1320] M. M. Chawla and C. P. Katti. Addendum: “A new fourth order method for computing eigenvalues of two-point boundary value problems” [BIT 20(4), 1980, pp. 511–514; MR 82e:65084]. BIT, 23(3):397–398, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=397. See

[1125].

Sakai:1983:SNCa

[1321] Manabu Sakai and Riaz A. Usmani. Some new consistency relations connecting spline values and integrals of the spline. BIT, 23(3):399–402, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=399.

Storoy:1983:RVL

[1322] Sverre Storøy. Ranking of vertices in the linear fractional programming problem. BIT, 23(3):403–405, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=403.

Anonymous:1983:Pc

[1323] Anonymous. Problems. BIT, 23(3):

406–407, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=406.

Anonymous:1983:NBJ

[1324] Anonymous. New books and journals. BIT, 23(3):407, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=3&spage=407.

Anonymous:1983:Nb

[1325] Anonymous. Notes. BIT, 23(3):408, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=3&spage=408.

Alagar:1983:EAM

[1326] V. S. Alagar, T. D. Bui, and Mai Thanh. Efficient algorithms for merging. BIT, 23(4):409–428, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=409.

Er:1983:AGT

[1327] M. C. Er. An analysis of the generalized Towers of Hanoi problem. BIT, 23(4):429–435, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=429.

Huang:1983:BST

[1328] Shou Hsuan Huang and C. K. Wong. Binary search trees with limited rotation. BIT, 23(4):436–455, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=436.

Nicholl:1983:CHS

[1329] T. M. Nicholl, D. T. Lee, Y. Z. Liao, and C. K. Wong. On the X-Y convex hull of a set of X-Y polygons. BIT, 23(4):456–471, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=456.

Raman:1983:PAF

[1330] Vasudevan Raman and S. Sitharama Iyengar. Properties and applications of forests of quadtrees for pictorial data representation. BIT, 23(4):472–486, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=472.

Dagnino:1983:EPI

[1331] Catterina Dagnino. Extended product integration rules. BIT, 23(4):487–499, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=487.

Lotstedt:1983:PBL

[1332] Per Lötstedt. Perturbation bounds for the linear least squares problem subject to linear inequality constraints. BIT, 23(4):500–519, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=500.

Qin:1983:SDS

[1333] Meng Zhao Qin. Stability of difference schemes for a class of partial differential equations. BIT, 23(4):520–528, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=520.

Mjelde:1983:MMR

[1334] K. M. Mjelde. Max-min resource allocation. BIT, 23(4):529–537, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=529.

Borwein:1983:VRC

[1335] J. M. Borwein and P. B. Borwein. A very rapidly convergent product expansion for π. BIT, 23(4):538–540, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=538.

Chawla:1983:USN

[1336] M. M. Chawla. Unconditionally stable Noumerov-type methods for second order differential equations. BIT, 23(4):541–542, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=541.

Sakai:1983:SNCb

[1337] Manabu Sakai. Some new consistency relations connecting spline values at mesh and mid points. BIT, 23(4): 543–546, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=543.

Weintraub:1983:EPG

[1338] S. Weintraub. Erratum: “A prime gap of 682 and a prime arithmetic sequence” [BIT 22(4), 1982, pp. 538; MR 84b:10014]. BIT, 23(4):546, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=546. See

[1270].

Anonymous:1983:Pd

[1339] Anonymous. Problems. BIT, 23(4):

547–548, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=547.

Anonymous:1983:Nd

[1340] Anonymous. Notes. BIT, 23(4):549, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=23&issue=4&spage=549.

Anonymous:1983:A

[1341] Anonymous. Acknowledgements. BIT, 23(4):550, December 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=23&issue=4&spage=550.

Anonymous:1983:I

[1342] Anonymous. Index 1983. BIT, 23 (4):551–552, 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Allison:1984:SPT

[1343] D. C. S. Allison and M. T. Noga. Some performance tests of convex hull algorithms. BIT, 24(1):2–13, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=1&spage=2.

Carlsson:1984:IWC

[1344] Svante Carlsson. Improving worstcase behavior of heaps. BIT, 24(1): 14–18, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=1&spage=14.

Ernvall:1984:ELM

[1345] J. Ernvall and O. Nevalainen. Estimating the length of minimal spanning trees in compression of files. BIT, 24(1):19–32, March 1984. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=24&
issue=1&spage=19.

Fenner:1984:LFA

[1346] T. I. Fenner and G. Loizou. Loopfree algorithms for traversing binary trees. BIT, 24(1):33–44, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=33.

Hsu:1984:CLC

[1347] W. J. Hsu and M. W. Du. Computing a longest common subsequence for a set of strings. BIT, 24(1):45– 59, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=45.

Malmquist:1984:CPS

[1348] J. P. Malmquist and E. L. Robertson. On the complexity of partitioning sparse matrix representations. BIT, 24(1):60–68, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=24&
issue=1&spage=60.

Bell:1984:AAS

[1349] G. E. Bell and G. M. Phillips.

Aitken acceleration of some alternating series. BIT, 24(1):70–77, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=70.

Chen:1984:AGH

[1350] J. Chen, L. W. Johnson, and R. D. Riess. An algorithm for Gauss harmonic formulas. BIT, 24(1):78–84, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=78.

Grigorieff:1984:ZSS

[1351] R. D. Grigorieff and P. J. Paes-Leme. On the zero-stability of the 3-step BDF-formula on nonuniform grids. BIT, 24(1):85–91, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=85.

Linz:1984:USL

[1352] Peter Linz. Uncertainty in the solution of linear operator equations. BIT, 24(1):92–101, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=24&
issue=1&spage=92.

Neumann:1984:BEA

[1353] M. Neumann and R. J. Plemmons.

Backward error analysis for linear systems associated with inverses of H-matrices. BIT, 24(1):102–112, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=102.

Bohman:1984:NCS

[1354] Jan Bohman and Carl-Erik Fröberg.

On numerical computation of singular integrals. BIT, 24(1):113–116, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=113.

Chawla:1984:NME

[1355] M. M. Chawla. Numerov made explicit has better stability. BIT, 24(1):117– 118, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=117.

Navlakha:1984:ATD

[1356] Jainendra K. Navlakha. An analytical technique for 3-dimensional interpolation. BIT, 24(1):119–122, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=1&spage=119.

Smoktunowicz:1984:BCI

[1357] Alicja Smoktunowicz and Jolanta Sokolnicka. Binary cascades iterative refinement in doubled-mantissa arithmetics. BIT, 24(1):123–127, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=1&spage=123.

Anonymous:1984:Pa

[1358] Anonymous. Problems. BIT, 24 (1):128–130, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=1&spage=128.

Anonymous:1984:NBJa

[1359] Anonymous. New books and journals. BIT, 24(1):131, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=24&
                                        issue=1&spage=131.

Anonymous:1984:Na

[1360] Anonymous. Notes. BIT, 24(1): 132, March 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=1&spage=132.

Ghosh:1984:PSA

[1361] Ratan K. Ghosh and G. P. Bhattacharjee. A parallel search algorithm for directed acyclic graphs. BIT, 24(2):133– 150, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=133.

Maus:1984:DTC

[1362] Arne Maus. Delaunay triangulation and the convex hull of n points in expected linear time. BIT, 24(2):151– 163, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=151.

Makinen:1984:CFS

[1363] Erkki Mäkinen. On context-free and Szilard languages. BIT, 24(2):164– 170, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=164.

Niemi:1984:SQL

[1364] Timo Niemi. Specification of a query language by the attribute method. BIT, 24(2):171–186, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=24&
issue=2&spage=171.

Tseng:1984:NPS

[1365] S. S. Tseng and R. C. T. Lee. A new parallel sorting algorithm based upon min-mid-max operations. BIT, 24(2): 187–195, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=2&spage=187.

Zaks:1984:NAG

[1366] Shmuel Zaks. A new algorithm for generation of permutations. BIT, 24(2): 196–204, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=2&spage=196.

Lotstedt:1984:SML

[1367] Per Lötstedt. Solving the minimal least squares problem subject to bounds on the variables. BIT, 24(2):205– 224, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=2&spage=205.

Thomas:1984:PPH

[1368] R. M. Thomas. Phase properties of high order, almost P - stable formulae. BIT, 24(2):225– 238, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=2&spage=225.

Berntsen:1984:UGQ

[1369] Jarle Berntsen and Terje O. Espelid. On the use of Gauss quadrature in adaptive automatic integration schemes. BIT, 24(2):239– 242, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=239.

Buell:1984:RCQ

[1370] Duncan A. Buell and Richard H.

Hudson. On runs of consecutive quadratic residues and quadratic nonresidues. BIT, 24(2):243–247, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=243.

Cash:1984:ESM

[1371] J. R. Cash. Efficient P -stable methods for periodic initial value problems. BIT, 24(2):248–252, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=248. See

erratum [1398].

Smith:1984:NAC

[1372] H. V. Smith. The numerical approximation of a class of finite integrals. BIT, 24(2):253– 256, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=2&spage=253.

Anonymous:1984:Pb

[1373] Anonymous. Problems. BIT, 24(2): 257–258, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=2&spage=257.

Anonymous:1984:NBJb

[1374] Anonymous. New books and journals. BIT, 24(2):259, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=24&
                                        issue=2&spage=259.

Anonymous:1984:Nb

[1375] Anonymous. Notes. BIT, 24(2): 260, June 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=2&spage=260.

Chan:1984:MFD

[1376] Mee Yee Chan. Multidisk file design: An analysis of folding buckets to disks. BIT, 24(3):262–268, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=261.

Chang:1984:ALD

[1377] R. C. Chang and R. C. T. Lee. On the average length of Delaunay triangulations. BIT, 24(3):269–273, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=269.

Gupta:1984:PSA

[1378] P. Gupta and G. P. Bhattacharjee. A parallel selection algorithm. BIT, 24(3):274–287, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=274.

Nordstrom:1984:PSP

[1379] Bengt Nordström and Jan Smith.

Propositions and specifications of programs in Martin-Löf’s type theory. BIT, 24(3):288–301, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=288.

Prabhu:1984:PPT

[1380] G. M. Prabhu and Narsingh Deo. On the power of a perturbation for testing nonisomorphism of graphs. BIT, 24 (3):302–307, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=302.

Tseng:1984:PAS

[1381] S. S. Tseng and R. C. T. Lee. A parallel algorithm to solve the stable marriage problem. BIT, 24(3):308–316, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=308.

Torn:1984:HOI

[1382] Aimo A. Torn. Hashing with overflow indexing. BIT, 24(3):317–332, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=317.

Yeh:1984:GAT

[1383] D. Y. Yeh and D. T. Lee. Graph algorithms on a tree-structured parallel computer. BIT, 24(3):333–340, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=333.

Agarwal:1984:SHP

[1384] Ravi P. Agarwal and Ramesh C. Gupta. On the solution of Holt’s problem. BIT, 24(3):341–346, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=341.

Dekker:1984:EBS

[1385] K. Dekker. Error bounds for the solution to the algebraic equations in Runge–Kutta methods. BIT, 24(3): 347–356, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=347.

Missirlis:1984:EFO

[1386] N. M. Missirlis. The extrapolated first order method for solving systems with complex eigenvalues. BIT, 24(3):357–365, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=357.

Neumaier:1984:IVS

[1387] A. Neumaier. An interval version of the secant method. BIT, 24(3):366–372, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=366.

Sammon:1984:IRK

[1388] P. H. Sammon and P. Forsyth, Jr. Instability in Runge–Kutta schemes for simulation of oil recovery. BIT, 24(3):373–379, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=373.

Brugner:1984:TTT

[1389] Georg Brügner. Three-triangle-tangram. BIT, 24(3):380–382, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=380.

Hairer:1984:NS

[1390] Ernst Hairer. A note on Dstability. BIT, 24(3):383–386, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=383.

Huang:1984:ANR

[1391] Shou Hsuan Stephen Huang and C. K. Wong. Average number of rotations and access cost in iR-trees. BIT, 24(3):387–390, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=3&spage=387.

Soderlind:1984:EBF

[1392] Gustaf Söderlind. An error bound for fixed-point iterations. BIT, 24(3): 391–393, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=391.

Ypma:1984:LSS

[1393] T. J. Ypma. Linear stability of stiff differential equation solvers. BIT, 24 (3):394–396, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=394.

Anonymous:1984:Pc

[1394] Anonymous. Problems. BIT, 24(3):

397–400, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=397.

Anonymous:1984:BR

[1395] Anonymous. Books received. BIT, 24(3):400, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=400.

Anonymous:1984:E

[1396] Anonymous. Erratum. BIT, 24(3):400, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=400.

Anonymous:1984:Nc

[1397] Anonymous. Notes. BIT, 24(3):400, September 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=3&spage=400.

Cash:1984:EES

[1398] J. R. Cash. Erratum: “Efficient P - stable methods for periodic initial value problems” [BIT 24(2), 1984, pp. 248– 252]. BIT, 24(3):400, 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [1371].

Anderssen:1984:CFD

[1399] R. S. Anderssen and F. R. de Hoog. On the correction of finite difference eigenvalue approximations for Sturm– Liouville problems with general boundary conditions. BIT, 24(4):401–412, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=401.

Axelsson:1984:EEI

[1400] O. Axelsson. Error estimates over infinite intervals of some discretizations of evolution equations. BIT, 24(4): 413–424, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=413.

Butcher:1984:ARK

[1401] J. C. Butcher. An application of the Runge–Kutta space. BIT, 24(4): 425–440, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=425.

Cybenko:1984:NSL

[1402] George Cybenko. The numerical stability of the lattice algorithm for least squares linear prediction problems. BIT, 24(4):441–455, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=441.

Deuflhard:1984:CPS

[1403] P. Deuflhard. Computation of periodic solutions of nonlinear ODEs. BIT, 24 (4):456–466, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=456.

Elden:1984:NCG

[1404] Lars Eldén. A note on the computation of the generalized cross-validation function for ill-conditioned least squares problems. BIT, 24(4):467–472, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=467.

Gautschi:1984:SOP

[1405] Walter Gautschi. On some orthogonal polynomials of interest in theoretical chemistry. BIT, 24(4):473–483, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=473.

Gear:1984:MLM

[1406] C. W. Gear and D. R. Wells. Multirate linear multistep methods. BIT, 24(4):484–502, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=484.

Goldstein:1984:NCA

[1407] A. A. Goldstein. A note on the complexity of an algorithm for Chebyshev approximation. BIT, 24(4):503–509, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=503.

Gustafson:1984:SCC

[1408] Sven-Åke Gustafson. On the stability of a class of convergence acceleration methods for power series. BIT, 24(4):510–519, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=510.

Hairer:1984:ESS

[1409] E. Hairer and H. Türke. The equivalence of B-stability and Astability. BIT, 24(4):520–528, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=520.

Iserles:1984:PFD

[1410] Arieh Iserles and Syvert P. Nørsett.

A proof of the first Dahlquist barrier by order stars. BIT, 24(4):529–537, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=529.

Jeltsch:1984:DFB

[1411] Rolf Jeltsch and Olavi Nevanlinna.

Dahlquist’s first barrier for multistage multistep formulas. BIT, 24(4): 538–555, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=538.

Kjellberg:1984:TOD

[1412] Göran Kjellberg. Two observations on Durand–Kerner’s root-finding method. BIT, 24(4):556–559, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=556.

Kreiss:1984:CDS

[1413] H. O. Kreiss. Central difference schemes and stiff boundary value problems. BIT, 24(4):560–567, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=560.

Kaagstrom:1984:GSV

[1414] Bo Kågström. The generalized singular value decomposition and the general (A−λB)-problem. BIT, 24(4):568–583, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=568.

LeVeque:1984:RCK

[1415] Randall J. LeVeque and Lloyd N. Trefethen. On the resolvent condition in the Kreiss Matrix Theorem. BIT, 24(4):584–591, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=584.

Liniger:1984:FDL

[1416] Werner Liniger. On factored discretizations of the Laplacian for the fast solution of Poisson’s equation on general regions. BIT, 24(4):592–608, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=592.

Mattheij:1984:SBV

[1417] R. M. M. Mattheij and R. E. O’Malley, Jr. On solving boundary value problems for multiscale systems using asymptotic approximations and multiple-shooting. BIT, 24(4):609–622, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=609.

Meurant:1984:BPC

[1418] Gérard Meurant. The block preconditioned conjugate gradient method on vector computers. BIT, 24(4):623–633, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=623.

Norsett:1984:ESM

[1419] Syvert P. Nørsett and Per G. Thomsen. Embedded SDIRK-methods of basic order three. BIT, 24(4):634–646, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=24&issue=4&spage=634.

Reichel:1984:AOP

[1420] Lothar Reichel. An asymptotically orthonormal polynomial family. BIT, 24(4):647–655, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=647.

Spijker:1984:RBS

[1421] M. N. Spijker. On the relation between stability and contractivity. BIT, 24(4):656–666, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=656.

Soderlind:1984:NDD

[1422] Gustaf Söderlind. On nonlinear difference and differential equations. BIT, 24(4):667–680, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=667.

VanDooren:1984:DCS

[1423] P. Van Dooren. Deadbeat control: a special inverse eigenvalue problem. BIT, 24(4):681–699, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=24&issue=4&spage=681.

Zlatev:1984:APC

[1424] Zahari Zlatev. Application of predictor-corrector schemes with several correctors in solving air pollution problems. BIT, 24(4):700–715, 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Arnborg:1985:EAC

[1425] Stefan Arnborg. Efficient algorithms for combinatorial problems on graphs with bounded decomposability — a survey. BIT, 25(1):1– 23, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=1.

Franksen:1985:NDM

[1426] Ole Emmanuel Franksen. The nature of data — from measurements to systems. BIT, 25(1):24– 50, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=24. In-

vited paper.

Nossum:1985:ATP

[1427] Rolf Nossum. Automated theorem proving methods. BIT, 25(1):51– 64, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=51.

Tienari:1985:CNN

[1428] Martti Tienari. Computer networking of Nordic universities. BIT, 25(1):65– 69, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=65. In-

vited paper.

Bui:1985:SIF

[1429] T. D. Bui and Mai Thanh. Significant improvements to the Ford-Johnson algorithm for sorting. BIT, 25(1):70– 75, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=70.

Chazelle:1985:PGD

[1430] Bernard Chazelle, Leo J. Guibas, and D. T. Lee. The power of geometric duality. BIT, 25(1):76–90, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=76.

Daini:1985:LDG

[1431] Ola-Olu A. Daini. A languagedriven generalized numerical database translator. BIT, 25(1):91–105, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=91.

Evans:1985:APP

[1432] D. J. Evans and Nadia Y. Yousif. Analysis of the performance of the parallel quicksort method. BIT, 25(1):106– 112, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=106.

Flajolet:1985:ACD

[1433] Philippe Flajolet. Approximate counting: a detailed analysis. BIT, 25 (1):113–134, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=113.

Tamminen:1985:SAC

[1434] Markku Tamminen. On search by address computation. BIT, 25(1):135– 147, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=135. Dis-

cusses hashing applications to file searching.

Yang:1985:BMC

[1435] W. P. Yang and M. W. Du. A backtracking method for constructing perfect hash functions from a set of mapping functions. BIT, 25(1):148– 164, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=148.

Axelsson:1985:SPI

[1436] O. Axelsson. A survey of preconditioned iterative methods for linear systems of algebraic equations. BIT, 25(1):165–187, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=25&
                                        issue=1&spage=165.

Dahlquist:1985:YNI

[1437] Germund Dahlquist. 33 years of numerical instability. I. BIT, 25(1):188– 204, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=188.

Riesel:1985:MFM

[1438] Hans Riesel. Modern factorization methods. BIT, 25(1):205–222, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=205.

DiLena:1985:CEU

[1439] Giovanni Di Lena and Roberto I. Peluso. On conditions for the existence and uniqueness of solutions to the algebraic equations in Runge– Kutta methods. BIT, 25(1):223– 232, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=223.

Gupta:1985:MRK

[1440] Suchitra Gupta. Multiderivative Runge–Kutta processes for two-point boundary value problems. BIT, 25 (1):233–241, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=233.

Ioakimidis:1985:SQT

[1441] N. I. Ioakimidis and E. G. Anastasselou. A simple quadrature-type method for the computation of real zeros of analytic functions in finite intervals. BIT, 25(1):242–249, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=242.

Jain:1985:FOI

[1442] M. K. Jain. Fifth order implicit multipoint method for solving equations. BIT, 25(1):250–255, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=250.

Neumaier:1985:IIZ

[1443] A. Neumaier. Interval iteration for zeros of systems of equations. BIT, 25(1):256–273, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=256.

Wallace:1985:PRR

[1444] Stein W. Wallace. Pivoting rules and redundancy schemes in extreme point enumeration. BIT, 25(1):274– 280, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=274.

Amini:1985:SAM

[1445] S. Amini. Stability analysis of methods employing reducible rules for Volterra integral equations of the first kind. BIT, 25(1):281–284, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=281.

Dekker:1985:NCS

[1446] Kees Dekker and Ernst Hairer. A necessary condition for BSIstability. BIT, 25(1):285–288, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=25&
                                        issue=1&spage=285.

Ernvall:1985:NCH

[1447] Jarmo Ernvall, Jyrki Katajainen, and Martti Penttonen. NP-completeness of the Hamming salesman problem. BIT, 25(1):289–292, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=25&
issue=1&spage=289.

Makinen:1985:NDF

[1448] Erkki Mäkinen. A note on depthfirst derivations. BIT, 25(1):293– 296, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=1&spage=293.

Reimer:1985:ASI

[1449] M. Reimer. All symmetric interpolatory block-implicit methods of order less than six are Astable. BIT, 25(1):297–298, March 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=25&
issue=1&spage=297.

Anonymous:1985:Pa

[1450] Anonymous. Problems. BIT, 25(1):

299–300, 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1985:BR

[1451] Anonymous. Books received. BIT, 25(1):300, 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Auger:1985:PAM

[1452] Ivan E. Auger and M. S. Krishnamoorthy. A parallel algorithm for the monadic unification problem. BIT, 25(2):302– 306, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=302.

Francis:1985:OAP

[1453] N. D. Francis and D. I. Fleming. Optimum allocation of places to students in a national university system. BIT, 25(2):307–317, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=307.

Krogdahl:1985:MIS

[1454] Stein Krogdahl. Multiple inheritance in Simula-like languages. BIT, 25(2): 318–326, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=318.

Pettorossi:1985:THP

[1455] Alberto Pettorossi. Towers of Hanoi problems: deriving iterative solutions by program transformations. BIT, 25(2):327–334, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=327. See

correction [3457].

Regnier:1985:AGF

[1456] Mireille Regnier. Analysis of grid file algorithms. BIT, 25(2):335– 357, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=335.

Thorelli:1985:LLM

[1457] Lars-Erik Thorelli. A language for linking modules into systems. BIT, 25(2): 358–378, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=358.

vanAsselt:1985:FNR

[1458] E. J. van Asselt. On M-functions and nonlinear relaxation methods. BIT, 25(2):380–385, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=380.

Cooper:1985:ASE

[1459] G. J. Cooper. Algebraic stability and the existence of solutions of implicit Runge–Kutta equations. BIT, 25(2): 386–390, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=2&spage=386.

Sand:1985:CVS

[1460] Jørgen Sand. A_{0}-contractivity of the variable-step BDFs of order less than four. BIT, 25(2):391– 398, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=2&spage=391.

Schneider:1985:GSI

[1461] Claus Schneider. Generalized singlyimplicit Runge–Kutta methods with arbitrary knots. BIT, 25(2):399– 412, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=2&spage=399.

Venturino:1985:EBG

[1462] Ezio Venturino. Error bounds of the Galerkin method for singular integral equations of the second kind. BIT, 25(2):413–419, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=2&spage=413.

Berkovits:1985:AHJ

[1463] Shimshon Berkovits and Joel E. Sachs. Analysis of the Herlestam and Johannesson discrete logarithm scheme in GF(2^{n}) for large n^{1}. BIT, 25(2):420– 424, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=420.

Valenca:1985:MSU

[1464] M. R. Valença. Multiple shooting using interval analysis. BIT, 25(2):425– 427, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=425.

Anonymous:1985:Pb

[1465] Anonymous. Problems. BIT, 25(2):

428–430, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=428.

Anonymous:1985:NBJa

[1466] Anonymous. New books and journals. BIT, 25(2):431, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=25&
issue=2&spage=431.

Anonymous:1985:Na

[1467] Anonymous. Notes. BIT, 25(2):

432, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=2&spage=432.

Fenner:1985:AFM

[1468] T. I. Fenner and A. M. Frieze. An algorithm for finding a matroid basis which maximizes the product of the weights of the elements. BIT, 25(3):433–438, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=3&spage=433.

Koskimies:1985:NOP

[1469] Kai Koskimies. A note on onepass evaluation of attribute grammars. BIT, 25(3):439–450, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=3&spage=439.

Noga:1985:SLE

[1470] M. T. Noga and D. C. S. Allison. Sorting in linear expected time. BIT, 25(3):451–465, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=3&spage=451.

Nurmi:1985:FLS

[1471] Otto Nurmi. A fast line-sweep algorithm for hidden line elimination. BIT, 25(3):466–472, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=466.

Quinn:1985:NTP

[1472] Michael J. Quinn. A note on two parallel algorithms to solve the stable marriage problem. BIT, 25(3):473–476, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=473.

Ramamohanarao:1985:PMR

[1473] K. Ramamohanarao and R. Sacks-Davis. Partial match retrieval using recursive linear hashing. BIT, 25(3): 477–484, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=477.

Winter:1985:GSP

[1474] Pawel Winter. Generalized Steiner problem in outerplanar networks. BIT, 25(3):485–496, September 1985. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=25&issue=3&spage=
485.

Krishnaiah:1985:SME

[1475] U. Anantha Krishnaiah. E-stable methods for exponentially decreasing solutions of second order initial value problems. BIT, 25(3):497–506, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=497.

Barlow:1985:SAA

[1476] J. L. Barlow. Stability analysis of the G-algorithm and a note on its application to sparse least squares problems. BIT, 25(3):507–520, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=507.

Butcher:1985:NTS

[1477] J. C. Butcher. The nonexistence of ten-stage eighth order explicit Runge– Kutta methods. BIT, 25(3):521–540, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=521.

Ersoy:1985:RRD

[1478] O. Ersoy. Representation of real discrete Fourier transform in terms of a new set of functions based upon Möbius inversion. BIT, 25(3):541–556, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=541.

Anonymous:1985:Pc

[1479] Anonymous. Problems. BIT, 25(3):

557–559, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=557.

Anonymous:1985:Nc

[1480] Anonymous. Notes. BIT, 25(3):560, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=560.

Anonymous:1985:NBJb

[1481] Anonymous. New books and journals. BIT, 25(3):560, September 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=3&spage=560.

Armenakis:1985:ARF

[1482] A. C. Armenakis, L. E. Garey, and R. D. Gupta. An adaptation of a root finding method to searching ordered disk files. BIT, 25(4):561–568, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=561. See

revisitation [1759].

Hinrichs:1985:IGF

[1483] Klaus Hinrichs. Implementation of the grid file: design concepts and experience. BIT, 25(4):569–592, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=569.

Jakobsson:1985:CCS

[1484] Matti Jakobsson. Compression of character strings by an adaptive dictionary. BIT, 25(4):593–603, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=593.

Makinen:1985:PGG

[1485] Erkki Mäkinen. On permutative grammars generating context-free languages. BIT, 25(4):604–610, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=604.

Penttonen:1985:NCS

[1486] Martti Penttonen and Jyrki Katajainen. Notes on the complexity of sorting in abstract machines. BIT, 25(4):611–622, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=611.

Dixon:1985:OED

[1487] Jennifer Dixon. On the order of the error in discretization methods for weakly singular second kind Volterra integral equations with nonsmooth solutions. BIT, 25(4):623–634, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=623.

Jankowski:1985:ASC

[1488] M. Jankowski and H. Woźniakowski.

The accurate solution of certain continuous problems using only single precision arithmetic. BIT, 25(4):635–651, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=635.

Kraaijevanger:1985:CIM

[1489] J. F. B. M. Kraaijevanger. Bconvergence of the implicit midpoint rule and the trapezoidal rule. BIT, 25 (4):652–666, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=652.

Smeets:1985:NPF

[1490] Bernard Smeets. On the number of polynomials over GF(2) that factor into 2, 3 or 4 prime polynomials. BIT, 25(4):667–674, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=4&spage=667.

Cooper:1985:RRK

[1491] G. J. Cooper. Reducible Runge– Kutta methods. BIT, 25(4):675–680, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=4&spage=675.

Ioakimidis:1985:MCQ

[1492] N. I. Ioakimidis. A modification of the classical quadrature method for locating zeros of analytic functions. BIT, 25(4):681–686, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=4&spage=681.

Sand:1985:IMS

[1493] Jørgen Sand. Integration methods for solving equations. BIT, 25(4):687–688, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=25&issue=4&spage=687.

Anonymous:1985:Pd

[1494] Anonymous. Problems. BIT, 25(4): 689–690, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=689.

Anonymous:1985:NBJc

[1495] Anonymous. New books and journals. BIT, 25(4):691, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=691.

Anonymous:1985:BPY

[1496] Anonymous. BIT Prize for Young Nordic Scientists. BIT, 25(4):692, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=692.

Anonymous:1985:Nd

[1497] Anonymous. Notes. BIT, 25(4):692, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=692.

Anonymous:1985:A

[1498] Anonymous. Acknowledgements. BIT, 25(4):693, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=4&spage=693.

Anonymous:1985:I

[1499] Anonymous. Index 1985. BIT, 25 (4):694–696, 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Chan:1986:GCP

[1500] Becky Chan and Selim G. Akl. Generating combinations in parallel. BIT, 26(1):1–6, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=26&
                                        issue=1&spage=1.

Chang:1986:MST

[1501] R. C. Chang and R. C. T. Lee. An O(Nlog N) minimal spanning tree algorithm for N points in the plane. BIT, 26(1):7–16, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=26&
                                        issue=1&spage=7.

Louchard:1986:BMA

[1502] G. Louchard. Brownian motion and algorithm complexity. BIT, 26(1):17– 34, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=1&spage=17.

Quinn:1986:UBS

[1503] Michael J. Quinn and Narsingh Deo. An upper bound for the speedup of parallel best-bound branch-and-bound algorithms. BIT, 26(1):35– 43, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=35.

Winter:1986:AES

[1504] Pawel Winter. An algorithm for the enumeration of spanning trees. BIT, 26(1):44–62, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=44.

Chawla:1986:PLA

[1505] M. M. Chawla and P. S. Rao. Phaselag analysis of explicit Nyström methods for “y = f(x, y)”. BIT, 26(1):63– 70, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=63.

Cuyt:1986:MPA

[1506] Annie Cuyt. Multivariate Padé approximants revisited. BIT, 26(1):71– 79, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=71.

Hofmann:1986:RDC

[1507] Bernd Hofmann and Volkmar Friedrich.

On regularization and discretization control for the numerical solution of inverse problems in parabolic equations. BIT, 26(1):80–92, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=26&
issue=1&spage=80.

Iyengar:1986:DFM

[1508] S. R. K. Iyengar and R. K. Jain. Derivative free multipoint iterative methods for simple and multiple roots. BIT, 26(1):93–99, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=93.

Norsett:1986:LEC

[1509] Syvert P. Nørsett and Per G. Thomsen. Local error control in SDIRK-methods. BIT, 26(1):100– 113, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=100.

Sand:1986:CCL

[1510] Jørgen Sand. Circle contractive linear multistep methods. BIT, 26(1):114– 122, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=114.

Borwein:1986:ECI

[1511] J. M. Borwein and P. B. Borwein. An explicit cubic iteration for π. BIT, 26(1):123–126, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=123.

Bing-Chao:1986:OWS

[1512] Huang Bing-Chao and Donald E.

Knuth. A one-way, stackless quicksort algorithm. BIT, 26(1):127–130, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=127.

Rabinowitz:1986:CIP

[1513] Philip Rabinowitz. The convergence of interpolatory product integration rules. BIT, 26(1):131–134, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=131.

Anonymous:1986:Pa

[1514] Anonymous. Problems. BIT, 26(1):

135, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=1&spage=135.

Anonymous:1986:Na

[1515] Anonymous. Notes. BIT, 26(1): 136, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=1&spage=136.

Du:1986:DAM

[1516] H. C. Du. Disk allocation methods for binary Cartesian product files. BIT, 26(2):138–147, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=2&spage=138.

Edenbrandt:1986:QTP

[1517] Anders Edenbrandt. Quotient tree partitioning of undirected graphs. BIT, 26(2):148–155, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=2&spage=148.

Kingston:1986:ACH

[1518] Jeffrey H. Kingston. The amortized complexity of Henriksen’s algorithm. BIT, 26(2):156–163, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=26&
                                        issue=2&spage=156.

Meldal:1986:ASN

[1519] Sigurd Meldal. An axiomatic semantics for nested concurrency. BIT, 26(2): 164–174, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=164.

Naur:1986:TTT

[1520] Peter Naur. Thinking and Turing’s test. BIT, 26(2):175–187, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=175.

Saikkonen:1986:DTR

[1521] Heikki Saikkonen and Stefan Rönn.

Distributed termination on a ring. BIT, 26(2):188–194, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=188.

Zhang:1986:NPD

[1522] Y. Zhang. A note on parallel depth first search. BIT, 26(2):195– 198, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=195.

Amini:1986:EQR

[1523] S. Amini. Efficient quadrature rules with a priori error estimates for integrands with end point singularities. BIT, 26(2):200–208, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=200.

Golub:1986:CTS

[1524] Gene H. Golub and R. Kannan. Convergence of a two-stage Richardson process for nonlinear equations. BIT, 26(2):209–216, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=26&
issue=2&spage=209.

Hanna:1986:APC

[1525] Martin S. Hanna, David G. Evans, and Peter N. Schweitzer. On the approximation of plane curves by parametric cubic splines. BIT, 26(2):217– 232, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=217.

Hunter:1986:MGQ

[1526] D. B. Hunter and G. E. Okecha. A modified Gaussian quadrature rule for integrals involving poles of any order. BIT, 26(2):233– 240, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=233.

Koch:1986:JTG

[1527] Per Erik Koch. Jackson theorems for generalized polynomials with special applications to trigonometric and hyperbolic functions. BIT, 26(2):241– 250, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=241.

Andrew:1986:ANM

[1528] A. L. Andrew. The accuracy of Numerov’s method for eigenvalues. BIT, 26(2):251–253, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=26&
issue=2&spage=251.

Brezinski:1986:IMP

[1529] Claude Brezinski. On interpolatory multivariate Padé-type approximants. BIT, 26(2):254–258, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=2&spage=254.

Herzberger:1986:BOC

[1530] J. Herzberger. Bounds for the Rorder of certain iterative numerical processes. BIT, 26(2):259– 262, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=2&spage=259.

Anonymous:1986:Pb

[1531] Anonymous. Problems. BIT, 26(2): 263, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=2&spage=263.

Anonymous:1986:Nb

[1532] Anonymous. Notes. BIT, 26(2): 264, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=2&spage=264.

Anonymous:1986:NBJa

[1533] Anonymous. New books and journals. BIT, 26(2):264, June 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=26&
                                        issue=2&spage=264.

Anonymous:1986:NBJb

[1534] Anonymous. New books and journals. BIT, 26(2):264, 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Berztiss:1986:TBT

[1535] Alfs Berztiss. A taxonomy of binary tree traversals. BIT, 26(3):266–276, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=266.

Chen:1986:PGP

[1536] G. H. Chen and Maw Sheng Chern.

Parallel generation of permutations and combinations. BIT, 26(3):277–283, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=277.

Krarup:1986:AAA

[1537] Jakob Krarup and Peter Pruzan.

Assessment of approximate algorithms: The error measure’s crucial role. BIT, 26(3):284–294, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=284.

Meldal:1986:PCE

[1538] Sigurd Meldal. Partial correctness of exits from concurrent structures. BIT, 26(3):295–302, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=295.

Schreiber:1986:SAM

[1539] Robert Schreiber. On systolic array methods for band matrix factorizations. BIT, 26(3):303–316, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=303.

Chawla:1986:OMS

[1540] M. M. Chawla, S. McKee, and G. Shaw. Order h^{2} method for a singular two-point boundary value problem. BIT, 26(3):318–326, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=318.

Forster:1986:CQS

[1541] Klaus-Jürgen Förster. On Chebyshev quadrature for a special class of weight functions. BIT, 26(3):327–332, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=327.

Koppinen:1986:CSC

[1542] M. Koppinen. Computation of simple characters of a Chevalley group. BIT, 26(3):333–338, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=333.

Norsett:1986:SBM

[1543] Syvert P. Nørsett and Per G. Thomsen. Switching between modified Newton and fix-point iteration for implicit ODE-solvers. BIT, 26(3):339–348, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=339.

Reichel:1986:PAU

[1544] Lothar Reichel. On polynomial approximation in the uniform norm by the discrete least squares method. BIT, 26 (3):349–368, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=349.

Saridakis:1986:GCO

[1545] Yiannis G. Saridakis. Generalized consistent orderings and the accelerated overrelaxation method. BIT, 26(3): 369–376, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=369.

Schmidt:1986:CII

[1546] Jochen W. Schmidt. Convex interval interpolation with cubic splines. BIT, 26(3):377–387, September 1986. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        0006-3835&volume=26&issue=3&spage=
                                        377.

Frenzen:1986:NEE

[1547] C. L. Frenzen. A note on an error estimate for least squares approximation. BIT, 26(3):388–391, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=3&spage=388.

Pardalos:1986:ACN

[1548] Panos M. Pardalos. An algorithm for a class of nonlinear fractional problems using ranking of the vertices. BIT, 26(3):392–395, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=3&spage=392.

Sakai:1986:CCP

[1549] Manabu Sakai. A characterization of a class of polynomial splines by convolution. BIT, 26(3):396–397, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=3&spage=396.

Anonymous:1986:Pc

[1550] Anonymous. Problems. BIT, 26(3): 398–399, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=398.

Anonymous:1986:NBJc

[1551] Anonymous. New books and journals. BIT, 26(3):399–400, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=399.

Anonymous:1986:Nc

[1552] Anonymous. Notes. BIT, 26(3):400, September 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=3&spage=400.

Brinck:1986:CPN

[1553] Keith Brinck. Computing parent nodes in threaded binary trees. BIT, 26(4): 401–409, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=401.

Chang:1986:DKL

[1554] C. C. Chang. On the design of a key-lock-pair mechanism in information protection systems. BIT, 26(4): 410–417, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=410.

Chaudhuri:1986:PAA

[1555] Pranay Chaudhuri and Ratan K. Ghosh. Parallel algorithms for analyzing activity networks. BIT, 26(4): 418–429, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=418.

Franklin:1986:RVR

[1556] W. Randolph Franklin and Varol Akman. Reconstructing visible regions from visible segments. BIT, 26(4): 430–441, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=430.

Huang:1986:OPQ

[1557] Shou-Hsuan Stephen Huang. Ordered priority queues. BIT, 26(4):442–450, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=442.

Schreiber:1986:SAU

[1558] Robert Schreiber and Wei-Pai Tang. On systolic arrays for updating the Cholesky factorization. BIT, 26(4): 451–466, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=451.

Yeh:1986:DPA

[1559] D. Yun Yeh. A dynamic programming approach to the complete set partitioning problem. BIT, 26(4):467–474, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=467.

Clenshaw:1986:UAR

[1560] C. W. Clenshaw and F. W. J. Olver.

Unrestricted algorithms for reciprocals and square roots. BIT, 26(4):475–492, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=475.

Concus:1986:CIB

[1561] P. Concus and G. Meurant. On computing INV block preconditionings for the conjugate gradient method. BIT, 26(4):493–504, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=493.

Jeltsch:1986:EBM

[1562] Rolf Jeltsch and Olavi Nevanlinna.

Error bounds for multistep methods revisited. BIT, 26(4):505–520, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=4&spage=505.

Milovanovic:1986:NED

[1563] G. V. Milovanović, B. P. Acharya, and T. N. Pattnaik. On numerical evaluation of double integrals of an analytic function of two complex variables. BIT, 26(4):521–526, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=4&spage=521.

Neuman:1986:CIS

[1564] Edward Neuman. Convex interpolating splines of arbitrary degree. III. BIT, 26(4):527–536, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=4&spage=527.

Guggenheimer:1986:IAD

[1565] H. Guggenheimer. Initial approximations in Durand–Kerner’s root finding method. BIT, 26(4):537–539, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=26&issue=4&spage=537.

Anonymous:1986:Pd

[1566] Anonymous. Problems. BIT, 26(4): 540–542, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=540.

Anonymous:1986:NBJd

[1567] Anonymous. New books and journals. BIT, 26(4):543, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=543.

Anonymous:1986:Nd

[1568] Anonymous. Notes. BIT, 26(4):544, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=544.

Anonymous:1986:A

[1569] Anonymous. Acknowledgements. BIT, 26(4):545, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=26&issue=4&spage=545.

Carlsson:1987:ACR

[1570] Svante Carlsson. Average-case results on Heapsort. BIT, 27(1):2– 17, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=2.

Doroslovacki:1987:GCT

[1571] Rade Doroslovački, Ivan Stojmenović, and Ratko Tošić. Generating and counting triangular systems. BIT, 27(1):18–24, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=27&
issue=1&spage=18.

vanLeeuwen:1987:APM

[1572] J. van Leeuwen and J. Wiedermann. Array processing machines: An abstract model. BIT, 27(1):25– 43, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=25.

Wegner:1987:GOW

[1573] Lutz M. Wegner. A generalized, oneway, stackless quicksort. BIT, 27(1): 44–48, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=44.

Ben-Israel:1987:SAP

[1574] Adi Ben-Israel. A Z-simplex algorithm with partial updates. BIT, 27(1):50– 61, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=50.

Burrage:1987:OCA

[1575] K. Burrage and W. H. Hundsdorfer. The order of B-convergence of algebraically stable Runge–Kutta methods. BIT, 27(1):62–71, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=27&
issue=1&spage=62.

Criscuolo:1987:CGT

[1576] Giuliana Criscuolo and Giuseppe Mastroianni. Convergence of Gauss type product formulas for the evaluation of two-dimensional Cauchy principal value integrals. BIT, 27(1):72– 84, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=72.

Ehrenmark:1987:ETF

[1577] Ulf Torsten Ehrenmark. On the error term of the Filon quadrature formulae. BIT, 27(1):85–97, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=85.

Mayer:1987:SBP

[1578] Daniel C. Mayer. Sharp bounds for the partition function of integer sequences. BIT, 27(1):98–110, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=1&spage=98.

Petkovic:1987:SIF

[1579] Miodrag S. Petković and Lidija V. Stefanović. On some iteration functions for the simultaneous computation of multiple complex polynomial zeros. BIT, 27(1):111–122, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=1&spage=111.

Rockswold:1987:SLM

[1580] Gary K. Rockswold. A_{0}-stable linear multistep formulas of the αtype. BIT, 27(1):123–128, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=27&
                                        issue=1&spage=123.

Anonymous:1987:Pa

[1581] Anonymous. Problems. BIT, 27 (1):129–130, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=1&spage=129.

Anonymous:1987:NBJa

[1582] Anonymous. New books and journals. BIT, 27(1):131, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=27&
issue=1&spage=131.

Anonymous:1987:Na

[1583] Anonymous. Notes. BIT, 27(1):

132, March 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=1&spage=132.

Burgdorff:1987:AMR

[1584] H. A. Burgdorff, S. Jajodia, F. N.

Springsteel, and Y. Zalcstein. Alternative methods for the reconstruction of trees from their traversals. BIT, 27(2): 133–140, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=133.

Chen:1987:NSA

[1585] G. H. Chen, Maw-Sheng Chern, and R. C. T. Lee. A new systolic architecture for convex hull and half-plane intersection problems. BIT, 27(2):141– 147, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=141.

Friesen:1987:PON

[1586] Donald K. Friesen and Michael A.

Langston. Bin packing: on optimizing the number of pieces packed. BIT, 27(2):148–156, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=148.

Loizou:1987:FWC

[1587] G. Loizou and P. Thanisch. On finding a worst-case optimal fourth normal form database decomposition. BIT, 27(2):157–162, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=27&
                                        issue=2&spage=157.

Makinen:1987:LDB

[1588] Erkki Mäkinen. Left distance binary tree representations. BIT, 27(2):163– 169, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=163.

Pawagi:1987:MWI

[1589] Shaunak Pawagi. Maximum weight independent set in trees. BIT, 27(2):170– 180, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=170. See

erratum [1630].

Butcher:1987:LNL

[1590] J. C. Butcher. Linear and nonlinear stability for general linear methods. BIT, 27(2):181–189, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=27&
issue=2&spage=181.

Butcher:1987:LNS

[1591] J. C. Butcher. Linear and nonlinear stability for general linear methods. BIT, 27(2):182–189, 1987. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cuminato:1987:UCC

[1592] J. A. Cuminato. On the uniform convergence of a collocation method for a class of singular integral equations. BIT, 27(2):190– 202, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=190. See

erratum [1629].

Gautschi:1987:CHT

[1593] Walter Gautschi and Jet Wimp. Computing the Hilbert transform of a Jacobi weight function. BIT, 27(2):203– 215, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=203.

Lubich:1987:MDI

[1594] Ch. Lubich and A. Ostermann. Multigrid dynamic iteration for parabolic equations. BIT, 27(2):216–234, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=216.

McAlevey:1987:PIR

[1595] L. G. McAlevey. Product integration rules for Volterra integral equations of the first kind. BIT, 27(2):235– 247, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=235.

Riesel:1987:MTN

[1596] Hans Riesel. On the metric theory of nearest integer continued fractions. BIT, 27(2):248– 263, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=248. See

errata [1651].

Strehmel:1987:CRL

[1597] K. Strehmel and R. Weiner. Bconvergence results for linearly implicit one step methods. BIT, 27(2):264– 281, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=2&spage=264.

Sakai:1987:RRS

[1598] Manabu Sakai and Riaz A. Usmani. A recursion relation for simple rational B-splines. BIT, 27(2):282– 284, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=282.

Anonymous:1987:Pb

[1599] Anonymous. Problems. BIT, 27(2):

285–287, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=285.

Anonymous:1987:Nb

[1600] Anonymous. Notes. BIT, 27(2):

288, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=2&spage=288.

Anonymous:1987:NBJb

[1601] Anonymous. New books and journals. BIT, 27(2):288, June 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=27&
issue=2&spage=288.

Belik:1987:DAM

[1602] Ferenc Belik. Deadlock avoidance with a modified banker’s algorithm. BIT, 27(3):290–305, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=3&spage=290.

Chang:1987:PTD

[1603] C. C. Chang and C. Y. Chen. Performance of two-disk partition data allocations. BIT, 27(3):306–314, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=3&spage=306.

Hasham:1987:BMM

[1604] A. Hasham and J.-R. Sack. Bounds for min-max heaps. BIT, 27(3):315–323, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=3&spage=315.

Munro:1987:SMI

[1605] J. Ian Munro and Patricio V. Poblete. Searchability in merging and implicit data structures. BIT, 27(3):324–329, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=3&spage=324.

Makinen:1987:TS

[1606] Erkki Mäkinen. On top-down splaying. BIT, 27(3):330–339, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=330.

Noga:1987:SPD

[1607] M. T. Noga. Sorting in parallel by double distributive partitioning. BIT, 27(3):340–348, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=340. See

[3751].

Auzinger:1987:ICI

[1608] W. Auzinger and J. P. Monnet. IDeC — convergence independent of error asymptotics. BIT, 27(3):350–367, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=350.

Dubeau:1987:CRP

[1609] François Dubeau and Jean Savoie.

On consistency relations for polynomial splines on a uniform partition. BIT, 27(3):368–373, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=368.

Fuchs:1987:SSC

[1610] P. M. Fuchs. On the stability of spline-collocation methods of multivalue type. BIT, 27(3):374–388, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=374.

Gautschi:1987:NES

[1611] Walter Gautschi, M. A. Kovačević, and Gradimir V. Milovanović. The numerical evaluation of singular integrals with coth-kernel. BIT, 27(3):389–402, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=389.

Muir:1987:RAS

[1612] P. H. Muir and W. H. Enright. Relationships among some classes of implicit Runge–Kutta methods and their stability functions. BIT, 27(3):403–423, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=403.

Spijker:1987:NCN

[1613] M. N. Spijker. A note on contractivity in the numerical solution of initial value problems. BIT, 27(3):424–437, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=424.

Anonymous:1987:Pc

[1614] Anonymous. Problems. BIT, 27(3):

438–439, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=438.

Anonymous:1987:NBJc

[1615] Anonymous. New books and journals. BIT, 27(3):439, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=439.

Anonymous:1987:Nc

[1616] Anonymous. Notes. BIT, 27(3):440, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=3&spage=440.

Bassiouni:1987:ARR

[1617] M. A. Bassiouni and U. Khamare.

Algorithms for reducing rollbacks in concurrency control by certification. BIT, 27(4):441–457, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=441.

Joe:1987:CLV

[1618] B. Joe and R. B. Simpson. Corrections to Lee’s visibility polygon algorithm. BIT, 27(4):458–473, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=458. See

[3750].

Lingas:1987:AML

[1619] Andrzej Lingas, Christos Levcopoulos, and Jörg Sack. Algorithms for minimum length partitions of polygons. BIT, 27(4):474–479, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=474.

Melhem:1987:VCS

[1620] Rami Melhem. Verification of a class of self-timed computational networks. BIT, 27(4):480–500, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=480.

Bao:1987:IIC

[1621] Paul G. Bao and Jon Rokne. Inclusion isotonicity of circular complex centered forms. BIT, 27(4):502–509, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=502.

Butcher:1987:EAS

[1622] J. C. Butcher. The equivalence of algebraic stability and AN-stability. BIT, 27(4):510–533, December 1987. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=27&issue=4&spage=
510.

Hansen:1987:TSM

[1623] Per Christian Hansen. The truncated SVD as a method for regularization. BIT, 27(4):534–553, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=534.

Miekkala:1987:SCS

[1624] Ulla Miekkala and Olavi Nevanlinna.

Sets of convergence and stability regions. BIT, 27(4):554–584, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=554.

Ruhe:1987:CNM

[1625] Axel Ruhe. Closest normal matrix finally found! BIT, 27(4):585–598, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=585.

Thomas:1987:EFO

[1626] R. M. Thomas. Efficient fourth order P -stable formulae. BIT, 27(4):599–614, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=4&spage=599.

Usmani:1987:QSI

[1627] Riaz A. Usmani. On quadratic spline interpolation. BIT, 27(4):615–622, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=4&spage=615.

Raina:1987:EBO

[1628] B. L. Raina, Nancy Kaul, and Sunita Warikoo. Error bound for an osculatory quadrature formula. BIT, 27(4):623–626, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=27&issue=4&spage=623.

Cuminato:1987:EUC

[1629] J. A. Cuminato. Erratum: “On the uniform convergence of a collocation method for a class of singular integral equations” [BIT 27(2), 1987, pp. 190–202]. BIT, 27(4):627, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=627. See

[1592].

Pawagi:1987:EMW

[1630] S. Pawagi. Erratum: “Maximum weight independent set in trees” [BIT 27(2), 1987, pp. 170– 180]. BIT, 27(4):627, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=627. See

[1589].

Anonymous:1987:Pd

[1631] Anonymous. Problems. BIT, 27(4):

628–629, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=628.

Anonymous:1987:NBJd

[1632] Anonymous. New books and journals. BIT, 27(4):630, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=630.

Anonymous:1987:Nd

[1633] Anonymous. Notes. BIT, 27(4):631, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=631.

Anonymous:1987:A

[1634] Anonymous. Acknowledgements. BIT, 27(4):632, December 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=27&issue=4&spage=632.

Chaudhuri:1988:FPG

[1635] Pranay Chaudhuri. Fast parallel graph searching with applications. BIT, 28(1):1–18, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=28&
issue=1&spage=1.

Karlsson:1988:GMG

[1636] Rolf G. Karlsson. Greedy matching on a grid. BIT, 28(1):19–26, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=19.

Manolopoulos:1988:EDH

[1637] Y. P. Manolopoulos and J. G. Kollias. Estimating disk head movement in batched searching. BIT, 28(1):27– 36, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=27.

Rao:1988:DDS

[1638] Nageswara S. V. Rao, Vijay K.

Vaishnavi, and S. Sitharama Iyengar. On the dynamization of data structures. BIT, 28(1):37–53, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=37.

Wang:1988:OAT

[1639] Ling-Ling Wang and Wen-Hsiang Tsai.

Optimal assignment of task modules with precedence for distributed processing by graph matching and state-space search. BIT, 28(1):54– 68, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=54.

Bales:1988:ARA

[1640] Laurence A. Bales, Ohannes A.

Karakashian, and Steven M. Serbin. On the A_{0}-acceptability of rational approximations to the exponential function with only real poles. BIT, 28(1): 69–79, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=69.

Baumann:1988:OCF

[1641] Eckart Baumann. Optimal centered forms. BIT, 28(1):80–87, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=80.

Chawla:1988:FOM

[1642] M. M. Chawla, R. Subramanian, and H. L. Sathi. A fourth order method for a singular two-point boundary value problem. BIT, 28(1):88– 97, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=88.

Cuyt:1988:MQL

[1643] Annie Cuyt. A multivariate QDlike algorithm. BIT, 28(1):98–112, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=98.

Eirola:1988:ICO

[1644] Timo Eirola. Invariant curves of one-step methods. BIT, 28(1):113– 122, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=113.

Ekblom:1988:NAH

[1645] Håkan Ekblom. A new algorithm for the Huber estimator in linear models. BIT, 28(1):123–132, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=123.

Higham:1988:SPP

[1646] Nicholas J. Higham. The symmetric Procrustes problem. BIT, 28(1):133– 143, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=133.

Kreiss:1988:CSS

[1647] Gunilla Kreiss. Convergence to steady state of solutions of the Euler equations. I. BIT, 28(1):144– 162, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=144.

Smolarski:1988:OIM

[1648] Dennis C. Smolarski and Paul E. Saylor. An optimum iterative method for solving any linear system with a square matrix. BIT, 28(1):163– 178, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=163.

Mannila:1988:NLE

[1649] Heikki Mannila and Derick Wood.

A note on the largest empty rectangle problem. BIT, 28(1):179–183, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=179.

Schneid:1988:SPC

[1650] Josef Schneid. Stability properties of collocation methods. BIT, 28(1):184– 187, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=184.

Riesel:1988:EMT

[1651] Hans Riesel. Errata: “On the metric theory of nearest integer continued fractions” [BIT 27(2), 1987, pp. 248–263; MR 88k:11048]. BIT, 28(1): 188, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=188. See

[1596].

Anonymous:1988:Pa

[1652] Anonymous. Problems. BIT, 28 (1):189–191, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=1&spage=189.

Anonymous:1988:Na

[1653] Anonymous. Notes. BIT, 28(1): 192, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=1&spage=192.

Anonymous:1988:NBJa

[1654] Anonymous. New books and journals. BIT, 28(1):192, March 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=28&
issue=1&spage=192.

Anonymous:1988:NBP

[1655] Anonymous. Notes: BIT Prize for Young Nordic Scientists. BIT, 28 (1):192, 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Aurenhammer:1988:LST

[1656] F. Aurenhammer. On-line sorting of twisted sequences in linear time. BIT, 28(2):194–204, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=194.

Chang:1988:APC

[1657] C. C. Chang. Application of principal component analysis to multi-disk concurrent accessing. BIT, 28(2):205– 214, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=205.

Huang:1988:VPA

[1658] Shou-Hsuan Stephen Huang. A VLSI partition algorithm. BIT, 28(2):215– 226, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=215.

Karlsson:1988:SAG

[1659] Rolf G. Karlsson and Mark H. Overmars. Scanline algorithms on a grid. BIT, 28(2):227–241, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=227.

Nicolaou:1988:REP

[1660] P. P. Nicolaou and C. C. Lefas. On the recovery of error patterns from signatures obtained in digital systems testing. BIT, 28(2):242– 252, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=242.

Andrew:1988:CFE

[1661] Alan L. Andrew. Correction of finite element eigenvalues for problems with natural or periodic boundary conditions. BIT, 28(2):254– 269, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=254.

Gustafsson:1988:PSC

[1662] Kjell Gustafsson, Michael Lundh, and Gustaf Söderlind. A PI stepsize control for the numerical solution of ordinary differential equations. BIT, 28(2): 270–287, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=270.

Irodotou-Ellina:1988:QSC

[1663] M. Irodotou-Ellina and E. N. Houstis.

An O(h^{6}) quintic spline collocation method for fourth order two-point boundary value problems. BIT, 28(2): 288–301, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=288.

Jain:1988:MSB

[1664] M. K. Jain. A modification of the Stiefel–Bettis method for nonlinearly damped oscillators. BIT, 28(2):302– 307, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=302.

Kaasschieter:1988:PTC

[1665] E. F. Kaasschieter. A practical termination criterion for the conjugate gradient method. BIT, 28(2):308– 322, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=308.

Pardalos:1988:QPD

[1666] Panos M. Pardalos. Quadratic problems defined on a convex hull of points. BIT, 28(2):323–328, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=323.

Sakai:1988:SPA

[1667] Manabu Sakai and Riaz A. Usmani. A shape preserving area true approximation of histogram by rational splines. BIT, 28(2):329– 339, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=329.

Schmidt:1988:PCP

[1668] Jochen W. Schmidt and Walter Hess. Positivity of cubic polynomials on intervals and positive spline interpolation. BIT, 28(2):340– 352, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=340.

Chen:1988:OTA

[1669] G. H. Chen, M. T. Kuo, and J. P.

Sheu. An optimal time algorithm for finding a maximum weight independent set in a tree. BIT, 28(2):353– 356, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=353.

Cools:1988:EPC

[1670] Ronald Cools and Ann Haegemans. An embedded pair of cubature formulae of degree 5 and 7 for the triangle. BIT, 28(2):357– 359, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=357.

Forster:1988:CQV

[1671] Klaus-Jürgen Förster. On Chebyshev quadrature and variance of quadrature formulas. BIT, 28(2):360– 363, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=2&spage=360.

Knott:1988:LOA

[1672] Gary D. Knott. Linear open addressing and Peterson’s theorem rehashed. BIT, 28(2):364– 371, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=364.

Anonymous:1988:Pb

[1673] Anonymous. Problems. BIT, 28(2): 372–374, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=372.

Anonymous:1988:Nb

[1674] Anonymous. Notes. BIT, 28(2): 375, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=2&spage=375.

Anonymous:1988:NBJb

[1675] Anonymous. New books and journals. BIT, 28(2):375, June 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=28&
                                        issue=2&spage=375.

Dijkstra:1988:BOT

[1676] Edsger W. Dijkstra. On binary operators and their derived relations. BIT, 28(3):377–382, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=377.

Einarsson:1988:TAM

[1677] Bo Einarsson. Tutorial on analytical methods as a complement to numerical computing. BIT, 28(3):383–390, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=383.

Erdos:1988:ACC

[1678] Paul Erdős and Hans Riesel. On admissible constellations of consecutive primes. BIT, 28(3):391–396, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=391.

Ershov:1988:BCA

[1679] Andrei P. Ershov. Basic concepts of algorithms and programming to be taught in a school course in informatics. BIT, 28(3):397–405, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=397.

Froberg:1988:CPR

[1680] Carl-Erik Fröberg. On a combinatorial problem related to permanents. BIT, 28(3):406–411, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=406.

Mayoh:1988:F

[1681] Brian Mayoh. On formalisms. BIT, 28 (3):412–426, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=412.

Ravn:1988:DME

[1682] Anders P. Ravn, Hans Rischel, and Hans Henrik Løvengreen. A design method for embedded software systems. BIT, 28(3):427–438, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=427.

Smillie:1988:ATN

[1683] K. W. Smillie. Array theory and the Nial programming language. BIT, 28(3):439–449, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=439.

Sveinsdottir:1988:DCT

[1684] Edda Sveinsdottir and Erik Frøkjær. Datalogy — the Copenhagen tradition of computer science. BIT, 28(3): 450–472, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=450.

Turski:1988:TCI

[1685] W. M. Turski. Time considered irrelevant for real-time systems. BIT, 28(3):473–486, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=473.

Aasa:1988:ECS

[1686] Annika Aasa, Sören Holmström, and Christina Nilsson. An efficiency comparison of some representations of purely functional arrays. BIT, 28(3):489–503, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=489.

deBakker:1988:MSC

[1687] J. W. de Bakker and J.-J. Ch.

Meyer. Metric semantics for concurrency. BIT, 28(3):504–529, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=504.

Blum:1988:SCP

[1688] Edward K. Blum. The semantics and complexity of parallel programs for vector computations. Part I: a case study using ADA. BIT, 28(3):530–551, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=530.

Clausen:1988:AAA

[1689] Jens Clausen and Jakob Krarup. Arranging apples in an array. BIT, 28 (3):552–568, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=552.

Gries:1988:GRC

[1690] David Gries and Jin Yun Xue. Generating a random cyclic permutation. BIT, 28(3):569–572, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=569.

Johansen:1988:IIU

[1691] Peter Johansen. Inductive inference of ultimately periodic sequences. BIT, 28(3):573–580, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=3&spage=573.

Kurki-Suonio:1988:DRS

[1692] R. Kurki-Suonio and T. Kankaanpää. On the design of reactive systems. BIT, 28(3):581–604, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=581.

Nordstrom:1988:TGR

[1693] Bengt Nordström. Terminating general recursion. BIT, 28(3):605–619, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=605.

Staunstrup:1988:HLD

[1694] J. Staunstrup and M. R. Greenstreet. From high-level descriptions to VLSI circuits. BIT, 28(3):620–638, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=620.

Stewart:1988:SLD

[1695] Alan Stewart. SIMD language design using prescriptive semantics. BIT, 28(3):639–649, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=639.

Bales:1988:SRA

[1696] Laurence A. Bales, Ohannes A.

Karakashian, and Steve M. Serbin. On the stability of rational approximations to the cosine with only imaginary poles. BIT, 28(3):651–658, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=651.

Bjorck:1988:BAS

[1697] Åke Björck. A bidiagonalization algorithm for solving large and sparse ill-posed systems of linear equations. BIT, 28(3):659–670, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=659.

Christiansen:1988:UDS

[1698] Hanne Dalgas Christiansen. Uniformly distributed sums of Bernoulli variables. BIT, 28(3):671–677, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=671.

Hairer:1988:ERK

[1699] E. Hairer, Ch. Lubich, and M. Roche. Error of Runge–Kutta methods for stiff problems studied via differential algebraic equations. BIT, 28(3):678–700, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=678.

Korner:1988:NBR

[1700] Frank Körner. A new branching rule for the branch and bound algorithm for solving nonlinear integer programming problems. BIT, 28(3):701–708, September 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=3&spage=701.

Phua:1988:SSL

[1701] Kang Hoh Phua. Solving sparse linear systems by an ABS-method that corresponds to LU-decomposition. BIT, 28 (3):709–718, 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1988:Pc

[1702] Anonymous. Problems. BIT, 28(3):

719–724, 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bruhn:1988:PNH

[1703] Erik Bruhn, Carl-Erik Fröberg, and Christian Gram. Peter Naur on his 60th birthday, October 25, 1988. BIT, 28(3):i–iii, 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Beck:1988:SIS

[1704] István Beck and Stein Krogdahl.

A select and insert sorting algorithm. BIT, 28(4):725–735, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=725.

Faloutsos:1988:SFI

[1705] Christos Faloutsos. Signature files: an integrated access method for text and attributes, suitable for optical disk storage. BIT, 28(4):736–754, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=736.

Gibson:1988:GBA

[1706] J. K. Gibson. A generalisation of Brickell’s algorithm for fast modular multiplication. BIT, 28(4):755–763, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=755.

Richards:1988:DCS

[1707] Dana Richards and Pravin Vaidya. On the distribution of comparisons in sorting algorithms. BIT, 28(4):764–774, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=764.

Skiena:1988:ELM

[1708] Steven S. Skiena. Encroaching lists as a measure of presortedness. BIT, 28(4):775–784, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=775.

Stojmenovic:1988:PDC

[1709] Ivan Stojmenović and Michael A.

Langston. On a proposed divideand-conquer minimal spanning tree algorithm. BIT, 28(4):785–789, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=785.

Cools:1988:WDM

[1710] Ronald Cools and Ann Haegemans. Why do so many cubature formulae have so many positive weights. BIT, 28(4):791–802, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=791.

Denzer:1988:OML

[1711] V. Denzer and A. Ecker. Optimal multipliers for linear congruential pseudorandom number generators with prime moduli. BIT, 28(4):803–808, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=803.

Gekeler:1988:IRK

[1712] E. W. Gekeler. On implicit Runge– Kutta methods with higher derivatives. BIT, 28(4):809–816, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=809.

Jain:1988:SSM

[1713] R. K. Jain. Single step methods for general second order singular initial value problems with spherical symmetry. BIT, 28(4):817–824, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=817.

Liu:1988:SSE

[1714] M. Z. Liu and J. F. B. M. Kraaijevanger. On the solvability of the systems of equations arising in implicit Runge–Kutta methods. BIT, 28(4):825–838, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=825.

Riesel:1988:SSC

[1715] Hans Riesel. Some soluble cases of the discrete logarithm problem. BIT, 28(4):839–851, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=839.

Rokne:1988:NCF

[1716] Jon Rokne and Paul G. Bao. The number of centered forms for a polynomial. BIT, 28(4):852–866, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=852.

Sakai:1988:NSW

[1717] M. Sakai and Riaz A. Usmani. Nonpolynomial splines and weakly singular two-point boundary value problems. BIT, 28(4):867–876, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=28&issue=4&spage=867.

Sanz-Serna:1988:RKS

[1718] J. M. Sanz-Serna. Runge–Kutta schemes for Hamiltonian systems. BIT, 28(4):877–883, December 1988. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=28&issue=4&spage=
877.

Schmitt:1988:SIR

[1719] Bernhard A. Schmitt. Stability of implicit Runge–Kutta methods for nonlinear stiff differential equations. BIT, 28(4):884–897, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=884.

Thomas:1988:ESO

[1720] R. M. Thomas. Efficient sixth order methods for nonlinear oscillation problems. BIT, 28(4):898–903, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=898.

Anonymous:1988:Pd

[1721] Anonymous. Problems. BIT, 28(4): 904–906, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=904.

Anonymous:1988:NBJc

[1722] Anonymous. New books and journals. BIT, 28(4):907, December 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=28&issue=4&spage=907.

Anonymous:1988:A

[1723] Anonymous. Acknowledgements. BIT, 28(4):908–909, December 1988. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=28&issue=4&spage=
908.

Anonymous:1988:I

[1724] Anonymous. Index 1988. BIT, 28 (4):909–912, 1988. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gori:1989:AAC

[1725] M. Gori and G. Soda. An algebraic approach to Cichelli’s perfect hashing. BIT, 29(1):2–13, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=2.

Gupta:1989:PDG

[1726] P. Gupta and G. P. Bhattacharjee. A parallel derangement generation algorithm. BIT, 29(1):14– 22, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=14.

Lin:1989:SGC

[1727] Chau-Jy Lin and Jong Chuang Tsay. A systolic generation of combinations. BIT, 29(1):23–36, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=23.

Selmer:1989:SFP

[1728] Ernst S. Selmer. On Shellsort and the Frobenius problem. BIT, 29(1):37– 40, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=37.

Veklerov:1989:DSM

[1729] Eugene Veklerov. On the decomposability of the stable marriage problem. BIT, 29(1):41–46, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=41.

Christiansen:1989:ECO

[1730] Edmund Christiansen and Henrik Gordon Petersen. Estimation of convergence orders in repeated Richardson extrapolation. BIT, 29(1):48– 59, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=48.

Ekblom:1989:ANH

[1731] Håkan Ekblom and Kaj Madsen. Algorithms for nonlinear Huber estimation. BIT, 29(1):60–76, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=29&
issue=1&spage=60.

Hairer:1989:ERM

[1732] E. Hairer, Ch. Lubich, and M. Roche.

Error of Rosenbrock methods for stiff problems studied via differential algebraic equations. BIT, 29(1):77– 90, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=77.

Keeling:1989:IRK

[1733] Stephen L. Keeling. On implicit Runge–Kutta methods with a stability function having distinct real poles. BIT, 29(1):91–109, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=91.

Meadows:1989:SNR

[1734] Catherine Meadows and George Purdy.

Summing over a network without revealing summands. BIT, 29(1):110– 125, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=110.

Muir:1989:NEE

[1735] P. H. Muir and P. W. Beame. A note on error expressions for reflected and averaged implicit Runge– Kutta methods. BIT, 29(1):126– 139, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=1&spage=126.

Sakai:1989:PIR

[1736] M. Sakai and J. W. Schmidt. Positive interpolation with rational splines. BIT, 29(1):140–147, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=1&spage=140.

Sorevik:1989:FSI

[1737] Tor Sørevik and Terje O. Espelid. Fully symmetric integration rules for the 4-cube. BIT, 29(1):148–153, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=1&spage=148.

Neuman:1989:SPI

[1738] Edward Neuman. Short proof of an identity of Gaffney. BIT, 29(1):155– 156, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=1&spage=155.

Anonymous:1989:Pa

[1739] Anonymous. Problems. BIT, 29 (1):157–158, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=157.

Anonymous:1989:NBJ

[1740] Anonymous. New books and journals. BIT, 29(1):159, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=29&
issue=1&spage=159.

Anonymous:1989:Na

[1741] Anonymous. Notes. BIT, 29(1):

160, March 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=1&spage=160.

Anonymous:1989:I

[1742] Anonymous. Index 1980–1988. BIT, 29 (1):161–181, 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gropp:1989:RMR

[1743] W. D. Gropp and I. C. F. Ipsen.

Recursive mesh refinement on hypercubes. BIT, 29(2):186–211, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=186.

Huang:1989:CSA

[1744] Nen Fu Huang and Tsing Her Huang.

On the complexity of some arborescences finding problems on a multihop radio network. BIT, 29(2):212– 216, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=212.

Huang:1989:LSA

[1745] Shing Tsaan Huang and Ming-Shin Tsai. A linear systolic algorithm for the connected component problem. BIT, 29(2):217– 226, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=217.

Jokinen:1989:ARI

[1746] Matti O. Jokinen. Automatically replacing indices into parallel arrays with pointers to records. BIT, 29(2):227– 238, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=227.

Quinn:1989:ABT

[1747] Michael J. Quinn. Analysis and benchmarking of two parallel sorting algorithms: Hyperquicksort and Quick Merge. BIT, 29(2):239– 250, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=239.

Strothotte:1989:NCM

[1748] Thomas Strothotte, Patrik Eriksson, and Sören Vallner. A note on constructing min-max heaps. BIT, 29(2): 251–256, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=251.

Crisci:1989:SOS

[1749] M. R. Crisci, E. Russo, and A. Vecchio. On the stability of the one-step exact collocation methods for the numerical solution of the second kind Volterra integral equation. BIT, 29(2): 258–269, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=258.

Eiermann:1989:AGI

[1750] Martin C. Eiermann. Automatic, guaranteed integration of analytic functions. BIT, 29(2):270–282, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=270.

Espelid:1989:DNE

[1751] Terje O. Espelid and Tor Sørevik. A discussion of a new error estimate for adaptive quadrature. BIT, 29(2):283– 294, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=283.

Fuchs:1989:SSC

[1752] P. M. Fuchs. A-stable splinecollocation methods of multivalue type. BIT, 29(2):295–310, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=29&
                                        issue=2&spage=295.

Hadjidimos:1989:PDC

[1753] A. Hadjidimos and M. Neumann. Precise domains of convergence for the block SSOR method associated with p-cyclic matrices. BIT, 29(2):311– 320, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=311.

Kuang:1989:SIR

[1754] Jiao Xun Kuang and Jia Xiang Xiang. On the D-suitability of implicit Runge–Kutta methods. BIT, 29(2): 321–327, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=321.

Nevanlinna:1989:RPLa

[1755] Olavi Nevanlinna. Remarks on Picard– Lindelöf iteration: Part I. BIT, 29(2): 328–346, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=328.

Rabinowitz:1989:IPR

[1756] Philip Rabinowitz. On an interpolatory product rule for evaluating Cauchy principal value integrals. BIT, 29(2): 347–355, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=347.

Scraton:1989:PUE

[1757] R. E. Scraton. The practical use of the Euler transformation. BIT, 29(2): 356–360, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=356.

Slough:1989:EAT

[1758] W. Slough and K. Efe. Efficient algorithms for tree reconstruction. BIT, 29(2):361–363, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=2&spage=361.

Manolopoulos:1989:ARF

[1759] Y. Manolopoulos and G. Poulakis.

An adaptation of a root-finding method to searching ordered disk files revisited. BIT, 29(2):364– 369, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=364. See

[1482].

Lindberg:1989:VSC

[1760] Bengt Lindberg. Variable stepsize in Cowell’s method. BIT, 29(2):369– 371, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=369.

Rohn:1989:ARL

[1761] J. Rohn. An asymptotic result for linear interval systems. BIT, 29(2):372– 374, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=372.

Anonymous:1989:Pb

[1762] Anonymous. Problems. BIT, 29(2): 375–376, June 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=2&spage=375.

Anonymous:1989:Nb

[1763] Anonymous. Notes. BIT, 29(2):376, 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Baeza-Yates:1989:TAW

[1764] Ricardo A. Baeza-Yates. A trivial algorithm whose analysis is not: a continuation. BIT, 29(3):378–394, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=378. See

[3752].

Galle:1989:BSS

[1765] Per Galle. Branch & sample:

a simple strategy for constraint satisfaction. BIT, 29(3):395–408, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=395.

Melhem:1989:ASN

[1766] Rami Melhem and Concettina Guerra.

The application of a sequence notation to the design of systolic computations. BIT, 29(3):409–427, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=409.

Scheuermann:1989:HFR

[1767] Peter Scheuermann, Young Chul Park, and Edward Omiecinski. A heuristic file reorganization algorithm based on record clustering. BIT, 29(3):428–447, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=428.

Tseng:1989:APP

[1768] S. S. Tseng. The average performance of a parallel stable marriage algorithm. BIT, 29(3):448–456, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=448.

Bi:1989:CRF

[1769] Zhi Yong Bi and Paul H. Calamai. A class of root finding methods. BIT, 29(3):458–463, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=458.

Bialecki:1989:MSQ

[1770] Bernard Bialecki. A modified sinc quadrature rule for functions with poles near the arc of integration. BIT, 29(3):464–476, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=464.

Galanis:1989:OFS

[1771] S. Galanis, A. Hadjidimos, and D. Noutsos. Optimum first and second order extrapolations of successive overrelaxation type methods for certain classes of matrices. BIT, 29(3):477–490, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=477.

Hansen:1989:RGT

[1772] Per Christian Hansen. Regularization, GSVD and truncated GSVD. BIT, 29(3):491–504, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=491.

Hundsdorfer:1989:ESC

[1773] W. H. Hundsdorfer and J. Schneid.

On the equivalence of BS-stability and B-consistency. BIT, 29(3):505–511, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=505.

Hunter:1989:ECP

[1774] D. B. Hunter and H. V. Smith.

The evaluation of Cauchy principal value integrals involving unknown poles. BIT, 29(3):512–517, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=512.

Jain:1989:SIR

[1775] R. K. Jain, Rajive Kumar, and Rakesh Goel. Superstable implicit Runge– Kutta methods for second order initial value problems. BIT, 29(3):518–526, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=518.

Lyness:1989:NLR

[1776] J. N. Lyness and T. Sørevik. The number of lattice rules. BIT, 29(3):527–534, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=527.

Nevanlinna:1989:RPLb

[1777] Olavi Nevanlinna. Remarks on Picard– Lindelöf iteration: Part II. BIT, 29 (3):535–562, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=3&spage=535.

Steidl:1989:SRR

[1778] Gabriele Steidl. On symmetric radix representation of Gaussian integers. BIT, 29(3):563–571, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=563.

Makinen:1989:CBT

[1779] Erkki Mäkinen. Constructing a binary tree from its traversals. BIT, 29(3): 572–575, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=572.

Anonymous:1989:JVG

[1780] Anonymous. Jan V. Garwick:

in memoriam. BIT, 29(3):576, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=576.

Axelsson:1989:PCG

[1781] Owe Axelsson. Preconditioned conjugate gradient methods. Preface. BIT, 29(4):577–582, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=577.

Ashby:1989:APP

[1782] Steven F. Ashby, Thomas A. Manteuffel, and Paul E. Saylor. Adaptive polynomial preconditioning for Hermitian indefinite linear systems. BIT, 29 (4):583–609, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=583.

Greenbaum:1989:OPG

[1783] A. Greenbaum and G. H. Rodrigue. Optimal preconditioners of a given sparsity pattern. BIT, 29(4):610–634, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=610.

Duff:1989:EOP

[1784] Iain S. Duff and Gérard A. Meurant. The effect of ordering on preconditioned conjugate gradients. BIT, 29(4):635–657, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=635.

Beauwens:1989:AFP

[1785] Robert Beauwens. Approximate factorizations with S/P consistently ordered M-factors. BIT, 29(4):658–681, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=658.

Notay:1989:IFS

[1786] Yvan Notay. Incomplete factorization of singular linear systems. BIT, 29(4):682–702, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=682.

Paolini:1989:DSV

[1787] Gaia Valeria Paolini and Giuseppe Radicati di Brozolo. Data structures to vectorize CG algorithms for general sparsity patterns. BIT, 29(4):703–718, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=703.

Wittum:1989:TID

[1788] Gabriel Wittum and Frank Liebau.

On truncated incomplete decompositions. BIT, 29(4):719–740, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=719.

Evans:1989:CBP

[1789] D. J. Evans and C. Li. Connections between the PCG method applied to systems with red/black form and its reduced system. BIT, 29(4):741–747, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=741.

Jung:1989:TLH

[1790] M. Jung, U. Langer, and U. Semmler. Two-level hierarchically preconditioned conjugate gradient methods for solving linear elasticity finite element equations. BIT, 29(4):748–768, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=4&spage=748.

Axelsson:1989:SMP

[1791] O. Axelsson and P. S. Vassilevski. A survey of multilevel preconditioned iterative methods. BIT, 29(4):769–793, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=4&spage=769.

Carey:1989:BFS

[1792] G. F. Carey and E. Barragy. Basis function selection and preconditioning high degree finite element and spectral methods. BIT, 29(4):794–804, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=4&spage=794.

Kolotilina:1989:BSP

[1793] L. Yu. Kolotilina and A. Yu. Yeremin. Block SSOR preconditionings for high order 3D FE systems. BIT, 29(4):805–823, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=805.

Kaasschieter:1989:GFE

[1794] E. F. Kaasschieter. A general finite element preconditioning for the conjugate gradient method. BIT, 29(4):824–849, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=824.

Ewing:1989:PCG

[1795] Richard E. Ewing. Preconditioned conjugate gradient methods for large-scale fluid flow applications. BIT, 29(4):850–866, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=850.

Axelsson:1989:IBM

[1796] O. Axelsson, V. Eijkhout, B. Polman, and P. Vassilevski. Incomplete block-matrix factorization iterative methods for convection-diffusion problems. BIT, 29(4):867–889, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=867.

Elman:1989:RSI

[1797] Howard C. Elman. Relaxed and stabilized incomplete factorizations for nonselfadjoint linear systems. BIT, 29 (4):890–915, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=4&spage=890.

Aarden:1989:PCT

[1798] J. M. C. Aarden and K. E. Karlsson. Preconditioned CG-type methods for solving the coupled system of fundamental semiconductor equations. BIT, 29(4):916–937, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=4&spage=916.

Bank:1989:ABF

[1799] R. E. Bank, T. F. Chan, W. M. Coughran, Jr., and R. K. Smith. The alternate-block-factorization procedure for systems of partial differential equations. BIT, 29(4):938–954, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=29&issue=4&spage=938.

Anonymous:1989:A

[1800] Anonymous. Acknowledgements. BIT, 29(4):955–956, December 1989. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=29&issue=4&spage=
955.

Anonymous:1989:BR

[1801] Anonymous. Books received. BIT, 29(4):956, December 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=4&spage=956.

Andersen:1990:GTM

[1802] N. Andersen. A general transposition method for a matrix on auxiliary store. BIT, 30(1):2–16, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=2.

Bauslaugh:1990:GAP

[1803] Bruce Bauslaugh and Frank Ruskey.

Generating alternating permutations lexicographically. BIT, 30(1):17–26, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=17.

Chang:1990:RCT

[1804] Chin-Chen Chang and Chu-Hsing Lin.

A reciprocal confluence tree unit and its applications. BIT, 30(1):27–33, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=1&spage=27.

Jajodia:1990:LOJ

[1805] Sushil Jajodia and Frederick N. Springsteel. Lossless outer joins with incomplete information. BIT, 30(1):34– 41, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=1&spage=34.

Lang:1990:ARB

[1806] S. D. Lang. Analysis of recursive batched interpolation search. BIT, 30(1):42–50, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=1&spage=42.

Robison:1990:IRN

[1807] Arch D. Robison. An improved rewriting-number algorithm. BIT, 30(1):51–61, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=1&spage=51.

Sprugnoli:1990:CLB

[1808] Renzo Sprugnoli. Counting labels in binary trees. BIT, 30(1):62– 69, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=62.

Stewart:1990:ATS

[1809] Alan Stewart. An axiomatic treatment of SIMD assignment. BIT, 30(1):70– 82, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=70.

Syslo:1990:TPE

[1810] Maciej M. Sysło and Pawel Winter. In-trees and plane embeddings of outerplanar graphs. BIT, 30(1):83– 90, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=83.

Argyros:1990:EMS

[1811] Ioannis K. Argyros. Error for the modified secant method. BIT, 30(1):92– 100, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=92.

Barrlund:1990:PBP

[1812] Anders Barrlund. Perturbation bounds on the polar decomposition. BIT, 30(1):101–113, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=

article&issn=0006-3835&volume=30&
issue=1&spage=101.

Hsu:1990:QMC

[1813] L. C. Hsu, R. J. Tomkins, and Chung-Lie Wang. A quadrature method for a class of strongly oscillatory infinite integrals. BIT, 30(1):114– 125, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=114.

Lie:1990:LEE

[1814] Ivar Lie. Local error estimation for multistep collocation methods. BIT, 30(1):126–144, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=30&
issue=1&spage=126.

Peherstorfer:1990:WFA

[1815] Franz Peherstorfer. Weight functions admitting repeated positive Kronrod quadrature. BIT, 30(1):145– 151, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=145.

Ehrenmark:1990:NEE

[1816] Ulf Torsten Ehrenmark. A note on an extension of extrapolative techniques for a class of infinite oscillatory integrals. BIT, 30(1):152– 155, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=152.

Legault:1990:CCP

[1817] R. Legault and T. D. Bui. Comparing the contractivity properties of semi-implicit methods. BIT, 30(1):156– 160, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=156.

Rohn:1990:NSS

[1818] J. Rohn. On nonconvexity of the solution set of a system of linear interval equations. BIT, 30(1):161– 165, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=161.

Schneid:1990:NCC

[1819] J. Schneid. A necessary condition for B-convergence of Runge– Kutta methods. BIT, 30(1):166– 170, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=166.

Stojanovic:1990:NSS

[1820] Mirjana Stojanović. Numerical solution of a singularly perturbed problem via exponential splines. BIT, 30(1):171– 176, March 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=1&spage=171.

Douglas:1990:MPA

[1821] Craig C. Douglas and Willard L. Miranker. The multilevel principle applied to sorting. BIT, 30(2):177– 195, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=177.

Huang:1990:DCA

[1822] Nen Fu Huang. A divide-and-conquer algorithm for constructing relative neighborhood graph. BIT, 30(2):196– 206, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=196.

Huang:1990:CWT

[1823] Shou-Hsuan Huang and Venkatraman Viswanathan. On the construction of weighted time-optimal B-trees. BIT, 30(2):207–215, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=30&
issue=2&spage=207.

Manolopoulos:1990:ODP

[1824] Yannis Manolopoulos and J. (Yannis) G. Kollias. Optimal data placement in two-headed disk systems. BIT, 30(2):216–219, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=30&
issue=2&spage=216.

Munro:1990:SSS

[1825] J. Ian Munro, Venkatesh Raman, and Jeffrey S. Salowe. Stable in situ sorting and minimum data movement. BIT, 30(2):220–234, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=220.

Winters:1990:MPH

[1826] Vincent G. Winters. Minimal perfect hashing in polynomial time. BIT, 30(2):235–244, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=235.

Zubair:1990:EBC

[1827] M. Zubair and S. N. Gupta. Embeddings on a Boolean cube. BIT, 30(2): 245–256, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=245.

Bhavsar:1990:EDL

[1828] Virendra C. Bhavsar, Uday G. Gujar, Joseph D. Horton, and Lambros A. Lambrou. Evaluation of the discrepancy of the linear congruential pseudorandom number sequences. BIT, 30(2): 257–267, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=2&spage=257.

Egecioglu:1990:PMF

[1829] Ö. Eğecioğlu, E. Gallopoulos, and Ç. K. Koç. A parallel method for fast and practical high-order Newton interpolation. BIT, 30(2):268– 288, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=2&spage=268.

Flaam:1990:ROP

[1830] Sjur D. Flåm and Jochem Zowe. Relaxed outer projections, weighted averages and convex feasibility. BIT, 30(2): 289–300, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=2&spage=289.

Gurwitz:1990:WMA

[1831] Chaya Gurwitz. Weighted median algorithms for L_{1} approximation. BIT, 30(2):301–310, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=30&
issue=2&spage=301.

Kao:1990:ENC

[1832] Thomas Chiungtung Kao and Gary D.

Knott. An efficient and numerically correct algorithm for the 2D convex hull problem. BIT, 30(2):311– 331, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=311.

Reichel:1990:NIL

[1833] Lothar Reichel. Newton interpolation at Leja points. BIT, 30(2):332– 346, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=332.

Sidi:1990:RAE

[1834] Avram Sidi. On rates of acceleration of extrapolation methods for oscillatory infinite integrals. BIT, 30(2):347– 357, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=347.

Strehmel:1990:EBP

[1835] K. Strehmel, R. Weiner, and I. Dannehl. On error behaviour of partitioned linearly implicit Runge–Kutta methods for stiff and differential algebraic systems. BIT, 30(2):358– 375, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=358.

Anonymous:1990:BP

[1836] Anonymous. Bit prize 1990. BIT, 30 (2):376, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=2&spage=376.

Anonymous:1990:NBJa

[1837] Anonymous. New books and journals. BIT, 30(2):376, June 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=30&
issue=2&spage=376.

Fjallstrom:1990:SCH

[1838] Per-Olof Fjällström, Jyrki Katajainen, Christos Levcopoulos, and Ola Petersson. A sublogarithmic convex hull algorithm. BIT, 30(3):378–384, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=378.

Overmars:1990:SLS

[1839] Mark H. Overmars, Haijo Schipper, and Micha Sharir. Storing line segments in partition trees. BIT, 30(3): 385–403, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=385.

Ramamohanarao:1990:MAH

[1840] Kotagiri Ramamohanarao, John Shepherd, and Ron Sacks-Davis. Multiattribute hashing with multiple file copies for high performance partialmatch retrieval. BIT, 30(3):404–423, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=404.

Stojmenovic:1990:OAG

[1841] Ivan Stojmenović. An optimal algorithm for generating equivalence relations on a linear array of processors. BIT, 30(3):424–436, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=424.

Tarhio:1990:UAG

[1842] Jorma Tarhio. Uncle-attributed grammars. BIT, 30(3):437–449, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=437.

Varadarajan:1990:PME

[1843] Ravi Varadarajan. Partitioning multiedge graphs. BIT, 30(3):450–463, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=450.

Elden:1990:ACF

[1844] Lars Eldén. Algorithms for the computation of functionals defined on the solution of a discrete illposed problem. BIT, 30(3):466–483, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=466.

ElTarazi:1990:QSI

[1845] M. N. El Tarazi. Quadratic spline interpolation on uniform meshes. BIT, 30 (3):484–489, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=484.

Govaerts:1990:BEO

[1846] W. Govaerts and J. D. Pryce. Block elimination with one refinement solves bordered linear systems accurately. BIT, 30(3):490–507, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=490.

Herzberger:1990:UEB

[1847] J. Herzberger. Using error bounds for hyperpower methods to calculate inclusions for the inverse of a matrix. BIT, 30(3):508–515, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=508.

vanderMarel:1990:SRP

[1848] Roeland P. van der Marel. Stability radius of polynomials occurring in the numerical solution of initial value problems. BIT, 30(3):516–528, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=516.

Petras:1990:EEF

[1849] Knut Petras. Error estimates for Filon quadrature formulae. BIT, 30 (3):529–541, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=3&spage=529.

Sowa:1990:SRK

[1850] Johan Sowa. Stability of a Runge– Kutta method for the Navier–Stokes equation. BIT, 30(3):542–560, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=542.

Grazzini:1990:PES

[1851] E. Grazzini and F. Pippolini. Performance evaluation of shared and separate inverted files. BIT, 30(3):561–565, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=561.

Sakai:1990:ACI

[1852] Manabu Sakai and Riaz A. Usmani. An application of Chawla’s identity to a different scheme for singular problems. BIT, 30(3):566–568, September 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=3&spage=566.

Chen:1990:MPT

[1853] Chiu Yuan Chen and Ruei Chuan Chang. On the minimality of polygon triangulation. BIT, 30(4):570–582, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=4&spage=570.

Kokosinski:1990:GPT

[1854] Zbigniew Kokosiński. On generation of permutations through decomposition of symmetric groups into cosets. BIT, 30(4):583–591, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=583.

Kuo:1990:IMO

[1855] R. T. Kuo and S. S. Tseng. On the invariance of male optimal stable matching. BIT, 30(4):592–598, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=592.

Mannila:1990:UDT

[1856] Heikki Mannila and Esko Ukkonen.

Unifications, deunifications, and their complexity. BIT, 30(4):599–619, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=599.

Manolopoulos:1990:AEB

[1857] Yannis Manolopoulos and Christos Faloutsos. Analysis for the end of block wasted space. BIT, 30(4):620–630, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=620.

Tan:1990:MSM

[1858] Jimmy J. M. Tan. A maximum stable matching for the roommates problem. BIT, 30(4):631–640, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=631.

Caira:1990:CPR

[1859] R. Caira, C. Costabile, and F. Costabile. A class of pseudo Runge–Kutta methods. BIT, 30(4):642–649, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=642.

Csendes:1990:IMB

[1860] Tibor Csendes. Interval method for bounding level sets: revisited and tested with global optimization problems. BIT, 30(4):650–657, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=650.

Hansen:1990:DPC

[1861] Per Christian Hansen. The discrete Picard condition for discrete ill-posed problems. BIT, 30(4):658–672, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=658.

Lambert:1990:LEL

[1862] J. D. Lambert. On the local error and the local truncation error of linear multistep methods. BIT, 30(4):673–681, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=673.

Madsen:1990:FAR

[1863] Kaj Madsen and Hans Bruun Nielsen.

Finite algorithms for robust linear regression. BIT, 30(4):682–699, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=682.

Owren:1990:SSR

[1864] Brynjulf Owren and Kristian Seip.

Some stability results for explicit Runge–Kutta methods. BIT, 30(4): 700–706, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=700.

Pan:1990:MLD

[1865] C.-T. Pan. A modification to the Linpack downdating algorithm. BIT, 30 (4):707–722, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=707.

Price:1990:EPF

[1866] C. J. Price and I. D. Coope. An exact penalty function algorithm for semi-infinite programmes. BIT, 30(4): 723–734, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=4&spage=723.

Sakai:1990:OAP

[1867] Manabu Sakai and Riaz A. Usmani. On orders of approximation of plane curves by parametric cubic splines. BIT, 30(4):735–741, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=4&spage=735.

Zuhe:1990:SMP

[1868] Shen Zuhe, A. Neumaier, and M. C. Eiermann. Solving minimax problems by interval methods. BIT, 30(4):742–751, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=30&issue=4&spage=742.

Argyros:1990:SEN

[1869] Ioannis K. Argyros. On the solution of equations with nondifferentiable and Pták error estimates. BIT, 30(4):752–754, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=752.

Anonymous:1990:A

[1870] Anonymous. Acknowledgements. BIT, 30(4):755–756, December 1990. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=30&issue=4&spage=
755.

Anonymous:1990:NBJb

[1871] Anonymous. New books and journals. BIT, 30(4):756, December 1990. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=30&issue=4&spage=756.

Gunther:1991:DAD

[1872] Oliver Günther and Eugene Wong.

A dual approach to detect polyhedral intersections in arbitrary dimensions. BIT, 31(1):2–14, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=31&
issue=1&spage=2.

Johnsen:1991:GBT

[1873] Ben Johnsen. Generating binary trees with uniform probability. BIT, 31(1): 15–31, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=1&spage=15.

Konvalina:1991:BSS

[1874] John Konvalina and Yi-Hsin Liu. Bit strings without q-separation. BIT, 31(1):32–35, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=1&spage=32.

Vaishnavi:1991:HAD

[1875] Vijay K. Vaishnavi and Timothy D. Korson. Height-analysis of k-dimensional leaf and node height-balanced trees: a new approach. BIT, 31(1):36–49, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=31&
                                        issue=1&spage=36.

vandeVel:1991:MHS

[1876] H. van de Vel and Sun Shijie. A modification of Hochbaum and Shmoys’ algorithm for scheduling problems. BIT, 31(1):50–52, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=1&spage=50.

Yuan:1991:EFT

[1877] Shyan-Ming Yuan. An efficient faulttolerant decentralized commit protocol for single site failure. BIT, 31(1):53– 68, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=1&spage=53.

Zhang:1991:OPQ

[1878] Wei Xiong Zhang and Nageswara S. V. Rao. Optimal parallel quicksort on EREW PRAM. BIT, 31(1): 69–74, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=1&spage=69.

Bjorstad:1991:SSO

[1879] Petter E. Bjørstad and Jan Mandel.

On the spectra of sums of orthogonal projections with applications to parallel computing. BIT, 31(1):76– 88, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=1&spage=76.

Burmeister:1991:DAJ

[1880] W. Burmeister, Ch. Grossmann, and S. Scholz. Directional approximation of the Jacobians in ROW methods. BIT, 31(1):89–101, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=1&spage=89.

Favati:1991:LEE

[1881] Paola Favati, Grazia Lotti, and Francesco Romani. Local error estimates in quadrature. BIT, 31(1):102– 111, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=1&spage=102.

Fraigniaud:1991:DKP

[1882] Pierre Fraigniaud. The Durand–Kerner polynomials roots-finding method in case of multiple roots. BIT, 31(1):112– 123, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=1&spage=112.

Hundsdorfer:1991:CLM

[1883] W. H. Hundsdorfer and B. I. Steininger. Convergence of linear multistep and one-leg methods for stiff non-linear initial value problems. BIT, 31(1):124–143, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=1&spage=124.

Orel:1991:RPA

[1884] Bojan Orel. Real pole approximations to the exponential function. BIT, 31(1):144–159, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=31&
issue=1&spage=144.

Raptis:1991:FSP

[1885] A. D. Raptis and T. E. Simos. A fourstep phase-fitted method for the numerical integration of second order initial value problems. BIT, 31(1):160– 168, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=1&spage=160.

Stoffer:1991:ICV

[1886] Daniel Stoffer and Kaspar Nipp. Invariant curves for variable step size integrators. BIT, 31(1):169–180, March 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=1&spage=169. See

erratum [1964].

Arikati:1991:ERP

[1887] Srinivasa R. Arikati, C. Pandu Rangan, and Glenn K. Manacher. Efficient reduction for path problems on circular-arc graphs. BIT, 31(2):181– 193, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=181.

Atkinson:1991:RSS

[1888] M. D. Atkinson. The recursive structure of some ordering problems. BIT, 31(2):194–201, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=2&spage=194.

vanKreveld:1991:FSR

[1889] Marc J. van Kreveld and Mark T. de Berg. Finding squares and rectangles in sets of points. BIT, 31(2):202– 219, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=2&spage=202.

Liu:1991:NOM

[1890] L. T. Liu, Gen-Huey Chen, and Yue Li Wang. A note on optimal multiway split trees. BIT, 31(2):220– 229, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=2&spage=220.

Wang:1991:FMN

[1891] Cao An Wang. Finding minimal nested polygons. BIT, 31(2):230– 236, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=2&spage=230.

Bjorck:1991:CWP

[1892] Åke Björck. Component-wise perturbation analysis and error bounds for linear least squares solutions. BIT, 31(2): 237–244, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=237.

Ekolin:1991:FDM

[1893] Gunnar Ekolin. Finite difference methods for a nonlocal boundary value problem for the heat equation. BIT, 31(2): 245–261, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=245.

Elliott:1991:IEB

[1894] David Elliott and George M. Phillips.

Improved error bounds for nearminimax approximations. BIT, 31(2): 262–275, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=262.

Ioakimidis:1991:NEA

[1895] N. I. Ioakimidis, K. E. Papadakis, and E. A. Perdios. Numerical evaluation of analytic functions by Cauchy’s theorem. BIT, 31(2):276– 285, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=276.

Knopfmacher:1991:NSE

[1896] Arnold Knopfmacher and John Knopfmacher. The number of steps in the Euclidean algorithm over complex quadratic fields. BIT, 31(2):286– 292, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=2&spage=286.

Lubich:1991:RCS

[1897] Christian Lubich and Olavi Nevanlinna. On resolvent conditions and stability estimates. BIT, 31(2):293– 313, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=2&spage=293.

Muller:1991:ACS

[1898] Michael Müller. Algebraic characterization of I-stable Runge–Kutta methods. BIT, 31(2):314–320, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=31&
                                        issue=2&spage=314.

Rabinowitz:1991:PIS

[1899] Philip Rabinowitz. Product integration of singular integrands using Hermite– Fejér interpolation. BIT, 31(2):321– 327, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=321.

Schmidt:1991:CII

[1900] Jochen W. Schmidt. Convex interval interpolation with cubic splines. II. BIT, 31(2):328–340, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=31&
issue=2&spage=328.

Sun:1991:PBC

[1901] Ji Guang Sun. Perturbation bounds for the Cholesky and QR factorizations. BIT, 31(2):341–352, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=341.

Makinen:1991:LTS

[1902] Erkki Mäkinen. A linear time and space algorithm for finding isomorphic subtrees of a binary tree. BIT, 31(2):353– 357, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=353.

Barrlund:1991:PBD

[1903] Anders Barrlund. Perturbation bounds for the LDL^{H} and LU decompositions. BIT, 31(2):358– 363, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=358.

Carnicer:1991:LBN

[1904] J. M. Carnicer. A lower bound for the number of stages of an explicit continuous Runge–Kutta method to obtain convergence of given order. BIT, 31(2): 364–368, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=364.

Kvaalen:1991:FBM

[1905] Eric Kvaalen. A faster Broyden method. BIT, 31(2):369–372, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=2&spage=369.

Owren:1991:URR

[1906] Brynjulf Owren and Kristian Seip. A uniqueness result related to the stability of explicit Runge–Kutta methods. BIT, 31(2):373–374, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=31&
issue=2&spage=373.

Zha:1991:IFP

[1907] Hong Yuan Zha. Implicit QR factorization of a product of three matrices. BIT, 31(2):375–379, June 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=31&
issue=2&spage=375.

Gambosi:1991:SUP

[1908] Giorgio Gambosi, Giuseppe F. Italiano, and Maurizio Talamo. The set union problem with dynamic weighted backtracking. BIT, 31(3):381–393, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=381.

Grape:1991:CPS

[1909] Per Grape. Complete problems with L-samplable distributions. BIT, 31(3): 394–405, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=394.

Puustjarvi:1991:DMT

[1910] Juha Puustjärvi. Distributed management of transactions in heterogeneous distributed database systems. BIT, 31(3):406–420, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=406.

Schipper:1991:DPT

[1911] Haijo Schipper and Mark H. Overmars. Dynamic partition trees. BIT, 31(3): 421–436, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=421.

Gautschi:1991:QFH

[1912] Walter Gautschi. Quadrature formulae on half-infinite intervals. BIT, 31 (3):437–446, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=437.

Higham:1991:IRE

[1913] Nicholas J. Higham. Iterative refinement enhances the stability of QR factorization methods for solving linear equations. BIT, 31(3):447–468, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=447.

vanderHouwen:1991:SCB

[1914] P. J. van der Houwen, B. P. Sommeijer, and Nguyen Huu Cong. Stability of collocation-based Runge–Kutta– Nyström methods. BIT, 31(3):469–481, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=469.

Kraaijevanger:1991:CRK

[1915] J. F. B. M. Kraaijevanger. Contractivity of Runge–Kutta methods. BIT, 31(3):482–528, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=482.

Worley:1991:IL

[1916] R. T. Worley. On integration lattices. BIT, 31(3):529–539, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=529.

Kohler:1991:GCN

[1917] Peter Köhler. On a generalization of compound Newton–Cotes quadrature formulas. BIT, 31(3):540–544, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=540.

Lubich:1991:PRK

[1918] Ch. Lubich. On projected Runge– Kutta methods for differential-algebraic equations. BIT, 31(3):545–550, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=3&spage=545.

Spijker:1991:CLT

[1919] M. N. Spijker. On a conjecture by LeVeque and Trefethen related to the Kreiss matrix theorem. BIT, 31(3):551–555, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=3&spage=551.

Anonymous:1991:BPH

[1920] Anonymous. BIT Prize 1991 [to Heiki Mannila]. BIT, 31(3):556, September 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=3&spage=556.

Bartling:1991:PAA

[1921] Frank Bartling and Klaus Hinrichs. Probabilistic analysis of an algorithm for solving the k-dimensional all-nearest-neighbors problem by projection. BIT, 31(4):558–565, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=4&spage=558.

Chen:1991:OCD

[1922] C. Y. Chen and H. F. Lin. Optimality criteria of the disk modulo allocation method for Cartesian product files. BIT, 31(4):566–575, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=566.

Eloranta:1991:MNT

[1923] Jaana Eloranta. Minimizing the number of transitions with respect to observation equivalence. BIT, 31(4): 576–590, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=576.

Lingas:1991:UMP

[1924] Andrzej Lingas. An unfeasible matching problem. BIT, 31(4):591–597, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=591.

Teo:1991:TUB

[1925] Kim-Heng H. Teo and Tai-Ching C.

Tuan. A tight upper bound for the number of intersections between two rectangular paths. BIT, 31(4):598–606, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=598.

Abrahamsson:1991:EMM

[1926] Leif Abrahamsson and Heinz-Otto O.

Kreiss. Explicit methods for mildly stiff oscillatory systems. BIT, 31(4): 608–619, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=4&spage=608.

Banerjee:1991:AEF

[1927] Uday Banerjee. Approximation of the eigenvalues of a fourth order differential equation with non-smooth coefficients. BIT, 31(4):620–631, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=4&spage=620.

Chan:1991:CSC

[1928] Raymond H. Chan and Xiao Qing Jin. Circulant and skew-circulant preconditioners for skew-Hermitian type Toeplitz systems. BIT, 31(4):632–646, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=4&spage=632.

Enright:1991:PDC

[1929] W. H. Enright and D. J. Higham. Parallel defect control. BIT, 31(4): 647–663, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=31&issue=4&spage=647.

Hansbo:1991:EEN

[1930] Anita Hansbo. Error estimates for the numerical solution of a time-periodic linear parabolic problem. BIT, 31(4):664–685, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=664.

Matos:1991:SNA

[1931] Ana C. Matos. Some new acceleration methods for periodic-linearly convergent power series. BIT, 31(4):686–696, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=686.

Watson:1991:SEV

[1932] G. A. Watson and K. F. C.

Yiu. On the solution of the errors in variables problem using the l_{1} norm. BIT, 31(4):697–710, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=697.

Zha:1991:PPS

[1933] Hong Yuan Zha. The product-product singular value decomposition of matrix triplets. BIT, 31(4):711–726, December 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=31&issue=4&spage=711.

Anonymous:1991:A

[1934] Anonymous. Acknowledgements. BIT, 31(4):727–728, December 1991. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0006-3835&volume=31&issue=4&spage=
                                        727.

Anonymous:1991:I

[1935] Anonymous. Index 1991. BIT, 31 (4):729–732, 1991. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Nordstrom:1992:P

[1936] Bengt Nordström. Preface. BIT, 32(1): 2–4, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=1&spage=2.

Colson:1992:LPR

[1937] Loïc Colson. On list primitive recursion and the complexity of computing inf. BIT, 32(1):5–9, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=1&spage=5.

Coquand:1992:PTT

[1938] Thierry Coquand. The paradox of trees in type theory. BIT, 32(1):10– 14, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=1&spage=10.

Despeyroux:1992:TIP

[1939] Joëlle Despeyroux. Theo: an interactive proof development system. BIT, 32(1):15–29, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=32&
issue=1&spage=15.

Fredholm:1992:PID

[1940] Daniel Fredholm and Svetozar Serafimovski. Partial inductive definitions as type-systems for λ-terms. BIT, 32(1): 30–44, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=1&spage=30.

Hallnas:1992:SDI

[1941] Lars Hallnäs. On systems of definitions, induction and recursion. BIT, 32(1): 45–63, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=1&spage=45.

Mints:1992:CSI

[1942] G. Mints. Complexity of subclasses of the intuitionistic propositional calculus. BIT, 32(1):64– 69, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=1&spage=64.

Palmgren:1992:RML

[1943] Erik Palmgren and Viggo Stoltenberg-Hansen. Remarks on Martin-Löf’s partial type theory. BIT, 32(1):70– 83, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=1&spage=70.

Salvesen:1992:SST

[1944] Anne Salvesen. On specifications, subset types and interpretation of proposition in type theory. BIT, 32(1):84– 101, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=1&spage=84.

Bellen:1992:LEE

[1945] A. Bellen, Z. Jackiewicz, and M. Zennaro. Local error estimation for singlyimplicit formulas by two-step Runge– Kutta methods. BIT, 32(1):104– 117, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=1&spage=104.

Butcher:1992:GPA

[1946] J. C. Butcher and F. H. Chipman. Generalized Padé approximations to the exponential function. BIT, 32(1):118– 130, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=1&spage=118.

Calvo:1992:OCC

[1947] M. P. Calvo and J. M. Sanz-Serna. Order conditions for canonical Runge–Kutta–Nyström methods. BIT, 32(1):131–142, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=32&
issue=1&spage=131.

Elhay:1992:JMS

[1948] Sylvan Elhay, Gene H. Golub, and Jaroslav Kautský. Jacobi matrices for sums of weight functions. BIT, 32(1):143–166, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=32&
issue=1&spage=143.

Lin:1992:FCF

[1949] Qun Lin and J. G. Rokne. A family of centered forms for a polynomial. BIT, 32(1):167–176, March 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=1&spage=167.

Gilbert:1992:AT

[1950] John R. Gilbert. Algorithm theory. BIT, 32(2):177–179, 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:1992:P

[1951] Anonymous. Preface. BIT, 32(2):178– 179, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=2&spage=178.

Boppana:1992:AMI

[1952] Ravi Boppana and Magnús M. Halldórsson. Approximating maximum independent sets by excluding subgraphs. BIT, 32(2):180– 196, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=2&spage=180.

Arnborg:1992:CRP

[1953] Stefan Arnborg and Andrzej Proskurowski. Canonical representations of partial 2- and 3-trees. BIT, 32(2):197– 214, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=2&spage=197.

Pantziou:1992:EPA

[1954] Grammati E. Pantziou, Paul G. Spirakis, and Christos D. Zaroliagis. Efficient parallel algorithms for shortest paths in planar digraphs. BIT, 32(2): 215–236, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=215.

Eppstein:1992:FSS

[1955] David Eppstein. Finding the k smallest spanning trees. BIT, 32(2):237– 248, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=237.

Hershberger:1992:ASD

[1956] John Hershberger and Subhash Suri.

Applications of a semi-dynamic convex hull algorithm. BIT, 32(2):249– 267, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=249.

vanKreveld:1992:IQS

[1957] Marc van Kreveld, Mark Overmars, and Pankaj K. Agarwal. Intersection queries in sets of disks. BIT, 32(2): 268–279, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=268.

Levcopoulos:1992:FAG

[1958] Christos Levcopoulos and Andrzej Lingas. Fast algorithms for greedy triangulation. BIT, 32(2):280–296, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=280.

Baeza-Yates:1992:IBE

[1959] Ricardo Baeza-Yates, Gaston H. Gonnet, and Nivio Ziviani. Improved bounds for the expected behaviour of AVL trees. BIT, 32(2):297– 315, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=297.

Papadakis:1992:ASU

[1960] Thomas Papadakis, J. Ian Munro, and Patricio V. Poblete. Average search and update costs in skip lists. BIT, 32(2):316–332, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=316.

Alefeld:1992:SEM

[1961] Götz E. Alefeld and Florian A. Potra. Some efficient methods for enclosing simple zeros of nonlinear equations. BIT, 32(2):334– 344, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=334.

Saito:1992:BSA

[1962] Satoshi Saito, Hiroshi Sugiura, and Taketomo Mitsui. Butcher’s simplifying assumption for symplectic integrators. BIT, 32(2):345– 349, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=345.

Sand:1992:SBE

[1963] Jørgen Sand and Stig Skelboe. Stability of backward Euler multirate methods and convergence of waveform relaxation. BIT, 32(2):350– 366, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=350.

Stoffer:1992:EIC

[1964] D. Stoffer and K. Nipp. Erratum: “Invariant curves for variable step size integrators” [BIT 31(1), 1991, pp. 169– 180; MR 91m:65204]. BIT, 32(2):367– 368, June 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=2&spage=367. See

[1886].

Blieberger:1992:FSH

[1965] J. Blieberger and U. Schmid. FCFSscheduling in a hard real-time environment under rush-hour conditions. BIT, 32(3):370–383, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=370.

Hoshi:1992:PUQ

[1966] Mamoru Hoshi and Philippe Flajolet. Page usage in a quadtree index. BIT, 32(3):384–402, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=384.

Jan:1992:MVE

[1967] Rong Hong Jan, Lih-Hsing Hsu, and Yueh-Ying Lee. The most vital edges with respect to the number of spanning trees in two-terminal seriesparallel graphs. BIT, 32(3):403–412, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=403.

Levene:1992:INJ

[1968] Mark Levene and George Loizou. Inferring null join dependencies in relational databases. BIT, 32(3):413–429, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=413.

Mairson:1992:ETE

[1969] Harry G. Mairson. The effect of table expansion on the program complexity of perfect hash functions. BIT, 32(3):430–440, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=430.

DAzevedo:1992:TCE

[1970] E. F. D’Azevedo, P. A. Forsyth, and Wei Pai Tang. Towards a cost-effective ILU preconditioner with high level fill. BIT, 32(3):442–463, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=442.

Bar-On:1992:NAP

[1971] Ilan Bar-On and Avram Sidi. New algorithms for polynomial and trigonometric interpolation on parallel computers. BIT, 32(3):464–480, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=464.

Chan:1992:SQC

[1972] Tony F. Chan, Julia A. Olkin, and Donald W. Cooley. Solving quadratically constrained least squares using black box solvers. BIT, 32(3):481–494, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=481.

Rogina:1992:BSA

[1973] Mladen Rogina. Basis of splines associated with some singular differential operators. BIT, 32(3):496–505, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=496.

Wolke:1992:IRL

[1974] Ralf Wolke. Iteratively reweighted least squares: a comparison of several single step algorithms for linear models. BIT, 32(3):506–524, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=506.

Kjurkchiev:1992:NEO

[1975] N. Kjurkchiev. Note on the estimation of the order of convergence of some iterative methods. BIT, 32(3):525–528, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=3&spage=525.

Guthmann:1992:EPT

[1976] Andreas Guthmann. Effective primality tests for integers of the forms N = k · 3^{n} + 1 and N = k · 2^{m}3^{n} + 1. BIT, 32(3):529–534, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=529.

Lubich:1992:CAP

[1977] Ch. Lubich. Chebyshev acceleration of Picard–Lindelöf iteration. BIT, 32(3):535–538, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=535.

Saito:1992:FSI

[1978] Satoshi Saito, Hiroshi Sugiura, and Taketomo Mitsui. Family of symplectic implicit Runge–Kutta formulae. BIT, 32(3):539–543, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=539.

Anonymous:1992:NBJa

[1979] Anonymous. New books and journals. BIT, 32(3):544, September 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=3&spage=544.

Chen:1992:AAS

[1980] Tung Shou Chen, Wei Pang Yang, and R. C. T. Lee. Amortized analysis of some disk scheduling algorithms: SSTF, SCAN, and N-StepSCAN. BIT, 32(4):546–558, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=546.

Coolsaet:1992:OAT

[1981] K. Coolsaet, H. De Meyer, and V. Fack. Optimal algorithms for total exchange without buffering on the hypercube. BIT, 32(4):559–569, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=559.

deGraaf:1992:EHH

[1982] Jeannette M. de Graaf and Walter A. Kosters. Expected heights in heaps. BIT, 32(4):570–579, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=570.

Katajainen:1992:SMS

[1983] Jyrki Katajainen and Tomi Pasanen. Stable minimum space partitioning in linear time. BIT, 32(4):580–585, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=580.

Olariu:1992:CTC

[1984] Stephan Olariu, James L. Schwing, and Jingyuan Zhang. A constanttime channel-assignment algorithm on reconfigurable meshes. BIT, 32(4): 586–597, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=586.

Perkins:1992:MDU

[1985] A. Louise Perkins. A mixed directedundirected data structure for a parallel implementation of a domain decomposition algorithm. BIT, 32(4):598–608, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=598.

Yu:1992:WMI

[1986] Chang Wu Yu and Gen-Huey Chen.

The weighted maximum independent set problem in permutation graphs. BIT, 32(4):609–618, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=609.

Hernandez:1992:SRK

[1987] Diego Bricio Hernández and Renato Spigler. A-stability of Runge– Kutta methods for systems with additive noise. BIT, 32(4):620–633, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=620.

intHout:1992:NIP

[1988] K. J. in ’t Hout. A new interpolation procedure for adapting Runge– Kutta methods to delay differential equations. BIT, 32(4):634–649, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=634.

Jin:1992:CPS

[1989] Xiao-Qing Jin and Raymond H. Chan. Circulant preconditioners for second order hyperbolic equations. BIT, 32(4):650–664, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=650.

Lyness:1992:AFO

[1990] J. N. Lyness and T. Sørevik. An algorithm for finding optimal integration lattices of composite order. BIT, 32(4):665–675, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=665.

Muller:1992:IEP

[1991] Michael Müller. An inverse eigenvalue problem: computing B-stable Runge–Kutta methods having real poles. BIT, 32(4):676–688, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=676.

Skelboe:1992:MPI

[1992] Stig Skelboe. Methods for parallel integration of stiff systems of ODEs. BIT, 32(4):689–701, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=689.

Sun:1992:CPB

[1993] Ji Guang Sun. Componentwise perturbation bounds for some matrix decompositions. BIT, 32(4):702–714, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=32&issue=4&spage=702.

Anonymous:1992:A

[1994] Anonymous. Acknowledgements. BIT, 32(4):715–716, December 1992. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0006-3835&volume=32&issue=4&spage=
                                        715.

Anonymous:1992:NBJb

[1995] Anonymous. New books and journals. BIT, 32(4):716, December 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=32&issue=4&spage=716.

Atkinson:1993:PPP

[1996] M. D. Atkinson and Murali Thiyagarajah. The permutational power of a priority queue. BIT, 33(1):1–6, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=1&spage=1.

Bhagavathi:1993:SRM

[1997] D. Bhagavathi, P. J. Looges, S. Olariu, J. L. Schwing, and J. Zhang. Selection on rectangular meshes with multiple broadcasting. BIT, 33(1):7–14, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=1&spage=7.

Cesarini:1993:SAH

[1998] F. Cesarini and G. Soda. Single access hashing with overflow separators for dynamic files. BIT, 33(1):15–28, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=15.

Gambosi:1993:EID

[1999] Giorgio Gambosi, Marco Protasi, and Maurizio Talamo. An efficient implicit data structure for relation testing and searching in partially ordered sets. BIT, 33(1):29–45, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=29.

Liaw:1993:OAA

[2000] Horng Twu Liaw and Chin Laung Lei.

An optimal algorithm to assign cryptographic keys in a tree structure for access control. BIT, 33(1):46–56, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=46.

Stewart:1993:PGH

[2001] Iain A. Stewart. On parallelizing a greedy heuristic for finding small dominant sets. BIT, 33(1):57–62, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=57.

Carstensen:1993:QLC

[2002] Carsten Carstensen. On quadratic-like convergence of the means for two methods for simultaneous rootfinding of polynomials. BIT, 33(1):63–73, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=63.

Crouzeix:1993:SRA

[2003] M. Crouzeix, S. Larsson, S. Piskarëv, and V. Thomée. The stability of rational approximations of analytic semigroups. BIT, 33(1):74–84, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=74.

Dahlquist:1993:MEF

[2004] Germund Dahlquist. A “multigrid” extension of the FFT for the numerical inversion of Fourier and Laplace transforms. BIT, 33(1):85–112, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=85.

Delvos:1993:HIT

[2005] Franz-Jürgen Delvos. Hermite interpolation with trigonometric polynomials. BIT, 33(1):113–123, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=113.

Higham:1993:PTB

[2006] Nicholas J. Higham. Perturbation theory and backward error analysis for AX − XB = C. BIT, 33(1):124–136, March 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=124.

Jay:1993:CCR

[2007] Laurent Jay. Convergence of a class of Runge–Kutta methods for differential-algebraic systems of index 2. BIT, 33(1):137–150, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=137.

Lewanowicz:1993:EBN

[2008] Stanisław Lewanowicz. Error bounds for a near-minimax approximation. BIT, 33(1):151–157, March 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=151.

Park:1993:RAH

[2009] Haesun Park and Vjeran Hari. A real algorithm for the Hermitian eigenvalue decomposition. BIT, 33 (1):158–171, March 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=1&spage=158.

Skeel:1993:VSS

[2010] Robert D. Skeel. Variable step size destabilizes the Störmer/Leapfrog/ Verlet method. BIT, 33(1):172–175, March 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.mai.liu.se/BIT/contents/bit33.
html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0006-3835&volume=33&issue=1&spage=
172.

Imielinska:1993:GHA

[2011] Celina Imielińska and Bahman Kalantari. A generalized hypergreedy algorithm for weighted perfect matching. BIT, 33(2):177–189, June 1993.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=2&spage=177.

Lundberg:1993:PBM

[2012] Lars Lundberg. Performance bounds on multiprocessor scheduling strategies for statically allocated programs. BIT, 33(2):190–213, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=2&spage=190.

Olsson:1993:ELP

[2013] Roland Olsson. Execution of logic programs by iterative-deepening A^{∗} SLDtree search. BIT, 33(2):214–231, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=2&spage=214.

Sykora:1993:CNH

[2014] Ondrej Sýkora and Imrich Vřto.

On crossing numbers of hypercubes and cube connected cycles. BIT, 33(2):232–237, June 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=232.

Wang:1993:EUC

[2015] Cao An Wang. Efficiently updating constrained Delaunay triangulations. BIT, 33(2):238–252, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=238.

Barrera:1993:IBS

[2016] Tony Barrera and Pelle Olsson. An integer based square root algorithm. BIT, 33(2):253–261, June 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=253.

Dax:1993:RRM

[2017] Achiya Dax. A row relaxation method for large minimax problems. BIT, 33(2):262–276, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=262.

Gautschi:1993:RRO

[2018] Walter Gautschi. Is the recurrence relation for orthogonal polynomials always stable? BIT, 33(2):277–284, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=2&spage=277.

Higham:1993:SO

[2019] Desmond J. Higham and Lloyd N.

Trefethen. Stiffness of ODEs. BIT, 33(2):285–303, June 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=2&spage=285.

Li:1993:PBG

[2020] Ren Cang Li. A perturbation bound for the generalized polar decomposition. BIT, 33(2):304–308, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=2&spage=304.

Moore:1993:HOA

[2021] Peter K. Moore and Joseph E. Flaherty.

High-order adaptive finite elementsingly implicit Runge–Kutta methods for parabolic differential equations. BIT, 33(2):309–331, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=309.

Schneider:1993:NEM

[2022] Stefan Schneider. Numerical experiments with a multistep Radau method. BIT, 33(2):332–350, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=332.

Anonymous:1993:NBJa

[2023] Anonymous. New books and journals. BIT, 33(2):351, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=2&spage=351.

Chou:1993:MSI

[2024] Hong Chich Chou and Chung Ping Chung. Modeling of superscalar instruction scheduling and analysis of a heuristic scheduling algorithm. BIT, 33(3):354–371, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=354.

Dutton:1993:WHS

[2025] Ronald D. Dutton. Weak-heap sort. BIT, 33(3):372–381, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=3&spage=372.

Jan:1993:KLP

[2026] J. K. Jan and C. C. Chang. A key-lock-pair oriented access control scheme for the growth of users and files. BIT, 33(3):382–389, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=3&spage=382.

Liang:1993:AFC

[2027] Chiou Kuo Liang. An O(n^{2}) algorithm for finding the compact sets of a graph. BIT, 33(3):390–395, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=3&spage=390.

Olariu:1993:OCH

[2028] S. Olariu, J. L. Schwing, and J. Zhang. Optimal convex hull algorithms on enhanced meshes. BIT, 33(3):396–410, September 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=396.

Poblete:1993:NHE

[2029] Patricio V. Poblete. A note on the height of the kth element in a heap. BIT, 33(3):411–412, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=411.

Yu:1993:PAP

[2030] Chang Wu Yu and Gen Huey Chen. Parallel algorithms for permutation graphs. BIT, 33(3):413–419, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=413.

Yu:1993:LTA

[2031] Ming Shing Yu and Cheng Hsing Yang. A linear time algorithm for the maximum matching problem on cographs. BIT, 33(3):420–432, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=420.

Amodio:1993:SSB

[2032] P. Amodio, F. Mazzia, and D. Trigiante. Stability of some boundary value methods for the solution of initial value problems. BIT, 33 (3):434–451, September 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=3&spage=434.

Butcher:1993:DIG

[2033] J. C. Butcher and Z. Jackiewicz. Diagonally implicit general linear methods for ordinary differential equations. BIT, 33(3):452–472, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=3&spage=452.

Calvetti:1993:FIV

[2034] D. Calvetti and L. Reichel. Fast inversion of Vandermonde-like matrices involving orthogonal polynomials. BIT, 33(3):473–484, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=3&spage=473.

Miekkala:1993:GPS

[2035] Ulla Miekkala. Graph properties for splitting with grounded Laplacian matrices. BIT, 33(3):485–495, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=485.

Schmidt:1993:CMP

[2036] Jochen W. Schmidt and Walter Heß. S-convex, monotone, and positive interpolation with rational bicubic splines of C^{2}-continuity. BIT, 33 (3):496–511, September 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=496.

Schwetlick:1993:SSU

[2037] H. Schwetlick and V. Kunert. Spline smoothing under constraints on derivatives. BIT, 33(3):512–528, September 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=3&spage=512.

Chang:1993:RCO

[2038] Chin-Chen Chang, Huey-Cheue Kowng, and Tzong-Chen Wu. A refinement of a compression-oriented addressing scheme. BIT, 33(4):529–535, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=529.

Chen:1993:GPA

[2039] Jong-Jen Chen, Chiau-Shin Wang, and Ching-Roung Chou. A general performance analysis method for uniform memory architectures. BIT, 33 (4):536–560, December 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=536.

Harris:1993:ODM

[2040] Evan P. Harris and Kotagiri Ramamohanarao. Optimal dynamic multiattribute hashing for range queries. BIT, 33(4):561–579, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=561.

Jansen:1993:CDC

[2041] Klaus Jansen and Gerhard J. Woeginger. The complexity of detecting crossing-free configurations in the plane. BIT, 33(4):580–595, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=580.

Lysne:1993:EPP

[2042] Olav Lysne. The equational part of proofs by structural induction. BIT, 33(4):596–618, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=596.

Al-Rabeh:1993:OOD

[2043] A. H. Al-Rabeh. Optimal order diagonally implicit Runge–Kutta methods. BIT, 33(4):619–633, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=619.

Bialecki:1993:FAS

[2044] Bernard Bialecki and D. Scott Dillery. Fourier analysis of Schwarz alternating methods for piecewise Hermite bicubic orthogonal spline collocation. BIT, 33(4):634–646, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=634.

DiLena:1993:RRP

[2045] G. Di Lena, R. I. Peluso, and G. Piazza. Results on the relative perturbation of the singular values of a matrix. BIT, 33(4):647–653, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=647.

Hernandez:1993:CSI

[2046] Diego Bricio Hernández and Renato Spigler. Convergence and stability of implicit Runge–Kutta methods for systems with multiplicative noise. BIT, 33(4):654–669, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=654.

Schneider:1993:CRG

[2047] Stefan Schneider. Convergence results for general linear methods on singular perturbation problems. BIT, 33 (4):670–686, December 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit33.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=33&issue=4&spage=670.

Soderkvist:1993:PAO

[2048] Inge Söderkvist. Perturbation analysis of the orthogonal Procrustes problem. BIT, 33(4):687–694, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit33.html;

                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=687.

Xu:1993:SAJ

[2049] Shu Fang Xu. A stability analysis of the Jacobi matrix inverse eigenvalue problem. BIT, 33(4):695–702, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit33.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=695.

Anonymous:1993:A

[2050] Anonymous. Acknowledgements. BIT, 33(4):703, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=703.

Anonymous:1993:NBJb

[2051] Anonymous. New books and journals. BIT, 33(4):704, December 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=33&issue=4&spage=704.

Bjorck:1994:SAL

[2052] Åke Björck and C. C. Paige. Solution of augmented linear systems using orthogonal factorizations. BIT, 34(1):1–24, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=1&spage=1.

Butcher:1994:TAD

[2053] J. C. Butcher. A transformation for the analysis of DIMSIMs. BIT, 34(1):25–32, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=1&spage=25.

Christara:1994:QSC

[2054] Christina C. Christara. Quadratic spline collocation methods for elliptic partial differential equations. BIT, 34(1):33–61, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=1&spage=33.

Espelid:1994:IVS

[2055] Terje O. Espelid. On integrating vertex singularities using extrapolation. BIT, 34(1):62–79, March 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=1&spage=62.

Hairer:1994:NES

[2056] E. Hairer, A. Murua, and J. M. Sanz-Serna. The non-existence of symplectic multi-derivative Runge–Kutta methods. BIT, 34(1):80–87, March 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=1&spage=80.

Higham:1994:RKS

[2057] Desmond J. Higham. Runge–Kutta stability on a Floquet problem. BIT, 34(1):88–98, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=1&spage=88.

Huckle:1994:FTT

[2058] Thomas Huckle. Fast transforms for tridiagonal linear equations. BIT, 34(1):99–112, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=1&spage=99.

Kraaijevanger:1994:TCR

[2059] J. F. B. M. Kraaijevanger. Two counterexamples related to the Kreiss matrix theorem. BIT, 34(1):113– 119, March 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.mai.liu.se/BIT/contents/bit34.
html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0006-3835&volume=34&issue=1&spage=
113.

Mastroianni:1994:SNM

[2060] Giuseppe Mastroianni and Siegfried Prössdorf. Some nodes matrices appearing in the numerical analysis for singular integral equations. BIT, 34(1):120–128, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=1&spage=120.

Mazya:1994:SAT

[2061] Vladimir Maz’ya and Vladimir Karlin. Semi-analytic time-marching algorithms for semi-linear parabolic equations. BIT, 34(1):129–147, March 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=1&spage=129.

Naur:1994:PVF

[2062] Peter Naur. Proof versus formalization. BIT, 34(1):148–164, March 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=1&spage=148.

Ruhe:1994:RKA

[2063] Axel Ruhe. The rational Krylov algorithm for nonsymmetric eigenvalue problems. III: Complex shifts for real matrices. BIT, 34(1):165–176, March 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=1&spage=165.

Ruhe:1994:CEF

[2064] Axel Ruhe. Carl-Erik Fröberg on his 75th birthday June 23, 1993. BIT, 34(1):i–ii, March 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=1&spage=??.

Bruaset:1994:NSO

[2065] A. M. Bruaset and A. Tveito. A numerical study of optimized sparse preconditioners. BIT, 34(2):177–204, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=177.

Chakravarti:1994:FAM

[2066] P. C. Chakravarti and M. Barrientos. On a fast and accurate method for computing Fourier transforms. BIT, 34(2):205–214, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=2&spage=205.

Crisci:1994:PFM

[2067] M. R. Crisci. Parallel frontal methods for ODE’s. BIT, 34(2):215–227, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=2&spage=215.

vanDorsselaer:1994:SEB

[2068] J. L. M. van Dorsselaer and W. Hundsdorfer. Stability estimates based on numerical ranges with an application to a spectral method. BIT, 34(2):228–238, June 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=2&spage=228.

Gulliksson:1994:IRC

[2069] M. Gulliksson. Iterative refinement for constrained and weighted linear least squares. BIT, 34(2):239–253, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;

                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=239.

Han:1994:AEE

[2070] G. Q. Han. Asymptotic error expansion for the Nyström method for nonlinear Fredholm integral equations of the second kind. BIT, 34(2):254–261, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=254.

Koto:1994:SPS

[2071] T. Koto. A stability property of A-stable natural Runge–Kutta methods for systems of delay differential equations. BIT, 34(2):262–267, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=262.

Liu:1994:RBT

[2072] A. Liu and B. Joe. Relationship between tetrahedron shape measures. BIT, 34(2):268–287, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=268.

Novak:1994:OSQ

[2073] E. Novak and K. Petras. Optimal stochastic quadrature formulas for convex functions. BIT, 34(2):288–294, June 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=2&spage=288.

Schneid:1994:CCS

[2074] J. Schneid. A characterization of compactly supported orthonormal wavelets. BIT, 34(2):295–303, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=2&spage=295.

Walker:1994:SMA

[2075] W. J. Walker. A summability method for the arithmetic Fourier transform. BIT, 34(2):304–309, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=2&spage=304.

Hairer:1994:SRK

[2076] E. Hairer and G. Wanner. Symplectic Runge–Kutta methods with real eigenvalues. BIT, 34(2):310–312, June 1994. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=310.

Jin:1994:NBC

[2077] X.-Q. Jin. A note on best conditioned preconditioners. BIT, 34 (2):313–317, June 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=313.

Kjurkchiev:1994:SRD

[2078] N. Kjurkchiev and K. Mahdi. Some remarks on Dvorcuk’s root-finding method. BIT, 34(2):318–322, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=2&spage=318.

Anonymous:1994:B

[2079] Anonymous. Books. BIT, 34(2):323, 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html.

Anonymous:1994:NBJ

[2080] Anonymous. New books and journals. BIT, 34(2):323, June 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=34&
issue=2&spage=323.

Bjorhus:1994:DID

[2081] M. Bjørhus. On dynamic iteration for delay differential equations. BIT, 34(3):325–336, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=3&spage=325.

Bochev:1994:QIS

[2082] P. B. Bochev and C. Scovel. On quadratic invariants and symplectic structure. BIT, 34(3):337–345, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=3&spage=337.

DellaVecchia:1994:TNF

[2083] B. Della Vecchia. Two new formulas for the numerical evaluation of the Hilbert Transform. BIT, 34 (3):346–360, September 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=3&spage=346.

Higueras:1994:RKS

[2084] Inmaculada Higueras. Runge–Kutta schemes that maintain invariant solutions for differential algebraic systems. BIT, 34(3):361–366, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=3&spage=361.

Jin:1994:HPT

[2085] X.-Q. Jin. Hartley preconditioners for Toeplitz systems generated by positive continuous functions. BIT, 34 (3):367–371, September 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=3&spage=367.

Jonasson:1994:CSL

[2086] Kristján Jónasson and Kaj Madsen. Corrected sequential linear programming for sparse minimax optimization. BIT, 34(3):372–387, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=3&spage=372.

Kocak:1994:BSG

[2087] Zeynep F. Koçak and George M. Phillips. B-splines with geometric knot spacings. BIT, 34(3):388–399, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=3&spage=388.

Kuang:1994:ASO

[2088] J. X. Kuang, J. X. Xiang, and H. J. Tian. The asymptotic stability of one-parameter methods for neutral differential equations. BIT, 34 (3):400–408, September 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=3&spage=400.

Meerbergen:1994:SIC

[2089] K. Meerbergen, A. Spence, and D. Roose. Shift-invert and Cayley transforms for detection of rightmost eigenvalues of nonsymmetric matrices. BIT, 34(3):409–423, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=3&spage=409.

Soderkvist:1994:CNA

[2090] Inge Söderkvist and Per-Åke Wedin.

On condition numbers and algorithms for determining a rigid body movement. BIT, 34(3):424–436, September 1994. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=3&spage=424.

Wei:1994:SUC

[2091] Musheng Wei and Birong Zhang. Structures and uniqueness conditions of MK-weighted pseudoinverses. BIT, 34(3):437–450, September 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=3&spage=437.

Rodseth:1994:NPT

[2092] Ö. J. Rödseth. A note on primality tests for N = h · 2^{n} −

1. BIT, 34(3):451–454, September 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=34&issue=3&spage=451.

Alishenas:1994:MVS

[2093] Taifun Alishenas and Örn Ólafsson. Modeling and velocity stabilization of constrained mechanical systems. BIT, 34(4):455–483, December 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit34.html;
                                         http://www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=455.

Aspvall:1994:FMH

[2094] Bengt Aspvall and Pinar Heggernes.

Finding minimum height elimination trees for interval graphs in polynomial time. BIT, 34(4):484–509, December 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=484.

Bjorck:1994:ISB

[2095] Åke Björck, Eric Grimme, and Paul Van Dooren. An implicit shift bidiagonalization algorithm for ill-posed systems. BIT, 34(4):510–534, December 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=510.

Cuyt:1994:CMH

[2096] Annie Cuyt. On the convergence of the multivariate “homogeneous” qdalgorithm. BIT, 34(4):535–545, December 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=535.

Favati:1994:TPE

[2097] Paola Favati, Grazia Lotti, and Francesco Romani. Theoretical and practical efficiency measures for symmetric interpolatory quadrature formulas. BIT, 34(4):546–557, December 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=546.

Gander:1994:LSF

[2098] Walter Gander, Gene H. Golub, and Rolf Strebel. Least-squares fitting of circles and ellipses. BIT, 34 (4):558–578, December 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=558.

Serra:1994:PSA

[2099] Stefano Serra. Preconditioning strategies for asymptotically ill-conditioned block Toeplitz systems. BIT, 34 (4):579–594, December 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit34.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=579.

Anonymous:1994:A

[2100] Anonymous. Acknowledgements. BIT, 34(4):595, December 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.

liu.se/BIT/contents/bit34.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=34&issue=4&spage=595.

Anonymous:1994:I

[2101] Anonymous.  Index.  BIT, 34

(4):596–598, December 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit34.html.

Barrlund:1995:CSP

[2102] Anders Barrlund. Comparing stability properties of three methods in DAEs or ODEs with invariants. BIT Numerical Mathematics, 35(1):1–18, March 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=1.

Cahlon:1995:SVI

[2103] Baruch Cahlon. On the stability of Volterra integral equations with a lagging argument. BIT Numerical Mathematics, 35(1):19–29, March 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=19.

Clift:1995:WGB

[2104] Simon S. Clift and Wei-Pai Tang.

Weighted graph based ordering techniques for preconditioned conjugate gradient methods. BIT Numerical Mathematics, 35(1):30–47, March 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=30.

Evans:1995:AIM

[2105] D. J. Evans and Guang-Yao Lei. Approximate inverses of multidiagonal matrices and application to the block PCG method. BIT Numerical Mathematics, 35(1):48–63, March 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=48.

Hunter:1995:SEE

[2106] D. B. Hunter. Some error expansions for Gaussian quadrature. BIT Numerical Mathematics, 35(1):64–82, March 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=64.

Jeltsch:1995:MAS

[2107] Rolf Jeltsch, Rosemary A. Renaut, and Kosie J. H. Smit. The maximal accuracy of stable difference schemes for the wave equation. BIT Numerical Mathematics, 35(1):83–115, March 1995. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=83.

Rhee:1995:CCP

[2108] Noah H. Rhee and Vjeran Hari.

On the cubic convergence of the Paardekooper method. BIT Numerical Mathematics, 35(1):116–132, March 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=116.

Zhu:1995:PEE

[2109] Chang Zhong Zhu and Charles B.

Dunham. Parameter error estimate of near alternation approximation. BIT Numerical Mathematics, 35(1):133–142, March 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=1&spage=133.

Arevalo:1995:CMD

[2110] Carmen Arévalo and Gustaf Söderlind.

Convergence of multistep discretizations of DAEs. BIT Numerical Mathematics, 35(2):143–168, June 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;

                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=2&spage=143.

Brezinski:1995:LAB

[2111] C. Brezinski and M. Redivo-Zaglia. Look-ahead in BI-CGSTAB and other product methods for linear systems. BIT Numerical Mathematics, 35(2):169–201, June 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=2&spage=169.

Butcher:1995:FOR

[2112] J. C. Butcher. On fifth order Runge–Kutta methods. BIT Numerical Mathematics, 35(2):202–209, June 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=2&spage=202.

Farestam:1995:FAF

[2113] S. Farestam and R. B. Simpson. A framework for advancing front techniques of finite element mesh generation. BIT Numerical Mathematics, 35(2):210–232, June 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=2&spage=210.

Goodman:1995:CFG

[2114] T. N. T. Goodman, C. A. Micchelli, G. Rodriguez, and S. Seatzu. On the Cholesky factorization of the Gram matrix of locally supported functions. BIT Numerical Mathematics, 35(2):233–257, June 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=2&spage=233.

McLachlan:1995:CMP

[2115] Robert I. McLachlan. Composition methods in the presence of small parameters. BIT Numerical Mathematics, 35(2):258–268, June 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=2&spage=258.

Portillo:1995:LDS

[2116] A. Portillo and J. M. Sanz-Serna.

Lack of dissipativity is not symplecticness. BIT Numerical Mathematics, 35(2):269–276, June 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=2&spage=269.

Rabinowitz:1995:UCC

[2117] Philip Rabinowitz and Elisabetta Santi. On the uniform convergence of Cauchy principal values of quasiinterpolating splines. BIT Numerical Mathematics, 35(2):277–290, June 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=2&spage=277.

Bjorhus:1995:NCD

[2118] Morten Bjørhus. A note on the convergence of discretized dynamic iteration. BIT Numerical Mathematics, 35(2):291–296, June 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=2&spage=291.

Arevalo:1995:IAB

[2119] Carmen Arévalo and Per Lötstedt. Improving the accuracy of BDF methods for index 3 differential-algebraic equations. BIT Numerical Mathematics, 35(3):297–308, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=3&spage=297.

Drkosova:1995:NSG

[2120] J. Drkošová, A. Greenbaum, M. Rozložník, and Z. Strakoš. Numerical stability of GMRES. BIT Numerical Mathematics, 35(3):309–330, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=309.

Lee:1995:TLO

[2121] David Lee and Henryk Woźniakowski.

Testing linear operators. BIT Numerical Mathematics, 35(3):331– 351, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit35.html;   http://www.
springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=35&
issue=3&spage=331.

Luther:1995:HAT

[2122] Wolfram Luther. Highly accurate tables for elementary functions. BIT Numerical Mathematics, 35(3): 352–360, September 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=352.

Schwetlick:1995:LSA

[2123] Hubert Schwetlick and Torsten Schütze.

Least squares approximation by splines with free knots. BIT Numerical Mathematics, 35(3):361–384, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=361.

Sun:1995:NBP

[2124] Ji-Guang Sun. A note on backward perturbations for the Hermitian eigenvalue problem. BIT Numerical Mathematics, 35(3):385–393, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=385.

Tuomela:1995:NHO

[2125] Jukka Tuomela. A note on high order schemes for the one-dimensional wave equation. BIT Numerical Mathematics, 35(3):394–405, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=394.

Verner:1995:OEC

[2126] J. H. Verner and M. Zennaro. The orders of embedded continuous explicit Runge–Kutta methods. BIT Numerical Mathematics, 35(3):406–416, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=406.

Xue:1995:ERP

[2127] Jungong Xue and Erxiong Jiang. Entrywise relative perturbation theory for nonsingular M-matrices and applications. BIT Numerical Mathematics, 35(3):417–427, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=417.

Yalamov:1995:SCR

[2128] Plamen Y. Yalamov. On the stability of the cyclic reduction without back substitution for tridiagonal systems. BIT Numerical Mathematics, 35(3):428–447, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=3&spage=428.

Zha:1995:NCS

[2129] Hongyuan Zha and Zhenyue Zhang.

A note on constructing a symmetric matrix with specified diagonal entries and eigenvalues. BIT Numerical Mathematics, 35(3):448–452, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=3&spage=448. See

remark [2316].

Gulliksson:1995:MGS

[2130] Mårten Gulliksson. On the modified Gram–Schmidt algorithm for weighted and constrained linear least squares problems. BIT Numerical Mathematics, 35(4):453–468, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=4&spage=453.

Hadjidimos:1995:ESC

[2131] A. Hadjidimos, D. Noutsos, and M. Tzoumas. Exact SOR convergence regions for a general class of p-cyclic matrices. BIT Numerical Mathematics, 35(4):469–487, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit35.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=35&issue=4&spage=469.

Hao:1995:SAF

[2132] Dinh Nho Hào, H.-J. Reinhardt, and A. Schneider. Stable approximation of fractional derivatives of rough functions. BIT Numerical Mathematics, 35(4):488–503, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=488.

Hu:1995:SAN

[2133] Guang-Da Hu and Taketomo Mitsui. Stability analysis of numerical methods for systems of neutral delaydifferential equations. BIT Numerical Mathematics, 35(4):504–515, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=504.

Jia:1995:BIO

[2134] Zhongxiao Jia. A block incomplete orthogonalization method for large nonsymmetric eigenproblems. BIT Numerical Mathematics, 35(4): 516–539, December 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=516.

Jiang:1995:NKB

[2135] Tianzi Jiang. On a new kind of Birkhoff type trigonometric interpolation. BIT Numerical Mathematics, 35(4):540–547, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=540.

Karlin:1995:TMA

[2136] Vladimir Karlin and Vladimir Maz′ya. Time-marching algorithms for initialboundary value problems based upon “approximate approximations”. BIT Numerical Mathematics, 35(4): 548–560, December 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=548.

Lin:1995:IIM

[2137] Qun Lin and J. G. Rokne. An interval iteration for multiple roots of transcendental equations. BIT Numerical Mathematics, 35(4):561–571, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=561.

Munthe-Kaas:1995:LBT

[2138] Hans Munthe-Kaas. Lie–Butcher theory for Runge–Kutta methods. BIT Numerical Mathematics, 35(4): 572–587, December 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=572.

Saunders:1995:SSR

[2139] Michael A. Saunders. Solution of sparse rectangular systems using LSQR and Craig. BIT Numerical Mathematics, 35(4):588–604, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=588.

Bakaev:1995:BAH

[2140] N. Yu. Bakaev. On the bounds of approximations of holomorphic semigroups. BIT Numerical Mathematics, 35(4):605–608, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=605.

Bohman:1995:HIC

[2141] J. Bohman and C.-E. Fröberg.

Heuristic investigation of chaotic mapping producing fractal objects. BIT Numerical Mathematics, 35(4): 609–615, December 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=609.

Anonymous:1995:A

[2142] Anonymous. Acknowledgements.

BIT Numerical Mathematics, 35 (4):616, December 1995. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit35.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=35&issue=4&spage=616.

Anonymous:1995:I

[2143] Anonymous. Index. BIT Numerical Mathematics, 35(4):617–619, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit35.html.

Arevalo:1996:SMM

[2144] C. Arévalo, C. Führer, and G. Söderlind. Stabilized multistep methods for index 2 Euler–Lagrange DAEs. BIT Numerical Mathematics, 36(1):1–13, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=1.

Barlow:1996:AST

[2145] Jesse L. Barlow, Peter A. Yoon, and Hongyuan Zha. An algorithm and a stability theory for downdating the ULV decomposition. BIT Numerical Mathematics, 36(1):14–40, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=14.

Bennett:1996:GIA

[2146] J. M. Bennett and J. J. Edwards.

A graph isomorphism algorithm using pseudoinverses. BIT Numerical Mathematics, 36(1):41–53, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=41.

Burrage:1996:PWR

[2147] K. Burrage, Z. Jackiewicz, S. P.

Nørsett, and R. A. Renaut. Preconditioning waveform relaxation iterations for differential systems. BIT Numerical Mathematics, 36(1):54–76, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=54.

Cooper:1996:SLS

[2148] G. J. Cooper. Some linear stability results for iterative schemes for implicit Runge–Kutta methods. BIT Numerical Mathematics, 36(1):77–85, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=77.

Gustafsson:1996:IFP

[2149] Ivar Gustafsson. An incomplete factorization preconditioning method based on modification of element matrices. BIT Numerical Mathematics, 36(1):86–100, March 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=86.

Jin:1996:PCW

[2150] Xiao-Qing Jin. A preconditioner for constrained and weighted least squares problems with Toeplitz structure. BIT Numerical Mathematics, 36(1):101–109, March 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=101.

Lin:1996:HOQ

[2151] Fu-Rong Lin and Michael K. Ng. Higher-order quadratures for circulant preconditioned Wiener–Hopf equations. BIT Numerical Mathematics, 36(1):110–121, March 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=110.

Reich:1996:EEC

[2152] Sebastian Reich. Enhancing energy conserving methods. BIT Numerical Mathematics, 36(1):122–134, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=122.

Serra:1996:ESP

[2153] Stefano Serra. On the extreme spectral properties of Toeplitz matrices generated by L^{1} functions with several minima/maxima. BIT Numerical Mathematics, 36(1):135–142, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=135.

Sidi:1996:FRC

[2154] Avram Sidi. Further results on convergence and stability of a generalization of the Richardson extrapolation process. BIT Numerical Mathematics, 36(1):143–157, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=143.

Tuomela:1996:CAO

[2155] Jukka Tuomela. On the construction of arbitrary order schemes for the many-dimensional wave equation. BIT Numerical Mathematics, 36(1):158–165, March 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=158.

Yoo:1996:ADM

[2156] K. Yoo and H. Park. Accurate downdating of a modified Gram–Schmidt QR decomposition. BIT Numerical Mathematics, 36(1):166–181, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=166.

Kjurkchiev:1996:NVF

[2157] N. Kjurkchiev. A note on the Le Verrier–Fadeev’s method. BIT Numerical Mathematics, 36(1):182–186, March 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.mai.liu.se/BIT/contents/bit36.
html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0006-3835&volume=36&issue=1&spage=
182.

Anonymous:1996:NBJ

[2158] Anonymous. New books and journals. BIT Numerical Mathematics, 36(1):187, March 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=1&spage=187.

Arioli:1996:REA

[2159] M. Arioli and C. Fassino. Roundoff error analysis of algorithms based on Krylov subspace methods. BIT Numerical Mathematics, 36(2):189–205, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=189.

Bar-On:1996:CRN

[2160] Ilan Bar-On, Bruno Codenotti, and Mauro Leoncini. Checking robust nonsingularity of tridiagonal matrices in linear time. BIT Numerical Mathematics, 36(2):206–220, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=206.

Brunner:1996:ICM

[2161] Hermann Brunner, Lin Qun, and Yan Ningning. The iterative correction method for Volterra integral equations. BIT Numerical Mathematics, 36(2):221–228, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=221.

Chan:1996:CLR

[2162] R. P. K. Chan and P. Chartier. A composition law for Runge–Kutta methods applied to index-2 differentialalgebraic equations. BIT Numerical Mathematics, 36(2):229–246, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=229.

Elden:1996:PEA

[2163] L. Eldén and H. Park. Perturbation and error analyses for block downdating of a Cholesky decomposition. BIT Numerical Mathematics, 36(2):247–263, June 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=247.

Gutknecht:1996:OLA

[2164] Martin H. Gutknecht and Marlis Hochbruck. Optimized look-ahead recurrences for adjacent rows in the Padé table. BIT Numerical Mathematics, 36(2):264–286, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=264.

Hanke:1996:LCM

[2165] Martin Hanke. Limitations of the L-curve method in ill-posed problems. BIT Numerical Mathematics, 36(2):287–301, June 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=287.

Luk:1996:FRR

[2166] W. S. Luk. Finding roots of a real polynomial simultaneously by means of Bairstow’s method. BIT Numerical Mathematics, 36(2):302–308, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=302.

Marthinsen:1996:CEN

[2167] Arne Marthinsen. Continuous extensions to Nyström methods for second order initial value problems. BIT Numerical Mathematics, 36(2):309–332, June 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=309.

Rahola:1996:DFT

[2168] Jussi Rahola. Diagonal forms of the translation operators in the fast multipole algorithm for scattering problems. BIT Numerical Mathematics, 36(2):333–358, June 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=333.

Sakai:1996:FPR

[2169] M. Sakai and R. A. Usmani. On fair parametric rational cubic curves. BIT Numerical Mathematics, 36 (2):359–377, June 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=359.

Sand:1996:SSB

[2170] Jørgen Sand. On some stability bounds subject to Hille–Yosida resolvent conditions. BIT Numerical Mathematics, 36(2):378–386, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=378.

Venter:1996:AQF

[2171] Antoinetta Venter and Dirk P. Laurie. Automatic quadrature of functions of the form g(|f(x)|). BIT Numerical Mathematics, 36(2):387–394, June 1996. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=387.

Kalantari:1996:HOI

[2172] B. Kalantari and I. Kalantari. High order iterative methods for approximating square roots. BIT Numerical Mathematics, 36(2):395–399, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=2&spage=395.

Baglama:1996:IMC

[2173] J. Baglama, D. Calvetti, and L. Reichel. Iterative methods for the computation of a few eigenvalues of a large symmetric matrix. BIT Numerical Mathematics, 36(3):400–421, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=3&spage=400. In-

ternational Linear Algebra Year (Toulouse, 1995).

Braconnier:1996:CFV

[2174] Thierry Braconnier and Nicholas J.

Higham. Computing the field of values and pseudospectra using the Lánczos method with continuation. BIT Numerical Mathematics, 36(3): 422–440, September 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=422. In-

ternational Linear Algebra Year (Toulouse, 1995).

Bruhl:1996:CTA

[2175] Martin Brühl. A curve tracing algorithm for computing the pseudospectrum. BIT Numerical Mathematics, 36(3):441–454, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=441. In-

ternational Linear Algebra Year (Toulouse, 1995).

Chaitin-Chatelin:1996:CFP

[2176] F. Chaitin-Chatelin and S. Gratton. Convergence in finite precision of successive iteration methods under high nonnormality. BIT Numerical Mathematics, 36(3):455–469, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=455. In-

ternational Linear Algebra Year (Toulouse, 1995).

Cullum:1996:AVN

[2177] Jane Cullum. Arnoldi versus nonsymmetric Lánczos algorithms for solving matrix eigenvalue problems. BIT Numerical Mathematics, 36(3): 470–493, September 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=3&spage=470. In-

ternational Linear Algebra Year (Toulouse, 1995).

Edelman:1996:CGL

[2178] Alan Edelman and Steven T. Smith.

On conjugate gradient-like methods for eigen-like problems. BIT Numerical Mathematics, 36(3):494–508, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=3&spage=494. In-

ternational Linear Algebra Year (Toulouse, 1995).

Fattebert:1996:IIM

[2179] Jean-Luc Fattebert. An inverse iteration method using multigrid for quantum chemistry. BIT Numerical Mathematics, 36(3):509–522, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=3&spage=509. In-

ternational Linear Algebra Year (Toulouse, 1995).

Gratton:1996:CNL

[2180] S. Gratton. On the condition number of linear least squares problems in a weighted Frobenius norm. BIT Numerical Mathematics, 36(3): 523–530, September 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=523. In-

ternational Linear Algebra Year (Toulouse, 1995).

Nevanlinna:1996:MRP

[2181] Olavi Nevanlinna. Meromorphic resolvents and power bounded operators. BIT Numerical Mathematics, 36(3):531–541, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=531. In-

ternational Linear Algebra Year (Toulouse, 1995).

Parlett:1996:IST

[2182] B. N. Parlett. Invariant subspaces for tightly clustered eigenvalues of tridiagonals. BIT Numerical Mathematics, 36(3):542–562, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=542. In-

ternational Linear Algebra Year (Toulouse, 1995).

Saad:1996:SLE

[2183] Y. Saad, A. Stathopoulos, J. Chelikowsky, K. Wu, and S. Öğüt. Solution of large eigenvalue problems in electronic structure calculations. BIT Numerical Mathematics, 36(3):563–578, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=3&spage=563. In-

ternational Linear Algebra Year (Toulouse, 1995).

Simoncini:1996:ARM

[2184] V. Simoncini and M. Sadkane. Arnoldi– Riccati method for large eigenvalue problems. BIT Numerical Mathematics, 36(3):579–594, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=3&spage=579. In-

ternational Linear Algebra Year (Toulouse, 1995).

Sleijpen:1996:JDT

[2185] Gerard L. G. Sleijpen, Albert G. L.

Booten, Diederik R. Fokkema, and Henk A. Van der Vorst. Jacobi– Davidson type methods for generalized eigenproblems and polynomial eigenproblems. BIT Numerical Mathematics, 36(3):595–633, September 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=3&spage=595. In-

ternational Linear Algebra Year (Toulouse, 1995).

Auzinger:1996:ECT

[2186] W. Auzinger, R. Frank, and G. Kirlinger. Extending convergence theory for nonlinear stiff problems. I. BIT Numerical Mathematics, 36(4): 635–652, December 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=635.

Bendtsen:1996:CSA

[2187] C. Bendtsen. On the construction of stiffly accurate and B-stable Runge– Kutta methods. BIT Numerical Mathematics, 36(4):653–663, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=653.

Campbell:1996:GMP

[2188] S. L. Campbell, I. C. F. Ipsen, C. T. Kelley, and C. D. Meyer. GMRES and the minimal polynomial. BIT Numerical Mathematics, 36(4):664–675, December 1996.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=664.

Costabile:1996:AEE

[2189] F. Costabile, M. I. Gualtieri, and S. Serra. Asymptotic expansion and extrapolation for Bernstein polynomials with applications. BIT Numerical Mathematics, 36(4):676–687, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=676.

Jackiewicz:1996:GLM

[2190] Z. Jackiewicz and R. Vermiglio.

General linear methods with external stages of different orders. BIT Numerical Mathematics, 36(4): 688–712, December 1996. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.
liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=688.

Jackson:1996:AOR

[2191] K. R. Jackson, A. Kværnø, and S. P.

Nørsett. An analysis of the order of Runge–Kutta methods that use an iterative scheme to compute their internal stage values. BIT Numerical Mathematics, 36(4):713–765, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=713.

MacDonald:1996:NSS

[2192] D. A. MacDonald. A note on the summation of slowly convergent alternating series. BIT Numerical Mathematics, 36(4):766–774, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=766.

Nevanlinna:1996:CKM

[2193] Olavi Nevanlinna. Convergence of Krylov methods for sums of two operators. BIT Numerical Mathematics, 36(4):775–785, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.mai.
                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=775.

Shen:1996:SCL

[2194] Wen Shen, Aslak Tveito, and Ragnar Winther. A system of conservation laws including a stiff relaxation term; the 2D case. BIT Numerical Mathematics, 36(4):786–813, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=786.

Simoncini:1996:NS

[2195] V. Simoncini. On the numerical solution of AX − XB = C. BIT Numerical Mathematics, 36(4):814–830, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=814.

Taasan:1996:FLA

[2196] Shlomo Ta’asan and Hong Zhang.

Fourier–Laplace analysis of the multigrid waveform relaxation method for hyperbolic equations. BIT Numerical Mathematics, 36(4):831–841, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=831.

Varah:1996:LSD

[2197] J. M. Varah. Least squares data fitting with implicit functions. BIT Numerical Mathematics, 36(4):842–854, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit36.html;
 http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=36&issue=4&spage=842.

Koto:1996:SPS

[2198] Toshiyuki Koto. A stability property of A-stable collocation-based Runge– Kutta methods for neutral delay differential equations. BIT Numerical Mathematics, 36(4):855–859, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=855.

Riesel:1996:SDS

[2199] Hans Riesel. Summation of double series using the Euler–Maclaurin sum formula. BIT Numerical Mathematics, 36(4):860–862, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit36.html;
                                         http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=860.

Anonymous:1996:A

[2200] Anonymous. Acknowledgements. BIT Numerical Mathematics, 36(4):863, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=36&issue=4&spage=863.

Brunner:1997:DDV

[2201] Hermann Brunner. On the discretization of differential and Volterra integral equations with variable delay. BIT Numerical Mathematics, 37(1):1–12, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit37.html.

Edlund:1997:LEB

[2202] Ove Edlund. Linear M-estimation with bounded variables. BIT Numerical Mathematics, 37(1):13–23, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Higham:1997:TSP

[2203] Desmond J. Higham. Time-stepping and preserving orthonormality. BIT Numerical Mathematics, 37(1):24–36, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.mai.liu.se/BIT/contents/bit37.
html.

Hill:1997:DRK

[2204] Adrian T. Hill. Dissipativity of Runge–Kutta methods in Hilbert spaces. BIT Numerical Mathematics, 37(1):37–42, March 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit37.html.

Komori:1997:RTA

[2205] Yoshio Komori, Taketomo Mitsui, and Hiroshi Sugiura. Rooted tree analysis of the order conditions of ROWtype scheme for stochastic differential equations. BIT Numerical Mathematics, 37(1):43–66, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Li:1997:RPB

[2206] Ren-Cang Li. Relative perturbation bounds for the unitary polar factor. BIT Numerical Mathematics, 37(1):67–75, March 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Quak:1997:ASW

[2207] Ewald Quak and Norman Weyrich. Algorithms for spline wavelet packets on an interval. BIT Numerical Mathematics, 37(1):76–95, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Saunders:1997:CPL

[2208] Michael A. Saunders. Computing projections with LSQR. BIT Numerical Mathematics, 37(1):96–104, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit37.html.

Schutze:1997:CAS

[2209] Torsten Schütze and Hubert Schwetlick. Constrained approximation by splines with free knots. BIT Numerical Mathematics, 37(1):105–137, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Simpson:1997:DMA

[2210] R. Bruce Simpson. A data modeling abstraction for describing triangular mesh algorithms. BIT Numerical Mathematics, 37(1):138–163, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Sterner:1997:SIR

[2211] E. Sterner. Semi-implicit Runge– Kutta schemes for the Navier–Stokes equations. BIT Numerical Mathematics, 37(1):164–178, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Sun:1997:OBP

[2212] Ji-Guang Sun. On optimal backward perturbation bounds for the linear least squares problem. BIT Numerical Mathematics, 37(1):179–188, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Walz:1997:ITS

[2213] Guido Walz. Identities for trigonometric B-splines with an application to curve design. BIT Numerical Mathematics, 37(1):189–201, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Zha:1997:COC

[2214] Hongyuan Zha and Zhenyue Zhang.

Computing the optimal commuting matrix pairs. BIT Numerical Mathematics, 37(1):202–220, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Bendtsen:1997:IRK

[2215] C. Bendtsen. On implicit Runge– Kutta methods with high stage order. BIT Numerical Mathematics, 37(1):221–226, March 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Brass:1997:CDB

[2216] Helmut Brass and Knut Petras. On a conjecture of D. B. Hunter. BIT Numerical Mathematics, 37(1): 227–231, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit37.html.

Phillips:1997:CAG

[2217] George M. Phillips. A de Casteljau algorithm for generalized Bernstein polynomials. BIT Numerical Mathematics, 37(1):232–236, March 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Bakaev:1997:GFE

[2218] N. Yu. Bakaev. On the Galerkin finite element approximations to multi-dimensional differential and integro-differential parabolic equations. BIT Numerical Mathematics, 37(2):237–255, June 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit37.html.

Dahlquist:1997:SFDa

[2219] Germund Dahlquist. On summation formulas due to Plana, Lindelöf and Abel, and related Gauss– Christoffel rules. I. BIT Numerical Mathematics, 37(2):256–295, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Dryja:1997:NDD

[2220] M. Dryja and W. Hackbusch. On the nonlinear domain decomposition method. BIT Numerical Mathematics, 37(2):296–311, June 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Feng:1997:CCS

[2221] Yu Yu Feng and Jernej Kozak. On G^{2} continuous cubic spline interpolation. BIT Numerical Mathematics, 37(2):312–332, June 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Gemignani:1997:PRC

[2222] Luca Gemignani. Polynomial root computation by means of the LR algorithm. BIT Numerical Mathematics, 37(2):333–345, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Hoffmann:1997:ARK

[2223] W. Hoffmann and J. J. B. de Swart. Approximating Runge–Kutta matrices by triangular matrices. BIT Numerical Mathematics, 37(2):346–354, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit37.html.

Lee:1997:TLO

[2224] David Lee and Henryk Woźniakowski. Testing linear operators — an average case study. BIT Numerical Mathematics, 37(2):355–376, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Leyk:1997:BSI

[2225] Zbigniew Leyk. Breakdowns and stagnation in iterative methods. BIT Numerical Mathematics, 37(2): 377–403, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit37.html.

Plato:1997:GSL

[2226] Robert Plato. The Galerkin scheme for Lavrentiev’s m-times iterated method to solve linear accretive Volterra integral equations of the first kind. BIT Numerical Mathematics, 37(2):404–423, June 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit37.html.

Schneider:1997:CGL

[2227] Stefan Schneider. Convergence of general linear methods on differentialalgebraic systems of index 3. BIT Numerical Mathematics, 37(2):424– 441, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.mai.liu.se/BIT/contents/bit37.
html.

Spijker:1997:EGA

[2228] M. N. Spijker and F. A. J. Straetemans. Error growth analysis via stability regions for discretizations of initial value problems. BIT Numerical Mathematics, 37(2):442–464, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Ramaswami:1997:HOS

[2229] Geetha Ramaswami. Higher order symplectic RK and RKN methods using perturbed collocation. BIT Numerical Mathematics, 37(2):465–471, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Anonymous:1997:NB

[2230] Anonymous. New books. BIT Numerical Mathematics, 37(2):472, June 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Alvarado:1997:MEM

[2231] Fernando L. Alvarado. Matrix enlarging methods and their application. BIT Numerical Mathematics, 37(3):473–505, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Ashcraft:1997:UDD

[2232] Cleve Ashcraft and Joseph W. H. Liu. Using domain decomposition to find graph bisectors. BIT Numerical Mathematics, 37(3):506–534, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Hansen:1997:SSL

[2233] P. C. Hansen, Tz. Ostromsky, A. Sameh, and Z. Zlatev. Solving sparse linear least-squares problems on some supercomputers by using large dense blocks. BIT Numerical Mathematics, 37(3):535–558, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Kumfert:1997:TIA

[2234] Gary Kumfert and Alex Pothen. Two improved algorithms for envelope and wavefront reduction. BIT Numerical Mathematics, 37(3):559–590, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Burdakov:1997:GAO

[2235] Oleg Burdakov. A greedy algorithm for the optimal basis problem. BIT Numerical Mathematics, 37(3):591–599, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Dax:1997:PPA

[2236] Achiya Dax. A proximal point algorithm for minimax problems. BIT Numerical Mathematics, 37(3): 600–622, September 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Brenner:1997:MMC

[2237] S. C. Brenner and L.-Y. Sung. Multigrid methods for the computation of singular solutions and stress intensity factors. II. Crack singularities. BIT Numerical Mathematics, 37(3):623–643, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Chan:1997:WSA

[2238] T. F. Chan, W. P. Tang, and W. L. Wan. Wavelet sparse approximate inverse preconditioners. BIT Numerical Mathematics, 37(3):644–660, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Douglas:1997:CMU

[2239] C. C. Douglas, S. Malhotra, and M. H. Schultz. A characterization of mapping unstructured grids onto structured grids and using multigrid as a preconditioner. BIT Numerical Mathematics, 37(3):661–677, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Douglas:1997:ADD

[2240] J. Douglas, Jr. and C.-S. Huang. An accelerated domain decomposition procedure based on Robin transmission conditions. BIT Numerical Mathematics, 37(3):678–686, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Golub:1997:MMQ

[2241] G. H. Golub and G. Meurant. Matrices, moments and quadrature. II. How to compute the norm of the error in iterative methods. BIT Numerical Mathematics, 37(3):687–705, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Greenbaum:1997:NBM

[2242] A. Greenbaum, M. Rozložník, and Z. Strakoš. Numerical behaviour of the modified Gram–Schmidt GM-RES implementation. BIT Numerical Mathematics, 37(3):706–719, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Lacour:1997:TDA

[2243] C. Lacour and Y. Maday. Two different approaches for matching nonconforming grids: the mortar element method and the FETI method. BIT Numerical Mathematics, 37(3): 720–738, September 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Lotzbeyer:1997:PAM

[2244] H. Lötzbeyer and U. Rüde. Patchadaptive multilevel iteration. BIT Numerical Mathematics, 37(3):739– 758, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit37.html. Direct meth-

ods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Rottner:1997:NSF

[2245] Th. Rottner, I. Lenhardt, G. Alefeld, and K. Schweizerhof. Nonlinear structural finite element analysis using the preconditioned Lánczos method on serial and parallel computers. BIT Numerical Mathematics, 37(3):759–769, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html. Di-

rect methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Burrage:1997:BMS

[2246] K. Burrage, P. M. Burrage, and J. A. Belward. A bound on the maximum strong order of stochastic Runge–Kutta methods for stochastic ordinary differential equations. BIT Numerical Mathematics, 37(4):771–780, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Christara:1997:MMM

[2247] Christina C. Christara and Barry Smith. Multigrid and multilevel methods for quadratic spline collocation. BIT Numerical Mathematics, 37(4):781–803, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Dahlquist:1997:SFDb

[2248] Germund Dahlquist. On summation formulas due to Plana, Lindelöf and Abel, and related Gauss–Christoffel rules. II. BIT Numerical Mathematics, 37(4):804–832, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

vanDorsselaer:1997:PMP

[2249] Jos L. M. van Dorsselaer. Pseudospectra for matrix pencils and stability of equilibria. BIT Numerical Mathematics, 37(4):833–845, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Entacher:1997:QMC

[2250] Karl Entacher. Quasi-Monte Carlo methods for numerical integration of multivariate Haar series. BIT Numerical Mathematics, 37(4):846– 861, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit37.html.

Karlson:1997:EOB

[2251] Rune Karlson and Bertil Waldén. Estimation of optimal backward perturbation bounds for the linear least squares problem. BIT Numerical Mathematics, 37(4):862–869, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Koto:1997:NSR

[2252] T. Koto. NP-stability of Runge– Kutta methods based on classical quadrature. BIT Numerical Mathematics, 37(4):870–884, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Ng:1997:PEP

[2253] Michael K. Ng. Preconditioning of elliptic problems by approximation in the transform domain. BIT Numerical Mathematics, 37(4):885–900, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit37.html.

Oja:1997:LDR

[2254] Peeter Oja. Low degree rational spline interpolation. BIT Numerical Mathematics, 37(4):901–909, December 1997.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Pena:1997:BSP

[2255] J. M. Peña. Backward stability of a pivoting strategy for sign-regular linear systems. BIT Numerical Mathematics, 37(4):910–924, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

DeSamblanx:1997:IAR

[2256] G. De Samblanx, K. Meerbergen, and A. Bultheel. The implicit application of a rational filter in the RKS method. BIT Numerical Mathematics, 37(4):925–947, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Schmidt:1997:FUB

[2257] Jochen W. Schmidt and Walter Hess.

Fair upper bounds for the curvature in univariate convex interpolation. BIT Numerical Mathematics, 37(4):948–960, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit37.html.

Stewart:1997:WML

[2258] G. W. Stewart. On the weighting method for least squares problems with linear equality constraints. BIT Numerical Mathematics, 37(4): 961–967, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit37.html.

Tuomela:1997:SPQ

[2259] Jukka Tuomela. On singular points of quasilinear differential and differentialalgebraic equations. BIT Numerical Mathematics, 37(4):968–977, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit37.html.

Zhou:1997:MPE

[2260] Aihui Zhou. Multi-parameter error resolution for the collocation method of Volterra integral equations. BIT Numerical Mathematics, 37(4):978–987, December 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit37.html.

Anonymous:1997:PA

[2261] Anonymous. Publisher’s announcement. BIT Numerical Mathematics, 37(4):988, December 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit37.html.

Anonymous:1997:A

[2262] Anonymous. Acknowledgements. BIT Numerical Mathematics, 37(4): 989, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit37.html.

Anonymous:1997:I

[2263] Anonymous. Index. BIT Numerical Mathematics, 37(4):990–993, December 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit37.html.

Bhatia:1998:ESM

[2264] Rajendra Bhatia, Fuad Kittaneh, and Ren-Cang C. Li. Eigenvalues of symmetrizable matrices. BIT Numerical Mathematics, 38(1):1–11, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Burnton:1998:GRK

[2265] Christopher Burnton and Rudolf Scherer. Gauss–Runge–Kutta–Nyström methods. BIT Numerical Mathematics, 38(1):12–21, March 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Gallestey:1998:CSV

[2266] E. Gallestey. Computing spectral value sets using the subharmonicity of the norm of rational matrices. BIT Numerical Mathematics, 38(1):22–33, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Hansen:1998:RDP

[2267] Per Christian Hansen. Rank-deficient prewhitening with quotient SVD and ULV decompositions. BIT Numerical Mathematics, 38(1):34–43, March 1998.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Higham:1998:ADL

[2268] D. J. Higham and A. M. Stuart. Analysis of the dynamics of local error control via a piecewise continuous residual. BIT Numerical Mathematics, 38(1):44–57, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Huhtanen:1998:ARP

[2269] M. Huhtanen. An alternative to the Ritz pairs with the EN subspace method. BIT Numerical Mathematics, 38(1):58–71, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

intHout:1998:POD

[2270] Karel in ’t Hout and Christian Lubich. Periodic orbits of delay differential equations under discretization. BIT Numerical Mathematics, 38(1):72–91, March 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Munthe-Kaas:1998:RKM

[2271] Hans Munthe-Kaas. Runge–Kutta methods on Lie groups. BIT Numerical Mathematics, 38(1):92–111, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Petkovic:1998:PES

[2272] M. Petković, – D. Herceg, and S. Ilić.

Point estimation and some applications to iterative methods. BIT Numerical Mathematics, 38(1):112–126, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Pinar:1998:PCM

[2273] Mustafa C. Pinar and Samir Elhedhli.

A penalty continuation method for the l_{∞} solution of overdetermined linear systems. BIT Numerical Mathematics, 38(1):127–150, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Popa:1998:EBP

[2274] Constantin Popa. Extensions of block-projections methods with relaxation parameters to inconsistent and rank-deficient least-squares problems. BIT Numerical Mathematics, 38(1):151–176, March 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Shapira:1998:CUM

[2275] Yair Shapira. Coloring update methods. BIT Numerical Mathematics, 38(1):177–185, March 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Yamamura:1998:ISN

[2276] Kiyotaka Yamamura, Hitomi Kawata, and Ai Tokue. Interval solution of nonlinear equations using linear programming. BIT Numerical Mathematics, 38(1):186–199, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Zha:1998:MGS

[2277] Hongyuan Zha and Zhenyue Zhang. Modifying the generalized singular value decomposition with application in direction-of-arrival finding. BIT Numerical Mathematics, 38(1): 200–216, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit38.html.

Britz:1998:EPN

[2278] Dieter Britz. An error propagation in the numerical literature. BIT Numerical Mathematics, 38(1):217–218, March 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit38.html.

Arunsawatwong:1998:SZR

[2279] Suchin Arunsawatwong. Stability of Zakian I_{MN} recursions for linear delay differential equations. BIT Numerical Mathematics, 38(2):219–233, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit38.html.

Benzi:1998:NET

[2280] Michele Benzi and Miroslav Tůma.

Numerical experiments with two approximate inverse preconditioners. BIT Numerical Mathematics, 38(2): 234–241, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Cai:1998:ACI

[2281] Xing Cai and Even Mehlum. Application of Cauchy integrals and singular integral equations in scattered data problems. BIT Numerical Mathematics, 38(2):242–255, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Criscuolo:1998:NEC

[2282] G. Criscuolo and L. Scuderi. The numerical evaluation of Cauchy principal value integrals with non-standard weight functions. BIT Numerical Mathematics, 38(2):256–274, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Elfving:1998:SIP

[2283] Tommy Elfving. A stationary iterative pseudoinverse algorithm. BIT Numerical Mathematics, 38(2):275–282, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Entacher:1998:QMC

[2284] Karl Entacher. Quasi-Monte Carlo methods for numerical integration of multivariate Haar series. II. BIT Numerical Mathematics, 38(2): 283–292, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Gustafsson:1998:IBC

[2285] Bertil Gustafsson. On the implementation of boundary conditions for the method of lines. BIT Numerical Mathematics, 38(2):293–314, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Iserles:1998:RKM

[2286] Arieh Iserles, Geetha Ramaswami, and Mark Sofroniou. Runge–Kutta methods for quadratic ordinary differential equations. BIT Numerical Mathematics, 38(2):315–346, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Li:1998:GSO

[2287] Changjun Li, Baojia Li, and David J. Evans. A generalized successive overrelaxation method for least squares problems. BIT Numerical Mathematics, 38(2):347–355, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Malinen:1998:PIT

[2288] J. Malinen. Properties of iteration of Toeplitz operators with Toeplitz preconditioners. BIT Numerical Mathematics, 38(2):356–371, June 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Medovikov:1998:HOE

[2289] Alexei A. Medovikov. High order explicit methods for parabolic equations. BIT Numerical Mathematics, 38(2):372–390, June 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Schmitt:1998:EAK

[2290] Bernhard A. Schmitt and Rüdiger Weiner. Equilibrium attractivity of Krylov-W -methods for nonlinear stiff ODEs. BIT Numerical Mathematics, 38(2):391–414, June 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Arnold:1998:HER

[2291] M. Arnold. Half-explicit Runge– Kutta methods with explicit stages for differential-algebraic systems of index 2. BIT Numerical Mathematics, 38(3):415–438, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Aubry:1998:PSR

[2292] A. Aubry and P. Chartier. Pseudosymplectic Runge–Kutta methods. BIT Numerical Mathematics, 38(3): 439–461, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/BIT/
                                        contents/bit38.html.

Bakaev:1998:VSA

[2293] N. Yu. Bakaev. On the variable stepsize approximations of a Cauchy problem for the evolution equation. BIT Numerical Mathematics, 38(3): 462–485, September 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Chang:1998:SLF

[2294] X.-W. Chang and C. C. Paige. On the sensitivity of the LU factorization. BIT Numerical Mathematics, 38(3):486–501, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Eisenstat:1998:RPR

[2295] S. C. Eisenstat and I. C. F. Ipsen. Relative perturbation results for eigenvalues and eigenvectors of diagonalisable matrices. BIT Numerical Mathematics, 38(3):502–509, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Favati:1998:RFI

[2296] P. Favati and B. Meini. Relaxed functional iteration techniques for the numerical solution of M/G/1 type Markov chains. BIT Numerical Mathematics, 38(3):510–526, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Fisher:1998:MRM

[2297] B. Fisher, A. Ramage, D. Silvester, and A. J. Wathen. Minimum residual methods for augmented systems. BIT Numerical Mathematics, 38(3): 527–543, September 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Holmgren:1998:FPPa

[2298] S. Holmgren and K. Otto. A framework for polynomial preconditioners based on fast transforms I: Theory. BIT Numerical Mathematics, 38(3):544–559, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Kamm:1998:TLS

[2299] J. Kamm and J. G. Nagy. A total least squares method for Toeplitz systems of equations. BIT Numerical Mathematics, 38(3):560–582, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Kravanja:1998:CZA

[2300] P. Kravanja, R. Cools, and A. Haegemans. Computing zeros of analytic mappings: a logarithmic residue approach. BIT Numerical Mathematics, 38(3):583–596, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Thomee:1998:NMB

[2301] V. Thomée and A. S. Vasudeva Murthy. A numerical method for the Benjamin– Ono equation. BIT Numerical Mathematics, 38(3):597–611, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit38.html.

Psihoyios:1998:SRG

[2302] G.-Y. Psihoyios and J. R. Cash. A stability result for general linear methods with characteristic function having real poles only. BIT Numerical Mathematics, 38(3):612–617, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Anonymous:1998:AHA

[2303] Anonymous. Announcement: Householder award X. BIT Numerical Mathematics, 38(3):618, September 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Alefeld:1998:MIN

[2304] G. E. Alefeld, F. A. Potra, and W. Völker. Modifications of the interval-Newton-method with improved asymptotic efficiency. BIT Numerical Mathematics, 38(4):619– 635, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Arioli:1998:KSM

[2305] M. Arioli, V. Pták, and Z. Strakos.

Krylov sequences of maximal length and convergence of GMRES. BIT Numerical Mathematics, 38(4):636– 643, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Becker:1998:SOB

[2306] J. Becker. A second order backward difference method with variable steps for a parabolic problem. BIT Numerical Mathematics, 38(4):644–662, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Cao:1998:MOM

[2307] Y. Cao and Q. Li. The maximum order of multistep formula for solving index-2 differential-algebraic equations. BIT Numerical Mathematics, 38(4):663–673, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Chen:1998:NMH

[2308] B. Chen and M. C. Pinar. On Newton’s method for Huber’s robust M-estimation problems in linear regression. BIT Numerical Mathematics, 38(4):674–684, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Dhillon:1998:CII

[2309] I. Dhillon. Current inverse iteration software can fail. BIT Numerical Mathematics, 38(4):685–704, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

                                        liu.se/BIT/contents/bit38.html.

Floater:1998:TAS

[2310] M. S. Floater and A. Iske. Thinning algorithms for scattered data interpolation. BIT Numerical Mathematics, 38(4):705–720, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Holmgren:1998:FPPb

[2311] S. Holmgren and K. Otto. A framework for polynomial preconditioners based on fast transforms II: PDE applications. BIT Numerical Mathematics, 38(4):721–736, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://www.mai.liu.
                                        se/BIT/contents/bit38.html.

Koto:1998:CSP

[2312] T. Koto. A criterion for p-stability properties of Runge–Kutta methods. BIT Numerical Mathematics, 38(4): 737–750, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Lamba:1998:CRM

[2313] H. Lamba and A. M. Stuart. Convergence results for the MATLAB ode23 routine. BIT Numerical Mathematics, 38(4):751–780, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Yoon:1998:ERD

[2314] P. A. Yoon and J. L. Barlow. An efficient rank detection procedure for modifying the ULV decomposition. BIT Numerical Mathematics, 38(4): 781–801, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Aubry:1998:NPS

[2315] A. Aubry and P. Chartier. A note on pseudo-symplectic Runge–Kutta methods. BIT Numerical Mathematics, 38 (4):802–806, December 1998. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html.

Ikramov:1998:RNC

[2316] K. D. Ikramov. A remark on “A note on constructing a symmetric matrix with specified diagonal entries and eigenvalues” [BIT 35(3), September 1995, pp. 448–451]. BIT Numerical Mathematics, 38(4):807, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.mai.liu.
se/BIT/contents/bit38.html. See

[2129].

Anonymous:1998:A

[2317] Anonymous. Acknowledgements. BIT Numerical Mathematics, 38(4): 808, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/
contents/bit38.html.

Anonymous:1998:I

[2318] Anonymous. Index. BIT Numerical Mathematics, 38(4):809–812, December 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.mai.

liu.se/BIT/contents/bit38.html.

Bellen:1999:CAS

[2319] A. Bellen, N. Guglielmi, and M. Zennaro. On the contractivity and asymptotic stability of systems of delay differential equations of neutral type. BIT Numerical Mathematics, 39(1):1– 24, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=1.

Cirilli:1999:AEA

[2320] Stéphane Cirilli, Ernst Hairer, and Benedict Leimkuhler. Asymptotic error analysis of the adaptive Verlet method. BIT Numerical Mathematics, 39(1):25– 33, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=25.

Cox:1999:ASN

[2321] Anthony J. Cox and Nicholas J.

Higham. Accuracy and stability of the null space method for solving the equality constrained least squares problem. BIT Numerical Mathematics, 39(1):34– 50, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=34.

Dahlquist:1999:SFD

[2322] Germund Dahlquist. On summation formulas due to Plana, Lindelof and Abel, and related Gauss– Christoffel rules, III. BIT Numerical Mathematics, 39(1):51–78, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=51.

Gonzalez:1999:OCR

[2323] Cesáreo González and Alexander Ostermann. Optimal convergence results for Runge–Kutta discretizations of linear nonautonomous parabolic problems. BIT Numerical Mathematics, 39(1):79–95, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=79.

Kalantari:1999:OCD

[2324] Bahman Kalantari. On the order of convergence of a determinantal family of root-finding methods. BIT Numerical Mathematics, 39(1):96–109, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=96.

Koto:1999:NSB

[2325] T. Koto. Naimark–Sacker bifurcations in the Euler method for a delay differential equation. BIT Numerical Mathematics, 39(1):110–115, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=110.

Owren:1999:RKM

[2326] B. Owren and A. Marthinsen. Runge– Kutta methods adapted to manifolds and based on rigid frames. BIT Numerical Mathematics, 39(1):116– 142, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=116.

Rump:1999:ICN

[2327] Siegfried M. Rump. Ill-conditionedness need not be componentwise near to ill-posedness for least squares problems. BIT Numerical Mathematics, 39(1):143–151, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=143.

Serra:1999:SCA

[2328] S. Serra. Spectral and computational analysis of block Toeplitz matrices having nonnegative definite matrixvalued generating functions. BIT Numerical Mathematics, 39(1):152– 175, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=1&spage=152.

Watson:1999:LSF

[2329] G. A. Watson. Least squares fitting of circles and ellipses to measured data. BIT Numerical Mathematics, 39(1):176–191, March 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=39&
issue=1&spage=176.

Brunner:1999:OML

[2330] D. Brunner and A. Uhl. Optimal multipliers for linear congruential pseudo random number generators with prime moduli: Parallel computation and properties. BIT Numerical Mathematics, 39(2):193– 209, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=193.

Cox:1999:BEB

[2331] Anthony J. Cox and Nicholas J. Higham. Backward error bounds for constrained least squares problems. BIT Numerical Mathematics, 39(2): 210–227, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=210.

Eriksson:1999:QNM

[2332] Jerry Eriksson. Quasi-Newton methods for nonlinear least squares focusing on curvatures. BIT Numerical Mathematics, 39(2):228–254, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=228.

Hernandez:1999:HSN

[2333] M. A. Hernández and M. A. Salanova. How to solve nonlinear equations when a third-order method is not applicable. BIT Numerical Mathematics, 39(2): 255–269, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=255.

Huang:1999:SAR

[2334] Chengming Huang, Hongyuan Fu, Shoufu Li, and Guangnan Chen. Stability analysis of Runge–Kutta methods for non-linear delay differential equations. BIT Numerical Mathematics, 39(2):270–280, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=39&
issue=2&spage=270.

Iserles:1999:IMM

[2335] A. Iserles, A. Marthinsen, and S. P.

Nørsett. On the implementation of the method of Magnus series for linear differential equations. BIT Numerical Mathematics, 39(2):281– 304, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=281.

Lyche:1999:SSF

[2336] T. Lyche and K. Mørken. The sensitivity of a spline function to perturbations of the knots. BIT Numerical Mathematics, 39(2):305– 322, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=305.

Popa:1999:BPA

[2337] Constantin Popa. Block-projections algorithms with blocks containing mutually orthogonal rows and columns. BIT Numerical Mathematics, 39(2): 323–338, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=2&spage=323.

Shi:1999:BNP

[2338] Xiquan Shi and Renhong Wang. The Bezout number for piecewise algebraic curves. BIT Numerical Mathematics, 39(2):339–349, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=39&
                                        issue=2&spage=339.

Simoncini:1999:RNL

[2339] V. Simoncini. Remarks on nonlinear spectral perturbation. BIT Numerical Mathematics, 39(2):350– 365, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=2&spage=350.

Herzberger:1999:SNB

[2340] Jürgen Herzberger. Scientific notes — bounds for the positive root of a class of polynomials with applications. BIT Numerical Mathematics, 39(2): 366–372, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=366.

Song:1999:NEB

[2341] Yongzhong Song. A note on an error bound for the AOR method. BIT Numerical Mathematics, 39(2):373– 383, June 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=2&spage=373.

Amodio:1999:NAB

[2342] P. Amodio and F. Mazzia. A new approach to backward error analysis of LU factorization. BIT Numerical Mathematics, 39(3):385–402, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=3&spage=385. See

note [2452].

Bar-On:1999:RSB

[2343] Ilan Bar-On and Mauro Leoncini.

Reliable solution of bidiagonal systems with applications. BIT Numerical Mathematics, 39(3):403–416, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=3&spage=403.

Benzi:1999:BEM

[2344] Michele Benzi and Gene H. Golub. Bounds for the entries of matrix functions with applications to preconditioning. BIT Numerical Mathematics, 39(3):417–438, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=3&spage=417.

Delvos:1999:LIA

[2345] Franz-Jürgen Delvos and Ludger Knoche. Lacunary interpolation by antiperiodic trigonometric polynomials. BIT Numerical Mathematics, 39(3):439–450, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=39&
                                        issue=3&spage=439.

Holdahl:1999:USS

[2346] Runar Holdahl, Helge Holden, and Knut-Andreas Lie. Unconditionally stable splitting methods for the shallow water equations. BIT Numerical Mathematics, 39(3):451–472, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=3&spage=451.

Hundsdorfer:1999:SAF

[2347] Willem Hundsdorfer. Stability of approximate factorization with θmethods. BIT Numerical Mathematics, 39(3):473–483, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=39&
issue=3&spage=473.

Lucht:1999:ISD

[2348] W. Lucht, K. Strehmel, and C. Eichler-Liebenow. Indexes and special discretization methods for linear partial differential algebraic equations. BIT Numerical Mathematics, 39(3):484–512, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=39&
issue=3&spage=484.

Potts:1999:PIC

[2349] Daniel Potts and Gabrielle Steidl.

Preconditioners for ill-conditioned Toeplitz matrices. BIT Numerical Mathematics, 39(3):513–533, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=3&spage=513.

Rump:1999:FPI

[2350] Siegfried M. Rump. Fast and parallel interval arithmetic. BIT Numerical Mathematics, 39(3):534–554, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=3&spage=534.

Zhang:1999:PID

[2351] Jian-Zhong Zhang and Cheng-Xian Xu. A projected indefinite dogleg path method for equality constrained optimization. BIT Numerical Mathematics, 39(3):555–578, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=3&spage=555.

Perez-Acosta:1999:CGQ

[2352] F. Pérez-Acosta and J. C. Santos-Léon. Convergence of Gaussian quadrature and Lagrange interpolation in Haar systems. BIT Numerical Mathematics, 39(3):579–584, September 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=3&spage=579.

Amini:1999:ADF

[2353] S. Amini and A. T. J. Profit. Analysis of a diagonal form of the fast multipole algorithm for scattering theory. BIT Numerical Mathematics, 39(4):585–602, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=39&issue=4&spage=585.

Calvetti:1999:ECL

[2354] D. Calvetti, G. H. Golub, and L. Reichel. Estimation of the L-curve via Lánczos bidiagonalization. BIT Numerical Mathematics, 39(4):603–619, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=603.

Hochbruck:1999:EIQ

[2355] Marlis Hochbruck and Christian Lubich. Exponential integrators for quantum-classical molecular dynamics. BIT Numerical Mathematics, 39(4):620–645, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=39&
issue=4&spage=620.

Kravanja:1999:LCZ

[2356] P. Kravanja, T. Sakurai, and M. Van Barel. On locating clusters of zeros of analytic functions. BIT Numerical Mathematics, 39(4):646–682, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=646.

Lie:1999:DSM

[2357] Knut-Andreas Lie. A dimensional splitting method for quasilinear hyperbolic equations with variable coefficients. BIT Numerical Mathematics, 39(4):683–700, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=683.

Lyness:1999:CAL

[2358] J. N. Lyness and S. Joe. A constructive approach to lattice rule canonical forms. BIT Numerical Mathematics, 39(4):701–715, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=701.

Marcinkowski:1999:MEM

[2359] Leszek Marcinkowski. The mortar element method with locally nonconforming elements. BIT Numerical Mathematics, 39(4):716–739, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=716.

Pan:1999:BSV

[2360] Ching-Tsuan Pan and Ping Tak Peter Tang. Bounds on singular values revealed by QR factorizations. BIT Numerical Mathematics, 39(4):740–756, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=740.

Park:1999:LRA

[2361] H. Park, L. Zhang, and J. B.

Rosen. Low rank approximation of a Hankel matrix by structured total least norm. BIT Numerical Mathematics, 39(4):757–779, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=757.

Celorrio:1999:EMF

[2362] Ricardo Celorrio and Francisco-Javier Sayas. The Euler–Maclaurin formula in presence of a logarithmic singularity. BIT Numerical Mathematics, 39(4):780–785, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=780.

Anonymous:1999:A

[2363] Anonymous. Acknowledgements. BIT Numerical Mathematics, 39(4):786, December 1999. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=39&issue=4&spage=786.

Anonymous:1999:AIV

[2364] Anonymous. Author index volume 39, 1999 BIT, numerical mathematics. BIT Numerical Mathematics, 39(4):787–790, December 1999. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=39&
issue=4&spage=787.

Bakaev:2000:ADP

[2365] Nikolai Yu. Bakaev. Analysis of discretizations of parabolic problems in pairs of spaces. BIT Numerical Mathematics, 40(1):1–23, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=1.

Chan:2000:FCO

[2366] Raymond H. Chan, Wing-Fai Ng, and Hai-Wei Sun. Fast construction of optimal circulant preconditioners for matrices from the fast dense matrix method. BIT Numerical Mathematics, 40(1):24– 40, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=24.

Engo:2000:CGI

[2367] Kenth Engö. On the construction of geometric integrators in the RKMK class. BIT Numerical Mathematics, 40(1):41– 61, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=41.

Favati:2000:RGE

[2368] Paola Favati, Mauro Leoncini, and Angeles Martinez. On the robustness of Gaussian elimination with partial pivoting. BIT Numerical Mathematics, 40(1):62–73, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=62.

Fraysse:2000:SBE

[2369] V. Frayssé, S. Gratton, and V. Toumazou. Structured backward error and condition number for linear systems of the type A^{∗}Ax = b. BIT Numerical Mathematics, 40(1):74–83, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=74.

Gander:2000:AQR

[2370] Walter Gander and Walter Gautschi.

Adaptive quadrature — revisited. BIT Numerical Mathematics, 40(1):84–101, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=84.

Novruzi:2000:NMS

[2371] A. Novruzi and J. R. Roche. Newton’s method in shape optimization: A three-dimensional case. BIT Numerical Mathematics, 40(1):102–120, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=102.

Owren:2000:NIL

[2372] B. Owren and B. Welfert. The Newton iteration on Lie groups. BIT Numerical Mathematics, 40(1):121– 145, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=121.

Reid:2000:ISL

[2373] J. K. Reid. Implicit scaling of linear least squares problems. BIT Numerical Mathematics, 40(1):146– 157, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=146.

Schenk:2000:ESF

[2374] O. Schenk, K. Gärtner, and W. Fichtner. Efficient sparse LU factorization with left–right looking strategy on shared memory multiprocessors. BIT Numerical Mathematics, 40(1):158– 176, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=1&spage=158.

Abdulle:2000:REC

[2375] Assyr Abdulle. On roots and error constants of optimal stability polynomials. BIT Numerical Mathematics, 40(1):177–182, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=40&
issue=1&spage=177.

Spellucci:2000:QDM

[2376] P. Spellucci and W. M. Hartmann. A QR decomposition for matrix pencils. BIT Numerical Mathematics, 40(1):183–189, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=40&
issue=1&spage=183.

Anonymous:2000:AIB

[2377] Anonymous. Author index: BIT Numerical Mathematics volumes 30– 39 (1990–1999). BIT Numerical Mathematics, 40(1):191–214, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=40&
issue=1&spage=191.

Froberg:2000:BNC

[2378] Carl-Erik Fröberg. BIT — A Nordic computer related journal. BIT Numerical Mathematics, 40(1):i–vi, March 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=1&spage=1.

Boumenir:2000:HAE

[2379] Amin Boumenir. Higher approximation of eigenvalues by the sampling method. BIT Numerical Mathematics, 40(2): 215–225, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=2&spage=215.

Burrage:2000:SMI

[2380] K. Burrage and S. Piskarev. Stochastic methods for ill-posed problems. BIT Numerical Mathematics, 40(2): 226–240, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=2&spage=226.

Dutt:2000:SDC

[2381] Alok Dutt, Leslie Greengard, and Vladimir Rokhlin. Spectral deferred correction methods for ordinary differential equations. BIT Numerical Mathematics, 40(2):241– 266, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=2&spage=241.

Hyvonen:2000:RBK

[2382] S. Hyvönen and O. Nevanlinna. Robust bounds for Krylov methods. BIT Numerical Mathematics, 40(2):267– 290, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=2&spage=267.

Jansson:2000:CCE

[2383] Christian Jansson. Convex-concave extensions. BIT Numerical Mathematics, 40(2):291–313, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=2&spage=291.

Lamba:2000:DSA

[2384] H. Lamba. Dynamical systems and adaptive timestepping in ODE solvers. BIT Numerical Mathematics, 40(2): 314–335, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=2&spage=314.

Lin:2000:FPI

[2385] Fu-Rong Lin and Michael K. Ng. Fast preconditioned iterative methods for convolution-type integral equations. BIT Numerical Mathematics, 40(2): 336–350, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=2&spage=336.

Olsson:2000:ARK

[2386] Hans Olsson and Gustaf Söderlind. The approximate Runge–Kutta computational process. BIT Numerical Mathematics, 40(2):351–373, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=40&
                                        issue=2&spage=351.

Takama:2000:AOC

[2387] Norio Takama, Yoshiaki Muroya, and Emiko Ishiwata. On the attainable order of collocation methods for delay differential equations with proportional delay. BIT Numerical Mathematics, 40(2):374–394, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=2&spage=374.

Dopico:2000:NTS

[2388] Froilán M. Dopico. A note on sin Θ theorems for singular subspace variations. BIT Numerical Mathematics, 40(2): 395–403, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=2&spage=395.

Higham:2000:SSM

[2389] Desmond J. Higham. A-stability and stochastic mean-square stability. BIT Numerical Mathematics, 40(2):404– 409, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=2&spage=404.

Tsatsomeros:2000:RTM

[2390] Michael J. Tsatsomeros and Lei Li.

A recursive test for P -matrices. BIT Numerical Mathematics, 40(2):410– 414, June 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=2&spage=410.

Beyn:2000:RKD

[2391] W.-J. Beyn and J. Schropp. Rungekutta discretizations of singularly perturbed gradient equations. BIT Numerical Mathematics, 40(3):415–433, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=3&spage=415.

Blanes:2000:IHO

[2392] S. Blanes, F. Casas, and J. Ros. Improved high order integrators based on the Magnus expansion. BIT Numerical Mathematics, 40(3):434–450, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=3&spage=434.

Brugnano:2000:ATM

[2393] L. Brugnano, K. Burrage, and P. M. Burrage. Adams-type methods for the numerical solution of stochastic ordinary differential equations. BIT Numerical Mathematics, 40(3):451–470, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=3&spage=451.

Fassbender:2000:EAS

[2394] Heike Fassbender. Error analysis of the symplectic Lánczos method for the symplectic eigenvalue problem. BIT Numerical Mathematics, 40(3):471–496, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=3&spage=471.

Frank:2000:DEB

[2395] J. E. Frank and P. J. Van Der Houwen. Diagonalizable extended backward differentiation formulas. BIT Numerical Mathematics, 40(3):497–512, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=3&spage=497.

Gulliksson:2000:PIR

[2396] M. E. Gulliksson, P.-Å. Å. Wedin, and Yimin Wei. Perturbation identities for regularized Tikhonov inverses and weighted pseudoinverses. BIT Numerical Mathematics, 40(3):513–524, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=3&spage=513.

Ipsen:2000:EBG

[2397] I. C. F. Ipsen. Expressions and bounds for the GMRES residual. BIT Numerical Mathematics, 40(3):524–535, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=3&spage=524.

Lin:2000:ICF

[2398] Chih-Jen Lin and Romesh Saigal. An incomplete Cholesky factorization for dense symmetric positive definite matrices. BIT Numerical Mathematics, 40(3):536–558, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=3&spage=536.

Reich:2000:FVM

[2399] Sebastian Reich. Finite volume methods for multi-symplectic PDES. BIT Numerical Mathematics, 40(3):559– 582, September 2000. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=3&spage=559.

Yun:2000:BPN

[2400] Jae Heon Yun. Block ILU preconditioners for a nonsymmetric blocktridiagonal M-matrix. BIT Numerical Mathematics, 40(3):583–605, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=3&spage=583.

Hu:2000:RBW

[2401] Guang-Da Hu and Guang-Di Hu. A relation between the weighted logarithmic norm of a matrix and the Lyapunov equation. BIT Numerical Mathematics, 40(3):606–610, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=3&spage=606.

Calvo:2000:RKM

[2402] M. Calvo, S. González-Pinto, and J. I. Montijano. Runge–Kutta methods for the numerical solution of stiff semi-linear systems. BIT Numerical Mathematics, 40(4):611–639, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=611.

Davies:2000:NSG

[2403] Philip I. Davies and Nicholas J. Higham. Numerically stable generation of correlation matrices and their factors. BIT Numerical Mathematics, 40(4):640–651, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=4&spage=640.

Faltinsen:2000:BEA

[2404] S. Faltinsen. Backward error analysis for Lie-group methods. BIT Numerical Mathematics, 40(4):652–670, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=4&spage=652.

Golub:2000:III

[2405] Gene H. Golub and Qiang Ye. Inexact inverse iteration for generalized eigenvalue problems. BIT Numerical Mathematics, 40(4):671–684, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=4&spage=671.

Gonzalez-Pinto:2000:SAF

[2406] S. González-Pinto, J. I. Montijano, and S. Pérez-Rodríguez. On the starting algorithms for fully implicit Runge–Kutta methods. BIT Numerical Mathematics, 40(4):685–714, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=685.

Gwinner:2000:LSS

[2407] J. Gwinner and A. Rott. Least squares solutions of bivariate surface fitting problems using tensor product splines. BIT Numerical Mathematics, 40(4):715–725, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=715.

Hairer:2000:SPM

[2408] E. Hairer. Symmetric projection methods for differential equations on manifolds. BIT Numerical Mathematics, 40(4):726–734, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=726.

Jahnke:2000:EBE

[2409] Tobias Jahnke and Christian Lubich. Error bounds for exponential operator splittings. BIT Numerical Mathematics, 40(4):735–744, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=735.

Li:2000:FOR

[2410] Zhi-Ping Li. Finite order rankone convex envelopes and computation of microstructures with laminates in laminates. BIT Numerical Mathematics, 40(4):745–761, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=4&spage=745.

Lu:2000:FOC

[2411] Ya Yan Lu. Fourth order conservative exponential methods for linear evolution equations. BIT Numerical Mathematics, 40(4):762–774, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=40&issue=4&spage=762.

Tracogna:2000:TSR

[2412] S. Tracogna and B. Welfert. Twostep Runge–Kutta: Theory and practice. BIT Numerical Mathematics, 40(4):775–799, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=40&
issue=4&spage=775.

Embree:2000:MWC

[2413] M. Embree. The most widely cited papers in BIT. BIT Numerical Mathematics, 40(4):800–801, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=800.

Anonymous:2000:A

[2414] Anonymous. Acknowledgements. BIT Numerical Mathematics, 40(4):802, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=40&issue=4&spage=802.

Anonymous:2000:AIV

[2415] Anonymous. Author index: Volume 40, 2000, BIT, Numerical Mathematics. BIT Numerical Mathematics, 40(4):803–806, December 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=40&
                                        issue=4&spage=803.

Arnold:2001:PDI

[2416] Martin Arnold and Michael Günther. Preconditioned dynamic iteration for coupled differential-algebraic systems. BIT Numerical Mathematics, 41(1):1– 25, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=1&spage=1.

Ashby:2001:RIP

[2417] S. F. Ashby, M. J. Holst, T. A. Manteuffel, and P. E. Saylor. The role of the inner product in stopping criteria for conjugate gradient iterations.

BIT Numerical Mathematics, 41(1):26– 52, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=1&spage=26.

Bai:2001:CIO

[2418] Zhong-Zhi Bai, Iain S. Duff, and Andrew J. Wathen. A class of incomplete orthogonal factorization methods. I: Methods and theories. BIT Numerical Mathematics, 41(1):53–70, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=1&spage=53.

Golub:2001:SLM

[2419] Gene H. Golub, X. Wu, and Jin-Yun Yuan. SOR-like methods for augmented systems. BIT Numerical Mathematics, 41(1):71–85, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=1&spage=71.

Gutknecht:2001:RST

[2420] Martin H. Gutknecht and Miroslav Rozložník. Residual smoothing techniques: Do they improve the limiting accuracy of iterative solvers? BIT Numerical Mathematics, 41(1):86–114, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=1&spage=86.

Li:2001:CMP

[2421] Chi-Kwong Li and Roy Mathias. Construction of matrices with prescribed singular values and eigenvalues. BIT Numerical Mathematics, 41(1):115– 126, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=1&spage=115.

Manni:2001:SPH

[2422] Carla Manni. On shape preserving C^{2} Hermite interpolation. BIT Numerical Mathematics, 41(1):127–148, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=1&spage=127.

Mastronardi:2001:FSR

[2423] N. Mastronardi, S. Chandrasekaran, and S. Van Huffel. Fast and stable reduction of diagonal plus semi-separable matrices to tridiagonal and bidiagonal form. BIT Numerical Mathematics, 41(1):149–157, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=1&spage=149.

Meinardus:2001:BAF

[2424] Günther Meinardus and Guido Walz. Best approximation by free knot splines. BIT Numerical Mathematics, 41(1):158–178, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=1&spage=158.

Sun:2001:CMM

[2425] Hai-Wei Sun, Xiao-Qing Jin, and Qian-Shun Chang. Convergence of the multigrid method for ill-conditioned block Toeplitz systems. BIT Numerical Mathematics, 41(1):179–190, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=1&spage=179.

Zubik-Kowal:2001:SNS

[2426] Barbara Zubik-Kowal. Stability in the numerical solution of linear parabolic equations with a delay term. BIT Numerical Mathematics, 41(1):191–206, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=1&spage=191.

Marthinsen:2001:NCC

[2427] A. Marthinsen and B. Owren. A note on the construction of Crouch– Grossman methods. BIT Numerical Mathematics, 41(1):207–214, January 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=1&spage=207.

Bakaev:2001:MNR

[2428] Nikolai Yu. Bakaev. Maximum norm resolvent estimates for elliptic finite element operators. BIT Numerical Mathematics, 41(2):215–239, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=2&spage=215.

Butcher:2001:GLM

[2429] J. C. Butcher. General linear methods for stiff differential equations. BIT Numerical Mathematics, 41(2):240– 264, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=2&spage=240.

Cullum:2001:PAN

[2430] J. K. Cullum and A. E. Ruehli. Pseudospectra analysis, nonlinear eigenvalue problems, and studying systems with time delays. BIT Numerical Mathematics, 41(2):265–281, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=2&spage=265.

Ford:2001:WBP

[2431] Judith Ford and Ke Chen. Waveletbased preconditioners for dense matrices with non-smooth local features.

BIT Numerical Mathematics, 41(2): 282–307, March 2001. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=2&spage=282.

Gulliksson:2001:ISF

[2432] M. Gulliksson, I. Söderkvist, and G. A. Watson. Implicit surface fitting using directional constraints. BIT Numerical Mathematics, 41(2):308– 321, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=2&spage=308.

intHout:2001:CRK

[2433] K. J. in ’t Hout. Convergence of Runge–Kutta methods for delay differential equations. BIT Numerical Mathematics, 41(2):322–344, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=2&spage=322.

Liang:2001:FVM

[2434] Shengde Liang, Xiuling Ma, and Aihui Zhou. Finite volume methods for eigenvalue problems. BIT Numerical Mathematics, 41(2):345–363, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=2&spage=345.

Pryce:2001:SSA

[2435] J. D. Pryce. A simple structural analysis method for DAEs. BIT Numerical Mathematics, 41(2):364– 394, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=2&spage=364.

Zanna:2001:ASL

[2436] Antonella Zanna, Kenth Engø, and Hans Z. Munthe-Kaas. Adjoint and selfadjoint Lie-group methods. BIT Numerical Mathematics, 41(2):395– 421, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=2&spage=395.

Jay:2001:NOC

[2437] L. O. Jay. A note on Qorder of convergence. BIT Numerical Mathematics, 41(2):422–429, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=2&spage=422.

Malyshev:2001:OBP

[2438] A. N. Malyshev. Optimal backward perturbation bounds for the LSS problem. BIT Numerical Mathematics, 41(2):430–432, March 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=41&
issue=2&spage=430.

Bertaccini:2001:CRB

[2439] Daniele Bertaccini and Michael K.

Ng. The convergence rate of block preconditioned systems arising from LMF-based ODE codes. BIT Numerical Mathematics, 41(3):433–450, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=433.

Costabile:2001:ERR

[2440] F. A. Costabile and F. Dell’accio.

Expansion over a rectangle of real functions in Bernoulli polynomials and applications. BIT Numerical Mathematics, 41(3):451–464, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=451.

Dobrowolski:2001:ADV

[2441] M. Dobrowolski and G. Stoyan. Algebraic and discrete Velte decompositions. BIT Numerical Mathematics, 41(3):465–479, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=465.

Douglas:2001:LCE

[2442] Jim Douglas, Jr. and Chieh-Sen Huang. A locally conservative Eulerian–Lagrangian finite difference method for a parabolic equation. BIT Numerical Mathematics, 41(3):480– 489, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=480.

Gautschi:2001:CHT

[2443] Walter Gautschi and Jörg Waldvogel. Computing the Hilbert transform of the generalized Laguerre and Hermite weight functions. BIT Numerical Mathematics, 41(3):490– 503, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=490.

Gunther:2001:MPR

[2444] M. Günther, A. Kværnø, and P. Rentrop. Multirate partitioned Runge–Kutta methods. BIT Numerical Mathematics, 41(3):504–514, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=504.

Langlois:2001:ALC

[2445] Philippe Langlois. Automatic linear correction of rounding errors. BIT Numerical Mathematics, 41(3):515– 539, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=515.

Rump:2001:RPS

[2446] Siegfried M. Rump. Rigorous and portable standard functions. BIT Numerical Mathematics, 41(3):540– 562, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=540.

Tasche:2001:WAC

[2447] Manfred Tasche and Hansmartin Zeuner. Worst and average case roundoff error analysis for FFT. BIT Numerical Mathematics, 41(3):563– 581, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=563.

Timonov:2001:FPS

[2448] Alexandre Timonov. Factorised preconditionings of successive approximations in finite precision. BIT Numerical Mathematics, 41(3):582– 598, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=582.

Tuomela:2001:NSI

[2449] J. Tuomela and T. Arponen. On the numerical solution of involutive ordinary differential systems II: Higher order methods. BIT Numerical Mathematics, 41(3):599–628, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=3&spage=599.

Engo:2001:BF

[2450] Kenth Engø. On the BCHformula in SO(3). BIT Numerical Mathematics, 41(3):629–632, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=3&spage=629.

Mead:2001:SPS

[2451] J. L. Mead, R. A. Renaut, and B. D. Welfert. Stability of a pivoting strategy for parallel Gaussian elimination. BIT Numerical Mathematics, 41(3): 633–639, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=3&spage=633.

Pena:2001:NPP

[2452] J. M. Peña. A note on a paper by P. Amodio and F. Mazzia. BIT Numerical Mathematics, 41(3):640– 643, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=3&spage=640. See

[2342].

Burrage:2001:ACS

[2453] K. Burrage, G. Hertono, Z. Jackiewicz, and B. D. Welfert. Acceleration of convergence of static and dynamic iterations. BIT Numerical Mathematics, 41(4):645–655, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=645.

Butcher:2001:REE

[2454] J. C. Butcher and Z. Jackiewicz.

A reliable error estimation for diagonally implicit multistage integration methods. BIT Numerical Mathematics, 41(4):656–665, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=656.

Capobianco:2001:AWH

[2455] M. R. Capobianco, G. Criscuolo, and R. Giova. Approximation of the weighted Hilbert transform on the real line by an interpolatory process. BIT Numerical Mathematics, 41(4):666–682, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=4&spage=666.

Casado:2001:HRC

[2456] L. G. Casado, I. García, and T. Csendes. A heuristic rejection criterion in interval global optimization algorithms. BIT Numerical Mathematics, 41(4):683–692, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=4&spage=683.

Gilbert:2001:CRC

[2457] J. R. Gilbert, X. S. Li, E. G. Ng, and B. W. Peyton. Computing row and column counts for sparse QR and LU factorization. BIT Numerical Mathematics, 41(4):693–710, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=4&spage=693.

Grune:2001:PAR

[2458] L. Grüne and P. E. Kloeden. Pathwise approximation of random ordinary differential equations. BIT Numerical Mathematics, 41(4):711–721, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=4&spage=711.

Karatsuba:2001:FCS

[2459] Ekatherina A. Karatsuba. Fast computation of ζ(3) and of some special integrals using the Ramanujan formula and polylogarithms. BIT Numerical Mathematics, 41(4):722–730, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=722.

Lang:2001:RAT

[2460] J. Lang and J. Verwer. ROS3P — an accurate third-order Rosenbrock solver designed for parabolic problems. BIT Numerical Mathematics, 41(4):731–738, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=41&
issue=4&spage=731.

Malyshev:2001:COB

[2461] A. Malyshev and M. Sadkane. Computation of optimal backward perturbation bounds for large sparse linear least squares problems. BIT Numerical Mathematics, 41(4):739–747, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=739.

Melnik:2001:OCF

[2462] K. N. Melnik and R. V. N. Melnik.

Optimal cubature formulae and recovery of fast-oscillating functions from an interpolational class. BIT Numerical Mathematics, 41(4):748–775, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=748.

Muir:2001:MIR

[2463] P. H. Muir and M. Adams. Monoimplicit Runge–Kutta–Nyström methods with application to boundary value ordinary differential equations. BIT Numerical Mathematics, 41(4):776–799, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=41&
issue=4&spage=776.

Schulze:2001:TTC

[2464] Jürgen Schulze. Towards a tighter coupling of bottom-up and top-down sparse matrix ordering methods. BIT Numerical Mathematics, 41(4):800– 841, September 2001. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=800.

Politi:2001:FEM

[2465] Tiziano Politi. A formula for the exponential matrix of a real skewsymmetric matrix of order 4. BIT Numerical Mathematics, 41(4):842–845, September 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=4&spage=842.

Baryamureeba:2001:IEW

[2466] Venansius Baryamureeba. The impact of equal weighting of low- and highconfidence observations on robust linear regression computations. BIT Numerical Mathematics, 41(5):847–855, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=847.

Beckermann:2001:SAE

[2467] B. Beckermann and A. B. J. Kuijlaars. On the sharpness of an asymptotic error estimate for conjugate gradients. BIT Numerical Mathematics, 41(5):856–867, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=856.

Berrut:2001:LRP

[2468] Jean-Paul Berrut and Richard Baltensperger. The linear rational pseudospectral method for boundary value problems. BIT Numerical Mathematics, 41(5):868–879, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=868.

deBoor:2001:WIB

[2469] Carl de Boor. What is the inverse of a basis? BIT Numerical Mathematics, 41(5):880–890, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=880.

Brunner:2001:SCM

[2470] H. Brunner, A. Pedas, and G. Vainikko. A spline collocation method for linear Volterra integro-differential equations with weakly singular kernels. BIT Numerical Mathematics, 41(5):891–900, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=891.

Bucker:2001:UHN

[2471] H. Martin Bücker and Manfred Sauren. On using Hölder norms in the quasiminimal residual approach. BIT Numerical Mathematics, 41(5):901–911, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=901.

delBuono:2001:RKT

[2472] N. del Buono and L. Lopez. Runge– Kutta type methods based on geodesics for systems of ODEs on the Stiefel manifold. BIT Numerical Mathematics, 41(5):912–923, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=41&
issue=5&spage=912.

Dekker:2001:PGD

[2473] K. Dekker. Partitioned-GMRES in domain decomposition with approximate subdomain solution. BIT Numerical Mathematics, 41(5):924–935, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=924.

Elmroth:2001:FSR

[2474] Erik Elmroth and Fred G. Gustavson. A faster and simpler recursive algorithm for the LAPACK routine DGELS. BIT Numerical Mathematics, 41(5):936–949, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=936.

Fischer:2001:TRI

[2475] Bernd Fischer, Martin Ludwig, and Andreas Meister. The thermoregulation of infants: Modeling and numerical simulation. BIT Numerical Mathematics, 41(5):950–966, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=950.

Gansterer:2001:LCD

[2476] W. N. Gansterer, J. Schneid, and C. W. Ueberhuber. A low-complexity divide-and-conquer method for computing eigenvalues and eigenvectors of symmetric band matrices. BIT Numerical Mathematics, 41(5):967–976, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=967.

Golub:2001:SAE

[2477] G. H. Golub and H. Melbö. A stochastic approach to error estimates for iterative linear solvers: Part 1. BIT Numerical Mathematics, 41(5):977–985, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=977.

Gustafsson:2001:DCM

[2478] Bertil Gustafsson and Wendy Kress. Deferred correction methods for initial value problems. BIT Numerical Mathematics, 41(5):986–995, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=986.

Hairer:2001:GIO

[2479] E. Hairer. Geometric integration of ordinary differential equations on manifolds. BIT Numerical Mathematics, 41(5):996–1007, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=996.

Hanke:2001:LBM

[2480] Martin Hanke. On Lánczos based methods for the regularization of discrete ill-posed problems. BIT Numerical Mathematics, 41(5):1008–1018, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1008.

Hasegawa:2001:EAC

[2481] Takemitsu Hasegawa. Error analysis of Clenshaw’s algorithm for evaluating derivatives of a polynomial. BIT Numerical Mathematics, 41(5):1019–1028, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1019.

Hamarik:2001:MER

[2482] U. Hämarik and U. Tautenhahn. On the monotone error rule for parameter choice in iterative and continuous regularization methods. BIT Numerical Mathematics, 41(5):1029–1038, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1029.

Kirlinger:2001:CBD

[2483] Gabriela Kirlinger. On the convergence of backward differentiation formulas for stiff initial value problems. BIT Numerical Mathematics, 41(5):1039–1048, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1039.

Kitagawa:2001:RUQ

[2484] T. Kitagawa, S. Nakata, and Y. Hosoda. Regularization using QR factorization and the estimation of the optimal parameter. BIT Numerical Mathematics, 41(5):1049–1058, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1049.

Kojdecki:2001:ESC

[2485] Marek Andrzej Kojdecki. Examples of saturated convergence rates for Tikhonov regularization. BIT Numerical Mathematics, 41(5):1059–1068, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1059.

Meerbergen:2001:LMS

[2486] Karl Meerbergen. The Lánczos method with semi-definite inner product. BIT Numerical Mathematics, 41(5):1069–1078, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1069.

Ng:2001:CPI

[2487] Michael K. Ng and Daniel Potts.

Circulant preconditioners for indefinite Toeplitz systems. BIT Numerical Mathematics, 41(5):1079–1088, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1079.

Orel:2001:EMM

[2488] Bojan Orel. Extrapolated Magnus methods. BIT Numerical Mathematics, 41(5):1089–1100, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=41&issue=5&spage=1089.

Reginska:2001:AWS

[2489] Teresa Regínska. Application of wavelet shrinkage to solving the sideways heat equation. BIT Numerical Mathematics, 41(5):1101–1110, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=1101.

Voss:2001:VIL

[2490] Heinrich Voss. A variant of the inverted Lánczos method. BIT Numerical Mathematics, 41(5):1111–1120, ???? 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:2001:A

[2491] Anonymous. Acknowledgements. BIT Numerical Mathematics, 41(5):1121, ???? 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:2001:IV

[2492] Anonymous. Index: Volume 41. BIT Numerical Mathematics, 41(5):1123– 1127, ???? 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Soderlind:2001:P

[2493] Gustaf Söderlind. Preface. BIT Numerical Mathematics, 41(5):??, December 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=41&issue=5&spage=??.

Benner:2002:PAE

[2494] Peter Benner, Volker Mehrmann, and Hongguo Xu. Perturbation analysis for the eigenvalue problem of a formal product of matrices. BIT Numerical Mathematics, 42(1):1–43, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=1.

Calvetti:2002:GCD

[2495] D. Calvetti, B. Lewis, and L. Reichel. GMRES, L-curves, and discrete ill-posed problems. BIT Numerical Mathematics, 42(1):44–65, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=44.

Chang:2002:SFG

[2496] Xiao-Wen Chang. Some features of Gaussian elimination with rook pivoting. BIT Numerical Mathematics, 42(1):66–83, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=66.

Dopico:2002:PTS

[2497] Froilán M. Dopico and Julio Moro.

Perturbation theory for simultaneous bases of singular subspaces. BIT Numerical Mathematics, 42(1):84–109, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=84.

Gautschi:2002:CBA

[2498] Walter Gautschi. Computation of Bessel and Airy functions and of related Gaussian quadrature formulae. BIT Numerical Mathematics, 42 (1):110–118, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=1&spage=110.

Lera:2002:GMA

[2499] D. Lera and Ya. D. Sergeyev. Global minimization algorithms for Hölder functions. BIT Numerical Mathematics, 42(1):119–133, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=42&
                                        issue=1&spage=119.

Lotstedt:2002:ISH

[2500] Per Lötstedt, Stefan Söderberg, Alison Ramage, and Lina Hemmingsson-Frändén. Implicit solution of hyperbolic equations with spacetime adaptivity. BIT Numerical Mathematics, 42(1):134–158, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=42&
                                        issue=1&spage=134.

Simoncini:2002:IRQ

[2501] Valeria Simoncini and Lars Eldén. Inexact Rayleigh quotient-type methods for eigenvalue computations. BIT Numerical Mathematics, 42(1):159– 182, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=159.

Venter:2002:DAI

[2502] A. Venter and D. P. Laurie. A doubly adaptive integration algorithm using stratified rules. BIT Numerical Mathematics, 42(1):183–193, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=183.

Viswanath:2002:HMT

[2503] Divakar Viswanath. How many timesteps for a cycle? Analysis of the Wisdom–Holman algorithm. BIT Numerical Mathematics, 42(1):194– 205, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=1&spage=194.

Wang:2002:UCT

[2504] Xinghua Wang, Chong Li, and Ming-Jun Lai. A unified convergence theory for Newton-type methods for zeros of nonlinear operators in Banach spaces. BIT Numerical Mathematics, 42(1):206–213, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=1&spage=206.

Yamamura:2002:FAS

[2505] Kiyotaka Yamamura and Shigeru Tanaka. Finding all solutions of systems of nonlinear equations using the dual simplex method. BIT Numerical Mathematics, 42(1):214– 230, March 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=1&spage=214.

Bacchelli:2002:MP

[2506] S. Bacchelli, M. Cotronei, and T. Sauer. Multifilters with and without prefilters. BIT Numerical Mathematics, 42(2):231–261, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=42&
                                        issue=2&spage=231.

Blanes:2002:HOO

[2507] S. Blanes, F. Casas, and J. Ros. High order optimized geometric integrators for linear differential equations. BIT Numerical Mathematics, 42(2):262– 284, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=2&spage=262.

Diele:2002:CMI

[2508] F. Diele and I. Sgura. The Cayley method and the inverse eigenvalue problem for Toeplitz matrices. BIT Numerical Mathematics, 42(2): 285–299, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=2&spage=285.

Duff:2002:INE

[2509] Iain S. Duff and Christof Vömel. Incremental norm estimation for dense and sparse matrices. BIT Numerical Mathematics, 42(2):300–322, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=2&spage=300.

Elden:2002:SQC

[2510] Lars Eldèn. Solving quadratically constrained least squares problems using a differential-geometric approach. BIT Numerical Mathematics, 42(2): 323–335, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=2&spage=323.

Han:2002:NAS

[2511] Tian Min Han and Yuhuan Han. A new approach of solving implicit equations arising from Adams–Moulton methods. BIT Numerical Mathematics, 42(2): 336–350, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=2&spage=336.

Hansbo:2002:SSN

[2512] Anita Hansbo. Strong stability and non-smooth data error estimates for discretizations of linear parabolic problems. BIT Numerical Mathematics, 42(2):351–379, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=2&spage=351.

Huang:2002:NSG

[2513] Chengming Huang, Shoufu Li, Hongyuan Fu, and Guangnan Chen. Nonlinear stability of general linear methods for delay differential equations. BIT Numerical Mathematics, 42(2): 380–392, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=2&spage=380.

Huhtanen:2002:GDO

[2514] Marko Huhtanen and Rasmus Munk Larsen. On generating discrete orthogonal bivariate polynomials. BIT Numerical Mathematics, 42(2):393– 407, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=2&spage=393.

Nipp:2002:NID

[2515] Kaspar Nipp. Numerical integration of differential algebraic systems and invariant manifolds. BIT Numerical Mathematics, 42(2):408– 439, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=2&spage=408.

Notaris:2002:PWI

[2516] Sotirios E. Notaris. Positivity of the weights of interpolatory quadrature formulae with Jacobi abscissae. BIT Numerical Mathematics, 42(2): 440–446, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=2&spage=440.

Tupper:2002:AMR

[2517] P. F. Tupper. Adaptive model reduction for chemical kinetics. BIT Numerical Mathematics, 42(2):447– 465, June 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=2&spage=447.

Batra:2002:SPE

[2518] P. Batra. Simultaneous point estimates for Newton’s method. BIT Numerical Mathematics, 42(3):467–476, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=3&spage=467.

Butcher:2002:NAA

[2519] J. C. Butcher and T. M. H. Chan. A new approach to the algebraic structures for integration methods. BIT Numerical Mathematics, 42(3):477–489, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=3&spage=477.

Diethelm:2002:NSB

[2520] K. Diethelm and N. J. Ford. Numerical solution of the Bagley– Torvik equation. BIT Numerical Mathematics, 42(3):490–507, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=3&spage=490.

Ding:2002:CAS

[2521] Xiaohua Ding and Mingzhu Liu. Convergence aspects of step-parallel iteration of Runge–Kutta methods for delay differential equations. BIT Numerical Mathematics, 42(3):508–518, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=42&issue=3&spage=508.

Ezquerro:2002:ANM

[2522] J. A. Ezquerro and M. A. Hernández.

On an application of Newton’s method to nonlinear operators with wconditioned second derivative. BIT Numerical Mathematics, 42(3):519–530, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=519.

Fierro:2002:TVS

[2523] Ricardo D. Fierro and Per Christian Hansen. Truncated VSV solutions to symmetric rank-deficient problems. BIT Numerical Mathematics, 42(3):531–540, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=42&
issue=3&spage=531.

Huckle:2002:MMM

[2524] Thomas Huckle and Jochen Staudacher. Matrix multilevel methods and preconditioning. BIT Numerical Mathematics, 42(3):541–560, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=541.

Iserles:2002:GED

[2525] Arieh Iserles. On the global error of discretization methods for highly-oscillatory ordinary differential equations. BIT Numerical Mathematics, 42(3):561–599, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=561.

Smoktunowicz:2002:BSC

[2526] Alicja Smoktunowicz. Backward stability of Clenshaw’s algorithm. BIT Numerical Mathematics, 42(3):600–610, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=600.

Song:2002:MCI

[2527] Yongzhong Song. Monotone convergence of iterative methods for singular linear systems. BIT Numerical Mathematics, 42(3):611–624, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=611.

Tian:2002:TSS

[2528] T. H. Tian and K. Burrage. Twostage stochastic Runge–Kutta methods for stochastic differential equations. BIT Numerical Mathematics, 42(3):625–643, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=42&
issue=3&spage=625.

Tornberg:2002:MDQ

[2529] Anna-Karin Tornberg. Multi-dimensional quadrature of singular and discontinuous functions. BIT Numerical Mathematics, 42(3):644–669, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=644.

Zagar:2002:CSI

[2530] Emil Žagar. On G^{2} continuous spline interpolation of curves in R^{d}. BIT Numerical Mathematics, 42(3):670–688, September 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=3&spage=670.

Calvo:2002:TDP

[2531] M. Calvo and R. D. Grigorieff. Time discretization of parabolic problems with the variable 3-step BDF. BIT Numerical Mathematics, 42(4):689–701, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=689.

Christara:2002:FFT

[2532] Christina C. Christara and Kit Sun Ng. Fast Fourier Transform solvers and preconditioners for quadratic spline collocation. BIT Numerical Mathematics, 42(4):702–739, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=702.

Dryja:2002:DNA

[2533] Maksymilian Dryja. The Dirichlet– Neumann algorithm for mortar saddle point problems. BIT Numerical Mathematics, 42(4):740–752, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=740.

Estatico:2002:CFS

[2534] Claudio Estatico. A class of filtering superoptimal preconditioners for highly ill-conditioned linear systems. BIT Numerical Mathematics, 42(4):753–778, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=753.

Gander:2002:OSW

[2535] Martin J. Gander and Hongkai Zhao. Overlapping Schwarz waveform relaxation for the heat equation in n dimensions. BIT Numerical Mathematics, 42(4):779–795, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=779.

Gerdts:2002:CIS

[2536] M. Gerdts and C. Büskens. Consistent initialization of sensitivity matrices for a class of parametric DAE systems. BIT Numerical Mathematics, 42(4):796–813, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=796.

Golub:2002:STD

[2537] Gene H. Golub and Jin-Yun Yuan.

Symmetric-triangular decomposition and its applications Part I: Theorems and algorithms. BIT Numerical Mathematics, 42(4):814–822, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=814.

Higueras:2002:LNN

[2538] Inmaculada Higueras and Gustaf Söderlind. Logarithmic norms and nonlinear DAE stability. BIT Numerical Mathematics, 42(4):823–841, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=823.

Oja:2002:MPA

[2539] Peeter Oja. Monotonicity preserving adaptive interpolating splines. BIT Numerical Mathematics, 42(4):842–855, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=842.

Pinar:2002:LHE

[2540] M. Ç. Pinar. Linear Huber Mestimator under ellipsoidal data uncertainty. BIT Numerical Mathematics, 42(4):856–866, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=42&
issue=4&spage=856.

Watson:2002:IAI

[2541] G. A. Watson. Incorporating angular information into parametric models. BIT Numerical Mathematics, 42(4):867–878, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=42&
issue=4&spage=867.

Kyurkchiev:2002:OCF

[2542] N. Kyurkchiev and A. Iliev. On the R-order of convergence of a family of methods for simultaneous extraction of part of all roots of algebraic polynomials. BIT Numerical Mathematics, 42(4):879–885, December 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=42&issue=4&spage=879.

Anonymous:2002:PA

[2543] Anonymous. Publisher’s announcement. BIT Numerical Mathematics, 42(4):886, 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:2002:A

[2544] Anonymous. Acknowledgements. BIT Numerical Mathematics, 42(4):887, 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:2002:AIV

[2545] Anonymous. Author index: Volume 42, 2002 BIT, Numerical Mathematics. BIT Numerical Mathematics, 42 (4):887–891, 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bryden:2003:BAS

[2546] Alan Bryden and Desmond J. Higham.

On the boundedness of asymptotic stability regions for the stochastic theta method. BIT Numerical Mathematics, 43(1):1–6, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=1&spage=1.

Chan:2003:INL

[2547] R. H. Chan, H. L. Chung, and S.- F. Xu. The inexact Newton-like method for inverse eigenvalue problem.

BIT Numerical Mathematics, 43(1):7– 20, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=7.

Engo:2003:PRK

[2548] Kenth Engø. Partitioned Runge– Kutta methods in Lie-group setting. BIT Numerical Mathematics, 43(1):21– 39, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=21.

Frank:2003:CPS

[2549] Jason Frank and Sebastian Reich. Conservation properties of smoothed particle hydrodynamics applied to the shallow water equation. BIT Numerical Mathematics, 43(1):41–55, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=41.

Grimm:2003:GTC

[2550] V. Grimm and R. Scherer. A generalized W -transformation for constructing symplectic partitioned Runge–Kutta methods. BIT Numerical Mathematics, 43(1):57–66, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=57.

Higueras:2003:IMI

[2551] I. Higueras and T. Roldán. IRK methods for index 2 and 3 DAEs: Starting algorithms. BIT Numerical Mathematics, 43(1):67–92, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=67.

Jay:2003:SII

[2552] L. O. Jay. Solution of index 2 implicit differential-algebraic equations by Lobatto Runge–Kutta methods. BIT Numerical Mathematics, 43(1):93– 106, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=93.

Krogstad:2003:LCL

[2553] Stein Krogstad. A low complexity Lie group method on the Stiefel manifold. BIT Numerical Mathematics, 43(1):107–122, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=1&spage=107.

Larcher:2003:WDH

[2554] G. Larcher, F. Pillichshammer, and K. Scheicher. Weighted discrepancy and high-dimensional numerical integration. BIT Numerical Mathematics, 43(1):123–137, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=1&spage=123.

Lin:2003:STP

[2555] F. R. Lin, X. Q. Jin, and S. L. Lei. Strang-type preconditioners for solving linear systems from delay differential equations. BIT Numerical Mathematics, 43(1):139–152, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=139.

Schutze:2003:BFK

[2556] Torsten Schütze and Hubert Schwetlick. Bivariate free knot splines. BIT Numerical Mathematics, 43(1):153– 178, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=153.

Shardlow:2003:WCN

[2557] Tony Shardlow. Weak convergence of a numerical method for a stochastic heat equation. BIT Numerical Mathematics, 43(1):179–193, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=179.

Vermiglio:2003:SNA

[2558] R. Vermiglio and L. Torelli. A stable numerical approach for implicit non-linear neutral delay differential equations. BIT Numerical Mathematics, 43(1):195–215, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=1&spage=195.

Zhang:2003:SSM

[2559] J. Z. Zhang, Y. Xue, and K. Zhang.

A structured secant method based on a new quasi-Newton equation for nonlinear least squares problems. BIT Numerical Mathematics, 43(1):217– 229, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=217.

Ruhe:2003:Ea

[2560] Axel Ruhe. Editorial. BIT Numerical Mathematics, 43(1):v–viii, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=1&spage=??.

Benzi:2003:AIP

[2561] Michele Benzi and Daniele Bertaccini. Approximate inverse preconditioning for shifted linear systems.

BIT Numerical Mathematics, 43(2): 231–244, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=231.

Boese:2003:AEZ

[2562] Fritz G. Boese and Wolfram J. Luther. Accurate enclosure of the zero set of multivariate polynomials. BIT Numerical Mathematics, 43(2):245– 261, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=245.

Calvetti:2003:TRL

[2563] Daniela Calvetti and Lothar Reichel. Tikhonov regularization of large linear problems. BIT Numerical Mathematics, 43(2):263–283, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=263.

Ching:2003:IMQ

[2564] W. K. Ching. Iterative methods for queuing systems with batch arrivals and negative customers. BIT Numerical Mathematics, 43(2):285– 296, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=285.

Costantini:2003:GSS

[2565] Paolo Costantini and Maria Lucia Sampoli. A general scheme for shape preserving planar interpolating curves. BIT Numerical Mathematics, 43(2): 297–317, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=297.

Espelid:2003:DAQ

[2566] Terje O. Espelid. Doubly adaptive quadrature routines based on Newton–Cotes rules. BIT Numerical Mathematics, 43(2):319–337, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=319.

Franken:2003:PRK

[2567] D. Fränken and K. Ochs. Passive Runge–Kutta methods — properties, parametric representation, and order conditions. BIT Numerical Mathematics, 43(2):339–361, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=339.

Hout:2003:AEG

[2568] K. J. In ’T Hout and M. N. Spijker. Analysis of error growth via stability regions in numerical initial value problems. BIT Numerical Mathematics, 43(2):363–385, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=363.

Larsson:2003:PSH

[2569] Elisabeth Larsson and Sverker Holmgren. Parallel solution of the Helmholtz equation in a multilayer domain. BIT Numerical Mathematics, 43(2):387– 411, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=387.

Lyness:2003:NLR

[2570] J. N. Lyness and S. Joe. The number of lattice rules of specified upper class and rank. BIT Numerical Mathematics, 43(2):413– 426, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=413.

Park:2003:LDR

[2571] Haesun Park, Moongu Jeon, and J. Ben Rosen. Lower dimensional representation of text data based on centroids and least squares. BIT Numerical Mathematics, 43(2):427– 448, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=427.

Hanke:2003:NTR

[2572] Martin Hanke. A note on Tikhonov regularization of large linear problems. BIT Numerical Mathematics, 43(2): 449–451, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=449.

Kalantari:2003:EFP

[2573] B. Kalantari and Y. Jin. On extraneous fixed-points of the basic family of iteration functions. BIT Numerical Mathematics, 43(2):453– 458, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=2&spage=453.

Simoncini:2003:RFO

[2574] V. Simoncini. Restarted full orthogonalization method for shifted linear systems. BIT Numerical Mathematics, 43(2):459–466, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=2&spage=459.

Anonymous:2003:ICI

[2575] Anonymous. Introduction to the contents of issue 43:2. BIT Numerical Mathematics, 43(2):v–vi, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=2&spage=??.

Ruhe:2003:Eb

[2576] Axel Ruhe.  Editorial.  BIT

Numerical Mathematics, 43(2):iii, June 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=2&spage=??.

Akrivis:2003:MIE

[2577] Georgios Akrivis and Fotini Karakatsani. Modified implicit-explicit BDF methods for nonlinear parabolic equations. BIT Numerical Mathematics, 43(3):467–483, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=43&
                                        issue=3&spage=467.

Andrew:2003:ACM

[2578] Alan L. Andrew. Asymptotic correction of more Sturm–Liouville eigenvalue estimates. BIT Numerical Mathematics, 43(3):485–503, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=3&spage=485.

Bojanczyk:2003:ECI

[2579] Adam Bojanczyk, Nicholas J. Higham, and Harikrishna Patel. The equality constrained indefinite least squares problem: theory and algorithms. BIT Numerical Mathematics, 43(3):505– 517, September 2003. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=3&spage=505.

Buchmann:2003:SDP

[2580] F. M. Buchmann and W. P. Petersen. Solving Dirichlet problems numerically using the Feynman– Kac representation. BIT Numerical Mathematics, 43(3):519–540, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=3&spage=519.

Calvetti:2003:SGL

[2581] Daniela Calvetti and Lothar Reichel.

Symmetric Gauss–Lobatto and modified anti-Gauss rules. BIT Numerical Mathematics, 43(3):541–554, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=3&spage=541.

Gonzalez-Pinto:2003:SAR

[2582] S. González-Pinto and D. Hernández-Abreu. Strong A-acceptability for rational functions. BIT Numerical Mathematics, 43(3):555–561, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=3&spage=555.

Huhtanen:2003:TSS

[2583] Marko Huhtanen. Two splittings of a square matrix. BIT Numerical Mathematics, 43(3):563–570, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=3&spage=563.

Kloeden:2003:RKM

[2584] P. E. Kloeden and J. Schropp. Runge-Kutta methods for monotone differential and delay equations. BIT Numerical Mathematics, 43(3):571–586, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=3&spage=571.

Laszlo:2003:SOO

[2585] L. László. Second-order optimality condition for ΔH-matrices. BIT Numerical Mathematics, 43(3):587–593, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=3&spage=587.

Makazaga:2003:NRK

[2586] J. Makazaga and A. Murua. New Runge–Kutta based schemes for ODEs with cheap global error estimation. BIT Numerical Mathematics, 43(3):595–610, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=3&spage=595.

Montefusco:2003:PSM

[2587] Laura B. Montefusco and Serena Papi.

A parameter selection method for wavelet shrinkage denoising. BIT Numerical Mathematics, 43(3):611–626, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=3&spage=611.

Wei:2003:REE

[2588] Musheng Wei and Qiaohua Liu.

Roundoff error estimates of the modified Gram–Schmidt algorithm with column pivoting. BIT Numerical Mathematics, 43(3):627–645, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=3&spage=627.

Yan:2003:SPA

[2589] Yubin Yan. Smoothing A properties and approximation of time derivatives for parabolic equations: variable time steps. BIT Numerical Mathematics, 43(3):647–669, September 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=3&spage=647.

Aimi:2003:GCR

[2590] A. Aimi, L. Bassotti, and M. Diligenti. Groups of congruences and restriction matrices. BIT Numerical Mathematics, 43(4):671–693, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=4&spage=671.

Butcher:2003:CPG

[2591] J. C. Butcher and W. M. Wright. The construction of practical general linear methods. BIT Numerical Mathematics, 43(4):695–721, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=4&spage=695.

Casas:2003:CEL

[2592] F. Casas and B. Owren. Cost efficient Lie group integrators in the RKMK class. BIT Numerical Mathematics, 43(4):723–742, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=4&spage=723.

Henn:2003:LMS

[2593] Stefan Henn. A Levenberg–Marquardt scheme for nonlinear image registration. BIT Numerical Mathematics, 43(4):743–759, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=4&spage=743.

Janssen:2003:PSO

[2594] M. Janssen and P. Van Hentenryck. Precisely A(α)-stable oneleg multistep methods. BIT Numerical Mathematics, 43(4):761–774, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=4&spage=761.

Kressner:2003:BAO

[2595] D. Kressner. Block algorithms for orthogonal symplectic factorizations. BIT Numerical Mathematics, 43(4):775–790, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=4&spage=775.

Kulpa:2003:IPI

[2596] Z. Kulpa and S. Markov. On the inclusion properties of interval multiplication: a diagrammatic study. BIT Numerical Mathematics, 43(4):791–810, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=4&spage=791.

Osterby:2003:FWR

[2597] Ole Østerby. Five ways of reducing the Crank–Nicolson oscillations. BIT Numerical Mathematics, 43(4):811–822, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=4&spage=811.

Rump:2003:EB

[2598] Siegfried M. Rump and Jens-Peter M. Zemke. On eigenvector bounds. BIT Numerical Mathematics, 43(4):823– 837, November 2003. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=4&spage=823.

Anonymous:2003:E

[2599] Anonymous. Editorial. BIT Numerical Mathematics, 43(4):??, November 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=4&spage=??.

Arioli:2003:NSA

[2600] Mario Arioli and Gianmarco Manzini. Null space algorithm and spanning trees in solving Darcy’s equation. BIT Numerical Mathematics, 43(5): 839–848, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=5&spage=839.

Arnal:2003:PNT

[2601] Josep Arnal, Violeta Migallón, and José Penadés. Parallel Newton twostage multisplitting iterative methods for nonlinear systems. BIT Numerical Mathematics, 43(5):849–861, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=849.

Axelsson:2003:SAM

[2602] Owe Axelsson. A survey of algebraic multilevel iteration (AMLI) methods. BIT Numerical Mathematics, 43(5):863–879, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=5&spage=863.

Benzi:2003:OHS

[2603] Michele Benzi, Martin J. Gander, and Gene H. Golub. Optimization of the Hermitian and skew-Hermitian splitting iteration for saddlepoint problems. BIT Numerical Mathematics, 43(5):881–900, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=881.

Bertaccini:2003:BTP

[2604] Daniele Bertaccini and Michael K.

Ng. Band-Toeplitz preconditioned GMRES iterations for time-dependent PDEs. BIT Numerical Mathematics, 43(5):901–914, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=43&
issue=5&spage=901.

Chang:2003:ACC

[2605] Xiao-Wen Chang and Christopher C. Paige. An algorithm for combined code and carrier phase based GPS positioning. BIT Numerical Mathematics, 43(5):915–927, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=915.

Dax:2003:OQC

[2606] Achiya Dax. An open question on cyclic relaxation. BIT Numerical Mathematics, 43(5):929–943, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=929.

Efstathiou:2003:WRA

[2607] Evridiki Efstathiou and Martin J. Gander. Why restricted additive Schwarz converges faster than additive Schwarz. BIT Numerical Mathematics, 43(5):945–959, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=945.

Elman:2003:PTN

[2608] H. C. Elman, D. Loghin, and A. J. Wathen. Preconditioning techniques for Newton’s method for the incompressible Navier–Stokes equations. BIT Numerical Mathematics, 43(5):961–974, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=961.

Jacobsen:2003:SPL

[2609] M. Jacobsen, P. C. Hansen, and M. A. Saunders. Subspace preconditioned LSQR for discrete illposed problems. BIT Numerical Mathematics, 43(5):975–989, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=975.

Janovska:2003:GTA

[2610] D. Janovská and G. Opfer. Givens’ transformation applied to quaternion valued vectors. BIT Numerical Mathematics, 43(5):991–1002, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=43&issue=5&spage=991.

Nagy:2003:SDC

[2611] J. G. Nagy and K. M. Palmer. Steepest descent, CG, and iterative regularization of ill-posed problems. BIT Numerical Mathematics, 43(5):1003–1017, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=5&spage=1003.

Spiteri:2003:NCM

[2612] Pierre Spiteri. A new characterization of M-matrices and Hmatrices. BIT Numerical Mathematics, 43(5):1019–1032, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=5&spage=1019.

Strakos:2003:E

[2613] Zdenek Strakos and Daniel B. Szyld. Editorial. BIT Numerical Mathematics, 43(5):??, December 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=43&issue=5&spage=??.

Strakos:2003:ECC

[2614] Zdeněk Strakoš and Daniel B. Szyld. Editorial [Conference on Computational Linear Algebra with Applications]. BIT Numerical Mathematics, 43(5):iii–v, 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). Held in Milovy, August 4–10, 2002.

Alonso-Mallo:2004:AOR

[2615] I. Alonso-Mallo and B. Cano. Avoiding order reduction of Runge–Kutta discretizations for linear time-dependent parabolic problems. BIT Numerical Mathematics, 44(1):1–20, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=1.

Burrage:2004:ISR

[2616] Kevin Burrage and Tianhai Tian. Implicit stochastic Runge–Kutta methods for stochastic differential equations. BIT Numerical Mathematics, 44(1):21– 39, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=21.

Dieci:2004:PPP

[2617] Luca Dieci and Jorge Rebaza. Point-to-periodic and periodic-to-periodic connections. BIT Numerical Mathematics, 44(1):41–62, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=41. See

erratum [2655].

Lewanowicz:2004:GBP

[2618] Stanisław Lewanowicz and Paweł Woźny. Generalized Bernstein polynomials. BIT Numerical Mathematics, 44(1):63–78, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=1&spage=63.

Liesen:2004:WCG

[2619] Jörg Liesen and Petr Tichý. The worstcase GMRES for normal matrices. BIT Numerical Mathematics, 44(1):79–98, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=1&spage=79.

McLachlan:2004:NSS

[2620] Robert I. McLachlan, Matthew Perlmutter, and G. R. W. Quispel. On the nonlinear stability of symplectic integrators. BIT Numerical Mathematics, 44(1):99–117, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=1&spage=99.

Reginska:2004:RDI

[2621] Teresa Regińska. Regularization of discrete ill-posed problems. BIT Numerical Mathematics, 44(1):119–133, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=1&spage=119.

Sezgin:2004:MSS

[2622] F. Sezgin. A method of systematic search for optimal multipliers in congruential random number generators. BIT Numerical Mathematics, 44(1):135–149, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=135.

Silvester:2004:BBM

[2623] David J. Silvester and Milan D.

Mihajlović. A black-box multigrid preconditioner for the biharmonic equation. BIT Numerical Mathematics, 44(1):151–163, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=151.

Thomee:2004:MNE

[2624] V. Thomée and L. B. Wahlbin.

Maximum-norm estimates for finiteelement methods for a strongly damped wave equation. BIT Numerical Mathematics, 44(1):165–179, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=165.

Zhao:2004:SMF

[2625] Jinxi Zhao, Weiguo Wang, and Weiqing Ren. Stability of the matrix factorization for solving block tridiagonal symmetric indefinite linear systems. BIT Numerical Mathematics, 44(1):181– 188, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=181.

Yuan:2004:SCD

[2626] J. Y. Yuan, G. H. Golub, R. J. Plemmons, and W. A. G. Cecílio. Semiconjugate direction methods for real positive definite systems. BIT Numerical Mathematics, 44(1):189–207, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=189.

Anonymous:2004:ICIa

[2627] Anonymous. Introduction to the contents of issue 44:1. BIT Numerical Mathematics, 44(1):??, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=??.

Anonymous:2004:Ea

[2628] Anonymous. Editorial. BIT Numerical Mathematics, 44(1):??, January 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=1&spage=??.

Byers:2004:CCE

[2629] R. Byers and D. Kressner. On the condition of a complex eigenvalue under real perturbations. BIT Numerical Mathematics, 44(2):209– 214, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=2&spage=209.

Cano:2004:TIE

[2630] B. Cano and A. Duran. A technique to improve the error propagation when integrating relative equilibria. BIT Numerical Mathematics, 44(2): 215–235, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=2&spage=215.

Chen:2004:SNP

[2631] Xiao shan Chen, Wen Li, and Weiwei Sun. Some new perturbation bounds for the generalized polar decomposition. BIT Numerical Mathematics, 44(2):237–244, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=2&spage=237.

Daras:2004:BCG

[2632] Nicholas J. Daras. On the best choice for the generating polynomial of a Padé-type approximation. BIT Numerical Mathematics, 44(2):245– 257, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=245.

Demichelis:2004:FPI

[2633] V. Demichelis and P. Rabinowitz. Finite-part integrals and modified splines. BIT Numerical Mathematics, 44(2):259–267, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=259.

Jakobsen:2004:EBA

[2634] E. R. Jakobsen. On error bounds for approximation schemes for nonconvex degenerate elliptic equations. BIT Numerical Mathematics, 44(2): 269–285, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=269.

Sayas:2004:ANE

[2635] Francisco-Javier Sayas. Aubin– Nitsche estimates are equivalent to compact embeddings. BIT Numerical Mathematics, 44(2):287–290, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=287.

Skrzipek:2004:IVL

[2636] M.-R. Skrzipek. Inversion of Vandermondelike matrices. BIT Numerical Mathematics, 44(2):291–306, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=44&
issue=2&spage=291.

Sundberg:2004:STP

[2637] S. Sundberg and L. von Sydow. Semi-Toeplitz preconditioning for the linearized Navier–Stokes equations. BIT Numerical Mathematics, 44(2):307– 341, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=2&spage=307.

Thalhammer:2004:CBV

[2638] Mechthild Thalhammer. On the convergence behaviour of variable stepsize multistep methods for singularly perturbed problems. BIT Numerical Mathematics, 44(2):343– 361, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=2&spage=343.

Wang:2004:SHT

[2639] Li Wang and Zhong-Zhi Bai. Skew-Hermitian triangular splitting iteration methods for non-Hermitian positive definite linear systems of strong skew-Hermitian parts. BIT Numerical Mathematics, 44(2):363– 386, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=363.

Voss:2004:AMN

[2640] H. Voss. An Arnoldi method for nonlinear eigenvalue problems. BIT Numerical Mathematics, 44(2):387– 401, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=387.

Anonymous:2004:Eb

[2641] Anonymous.  Editorial.  BIT

Numerical Mathematics, 44(2):??, May 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=2&spage=??.

Arponen:2004:EGI

[2642] T. Arponen and B. Leimkuhler. An efficient geometric integrator for thermostatted anti-/ferromagnetic models. BIT Numerical Mathematics, 44 (3):403–424, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=3&spage=403.

Brouder:2004:TRD

[2643] Ch. Brouder. Trees, renormalization and differential equations. BIT Numerical Mathematics, 44(3):425–438, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=425.

Cotter:2004:AIA

[2644] Colin John Cotter and Sebastian Reich. Adiabatic invariance and applications: from molecular dynamics to numerical weather prediction. BIT Numerical Mathematics, 44(3):439– 455, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=439.

Hazra:2004:SPT

[2645] Subhendu Bikash Hazra and Volker Schulz. Simultaneous pseudo-timestepping for PDE-model based optimization problems. BIT Numerical Mathematics, 44(3):457–472, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=457.

Iserles:2004:MNS

[2646] A. Iserles. On the method of Neumann series for highly oscillatory equations. BIT Numerical Mathematics, 44(3):473–488, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=3&spage=473.

Kvaerno:2004:SDI

[2647] A. Kværnø. Singly diagonally implicit Runge–Kutta methods with an explicit first stage. BIT Numerical Mathematics, 44(3):489–502, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=3&spage=489.

Lubich:2004:CQR

[2648] Christian Lubich. Convolution quadrature revisited. BIT Numerical Mathematics, 44(3):503–514, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=3&spage=503.

McLachlan:2004:EGI

[2649] Robert I. McLachlan and G. Reinout W. Quispel. Explicit geometric integration of polynomial vector fields. BIT Numerical Mathematics, 44(3):515– 538, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=3&spage=515.

Tordesillas:2004:BLS

[2650] A. Tordesillas, S. D. C. Walsh, and B. S. Gardiner. Bridging the length scales: micromechanics of granular media. BIT Numerical Mathematics, 44(3):539–556, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=539.

Butcher:2004:USG

[2651] J. C. Butcher and Z. Jackiewicz.

Unconditionally stable general linear methods for ordinary differential equations. BIT Numerical Mathematics, 44(3):557–570, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=557.

Dullin:2004:ADD

[2652] Holger Dullin and Francesco Fassò.

An algorithm for detecting directional quasi-convexity. BIT Numerical Mathematics, 44(3):571–584, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=571.

Li:2004:ACE

[2653] Ren-Cang Li. Accuracy of computed eigenvectors via optimizing a Rayleigh quotient. BIT Numerical Mathematics, 44(3):585–593, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=585.

Moret:2004:RRA

[2654] I. Moret and P. Novati. RDrational approximations of the matrix exponential. BIT Numerical Mathematics, 44(3):595–615, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=595.

Dieci:2004:EPP

[2655] Luca Dieci and Jorge Rebaza. Erratum: “Point-to-periodic and periodic-to-periodic connections” [BIT 44 (2004), no. 1, 41–62; MR2057361]. BIT Numerical Mathematics, 44(3):617–618, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=617. See

[2617].

Anonymous:2004:ICIb

[2656] Anonymous. Introduction to the contents of issue 44:3. BIT Numerical Mathematics, 44(3):??, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=??.

Anonymous:2004:Ec

[2657] Anonymous. Editorial. BIT Numerical Mathematics, 44(3):??, August 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=3&spage=??.

Anonymous:2004:P

[2658] Anonymous. Preface. BIT Numerical Mathematics, 44(4):3–4, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=4&spage=3.

Ruhe:2004:ICI

[2659] Axel Ruhe. Introduction to the contents of issue 44:4. BIT Numerical Mathematics, 44(4):5, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=4&spage=5.

Atieg:2004:IOD

[2660] A. Atieg and G. A. Watson. Incomplete orthogonal distance regression. BIT Numerical Mathematics, 44(4):619–629, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=44&
issue=4&spage=619.

Brunner:2004:DCS

[2661] Hermann Brunner. On the divergence of collocation solutions in smooth piecewise polynomial spaces for Volterra integral equations. BIT Numerical Mathematics, 44(4):631–650, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=4&spage=631.

Bueno:2004:SST

[2662] M. Isabel Bueno and Froilán M. Dopico. Stability and sensitivity of tridiagonal LU factorization without pivoting. BIT Numerical Mathematics, 44(4):651–673, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=4&spage=651.

Englund:2004:CSI

[2663] Jonas Englund and Johan Helsing. A comparison of splittings and integral equation solvers for a nonseparable elliptic equation. BIT Numerical Mathematics, 44(4):675–697, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=4&spage=675.

Faou:2004:ECN

[2664] Erwan Faou, Ernst Hairer, and Truong-Linh Pham. Energy conservation with non-symplectic methods: Examples and counter-examples. BIT Numerical Mathematics, 44(4):699–709, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=4&spage=699.

Gautschi:2004:GGR

[2665] Walter Gautschi. Generalized Gauss– Radau and Gauss–Lobatto formulae. BIT Numerical Mathematics, 44(4): 711–720, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=4&spage=711.

Hochstenbach:2004:HRE

[2666] M. E. Hochstenbach. Harmonic and refined extraction methods for the singular value problem, with applications in least squares problems. BIT Numerical Mathematics, 44(4):721–754, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=4&spage=721.

Iserles:2004:QMH

[2667] A. Iserles and S. P. NØrsett. On quadrature methods for highly oscillatory integrals and their implementation. BIT Numerical Mathematics, 44(4):755–772, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=44&
                                        issue=4&spage=755.

Nilsson:2004:SHF

[2668] Martin Nilsson. Stability of the high frequency fast multipole method for Helmholtz’ equation in three dimensions. BIT Numerical Mathematics, 44(4):773–791, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=44&
                                        issue=4&spage=773.

Sima:2004:RTL

[2669] Diana M. Sima, Sabine Van Huffel, and Gene H. Golub. Regularized total least squares based on quadratic eigenvalue problem solvers. BIT Numerical Mathematics, 44(4):793–812, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=4&spage=793.

Xie:2004:ACP

[2670] Dexuan Xie. Analysis of a class of parallel multigrid smoothers. BIT Numerical Mathematics, 44(4):813–828, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=44&issue=4&spage=813.

Yan:2004:SGA

[2671] Yubin Yan. Semidiscrete Galerkin approximation for a linear stochastic parabolic partial differential equation driven by an additive noise. BIT Numerical Mathematics, 44(4):829–847, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=44&issue=4&spage=829.

Carbonell:2005:LLM

[2672] F. Carbonell, J. C. Jimenez, R. J.

Biscay, and H. de la Cruz. The local linearization method for numerical integration of random differential equations. BIT Numerical Mathematics, 45(1):1–14, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=45&
issue=1&spage=1.

Favati:2005:PBF

[2673] P. Favati, G. Lotti, and O. Menchi.

Preconditioners based on fit techniques for the iterative regularization in the image deconvolution problem. BIT Numerical Mathematics, 45(1):15– 35, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=15.

Jakobsen:2005:CRS

[2674] E. R. Jakobsen and K. H. Karlsen.

Convergence rates for semi-discrete splitting approximations for degenerate parabolic equations with source terms. BIT Numerical Mathematics, 45(1):37– 67, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=37.

Kalisch:2005:EAS

[2675] Henrik Kalisch. Error analysis of a spectral projection of the regularized Benjamin–Ono equation. BIT Numerical Mathematics, 45(1):69–89, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=69.

Lorenz:2005:AIH

[2676] Katina Lorenz, Tobias Jahnke, and Christian Lubich. Adiabatic integrators for highly oscillatory secondorder linear differential equations with time-varying eigendecomposition. BIT Numerical Mathematics, 45(1):91–115, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=91.

Milovanovic:2005:EES

[2677] G. V. Milovanović and M. M. Spalević. An error expansion for some Gauss– Turán quadratures and L^{1}-estimates of the remainder term. BIT Numerical Mathematics, 45(1):117–136, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=117.

Negoita:2005:CGL

[2678] C. Negoita and R. A. Renaut. On the convergence of the generalized linear least squares algorithm. BIT Numerical Mathematics, 45(1):137– 158, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=137.

Papadopoulos:2005:CIO

[2679] A. T. Papadopoulos, I. S. Duff, and A. J. Wathen. A class of incomplete orthogonal factorization methods. II. Implementation and results. BIT Numerical Mathematics, 45(1):159– 179, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=159.

Robbe:2005:CAB

[2680] M. Robbé and M. Sadkane. Convergence analysis of the block Householder block diagonalization algorithm. BIT Numerical Mathematics, 45(1):181– 195, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=1&spage=181.

Schmitt:2005:MIP

[2681] Bernhard A. Schmitt, Rüdiger Weiner, and Helmut Podhaisky. Multi-implicit peer two-step W -methods for parallel time integration. BIT Numerical Mathematics, 45(1):197–217, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=1&spage=197.

Ruhe:2005:ICIa

[2682] Axel Ruhe. Introduction to the contents of issue 45:1. BIT Numerical Mathematics, 45(1):??, March 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=45&
                                        issue=1&spage=??.

Aparicio:2005:NEE

[2683] Nairo D. Aparicio, Simon J. A. Malham, and Marcel Oliver. Numerical evaluation of the Evans function by Magnus integration. BIT Numerical Mathematics, 45(2):219– 258, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=2&spage=219.

Barlow:2005:IGS

[2684] Jesse L. Barlow, Alicja Smoktunowicz, and Hasan Erbay. Improved Gram– Schmidt type downdating methods. BIT Numerical Mathematics, 45(2):

259–285, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=2&spage=259.

Cohen:2005:NEC

[2685] D. Cohen, E. Hairer, and Ch. Lubich. Numerical energy conservation for multi-frequency oscillatory differential equations. BIT Numerical Mathematics, 45(2):287– 305, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=2&spage=287.

Du:2005:AAE

[2686] Qiang Du and Wenxiang Zhu. Analysis and applications of the exponential time differencing schemes and their contour integration modifications. BIT Numerical Mathematics, 45(2): 307–328, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=2&spage=307.

Elden:2005:MLF

[2687] L. Eldén, P. C. Hansen, and M. Rojas. Minimization of linear functionals defined on solutions of largescale discrete ill-posed problems. BIT Numerical Mathematics, 45(2):329– 340, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=2&spage=329.

Layton:2005:ICQ

[2688] Anita T. Layton and Michael L. Minion. Implications of the choice of quadrature nodes for Picard integral deferred corrections methods for ordinary differential equations. BIT Numerical Mathematics, 45(2):341– 373, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=2&spage=341.

Marcinkowski:2005:ASM

[2689] Leszek Marcinkowski. Additive Schwarz method for mortar discretization of elliptic problems with P_{1} nonconforming finite elements. BIT Numerical Mathematics, 45(2):375– 394, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=2&spage=375.

Wulling:2005:SWL

[2690] Wolfgang Wülling. The stabilization of weights in the Lánczos and conjugate gradient method. BIT Numerical Mathematics, 45(2):395– 414, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=2&spage=395.

Zilinskas:2005:UIC

[2691] A. Zilinskas and J. Zilinskas. On underestimating in interval computations. BIT Numerical Mathematics, 45(2): 415–427, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=2&spage=415.

Ruhe:2005:ICIb

[2692] Axel Ruhe. Introduction to the contents of issue 45:2. BIT Numerical Mathematics, 45(2):??, June 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=45&
issue=2&spage=??.

Akrivis:2005:LIF

[2693] G. Akrivis and S. Larsson. Linearly implicit finite element methods for the time-dependent Joule heating problem. BIT Numerical Mathematics, 45(3):429–442, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=429.

Al-Subaihi:2005:FPC

[2694] I. Al-Subaihi and G. A. Watson. Fitting parametric curves and surfaces by l_{∞} distance regression. BIT Numerical Mathematics, 45(3):443–461, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=3&spage=443.

Birken:2005:SPF

[2695] P. Birken and A. Meister. Stability of preconditioned finite volume schemes at low Mach numbers. BIT Numerical Mathematics, 45(3):463–480, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=3&spage=463.

Branden:2005:PBF

[2696] H. Brandén and P. Sundqvist. Preconditioners based on fundamental solutions. BIT Numerical Mathematics, 45(3):481–494, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=3&spage=481.

Freund:2005:KSA

[2697] Roland W. Freund. Krylov subspaces associated with higher-order linear dynamical systems. BIT Numerical Mathematics, 45(3):495–516, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=3&spage=495.

Kulikov:2005:GEE

[2698] G. Yu. Kulikov and S. K. Shindin.

Global error estimation and extrapolated multistep methods for index 1 differential-algebraic systems. BIT Numerical Mathematics, 45(3):517–542, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=517.

Kumar:2005:WTG

[2699] B. V. Rathish Kumar and M. Mehra.

Wavelet-Taylor Galerkin method for the Burgers equation. BIT Numerical Mathematics, 45(3):543–560, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=543.

Nedialkov:2005:SDA

[2700] Nedialko S. Nedialkov and John D.

Pryce. Solving differential-algebraic equations by Taylor series (I): Computing Taylor coefficients. BIT Numerical Mathematics, 45(3):561–591, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=561.

Gautschi:2005:NQC

[2701] Walter Gautschi. Numerical quadrature computation of the Macdonald function for complex orders. BIT Numerical Mathematics, 45(3):593–603, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=593.

Verheyden:2005:NPC

[2702] Koen Verheyden and Kurt Lust. A Newton–Picard collocation method for periodic solutions of delay differential equations. BIT Numerical Mathematics, 45(3):605–625, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=605.

Zemui:2005:FOS

[2703] Abraham Zemui. Fourth order symmetric finite difference schemes for the acoustic wave equation. BIT Numerical Mathematics, 45(3):627–651, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=627.

Ruhe:2005:ICIc

[2704] Axel Ruhe. Introduction to the contents of issue 45:3. BIT Numerical Mathematics, 45(3):??, September 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=3&spage=??.

Beliakov:2005:MPA

[2705] G. Beliakov. Monotonicity preserving approximation of multivariate scattered data. BIT Numerical Mathematics, 45(4):653–677, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=653.

Floater:2005:ALE

[2706] M. S. Floater. Arc length estimation and the convergence of polynomial curve interpolation. BIT Numerical Mathematics, 45(4):679–694, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=679.

Guzman:2005:QSP

[2707] J. Guzmán. Quadrature and Schatz’s pointwise estimates for finite element methods. BIT Numerical Mathematics, 45(4):695–707, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=695.

Grimm:2005:GIM

[2708] V. Grimm and G. R. W. Quispel. Geometric integration methods that preserve Lyapunov functions. BIT Numerical Mathematics, 45(4):709–723, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=709.

Hanslien:2005:MPE

[2709] Monica Hanslien, Kenneth H. Karlsen, and Aslak Tveito. A maximum principle for an explicit finite difference scheme approximating the Hodgkin–Huxley model. BIT Numerical Mathematics, 45(4):725–741, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=725.

Liu:2005:ASN

[2710] M. Z. Liu, Z. W. Yang, and G. D. Hu. Asymptotical stability of numerical methods with constant stepsize for pantograph equations. BIT Numerical Mathematics, 45(4):743–759, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=743.

Rang:2005:NRM

[2711] J. Rang and L. Angermann. New Rosenbrock W -methods of order 3 for partial differential algebraic equations of index 1. BIT Numerical Mathematics, 45(4):761–787, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=761.

Strakos:2005:EEP

[2712] Zdenek Strakos and Petr Tichý. Error estimation in preconditioned conjugate gradients. BIT Numerical Mathematics, 45(4):789–817, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=789.

Aahlander:2005:AGF

[2713] Krister Åhlander and Hans Munthe-Kaas. Applications of the Generalized Fourier Transform in numerical linear algebra. BIT Numerical Mathematics, 45(4):819–850, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=819.

Analysis:2005:AIV

[2714] Department of Numerical Analysis, Computer Science (NADA), Royal Institute of Technology (KTH), Stockholm, and Sweden. Author index volume 40-45, 2000-2005. BIT Numerical Mathematics, 45(4):851–870, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=45&issue=4&spage=851.

Anonymous:2005:E

[2715] Anonymous. Editorial. BIT Numerical Mathematics, 45(4):v–vi, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=4&spage=??.

Ruhe:2005:ICId

[2716] Axel Ruhe. Introduction to the contents of issue 45:4. BIT Numerical Mathematics, 45(4):vii, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=45&issue=4&spage=??.

Runborg:2006:P

[2717] Olof Runborg and Anna-Karin Tornberg. Preface. BIT Numerical Mathematics, 46(S-1):1–3, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Benamou:2006:SLN

[2718] J.-D. Benamou. A semi-Lagrangian numerical method for geometric optics type problems. BIT Numerical Mathematics, 46(S-1):5–17, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Carleson:2006:CPM

[2719] L. Carleson. Computations in pure mathematics (outline of lecture at the B. Engquist conference). BIT Numerical Mathematics, 46(S-1):19–20, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Halpern:2006:ABC

[2720] L. Halpern. Absorbing boundary conditions and optimized Schwarz waveform relaxation. BIT Numerical Mathematics, 46(S-1):21–34, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jeltsch:2006:CPF

[2721] Rolf Jeltsch and Manuel Torrilhon.

On curl-preserving finite volume discretizations for shallow water equations. BIT Numerical Mathematics, 46 (S-1):35–53, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Kreiss:2006:NVC

[2722] H.-O. Kreiss and J. Yström. A note on viscous conservation laws with complex characteristics. BIT Numerical Mathematics, 46(S-1):55–59, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ferm:2006:CSF

[2723] Lars Ferm, Per Lötstedt, and Paul Sjöberg. Conservative solution of the Fokker–Planck equation for stochastic chemical reactions. BIT Numerical Mathematics, 46(S-1):61–83, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Calvez:2006:LFT

[2724] V. Calvez and B. Perthame. A Lyapunov function for a two-chemical species version of the chemotaxis model. BIT Numerical Mathematics, 46(S-1):85–97, November 2006. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Olsson:2006:RKE

[2725] K. Henrik A. Olsson and Axel Ruhe. Rational Krylov for eigenvalue computation and model order reduction. BIT Numerical Mathematics, 46(S-1): 99–111, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Chatzipantelidis:2006:PFE

[2726] P. Chatzipantelidis, R. D. Lazarov, V. Thomée, and L. B. Wahlbin. Parabolic finite element equations in nonconvex polygonal domains. BIT Numerical Mathematics, 46(S-1):113– 143, November 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Tornberg:2006:NTB

[2727] A.-K. Tornberg, B. Engquist, B. Gustafsson, and P. Wahlund. A new type of boundary treatment for wave propagation. BIT Numerical Mathematics, 46(S-1):145–170, November 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=1&spage=145.

Ruhe:2006:P

[2728] Axel Ruhe. Preface. BIT Numerical Mathematics, 46(1):1, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:2006:ICIa

[2729] Axel Ruhe. Introduction to the contents of issue 46:1. BIT Numerical Mathematics, 46(1):3, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Butcher:2006:LMM

[2730] J. C. Butcher and A. T. Hill. Linear multistep methods as irreducible general linear methods. BIT Numerical Mathematics, 46(1):5–19, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gautschi:2006:CKL

[2731] Walter Gautschi. Computing the Kontorovich–Lebedev integral transforms and their inverses. BIT Numerical Mathematics, 46(1):21–40, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hansen:2006:ERI

[2732] P. C. Hansen, M. E. Kilmer, and R. H.

Kjeldsen. Exploiting residual information in the parameter choice for discrete ill-posed problems. BIT Numerical Mathematics, 46(1):41–59, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Huckle:2006:MMB

[2733] Thomas Huckle and Jochen Staudacher. Multigrid methods for block Toeplitz matrices with small size blocks. BIT Numerical Mathematics, 46(1):61–83, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Li:2006:MTM

[2734] Zhiping Li. The mesh transformation method and optimal finite element solutions. BIT Numerical Mathematics, 46(1):85–95, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Rossler:2006:RKM

[2735] Andreas Rößler. Runge–Kutta methods for Itô stochastic differential equations with scalar noise. BIT Numerical Mathematics, 46(1):97–110, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Shardlow:2006:MES

[2736] Tony Shardlow. Modified equations for stochastic differential equations. BIT Numerical Mathematics, 46(1): 111–125, March 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Siewer:2006:MS

[2737] Ralf Siewer. Martensen splines. BIT Numerical Mathematics, 46(1):127– 140, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Singer:2006:IQF

[2738] Sanja Singer. Indefinite QR factorization. BIT Numerical Mathematics, 46(1):141–161, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=1&spage=141.

Smyrlis:2006:MFS

[2739] Yiorgos-Sokratis Smyrlis. The method of fundamental solutions: a weighted least–squares approach. BIT Numerical Mathematics, 46(1):163–194, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=1&spage=163.

Waldvogel:2006:FCF

[2740] Jörg Waldvogel. Fast construction of the Fejér and Clenshaw– Curtis quadrature rules. BIT Numerical Mathematics, 46(1):195–202, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=1&spage=195.

Xu:2006:CNS

[2741] W. Xu, Y. Wei, and S. Qiao.

Condition numbers for structured least squares problems. BIT Numerical Mathematics, 46(1):203–225, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=46&issue=1&spage=203.

Szepessy:2006:PIC

[2742] Anders Szepessy. Preface and introduction to the contents of issue 46:2. BIT Numerical Mathematics, 46(2): 227, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bellen:2006:SCM

[2743] A. Bellen, H. Brunner, S. Maset, and L. Torelli. Superconvergence in collocation methods on quasi-graded meshes for functional differential equations with vanishing delays. BIT Numerical Mathematics, 46(2):229–247, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Boman:2006:EPC

[2744] Mats Boman. Estimates for the L[2]- projection onto continuous finite element spaces in a weighted L[p]-norm. BIT Numerical Mathematics, 46(2): 249–260, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Buckwar:2006:AMS

[2745] E. Buckwar, R. Horváth-Bokor, and R. Winkler. Asymptotic mean-square stability of two-step methods for stochastic ordinary differential equations. BIT Numerical Mathematics, 46(2):261–282, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Elman:2006:FAM

[2746] Howard C. Elman and Alison Ramage. Fourier analysis of multigrid for a model two-dimensional convectiondiffusion equation. BIT Numerical Mathematics, 46(2):283–306, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Helsing:2006:AIP

[2747] Johan Helsing. Approximate inverse preconditioners for some large dense random electrostatic interaction matrices. BIT Numerical Mathematics, 46(2):307–323, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Henn:2006:TRI

[2748] Stefan Henn. A translation and rotation invariant Gauss–Newton like scheme for image registration. BIT Numerical Mathematics, 46(2):325– 344, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Laszkiewicz:2006:AMF

[2749] B. Laszkiewicz and K. Zietak. Approximation of matrices and a family of Gander methods for polar decomposition. BIT Numerical Mathematics, 46(2):345–366, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Moon:2006:CRA

[2750] K.-S. Moon, E. von Schwerin, A. Szepessy, and R. Tempone. Convergence rates for an adaptive dual weighted residual finite element algorithm. BIT Numerical Mathematics, 46(2):367–407, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ostermann:2006:CEE

[2751] A. Ostermann, M. Thalhammer, and W. M. Wright. A class of explicit exponential general linear methods. BIT Numerical Mathematics, 46(2): 409–431, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Rump:2006:VPD

[2752] S. M. Rump. Verification of positive definiteness. BIT Numerical Mathematics, 46(2):433–452, June 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=2&spage=433.

Soderlind:2006:P

[2753] Gustaf Söderlind. Preface. BIT Numerical Mathematics, 46(3):453–454, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bartels:2006:SOL

[2754] R. H. Bartels, G. H. Golub, and F. F. Samavati. Some observations on local least squares. BIT Numerical Mathematics, 46(3):455–477, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Butcher:2006:TYS

[2755] J. C. Butcher. Thirty years of Gstability. BIT Numerical Mathematics, 46(3):479–489, September 2006. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gander:2006:GNA

[2756] Walter Gander. Generating numerical algorithms using a computer algebra system. BIT Numerical Mathematics, 46(3):491–504, September 2006. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gear:2006:TEM

[2757] C. W. Gear. Towards explicit methods for differential algebraic equations. BIT Numerical Mathematics, 46(3): 505–514, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hairer:2006:SLM

[2758] Ernst Hairer. Symmetric linear multistep methods. BIT Numerical Mathematics, 46(3):515–524, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:2006:NSD

[2759] D. J. Higham and X. Mao. Nonnormality and stochastic differential equations. BIT Numerical Mathematics, 46 (3):525–532, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Huhtanen:2006:MRA

[2760] Marko Huhtanen and Olavi Nevanlinna. A minimum residual algorithm for solving linear systems. BIT Numerical Mathematics, 46(3):533–548, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Iserles:2006:CHO

[2761] A. Iserles and S. P. Nørsett. On the computation of highly oscillatory multivariate integrals with stationary points. BIT Numerical Mathematics, 46(3):549–566, September 2006. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jeltsch:2006:RCR

[2762] Rolf Jeltsch. Reducibility and contractivity of Runge–Kutta methods revisited. BIT Numerical Mathematics, 46 (3):567–587, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Morigi:2006:ILR

[2763] S. Morigi, L. Reichel, and F. Sgallari. An iterative Lavrentiev regularization method. BIT Numerical Mathematics, 46(3):589–606, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=3&spage=589. In

memory of Germund Dahlquist (1925– 2005).

Reich:2006:LIT

[2764] Sebastian Reich. Linearly implicit time stepping methods for numerical weather prediction. BIT Numerical Mathematics, 46(3):607–616, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=3&spage=607. In

memory of Germund Dahlquist (1925– 2005).

Skelboe:2006:APT

[2765] Stig Skelboe. Adaptive partitioning techniques for ordinary differential equations. BIT Numerical Mathematics, 46(3):617–629, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=3&spage=617. In

memory of Germund Dahlquist (1925– 2005).

Soderlind:2006:LNH

[2766] Gustaf Söderlind. The logarithmic norm. History and modern theory. BIT Numerical Mathematics, 46(3):631–652, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=46&
issue=3&spage=631. In memory of

Germund Dahlquist (1925–2005).

Trefethen:2006:TQR

[2767] L. N. Trefethen, J. A. C. Weideman, and T. Schmelzer. Talbot quadratures and rational approximations. BIT Numerical Mathematics, 46(3):653–670, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=46&issue=3&spage=653. In

memory of Germund Dahlquist (1925– 2005).

Wanner:2006:DCP

[2768] G. Wanner. Dahlquist’s classical papers on stability theory. BIT Numerical Mathematics, 46(3):671–683, September 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=46&issue=3&spage=671. In

memory of Germund Dahlquist (1925– 2005).

Ruhe:2006:E

[2769] Axel Ruhe. Editorial. BIT Numerical Mathematics, 46(4):685–686, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:2006:ICIb

[2770] Axel Ruhe. Introduction to the contents of issue 46:4. BIT Numerical Mathematics, 46(4):687–688, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Alcock:2006:NBM

[2771] Jamie Alcock and Kevin Burrage. A note on the Balanced Method. BIT Numerical Mathematics, 46(4):689–710, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cullum:2006:MFP

[2772] J. K. Cullum and M. Tuma. Matrixfree preconditioning using partial matrix estimation. BIT Numerical Mathematics, 46(4):711–729, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Garcia-Archilla:2006:KMD

[2773] Bosco García-Archilla, Juan Sánchez, and Carles Simó. Krylov methods and determinants for detecting bifurcations in one parameter dependent partial differential equations. BIT Numerical Mathematics, 46(4):731–757, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Golub:2006:ATA

[2774] G. H. Golub and C. Greif. An Arnolditype algorithm for computing page rank. BIT Numerical Mathematics, 46 (4):759–771, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hausenblas:2006:NAP

[2775] Erika Hausenblas and Iuliana Marchis.

A numerical approximation of parabolic stochastic partial differential equations driven by a Poisson random measure. BIT Numerical Mathematics, 46(4): 773–811, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lopez:2006:PGP

[2776] L. Lopez and V. Simoncini. Preserving geometric properties of the exponential matrix by block Krylov subspace methods. BIT Numerical Mathematics, 46 (4):813–830, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lyche:2006:HSS

[2777] Tom Lyche and Jean-Louis Merrien. Hermite subdivision with shape constraints on a rectangular mesh. BIT Numerical Mathematics, 46(4):831– 859, December 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hoang:2006:FFR

[2778] Nguyen Si Hoang, Roger B. Sidje, and Nguyen Huu Cong. On functionallyfitted Runge–Kutta methods. BIT Numerical Mathematics, 46(4):861–874, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Thi:2006:PET

[2779] N. N. Pham Thi, W. Hundsdorfer, and B. P. Sommeijer. Positivity for explicit two-step methods in linear multistep and one-leg form. BIT Numerical Mathematics, 46(4):875–882, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=4&spage=875.

Wermuth:2006:PIL

[2780] N. Wermuth, M. Wiedenbeck, and D. R. Cox. Partial inversion for linear systems and partial closure of independence graphs. BIT Numerical Mathematics, 46(4):883–901, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=46&issue=4&spage=883.

Ruhe:2007:PIC

[2781] Axel Ruhe. Preface and introduction to the contents of issue 47:1. BIT Numerical Mathematics, 47(1):1– 2, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=1&spage=1.

Ascher:2007:ATI

[2782] U. M. Ascher, H. Huang, and K. van den Doel. Artificial time integration. BIT Numerical Mathematics, 47(1):3–25, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=1&spage=3.

Freitag:2007:CII

[2783] M. A. Freitag and A. Spence. Convergence of inexact inverse iteration with application to preconditioned iterative solves. BIT Numerical Mathematics, 47(1):27–44, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=1&spage=27.

Grigori:2007:RMT

[2784] L. Grigori, M. Cosnard, and E. G. Ng. On the row merge tree for sparse LU factorization with partial pivoting. BIT Numerical Mathematics, 47(1):45– 76, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=1&spage=45.

He:2007:FEM

[2785] Wen-Ming He and Jun-Zhi Cui. A finite element method for elliptic problems with rapidly oscillating coefficients. BIT Numerical Mathematics, 47(1):77– 102, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=1&spage=77.

Jensen:2007:IRM

[2786] T. K. Jensen and P. C. Hansen. Iterative regularization with minimumresidual methods. BIT Numerical Mathematics, 47(1):103–120, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=47&
                                        issue=1&spage=103.

Nedialkov:2007:SDA

[2787] Nedialko S. Nedialkov and John D. Pryce. Solving differential-algebraic equations by Taylor series (II): Computing the System Jacobian. BIT Numerical Mathematics, 47(1):121– 135, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=1&spage=121.

Savcenco:2007:MTS

[2788] V. Savcenco, W. Hundsdorfer, and J. G. Verwer. A multirate time stepping strategy for stiff ordinary differential equations. BIT Numerical Mathematics, 47(1):137–155, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=1&spage=137.

Sickenberger:2007:LEE

[2789] Thorsten Sickenberger, Ewa Weinmüller, and Renate Winkler. Local error estimates for moderately smooth problems: Part I: ODEs and DAEs. BIT Numerical Mathematics, 47(1):157– 187, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=1&spage=157.

Tocino:2007:PLT

[2790] A. Tocino. On preserving longtime features of a linear stochastic oscillator. BIT Numerical Mathematics, 47(1):189–196, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=47&
issue=1&spage=189.

Wright:2007:AMP

[2791] K. Wright. Adaptive methods for piecewise polynomial collocation for ordinary differential equations. BIT Numerical Mathematics, 47(1):197– 212, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=1&spage=197.

Aahlander:2007:SGF

[2792] Krister Åhlander and Daniel Henriksson. Sparse generalized Fourier transforms. BIT Numerical Mathematics, 47(1):213–237, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=1&spage=213.

Ruhe:2007:ICIa

[2793] Axel Ruhe. Introduction to the contents of issue 47:2. BIT Numerical Mathematics, 47(2):239–240, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=239.

Amodio:2007:HOG

[2794] P. Amodio and I. Sgura. High order generalized upwind schemes and numerical solution of singular perturbation problems. BIT Numerical Mathematics, 47(2):241– 257, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=2&spage=241.

Capobianco:2007:FRK

[2795] G. Capobianco, D. Conte, I. Del Prete, and E. Russo. Fast Runge–Kutta methods for nonlinear convolution systems of Volterra integral equations. BIT Numerical Mathematics, 47(2): 259–275, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=2&spage=259.

Chen:2007:AAM

[2796] Ke Chen and Joseph Savage. An accelerated algebraic multigrid algorithm for total-variation denoising. BIT Numerical Mathematics, 47(2): 277–296, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=2&spage=277.

Chen:2007:VCP

[2797] Chin-Yun Chen. Verified computed Peano constants and applications in numerical quadrature. BIT Numerical Mathematics, 47(2):297– 312, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=2&spage=297.

Disibuyuk:2007:GRB

[2798] Çetin Disibüyük and Halil Oruç. A generalization of rational Bernstein– Bézier curves. BIT Numerical Mathematics, 47(2):313–323, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=47&
                                        issue=2&spage=313.

Hoppensteadt:2007:NSV

[2799] F. C. Hoppensteadt, Z. Jackiewicz, and B. Zubik-Kowal. Numerical solution of Volterra integral and integrodifferential equations with rapidly vanishing convolution kernels. BIT Numerical Mathematics, 47(2):325– 350, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=325.

Iserles:2007:BSM

[2800] A. Iserles, G. R. W. Quispel, and P. S. P. Tse. B-series methods cannot be volume-preserving. BIT Numerical Mathematics, 47(2):351– 378, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=351.

Linss:2007:MNE

[2801] T. Linss. Maximum-norm error analysis of a non-monotone FEM for a singularly perturbed reaction-diffusion problem. BIT Numerical Mathematics, 47(2):379– 391, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=2&spage=379.

Muller-Gronbach:2007:IES

[2802] Thomas Müller-Gronbach and Klaus Ritter. An implicit Euler scheme with non-uniform time discretization for heat equations with multiplicative noise. BIT Numerical Mathematics, 47(2):393–418, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=47&
issue=2&spage=393.

Ostermann:2007:PER

[2803] Alexander Ostermann and Marnix van Daele. Positivity of exponential Runge–Kutta methods. BIT Numerical Mathematics, 47(2):419– 426, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=2&spage=419.

Pisinger:2007:BLS

[2804] G. Pisinger and A. Zimmermann.

Bivariate least squares approximation with linear constraints. BIT Numerical Mathematics, 47(2):427– 439, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=427.

Sommariva:2007:PGC

[2805] A. Sommariva and M. Vianello. Product Gauss cubature over polygons based on Green’s integration formula. BIT Numerical Mathematics, 47(2): 441–453, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=441.

Wendland:2007:SMS

[2806] Holger Wendland. On the stability of meshless symmetric collocation for boundary value problems. BIT Numerical Mathematics, 47(2):455– 468, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=455.

Xiang:2007:NAF

[2807] Shuhuang Xiang. Numerical analysis of a fast integration method for highly oscillatory functions. BIT Numerical Mathematics, 47(2):469– 482, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=2&spage=469.

Ruhe:2007:Ea

[2808] Axel Ruhe.  Editorial.  BIT

Numerical Mathematics, 47(3):483, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=483.

Ruhe:2007:ICIb

[2809] Axel Ruhe. Introduction to the contents of issue 47:3. BIT Numerical Mathematics, 47(3):485–486, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=485.

Chrysafinos:2007:API

[2810] Konstantinos Chrysafinos. Approximations of parabolic integro-differential equations using wavelet-Galerkin compression techniques. BIT Numerical Mathematics, 47(3):487–505, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=487.

Cottrell:2007:EDM

[2811] D. Cottrell and P. F. Tupper.

Energy drift in molecular dynamics simulations. BIT Numerical Mathematics, 47(3):507–523, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=507.

Kim:2007:INS

[2812] Pilwon Kim. Invariantization of numerical schemes using moving frames. BIT Numerical Mathematics, 47(3):525–546, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=47&
                                        issue=3&spage=525.

Kozak:2007:GIP

[2813] Jernej Kozak and Marjetka Krajnc. Geometric interpolation by planar cubic G^{1} splines. BIT Numerical Mathematics, 47(3):547–563, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=3&spage=547.

Layton:2007:SFE

[2814] William Layton. Superconvergence of finite element discretization of time relaxation models of advection. BIT Numerical Mathematics, 47(3):565–576, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=3&spage=565.

Li:2007:CCG

[2815] Ren-Cang Li. Convergence of CG and GMRES on a tridiagonal Toeplitz linear system. BIT Numerical Mathematics, 47(3):577–599, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=577.

Li:2007:ASM

[2816] Zhi-Ping Li and Carsten Carstensen.

An averaging scheme for macroscopic numerical simulation of nonconvex minimization problems. BIT Numerical Mathematics, 47(3):601–611, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=601.

Mazroui:2007:NMS

[2817] A. Mazroui, H. Mraoui, D. Sbibih, and A. Tijini. A new method for smoothing surfaces and computing Hermite interpolants. BIT Numerical Mathematics, 47(3):613–635, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=613.

Olver:2007:NAV

[2818] S. Olver. Numerical approximation of vector-valued highly oscillatory integrals. BIT Numerical Mathematics, 47(3):637–655, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=3&spage=637.

Rossler:2007:SOR

[2819] Andreas Rößler. Second order Runge– Kutta methods for Stratonovich stochastic differential equations. BIT Numerical Mathematics, 47(3):657– 680, September 2007. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=3&spage=657.

Yamamura:2007:EAF

[2820] Kiyotaka Yamamura and Koki Suda. An efficient algorithm for finding all solutions of separable systems of nonlinear equations. BIT Numerical Mathematics, 47(3):681–691, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=3&spage=681.

Ruhe:2007:Eb

[2821] Axel Ruhe. Editorial. BIT Numerical Mathematics, 47(4):693–695, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=4&spage=693.

Ruhe:2007:ICIc

[2822] Axel Ruhe. Introduction to the contents of issue 47:4. BIT Numerical Mathematics, 47(4):697, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=4&spage=697.

Annaby:2007:SBC

[2823] M. H. Annaby and M. M. Tharwat. Sinc-based computations of eigenvalues of Dirac systems. BIT Numerical Mathematics, 47(4):699–713, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=4&spage=699.

Burman:2007:NAV

[2824] Erik Burman. On nonlinear artificial viscosity, discrete maximum principle and hyperbolic conservation laws. BIT Numerical Mathematics, 47(4):715–733, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=47&
issue=4&spage=715.

Ferm:2007:NMC

[2825] Lars Ferm and Per Lötstedt. Numerical method for coupling the macro and meso scales in stochastic chemical kinetics. BIT Numerical Mathematics, 47(4):735–762, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=4&spage=735. In

memory of Germund Dahlquist (1925–

Granat:2007:CPD

[2826] Robert Granat, Bo Kågström, and Daniel Kressner. Computing periodic deflating subspaces associated with a specified set of eigenvalues. BIT Numerical Mathematics, 47(4):763–791, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=4&spage=763.

Huhtanen:2007:FMP

[2827] Marko Huhtanen. Factoring matrices into the product of two matrices. BIT Numerical Mathematics, 47(4):793–808, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=47&
                                        issue=4&spage=793.

Pisinger:2007:LLS

[2828] G. Pisinger and A. Zimmermann. Linear least squares problems with data over incomplete grids. BIT Numerical Mathematics, 47(4):809–824, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=4&spage=809.

Sablonniere:2007:QFA

[2829] P. Sablonnière. A quadrature formula associated with a univariate spline quasi interpolant. BIT Numerical Mathematics, 47(4):825–837, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=4&spage=825.

Wang:2007:IIR

[2830] X. H. Wang, C. C. Xie, and S. J. Yang. The Iyengar inequality, revisited. BIT Numerical Mathematics, 47(4):839–852, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=4&spage=839.

Xie:2007:NMP

[2831] Dexuan Xie and Shuzi Zhou. A new minimization protocol for solving nonlinear Poisson–Boltzmann mortar finite element equation. BIT Numerical Mathematics, 47(4):853–871, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=47&issue=4&spage=853.

Ye:2007:AAA

[2832] Qiang Ye, Hongyuan Zha, and Ren-Cang Li. Analysis of an alignment algorithm for nonlinear dimensionality reduction. BIT Numerical Mathematics, 47(4):873–885, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=47&issue=4&spage=873.

Ruhe:2008:E

[2833] Axel Ruhe.  Editorial.  BIT

Numerical Mathematics, 48(1):1–2, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=1&spage=1.

Ruhe:2008:ICIa

[2834] Axel Ruhe. Introduction to the contents of issue 48:1. BIT Numerical Mathematics, 48(1):3, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=48&
                                        issue=1&spage=3.

Conti:2008:FRP

[2835] C. Conti, M. Cotronei, and T. Sauer. Full rank positive matrix symbols: interpolation and orthogonality. BIT Numerical Mathematics, 48(1):5–27, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=1&spage=5.

Corless:2008:PCS

[2836] Robert M. Corless and Silvana Ilie. Polynomial cost for solving IVP for high-index DAE. BIT Numerical Mathematics, 48(1):29–49, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=1&spage=29.

Halidias:2008:NEM

[2837] Nikolaos Halidias and Peter E. Kloeden. A note on the Euler–Maruyama scheme for stochastic differential equations with a discontinuous monotone drift coefficient. BIT Numerical Mathematics, 48(1):51–59, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=1&spage=51.

Hoang:2008:SFF

[2838] Nguyen S. Hoang and Roger B. Sidje.

On the stability of functionally fitted Runge–Kutta methods. BIT Numerical Mathematics, 48(1):61–77, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=1&spage=61.

Lyness:2008:DRI

[2839] J. N. Lyness and S. Joe. Determination of the rank of an integration lattice. BIT Numerical Mathematics, 48(1):79– 93, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=1&spage=79.

Popova:2008:VPS

[2840] E. D. Popova and W. Krämer. Visualizing parametric solution sets. BIT Numerical Mathematics, 48(1):95– 115, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=1&spage=95.

Solo:2008:SEF

[2841] Aaron Solo. Sharp estimates for finite element approximations to parabolic problems with Neumann boundary data of low regularity. BIT Numerical Mathematics, 48(1):117–137, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=1&spage=117.

Wu:2008:STD

[2842] Xiaonan Wu, Gene H. Golub, José A. Cuminato, and Jin Yun Yuan. Symmetric-triangular decomposition and its applications part II: Preconditioners for indefinite systems. BIT Numerical Mathematics, 48(1):139– 162, March 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=1&spage=139.

Eirola:2008:PNA

[2843] Timo Eirola, Rolf Jeltsch, Claes Johnson, Erkki Somersalo, and Rolf Stenberg. Perspectives in Numerical Analysis 2008. BIT Numerical Mathematics, 48(2):163–165, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=163.

Babuska:2008:PFA

[2844] Ivo Babuska, Robert Lipton, and Michael Stuebner. The penetration function and its application to microscale problems. BIT Numerical Mathematics, 48(2):167–187, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=167.

daVeiga:2008:PPE

[2845] L. Beirão da Veiga, C. Chinosi, C. Lovadina, and R. Stenberg. A-priori and a-posteriori error analysis for a family of Reissner–Mindlin plate elements. BIT Numerical Mathematics, 48(2): 189–213, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=189.

Grimm:2008:RAT

[2846] V. Grimm and M. Hochbruck. Rational approximation to trigonometric operators. BIT Numerical Mathematics, 48(2):215–229, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=48&
                                        issue=2&spage=215.

Hairer:2008:ABL

[2847] E. Hairer, R. I. McLachlan, and A. Razakarivony. Achieving Brouwer’s law with implicit Runge–Kutta methods. BIT Numerical Mathematics, 48(2):231–243, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=2&spage=231.

Hanke:2008:SSE

[2848] Martin Hanke, Nuutti Hyvönen, Manfred Lehn, and Stefanie Reusswig. Source supports in electrostatics. BIT Numerical Mathematics, 48(2):245– 264, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=2&spage=245.

Hegland:2008:SGH

[2849] Markus Hegland, Andreas Hellander, and Per Lötstedt. Sparse grids and hybrid methods for the chemical master equation. BIT Numerical Mathematics, 48(2):265– 283, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=48&issue=2&spage=265.

Hoffman:2008:BIE

[2850] Johan Hoffman and Claes Johnson. Blow up of incompressible Euler solutions. BIT Numerical Mathematics, 48(2):285–307, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=285.

Kellogg:2008:LCS

[2851] R. Bruce Kellogg and Martin Stynes.

Layers and corner singularities in singularly perturbed elliptic problems. BIT Numerical Mathematics, 48(2): 309–314, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=309.

Pedas:2008:WCP

[2852] A. Pedas and G. Vainikko. What is the complexity of periodic weakly singular integral equations? BIT Numerical Mathematics, 48(2):315– 335, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=315.

Pohjanpelto:2008:RED

[2853] J. Pohjanpelto. Reduction of exterior differential systems with infinite dimensional symmetry groups. BIT Numerical Mathematics, 48(2):337– 355, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=337.

deSouza:2008:AAM

[2854] C. de Souza, F. Béchet, D. Leguillon, and É. Sanchez-Palencia. Anisotropic adaptive mesh procedure for computing very thin hyperbolic shells. BIT Numerical Mathematics, 48(2):357– 387, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=357.

Terrel:2008:AFD

[2855] A. R. Terrel, L. R. Scott, M. G. Knepley, and R. C. Kirby. Automated FEM discretizations for the Stokes equation. BIT Numerical Mathematics, 48(2): 389–404, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=389.

Tuomela:2008:KAM

[2856] J. Tuomela. Kinematic analysis of multibody systems. BIT Numerical Mathematics, 48(2):405–421, June 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=2&spage=405.

Ruhe:2008:ICIb

[2857] Axel Ruhe. Introduction to the contents of issue 48:3. BIT Numerical Mathematics, 48(3):423–424, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=423.

Asheim:2008:CFA

[2858] A. Asheim. A combined Filon/ asymptotic quadrature method for highly oscillatory problems. BIT Numerical Mathematics, 48(3):425–448, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=425.

Bialecki:2008:OSO

[2859] B. Bialecki, G. Fairweather, A. Karageorghis, and Q. N. Nguyen. Optimal superconvergent one step quadratic spline collocation methods. BIT Numerical Mathematics, 48(3):449–472, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=449.

Calabro:2008:ERQ

[2860] F. Calabrò and A. Corbo Esposito. An efficient and reliable quadrature algorithm for integration with respect to binomial measures. BIT Numerical Mathematics, 48(3):473–491, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=473.

Chen:2008:TPB

[2861] Xiao Shan Chen. Two perturbation bounds for singular values and eigenvalues. BIT Numerical Mathematics, 48(3):493–497, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=493.

Costabile:2008:ERK

[2862] F. Costabile and A. Napoli. Economical Runge–Kutta methods for numerical solution of stochastic differential equations. BIT Numerical Mathematics, 48(3):499–509, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=499.

Du:2008:IGS

[2863] Xiuhong Du and Daniel B. Szyld. Inexact GMRES for singular linear systems. BIT Numerical Mathematics, 48(3):511–531, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=48&
issue=3&spage=511.

Gao:2008:TSF

[2864] X. Gao, F. Liu, and A. Zhou.

Three-scale finite element eigenvalue discretizations. BIT Numerical Mathematics, 48(3):533–562, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=533.

Kaagstrom:2008:BAR

[2865] B. Kågström, D. Kressner, E. S.

Quintana-Ortí, and G. Quintana-Ortí. Blocked algorithms for the reduction to Hessenberg-triangular form revisited. BIT Numerical Mathematics, 48(3):563–584, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=563.

Leykekhman:2008:PEE

[2866] D. Leykekhman and L. B. Wahlbin.

A posteriori error estimates by recovered gradients in parabolic finite element equations. BIT Numerical Mathematics, 48(3):585–605, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=585.

Marcinkowski:2008:NNA

[2867] L. Marcinkowski and T. Rahman.

Neumann–Neumann algorithms for a mortar Crouzeix–Raviart element for 2nd order elliptic problems. BIT Numerical Mathematics, 48(3):607–626, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=607.

Reichel:2008:NZF

[2868] Lothar Reichel and Andriy Shyshkov. A new zero-finder for Tikhonov regularization. BIT Numerical Mathematics, 48(3):627–643, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=3&spage=627.

Bjorck:2008:E

[2869] Åke Björck. Editorial. BIT Numerical Mathematics, 48(4):645–647, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=645.

Ruhe:2008:ICIc

[2870] Axel Ruhe. Introduction to the contents of issue 48:4. BIT Numerical Mathematics, 48(4):649–650, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=649.

Bardsley:2008:SRN

[2871] Johnathan M. Bardsley. Stopping rules for a nonnegatively constrained iterative method for ill-posed Poisson imaging problems. BIT Numerical Mathematics, 48(4):651–664, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=651.

Costantini:2008:NSS

[2872] Paolo Costantini, Francesca Pelosi, and Maria L. Sampoli. New spline spaces with generalized tension properties. BIT Numerical Mathematics, 48(4):665–688, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=665.

Disibuyuk:2008:TPB

[2873] Çetin Disibüyük and Halil Oruç.

Tensor product q-Bernstein polynomials. BIT Numerical Mathematics, 48(4):689–700, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=48&
issue=4&spage=689.

Driscoll:2008:CSA

[2874] Tobin A. Driscoll, Folkmar Bornemann, and Lloyd N. Trefethen. The chebop system for automatic solution of differential equations. BIT Numerical Mathematics, 48(4):701–723, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=701.

Hoang:2008:ISS

[2875] N. S. Hoang and A. G. Ramm. An iterative scheme for solving nonlinear equations with monotone operators. BIT Numerical Mathematics, 48(4):725–741, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=725.

Khanamiryan:2008:QMH

[2876] M. Khanamiryan. Quadrature methods for highly oscillatory linear and nonlinear systems of ordinary differential equations: part I. BIT Numerical Mathematics, 48(4):743–761, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=743.

Luksan:2008:LMT

[2877] L. Luksan, C. Matonoha, and J. Vlcek. On Lagrange multipliers of trust-region subproblems. BIT Numerical Mathematics, 48(4):763–768, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=763.

Matejas:2008:RES

[2878] J. Matejas and V. Hari. Relative eigenvalue and singular value perturbations of scaled diagonally dominant matrices. BIT Numerical Mathematics, 48(4):769–781, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=48&issue=4&spage=769.

Ruhe:2009:PIC

[2879] Axel Ruhe. Preface and introduction to the contents of issue 49:1. BIT Numerical Mathematics, 49(1):1– 2, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=1.

Baboulin:2009:UDT

[2880] Marc Baboulin and Serge Gratton. Using dual techniques to derive componentwise and mixed condition numbers for a linear function of a linear least squares solution. BIT Numerical Mathematics, 49(1):3–19, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=3.

Cartis:2009:TRO

[2881] C. Cartis, N. I. M. Gould, and P. L. Toint. Trust-region and other regularisations of linear least-squares problems. BIT Numerical Mathematics, 49(1):21– 53, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=1&spage=21.

Marchi:2009:NCF

[2882] Stefano De Marchi, Marco Vianello, and Yuan Xu. New cubature formulae and hyperinterpolation in three variables. BIT Numerical Mathematics, 49(1):55–73, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=49&
                                        issue=1&spage=55.

Hakula:2009:CTD

[2883] H. Hakula and N. Hyvönen. On computation of test dipoles for factorization method. BIT Numerical Mathematics, 49(1):75–91, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=1&spage=75.

Hewitt:2009:ASG

[2884] L. L. Hewitt and A. T. Hill. Algebraically stable general linear methods and the G-matrix. BIT Numerical Mathematics, 49(1):93–111, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=93.

Jentzen:2009:PTS

[2885] Arnulf Jentzen and Peter E. Kloeden.

Pathwise Taylor schemes for random ordinary differential equations. BIT Numerical Mathematics, 49(1):113– 140, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=113.

Kitzhofer:2009:PFE

[2886] Georg Kitzhofer, Othmar Koch, and Ewa B. Weinmüller. Path following for essentially singular boundary value problems with application to the complex Ginzburg–Landau equation. BIT Numerical Mathematics, 49(1):141– 160, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=141.

Lamour:2009:CCI

[2887] René Lamour and Francesca Mazzia.

Computation of consistent initial values for properly stated index 3 DAEs. BIT Numerical Mathematics, 49(1): 161–175, March 2009. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=161.

McLachlan:2009:LPS

[2888] R. I. McLachlan, G. R. W. Quispel, and P. S. P. Tse. Linearizationpreserving self-adjoint and symplectic integrators. BIT Numerical Mathematics, 49(1):177–197, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=49&
issue=1&spage=177.

Neuhauser:2009:CSM

[2889] Christof Neuhauser and Mechthild Thalhammer. On the convergence of splitting methods for linear evolutionary Schrödinger equations involving an unbounded potential. BIT Numerical Mathematics, 49(1):199– 215, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=199.

Sickenberger:2009:LEE

[2890] Thorsten Sickenberger, Ewa Weinmüller, and Renate Winkler. Local error estimates for moderately smooth problems: Part II — SDEs and SDAEs with small noise. BIT Numerical Mathematics, 49(1):217–245, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=1&spage=217.

Ruhe:2009:ICIa

[2891] Axel Ruhe. Introduction to the contents of issue 49:2. BIT Numerical Mathematics, 49(2):247–248, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=247.

Adcock:2009:UMF

[2892] Ben Adcock. Univariate modified Fourier methods for second order boundary value problems. BIT Numerical Mathematics, 49(2):249– 280, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=249.

Bunck:2009:FAE

[2893] Benjamin F. Bunck. A fast algorithm for evaluation of normalized Hermite functions. BIT Numerical Mathematics, 49(2):281–295, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=281.

Emmrich:2009:CVT

[2894] Etienne Emmrich. Convergence of the variable two-step BDF time discretisation of nonlinear evolution problems governed by a monotone potential operator. BIT Numerical Mathematics, 49(2):297–323, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=297.

Ezquerro:2009:NIO

[2895] J. A. Ezquerro and M. A. Hernández. New iterations of R-order four with reduced computational cost. BIT Numerical Mathematics, 49(2):325– 342, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=325.

Geissert:2009:RWC

[2896] Matthias Geissert, Mihály Kovács, and Stig Larsson. Rate of weak convergence of the finite element method for the stochastic heat equation with additive noise. BIT Numerical Mathematics, 49(2):343– 356, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=343.

Jimenez:2009:RCL

[2897] Juan Carlos Jimenez and Felix Carbonell. Rate of convergence of Local Linearization schemes for random differential equations. BIT Numerical Mathematics, 49(2):357– 373, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=357.

Koley:2009:HOF

[2898] Ujjwal Koley, Siddhartha Mishra, Nils Henrik Risebro, and Magnus Svärd. Higher order finite difference schemes for the magnetic induction equations. BIT Numerical Mathematics, 49(2):375–395, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=375.

Lyness:2009:AEO

[2899] James N. Lyness and James W. Lottes.

Asymptotic expansions for oscillatory integrals using inverse functions. BIT Numerical Mathematics, 49(2):397– 417, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=397.

Rump:2009:CPS

[2900] Siegfried M. Rump, Paul Zimmermann, Sylvie Boldo, and Guillaume Melquiond. Computing predecessor and successor in rounding to nearest. BIT Numerical Mathematics, 49(2): 419–431, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=2&spage=419.

Vomel:2009:TFB

[2901] Christof Vömel and Jason Slemons. Twisted factorization of a banded matrix. BIT Numerical Mathematics, 49(2):433–447, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=49&
                                        issue=2&spage=433.

Wensch:2009:MIS

[2902] Jörg Wensch, Oswald Knoth, and Alexander Galant. Multirate infinitesimal step methods for atmospheric flow simulation. BIT Numerical Mathematics, 49(2):449–473, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=2&spage=449.

Ruhe:2009:ICIb

[2903] Axel Ruhe. Introduction to the contents of issue 49:3. BIT Numerical Mathematics, 49(3):475–476, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=3&spage=475.

Campos:2009:QFL

[2904] Rafael G. Campos and Francisco Mejía Díaz. Quadrature formulas for the Laplace and Mellin transforms. BIT Numerical Mathematics, 49(3):477– 486, September 2009. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=3&spage=477.

Castella:2009:SMC

[2905] F. Castella, P. Chartier, S. Descombes, and G. Vilmart. Splitting methods with complex times for parabolic equations. BIT Numerical Mathematics, 49(3):487–508, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=3&spage=487.

Du:2009:MDA

[2906] Kui Du, Graeme Fairweather, Que N.

Nguyen, and Weiwei Sun. Matrix decomposition algorithms for the C^{0}-quadratic finite element Galerkin method. BIT Numerical Mathematics, 49(3):509–526, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=49&
issue=3&spage=509.

Hansen:2009:HOS

[2907] Eskil Hansen and Alexander Ostermann. High order splitting methods for analytic semigroups exist. BIT Numerical Mathematics, 49(3):527–542, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=3&spage=527.

Kuo:2009:PSI

[2908] Frances Y. Kuo, Grzegorz W. Wasilkowski, and Henryk Woźniakowski. On the power of standard information for L_{∞} approximation in the randomized setting. BIT Numerical Mathematics, 49(3):543–564, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=3&spage=543.

Lamberti:2009:NIB

[2909] Paola Lamberti. Numerical integration based on bivariate quadratic spline quasi-interpolants on bounded domains. BIT Numerical Mathematics, 49(3):565–588, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=49&
                                        issue=3&spage=565.

Lamnii:2009:MMF

[2910] A. Lamnii, H. Mraoui, D. Sbibih, and A. Zidna. A multiresolution method for fitting scattered data on the sphere. BIT Numerical Mathematics, 49(3):589–610, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=3&spage=589.

Mazzia:2009:BCH

[2911] Francesca Mazzia and Alessandra Sestini. The BS class of Hermite spline quasi-interpolants on nonuniform knot distributions. BIT Numerical Mathematics, 49(3):611–628, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=3&spage=611.

Noschese:2009:SDN

[2912] Silvia Noschese and Lothar Reichel.

The structured distance to normality of banded Toeplitz matrices. BIT Numerical Mathematics, 49(3):629–640, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=3&spage=629.

Ruhe:2009:CV

[2913] Axel Ruhe. Collaborators for volume 49. BIT Numerical Mathematics, 49(4):641–643, December 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=49&
issue=4&spage=641.

Christiansen:2009:CLG

[2914] Snorre H. Christiansen and Tore Gunnar Halvorsen. Convergence of lattice gauge theory for Maxwell’s equations. BIT Numerical Mathematics, 49(4):645–667, December 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=

                                        article&issn=0006-3835&volume=49&
                                        issue=4&spage=645.

Hnetynkova:2009:REG

[2915] Iveta Hnetynková, Martin Plesinger, and Zdenek Strakos. The regularizing effect of the Golub–Kahan iterative bidiagonalization and revealing the noise level in the data. BIT Numerical Mathematics, 49(4):669–696, December 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=4&spage=669.

Izadi:2009:PEE

[2916] Mohammad Izadi. A posteriori error estimates for the coupling equations of scalar conservation laws. BIT Numerical Mathematics, 49(4):697–720, December 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=4&spage=697.

Pachon:2009:BRA

[2917] Ricardo Pachón and Lloyd N. Trefethen. Barycentric–Remez algorithms for best polynomial approximation in the chebfun system. BIT Numerical Mathematics, 49(4):721–741, December 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=49&issue=4&spage=721.

Teofanov:2009:SAF

[2918] Ljiljana Teofanov and Helena Zarin.

Superconvergence analysis of a finite element method for a two-parameter singularly perturbed problem. BIT Numerical Mathematics, 49(4):743–765, December 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=49&issue=4&spage=743.

Ruhe:2010:ICIa

[2919] Axel Ruhe. Introduction to the contents of issue 50:1. BIT Numerical Mathematics, 50(1):1–2, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=50&
issue=1&spage=1.

Calvo:2010:HOS

[2920] M. Calvo, J. M. Franco, J. I. Montijano, and L. Rández. On high order symmetric and symplectic trigonometrically fitted Runge–Kutta methods with an even number of stages. BIT Numerical Mathematics, 50(1):3– 21, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=3.

Chartier:2010:RFM

[2921] P. Chartier, E. Darrigrand, and E. Faou. A regular fast multipole method for geometric numerical integrations of Hamiltonian systems.

BIT Numerical Mathematics, 50(1):23– 40, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=23.

Chen:2010:PBP

[2922] Xiao Shan Chen. Perturbation bounds for the periodic Schur decomposition. BIT Numerical Mathematics, 50(1):41– 58, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=41.

Lechleiter:2010:IPD

[2923] Armin Lechleiter. Imaging of periodic dielectrics. BIT Numerical Mathematics, 50(1):59–83, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=59.

Mishra:2010:SNS

[2924] Siddhartha Mishra and Magnus Svärd. On stability of numerical schemes via frozen coefficients and the magnetic induction equations. BIT Numerical Mathematics, 50(1):85–108, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=85.

Nakatsukasa:2010:PBM

[2925] Yuji Nakatsukasa. Perturbation behavior of a multiple eigenvalue in generalized Hermitian eigenvalue problems. BIT Numerical Mathematics, 50(1):109–121, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=109.

Notaris:2010:ENG

[2926] Sotirios E. Notaris. The error norm of Gauss–Radau quadrature formulae for Chebyshev weight functions. BIT Numerical Mathematics, 50(1):123– 147, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=123.

Olver:2010:FNS

[2927] Sheehan Olver. Fast, numerically stable computation of oscillatory integrals with stationary points. BIT Numerical Mathematics, 50(1):149– 171, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=149.

Simonsson:2010:GAL

[2928] Lennart Simonsson and Lars Eldén.

Grassmann algorithms for low rank approximation of matrices with missing values. BIT Numerical Mathematics, 50(1):173–191, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=173.

Wu:2010:ECC

[2929] Jiming Wu, Xiaoping Zhang, and Dongjie Liu. An efficient calculation of the Clausen functions Cl_{n}(θ), (n ≥ 2). BIT Numerical Mathematics, 50(1):193–206, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=50&
issue=1&spage=193.

Zhang:2010:NPG

[2930] Naimin Zhang. A note on preconditioned GMRES for solving singular linear systems. BIT Numerical Mathematics, 50(1):207–220, March 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=1&spage=207.

Ruhe:2010:ICIb

[2931] Axel Ruhe. Introduction to the contents of issue 50:2. BIT Numerical Mathematics, 50(2):221–222, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=221.

Calvo:2010:PMP

[2932] M. Calvo, M. P. Laburta, J. I. Montijano, and L. Rández. Projection methods preserving Lyapunov functions. BIT Numerical Mathematics, 50(2): 223–241, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=223.

Coman:2010:IOT

[2933] Gheorghe Coman and Teodora Catinas. Interpolation operators on a triangle with one curved side. BIT Numerical Mathematics, 50(2):243– 267, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=243.

Conti:2010:ACB

[2934] Costanza Conti and Lucia Romani. Affine combination of Bspline subdivision masks and its non-stationary counterparts. BIT Numerical Mathematics, 50(2):269– 299, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=269.

Gusev:2010:NES

[2935] Sergei Gusev, Stefan Johansson, Bo Kågström, Anton Shiriaev, and Andras Varga. A numerical evaluation of solvers for the periodic Riccati differential equation. BIT Numerical Mathematics, 50(2):301– 329, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=301.

Holden:2010:FTM

[2936] Helge Holden, Nils Henrik Risebro, and Hilde Sande. Front tracking for a model of immiscible gas flow with large data. BIT Numerical Mathematics, 50(2): 331–376, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=331.

Li:2010:FPI

[2937] Chong-Jun Li, Vittoria Demichelis, and Catterina Dagnino. Finite-part integrals over polygons by an 8-node quadrilateral spline finite element. BIT Numerical Mathematics, 50(2):377– 394, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=377.

Needell:2010:RKS

[2938] Deanna Needell. Randomized Kaczmarz solver for noisy linear systems. BIT Numerical Mathematics, 50(2): 395–403, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=395.

Skelboe:2010:APT

[2939] Stig Skelboe. Adaptive partitioning techniques for index 1 IDEs. BIT Numerical Mathematics, 50(2):405– 423, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=405.

Sun:2010:SNS

[2940] Zhe Sun, Jinping Zeng, and Donghui Li. Semismooth Newton Schwarz iterative methods for the linear complementarity problem. BIT Numerical Mathematics, 50(2):425–449, June 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=2&spage=425.

Ruhe:2010:Ea

[2941] Axel Ruhe. Editorial. BIT Numerical Mathematics, 50(3):451–453, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=3&spage=451.

Auzinger:2010:SES

[2942] Winfried Auzinger and Felix Kramer.

On the stability and error structure of BDF schemes applied to linear parabolic evolution equations. BIT Numerical Mathematics, 50(3):455–480, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=3&spage=455.

Bazan:2010:GFA

[2943] Fermín S. Viloche Bazán and Leonardo S. Borges. GKB-FP: an algorithm for large-scale discrete ill-posed problems. BIT Numerical Mathematics, 50(3):481–507, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=50&
issue=3&spage=481.

DelaCruzCancino:2010:HOL

[2944] H. De la Cruz Cancino, R. J. Biscay, J. C. Jimenez, F. Carbonell, and T. Ozaki. High order local linearization methods: an approach for constructing A-stable explicit schemes for stochastic differential equations with additive noise. BIT Numerical Mathematics, 50(3):509–539, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=3&spage=509.

Debrabant:2010:RKM

[2945] Kristian Debrabant. Runge–Kutta methods for third order weak approximation of SDEs with multidimensional additive noise. BIT Numerical Mathematics, 50(3):541–558, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=3&spage=541.

Donatelli:2010:FTH

[2946] Marco Donatelli. Fast transforms for high order boundary conditions in deconvolution problems. BIT Numerical Mathematics, 50(3):559–576, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=3&spage=559.

Floater:2010:CRM

[2947] Michael S. Floater. A chain rule for multivariate divided differences. BIT Numerical Mathematics, 50(3):577–586, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=50&
issue=3&spage=577.

Kraft:2010:DWR

[2948] Karin Kraft and Stig Larsson. The dual weighted residuals approach to optimal control of ordinary differential equations. BIT Numerical Mathematics, 50(3):587–607, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=3&spage=587.

Ku:2010:PEE

[2949] Jaeun Ku. Pointwise error estimates for a stabilized Galerkin method: non-selfadjoint problems. BIT Numerical Mathematics, 50(3):609–630, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=3&spage=609.

Lopez-Fernandez:2010:QBM

[2950] María López-Fernández. A quadrature based method for evaluating exponential-type functions for exponential methods. BIT Numerical Mathematics, 50(3):631–655, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=3&spage=631.

Nagamine:2010:CMS

[2951] André Nagamine and José Alberto Cuminato. A collocation method for solving singular integrodifferential equations. BIT Numerical Mathematics, 50(3):657–688, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=3&spage=657.

Ruhe:2010:Eb

[2952] Axel Ruhe. Editorial. BIT Numerical Mathematics, 50(4):689–692, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=4&spage=689.

Brunner:2010:SCM

[2953] Hermann Brunner and Hui Liang.

Stability of collocation methods for delay differential equations with vanishing delays. BIT Numerical Mathematics, 50(4):693–711, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=4&spage=693.

Butcher:2010:TFS

[2954] J. C. Butcher and T. M. H.

Chan. The tree and forest spaces with applications to initial-value problem methods. BIT Numerical Mathematics, 50(4):713–728, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=4&spage=713.

Descombes:2010:ELE

[2955] Stéphane Descombes and Mechthild Thalhammer. An exact local error representation of exponential operator splitting methods for evolutionary problems and applications to linear Schrödinger equations in the semi-classical regime. BIT Numerical Mathematics, 50(4):729–749, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=4&spage=729.

Hoang:2010:DSM

[2956] N. S. Hoang. Dynamical Systems Method of gradient type for solving nonlinear equations with monotone operators. BIT Numerical Mathematics, 50(4):751–780, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=4&spage=751.

Holst:2010:CAF

[2957] Michael J. Holst, Mats G. Larson, Axel Målqvist, and Robert Söderlund. Convergence analysis of finite element approximations of the Joule heating problem in three spatial dimensions. BIT Numerical Mathematics, 50(4):781–795, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=4&spage=781.

Jahnke:2010:ESD

[2958] Tobias Jahnke and Derya Altintan. Efficient simulation of discrete stochastic reaction systems with a splitting method. BIT Numerical Mathematics, 50(4):797–822, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=50&issue=4&spage=797.

Li:2010:CSS

[2959] Jian Li, Lihua Shen, and Zhangxin Chen. Convergence and stability of a stabilized finite volume method for the stationary Navier– Stokes equations. BIT Numerical Mathematics, 50(4):823–842, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=4&spage=823.

Sablonniere:2010:EEE

[2960] P. Sablonnière, D. Sbibih, and M. Tahrichi. Error estimate and extrapolation of a quadrature formula derived from a quartic spline quasi-interpolant. BIT Numerical Mathematics, 50(4):843–862, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=50&issue=4&spage=843.

Shontz:2010:AWE

[2961] Suzanne M. Shontz and Stephen A.

Vavasis. Analysis of and workarounds for element reversal for a finite elementbased algorithm for warping triangular and tetrahedral meshes. BIT Numerical Mathematics, 50(4):863–884, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=4&spage=863.

Stuber:2010:NET

[2962] M. D. Stuber, V. Kumar, and P. I. Barton. Nonsmooth exclusion test for finding all solutions of nonlinear equations. BIT Numerical Mathematics, 50(4):885–917, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=50&issue=4&spage=885.

Ruhe:2011:B

[2963] Axel Ruhe. BIT50. BIT Numerical Mathematics, 51(1):1–5, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=51&
                                        issue=1&spage=1.

Adcock:2011:GPR

[2964] Ben Adcock. Gibbs phenomenon and its removal for a class of orthogonal expansions. BIT Numerical Mathematics, 51(1):7–41, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=1&spage=7.

Auzinger:2011:EAC

[2965] Winfried Auzinger, Herbert Lehner, and Ewa Weinmüller. An efficient asymptotically correct error estimator for collocation solutions to singular index-1 DAEs. BIT Numerical Mathematics, 51(1):43–65, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=1&spage=43.

Barnett:2011:NIR

[2966] Alex Barnett and Leslie Greengard. A new integral representation for quasi-periodic scattering problems in two dimensions. BIT Numerical Mathematics, 51(1):67–90, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=1&spage=67.

Cohen:2011:LEP

[2967] David Cohen and Ernst Hairer.

Linear energy-preserving integrators for Poisson systems. BIT Numerical Mathematics, 51(1):91–101, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=1&spage=91.

Gautschi:2011:OSO

[2968] Walter Gautschi. Optimally scaled and optimally conditioned Vandermonde and Vandermonde-like matrices. BIT Numerical Mathematics, 51(1):103– 125, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=1&spage=103.

Gleich:2011:SCT

[2969] David F. Gleich, Ying Wang, Xiangrui Meng, Farnaz Ronaghi, Margot Gerritsen, and Amin Saberi. Some computational tools for digital archive and metadata maintenance. BIT Numerical Mathematics, 51(1):127– 154, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=1&spage=127.

Gonzalez-Pinto:2011:GEE

[2970] S. González-Pinto and D. Hernández-Abreu. Global error estimates for a uniparametric family of stiffly accurate Runge–Kutta collocation methods on singularly perturbed problems. BIT Numerical Mathematics, 51(1):155– 175, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=1&spage=155.

Higham:2011:HSA

[2971] Desmond J. Higham, Somkid Intep, Xuerong Mao, and Lukasz Szpruch. Hybrid simulation of autoregulation within transcription and translation. BIT Numerical Mathematics, 51(1):177–196, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=

article&issn=0006-3835&volume=51&
issue=1&spage=177.

Hochstenbach:2011:FTR

[2972] Michiel E. Hochstenbach and Lothar Reichel. Fractional Tikhonov regularization for linear discrete illposed problems. BIT Numerical Mathematics, 51(1):197–215, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=51&
issue=1&spage=197.

Ramage:2011:MPA

[2973] Alison Ramage and Lina von Sydow.

A multigrid preconditioner for an adaptive Black–Scholes solver. BIT Numerical Mathematics, 51(1):217– 233, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=1&spage=217.

Reich:2011:DSF

[2974] Sebastian Reich. A dynamical systems framework for intermittent data assimilation. BIT Numerical Mathematics, 51(1):235–249, March 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=1&spage=235.

Ruhe:2011:ICIa

[2975] Axel Ruhe. Introduction to the contents of issue 51:2. BIT Numerical Mathematics, 51(2):251–252, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=251.

Connors:2011:PTD

[2976] Jeffrey M. Connors and Attou Miloua. Partitioned time discretization for parallel solution of coupled ODE systems. BIT Numerical Mathematics, 51(2): 253–273, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=253.

Cools:2011:CLU

[2977] Ronald Cools and Koen Poppe. Chebyshev lattices, a unifying framework for cubature with Chebyshev weight function. BIT Numerical Mathematics, 51(2):275–288, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=275.

Mukherjee:2011:OEE

[2978] Kaushik Mukherjee and Srinivasan Natesan. Optimal error estimate of upwind scheme on Shishkintype meshes for singularly perturbed parabolic problems with discontinuous convection coefficients. BIT Numerical Mathematics, 51(2):289– 315, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=289.

ORiordan:2011:PUN

[2979] E. O’Riordan and J. Quinn. Parameteruniform numerical methods for some linear and nonlinear singularly perturbed convection diffusion boundary turning point problems. BIT Numerical Mathematics, 51(2):317– 337, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=317.

Okayama:2011:ISC

[2980] Tomoaki Okayama, Takayasu Matsuo, and Masaaki Sugihara. Improvement of a Sinc-collocation method for Fredholm integral equations of the second kind. BIT Numerical Mathematics, 51(2): 339–366, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=339.

Rump:2011:VBS

[2981] Siegfried M. Rump. Verified bounds for singular values, in particular for the spectral norm of a matrix and its inverse. BIT Numerical Mathematics, 51(2):367–384, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=2&spage=367.

Scott:2011:ISI

[2982] Jennifer Scott and Miroslav Tuma. The importance of structure in incomplete factorization preconditioners. BIT Numerical Mathematics, 51(2): 385–404, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=2&spage=385.

Szpruch:2011:NSS

[2983] Lukasz Szpruch, Xuerong Mao, Desmond J. Higham, and Jiazhu Pan. Numerical simulation of a strongly nonlinear Ait–Sahalia-type interest rate model. BIT Numerical Mathematics, 51(2): 405–425, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=2&spage=405.

Verwer:2011:CSS

[2984] J. G. Verwer. Component splitting for semi-discrete Maxwell equations. BIT Numerical Mathematics, 51(2): 427–445, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=2&spage=427.

Whidborne:2011:CMT

[2985] James F. Whidborne and Nathalie Amar. Computing the maximum transient energy growth. BIT Numerical Mathematics, 51(2):447–457, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=447.

Zhang:2011:TAN

[2986] Qinghui Zhang. Theoretical analysis of numerical integration in Galerkin meshless methods. BIT Numerical Mathematics, 51(2):459– 480, June 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=2&spage=459.

Ruhe:2011:ICIb

[2987] Axel Ruhe. Introduction to the contents of issue 51:3. BIT Numerical Mathematics, 51(3):481–482, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=3&spage=481.

Banjai:2011:EAR

[2988] Lehel Banjai and Christian Lubich.

An error analysis of Runge–Kutta convolution quadrature. BIT Numerical Mathematics, 51(3):483–496, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=3&spage=483.

Borodachov:2011:OMS

[2989] Sergiy Borodachov and Tatyana Sorokina. An optimal multivariate spline method of recovery of twice differentiable functions. BIT Numerical Mathematics, 51(3):497–511, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=3&spage=497.

Bockmann:2011:BMI

[2990] Christine Böckmann and Athassawat Kammanee. Broyden method for inverse non-symmetric Sturm– Liouville problems. BIT Numerical Mathematics, 51(3):513–528, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=3&spage=513.

Debrabant:2011:BSA

[2991] Kristian Debrabant and Anne Kværnø. B-series analysis of iterated Taylor methods. BIT Numerical Mathematics, 51(3):529–553, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=51&
                                        issue=3&spage=529.

Dieci:2011:DED

[2992] Luca Dieci, Cinzia Elia, and Erik Van Vleck. Detecting exponential dichotomy on the real line: SVD and QR algorithms. BIT Numerical Mathematics, 51(3):555–579, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=3&spage=555.

Emmrich:2011:TDM

[2993] Etienne Emmrich. Time discretisation of monotone nonlinear evolution problems by the discontinuous Galerkin method. BIT Numerical Mathematics, 51(3):581–607, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=3&spage=581.

Feng:2011:NSC

[2994] Hui Feng, Xiaoping Zhang, and Jin Li. Numerical solution of a certain hypersingular integral equation of the first kind. BIT Numerical Mathematics, 51(3):609–630, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=3&spage=609.

Franz:2011:SNS

[2995] Sebastian Franz. SDFEM with non-standard higher-order finite elements for a convection-diffusion problem with characteristic boundary layers. BIT Numerical Mathematics, 51(3):631–651, September 2011. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=51&
                                        issue=3&spage=631.

Johnson:2011:CAA

[2996] Tomas Johnson and Warwick Tucker. On a computer-aided approach to the computation of Abelian integrals. BIT Numerical Mathematics, 51(3):653–667, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=51&
                                        issue=3&spage=653.

Klann:2011:WBM

[2997] E. Klann, R. Ramlau, and L. Reichel. Wavelet-based multilevel methods for linear ill-posed problems. BIT Numerical Mathematics, 51(3):669–694, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=3&spage=669.

Noutsos:2011:BPA

[2998] D. Noutsos and P. Vassalos. Band plus algebra preconditioners for twolevel Toeplitz systems. BIT Numerical Mathematics, 51(3):695–719, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=3&spage=695.

Popovic:2011:AFM

[2999] Jelena Popovic and Olof Runborg.

Analysis of a fast method for solving the high frequency Helmholtz equation in one dimension. BIT Numerical Mathematics, 51(3):721–755, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=3&spage=721.

Remogna:2011:QIO

[3000] Sara Remogna. Quasi-interpolation operators based on the trivariate seven-direction C^{2} quartic box spline. BIT Numerical Mathematics, 51(3):757–776, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=51&
issue=3&spage=757.

Ruhe:2011:E

[3001] Axel Ruhe. Editorial. BIT Numerical Mathematics, 51(4):777–780, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=777.

Beccari:2011:NUI

[3002] Carolina Vittoria Beccari, Giulio Casciola, and Lucia Romani. Nonuniform interpolatory curve subdivision with edge parameters built upon compactly supported fundamental splines. BIT Numerical Mathematics, 51(4):781–808, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=781.

Cifani:2011:DGM

[3003] Simone Cifani, Espen R. Jakobsen, and Kenneth H. Karlsen. The discontinuous Galerkin method for fractional degenerate convection-diffusion equations. BIT Numerical Mathematics, 51(4):809–844, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=51&
issue=4&spage=809.

Cumming:2011:MCC

[3004] Ben Cumming, Tim Moroney, and Ian Turner. A mass-conservative control volume-finite element method for solving Richards’ equation in heterogeneous porous media. BIT Numerical Mathematics, 51(4):845–864, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=845.

He:2011:ALB

[3005] Xin He, Maya Neytcheva, and Stefano Serra Capizzano. On an augmented Lagrangian-based preconditioning of Oseen type problems. BIT Numerical Mathematics, 51(4):865–888, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=51&
issue=4&spage=865.

Hochbruck:2011:EMM

[3006] Marlis Hochbruck and Alexander Ostermann. Exponential multistep methods of Adams-type. BIT Numerical Mathematics, 51(4):889–908, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=889.

Hundsdorfer:2011:SBP

[3007] W. Hundsdorfer, A. Mozartova, and M. N. Spijker. Special boundedness properties in numerical initial value problems. BIT Numerical Mathematics, 51(4):909–936, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=909.

Jarlebring:2011:ACF

[3008] Elias Jarlebring and Wim Michiels.

Analyzing the convergence factor of residual inverse iteration. BIT Numerical Mathematics, 51(4):937–957, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=4&spage=937.

Mikkelsen:2011:RRC

[3009] Carl Christian Kjelgaard Mikkelsen. Retracing the residual curve of a Lyapunov equation solver. BIT Numerical Mathematics, 51(4):959–975, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=4&spage=959.

Modin:2011:GIH

[3010] Klas Modin and Gustaf Söderlind. Geometric integration of Hamiltonian systems perturbed by Rayleigh damping. BIT Numerical Mathematics, 51(4):977–1007, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=4&spage=977.

Novakovic:2011:GBH

[3011] Vedran Novaković and Sanja Singer. A GPU-based hyperbolic SVD algorithm. BIT Numerical Mathematics, 51(4):1009–1030, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=51&issue=4&spage=1009.

Smith:2011:NEE

[3012] H. V. Smith and D. B. Hunter. The numerical evaluation of the error term in a quadrature formula of Clenshaw–Curtis type for the Gegenbauer weight function. BIT Numerical Mathematics, 51(4):1031–1038, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=1031.

VanDeun:2011:RIC

[3013] Joris Van Deun and Lloyd N. Trefethen. A robust implementation of the Carathéodory–Fejér method for rational approximation. BIT Numerical Mathematics, 51(4):1039–1050, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/content/
ag2514840142707r/;   http://www.
springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=51&
issue=4&spage=1039.

Xie:2011:NBI

[3014] Dexuan Xie and L. Ridgway Scott.

A new boundary integral equation for molecular electrostatics with charges over whole space. BIT Numerical Mathematics, 51(4):1051–1071, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=51&issue=4&spage=1051.

Ruhe:2012:PB

[3015] Axel Ruhe. Preface to BIT 52:1.

BIT Numerical Mathematics, 52(1):1– 2, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=1&spage=1.

Auzinger:2012:CRM

[3016] Winfried Auzinger and Magdalena Lapińska. Convergence of rational multistep methods of Adams–Padé type. BIT Numerical Mathematics, 52(1):3– 20, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=1&spage=3.

Gobbi:2012:CSA

[3017] Romina Gobbi and Renato Spigler. Comparing Shannon to autocorrelationbased wavelets for solving singularly perturbed elliptic BV problems. BIT Numerical Mathematics, 52(1):21–43, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=1&spage=21.

Hanke:2012:CSS

[3018] Martin Hanke, Lauri Harhanen, Nuutti Hyvönen, and Eva Schweickert. Convex source support in three dimensions. BIT Numerical Mathematics, 52(1):45– 63, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=1&spage=45.

Hannukainen:2012:CPM

[3019] Antti Hannukainen. Continuous preconditioners for the mixed Poisson problem. BIT Numerical Mathematics, 52(1):65–83, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=1&spage=65.

Kovacs:2012:WCF

[3020] Mihály Kovács, Stig Larsson, and Fredrik Lindgren. Weak convergence of finite element approximations of linear stochastic evolution equations with additive noise. BIT Numerical Mathematics, 52(1):85–108, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=1&spage=85.

Kruse:2012:CBS

[3021] Raphael Kruse. Characterization of bistability for stochastic multistep methods. BIT Numerical Mathematics, 52(1):109–140, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=1&spage=109.

Milstein:2012:SDP

[3022] G. N. Milstein and M. V. Tretyakov.

Solving the Dirichlet problem for Navier–Stokes equations by probabilistic approach. BIT Numerical Mathematics, 52(1):141–153, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=1&spage=141.

Peramaki:2012:CNS

[3023] Allan Perämäki. Convergence of a numerical solver for an R-linear Beltrami equation. BIT Numerical Mathematics, 52(1):155–178, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=1&spage=155.

Popova:2012:EDP

[3024] Evgenija D. Popova. Explicit description of 2D parametric solution sets. BIT Numerical Mathematics, 52 (1):179–200, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=1&spage=179.

Rump:2012:EEF

[3025] Siegfried M. Rump. Error estimation of floating-point summation and dot product. BIT Numerical Mathematics, 52(1):201–220, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=1&spage=201.

Sbibih:2012:NTB

[3026] D. Sbibih, A. Serghini, and A. Tijini. Normalized trivariate Bsplines on Worsey–Piper split and quasi-interpolants. BIT Numerical Mathematics, 52(1):221–249, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=1&spage=221.

Wen:2012:NSO

[3027] Liping Wen and Xiong Liu. Numerical stability of one-leg methods for neutral delay differential equations. BIT Numerical Mathematics, 52(1):251– 269, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=1&spage=251.

Ruhe:2012:ICIa

[3028] Axel Ruhe. Introduction to the contents of issue 52:2. BIT Numerical Mathematics, 52(2):271–272, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=271.

Abdi:2012:OBS

[3029] A. Abdi and J. C. Butcher. Order bounds for second derivative approximations. BIT Numerical Mathematics, 52(2):273–281, June 2012. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=2&spage=273.

Dangerfield:2012:BPN

[3030] C. E. Dangerfield, D. Kay, S. MacNamara, and K. Burrage. A boundary preserving numerical algorithm for the Wright–Fisher model with mutation. BIT Numerical Mathematics, 52(2): 283–304, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=283.

Donatelli:2012:MMT

[3031] Marco Donatelli, Stefano Serra-Capizzano, and Debora Sesana. Multigrid methods for Toeplitz linear systems with different size reduction. BIT Numerical Mathematics, 52(2):305– 327, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=305.

Jensen:2012:IOF

[3032] T. L. Jensen, J. H. Jørgensen, P. C. Hansen, and S. H. Jensen. Implementation of an optimal firstorder method for strongly convex total variation regularization. BIT Numerical Mathematics, 52(2):329– 356, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=329.

Jimenez:2012:CRS

[3033] J. C. Jimenez and H. de la Cruz Cancino. Convergence rate of strong Local Linearization schemes for stochastic differential equations with additive noise. BIT Numerical Mathematics, 52(2):357–382, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=357.

Khanamiryan:2012:QMH

[3034] Marianna Khanamiryan. Quadrature methods for highly oscillatory linear and non-linear systems of ordinary differential equations: part II. BIT Numerical Mathematics, 52(2):383– 405, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=383.

Klein:2012:LBR

[3035] Georges Klein and Jean-Paul Berrut.

Linear barycentric rational quadrature. BIT Numerical Mathematics, 52(2): 407–424, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=407.

Kuo:2012:GHQ

[3036] Frances Y. Kuo and Henryk Woźniakowski. Gauss–Hermite quadratures for functions from Hilbert spaces with Gaussian reproducing kernels. BIT Numerical Mathematics, 52(2):425–436, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=425.

Kupper:2012:RKM

[3037] Dominique Küpper, Anne Kværnø, and Andreas Rößler. A Runge– Kutta method for index 1 stochastic differential-algebraic equations with scalar noise. BIT Numerical Mathematics, 52(2):437–455, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=2&spage=437.

Narayan:2012:CCC

[3038] Akil Narayan and Jan S. Hesthaven. Computation of connection coefficients and measure modifications for orthogonal polynomials. BIT Numerical Mathematics, 52(2):457– 483, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=457.

Sadok:2012:NLC

[3039] Hassane Sadok and Daniel B. Szyld. A new look at CMRH and its relation to GMRES. BIT Numerical Mathematics, 52(2):485–501, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=485.

Smoktunowicz:2012:NAC

[3040] Alicja Smoktunowicz and Iwona Wróbel. Numerical aspects of computing the Moore–Penrose inverse of full column rank matrices. BIT Numerical Mathematics, 52(2):503– 524, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=503.

Xie:2012:IMP

[3041] Huiqing Xie. An iterative method for partial derivatives of eigenvectors of quadratic eigenvalue problems. BIT Numerical Mathematics, 52(2):525– 536, June 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=2&spage=525.

Ruhe:2012:ICIb

[3042] Axel Ruhe. Introduction to the contents of issue 52:3. BIT Numerical Mathematics, 52(3):537–538, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=3&spage=537.

Alcock:2012:SSO

[3043] Jamie Alcock and Kevin Burrage. Stable strong order 1.0 schemes for solving stochastic ordinary differential equations. BIT Numerical Mathematics, 52(3):539–557, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=3&spage=539.

Bartels:2012:LCS

[3044] Sören Bartels and Patrick Schreier. Local coarsening of simplicial finite element meshes generated by bisections. BIT Numerical Mathematics, 52(3):559–569, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0006-3835&volume=52&
                                        issue=3&spage=559.

Bergstrom:2012:ECG

[3045] Per Bergström, Ove Edlund, and Inge Söderkvist. Efficient computation of the Gauss–Newton direction when fitting NURBS using ODR. BIT Numerical Mathematics, 52(3):571–588, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0006-3835&
                                        volume=52&issue=3&spage=571.

Cioica:2012:AWM

[3046] Petru A. Cioica, Stephan Dahlke, Nicolas Döhring, Stefan Kinzel, Felix Lindner, et al. Adaptive wavelet methods for the stochastic Poisson equation. BIT Numerical Mathematics, 52(3):589–614, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=589.

Engquist:2012:EPC

[3047] B. Engquist, J. Häggblad, and O. Runborg. On energy preserving consistent boundary conditions for the Yee scheme in 2D. BIT Numerical Mathematics, 52(3):615–637, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=615.

Grimm:2012:RKS

[3048] Volker Grimm. Resolvent Krylov subspace approximation to operator functions. BIT Numerical Mathematics, 52(3):639–659, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=639.

Hiptmair:2012:SMS

[3049] Ralf Hiptmair and Shipeng Mao. Stable multilevel splittings of boundary edge element spaces. BIT Numerical Mathematics, 52(3):661–685, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=661.

Meurant:2012:GAP

[3050] Gérard Meurant. GMRES and the Arioli, Pták, and Strakos parametrization. BIT Numerical Mathematics, 52 (3):687–702, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=687.

Muntingh:2012:DDM

[3051] Georg Muntingh. Divided differences of multivariate implicit functions. BIT Numerical Mathematics, 52(3):703–723, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=52&
issue=3&spage=703.

Shakourifar:2012:SIC

[3052] Mohammad Shakourifar and Wayne Enright. Superconvergent interpolants for collocation methods applied to Volterra integro-differential equations with delay. BIT Numerical Mathematics, 52(3):725–740, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=725.

Wang:2012:FFI

[3053] Xiaojie Wang, Siqing Gan, and Desheng Wang. A family of fully implicit Milstein methods for stiff stochastic differential equations with multiplicative noise. BIT Numerical Mathematics, 52(3):741–772, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=741.

Wu:2012:ESM

[3054] Xinyuan Wu, Bin Wang, and Jianlin Xia. Explicit symplectic multidimensional exponential fitting modified Runge–Kutta–Nyström methods. BIT Numerical Mathematics, 52(3): 773–795, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=52&issue=3&spage=773.

Ruhe:2012:E

[3055] Axel Ruhe. Editorial. BIT Numerical Mathematics, 52(4):797–800, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0408-y;   http:
//link.springer.com/content/pdf/
10.1007/s10543-012-0408-y.

Boulton:2012:EEC

[3056] Lyonell Boulton and Michael Strauss. Eigenvalue enclosures and convergence for the linearized MHD operator. BIT Numerical Mathematics, 52(4):801–825, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-012-
                                        0389-x.

Bujanovic:2012:CTP

[3057] Zvonimir Bujanović and Zlatko Drmac. A contribution to the theory and practice of the block Kogbetliantz method for computing the SVD. BIT Numerical Mathematics, 52(4):827–849, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-012-0388-y.

Cai:2012:FNS

[3058] Haotao Cai. A fast numerical solution for the first kind boundary integral equation for the Helmholtz equation. BIT Numerical Mathematics, 52(4):851–875, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-012-
                                        0380-6;   http://link.springer.
                                        com/content/pdf/10.1007/s10543-
                                        012-0380-6.

Cohen:2012:OSE

[3059] David Cohen and Ludwig Gauckler. One-stage exponential integrators for nonlinear Schrödinger equations over long times. BIT Numerical Mathematics, 52(4):877–903, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0385-1.

Corti:2012:SFD

[3060] Paolo Corti and Siddhartha Mishra.

Stable finite difference schemes for the magnetic induction equation with Hall effect. BIT Numerical Mathematics, 52(4):905–932, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0383-3.

Effenberger:2012:CIN

[3061] Cedric Effenberger and Daniel Kressner. Chebyshev interpolation for nonlinear eigenvalue problems. BIT Numerical Mathematics, 52(4):933–951, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0381-5.

Han:2012:CCB

[3062] Xuli Han and Yuanpeng Zhu. Curve construction based on five trigonometric blending functions. BIT Numerical Mathematics, 52(4):953–979, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0386-0.

Heumann:2012:FDS

[3063] Holger Heumann, Ralf Hiptmair, Kun Li, and Jinchao Xu. Fully discrete semi-Lagrangian methods for advection of differential forms. BIT Numerical Mathematics, 52(4):981–1007, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-012-0382-4.

Mazure:2012:NCD

[3064] Marie-Laurence Mazure. On a new criterion to decide whether a spline space can be used for design. BIT Numerical Mathematics, 52(4):1009–1034, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-012-0390-4.

Rozloznik:2012:NSO

[3065] Miroslav Rozlozník, Miroslav Tuma, Alicja Smoktunowicz, and Jiří Kopal. Numerical stability of orthogonalization methods with a non-standard inner product. BIT Numerical Mathematics, 52(4):1035–1058, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-012-0398-9.

Rump:2012:IAF

[3066] Siegfried M. Rump. Interval arithmetic over finitely many endpoints. BIT Numerical Mathematics, 52(4):1059–1075, December 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0384-2.

Ruhe:2013:PBa

[3067] Axel Ruhe. Preface to BIT 53:1.

BIT Numerical Mathematics, 53(1):1– 2, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0427-3;   http:
//link.springer.com/content/pdf/
10.1007/s10543-013-0427-3.

Barth:2013:MMC

[3068] Andrea Barth, Annika Lang, and Christoph Schwab. Multilevel Monte Carlo method for parabolic stochastic partial differential equations. BIT Numerical Mathematics, 53(1):3–27, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0401-5.

Cano:2013:CIS

[3069] B. Cano. Conservation of invariants by symmetric multistep cosine methods for second-order partial differential equations. BIT Numerical Mathematics, 53(1):29–56, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0393-1.

Castrillon-Candas:2013:DAH

[3070] Julio E. Castrillón-Candás, Jun Li, and Victor Eijkhout. A discrete adapted hierarchical basis solver for radial basis function interpolation. BIT Numerical Mathematics, 53(1):57–86, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0397-x.

Dagnino:2013:EBA

[3071] Catterina Dagnino, Sara Remogna, and Paul Sablonnière. Error bounds on the approximation of functions and partial derivatives by quadratic spline quasiinterpolants on non-uniform criss-cross triangulations of a rectangular domain. BIT Numerical Mathematics, 53(1):87– 109, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0392-2.

DiPietro:2013:DSE

[3072] Daniele A. Di Pietro, Jean-Marc Gratien, and Christophe Prud’homme. A domain-specific embedded language in C++ for lowest-order discretizations of diffusive problems on general meshes. BIT Numerical Mathematics, 53(1):111–152, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0403-3.

Martin:2013:MFS

[3073] David R. Martin and Lothar Reichel. Minimization of functionals on the solution of a large-scale discrete illposed problem. BIT Numerical Mathematics, 53(1):153–173, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10543-012-
0396-y.

Sablonniere:2013:NIB

[3074] P. Sablonnière, D. Sbibih, and M. Tahrichi. Numerical integration based on bivariate quadratic spline quasi-interpolants on Powell–Sabin partitions. BIT Numerical Mathematics, 53(1):175–192, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0391-3.

Simeonov:2013:QBS

[3075] Plamen Simeonov and Ron Goldman. Quantum B-splines. BIT Numerical Mathematics, 53(1):193– 223, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0395-z.

Wu:2013:TBP

[3076] Gang Wu, Yimin Wei, Zhi gang Jia, Si tao Ling, and Lu Zhang. Towards backward perturbation bounds for approximate dual Krylov subspaces. BIT Numerical Mathematics, 53(1):225– 239, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0402-4.

Xiang:2013:EMV

[3077] Shuhuang Xiang and Hermann Brunner. Efficient methods for Volterra integral equations with highly oscillatory Bessel kernels. BIT Numerical Mathematics, 53(1):241–263, March 2013.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-012-
                                        0399-8.

Xue:2013:EVP

[3078] Huiyan Xue and Antonella Zanna. Explicit volume-preserving splitting methods for polynomial divergence-free vector fields. BIT Numerical Mathematics, 53(1):265–281, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-012-
                                        0394-0.

Ruhe:2013:PBb

[3079] Axel Ruhe. Preface to BIT 53:2. BIT Numerical Mathematics, 53(2): 283–284, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-013-0428-2;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-013-0428-2.pdf.

Baglama:2013:IRB

[3080] James Baglama and Lothar Reichel. An implicitly restarted block Lanczos bidiagonalization method using Leja shifts. BIT Numerical Mathematics, 53(2):285–310, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-012-
                                        0409-x.

Bebendorf:2013:HMA

[3081] M. Bebendorf, M. Bollhöfer, and M. Bratsch. Hierarchical matrix approximation with blockwise constraints. BIT Numerical Mathematics, 53(2):311–339, June 2013. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0413-1.

Bjork:2013:MCE

[3082] T. Björk, A. Szepessy, R. Tempone, and G. E. Zouraris. Monte Carlo Euler approximations of HJM term structure financial models. BIT Numerical Mathematics, 53(2):341–383, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0410-4.

Carnicer:2013:RIM

[3083] J. M. Carnicer, J. Delgado, and J. M. Peña. Richardson’s iterative method for surface interpolation. BIT Numerical Mathematics, 53(2):385– 396, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0411-3.

Floater:2013:AHS

[3084] Michael S. Floater and Bartlomiej P.

Siwek. Analysis of Hermite subdivision using piecewise polynomials. BIT Numerical Mathematics, 53(2): 397–409, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0418-9.

Hao:2013:AMB

[3085] Yong-Xia Hao, Chong-Jun Li, and Ren-Hong Wang. An approximation method based on MRA for the quasiplateau problem. BIT Numerical Mathematics, 53(2):411–442, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-012-
                                        0412-2.

Huang:2013:FST

[3086] Rong Huang. Factoring symmetric totally nonpositive matrices and inverses with a diagonal pivoting method. BIT Numerical Mathematics, 53(2): 443–458, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-012-0404-2.

Huckle:2013:DBR

[3087] T. Huckle and M. Sedlacek. Data based regularization for discrete deconvolution problems. BIT Numerical Mathematics, 53(2):459–473, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0421-9.

Jensen:2013:FEC

[3088] Max Jensen and Axel Målqvist. Finite element convergence for the Joule heating problem with mixed boundary conditions. BIT Numerical Mathematics, 53(2):475–496, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s10543-012-
0406-0.

Kovacs:2013:WCF

[3089] Mihály Kovács, Stig Larsson, and Fredrik Lindgren. Weak convergence of finite element approximations of linear stochastic evolution equations with additive noise II. Fully discrete schemes. BIT Numerical Mathematics, 53(2): 497–525, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-012-0405-1.

Pantelimon:2013:CFS

[3090] Ioana Pantelimon and Constantin Popa. Constraining by a family of strictly nonexpansive idempotent functions with applications in image reconstruction. BIT Numerical Mathematics, 53(2):527–544, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0414-0.

Shadimetov:2013:CIS

[3091] Kholmat M. Shadimetov and Abdullo R. Hayotov. Construction of interpolation splines minimizing seminorm in W (m, m−1)2(0, 1)W_{2}^{(m,m−1)}(0, 1) space. BIT Numerical Mathematics, 53(2):545–563, June 2013. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0407-z.

Ruhe:2013:PBc

[3092] Axel Ruhe. Preface to BIT 53:3. BIT Numerical Mathematics, 53(3): 565–566, September 2013. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0440-6;   http://link.springer.
                                        com/content/pdf/10.1007/s10543-
                                        013-0440-6.pdf.

Carroll:2013:EAR

[3093] John Carroll and Eoin O’Callaghan. Exponential almost Runge–Kutta methods for semilinear problems. BIT Numerical Mathematics, 53(3):567– 594, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10543-012-0416-
                                        y.

Guttel:2013:BBR

[3094] Stefan Güttel and Leonid Knizhnerman. A black-box rational Arnoldi variant for Cauchy–Stieltjes matrix functions. BIT Numerical Mathematics, 53(3):595–616, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0420-x.

Komori:2013:SRK

[3095] Yoshio Komori and Evelyn Buckwar. Stochastic Runge–Kutta methods with deterministic high order for ordinary differential equations. BIT Numerical Mathematics, 53(3):617–639, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0419-3.

Kreiss:2013:DSP

[3096] Gunilla Kreiss and Kenneth Duru. Discrete stability of perfectly matched layers for anisotropic wave equations in first and second order formulation. BIT Numerical Mathematics, 53(3):641–663, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0426-4.

Liu:2013:SPE

[3097] Huipo Liu, Wei Gong, Shuanghu Wang, and Ningning Yan. Superconvergence and a posteriori error estimates for the Stokes eigenvalue problems. BIT Numerical Mathematics, 53(3):665–687, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0422-8.

Saedpanah:2013:PEA

[3098] Fardin Saedpanah. A posteriori error analysis for a continuous space-time finite element method for a hyperbolic integro-differential equation. BIT Numerical Mathematics, 53(3):689–716, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0424-6.

Schmitt:2013:TSP

[3099] Bernhard A. Schmitt, Rüdiger Weiner, and Steffen Beck. Two-step peer methods with continuous output. BIT Numerical Mathematics, 53(3):717–739, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-012-0415-z.

Siewer:2013:NSC

[3100] Ralf Siewer. Nodal splines on compact intervals. BIT Numerical Mathematics, 53(3):741–753, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0429-1.

Vabishchevich:2013:SSH

[3101] Petr N. Vabishchevich. Splitting schemes for hyperbolic heat conduction equation. BIT Numerical Mathematics, 53(3):755–778, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0423-7.

Vatchev:2013:AFP

[3102] V. Vatchev and J. Del Castillo. Approximation of Fejér partial sums by interpolating functions. BIT Numerical Mathematics, 53(3):779–790, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-013-0425-5.

Wadbro:2013:UWC

[3103] Eddie Wadbro, Sara Zahedi, Gunilla Kreiss, and Martin Berggren. A uniformly well-conditioned, unfitted Nitsche method for interface problems. BIT Numerical Mathematics, 53(3):791–820, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-012-
0417-x.

Ruhe:2013:PBd

[3104] Axel Ruhe. Preface to BIT 53:4.

BIT Numerical Mathematics, 53(4): 821–825, December 2013. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0456-y;   http://link.springer.
com/content/pdf/10.1007/s10543-
013-0456-y.pdf.

Abdulle:2013:MSS

[3105] A. Abdulle, G. Vilmart, and K. C.

Zygalakis. Mean-square A-stable diagonally drift-implicit integrators of weak second order for stiff Itô stochastic differential equations. BIT Numerical Mathematics, 53(4):827–840, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0430-8.

Chand:2013:CAC

[3106] A. K. B. Chand and P. Viswanathan.

A constructive approach to cubic Hermite Fractal Interpolation Function and its constrained aspects. BIT Numerical Mathematics, 53(4):841–865, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0442-4.

DAmbrosio:2013:SIC

[3107] R. D’Ambrosio, E. Hairer, and C. J. Zbinden. G-symplecticity implies conjugate-symplecticity of the underlying one-step method. BIT Numerical Mathematics, 53(4):867–872, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0437-1.

Dagnino:2013:NIB

[3108] Catterina Dagnino, Paola Lamberti, and Sara Remogna. Numerical integration based on trivariate C^{2} quartic spline quasi-interpolants. BIT Numerical Mathematics, 53(4):873–896, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0431-7.

Martin-Dorel:2013:SIR

[3109] Érik Martin-Dorel, Guillaume Melquiond, and Jean-Michel Muller. Some issues related to double rounding. BIT Numerical Mathematics, 53(4):897–924, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0436-2.

Hiptmair:2013:STE

[3110] Ralf Hiptmair, Carlos Jerez-Hanckes, and Christoph Schwab. Sparse tensor edge elements. BIT Numerical Mathematics, 53(4):925–939, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0435-3.

Huang:2013:CRP

[3111] Tsung-Ming Huang, Zhongxiao Jia, and Wen-Wei Lin. On the convergence of Ritz pairs and refined Ritz vectors for quadratic eigenvalue problems. BIT Numerical Mathematics, 53(4):941–958, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0438-0.

Jia:2013:RDC

[3112] Zhongxiao Jia and Qian Zhang. Robust dropping criteria for F -norm minimization based sparse approximate inverse preconditioning. BIT Numerical Mathematics, 53(4):959–985, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0434-4.

Li:2013:CAH

[3113] Shishun Li and Zhengda Huang. Convergence analysis of HSS-multigrid methods for second-order nonselfadjoint elliptic problems. BIT Numerical Mathematics, 53(4):987–1012, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-013-0433-5.

Mazure:2013:ECS

[3114] Marie-Laurence Mazure. Extended Chebyshev spaces in rationality. BIT Numerical Mathematics, 53(4):1013– 1045, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10543-013-0432-
                                        6.

Zhang:2013:MEM

[3115] Yujie Zhang and Yangfeng Su. A memory-efficient model order reduction for time-delay systems. BIT Numerical Mathematics, 53(4):1047–1073, December 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-013-0439-z.

Benner:2014:SMC

[3116] Peter Benner, Miloud Sadkane, and Ahmed Salam. Structured matrix computations in non-Euclidean geometries: algorithms and applications. BIT Numerical Mathematics, 54(1):1– 5, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0479-z;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-014-0479-z.pdf.

Aurentz:2014:FCE

[3117] Jared L. Aurentz, Raf Vandebril, and David S. Watkins. Fast computation of eigenvalues of companion, comrade, and related matrices. BIT Numerical Mathematics, 54(1):7–30, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0449-x.

Bai:2014:MPC

[3118] Zhaojun Bai and Ren-Cang Li. Minimization principles and computation for the generalized linear response eigenvalue problem. BIT Numerical Mathematics, 54(1):31–54, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0472-6.

Bini:2014:GMS

[3119] Dario A. Bini, Bruno Iannazzo, Ben Jeuris, and Raf Vandebril. Geometric means of structured matrices. BIT Numerical Mathematics, 54(1):55– 83, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0450-4.

Boito:2014:IRS

[3120] P. Boito, Y. Eidelman, and L. Gemignani.

Implicit QR for rank-structured matrix pencils. BIT Numerical Mathematics, 54(1):85–111, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0478-0.

Caliari:2014:CSC

[3121] Marco Caliari, Peter Kandolf, and Alexander Ostermann. Comparison of software for computing the action of the matrix exponential. BIT Numerical Mathematics, 54(1):113– 128, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-013-0446-0.

Dykes:2014:SIM

[3122] L. Dykes, F. Marcellán, and L. Reichel. The structure of iterative methods for symmetric linear discrete illposed problems. BIT Numerical Mathematics, 54(1):129–145, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0476-2.

Fan:2014:REI

[3123] Hung-Yuan Fan, Peter Chang-Yi Weng, and Eric King wah Chu. Refining estimates of invariant and deflating subspaces for large and sparse matrices and pencils. BIT Numerical Mathematics, 54(1):147–169, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0469-1.

Lubich:2014:PSI

[3124] Christian Lubich and Ivan V. Oseledets. A projector-splitting integrator for dynamical low-rank approximation. BIT Numerical Mathematics, 54(1):171–188, March 2014.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0454-0.

Malyshev:2014:BMM

[3125] Alexander Malyshev and Miloud Sadkane. A bisection method for measuring the distance of a quadratic matrix polynomial to the quadratic matrix polynomials that are singular on the unit circle. BIT Numerical Mathematics, 54(1):189–200, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0445-1.

Mastronardi:2014:ASI

[3126] Nicola Mastronardi and Paul Van Dooren. An algorithm for solving the indefinite least squares problem with equality constraints. BIT Numerical Mathematics, 54(1):201– 218, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0452-2.

Mehl:2014:EPT

[3127] Christian Mehl, Volker Mehrmann, André C. M. Ran, and Leiba Rodman. Eigenvalue perturbation theory of symplectic, orthogonal, and unitary matrices under generic structured rank one perturbations. BIT Numerical Mathematics, 54(1):219–255, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10543-013-
                                        0451-3.

Munthe-Kaas:2014:SSL

[3128] H. Z. Munthe-Kaas, G. R. W. Quispel, and A. Zanna. Symmetric spaces and Lie triple systems in numerical analysis of differential equations. BIT Numerical Mathematics, 54(1):257– 282, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0473-5;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-014-0473-5.pdf.

Salam:2014:EBM

[3129] A. Salam and E. Al-Aidarous. Equivalence between modified symplectic Gram–Schmidt and Householder SR algorithms. BIT Numerical Mathematics, 54(1):283–302, March 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0441-5.

Ruhe:2014:PBa

[3130] Axel Ruhe. Preface to BIT 54:2. BIT Numerical Mathematics, 54(2): 303–304, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0504-2;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-014-0504-2.pdf.

Arnold:2014:AHD

[3131] Andreas Arnold and Tobias Jahnke. On the approximation of high-dimensional differential equations in the hierarchical Tucker format. BIT Numerical Mathematics, 54(2):305–341, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0444-2.

Barreras:2014:AED

[3132] A. Barreras and J. M. Peña. Accurate and efficient LDU decomposition of almost diagonally dominant Z-matrices. BIT Numerical Mathematics, 54(2): 343–356, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0461-1.

Bodendiek:2014:AOR

[3133] André Bodendiek and Matthias Bollhöfer. Adaptive-order rational Arnoldi-type methods in computational electromagnetism. BIT Numerical Mathematics, 54(2):357– 380, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0458-9.

Freitag:2014:NAC

[3134] Melina A. Freitag and Alastair Spence.

A new approach for calculating the real stability radius. BIT Numerical Mathematics, 54(2):381–400, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0457-x.

Goda:2014:CSP

[3135] Takashi Goda. Construction of scrambled polynomial lattice rules over F_{2} with small mean square weighted L_{∈} discrepancy. BIT Numerical Mathematics, 54(2):401–423, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0459-8.

Juntunen:2014:OSO

[3136] Mika Juntunen and Jeonghun Lee. Optimal second order rectangular elasticity elements with weakly symmetric stress. BIT Numerical Mathematics, 54(2):425–445, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0460-2.

Kressner:2014:LRT

[3137] Daniel Kressner, Michael Steinlechner, and Bart Vandereycken. Lowrank tensor completion by Riemannian optimization. BIT Numerical Mathematics, 54(2):447–468, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-013-
                                        0455-z.

Lamichhane:2014:FEM

[3138] Bishnu P. Lamichhane. A finite element method for a biharmonic equation based on gradient recovery operators. BIT Numerical Mathematics, 54(2): 469–484, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0462-0.

Ling:2014:ORI

[3139] Bo Ling and Yongping Liu. Optimal recovery of isotropic classes of r th differentiable multivariate functions. BIT Numerical Mathematics, 54(2): 485–500, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0453-1.

Novati:2014:GBA

[3140] Paolo Novati and Maria Rosaria Russo. A GCV based Arnoldi– Tikhonov regularization method. BIT Numerical Mathematics, 54(2):501– 521, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0447-z.

Sun:2014:NIA

[3141] Li Sun and Ke Chen. A new iterative algorithm for mean curvaturebased variational image denoising. BIT Numerical Mathematics, 54(2):523– 553, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0448-y.

Yan:2014:HON

[3142] Yubin Yan, Kamal Pal, and Neville J.

Ford. Higher order numerical methods for solving fractional differential equations. BIT Numerical Mathematics, 54(2):555–584, June 2014.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0443-3.

Ruhe:2014:P

[3143] Axel Ruhe. Preface to 54-3. BIT Numerical Mathematics, 54(3):585–586, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0518-9;   http:
//link.springer.com/content/pdf/
10.1007/s10543-014-0518-9.pdf.

Asheim:2014:RFN

[3144] Andreas Asheim. A remedy for the failure of the numerical steepest descent method on a class of oscillatory integrals. BIT Numerical Mathematics, 54(3):587–605, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0463-z.

Chen:2014:SPI

[3145] Hao Chen. A splitting preconditioner for the iterative solution of implicit Runge–Kutta and boundary value methods. BIT Numerical Mathematics, 54(3):607–621, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0467-3.

Chen:2014:SOL

[3146] Minghua Chen, Yantao Wang, Xiao Cheng, and Weihua Deng. Secondorder LOD multigrid method for multidimensional Riesz fractional diffusion equation. BIT Numerical Mathematics, 54(3):623–647, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0477-1.

Hansen:2014:RID

[3147] Per Christian Hansen, James G. Nagy, and Konstantinos Tigkos. Rotational image deblurring with sparse matrices. BIT Numerical Mathematics, 54(3):649–671, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-013-
0464-y.

Hansen:2014:IEL

[3148] Eskil Hansen and Tony Stillfjord. Implicit Euler and Lie splitting discretizations of nonlinear parabolic equations with delay. BIT Numerical Mathematics, 54(3):673–689, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0480-6.

Hoffmann:2014:CPB

[3149] Miklós Hoffmann, Imre Juhász, and Gyula Károlyi. A control point based curve with two exponential shape parameters. BIT Numerical Mathematics, 54(3):691–710, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10543-014-
                                        0468-2.

Huang:2014:ASD

[3150] Rong Huang, Jianzhou Liu, and Li Zhu. Accurate solutions of diagonally dominant tridiagonal linear systems. BIT Numerical Mathematics, 54(3):711–727, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0481-5.

Kieri:2014:ACS

[3151] Emil Kieri. Accelerated convergence for Schrödinger equations with non-smooth potentials. BIT Numerical Mathematics, 54(3):729–748, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-013-0465-x.

Lamnii:2014:RAO

[3152] M. Lamnii, A. Mazroui, and A. Tijini. Raising the approximation order of multivariate quasi-interpolants. BIT Numerical Mathematics, 54(3): 749–761, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10543-014-0470-
                                        8.

Macias-Diaz:2014:ENS

[3153] J. E. Macías-Díaz. On an exact numerical simulation of solitary-wave solutions of the Burgers–Huxley equation through Cardano’s method. BIT Numerical Mathematics, 54(3):763–776, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-013-0466-9.

Miyatake:2014:EPE

[3154] Yuto Miyatake. An energy-preserving exponentially-fitted continuous stage Runge–Kutta method for Hamiltonian systems. BIT Numerical Mathematics, 54(3):777–799, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0474-4.

Perret:2014:WSO

[3155] C. Perret and W. P. Petersen. A weak second-order split-step method for numerical simulations of stochastic differential equations. BIT Numerical Mathematics, 54(3):801–821, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0482-4.

Svard:2014:NTL

[3156] Magnus Svärd. A note on L^{∞} bounds and convergence rates of summation-by-parts schemes. BIT Numerical Mathematics, 54(3):823–830, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0471-7.

Wu:2014:SDS

[3157] Shu-Lin Wu and Mohammad D. Al-Khaleel. Semi-discrete Schwarz waveform relaxation algorithms for reaction diffusion equations. BIT Numerical Mathematics, 54(3):831–866, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0475-3.

Ruhe:2014:PBb

[3158] Axel Ruhe. Preface to BIT 54:4. BIT Numerical Mathematics, 54(4): 867–871, December 2014. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0535-8;   http://link.springer.
com/content/pdf/10.1007/s10543-
014-0535-8.pdf.

Auzinger:2014:EEB

[3159] Winfried Auzinger, Othmar Koch, and Amir Saboor Bagherzadeh. Error estimation based on locally weighted defect for boundary value problems in second order ordinary differential equations. BIT Numerical Mathematics, 54(4):873–900, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0488-y.

Banjai:2014:SRK

[3160] Lehel Banjai and Maryna Kachanovska. Sparsity of Runge–Kutta convolution weights for the three-dimensional wave equation. BIT Numerical Mathematics, 54(4):901–936, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0498-9.

Bueno-Orovio:2014:FSM

[3161] Alfonso Bueno-Orovio, David Kay, and Kevin Burrage. Fourier spectral methods for fractional-in-space reaction-diffusion equations. BIT Numerical Mathematics, 54(4):937–954, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0484-2.

Conti:2014:DHS

[3162] Costanza Conti, Jean-Louis Merrien, and Lucia Romani. Dual Hermite subdivision schemes of de Rhamtype. BIT Numerical Mathematics, 54(4):955–977, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0495-z.

Dagnino:2014:SFI

[3163] Catterina Dagnino, Sara Remogna, and Paul Sablonnière. On the solution of Fredholm integral equations based on spline quasi-interpolating projectors. BIT Numerical Mathematics, 54(4):979–1008, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0486-0.

Hurl:2014:SAC

[3164] Nicholas Hurl, William Layton, Yong Li, and Catalin Trenchea. Stability analysis of the Crank–Nicolson– Leapfrog method with the Robert– Asselin–Williams time filter. BIT Numerical Mathematics, 54(4):1009–1021, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0493-1.

Kiessling:2014:CEE

[3165] Jonas Kiessling and Raúl Tempone. Computable error estimates of a finite difference scheme for option pricing in exponential Lévy models. BIT Numerical Mathematics, 54(4):1023–1065, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0490-4.

Komori:2014:SEE

[3166] Yoshio Komori and Kevin Burrage. A stochastic exponential Euler scheme for simulation of stiff biochemical reaction systems. BIT Numerical Mathematics, 54(4):1067–1085, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0485-1.

Ku:2014:SMF

[3167] JaEun Ku. Supercloseness of the mixed finite element method for the primary function on unstructured meshes and its applications. BIT Numerical Mathematics, 54(4):1087–1097, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0494-0.

Lamnii:2014:CSQ

[3168] A. Lamnii, M. Lamnii, and H. Mraoui.

Cubic spline quasi-interpolants on Powell–Sabin partitions. BIT Numerical Mathematics, 54(4):1099–1118, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0489-x.

Miyatake:2014:GFF

[3169] Yuto Miyatake and Takayasu Matsuo. A general framework for finding energy dissipative/conservative H^{1}- Galerkin schemes and their underlying H^{1}-weak forms for nonlinear evolution equations. BIT Numerical Mathematics, 54(4):1119–1154, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0483-3.

Skelton:2014:ACP

[3170] Andrew Skelton and Allan R. Willms.

An algorithm for continuous piecewise linear bounding of discrete time series data. BIT Numerical Mathematics, 54(4):1155–1169, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0492-2.

Hiptmair:2015:PB

[3171] Ralf Hiptmair. Preface to BIT 55:1.

BIT Numerical Mathematics, 55(1):1– 3, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0548-y;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-015-0548-y.pdf.

Chan:2015:TDE

[3172] John Fun-Choi Chan and Peter Monk. Time dependent electromagnetic scattering by a penetrable obstacle. BIT Numerical Mathematics, 55(1):5–31, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0500-6.

Claeys:2015:SKG

[3173] Xavier Claeys, Ralf Hiptmair, and Elke Spindler. A second-kind Galerkin boundary element method for scattering at composite objects. BIT Numerical Mathematics, 55(1):33–57, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0496-y.

Dauge:2015:RME

[3174] Monique Dauge, Richard A. Norton, and Robert Scheichl. Regularity for Maxwell eigenproblems in photonic crystal fibre modelling. BIT Numerical Mathematics, 55(1):59–80, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0487-z.

Fliss:2015:RRT

[3175] Sonia Fliss, Dirk Klindworth, and Kersten Schmidt. Robin-to-Robin transparent boundary conditions for the computation of guided modes in photonic crystal wave-guides. BIT Numerical Mathematics, 55(1):81–115, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0521-1.

Ganesh:2015:EAC

[3176] M. Ganesh and S. C. Hawkins. An efficient O(N) algorithm for computing O(N^{∈}) acoustic wave interactions in large N-obstacle three dimensional configurations. BIT Numerical Mathematics, 55(1):117–139, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0491-3.

Gillman:2015:SAD

[3177] Adrianna Gillman, Alex H. Barnett, and Per-Gunnar Martinsson. A spectrally accurate direct solution technique for frequency-domain scattering problems with variable media. BIT Numerical Mathematics, 55(1):141– 170, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0499-8.

Graham:2015:WEB

[3178] I. G. Graham, M. Löhndorf, J. M. Melenk, and E. A. Spence. When is the error in the h-BEM for solving the Helmholtz equation bounded independently of k ? BIT Numerical Mathematics, 55(1):171–214, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0501-5.

Hohage:2015:CIE

[3179] Thorsten Hohage and Lothar Nannen. Convergence of infinite element methods for scalar waveguide problems. BIT Numerical Mathematics, 55(1):215– 254, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0525-x.

Hoppe:2015:AEE

[3180] Ronald H. W. Hoppe and Irwin Yousept. Adaptive edge element approximation of H (curl)-elliptic optimal control problems with control constraints. BIT Numerical Mathematics, 55(1):255–277, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0497-x.

Spence:2015:BAL

[3181] E. A. Spence. Bounding acoustic layer potentials via oscillatory integral techniques. BIT Numerical Mathematics, 55(1):279–318, March 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0506-0.

Beck:2015:PPI

[3182] Mélanie Beck and Martin J. Gander.

On the positivity of Poisson integrators for the Lotka–Volterra equations. BIT Numerical Mathematics, 55(2): 319–340, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0505-1.

Bolten:2015:GGT

[3183] Matthias Bolten, Marco Donatelli, Thomas Huckle, and Christos Kravvaritis. Generalized grid transfer operators for multigrid methods applied on Toeplitz matrices. BIT Numerical Mathematics, 55(2):341– 366, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0512-2.

Bremer:2015:HOA

[3184] James Bremer, Adrianna Gillman, and Per-Gunnar Martinsson. A high-order accurate accelerated direct solver for acoustic scattering from surfaces. BIT Numerical Mathematics, 55(2):367– 397, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0508-y.

Collier:2015:CMM

[3185] Nathan Collier, Abdul-Lateef Haji-Ali, Fabio Nobile, Erik von Schwerin, and Raúl Tempone. A continuation multilevel Monte Carlo algorithm. BIT Numerical Mathematics, 55(2): 399–432, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0511-3.

Gosse:2015:WBA

[3186] Laurent Gosse. A well-balanced and asymptotic-preserving scheme for the one-dimensional linear Dirac equation. BIT Numerical Mathematics, 55(2): 433–458, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0510-4.

Hiptmair:2015:CAS

[3187] R. Hiptmair, A. Paganini, and S. Sargheini. Comparison of approximate shape gradients. BIT Numerical Mathematics, 55(2):459– 485, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0515-z.

Oishi:2015:SAC

[3188] Cassio M. Oishi, Jin Y. Yuan, Jose A. Cuminato, and David E. Stewart. Stability analysis of Crank–Nicolson and Euler schemes for time-dependent diffusion equations. BIT Numerical Mathematics, 55(2):487–513, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0509-x.

Senosiain:2015:RNS

[3189] M. J. Senosiain and A. Tocino. A review on numerical schemes for solving a linear stochastic oscillator. BIT Numerical Mathematics, 55(2):515– 529, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0507-z.

Soderlind:2015:SSY

[3190] Gustaf Söderlind, Laurent Jay, and Manuel Calvo. Stiffness 1952–2012: Sixty years in search of a definition. BIT Numerical Mathematics, 55(2): 531–558, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0503-3.

Speleers:2015:IPB

[3191] Hendrik Speleers. Inner products of box splines and their derivatives. BIT Numerical Mathematics, 55(2):559– 567, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0513-1.

Sun:2015:DSN

[3192] Zhe Sun, Lei Wu, and Zhe Liu. A damped semismooth Newton method for the Brugnano–Casulli piecewise linear system. BIT Numerical Mathematics, 55(2):569–589, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0514-0.

Trogdon:2015:AGO

[3193] Thomas Trogdon. On the application of GMRES to oscillatory singular integral equations. BIT Numerical Mathematics, 55(2):591–620, June 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0502-4.

Bjorck:2015:AR

[3194] Åke Björck and Lars Eldén. Axel Ruhe 1942–2015. BIT Numerical Mathematics, 55(3):621–623, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0579-4;   http:
//link.springer.com/content/pdf/
10.1007/s10543-015-0579-4.pdf;
https://www.math.utah.edu/pub/bibnet/
authors/r/ruhe-axel.bib.

Awanou:2015:SEM

[3195] Gerard Awanou. Spline element method for Monge–Ampère equations. BIT Numerical Mathematics, 55(3):625–646, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0524-y.

Boglaev:2015:MIA

[3196] Igor Boglaev. Monotone iterative ADI method for semilinear parabolic problems. BIT Numerical Mathematics, 55(3):647–676, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0529-6.

Bosner:2015:EAS

[3197] Nela Bosner. Efficient algorithm for simultaneous reduction to the m-Hessenberg–triangular–triangular form. BIT Numerical Mathematics, 55(3): 677–703, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://link.springer.com/
article/10.1007/s10543-014-0516-
y.

Cohen:2015:LTA

[3198] David Cohen, Ludwig Gauckler, Ernst Hairer, and Christian Lubich. Longterm analysis of numerical integrators for oscillatory Hamiltonian systems under minimal non-resonance conditions. BIT Numerical Mathematics, 55(3):705–732, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0527-8.

Ghrist:2015:SOA

[3199] Michelle L. Ghrist, Bengt Fornberg, and Jonah A. Reeger. Stability ordinates of Adams predictor-corrector methods. BIT Numerical Mathematics, 55(3):733–750, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0528-7.

Lettieri:2015:SPH

[3200] Davide Lettieri, Carla Manni, and Francesca Pelosi. Shape preserving HC^{2} interpolatory subdivision. BIT Numerical Mathematics, 55(3):

751–779, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10543-014-0530-
                                        0.

Levin:2015:BML

[3201] David Levin. Between moving leastsquares and moving least-ℓ_{1}. BIT Numerical Mathematics, 55(3):781–796, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0522-0.

Sbibih:2015:SCC

[3202] Driss Sbibih, Abdelhafid Serghini, Ahmed Tijini, and Ahmed Zidna. Superconvergent C^{1} cubic spline quasiinterpolants on Powell–Sabin partitions. BIT Numerical Mathematics, 55(3):797–821, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0523-z.

Smith:2015:CPQ

[3203] H. V. Smith. Convergence properties of a quadrature formula of Clenshaw– Curtis type for the Gegenbauer weight function. BIT Numerical Mathematics, 55(3):823–842, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0520-2.

Speck:2015:MLS

[3204] Robert Speck, Daniel Ruprecht, Matthew Emmett, and Michael Minion. A multi-level spectral deferred correction method. BIT Numerical Mathematics, 55(3):843–867, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0517-x.

Zhang:2015:BPA

[3205] Lei-Hong Zhang, Wen-Wei Lin, and Ren-Cang Li. Backward perturbation analysis and residual-based error bounds for the linear response eigenvalue problem. BIT Numerical Mathematics, 55(3):869–896, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0519-8.

Elden:2015:PB

[3206] Lars Eldén. Preface to BIT 55:4.

BIT Numerical Mathematics, 55(4): 897–899, December 2015. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0593-6;   http://link.springer.
com/content/pdf/10.1007/s10543-
015-0593-6.pdf.

Behroozifar:2015:SMS

[3207] Mahmoud Behroozifar. Spectral method for solving high order nonlinear boundary value problems via operational matrices. BIT Numerical Mathematics, 55(4):901–925, December 2015.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-015-
                                        0544-2.

Butcher:2015:OCG

[3208] John C. Butcher and Gulshad Imran. Order conditions for G-symplectic methods. BIT Numerical Mathematics, 55(4):927–948, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0541-x.

DeMarchi:2015:FCO

[3209] Stefano De Marchi and Gabriele Santin. Fast computation of orthonormal basis for RBF spaces through Krylov space methods. BIT Numerical Mathematics, 55(4):949–966, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0537-6.

Deng:2015:LDG

[3210] Weihua Deng and Jan S. Hesthaven. Local discontinuous Galerkin methods for fractional ordinary differential equations. BIT Numerical Mathematics, 55(4):967–985, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0531-z.

Dieci:2015:OSD

[3211] Luca Dieci and Luciano Lopez. Onesided direct event location techniques in the numerical solution of discontinuous differential systems. BIT Numerical Mathematics, 55(4):987–1003, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0538-5.

Dumitrescu:2015:RBR

[3212] Bogdan Dumitrescu. On the relation between the randomized extended Kaczmarz algorithm and coordinate descent. BIT Numerical Mathematics, 55(4):1005–1015, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0526-9.

Collins:2015:PEA

[3213] J. B. Collins, D. Estep, and S. Tavener.

A posteriori error analysis for finite element methods with projection operators as applied to explicit time integration techniques. BIT Numerical Mathematics, 55(4):1017–1042, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0534-9.

Czekansky:2015:MHS

[3214] Johannes Czekansky and Tomas Sauer.

The multivariate Horner scheme revisited. BIT Numerical Mathematics, 55(4):1043–1056, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0533-x.

Kopec:2015:WBE

[3215] Marie Kopec. Weak backward error analysis for Langevin process. BIT Numerical Mathematics, 55(4):1057–1103, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0546-0.

Kopteva:2015:ECM

[3216] Natalia Kopteva and Martin Stynes. An efficient collocation method for a Caputo two-point boundary value problem. BIT Numerical Mathematics, 55(4):1105–1123, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-014-
                                        0539-4.

Pauli:2015:MMC

[3217] Stefan Pauli, Robert Nicholas Gantner, and Peter Arbenz. Multilevel Monte Carlo for the Feynman–Kac formula for the Laplace equation. BIT Numerical Mathematics, 55(4):1125–1143, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-014-0543-8.

Reichel:2015:TRF

[3218] Lothar Reichel and Xuebo Yu. Tikhonov regularization via flexible Arnoldi reduction. BIT Numerical Mathematics, 55(4):1145–1168, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0542-9.

Roos:2015:OPE

[3219] H.-G. Roos and M. Schopf. An optimal a priori error estimate in the maximum norm for the Il’in scheme in 2D. BIT Numerical Mathematics, 55(4):1169–1186, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0536-7.

Wang:2015:CFD

[3220] Yuan-Ming Wang. A compact finite difference method for solving a class of time fractional convection-subdiffusion equations. BIT Numerical Mathematics, 55(4):1187–1217, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-014-
0532-y.

Weiser:2015:FSC

[3221] Martin Weiser. Faster SDC convergence on non-equidistant grids by DIRK sweeps. BIT Numerical Mathematics, 55(4):1219–1241, December 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-014-0540-y.

Xie:2015:TML

[3222] Hehu Xie. A type of multi-level correction scheme for eigenvalue problems by nonconforming finite element methods. BIT Numerical Mathematics, 55(4):1243–1266, December 2015.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-015-
                                        0545-1.

Ait-Haddou:2016:PDR

[3223] Rachid Ait-Haddou. Polynomial degree reduction in the discrete L_{2}-norm equals best Euclidean approximation of h-Bézier coefficients. BIT Numerical Mathematics, 56(1):1–14, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-015-
                                        0558-9.

Alfeld:2016:LDO

[3224] Peter Alfeld and Tatyana Sorokina. Linear differential operators on bivariate spline spaces and spline vector fields. BIT Numerical Mathematics, 56(1):15–32, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-015-
                                        0557-x.

Beliakov:2016:PAC

[3225] Gleb Beliakov and Yuri Matiyasevich. A parallel algorithm for calculation of determinants and minors using arbitrary precision arithmetic. BIT Numerical Mathematics, 56(1):33– 50, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0547-z.

Das:2016:NSS

[3226] Pratibhamoy Das and Volker Mehrmann.

Numerical solution of singularly perturbed convection-diffusion-reaction problems with two small parameters. BIT Numerical Mathematics, 56(1):51– 76, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0559-8.

Tebbens:2016:CQQ

[3227] Jurjen Duintjer Tebbens and Gérard Meurant. On the convergence of Q-OR and Q-MR Krylov methods for solving nonsymmetric linear systems. BIT Numerical Mathematics, 56(1):77– 97, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0564-y.

Giesselmann:2016:RRE

[3228] Jan Giesselmann and Tristan Pryer.

Reduced relative entropy techniques for a priori analysis of multiphase problems in elastodynamics. BIT Numerical Mathematics, 56(1):99–127, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0560-2.

Hoang:2016:FFR

[3229] N. S. Hoang and R. B. Sidje. Functionally fitted Runge–Kutta–Nyström methods. BIT Numerical Mathematics, 56(1):129–150, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0561-1.

Leobacher:2016:NMS

[3230] Gunther Leobacher and Michaela Szölgyenyi. A numerical method for SDEs with discontinuous drift. BIT Numerical Mathematics, 56(1):151– 162, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0549-x.

Mali:2016:ESS

[3231] O. Mali and S. Repin. Estimates of the solution set for a class of elliptic problems with incompletely known data. BIT Numerical Mathematics, 56(1):163–188, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0552-2.

Moraes:2016:MHC

[3232] Alvaro Moraes, Raúl Tempone, and Pedro Vilanova. Multilevel hybrid Chernoff tau-leap. BIT Numerical Mathematics, 56(1):189–239, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0556-y.

Plestenjak:2016:NMN

[3233] Bor Plestenjak. Numerical methods for nonlinear two-parameter eigenvalue problems. BIT Numerical Mathematics, 56(1):241–262, March 2016.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0566-9.

Risebro:2016:MMC

[3234] Nils Henrik Risebro, Christoph Schwab, and Franziska Weber. Multilevel Monte Carlo front-tracking for random scalar conservation laws. BIT Numerical Mathematics, 56(1):263– 292, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0550-4. See correc-

tion [3354].

Rump:2016:IEB

[3235] Siegfried M. Rump, Florian Bünger, and Claude-Pierre Jeannerod. Improved error bounds for floating-point products and Horner’s scheme. BIT Numerical Mathematics, 56(1):293– 307, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0555-z.

Rump:2016:DUR

[3236] Siegfried M. Rump and Marko Lange.

On the definition of unit roundoff. BIT Numerical Mathematics, 56(1):309– 317, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0554-0.

Scholz:2016:RDB

[3237] Lena Scholz and Andreas Steinbrecher.

Regularization of DAEs based on the Signature method. BIT Numerical Mathematics, 56(1):319–340, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0565-x.

Sorevik:2016:TIL

[3238] Tor Sørevik and Morten A. Nome. Trigonometric interpolation on lattice grids. BIT Numerical Mathematics, 56(1):341–356, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0562-0;   http://link.springer.
com/content/pdf/10.1007/s10543-
015-0562-0.pdf.

Spalevic:2016:EEG

[3239] M. M. Spalević and A. S. Cvetković. Estimating the error of Gaussian quadratures with simple and multiple nodes by using their extensions with multiple nodes. BIT Numerical Mathematics, 56(1):357–374, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0551-3.

Weiner:2016:EKP

[3240] Rüdiger Weiner and Jürgen Bruder. Exponential Krylov peer integrators. BIT Numerical Mathematics, 56(1): 375–393, March 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://link.springer.com/
article/10.1007/s10543-015-0553-
1.

Bai:2016:EPA

[3241] Zhong-Zhi Bai and Lothar Reichel. Editorial preface: “Advances in numerical algebra and scientific computing”. BIT Numerical Mathematics, 56(2): 395–397, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-016-0612-2;   http:
//link.springer.com/content/pdf/
10.1007/s10543-016-0612-2.pdf.

Bai:2016:RCA

[3242] Zhong-Zhi Bai and Min Tao. Rigorous convergence analysis of alternating variable minimization with multiplier methods for quadratic programming problems with equality constraints. BIT Numerical Mathematics, 56(2): 399–422, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0563-z.

Cao:2016:SHP

[3243] Yang Cao, Zhi-Ru Ren, and Quan Shi. A simplified HSS preconditioner for generalized saddle point problems. BIT Numerical Mathematics, 56(2): 423–439, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0588-3.

Gander:2016:AND

[3244] Martin J. Gander, Qiang Niu, and Yingxiang Xu. Analysis of a new dimension-wise splitting iteration with selective relaxation for saddle point problems. BIT Numerical Mathematics, 56(2):441–465, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-016-
0606-0.

Huang:2016:PNI

[3245] Guangxin Huang, Lothar Reichel, and Feng Yin. Projected nonstationary iterated Tikhonov regularization. BIT Numerical Mathematics, 56(2):467– 487, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0568-7.

Ke:2016:SAM

[3246] Yi-Fen Ke and Chang-Feng Ma. Spectrum analysis of a more general augmentation block preconditioner for generalized saddle point matrices. BIT Numerical Mathematics, 56(2):489– 500, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0570-0.

Li:2016:TPM

[3247] Jing Li and Nai-Min Zhang. A tripleparameter modified SSOR method for solving singular saddle point problems. BIT Numerical Mathematics, 56(2): 501–521, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-016-0610-4.

Liang:2016:VAP

[3248] Zhao-Zheng Liang and Guo-Feng Zhang. Variants of the accelerated parameterized inexact Uzawa method for saddle-point problems. BIT Numerical Mathematics, 56(2):523– 542, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0582-9.

Meng:2016:GHM

[3249] Guo-Yan Meng, Rui-Ping Wen, and Qing-Shan Zhao. The generalized HSS method with a flexible shift-parameter for non-Hermitian positive definite linear systems. BIT Numerical Mathematics, 56(2):543–556, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0584-7.

Njeru:2016:ASL

[3250] Patrick Njue Njeru and Xue-Ping Guo. Accelerated SOR-like method for augmented linear systems. BIT Numerical Mathematics, 56(2):557– 571, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0571-z.

Ramage:2016:CFD

[3251] Alison Ramage and André M. Sonnet. Computational fluid dynamics for nematic liquid crystals. BIT Numerical Mathematics, 56(2):573– 586, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0586-5;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-015-0586-5.pdf.

Zhang:2016:VDP

[3252] Juli Zhang and Chuanqing Gu. A variant of the deteriorated PSS preconditioner for nonsymmetric saddle point problems. BIT Numerical Mathematics, 56(2):587–604, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-015-
                                        0590-9.

Girod:2016:NST

[3253] Alina Girod and Thorsten Hüls. Nonautonomous systems with transversal homoclinic structures under discretization. BIT Numerical Mathematics, 56(2):605–631, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-015-
                                        0567-8.

Ha:2016:ANS

[3254] Phi Ha and Volker Mehrmann. Analysis and numerical solution of linear delay differential-algebraic equations. BIT Numerical Mathematics, 56(2): 633–657, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0577-6.

Huang:2016:MFP

[3255] Zhongyi Huang and Ye Li. Monotone finite point method for non-equilibrium radiation diffusion equations. BIT Numerical Mathematics, 56(2):659– 679, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-015-0573-x.

Cognata:2016:WPS

[3256] Cristina La Cognata and Jan Nordström. Well-posedness, stability and conservation for a discontinuous interface problem. BIT Numerical Mathematics, 56(2):681–704, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0576-7.

Notaris:2016:ENC

[3257] Sotirios E. Notaris. The error norm of Clenshaw–Curtis and related quadrature formulae. BIT Numerical Mathematics, 56(2):705–728, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0569-6.

Ozaki:2016:SFP

[3258] Katsuhisa Ozaki, Florian Bünger, and Takeshi Ogita. Simple floating-point filters for the two-dimensional orientation problem. BIT Numerical Mathematics, 56(2):729–749, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0574-9.

Palitta:2016:MEB

[3259] Davide Palitta and Valeria Simoncini. Matrix-equation-based strategies for convection-diffusion equations. BIT Numerical Mathematics, 56(2):751– 776, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0575-8.

Scholz:2016:SAR

[3260] Lena Scholz and Andreas Steinbrecher. Structural-algebraic regularization for coupled systems of DAEs. BIT Numerical Mathematics, 56(2):777– 804, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0572-y.

Hansen:2016:OJH

[3261] Per Christian Hansen and Lars Eldén. Obituary: Jens Hugger 1958–2016. BIT Numerical Mathematics, 56(3):805, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-016-
                                        0623-z;   http://link.springer.
                                        com/content/pdf/10.1007/s10543-
                                        016-0623-z.pdf.

Annaby:2016:BTP

[3262] M. H. Annaby and R. M. Asharabi. Bounds for truncation and perturbation errors of nonuniform sampling series. BIT Numerical Mathematics, 56(3):807–832, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10543-015-0585-6;   http://link.
springer.com/article/10.1007/s10543-
015-0585-6.

Cuvelier:2016:EWA

[3263] François Cuvelier, Caroline Japhet, and Gilles Scarella. An efficient way to assemble finite element matrices in vector languages. BIT Numerical Mathematics, 56(3):833–864, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-015-0587-
4;   http://link.springer.com/article/
10.1007/s10543-015-0587-4.

Frommer:2016:EBE

[3264] Andreas Frommer and Marcel Schweitzer.

Error bounds and estimates for Krylov subspace approximations of Stieltjes matrix functions. BIT Numerical Mathematics, 56(3):865–892, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-015-0596-
3;   http://link.springer.com/article/
10.1007/s10543-015-0596-3.

Gazzola:2016:IDP

[3265] Silvia Gazzola and Paolo Novati. Inheritance of the discrete Picard condition in Krylov subspace methods. BIT Numerical Mathematics, 56(3):893–918, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/

                                        article/10.1007/s10543-015-0578-
                                        5;   http://link.springer.com/article/
                                        10.1007/s10543-015-0578-5.

Gazzola:2016:NFM

[3266] Silvia Gazzola and Lothar Reichel. A new framework for multiparameter regularization. BIT Numerical Mathematics, 56(3):919–949, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-015-0595-
                                        4;   http://link.springer.com/article/
                                        10.1007/s10543-015-0595-4.

Lundquist:2016:EFD

[3267] Tomas Lundquist and Jan Nordström. Efficient fully discrete summation-by-parts schemes for unsteady flow problems. BIT Numerical Mathematics, 56(3):951–966, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-015-0599-
                                        0;   http://link.springer.com/article/
                                        10.1007/s10543-015-0599-0.

Marcinkowski:2016:ASP

[3268] L. Marcinkowski, T. Rahman, A. Loneland, and J. Valdman. Additive Schwarz preconditioner for the finite volume element discretization of symmetric elliptic problems. BIT Numerical Mathematics, 56(3):967–993, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-015-0581-x;   http:

//link.springer.com/content/pdf/
10.1007/s10543-015-0581-x.pdf.

Muller:2016:NMR

[3269] Andreas Müller. A note on the motion representation and configuration update in time stepping schemes for the constrained rigid body. BIT Numerical Mathematics, 56(3):995–1015, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-015-0580-
y;   http://link.springer.com/article/
10.1007/s10543-015-0580-y. See er-

ratum [3270].

Muller:2016:ENM

[3270] Andreas Müller. Erratum to: A note on the motion representation and configuration update in time stepping schemes for the constrained rigid body. BIT Numerical Mathematics, 56(3):1017–1018, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-015-
0591-8;   http://link.springer.
com/content/pdf/10.1007/s10543-
015-0591-8.pdf. See [3269].

Potts:2016:RSL

[3271] Daniel Potts and Manfred Tasche. Reconstruction of sparse Legendre and Gegenbauer expansions. BIT Numerical Mathematics, 56(3):1019–1043, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-015-0598-

                                        1;   http://link.springer.com/article/
                                        10.1007/s10543-015-0598-1.

Reichel:2016:GAG

[3272] Lothar Reichel, Miodrag M. Spalević, and Tunan Tang. Generalized averaged Gauss quadrature rules for the approximation of matrix functionals. BIT Numerical Mathematics, 56(3):1045–1067, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-015-0592-
                                        7;   http://link.springer.com/article/
                                        10.1007/s10543-015-0592-7.

Rincon:2016:ACN

[3273] Mauro A. Rincon, Juliana C. Xavier, and Daniel G. Alfaro Vigo. Analysis and computation of a nonlinear Korteweg–de Vries system. BIT Numerical Mathematics, 56(3):1069–1099, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-015-0589-
                                        2;   http://link.springer.com/article/
                                        10.1007/s10543-015-0589-2.

Shardlow:2016:PAS

[3274] Tony Shardlow and Phillip Taylor. On the pathwise approximation of stochastic differential equations. BIT Numerical Mathematics, 56(3):1101–1129, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-015-0597-
                                        2;   http://link.springer.com/article/
                                        10.1007/s10543-015-0597-2.

Wang:2016:MEW

[3275] Lijin Wang, Jialin Hong, and Liying Sun. Modified equations for weakly convergent stochastic symplectic schemes via their generating functions. BIT Numerical Mathematics, 56(3):1131–1162, September 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-015-0583-
8;   http://link.springer.com/article/
10.1007/s10543-015-0583-8.

Elden:2016:PB

[3276] Lars Eldén. Preface to BIT 56:4.

BIT Numerical Mathematics, 56(4): 1163–1164, December 2016. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-016-
0640-y;   http://link.springer.
com/content/pdf/10.1007/s10543-
016-0640-y.pdf.

Bracco:2016:BHH

[3277] Cesare Bracco, Carlotta Giannelli, and Francesca Mazzia. Bivariate hierarchical Hermite spline quasiinterpolation. BIT Numerical Mathematics, 56(4):1165–1188, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0603-
3;   http://link.springer.com/article/
10.1007/s10543-016-0603-3.

Butcher:2016:SGL

[3278] J. C. Butcher, A. T. Hill, and T. J. T. Norton. Symmetric general linear methods. BIT Numerical Mathematics, 56(4):1189–1212, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0613-
1;   http://link.springer.com/article/
10.1007/s10543-016-0613-1.

Campos:2016:RAT

[3279] Carmen Campos and Jose E. Roman. Restarted Q-Arnoldi-type methods exploiting symmetry in quadratic eigenvalue problems. BIT Numerical Mathematics, 56(4):1213–1236, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0601-
5;   http://link.springer.com/article/
10.1007/s10543-016-0601-5.

Chesnel:2016:NAP

[3280] Lucas Chesnel and Xavier Claeys. A numerical approach for the Poisson equation in a planar domain with a small inclusion. BIT Numerical Mathematics, 56(4):1237–1256, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0615-
z;   http://link.springer.com/article/
10.1007/s10543-016-0615-z.

Groselj:2016:NRS

[3281] Jan Groselj. A normalized representation of super splines of arbitrary degree on Powell–Sabin triangulations. BIT Numerical Mathematics, 56(4):1257–1280, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-015-0600-
y;   http://link.springer.com/article/
10.1007/s10543-015-0600-y.

Hadjidimos:2016:SPP

[3282] Apostolos Hadjidimos. The saddle point problem and the Manteuffel algorithm. BIT Numerical Mathematics, 56(4):1281–1302, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0617-
x;   http://link.springer.com/article/
10.1007/s10543-016-0617-x.

Hansen:2016:HOS

[3283] Eskil Hansen and Alexander Ostermann. High-order splitting schemes for semilinear evolution equations. BIT Numerical Mathematics, 56(4): 1303–1316, December 2016. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10543-016-0604-2;   http://link.
springer.com/article/10.1007/s10543-
016-0604-2.

Kojima:2016:IPS

[3284] H. Kojima. Invariants preserving schemes based on explicit Runge– Kutta methods. BIT Numerical Mathematics, 56(4):1317–1337, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0608-
                                        y;   http://link.springer.com/article/
                                        10.1007/s10543-016-0608-y.

Liang:2016:CCS

[3285] Hui Liang and Hermann Brunner. On the convergence of collocation solutions in continuous piecewise polynomial spaces for Volterra integral equations. BIT Numerical Mathematics, 56(4):1339–1367, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0609-
                                        x;   http://link.springer.com/article/
                                        10.1007/s10543-016-0609-x.

Sakakibara:2016:ADS

[3286] Koya Sakakibara. Analysis of the dipole simulation method for two-dimensional Dirichlet problems in Jordan regions with analytic boundaries. BIT Numerical Mathematics, 56(4):1369–1400, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0605-
                                        1;   http://link.springer.com/article/
                                        10.1007/s10543-016-0605-1.

Shcherbakov:2016:RBF

[3287] Victor Shcherbakov. Radial basis function partition of unity operator splitting method for pricing multiasset American options. BIT Numerical Mathematics, 56(4):1401–1423, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0616-
y;   http://link.springer.com/article/
10.1007/s10543-016-0616-y.

Soltani:2016:TBD

[3288] Sara Soltani, Misha E. Kilmer, and Per Christian Hansen. A tensor-based dictionary learning approach to tomographic image reconstruction. BIT Numerical Mathematics, 56(4):1425–1454, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0607-
z;   http://link.springer.com/article/
10.1007/s10543-016-0607-z.

Stynes:2016:NCC

[3289] Martin Stynes, Eugene O’Riordan, and José Luis Gracia. Necessary conditions for convergence of difference schemes for fractional-derivative two-point boundary value problems. BIT Numerical Mathematics, 56(4):1455–1477, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0602-
4;   http://link.springer.com/article/
10.1007/s10543-016-0602-4.

Svard:2016:ESC

[3290] Magnus Svärd. Entropy solutions of the compressible Euler equations. BIT Numerical Mathematics, 56(4):1479–1496, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0611-
3;   http://link.springer.com/article/
10.1007/s10543-016-0611-3.

Zhou:2016:PMS

[3291] Weien Zhou, Liying Zhang, Jialin Hong, and Songhe Song. Projection methods for stochastic differential equations with conserved quantities. BIT Numerical Mathematics, 56(4):1497–1518, December 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0614-
0;   http://link.springer.com/article/
10.1007/s10543-016-0614-0.

Elden:2017:P

[3292] Lars Eldén. Preface. BIT Numerical Mathematics, 57(1):1–2, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-017-
0647-z;   http://link.springer.
com/content/pdf/10.1007/s10543-
017-0647-z.pdf.

Allouch:2017:SNM

[3293] Chafik Allouch, Driss Sbibih, and Mohamed Tahrichi. Superconvergent Nyström method for Urysohn integral equations. BIT Numerical Mathematics, 57(1):3–20, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10543-016-0629-6;   http://link.
springer.com/article/10.1007/s10543-
016-0629-6.

Andersson:2017:MSC

[3294] Adam Andersson and Raphael Kruse.

Mean-square convergence of the BDF2-Maruyama and backward Euler schemes for SDE satisfying a global monotonicity condition. BIT Numerical Mathematics, 57(1):21–53, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0624-
y;   http://link.springer.com/article/
10.1007/s10543-016-0624-y.

Auzinger:2017:PSM

[3295] Winfried Auzinger, Harald Hofstätter, and David Ketcheson. Practical splitting methods for the adaptive integration of nonlinear evolution equations. Part I: Construction of optimized schemes and pairs of schemes. BIT Numerical Mathematics, 57(1):55– 74, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-016-0626-9;   http:
//link.springer.com/content/pdf/
10.1007/s10543-016-0626-9.pdf.

Chen:2017:EST

[3296] Xiao Shan Chen. On estimating the separation of two periodic matrix sequences. BIT Numerical Mathematics, 57(1):75–91, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0621-
1;   http://link.springer.com/article/
10.1007/s10543-016-0621-1.

Coclite:2017:CFD

[3297] G. M. Coclite, J. Ridder, and N. H. Risebro. A convergent finite difference scheme for the Ostrovsky– Hunter equation on a bounded domain. BIT Numerical Mathematics, 57(1):93– 122, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0625-
                                        x;   http://link.springer.com/article/
                                        10.1007/s10543-016-0625-x.

delaCruz:2017:LLM

[3298] H. de la Cruz, J. C. Jimenez, and J. P. Zubelli. Locally linearized methods for the simulation of stochastic oscillators driven by random forces. BIT Numerical Mathematics, 57(1):123– 151, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0620-
                                        2;   http://link.springer.com/article/
                                        10.1007/s10543-016-0620-2.

Debrabant:2017:CAH

[3299] Kristian Debrabant and Anne Kværnø. Cheap arbitrary high order methods for single integrand SDEs. BIT Numerical Mathematics, 57(1):153– 168, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0619-
                                        8;   http://link.springer.com/article/
                                        10.1007/s10543-016-0619-8.

Ding:2017:IPS

[3300] Weiyang Ding, Liqun Qi, and Yimin Wei. Inheritance properties and sumof-squares decomposition of Hankel tensors: theory and algorithms. BIT Numerical Mathematics, 57(1):169– 190, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0622-
0;   http://link.springer.com/article/
10.1007/s10543-016-0622-0.

Petrova:2017:GGR

[3301] Guergana Petrova. Generalized Gauss– Radau and Gauss–Lobatto formulas with Jacobi weight functions. BIT Numerical Mathematics, 57(1):191– 206, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0627-
8;   http://link.springer.com/article/
10.1007/s10543-016-0627-8.

Quinn:2017:PUN

[3302] Jason Quinn. Parameter-uniform numerical methods for general nonlinear singularly perturbed reaction diffusion problems having a stable reduced solution. BIT Numerical Mathematics, 57(1):207–240, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10543-016-0628-
7;   http://link.springer.com/article/
10.1007/s10543-016-0628-7.

Tang:2017:EWS

[3303] Xiao Tang and Aiguo Xiao. Efficient weak second-order stochastic Runge–Kutta methods for Itô stochastic differential equations. BIT Numerical Mathematics, 57(1):241–260, March 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10543-016-0618-
                                        9;   http://link.springer.com/article/
                                        10.1007/s10543-016-0618-9.

Arraras:2017:MDS

[3304] A. Arrarás, K. J. in ’t Hout, W. Hundsdorfer, and L. Portero. Modified Douglas splitting methods for reactiondiffusion equations. BIT Numerical Mathematics, 57(2):261–285, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bai:2017:RHI

[3305] Zhong-Zhi Bai and Michele Benzi. Regularized HSS iteration methods for saddle-point linear systems. BIT Numerical Mathematics, 57(2):287– 311, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Butcher:2017:SMO

[3306] John C. Butcher, Gulshad Imran, and Helmut Podhaisky. A G-symplectic method with order 6. BIT Numerical Mathematics, 57(2):313–328, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Dagnino:2017:QIB

[3307] Catterina Dagnino and Sara Remogna. Quasi-interpolation based on the ZPelement for the numerical solution of integral equations on surfaces in R^{3}. BIT Numerical Mathematics, 57(2): 329–350, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See erratum [3367].

Huang:2017:MMG

[3308] G. Huang, A. Lanza, S. Morigi, L. Reichel, and F. Sgallari. Majorizationminimization generalized Krylov subspace methods for ℓ_{p}-ℓ_{q} optimization applied to image restoration. BIT Numerical Mathematics, 57(2): 351–378, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ji:2017:BFB

[3309] Hao Ji and Yaohang Li. A breakdownfree block conjugate gradient method. BIT Numerical Mathematics, 57(2): 379–403, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Mardal:2017:RPP

[3310] Kent-André Mardal, Bjørn Fredrik Nielsen, and Magne Nordaas. Robust preconditioners for PDE-constrained optimization with limited observations. BIT Numerical Mathematics, 57(2): 405–431, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/content/pdf/
10.1007/s10543-016-0635-8.pdf.

McKenzie:2017:SAB

[3311] Ross McKenzie and John Pryce. Structural analysis based dummy derivative selection for differential algebraic equations. BIT Numerical Mathematics, 57(2):433–462, June 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
content/pdf/10.1007/s10543-016-
0642-9.pdf.

Nipp:2017:ICD

[3312] Kaspar Nipp and Daniel Stoffer. Invariant curves for the discretised van der Pol equation. BIT Numerical Mathematics, 57(2):463–497, June 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Renaut:2017:EER

[3313] Rosemary A. Renaut, Michael Horst, Yang Wang, and Douglas Cochran. Efficient estimation of regularization parameters via downsampling and the singular value expansion. BIT Numerical Mathematics, 57(2):499–529, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Schichl:2017:IU

[3314] Hermann Schichl, Ferenc Domes, Tiago Montanher, and Kevin Kofler. Interval unions. BIT Numerical Mathematics, 57(2):531–556, June 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
content/pdf/10.1007/s10543-016-
0632-y.pdf.

Wang:2017:SMS

[3315] Xiaojie Wang, Ruisheng Qi, and Fengze Jiang. Sharp mean-square regularity results for SPDEs with fractional noise and optimal convergence rates for the numerical approximations. BIT Numerical Mathematics, 57(2):557–585, June 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Yang:2017:LIB

[3316] Chaoxia Yang. Linearly implicit BDF methods for nonlinear parabolic interface problems. BIT Numerical Mathematics, 57(2):587–606, June 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Elden:2017:PGK

[3317] Lars Eldén. Preface: Gunilla Kreiss new Editor-in-Chief. BIT Numerical Mathematics, 57(3):607, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0681-x;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0681-x.pdf.

Sadkane:2017:AMM

[3318] Miloud Sadkane and Roger B. Sidje.

An alternating maximization method for approximating the hump of the matrix exponential. BIT Numerical Mathematics, 57(3):609–628, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-016-0644-7.

Hallik:2017:QLR

[3319] Helle Hallik and Peeter Oja. Quadratic/ linear rational spline histopolation. BIT Numerical Mathematics, 57(3): 629–648, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0645-
1.

Zhao:2017:UAM

[3320] Xiaofei Zhao. Uniformly accurate multiscale time integrators for second order oscillatory differential equations with large initial data. BIT Numerical Mathematics, 57(3):649–683, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0646-0.

Yeganeh:2017:SDS

[3321] S. Yeganeh, R. Mokhtari, and J. S. Hesthaven. Space-dependent source determination in a time-fractional diffusion equation using a local discontinuous Galerkin method. BIT Numerical Mathematics, 57(3):685–707, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0648-y.

Luo:2017:MCM

[3322] Xingjun Luo, Chunmei Zeng, Suhang Yang, and Rong Zhang. Multiscale collocation methods for ill-posed integral equations with a modified posteriori parameter selection. BIT Numerical Mathematics, 57(3):709–730, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0649-x.

Hu:2017:DDS

[3323] Guang-Da Hu and Taketomo Mitsui. Delay-dependent stability of numerical methods for delay differential systems of neutral type. BIT Numerical Mathematics, 57(3):731–752, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0650-4;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0650-4.pdf.

Gauckler:2017:SPD

[3324] Ludwig Gauckler and Harry Yserentant.

Structure-preserving discretization of the chemical master equation. BIT Numerical Mathematics, 57(3):753–770, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0651-3.

Franz:2017:CLP

[3325] Sebastian Franz. Convergence of local projection stabilisation finite element methods for convection-diffusion problems on layer-adapted meshes. BIT Numerical Mathematics, 57(3):771–786, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0652-2.

Maalqvist:2017:FEC

[3326] Axel Målqvist and Tony Stillfjord. Finite element convergence analysis for the thermoviscoelastic Joule heating problem. BIT Numerical Mathematics, 57(3):787–810, September 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0653-
1;   https://link.springer.com/
content/pdf/10.1007/s10543-017-
0653-1.pdf.

Mehl:2017:SRR

[3327] Christian Mehl, Volker Mehrmann, and Punit Sharma. Stability radii for real linear Hamiltonian systems with perturbed dissipation. BIT Numerical Mathematics, 57(3):811–843, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0654-0.

Tan:2017:CMI

[3328] Guangning Tan, Nedialko S. Nedialkov, and John D. Pryce. Conversion methods for improving structural analysis of differential–algebraic equation systems. BIT Numerical Mathematics, 57(3):845–865, September 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0655-
z.

Muntingh:2017:SER

[3329] Georg Muntingh. Symbols and exact regularity of symmetric pseudosplines of any arity. BIT Numerical Mathematics, 57(3):867–900, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0656-y.

Montanher:2017:UIU

[3330] Tiago Montanher, Ferenc Domes, and Hermann Schichl. Using interval unions to solve linear systems of equations with uncertainties. BIT Numerical Mathematics, 57(3):901–926, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0657-x;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0657-x.pdf.

Lange:2017:EES

[3331] Marko Lange and Siegfried M. Rump.

Error estimates for the summation of real numbers with application to floating-point summation. BIT Numerical Mathematics, 57(3):927–941, September 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0658-9.

Gao:2017:GFC

[3332] Jing Gao and Arieh Iserles. A generalization of Filon–Clenshaw–Curtis quadrature for highly oscillatory integrals. BIT Numerical Mathematics, 57(4):943–961, December 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0682-
9.

Lang:2017:MSS

[3333] Annika Lang, Andreas Petersson, and Andreas Thalhammer. Mean-square stability analysis of approximations of stochastic differential equations in infinite dimensions. BIT Numerical Mathematics, 57(4):963–990, December 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0684-7;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0684-7.pdf.

Kemmochi:2017:EDN

[3334] Tomoya Kemmochi. Energy dissipative numerical schemes for gradient flows of planar curves. BIT Numerical Mathematics, 57(4):991–1017, December 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0685-6.

Fenu:2017:GTR

[3335] Caterina Fenu, Lothar Reichel, and Giuseppe Rodriguez. GCV for Tikhonov regularization by partial SVD. BIT Numerical Mathematics, 57 (4):1019–1039, December 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0662-
0.

Mirzaei:2017:DAS

[3336] Davoud Mirzaei. Direct approximation on spheres using generalized moving least squares. BIT Numerical Mathematics, 57(4):1041–1063, December 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0659-8.

Xie:2017:NME

[3337] Huiqing Xie. A new method for eigenvector derivatives of a quadratic eigenvalue problem. BIT Numerical Mathematics, 57(4):1065–1082, December 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0680-y.

Liu:2017:SBE

[3338] Ching-Sung Liu and Che-Rung Lee.

On the structured backward error of inexact Arnoldi methods for (skew)-Hermitian and (skew)- symmetric eigenvalue problems. BIT Numerical Mathematics, 57(4):1083– 1108, December 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0660-
2.

Skalna:2017:NMC

[3339] Iwona Skalna and Milan Hladík. A new method for computing a p-solution to parametric interval linear systems with affine-linear and nonlinear dependencies. BIT Numerical Mathematics, 57 (4):1109–1136, December 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0679-
4.

Nakatsukasa:2017:ASV

[3340] Yuji Nakatsukasa. Accuracy of singular vectors obtained by projectionbased SVD methods. BIT Numerical Mathematics, 57(4):1137–1152, December 2017. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0665-x;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0665-x.pdf.

Burkotova:2017:SBN

[3341] Jana Burkotová, Irena Rachunková, and Ewa B. Weinmüller. On singular BVPs with nonsmooth data: convergence of the collocation schemes. BIT Numerical Mathematics, 57(4): 1153–1184, December 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0686-
5.

Zulian:2017:PFE

[3342] Patrick Zulian, Teseo Schneider, Kai Hormann, and Rolf Krause. Parametric finite elements with bijective mappings. BIT Numerical Mathematics, 57 (4):1185–1203, December 2017. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0669-
6.

Kreiss:2018:P

[3343] Gunilla Kreiss and Lars Eldén. Preface. BIT Numerical Mathematics, 58(1):1– 3, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0700-6;   https:
//link.springer.com/content/pdf/
10.1007/s10543-018-0700-6.pdf.

Asadzadeh:2018:PEE

[3344] Mohammad Asadzadeh and Christoffer Standar. A posteriori error estimates for the one and one-half dimensional relativistic Vlasov–Maxwell system. BIT Numerical Mathematics, 58(1):5– 26, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0666-9;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0666-9.pdf.

Carnicer:2018:MPI

[3345] J. M. Carnicer and C. Godés. Multivariate polynomial interpolation using even and odd polynomials. BIT Numerical Mathematics, 58(1):27–49, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0674-9.

Falcao:2018:ESR

[3346] M. Irene Falcão, Fernando Miranda, and Ricardo Severino. Evaluation schemes in the ring of quaternionic polynomials. BIT Numerical Mathematics, 58(1):51–72, March 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0667-
8.

Ignat:2018:SMA

[3347] Liviu I. Ignat and Alejandro Pozo.

A splitting method for the augmented Burgers equation. BIT Numerical Mathematics, 58(1):73–102, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0673-x.

Kumar:2018:ASD

[3348] Rakesh Kumar. Adaptive semidiscrete formulation of BSQI–WENO scheme for the modified Burgers’ equation. BIT Numerical Mathematics, 58(1):103–132, March 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-017-0675-
                                        8.

Mele:2018:RTI

[3349] Giampaolo Mele and Elias Jarlebring. On restarting the tensor infinite Arnoldi method. BIT Numerical Mathematics, 58(1):133–162, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0671-z;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-017-0671-z.pdf.

Niu:2018:EPS

[3350] Xinyan Niu, Jianbo Cui, Jialin Hong, and Zhihui Liu. Explicit pseudosymplectic methods for stochastic Hamiltonian systems. BIT Numerical Mathematics, 58(1):163–178, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0668-7.

Notaris:2018:AGQ

[3351] Sotirios E. Notaris. Anti-Gaussian quadrature formulae based on the zeros of Stieltjes polynomials. BIT Numerical Mathematics, 58(1):179– 198, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0672-y.

Okayama:2018:TAS

[3352] Tomoaki Okayama. Theoretical analysis of sinc-collocation methods and Sinc–Nyström methods for systems of initial value problems. BIT Numerical Mathematics, 58(1):199–220, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0663-z.

Zhang:2018:PES

[3353] Jin Zhang and Xiaowei Liu. Pointwise estimates of SDFEM on Shishkin triangular meshes for problems with exponential layers. BIT Numerical Mathematics, 58(1):221–246, March 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https://link.

springer.com/article/10.1007/s10543-
017-0661-1.

Risebro:2018:CMM

[3354] Nils Henrik Risebro, Christoph Schwab, and Franziska Weber. Correction to: Multilevel Monte Carlo fronttracking for random scalar conservation laws. BIT Numerical Mathematics, 58(1):247–255, March 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-017-0670-
                                        0;   https://link.springer.com/
                                        content/pdf/10.1007/s10543-017-

0670-0.pdf. See [3234].

Anmarkrud:2018:GOC

[3355] Sverre Anmarkrud, Kristian Debrabant, and Anne Kværnø. General order conditions for stochastic partitioned Runge–Kutta methods. BIT Numerical Mathematics, 58(2):257– 280, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0693-6.

Fasino:2018:GNI

[3356] Dario Fasino and Antonio Fazzi. A Gauss–Newton iteration for total least squares problems. BIT Numerical Mathematics, 58(2):281– 299, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0678-5.

Guzel:2018:TFI

[3357] Ahmet Guzel and William Layton. Time filters increase accuracy of the fully implicit method. BIT Numerical Mathematics, 58(2):301– 315, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-018-0695-z.

Hrycak:2018:ECP

[3358] Tomasz Hrycak and Sebastian Schmutzhard. Evaluation of Chebyshev polynomials by a three-term recurrence in floating-point arithmetic. BIT Numerical Mathematics, 58(2):317–330, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0683-8.

Klinge:2018:OZS

[3359] Marcel Klinge, Rüdiger Weiner, and Helmut Podhaisky. Optimally zero stable explicit peer methods with variable nodes. BIT Numerical Mathematics, 58(2):331–345, June 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0691-
8.

Lee:2018:RTF

[3360] Jeonghun J. Lee. Robust three-field finite element methods for Biot’s consolidation model in poroelasticity. BIT Numerical Mathematics, 58(2):347– 372, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0688-3.

Nannen:2018:CSR

[3361] Lothar Nannen and Markus Wess.

Computing scattering resonances using perfectly matched layers with frequency dependent scaling functions. BIT Numerical Mathematics, 58(2): 373–395, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0694-0;   https:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-018-0694-0.pdf.

Norton:2018:CSG

[3362] Terence J. T. Norton. Composite symmetric general linear methods (COSY-GLMs) for the long-time integration of reversible Hamiltonian systems. BIT Numerical Mathematics, 58(2):397– 421, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0692-7;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-017-0692-7.pdf.

Rahimian:2018:UEL

[3363] Abtin Rahimian, Alex Barnett, and Denis Zorin. Ubiquitous evaluation of layer potentials using quadrature by kernel–independent expansion. BIT Numerical Mathematics, 58(2):423– 456, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0689-2.

Schmitt:2018:LHT

[3364] Jeremy Schmitt, Tatiana Shingel, and Melvin Leok. Lagrangian and Hamiltonian Taylor variational integrators. BIT Numerical Mathematics, 58(2): 457–488, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-017-0690-9.

Strachota:2018:EEF

[3365] Pavel Strachota and Michal Benes. Error estimate of the finite volume scheme for the Allen–Cahn equation. BIT Numerical Mathematics, 58(2): 489–507, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0687-4.

Wahlsten:2018:EUG

[3366] Markus Wahlsten and Jan Nordström.

The effect of uncertain geometries on advection–diffusion of scalar quantities. BIT Numerical Mathematics, 58(2): 509–529, June 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-017-0676-7;   https:
//link.springer.com/content/pdf/
10.1007/s10543-017-0676-7.pdf.

Dagnino:2018:EQI

[3367] Catterina Dagnino and Sara Remogna.

Erratum to: Quasi-interpolation based on the ZP-element for the numerical solution of integral equations on surfaces in R^{3}. BIT Numerical Mathematics, 58(2):531, June 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-017-0677-
6;   https://link.springer.com/
content/pdf/10.1007/s10543-017-

0677-6.pdf. See [3307].

Ahmadinia:2018:ALD

[3368] M. Ahmadinia, Z. Safari, and S. Fouladi. Analysis of local discontinuous Galerkin method for time-space fractional convection-diffusion equations. BIT Numerical Mathematics, 58(3):533–554, September 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-018-0697-
                                        x.

Bentbib:2018:SML

[3369] A. H. Bentbib, M. El Guide, K. Jbilou, E. Onunwor, and L. Reichel. Solution methods for linear discrete illposed problems for color image restoration. BIT Numerical Mathematics, 58(3):555–576, September 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-018-0706-

0. See correction [3370].

Bentbib:2018:CSM

[3370] A. H. Bentbib, M. El Guide, K. Jbilou, E. Onunwor, and L. Reichel. Correction to: Solution methods for linear discrete ill-posed problems for color image restoration. BIT Numerical Mathematics, 58(3):577– 578, September 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-018-0708-
                                        y;   https://link.springer.com/
                                        content/pdf/10.1007/s10543-018-

0708-y.pdf. See [3369].

Bihlo:2018:WBM

[3371] Alexander Bihlo and Scott MacLachlan. Well-balanced mesh-based and meshless schemes for the shallowwater equations. BIT Numerical Mathematics, 58(3):579–598, September 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0696-y.

Chen:2018:RPB

[3372] Yanmei Chen, Xiaofei Peng, and Wen Li. Relative perturbation bounds for eigenpairs of diagonalizable matrices. BIT Numerical Mathematics, 58(3):599–612, September 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0701-
5.

Guessab:2018:NIU

[3373] Allal Guessab and Boris Semisalov.

Numerical integration using integrals over hyperplane sections of simplices in a triangulation of a polytope. BIT Numerical Mathematics, 58(3):613–660, September 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0703-3.

Huang:2018:OEA

[3374] Chaobao Huang, Martin Stynes, and Na An. Optimal L^{∞}(L^{2}) error analysis of a direct discontinuous Galerkin method for a time-fractional reaction–diffusion problem. BIT Numerical Mathematics, 58(3):661–690, September 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0707-z.

Kopecz:2018:UPC

[3375] Stefan Kopecz and Andreas Meister. Unconditionally positive and conservative third order modified Patankar–Runge–Kutta discretizations of production-destruction systems. BIT Numerical Mathematics, 58(3): 691–728, September 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0705-
1.

Lin:2018:EPO

[3376] Xue-Lei Lin, Michael K. Ng, and Hai-Wei Sun. Efficient preconditioner of one-sided space fractional diffusion equation. BIT Numerical Mathematics, 58(3):729–748, September 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0699-
8.

Steyer:2018:LSS

[3377] Andrew J. Steyer and Erik S. Van Vleck. A Lyapunov and Sacker–Sell spectral stability theory for one-step methods. BIT Numerical Mathematics, 58(3):749–781, September 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0704-
2.

Wang:2018:EFO

[3378] Pengde Wang and Chengming Huang. An efficient fourth-order in space difference scheme for the nonlinear fractional Ginzburg–Landau equation. BIT Numerical Mathematics, 58(3):783–805, September 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0698-9.

Zhu:2018:PSD

[3379] Shuai Zhu and Shilie Weng. A parallel spectral deferred correction method for first-order evolution problems. BIT Numerical Mathematics, 58(3):807–834, September 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0702-
4.

Alqahtani:2018:MOP

[3380] Hessah Alqahtani and Lothar Reichel. Multiple orthogonal polynomials applied to matrix function evaluation. BIT Numerical Mathematics, 58(4):835–849, December 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0709-
x.

Babb:2018:APS

[3381] Tracy Babb, Adrianna Gillman, Sijia Hao, and Per-Gunnar Martinsson. An accelerated Poisson solver based on multidomain spectral discretization. BIT Numerical Mathematics, 58(4):851–879, December 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0714-
0;   https://link.springer.com/
content/pdf/10.1007/s10543-018-
0714-0.pdf.

Bolin:2018:WCG

[3382] David Bolin, Kristin Kirchner, and Mihály Kovács. Weak convergence of Galerkin approximations for fractional elliptic stochastic PDEs with spatial white noise. BIT Numerical Mathematics, 58(4):881–906, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0719-8;   https:
//link.springer.com/content/pdf/
10.1007/s10543-018-0719-8.pdf.

Caliari:2018:BEA

[3383] Marco Caliari, Peter Kandolf, and Franco Zivcovich. Backward error analysis of polynomial approximations for computing the action of the matrix exponential. BIT Numerical Mathematics, 58(4):907–935, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0718-9.

Ekstrom:2018:EFM

[3384] Sven-Erik Ekström, Isabella Furci, and Stefano Serra-Capizzano. Exact formulae and matrix-less eigensolvers for block banded symmetric Toeplitz matrices. BIT Numerical Mathematics, 58(4):937–968, December 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0715-
z;   https://link.springer.com/
content/pdf/10.1007/s10543-018-
0715-z.pdf.

Hairer:2018:EBB

[3385] Ernst Hairer and Christian Lubich. Energy behaviour of the Boris method for charged-particle dynamics. BIT Numerical Mathematics, 58(4):969–979, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0713-1.

Hashemi:2018:SPT

[3386] Behnam Hashemi and Yuji Nakatsukasa. On the spectral problem for trivariate functions. BIT Numerical Mathematics, 58(4):981–1008, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0710-4.

Holm:2018:SDH

[3387] Darryl D. Holm and Tomasz M. Tyranowski. Stochastic discrete Hamiltonian variational integrators. BIT Numerical Mathematics, 58(4):1009–1048, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0720-2;   https:
//link.springer.com/content/pdf/
10.1007/s10543-018-0720-2.pdf.

Kunkel:2018:RSD

[3388] Peter Kunkel and Volker Mehrmann.

Regular solutions of DAE hybrid systems and regularization techniques. BIT Numerical Mathematics, 58(4): 1049–1077, December 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0712-
2.

Lorenz:2018:RKM

[3389] Dirk A. Lorenz, Sean Rose, and Frank Schöpfer. The randomized Kaczmarz method with mismatched adjoint. BIT Numerical Mathematics, 58 (4):1079–1098, December 2018. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-018-0717-
x.

Oksa:2018:AQC

[3390] Gabriel Oksa, Yusaku Yamamoto, Martin Becka, and Marián Vajtersic. Asymptotic quadratic convergence of the parallel block-Jacobi EVD algorithm with dynamic ordering for Hermitian matrices. BIT Numerical Mathematics, 58(4):1099–1123, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0711-3.

Soderlind:2018:ZSM

[3391] Gustaf Söderlind, Imre Fekete, and István Faragó. On the zero-stability of multistep methods on smooth nonuniform grids. BIT Numerical Mathematics, 58(4):1125–1143, December 2018. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-018-0716-y;   https:
//link.springer.com/content/pdf/
10.1007/s10543-018-0716-y.pdf.

Bai:2019:BSI

[3392] Zhong-Zhi Bai and Kang-Ya Lu. On banded M-splitting iteration methods for solving discretized spatial fractional diffusion equations. BIT Numerical Mathematics, 59(1):1–33, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0727-8.

Cuyt:2019:MDF

[3393] Annie Cuyt and Oliver Salazar Celis.

Multivariate data fitting with error control. BIT Numerical Mathematics, 59(1):35–55, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0721-1.

Eirola:2019:KIH

[3394] Timo Eirola and Antti Koskela. Krylov integrators for Hamiltonian systems. BIT Numerical Mathematics, 59(1):57– 76, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-018-0732-y;   https:
//link.springer.com/content/pdf/
10.1007/s10543-018-0732-y.pdf.

Guo:2019:RTT

[3395] Ping Guo and Chong-Jun Li. Razumikhintype technique on stability of exact and numerical solutions for the nonlinear stochastic pantograph differential equations. BIT Numerical Mathematics, 59(1):77–96, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-018-
                                        0723-z.

Gustafsson:2019:SFE

[3396] Tom Gustafsson, Rolf Stenberg, and Juha Videman. A stabilised finite element method for the plate obstacle problem. BIT Numerical Mathematics, 59(1):97–124, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-018-
                                        0728-7.

Huhtanen:2019:OQS

[3397] Marko Huhtanen and Vesa Kotila. Optimal quotients for solving large eigenvalue problems. BIT Numerical Mathematics, 59(1):125–154, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-018-
                                        0725-x.

Majidian:2019:SAF

[3398] Hassan Majidian. Stable application of Filon–Clenshaw–Curtis rules to singular oscillatory integrals by exponential transformations. BIT Numerical Mathematics, 59(1):155–181, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-018-
                                        0730-0.

Qasemi:2019:TFD

[3399] Samaneh Qasemi, Davood Rostamy, and Nazdar Abdollahi. The timefractional diffusion inverse problem subject to an extra measurement by a local discontinuous Galerkin method. BIT Numerical Mathematics, 59(1):183–212, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0731-z.

Sato:2019:SCC

[3400] Shun Sato. Stability and convergence of a conservative finite difference scheme for the modified Hunter– Saxton equation. BIT Numerical Mathematics, 59(1):213–241, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0726-9.

Soleymani:2019:PMA

[3401] Fazlollah Soleymani. Pricing multiasset option problems: a Chebyshev pseudo-spectral method. BIT Numerical Mathematics, 59(1):243– 270, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-018-0722-0.

Wang:2019:FAB

[3402] Min Wang and Yangfeng Su. A further analysis of backward error in polynomial deflation. BIT Numerical Mathematics, 59(1):271–297, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0724-y.

Yang:2019:MRH

[3403] Ai-Li Yang, Yang Cao, and Yu-Jiang Wu. Minimum residual Hermitian and skew-Hermitian splitting iteration method for non-Hermitian positive definite linear systems. BIT Numerical Mathematics, 59(1):299– 319, March 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-018-0729-6.

Axelsson:2019:NVP

[3404] Owe Axelsson and Davod Khojasteh Salkuyeh. A new version of a preconditioning method for certain two-by-two block matrices with square blocks. BIT Numerical Mathematics, 59(2): 321–342, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-018-0741-x.

Cardoso:2019:CMG

[3405] João R. Cardoso and Amir Sadeghi. Computation of matrix gamma function. BIT Numerical Mathematics, 59(2):343–370, June 2019. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-018-
                                        00744-1.

Carnicer:2019:CON

[3406] J. M. Carnicer, Y. Khiar, and J. M. Peña. Central orderings for the Newton interpolation formula. BIT Numerical Mathematics, 59(2):371– 386, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-018-00743-2.

Haddock:2019:MMI

[3407] Jamie Haddock and Deanna Needell.

On Motzkin’s method for inconsistent linear systems. BIT Numerical Mathematics, 59(2):387–401, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0737-6.

Hrycak:2019:AEC

[3408] Tomasz Hrycak and Sebastian Schmutzhard.

Accurate evaluation of Chebyshev polynomials in floating-point arithmetic. BIT Numerical Mathematics, 59(2):403–416, June 2019. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0738-5.

Javanbakht:2019:NCB

[3409] Masoumeh Javanbakht and Tomas Sauer. Numerical computation of H-bases. BIT Numerical Mathematics, 59(2):417–442, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0733-x.

Maier:2019:ECW

[3410] Roland Maier and Daniel Peterseim.

Explicit computational wave propagation in micro-heterogeneous media. BIT Numerical Mathematics, 59(2): 443–462, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-018-0735-8.

Mazza:2019:SPF

[3411] M. Mazza and J. Pestana. Spectral properties of flipped Toeplitz matrices and related preconditioning. BIT Numerical Mathematics, 59(2):463– 482, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-018-0740-y;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-018-0740-y.pdf.

Montijano:2019:FSS

[3412] J. I. Montijano, H. Podhaisky, L. Rández, and M. Calvo. A family of L-stable singly implicit peer methods for solving stiff IVPs. BIT Numerical Mathematics, 59(2):483– 502, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-018-0734-9.

Popova:2019:ROI

[3413] Evgenija D. Popova. Rank one interval enclosure of the parametric united solution set. BIT Numerical Mathematics, 59(2):503–521, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-018-
                                        0739-4.

Ramezani:2019:ERH

[3414] Z. Ramezani and F. Toutounian. Extended and rational Hessenberg methods for the evaluation of matrix functions. BIT Numerical Mathematics, 59(2):523–545, June 2019. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0742-9.

Ranocha:2019:MPD

[3415] Hendrik Ranocha. Mimetic properties of difference operators: product and chain rules as for functions of bounded variation and entropy stability of second derivatives. BIT Numerical Mathematics, 59(2):547–563, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-018-
0736-7.

Aihara:2019:VRS

[3416] Kensuke Aihara, Ryosuke Komeyama, and Emiko Ishiwata. Variants of residual smoothing with a small residual gap. BIT Numerical Mathematics, 59(3):565–584, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00751-w.

Ait-Haddou:2019:NSR

[3417] Rachid Ait-Haddou. New stability results for explicit Runge–Kutta methods. BIT Numerical Mathematics, 59(3):585–612, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00752-9.

Anton:2019:WBE

[3418] Cristina Anton. Weak backward error analysis for stochastic Hamiltonian systems. BIT Numerical Mathematics, 59(3):613–646, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00747-6.

Blomker:2019:NCP

[3419] Dirk Blömker and Minoo Kamrani. Numerically computable a posterioribounds for the stochastic Allen–Cahn equation. BIT Numerical Mathematics, 59(3):647–673, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00745-8.

Chiang:2019:RIN

[3420] Chun-Yueh Chiang, Matthew M. Lin, and Xiao-Qing Jin. Riemannian inexact Newton method for structured inverse eigenvalue and singular value problems. BIT Numerical Mathematics, 59(3):675–694, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00754-7.

Diao:2019:SGE

[3421] Huai-An Diao and Qing-Le Meng. Structured generalized eigenvalue condition numbers for parameterized quasiseparable matrices. BIT Numerical Mathematics, 59(3):695–720, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00748-5.

Gazzola:2019:FGT

[3422] Silvia Gazzola and Malena Sabaté Landman. Flexible GMRES for total variation regularization. BIT Numerical Mathematics, 59(3):721–746, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00750-x;   https:
//link.springer.com/content/pdf/
10.1007/s10543-019-00750-x.pdf.

Li:2019:PPN

[3423] Libo Li and Dai Taguchi. On a positivity preserving numerical scheme for jump-extended CIR process: the alphastable case. BIT Numerical Mathematics, 59(3):747–774, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00753-8.

Ridder:2019:CFD

[3424] J. Ridder and A. M. Ruf. A convergent finite difference scheme for the Ostrovsky–Hunter equation with Dirichlet boundary conditions. BIT Numerical Mathematics, 59(3): 775–796, September 2019. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00746-7.

Xu:2019:BDN

[3425] Shixin Xu, Minxin Chen, Chun Liu, Ran Zhang, and Xingye Yue. Behavior of different numerical schemes for random genetic drift. BIT Numerical Mathematics, 59(3):797–821, September 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00749-4.

Breiten:2019:RBI

[3426] Tobias Breiten and Emil Ringh. Residual-based iterations for the generalized Lyapunov equation. BIT Numerical Mathematics, 59(4):823–852, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00760-9;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10543-019-00760-9.pdf.

Du:2019:SAE

[3427] Jie Du, Yang Yang, and Eric Chung. Stability analysis and error estimates of local discontinuous Galerkin methods for convection–diffusion equations on overlapping meshes. BIT Numerical Mathematics, 59(4):853–876, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00757-4.

Karvonen:2019:GKQ

[3428] Toni Karvonen and Simo Särkkä. Gaussian kernel quadrature at scaled Gauss–Hermite nodes. BIT Numerical Mathematics, 59(4):877–902, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00758-3;   https:
//link.springer.com/content/pdf/
10.1007/s10543-019-00758-3.pdf.

Kongskov:2019:DTG

[3429] Rasmus Dalgas Kongskov, Yiqiu Dong, and Kim Knudsen. Directional total generalized variation regularization. BIT Numerical Mathematics, 59(4):903–928, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00755-6.

Kumar:2019:MAV

[3430] Chaman Kumar and Sotirios Sabanis. On Milstein approximations with varying coefficients: the case of super-linear diffusion coefficients. BIT Numerical Mathematics, 59(4):929–968, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00756-5;   https://link.springer.
com/content/pdf/10.1007/s10543-
019-00756-5.pdf.  See correction

[3455].

Pozza:2019:IAR

[3431] Stefano Pozza and Valeria Simoncini.

Inexact Arnoldi residual estimates and decay properties for functions of non-Hermitian matrices. BIT Numerical Mathematics, 59(4):969–986, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00763-6.

Reisinger:2019:SEA

[3432] Christoph Reisinger and Zhenru Wang. Stability and error analysis of an implicit Milstein finite difference scheme for a two-dimensional Zakai SPDE. BIT Numerical Mathematics, 59(4): 987–1029, December 2019. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00761-8;   https://link.springer.
                                        com/content/pdf/10.1007/s10543-
                                        019-00761-8.pdf.

Renaut:2019:UPR

[3433] Rosemary A. Renaut, Anthony W. Helmstetter, and Saeed Vatankhah. Unbiased predictive risk estimation of the Tikhonov regularization parameter: convergence with increasing rank approximations of the singular value decomposition. BIT Numerical Mathematics, 59(4):1031–1061, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00762-7.

Sato:2019:LGS

[3434] Shun Sato. Linear gradient structures and discrete gradient methods for conservative/dissipative differential– algebraic equations. BIT Numerical Mathematics, 59(4):1063–1091, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00759-2.

Sun:2019:CCW

[3435] Jiguang Sun, Qian Zhang, and Zhimin Zhang. A curl-conforming weak Galerkin method for the quad-curl problem. BIT Numerical Mathematics, 59(4):1093–1114, December 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00764-5.

Cai:2020:PAE

[3436] Yunfeng Cai, Zhigang Jia, and Zheng-Jian Bai. Perturbation analysis of an eigenvector-dependent nonlinear eigenvalue problem with applications. BIT Numerical Mathematics, 60(1):1– 29, March 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00765-4.

Chandhini:2020:DIR

[3437] G. Chandhini, K. S. Prashanthi, and V. Antony Vijesh. Direct and integrated radial functions based quasilinearization schemes for nonlinear fractional differential equations. BIT Numerical Mathematics, 60(1):31–65, March 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00766-3.

Kettil:2020:NUD

[3438] G. Kettil, A. Målqvist, A. Mark, M. Fredlund, K. Wester, and F. Edelvik.

Numerical upscaling of discrete network models. BIT Numerical Mathematics, 60(1):67–92, March 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00767-2;   https://link.springer.
com/content/pdf/10.1007/s10543-
019-00767-2.pdf.

Akrivis:2020:MAV

[3439] Georgios Akrivis and Emmanouil Katsoprinakis. Maximum angles of A(ϑ)- stability of backward difference formulae. BIT Numerical Mathematics, 60(1):93–99, March 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00768-1.

Sabermahani:2020:FOG

[3440] Sedigheh Sabermahani, Yadollah Ordokhani, and Sohrab Ali Yousefi. Fractional-order general Lagrange scaling functions and their applications. BIT Numerical Mathematics, 60(1):101–128, March 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00769-0.

Bosner:2020:ACS

[3441] Tina Bosner, Bojan Crnković, and Jerko Skifić. Application of CCC– Schoenberg operators on image resampling. BIT Numerical Mathematics, 60(1):129–155, March 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00770-7.

Jawecki:2020:CUE

[3442] Tobias Jawecki, Winfried Auzinger, and Othmar Koch. Computable upper error bounds for Krylov approximations to matrix exponentials and associated φ-functions. BIT Numerical Mathematics, 60 (1):157–197, March 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00771-6;   https://link.springer.
com/content/pdf/10.1007/s10543-
019-00771-6.pdf.

Deriaz:2020:CFD

[3443] Erwan Deriaz. Compact finite difference schemes of arbitrary order for the Poisson equation in arbitrary dimensions. BIT Numerical Mathematics, 60(1):199–233, March 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00772-5.

Burman:2020:DBV

[3444] Erik Burman, Peter Hansbo, and Mats G. Larson. Dirichlet boundary value correction using Lagrange multipliers. BIT Numerical Mathematics, 60(1):235–260, March 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00773-4;   https://link.springer.

                                        com/content/pdf/10.1007/s10543-
                                        019-00773-4.pdf.

Altintan:2020:HME

[3445] Derya Altintan and Heinz Koeppl. Hybrid master equation for jump-diffusion approximation of biomolecular reaction networks. BIT Numerical Mathematics, 60(2):261–294, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00781-4;   https://link.springer.
                                        com/content/pdf/10.1007/s10543-
                                        019-00781-4.pdf.

Campos:2020:PJD

[3446] Carmen Campos and Jose E. Roman. A polynomial Jacobi–Davidson solver with support for non-monomial bases and deflation. BIT Numerical Mathematics, 60(2):295–318, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00778-z.

Chand:2020:AZF

[3447] A. K. B. Chand, N. Vijender, P. Viswanathan, and A. V. Tetenov. Affine zipper fractal interpolation functions. BIT Numerical Mathematics, 60(2):319–344, June 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10543-019-
                                        00774-3.

Christiansen:2020:PFI

[3448] Snorre H. Christiansen and Martin W. Licht. Poincaré=-Friedrichs inequalities of complexes of discrete distributional differential forms. BIT Numerical Mathematics, 60(2):345– 371, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00784-1.

Gillette:2020:NFE

[3449] Andrew Gillette, Kaibo Hu, and Shuo Zhang. Nonstandard finite element de Rham complexes on cubical meshes. BIT Numerical Mathematics, 60(2): 373–409, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00779-y.

Gracia:2020:CAF

[3450] José Luis Gracia, Eugene O’Riordan, and Martin Stynes. Convergence analysis of a finite difference scheme for a two-point boundary value problem with a Riemann–Liouville– Caputo fractional derivative. BIT Numerical Mathematics, 60(2):411– 439, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//link.springer.com/article/10.
1007/s10543-019-00777-0.

DeMarchi:2020:JVS

[3451] S. De Marchi, F. Marchetti, and E. Perracchione. Jumping with variably scaled discontinuous kernels (VS-DKs). BIT Numerical Mathematics, 60(2):441–463, June 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10543-019-
                                        00786-z;   https://link.springer.
                                        com/content/pdf/10.1007/s10543-
                                        019-00786-z.pdf.  See correction

[3456].

Porcu:2020:OCG

[3452] Roberto Porcù and Maurizio M. Chiaramonte. Orthogonality constrained gradient reconstruction for superconvergent linear functionals. BIT Numerical Mathematics, 60(2):465– 479, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00775-2.

Srivastava:2020:LPS

[3453] Shweta Srivastava and Sashikumaar Ganesan. Local projection stabilization with discontinuous Galerkin method in time applied to convection dominated problems in time-dependent domains. BIT Numerical Mathematics, 60(2): 481–507, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00783-2.

Tu:2020:NHO

[3454] Xiongbiao Tu, Ander Murua, and Yifa Tang. New high order symplectic integrators via generating functions with its application in many-body problem. BIT Numerical Mathematics, 60(2): 509–535, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00785-0.

Kumar:2020:CMA

[3455] Chaman Kumar and Sotirios Sabanis. Correction to: On Milstein approximations with varying coefficients: the case of super-linear diffusion coefficients. BIT Numerical Mathematics, 60(2):537, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00780-5;   https://link.springer.
com/content/pdf/10.1007/s10543-
019-00780-5.pdf. See [3430].

DeMarchi:2020:CJV

[3456] S. De Marchi, F. Marchetti, and E. Perracchione. Correction to: Jumping with variably scaled discontinuous kernels (VSDKs). BIT Numerical Mathematics, 60(2):539–540, June 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-020-
00800-9;   https://link.springer.
com/content/pdf/10.1007/s10543-
020-00800-9.pdf. See [3451].

Pettorossi:2020:CTH

[3457] Alberto Pettorossi. Correction to:

Towers of Hanoi problems: Deriving iterative solutions by program transformations. BIT Numerical Mathematics, 60(2):541, June 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10543-019-
00790-3;   https://link.springer.
com/content/pdf/10.1007/s10543-
019-00790-3.pdf. See [1455].

Brehier:2020:WCR

[3458] Charles-Edouard Bréhier and Ludovic Goudenège. Weak convergence rates of splitting schemes for the stochastic Allen–Cahn equation. BIT Numerical Mathematics, 60(3):543–582, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00788-x.

Cartis:2020:MEN

[3459] Coralia Cartis, Nicholas I. M. Gould, and Marius Lange. On monotonic estimates of the norm of the minimizers of regularized quadratic functions in Krylov spaces. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00791-2.

Ceruti:2020:TIS

[3460] Gianluca Ceruti and Christian Lubich. Time integration of symmetric and anti-symmetric low-rank matrices and Tucker tensors. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-019-00799-8.

Gander:2020:EBS

[3461] Martin J. Gander and Gerhard Wanner. Exact BDF stability angles with Maple. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00796-x.

Giesselmann:2020:PEA

[3462] Jan Giesselmann, Fabian Meyer, and Christian Rohde. A posteriori error analysis and adaptive nonintrusive numerical schemes for systems of random conservation laws. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00794-z.

Kahl:2020:ALF

[3463] Karsten Kahl and Nils Kintscher. Automated local Fourier analysis (aLFA). BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00797-w.

Mazure:2020:GCP

[3464] Marie-Laurence Mazure. Geometrically continuous piecewise Chebyshevian NU(R)BS. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00795-y.

Padgett:2020:AAF

[3465] Joshua L. Padgett. Analysis of an approximation to a fractional extension problem. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00787-y.

Viviani:2020:MVS

[3466] Milo Viviani. A minimal-variable symplectic method for isospectral flows. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00792-1.

Wang:2020:MSC

[3467] Xiaojie Wang, Jiayi Wu, and Bozhang Dong. Mean-square convergence rates of stochastic theta methods for SDEs under a coupled monotonicity condition. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00793-0.

Winters:2020:ESN

[3468] Andrew R. Winters, Christof Czernik, and Gregor J. Gassner. Entropy stable numerical approximations for the isothermal and polytropic Euler equations. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00789-w.

Yao:2020:GGN

[3469] Teng-Teng Yao, Zheng-Jian Bai, and Zhi Zhao. A geometric Gauss– Newton method for least squares inverse eigenvalue problems. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00798-9.

Zhu:2020:CAG

[3470] Shengfeng Zhu, Xianliang Hu, and Qifeng Liao. Convergence analysis of Galerkin finite element approximations to shape gradients in eigenvalue optimization. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-019-00782-3.

Kounchev:2020:CSO

[3471] O. Kounchev, H. Render, and T. Tsachev. On a class of L-splines of order 4: fast algorithms for interpolation and smoothing. BIT Numerical Mathematics, 60(4):879–899, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00801-8.

Trefethen:2020:QIC

[3472] Lloyd N. Trefethen. Quantifying the ill-conditioning of analytic continuation. BIT Numerical Mathematics, 60(4):901–915, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00802-7.

Chen:2020:MAV

[3473] Chuchu Chen, Jialin Hong, and Diancong Jin. Modified averaged vector field methods preserving multiple invariants for conservative stochastic differential equations. BIT Numerical Mathematics, 60(4):917–957, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00803-6.

Debrabant:2020:SMM

[3474] Kristian Debrabant, Giovanni Samaey, and Przemysław Zieliński. Study of micro-macro acceleration schemes for linear slow-fast stochastic differential equations with additive noise. BIT Numerical Mathematics, 60(4):959–998, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00804-5.

Lind:2020:SGS

[3475] Martin Lind, Adrian Muntean, and Omar Richardson. A semidiscrete Galerkin scheme for a coupled two-scale elliptic-parabolic system: well-posedness and convergence approximation rates. BIT Numerical Mathematics, 60(4):999–1031, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00805-4.

Ravibabu:2020:LSL

[3476] Mashetti Ravibabu and Arindama Singh. The least squares and line search in extracting eigenpairs in Jacobi–Davidson method. BIT Numerical Mathematics, 60(4):1033–1055, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00806-3.

Becker:2020:LUB

[3477] Sebastian Becker, Benjamin Gess, and Peter E. Kloeden. Lower and upper bounds for strong approximation errors for numerical approximations of stochastic heat equations. BIT Numerical Mathematics, 60(4):1057–1073, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00807-2.

Jesus:2020:NMF

[3478] Carla Jesus and Ercília Sousa. Numerical method with fractional splines for a subdiffusion problem. BIT Numerical Mathematics, 60(4):1075–1111, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00808-1.

Caliari:2020:AME

[3479] Marco Caliari, Fabio Cassini, and Franco Zivcovich. Approximation of the matrix exponential for matrices with a skinny field of values. BIT Numerical Mathematics, 60(4):1113–1131, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00809-0.

Deng:2020:SIE

[3480] Chang-Song Deng and Wei Liu. Semiimplicit Euler–Maruyama method for non-linear time-changed stochastic differential equations. BIT Numerical Mathematics, 60(4):1133–1151, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00810-7.

Musharbash:2020:SDL

[3481] Eleonora Musharbash, Fabio Nobile, and Eva Vidlicková. Symplectic dynamical low rank approximation of wave equations with random parameters. BIT Numerical Mathematics, 60(4):1153–1201, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00811-6.

Fu:2020:VFF

[3482] Taibai Fu, Zhoushun Zheng, and Beiping Duan. Variational formulation for fractional inhomogeneous boundary value problems. BIT Numerical Mathematics, 60(4):1203–1219, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00812-5.

Kurschner:2020:IML

[3483] Patrick Kürschner and Melina A. Freitag. Inexact methods for the low rank solution to large scale Lyapunov equations. BIT Numerical Mathematics, 60(4):1221–1259, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00813-4.

Anonymous:2020:RPS

[3484] Anonymous. Read and publish (Springer Compact) agreements. BIT Numerical Mathematics, 60(4):??, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anonymous:2020:CIP

[3485] Anonymous. COVID-19 and impact on peer review. BIT Numerical Mathematics, 60(4):??, December 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Li:2021:ODD

[3486] Xiao Li, Lili Ju, and Thi-Thao-Phuong Hoang. Overlapping domain decomposition based exponential time differencing methods for semilinear parabolic equations. BIT Numerical Mathematics, 61(1):1–36, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00817-0.

Henning:2021:NOH

[3487] Patrick Henning and Johan Wärnegård.

A note on optimal H^{1}-error estimates for Crank–Nicolson approximations to the nonlinear Schrödinger equation. BIT Numerical Mathematics, 61(1):37– 59, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00814-3.

Baumstark:2021:APT

[3488] Simon Baumstark and Katharina Schratz. Asymptotic preserving trigonometric integrators for the quantum Zakharov system. BIT Numerical Mathematics, 61(1):61–81, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00815-2.

afKlinteberg:2021:AQN

[3489] Ludvig af Klinteberg and Alex H. Barnett. Accurate quadrature of nearly singular line integrals in two and three dimensions by singularity swapping. BIT Numerical Mathematics, 61(1):83– 118, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00820-5.

Eftekhari:2021:ESG

[3490] Armin Eftekhari, Bart Vandereycken, and Konstantinos C. Zygalakis. Explicit stabilised gradient descent for faster strongly convex optimisation. BIT Numerical Mathematics, 61(1):119–139, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00819-y.

Zhang:2021:FDS

[3491] Yabin Zhang and Adrianna Gillman.

A fast direct solver for two dimensional quasi-periodic multilayered media scattering problems. BIT Numerical Mathematics, 61(1):141–171, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00818-z.

Gander:2021:DNW

[3492] Martin J. Gander, Felix Kwok, and Bankim C. Mandal. Dirichlet– Neumann waveform relaxation methods for parabolic and hyperbolic problems in multiple subdomains. BIT Numerical Mathematics, 61(1):173– 207, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00823-2.

Anciaux-Sedrakian:2021:PBA

[3493] A. Anciaux-Sedrakian, L. Grigori, and S. Yousef. An a posteriori-based adaptive preconditioner for controlling a local algebraic error norm. BIT Numerical Mathematics, 61(1):209– 235, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00822-3.

Massei:2021:RKS

[3494] Stefano Massei and Leonardo Robol.

Rational Krylov for Stieltjes matrix functions: convergence and pole selection. BIT Numerical Mathematics, 61(1):237–273, March 2021. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-020-00826-
z.

Chaudhry:2021:EEU

[3495] Jehanzeb H. Chaudhry, Donald Estep, and Simon J. Tavener. Error estimation and uncertainty quantification for first time to a threshold value. BIT Numerical Mathematics, 61(1):275– 307, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00825-0.

Bystricky:2021:AIE

[3496] Lukas Bystricky, Sara Pålsson, and Anna-Karin Tornberg. An accurate integral equation method for Stokes flow with piecewise smooth boundaries. BIT Numerical Mathematics, 61(1):309– 335, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00816-1.

Moorman:2021:RKA

[3497] Jacob D. Moorman, Thomas K. Tu, and Deanna Needell. Randomized Kaczmarz with averaging. BIT Numerical Mathematics, 61(1):337–359, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00824-1.

Lange:2021:VIN

[3498] Marko Lange and Siegfried M.

Rump. Verified inclusions for a nearest matrix of specified rank deficiency via a generalization of Wedin’s sin(θ) theorem. BIT Numerical Mathematics, 61(1):361–380, March 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00827-y.

Debrabant:2021:RKL

[3499] Kristian Debrabant, Anne Kværnø, and Nicky Cordua Mattsson. Runge– Kutta Lawson schemes for stochastic differential equations. BIT Numerical Mathematics, 61(2):381–409, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00839-8.

Gocwin:2021:OAQ

[3500] Maciej Goćwin. On optimal adaptive quadratures for automatic integration. BIT Numerical Mathematics, 61(2): 411–439, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00831-2.

Grunert:2021:NCS

[3501] Katrin Grunert, Anders Nordli, and Susanne Solem. Numerical conservative solutions of the Hunter–Saxton equation. BIT Numerical Mathematics, 61(2):441–471, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00835-y.

Hashimoto:2021:IRK

[3502] Yuka Hashimoto and Takashi Nodera. Inexact rational Krylov method for evolution equations. BIT Numerical Mathematics, 61(2):473–502, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00829-w.

Ito:2021:ABE

[3503] Shin ichi Ito, Takeru Matsuda, and Yuto Miyatake. Adjoint-based exact Hessian computation. BIT Numerical Mathematics, 61(2):503– 522, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00833-0.

Lasserre:2021:SFI

[3504] Jean B. Lasserre. Simple formula for integration of polynomials on a simplex. BIT Numerical Mathematics, 61(2): 523–533, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00828-x.

Luan:2021:EER

[3505] Vu Thai Luan. Efficient exponential Runge–Kutta methods of high order: construction and implementation. BIT Numerical Mathematics, 61(2): 535–560, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00834-z.

Pessia:2021:TTR

[3506] Alberto Pessia and Jing Tang. A threeterm recurrence relation for accurate evaluation of transition probabilities of the simple birth-and-death process. BIT Numerical Mathematics, 61(2): 561–585, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00836-x.

Saray:2021:AIO

[3507] Behzad Nemati Saray. Abel’s integral operator: sparse representation based on multiwavelets. BIT Numerical Mathematics, 61(2):587– 606, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00832-1.

Xia:2021:ALP

[3508] Jingmin Xia, Patrick E. Farrell, and Florian Wechsung. Augmented Lagrangian preconditioners for the Oseen–Frank model of nematic and cholesteric liquid crystals. BIT Numerical Mathematics, 61(2):607– 644, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00838-9.

Xu:2021:MMG

[3509] Fei Xu, Hehu Xie, and Meiling Yue.

A multigrid method for the ground state solution of Bose–Einstein condensates based on Newton iteration. BIT Numerical Mathematics, 61(2): 645–663, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00830-3.

Zeng:2021:OQE

[3510] Huilan Zeng, Chen-Song Zhang, and Shuo Zhang. Optimal quadratic element on rectangular grids for H^{1} problems. BIT Numerical Mathematics, 61(2):665–689, June 2021. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-020-00821-
4.

Zhou:2021:BBV

[3511] Yongtao Zhou and Martin Stynes. Block boundary value methods for linear weakly singular Volterra integrodifferential equations. BIT Numerical Mathematics, 61(2):691–720, June 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-020-00840-1.

Becher:2021:VTD

[3512] Simon Becher and Gunar Matthies. Variational time discretizations of higher order and higher regularity. BIT Numerical Mathematics, 61(3): 721–755, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-021-00851-
6.

Catinas:2021:NIO

[3513] Teodora Catinas. Nielson interpolation operators on an arbitrary triangle with one curved side. BIT Numerical Mathematics, 61(3):757–770, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00842-7.

Crevat:2021:APS

[3514] Joachim Crevat and Francis Filbet.

Asymptotic preserving schemes for the FitzHugh–Nagumo transport equation with strong local interactions. BIT Numerical Mathematics, 61(3):771–804, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00844-5.

Gonzalez-Pinto:2021:PBM

[3515] S. González-Pinto, E. Hairer, and D. Hernández-Abreu. Power boundedness in the maximum norm of stability matrices for ADI methods. BIT Numerical Mathematics, 61(3):805–827, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00845-4.

Jin:2021:SAT

[3516] Sixian Jin and Kei Kobayashi.

Strong approximation of time-changed stochastic differential equations involving drifts with random and non-random integrators. BIT Numerical Mathematics, 61(3):829–857, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00852-5.

Kim:2021:SPA

[3517] Hyun-Min Kim and Jie Meng. Structured perturbation analysis for an infinite size quasi-Toeplitz matrix equation with applications. BIT Numerical Mathematics, 61(3):859–879, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00847-2.

Kormann:2021:GFH

[3518] Katharina Kormann and Anna Yurova. A generalized Fourier–Hermite method for the Vlasov–Poisson system. BIT Numerical Mathematics, 61(3):881– 909, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-021-00853-
                                        4.

Li:2021:RBK

[3519] Hanyu Li and Yuanyang Zhu. Randomized block Krylov subspace methods for trace and log-determinant estimators. BIT Numerical Mathematics, 61(3):911–939, September 2021. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-021-00850-
                                        7.

Mardal:2021:ADP

[3520] Kent-Andre Mardal, Marie E. Rognes, and Travis B. Thompson. Accurate discretization of poroelasticity without Darcy stability. BIT Numerical Mathematics, 61(3):941–976, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00849-0.

Wang:2021:LTN

[3521] Bin Wang and Xinyuan Wu. A long-term numerical energy-preserving analysis of symmetric and/or symplectic extended RKN integrators for efficiently solving highly oscillatory Hamiltonian systems. BIT Numerical Mathematics, 61(3):977–1004, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00846-3.

Wang:2021:PEE

[3522] Fei Wang, Sheheryar Shah, and Wenqiang Xiao. A priori error estimates of discontinuous Galerkin methods for a quasi-variational inequality. BIT Numerical Mathematics, 61(3):1005–1022, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00848-1.

Wang:2021:HOC

[3523] Yuan-Ming Wang. A high-order compact finite difference method on nonuniform time meshes for variable coefficient reaction-subdiffusion problems with a weak initial singularity. BIT Numerical Mathematics, 61(3):1023–1059, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-020-00841-0.

Zhao:2021:ESO

[3524] Yong-Liang Zhao, Meng Li, and Xian-Ming Gu. An efficient secondorder energy stable BDF scheme for the space fractional Cahn–Hilliard equation. BIT Numerical Mathematics, 61(3):1061–1092, September 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00843-6.

Bai:2021:GKV

[3525] Xianglan Bai, Alessandro Buccini, and Lothar Reichel. Golub–Kahan vs. Monte Carlo: a comparison of bidiagonalization and a randomized SVD method for the solution of linear discrete ill-posed problems. BIT Numerical Mathematics, 61(4):1093–1114, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00857-0.

Cai:2021:UAN

[3526] Yongyong Cai and Yichen Guo. Uniformly accurate nested Picard integrators for a system of oscillatory ordinary differential equations. BIT Numerical Mathematics, 61(4):1115–1152, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00862-3.

Chaudhry:2021:PEA

[3527] Jehanzeb H. Chaudhry, Donald Estep, and Simon J. Tavener. A posteriori error analysis for Schwarz overlapping domain decomposition methods. BIT Numerical Mathematics, 61(4):1153–1191, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00864-1.

Claeys:2021:QSB

[3528] Xavier Claeys, Lorenzo Giacomel, and Carolina Urzúa-Torres. Quotient-space boundary element methods for scattering at complex screens. BIT Numerical Mathematics, 61(4):1193–1221, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00859-y.

Deligiannidis:2021:RWA

[3529] G. Deligiannidis, S. Maurer, and M. V.

Tretyakov. Random walk algorithm for the Dirichlet problem for parabolic integro-differential equation. BIT Numerical Mathematics, 61(4):1223–1269, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00863-2.

Galtung:2021:NSV

[3530] Sondre Tesdal Galtung and Katrin Grunert. A numerical study of variational discretizations of the Camassa– Holm equation. BIT Numerical Mathematics, 61(4):1271–1309, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00856-1.

Lin:2021:DPO

[3531] Fu-Rong Lin, Hai-Dong Qu, and Zi-Hang She. DNT preconditioner for one-sided space fractional diffusion equations. BIT Numerical Mathematics, 61(4):1311–1335, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00858-z.

Majidian:2021:ECO

[3532] Hassan Majidian. Efficient computation of oscillatory integrals by exponential transformations. BIT Numerical Mathematics, 61(4):1337–1365, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00855-2.

Soszynska:2021:ATS

[3533] Martyna Soszyńska and Thomas Richter. Adaptive time-step control for a monolithic multirate scheme coupling the heat and wave equation. BIT Numerical Mathematics, 61(4):1367–1396, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00854-3.

Su:2021:EIS

[3534] Chunmei Su and Gulcin M. Muslu. An exponential integrator sine pseudospectral method for the generalized improved Boussinesq equation. BIT Numerical Mathematics, 61(4):1397–1419, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00865-0.

Sun:2021:AIN

[3535] Jing Sun, Daxin Nie, and Weihua Deng. Algorithm implementation and numerical analysis for the two-dimensional tempered fractional Laplacian. BIT Numerical Mathematics, 61(4):1421–1452, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00860-5.

Zheng:2021:TFD

[3536] Xiangcheng Zheng and Hong Wang.

A time-fractional diffusion equation with space-time dependent hiddenmemory variable order: analysis and approximation. BIT Numerical Mathematics, 61(4):1453–1481, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00861-4.

Hashimoto:2021:CIR

[3537] Yuka Hashimoto and Takashi Nodera.

Correction to: Inexact rational Krylov method for evolution equations. BIT Numerical Mathematics, 61(4):1483–1487, December 2021. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00894-9.

Arnold:2022:WMS

[3538] Anton Arnold, Christian Klein, and Bernhard Ujvari. WKB-method for the 1D Schrödinger equation in the semi-classical limit: enhanced phase treatment. BIT Numerical Mathematics, 62(1):1–22, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00868-x.

Ceruti:2022:URI

[3539] Gianluca Ceruti and Christian Lubich. An unconventional robust integrator for dynamical low-rank approximation. BIT Numerical Mathematics, 62(1):23– 44, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00873-0.

deAlmagro:2022:CLT

[3540] Rodrigo T. Sato Martín de Almagro. Convergence of Lobatto-type Runge–Kutta methods for partitioned differential-algebraic systems of index 2. BIT Numerical Mathematics, 62(1): 45–67, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00871-2.

Dmytryshyn:2022:RPM

[3541] Andrii Dmytryshyn. Recovering a perturbation of a matrix polynomial from a perturbation of its first companion linearization. BIT Numerical Mathematics, 62(1):69–88, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00878-9.

Gaddam:2022:TNA

[3542] Sharat Gaddam, Thirupathi Gudi, and Kamana Porwal. Two new approaches for solving elliptic obstacle problems using discontinuous Galerkin methods. BIT Numerical Mathematics, 62(1):89– 124, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00869-w.

Kovac:2022:HCT

[3543] Erna Begović Kovac. Hybrid CURtype decomposition of tensors in the Tucker format. BIT Numerical Mathematics, 62(1):125–138, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00876-x.

Koyaguerebo-Ime:2022:AHO

[3544] Saint-Cyr E. R. Koyaguerebo-Imé and Yves Bourgault. Arbitrary high order A-stable and B-convergent numerical methods for ODEs via deferred correction. BIT Numerical Mathematics, 62(1):139–170, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00875-y.

Ma:2022:RKT

[3545] Anna Ma and Denali Molitor. Randomized Kaczmarz for tensor linear systems. BIT Numerical Mathematics, 62(1):171–194, March 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-021-00877-
                                        w.

Massei:2022:SAM

[3546] Stefano Massei. Some algorithms for maximum volume and cross approximation of symmetric semidefinite matrices. BIT Numerical Mathematics, 62(1):195–220, March 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-021-00872-
                                        1.

Musolas:2022:LRM

[3547] Antoni Musolas, Estelle Massart, and Youssef Marzouk. Low-rank multi-parametric covariance identification. BIT Numerical Mathematics, 62(1):221–249, March 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-021-00867-
                                        y.

Reddy:2022:FDP

[3548] G. Murali Mohan Reddy. Fully discrete a posteriori error estimates for parabolic integro-differential equations using the two-step backward differentiation formula. BIT Numerical Mathematics, 62(1):251–277, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00866-z.

Santin:2022:SBA

[3549] Gabriele Santin, Toni Karvonen, and Bernard Haasdonk. Sampling based approximation of linear functionals in reproducing kernel Hilbert spaces. BIT Numerical Mathematics, 62(1):279– 310, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00870-3.

Zhao:2022:RUD

[3550] Zhi Zhao, Xiao-Qing Jin, and Teng-Teng Yao. A Riemannian underdetermined BFGS method for least squares inverse eigenvalue problems. BIT Numerical Mathematics, 62(1): 311–337, March 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-021-00874-
z.

Annaby:2022:PUE

[3551] M. H. Annaby and S. R. Elsayed-Abdullah. Pointwise and uniform error estimates associated with Abel–Whittaker interpolation series and its dual. BIT Numerical Mathematics, 62(2):339–356, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00888-7.

Benzi:2022:RKM

[3552] Michele Benzi and Igor Simunec. Rational Krylov methods for fractional diffusion problems on graphs. BIT Numerical Mathematics, 62(2):357– 385, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00881-0.

Billaud-Friess:2022:NSA

[3553] Marie Billaud-Friess, Antonio Falcó, and Anthony Nouy. A new splitting algorithm for dynamical low-rank approximation motivated by the fibre bundle structure of matrix manifolds. BIT Numerical Mathematics, 62(2): 387–408, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00884-x.

Branden:2022:GIC

[3554] Henrik Brandén. Grid independent convergence using multilevel circulant preconditioning: Poisson’s equation. BIT Numerical Mathematics, 62(2): 409–429, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00885-w.

Cano:2022:HAO

[3555] B. Cano and N. Reguera. How to avoid order reduction when Lawson methods integrate nonlinear initial boundary value problems. BIT Numerical Mathematics, 62(2):431– 463, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00879-8.

Hon:2022:BTP

[3556] Sean Hon, Stefano Serra-Capizzano, and Andy Wathen. Band-Toeplitz preconditioners for ill-conditioned Toeplitz systems. BIT Numerical Mathematics, 62(2):465–491, June 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-021-00889-
6.

Hong:2022:PPS

[3557] Jialin Hong, Lihai Ji, and Jingjing Zhang. Positivity-preserving symplectic methods for the stochastic Lotka– Volterra predator-prey model. BIT Numerical Mathematics, 62(2):493– 520, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00891-y.

Naito:2022:HOW

[3558] Riu Naito and Toshihiro Yamada. A higher order weak approximation of McKean–Vlasov type SDEs. BIT Numerical Mathematics, 62(2):521– 559, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00880-1.

Noutsos:2022:BTC

[3559] Dimitrios Noutsos and Grigorios Tachyridis. Band-times-circulant preconditioners for non-symmetric Toeplitz systems. BIT Numerical Mathematics, 62(2):561–590, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https://link.

                                        springer.com/article/10.1007/s10543-
                                        021-00883-y.

Oviedo:2022:AMG

[3560] Harry Oviedo, Oscar Dalmau, and Rafael Herrera. An accelerated minimal gradient method with momentum for strictly convex quadratic optimization. BIT Numerical Mathematics, 62(2):591–606, June 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-021-00886-
                                        9.

Wahlsten:2022:EHM

[3561] Markus Wahlsten, Oskar Ålund, and Jan Nordström. An efficient hybrid method for uncertainty quantification. BIT Numerical Mathematics, 62(2): 607–629, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00882-z.

Yin:2022:ESF

[3562] Baoli Yin, Yang Liu, and Zhimin Zhang. Efficient shifted fractional trapezoidal rule for subdiffusion problems with nonsmooth solutions on uniform meshes. BIT Numerical Mathematics, 62(2):631– 666, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00890-z.

Zheng:2022:FSP

[3563] Zhong Zheng, Jing Chen, and Yue-Fen Chen. A fully structured preconditioner for a class of complex symmetric indefinite linear systems. BIT Numerical Mathematics, 62(2):667– 680, June 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00887-8.

Barbarino:2022:SAR

[3564] Giovanni Barbarino. A systematic approach to reduced GLT. BIT Numerical Mathematics, 62(3):681–743, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00896-7.

Bazavov:2022:CFL

[3565] Alexei Bazavov. Commutatorfree Lie group methods with minimum storage requirements and reuse of exponentials. BIT Numerical Mathematics, 62(3):745–771, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00892-x.

DellAccio:2022:NDS

[3566] Francesco Dell’Accio, Filomena Di Tommaso, Najoua Siar, and Marco Vianello. Numerical differentiation on scattered data through multivariate polynomial interpolation. BIT Numerical Mathematics, 62(3):773–801, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00897-6.

Eisenmann:2022:EEB

[3567] Monika Eisenmann, Mihály Kovács, Raphael Kruse, and Stig Larsson. Error estimates of the backward Euler–Maruyama method for multivalued stochastic differential equations. BIT Numerical Mathematics, 62(3):803–848, September 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-021-00893-
w.

Gemignani:2022:CTS

[3568] Luca Gemignani and Federico Poloni. Comparison theorems for splittings of m-matrices in (block) Hessenberg form. BIT Numerical Mathematics, 62(3):849–867, September 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-021-00899-
4.

Gunther:2022:MLI

[3569] Michael Günther and Adrian Sandu. Multirate linearly-implicit GARK schemes. BIT Numerical Mathematics, 62(3):869–901, September 2022. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-021-00898-
5.

Kemmochi:2022:SAV

[3570] Tomoya Kemmochi and Shun Sato.

Scalar auxiliary variable approach for conservative/dissipative partial differential equations with unbounded energy functionals. BIT Numerical Mathematics, 62(3):903–930, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00904-w.

Lanza:2022:AFR

[3571] A. Lanza, M. Pragliola, and F. Sgallari. Automatic fidelity and regularization terms selection in variational image restoration. BIT Numerical Mathematics, 62(3):931–964, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00901-z.

Lee:2022:EAE

[3572] Kookjin Lee, Howard C. Elman, Catherine E. Powell, and Dongeun Lee. Enhanced alternating energy minimization methods for stochastic Galerkin matrix equations. BIT Numerical Mathematics, 62(3):965–994, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00903-x.

Li:2022:NSG

[3573] Lei Li and Dongling Wang. Numerical stability of Grünwald–Letnikov method for time fractional delay differential equations. BIT Numerical Mathematics, 62(3):995–1027, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00900-0.

Lu:2022:HME

[3574] Peipei Lu, Andreas Rupp, and Guido Kanschat. Homogeneous multigrid for embedded discontinuous Galerkin methods. BIT Numerical Mathematics, 62(3):1029–1048, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00902-y.

Wu:2022:EET

[3575] Yifei Wu and Xiaofei Zhao. Embedded exponential-type low-regularity integrators for KdV equation under rough data. BIT Numerical Mathematics, 62(3):1049–1090, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00895-8.

Buggenhout:2022:CHH

[3576] Niel Van Buggenhout and Marc Van Barel. Correction to: “How and how not to check Gaussian quadrature formulae”. BIT Numerical Mathematics, 62(3):1091–1092, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00921-3.

Muniz:2022:CHS

[3577] Michelle Muniz, Matthias Ehrhardt, Michael Günther, and Renate Winkler. Correction to: Higher strong order methods for linear Itô SDEs on matrix Lie groups. BIT Numerical Mathematics, 62(3):1093, September 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00911-5.  See

[3578].

Muniz:2022:HSO

[3578] Michelle Muniz, Matthias Ehrhardt, Michael Günther, and Renate Winkler. Higher strong order methods for linear Itô SDEs on matrix Lie groups. BIT Numerical Mathematics, 62(4):1095–1119, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-021-00905-9. See cor-

rection [3577].

Debrabant:2022:LSH

[3579] Kristian Debrabant, Anne Kværnø, and Nicky Cordua Mattsson. Lawson schemes for highly oscillatory stochastic differential equations and conservation of invariants. BIT Numerical Mathematics, 62(4):1121–1147, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00906-8.

Ceruti:2022:RAR

[3580] Gianluca Ceruti, Jonas Kusch, and Christian Lubich. A rank-adaptive robust integrator for dynamical lowrank approximation. BIT Numerical Mathematics, 62(4):1149–1174, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-021-00907-7.

Liang:2022:FRB

[3581] Zhao-Zheng Liang and Yan Dou. Fast rotated BSOR method for block two-by-two linear systems with application to PDE-constrained optimal control problems. BIT Numerical Mathematics, 62(4):1175–1206, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00908-0.

Eidnes:2022:OTD

[3582] Sølve Eidnes. Order theory for discrete gradient methods. BIT Numerical Mathematics, 62(4):1207–1255, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00909-z.

Bentbib:2022:CIV

[3583] A. H. Bentbib, A. Khouia, and H. Sadok. Color image and video restoration using tensor CP decomposition. BIT Numerical Mathematics, 62(4):1257–1278, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00910-6.

Merkle:2022:MMC

[3584] Robin Merkle and Andrea Barth. Multilevel Monte Carlo estimators for elliptic PDEs with Lévy-type diffusion coefficient. BIT Numerical Mathematics, 62(4):1279–1317, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00912-4.

Shan:2022:FOD

[3585] Xiujie Shan, Jiebao Sun, Zhichang Guo, Wenjuan Yao, and Zhenyu Zhou. Fractional-order diffusion model for multiplicative noise removal in texture-rich images and its fast explicit diffusion solving. BIT Numerical Mathematics, 62(4):1319–1354, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00913-3.

Boehm:2022:ENA

[3586] Udo Boehm, Sonja Cox, Gregor Gantner, and Rob Stevenson. Efficient numerical approximation of a non-regular Fokker–Planck equation associated with first-passage time distributions. BIT Numerical Mathematics, 62(4):1355–1382, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00914-2.

Asadzadeh:2022:MSE

[3587] M. Asadzadeh and B. N. Saray. On a multiwavelet spectral element method for integral equation of a generalized Cauchy problem. BIT Numerical Mathematics, 62(4):1383–1416, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00915-1.

Bogoya:2022:FSE

[3588] Manuel Bogoya, Sergei M. Grudsky, Stefano Serra–Capizzano, and Cristina Tablino–Possio. Fine spectral estimates with applications to the optimally fast solution of large FDE linear systems. BIT Numerical Mathematics, 62(4):1417–1431, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00916-0.

Grubisic:2022:HOA

[3589] Luka Grubisić and Harri Hakula. High order approximations of the operator Lyapunov equation have low rank. BIT Numerical Mathematics, 62(4):1433–1459, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00917-z.

Chkifa:2022:NIU

[3590] Moulay Abdellah Chkifa. Newton interpolation using ℜ-Leja sequences. BIT Numerical Mathematics, 62(4):1461–1485, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00918-y.

Zeifang:2022:SIM

[3591] Jonas Zeifang, Jochen Schütz, and David C. Seal. Stability of implicit multiderivative deferred correction methods. BIT Numerical Mathematics, 62(4):1487–1503, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00919-x.

Leobacher:2022:WPN

[3592] Gunther Leobacher, Christoph Reisinger, and Wolfgang Stockinger. Wellposedness and numerical schemes for one-dimensional McKean–Vlasov equations and interacting particle systems with discontinuous drift. BIT Numerical Mathematics, 62(4):1505–1549, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00920-4.

Druskin:2022:MOR

[3593] Vladimir Druskin, Stefan Güttel, and Leonid Knizhnerman. Model order reduction of layered waveguides via rational Krylov fitting. BIT Numerical Mathematics, 62(4):1551–1572, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00922-2.

Liu:2022:CRB

[3594] Zhihui Liu. L^{p}-convergence rate of backward Euler schemes for monotone SDEs. BIT Numerical Mathematics, 62(4):1573–1590, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00923-1.

Yang:2022:NCS

[3595] Guoguo Yang, Kevin Burrage, Yoshio Komori, and Xiaohua Ding. A new class of structure-preserving stochastic exponential Runge–Kutta integrators for stochastic differential equations. BIT Numerical Mathematics, 62(4):1591–1623, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00924-0.

Calvo:2022:UAS

[3596] Maria Cabrera Calvo and Katharina Schratz. Uniformly accurate splitting schemes for the Benjamin– Bona–Mahony equation with dispersive parameter. BIT Numerical Mathematics, 62(4):1625–1647, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00925-z.

Benner:2022:PLN

[3597] Peter Benner, Bruno Iannazzo, Beatrice Meini, and Davide Palitta. Palindromic linearization and numerical solution of nonsymmetric algebraic T - Riccati equations. BIT Numerical Mathematics, 62(4):1649–1672, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00926-y.

Klein:2022:UDE

[3598] Simon-Christian Klein. Using the Dafermos entropy rate criterion in numerical schemes. BIT Numerical Mathematics, 62(4):1673–1701, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00927-x.

Aduamoah:2022:PMI

[3599] Mildred Aduamoah, Benjamin D.

Goddard, John W. Pearson, and Jonna C. Roden. Pseudospectral methods and iterative solvers for optimization problems from multiscale particle dynamics. BIT Numerical Mathematics, 62(4):1703–1743, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00928-w.

Patrizi:2022:EGR

[3600] Francesco Patrizi. Effective grading refinement for locally linearly independent LR B-splines. BIT Numerical Mathematics, 62(4):1745–1764, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00929-9.

Scott:2022:SLL

[3601] Jennifer Scott and Miroslav Tuma. Solving large linear least squares problems with linear equality constraints. BIT Numerical Mathematics, 62(4):1765–1787, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00930-2.

Bourgault:2022:VSD

[3602] Yves Bourgault and André Garon. Variable-step deferred correction methods based on backward differentiation formulae for ordinary differential equations. BIT Numerical Mathematics, 62(4):1789–1822, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00931-1.

daSilva:2022:VSD

[3603] Clauson Carvalho da Silva and Christian Lessig. Variational symplectic diagonally implicit Runge–Kutta methods for isospectral systems. BIT Numerical Mathematics, 62(4):1823–1840, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00932-0.

AlZohbi:2022:CSE

[3604] Maryam Al Zohbi, Ahmad El Hajj, and Mustapha Jazar. Convergent semi-explicit scheme to a non-linear eikonal system. BIT Numerical Mathematics, 62(4):1841–1872, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00933-z.

Gholami:2022:ABP

[3605] Ali Gholami and Silvia Gazzola. Automatic balancing parameter selection for Tikhonov–TV regularization. BIT Numerical Mathematics, 62(4):1873–1898, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00934-y.

An:2022:QEH

[3606] Congpei An and Hao-Ning Wu. On the quadrature exactness in hyperinterpolation. BIT Numerical Mathematics, 62(4):1899–1919, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00935-x.

Penent:2022:NEO

[3607] Guillaume Penent and Nicolas Privault. Numerical evaluation of ODE solutions by Monte Carlo enumeration of Butcher series. BIT Numerical Mathematics, 62(4):1921–1944, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00936-w.

Motamed:2022:HLR

[3608] Mohammad Motamed. A hierarchically low-rank optimal transport dissimilarity measure for structured data. BIT Numerical Mathematics, 62(4):1945–1982, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00937-9.

Luo:2022:DEP

[3609] Wei-Hua Luo, Xian-Ming Gu, and Bruno Carpentieri. A dimension expanded preconditioning technique for saddle point problems. BIT Numerical Mathematics, 62(4):1983–2004, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00938-8.

Xu:2022:EAM

[3610] Fei Xu, Qiumei Huang, and Manting Xie. An efficient adaptive multigrid method for the elasticity eigenvalue problem. BIT Numerical Mathematics, 62(4):2005–2033, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-022-00939-7.

Luo:2022:CDE

[3611] Wei-Hua Luo, Xian-Ming Gu, and Bruno Carpentieri. Correction to: a dimension expanded preconditioning technique for saddle point problems. BIT Numerical Mathematics, 62(4):2035, December 2022. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-022-00940-0.

Ivicic:2023:PBS

[3612] Ivana Kuzmanović Ivicić and Suzana Miodragović. Perturbation bounds for stable gyroscopic systems. BIT Numerical Mathematics, 63(1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00943-5.

Doan:2023:LRI

[3613] Cao-Kha Doan, Thi-Thao-Phuong Hoang, Lili Ju, and Katharina Schratz. Low regularity integrators for semilinear parabolic equations with maximum bound principles. BIT Numerical Mathematics, 63(1): ??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00946-2.

Almuslimani:2023:FAE

[3614] Ibrahim Almuslimani. A fully adaptive explicit stabilized integrator for advection-diffusion-reaction problems. BIT Numerical Mathematics, 63(1): ??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00945-3.

Hochbruck:2023:DLR

[3615] Marlis Hochbruck, Markus Neher, and Stefan Schrammer. Dynamical low-rank integrators for secondorder matrix differential equations. BIT Numerical Mathematics, 63(1): ??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00941-7.

Appelo:2023:EBD

[3616] Daniel Appelö, Lu Zhang, Thomas Hagstrom, and Fengyan Li. An energy-based discontinuous Galerkin method with tame CFL numbers for the wave equation. BIT Numerical Mathematics, 63(1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00954-2.

Glaubitz:2023:TSR

[3617] Jan Glaubitz and Jonah A. Reeger. Towards stability results for global radial basis function based quadrature formulas. BIT Numerical Mathematics, 63 (1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00956-0.

Wang:2023:FAS

[3618] Yuan-Ming Wang and Bo Xie. A fractional Adams–Simpson-type method for nonlinear fractional ordinary differential equations with non-smooth data. BIT Numerical Mathematics, 63 (1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00952-4.

Hajduk:2023:AAF

[3619] Hennes Hajduk and Andreas Rupp.

Analysis of algebraic flux correction schemes for semi-discrete advection problems. BIT Numerical Mathematics, 63(1):??, March 2023. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-023-00957-
z.

Hochbruck:2023:RAD

[3620] Marlis Hochbruck, Markus Neher, and Stefan Schrammer. Rankadaptive dynamical low-rank integrators for first-order and secondorder matrix differential equations. BIT Numerical Mathematics, 63(1): ??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00942-6.

Golubovic:2023:QPE

[3621] Zora Lj. Golubović and Gradimir V.

Milovanović. Quadrature processes for efficient calculation of the Clausen functions. BIT Numerical Mathematics, 63(1):??, March 2023. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-023-00944-

Ye:2023:CTF

[3622] Changqing Ye and Eric T. Chung. Convergence of trigonometric and finitedifference discretization schemes for FFT-based computational micromechanics. BIT Numerical Mathematics, 63(1):??, March 2023. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-023-00950-
                                        6.

Chaudhry:2023:EET

[3623] Jehanzeb H. Chaudhry, Donald Estep, Trevor Giannini, Zachary Stevens, and Simon J. Tavener. Error estimation for the time to a threshold value in evolutionary partial differential equations. BIT Numerical Mathematics, 63 (1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00947-1.

Carrel:2023:LRP

[3624] Benjamin Carrel, Martin J. Gander, and Bart Vandereycken. Low-rank Parareal: a low-rank parallel-in-time integrator. BIT Numerical Mathematics, 63(1):??, March 2023. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-023-00953-
                                        3.

Lubich:2023:LSI

[3625] Christian Lubich and Yanyan Shi. On a large-stepsize integrator for charged-particle dynamics. BIT Numerical Mathematics, 63(1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00951-5.

Cappellazzo:2023:KTE

[3626] G. Cappellazzo, W. Erb, F. Marchetti, and D. Poggiali. On Kosloff Tal– Ezer least-squares quadrature formulas. BIT Numerical Mathematics, 63 (1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00948-0.

Ivanenko:2023:STC

[3627] D. Ivanenko, A. Kohatsu-Higa, and A. Kulik. Small time chaos approximations for heat kernels of multidimensional diffusions. BIT Numerical Mathematics, 63(1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00949-z.

Jacobs:2023:ORA

[3628] Byron A. Jacobs, Fredrik Laurén, and Jan Nordström. On the order reduction of approximations of fractional derivatives: an explanation and a cure. BIT Numerical Mathematics, 63(1): ??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00961-3.

Kressner:2023:SQE

[3629] Daniel Kressner and Ivana Sain Glibić. Singular quadratic eigenvalue problems: linearization and weak condition numbers. BIT Numerical Mathematics, 63(1):??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00960-4.

Zolfaghari:2023:HOM

[3630] Reza Zolfaghari and Nedialko S. Nedialkov. An Hermite–Obreschkoff method for stiff high-index DAE. BIT Numerical Mathematics, 63(1): ??, March 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00955-1.

Bini:2023:GMQ

[3631] Dario A. Bini, Bruno Iannazzo, and Jie Meng. Geometric means of quasi-Toeplitz matrices. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00962-2.

Huybrechs:2023:AIE

[3632] Daan Huybrechs and Lloyd N. Trefethen. AAA interpolation of equispaced data. BIT Numerical Mathematics, 63(2):??, June 2023. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-023-00959-
                                        x.

Sommariva:2023:LCP

[3633] Alvise Sommariva and Marco Vianello.

Low cardinality positive interior cubature on NURBS-shaped domains. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00958-y.

Yu:2023:SUS

[3634] Annan Yu and Alex Townsend. On the stability of unevenly spaced samples for interpolation and quadrature. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00965-z.

DellAccio:2023:PIR

[3635] Francesco Dell’Accio, Domenico Mezzanotte, Federico Nudo, and Donatella Occorsio. Product integration rules by the constrained mock-Chebyshev least squares operator. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00968-w.

Liu:2023:PGA

[3636] Yanhui Liu, Xueying Zeng, and Weiguo Wang. Proximal gradient algorithm for nonconvex low tubal rank tensor recovery. BIT Numerical Mathematics, 63 (2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00964-0.

Jahnke:2023:NMS

[3637] Tobias Jahnke and Michael Kirn. On numerical methods for the seminonrelativistic limit system of the nonlinear Dirac equation. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00971-1.

Steinebach:2023:CRW

[3638] Gerd Steinebach. Construction of Rosenbrock–Wanner method Rodas5P and numerical benchmarks within the Julia Differential Equations package. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00967-x.

Rump:2023:CUF

[3639] Siegfried M. Rump. Computation of the unit in the first place (ufp) and the unit in the last place (ulp) in precisionp base β. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00970-2.

Kunkel:2023:DIO

[3640] Peter Kunkel and Volker Mehrmann. Discretization of inherent ODEs and the geometric integration of DAEs with symmetries. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00966-y.

Casulli:2023:LRT

[3641] Angelo Casulli and Leonardo Robol.

Low-rank tensor structure preservation in fractional operators by means of exponential sums. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00974-y.

Bartels:2023:FFP

[3642] Tinko Bartels, Vissarion Fisikopoulos, and Martin Weiser. Fast floatingpoint filters for robust predicates. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00975-x.

Plaskota:2023:MCI

[3643] Leszek Plaskota, Paweł Przybyłowicz, and Lukasz Stepień. Monte Carlo integration of C^{r} functions with adaptive variance reduction: an asymptotic analysis. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00972-0.

Kelly:2023:SCA

[3644] Cónall Kelly, Gabriel J. Lord, and Fandi Sun. Strong convergence of an adaptive time-stepping Milstein method for SDEs with monotone coefficients. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00969-9.

Lederer:2023:AEP

[3645] Philip L. Lederer. Asymptotically exact a posteriori error estimates for the BDM finite element approximation of mixed Laplace eigenvalue problems. BIT Numerical Mathematics, 63 (2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00976-w.

Eckhardt:2023:FDH

[3646] Daniel Eckhardt and Barbara Verfürth. Fully discrete heterogeneous multiscale method for parabolic problems with multiple spatial and temporal scales. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00973-z.

Park:2023:FRA

[3647] Taejun Park and Yuji Nakatsukasa. A fast randomized algorithm for computing an approximate null space. BIT Numerical Mathematics, 63(2): ??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00979-7.

Khoa:2023:ASV

[3648] Vo Anh Khoa, Nguyen Dat Thuc, and Ajith Gunaratne. Analysis and simulation of a variational stabilization for the Helmholtz equation with noisy Cauchy data. BIT Numerical Mathematics, 63 (2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00978-8.

Zheng:2023:MMC

[3649] Chao Zheng. Multilevel Monte Carlo using approximate distributions of the CIR process. BIT Numerical Mathematics, 63(2):??, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00980-0.

Liao:2023:SCV

[3650] Hong lin Liao, Tao Tang, and Tao Zhou. Stability and convergence of the variable-step time filtered backward Euler scheme for parabolic equations. BIT Numerical Mathematics, 63(3):??, September 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00982-y.

Zhou:2023:BEM

[3651] Hao Zhou, Yaozhong Hu, and Yanghui Liu. Backward Euler method for stochastic differential equations with non-Lipschitz coefficients driven by fractional Brownian motion. BIT Numerical Mathematics, 63(3):??, September 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00981-z.

Mitchell:2023:DCM

[3652] William Mitchell, Abbie Natkin, Paige Robertson, Marika Sullivan, Xuechen Yu, and Chenxin Zhu. Decomposition and conformal mapping techniques for the quadrature of nearly singular integrals. BIT Numerical Mathematics, 63(3):??, September 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00984-w.

Lee:2023:APP

[3653] Jeonghun J. Lee. Analysis and preconditioning of parameter-robust finite element methods for Biot’s consolidation model. BIT Numerical Mathematics, 63(3):??, September 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00983-x.

Cai:2023:SCR

[3654] Meng Cai, Ruisheng Qi, and Xiaojie Wang. Strong convergence rates of an explicit scheme for stochastic Cahn– Hilliard equation with additive noise. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00987-7.

Rempke:2023:PLI

[3655] Arne Rempke. Parallel line identification for line-implicit-solvers. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00977-9.

Linders:2023:REG

[3656] Viktor Linders, Hendrik Ranocha, and Philipp Birken. Resolving entropy growth from iterative methods. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00992-w.

Han:2023:SLO

[3657] Weimin Han, Feifei Jing, and Yuan Yao. Stabilized low-order mixed finite element methods for a Navier– Stokes hemivariational inequality. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00985-9.

Du:2023:SFE

[3658] Zhijie Du and Huoyuan Duan. A stabilized finite element method on nonaffine grids for time-harmonic Maxwell’s equations. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00988-6.

Duan:2023:PAS

[3659] Xue-Feng Duan, Yong-Shen Zhang, Qing-Wen Wang, and Chun-Mei Li. Paige’s Algorithm for solving a class of tensor least squares problem. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00990-y.

Faber:2023:FC

[3660] Vance Faber, Jörg Liesen, and Petr Tichý. On the Forsythe conjecture. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00991-x.

Celledoni:2023:DNN

[3661] Elena Celledoni, Helge Glöckner, Jørgen N. Riseth, and Alexander Schmeding. Deep neural networks on diffeomorphism groups for optimal shape reparametrization. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00989-5.

Lang:2023:GCA

[3662] Annika Lang and Mike Pereira. Galerkin–Chebyshev approximation of Gaussian random fields on compact Riemannian manifolds. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00986-8.

Galindo-Olarte:2023:SAE

[3663] Andrés Galindo-Olarte, Juntao Huang, Jennifer Ryan, and Yingda Cheng. Superconvergence and accuracy enhancement of discontinuous Galerkin solutions for Vlasov–Maxwell equations. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00993-9.

Denich:2023:GRI

[3664] Eleonora Denich and Paolo Novati.

Gaussian rule for integrals involving Bessel functions. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00997-5.

Jay:2023:NAJ

[3665] Laurent O. Jay and Olga Sokratova. A note on approximate Jacobians of implicit Runge–Kutta methods and convergence of modified Newton iterations. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00994-8.

Zhang:2023:NSP

[3666] Xiaodi Zhang and Shitian Dong. New structure-preserving mixed finite element method for the stationary MHD equations with magneticcurrent formulation. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00995-7.

Han:2023:EHV

[3667] Jiayu Han and Zhimin Zhang. An hp-version interior penalty discontinuous Galerkin method for the quadcurl eigenvalue problem. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00996-6.

Diab:2023:FSR

[3668] Malak Diab, Andreas Frommer, and Karsten Kahl. A flexible short recurrence Krylov subspace method for matrices arising in the time integration of port-Hamiltonian systems and ODEs/ DAEs with a dissipative Hamiltonian. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-00999-3.

Bernier:2023:SCS

[3669] J. Bernier, S. Blanes, F. Casas, and A. Escorihuela-Tomàs. Symmetricconjugate splitting methods for linear unitary problems. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-00998-4.

Deng:2023:PPT

[3670] Shounian Deng, Chen Fei, Weiyin Fei, and Xuerong Mao. Positivitypreserving truncated Euler–Maruyama method for generalised Ait–Sahaliatype interest model. BIT Numerical Mathematics, 63(4):??, December 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-01000-x.

Zhang:2024:RKA

[3671] Zeyi Zhang and Dong Shen. Randomized Kaczmarz algorithm with averaging and block projection. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-01002-9.

Hellman:2024:WPF

[3672] Fredrik Hellman, Axel Målqvist, and Malin Mosquera. Well-posedness and finite element approximation of mixed dimensional partial differential equations. BIT Numerical Mathematics, 64 (1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-01001-w.

Amat:2024:RCA

[3673] Sergio Amat, David Levin, Juan Ruiz-Álvarez, and Dionisio F. Yáñez. A regularization-correction approach for adapting subdivision schemes to the presence of discontinuities. BIT Numerical Mathematics, 64(1): ??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-01003-8.

Zeng:2024:IAT

[3674] Chao Zeng, Michael K. Ng, and Tai-Xiang Jiang. Incremental algorithms for truncated higher-order singular value decompositions. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-01004-7.

Malyshev:2024:SRL

[3675] Alexander Malyshev and Miloud Sadkane. On the stability radius for linear time-delay systems. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-023-01006-5.

Wang:2024:IUE

[3676] Bin Wang and Yaolin Jiang. Improved uniform error bounds on parareal exponential algorithm for highly oscillatory systems. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-023-01005-6.

Ku:2024:PEE

[3677] JaEun Ku and Martin Stynes. A posteriori error estimates for a dual finite element method for singularly perturbed reaction–diffusion problems. BIT Numerical Mathematics, 64(1): ??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01008-x.

Ren:2024:CQU

[3678] Hao Ren, Junjie Ma, and Huilan Liu. A convolution quadrature using derivatives and its application. BIT Numerical Mathematics, 64(1): ??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01009-w.

Ma:2024:CSF

[3679] Zheng Ma, Martin Stynes, and Chengming Huang. Convergence and superconvergence of a fractional collocation method for weakly singular Volterra integro-differential equations. BIT Numerical Mathematics, 64(1): ??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01011-2.

Prajapati:2024:SEB

[3680] Anshul Prajapati and Punit Sharma.

Structured eigenvalue backward errors for rational matrix functions with symmetry structures. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01010-3.

afKlinteberg:2024:SSQ

[3681] Ludvig af Klinteberg. Singularity swap quadrature for nearly singular line integrals on closed curves in two dimensions. BIT Numerical Mathematics, 64 (1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01013-0.

Massey:2024:ASS

[3682] Pedro Massey. Admissible subspaces and the subspace iteration method. BIT Numerical Mathematics, 64(1): ??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01012-1.

Adjerid:2024:UIF

[3683] Slimane Adjerid, Tao Lin, and Haroun Meghaichi. A unified immersed finite element error analysis for onedimensional interface problems. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01014-z.

Akrivis:2024:VTS

[3684] Georgios Akrivis, Minghua Chen, Jianxing Han, Fan Yu, and Zhimin Zhang. The variable two-step BDF method for parabolic equations. BIT Numerical Mathematics, 64(1):??, March 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01007-y.

Abbasi:2024:BPR

[3685] Mokhtar Abbasi and Touraj Nikazad.

Bounded perturbations resilient iterative methods for linear systems and least squares problems: operator-based approaches, analysis, and performance evaluation. BIT Numerical Mathematics, 64(2): ??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01015-y.

Erb:2024:HLR

[3686] W. Erb, T. Hangelbroek, F. J. Narcowich, C. Rieger, and J. D. Ward. Highly localized RBF Lagrange functions for finite difference methods on spheres. BIT Numerical Mathematics, 64(2):??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01016-x.

Cameron:2024:AHM

[3687] Thomas R. Cameron and Stef Graillat. Accurate Horner methods in real and complex floating-point arithmetic. BIT Numerical Mathematics, 64(2):

??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01017-w.

Li:2024:LFD

[3688] Qing Li and Steinar Evje. Learning the flux and diffusion function for degenerate convection-diffusion equations using different types of observations. BIT Numerical Mathematics, 64 (2):??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01018-9.

Uschmajew:2024:DLR

[3689] André Uschmajew and Andreas Zeiser. Dynamical low-rank approximation of the Vlasov–Poisson equation with piecewise linear spatial boundary. BIT Numerical Mathematics, 64(2): ??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01019-8.

Altmann:2024:SEI

[3690] R. Altmann, R. Maier, and B. Unger. Semi-explicit integration of second order for weakly coupled poroelasticity. BIT Numerical Mathematics, 64 (2):??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01021-0.

Xu:2024:MLD

[3691] Fei Xu, Bingyi Wang, and Manting Xie. Multilevel local defectcorrection method for the nonselfadjoint Steklov eigenvalue problems. BIT Numerical Mathematics, 64 (2):??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01022-z.

Rump:2024:QMI

[3692] Siegfried M. Rump and Takeshi Ogita.

On a quality measure for interval inclusions. BIT Numerical Mathematics, 64 (2):??, June 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01020-1.

Gustafsson:2024:SFE

[3693] Tom Gustafsson and Juha Videman. Stabilised finite element method for Stokes problem with nonlinear slip condition. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01025-w.

Feng:2024:CMP

[3694] Yuanyuan Feng and Lei Li. A class of monotonicity-preserving variablestep discretizations for Volterra integral equations. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01027-8.

Seguin:2024:LRR

[3695] Axel Séguin, Gianluca Ceruti, and Daniel Kressner. From low-rank retractions to dynamical low-rank approximation and back. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01028-7.

Jarman:2024:SIM

[3696] Benjamin Jarman, Lara Kassab, Deanna Needell, and Alexander Sietsema. Stochastic iterative methods for online rank aggregation from pairwise comparisons. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01024-x.

Chen:2024:EHV

[3697] Yanping Chen, Zhenrong Chen, and Yunqing Huang. An hp-version of the discontinuous Galerkin method for fractional integro-differential equations with weakly singular kernels. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01026-9.

Pradhan:2024:OCA

[3698] Gouranga Pradhan and Bhupen Deka. Optimal convergence analysis of the virtual element methods for viscoelastic wave equations with variable coefficients on polygonal meshes. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01030-z.

Erdogan:2024:WCT

[3699] Utku Erdogan and Gabriel J. Lord.

Weak convergence of tamed exponential integrators for stochastic differential equations. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01029-6.

Ceruti:2024:RSO

[3700] Gianluca Ceruti, Lukas Einkemmer, Jonas Kusch, and Christian Lubich. A robust second-order low-rank BUG integrator based on the midpoint rule. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01032-x.

Delville-Atchekzai:2024:SHX

[3701] Roxane Delville-Atchekzai, Xavier Claeys, and Matthieu Lecouvez. Substructuring the Hiptmair–Xu preconditioner for positive definite H(curl, Ω) problems. BIT Numerical Mathematics, 64(3):??, September 2024. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-024-01031-
y.

Kressner:2024:AEC

[3702] Daniel Kressner and Bor Plestenjak. Analysis of eigenvalue condition numbers for a class of randomized numerical methods for singular matrix pencils. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01033-w.

Bonizzoni:2024:SLO

[3703] Francesca Bonizzoni, Philip Freese, and Daniel Peterseim. Super-localized orthogonal decomposition for convectiondominated diffusion problems. BIT Numerical Mathematics, 64(3):??, September 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01035-8.

Cools:2024:BDC

[3704] Kristof Cools and Carolina Urzúa-Torres. Block diagonal Calderón preconditioning for scattering at multiscreens. BIT Numerical Mathematics, 64(4):??, December 2024. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-024-01034-
                                        9.

Przybylowicz:2024:LEB

[3705] Paweł Przybyłowicz, Verena Schwarz, and Michaela Szölgyenyi. Lower error bounds and optimality of approximation for jump-diffusion SDEs with discontinuous drift. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01036-7.

Demir:2024:PRA

[3706] Medine Demir and Volker John.

Pressure-robust approximation of the incompressible Navier–Stokes equations in a rotating frame of reference. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01037-6.

Engstrom:2024:LCN

[3707] Emil Engström and Eskil Hansen. Linearly convergent nonoverlapping domain decomposition methods for quasilinear parabolic equations. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01038-5.

Guglielmi:2024:SML

[3708] Nicola Guglielmi and Stefano Sicilia. Stabilization of a matrix via a low-rank-adaptive ODE. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01039-4.

Frachon:2024:DFC

[3709] Thomas Frachon, Erik Nilsson, and Sara Zahedi. Divergence-free cut finite element methods for Stokes flow. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01040-x.

Zou:2024:SGU

[3710] Guang an Zou, Yuanhong Wei, and Xiaofeng Yang. A stabilized Gauge–Uzawa discontinuous Galerkin method for the magnetohydrodynamic equations. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01044-7.

Wang:2024:ADP

[3711] Lu Wang and Hui Liang. Analysis of direct piecewise polynomial collocation methods for the Bagley–Torvik equation. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01043-8.

Seguin:2024:HIR

[3712] Axel Séguin and Daniel Kressner. Hermite interpolation with retractions on manifolds. BIT Numerical Mathematics, 64(4):??, December 2024. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01023-y.

Kang:2025:ENM

[3713] Hongchao Kang, Ao Liu, and Wentao Cai. Efficient numerical methods of integrals with products of two Bessel functions and their error analysis. BIT Numerical Mathematics, 65 (1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01045-6. See cor-

rection [3725].

Barbarino:2025:MLS

[3714] Giovanni Barbarino, Melker Claesson, Sven-Erik Ekström, Carlo Garoni, David Meadon, and Hendrik Speleers. Matrix-less spectral approximation for large structured matrices. BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01041-w.

Bruno:2025:FPS

[3715] Ludovico Bruni Bruno and Wolfgang Erb. The Fekete problem in segmental polynomial interpolation. BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01047-4.

Xu:2025:CVC

[3716] Xuefeng Xu. Convergence of VWcycle and WV-cycle multigrid methods. BIT Numerical Mathematics, 65 (1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01042-9.

Albrecht:2025:CGK

[3717] Kristof Albrecht and Armin Iske. On the convergence of generalized kernel-based interpolation by greedy data selection algorithms. BIT Numerical Mathematics, 65(1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01048-3.

Redmann:2025:EDR

[3718] Martin Redmann and Sebastian Riedel. Exact dimension reduction for rough differential equations. BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01046-5.

Bochacik:2025:CSR

[3719] Tomasz Bochacik, Paweł Przybyłowicz, and Łukasz Stepień. Convergence and stability of randomized implicit two-stage Runge–Kutta schemes. BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-024-01050-9.

Zhong:2025:EAI

[3720] Yimin Zhong, Kui Ren, Olof Runborg, and Richard Tsai. Error analysis for the implicit boundary integral method.

BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01051-8.

Huang:2025:CRR

[3721] Xin Huang, Petr Plecháč, Mattias Sandberg, and Anders Szepessy. Convergence rates for random feature neural network approximation in molecular dynamics. BIT Numerical Mathematics, 65(1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01052-1.

Sloan:2025:DRI

[3722] Ian H. Sloan and Vesa Kaarnioja.

Doubling the rate: improved error bounds for orthogonal projection with application to interpolation. BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-024-01049-2.

Wenzel:2025:AMA

[3723] Tizian Wenzel, Daniel Winkle, Gabriele Santin, and Bernard Haasdonk. Adaptive meshfree approximation for linear elliptic partial differential equations with PDE-greedy kernel methods. BIT Numerical Mathematics, 65 (1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01053-0.

Li:2025:BIM

[3724] Shishun Li, Jing-Yuan Wang, and Xiao-Chuan Cai. Block implicit methods with L-stability for parabolic problems. BIT Numerical Mathematics, 65 (1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01056-x.

Kang:2025:CEN

[3725] Hongchao Kang, Ao Liu, and Wentao Cai. Correction: Efficient numerical methods of integrals with products of two Bessel functions and their error analysis. BIT Numerical Mathematics, 65(1):??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01057-w.  See

[3713].

Chiwere:2025:BIF

[3726] Michael Chiwere and Grady B. Wright. Barycentric interpolation formulas for the sphere and the disk. BIT Numerical Mathematics, 65(1): ??, March 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01055-y.

Davis:2025:RMF

[3727] Owen Davis, Mohammad Motamed, and Raúl Tempone. Residual multifidelity neural network computing. BIT Numerical Mathematics, 65(2): ??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01058-9.

Ling:2025:EQC

[3728] Sitao Ling and Zhehan Hu. Efficient quaternion CUR decomposition based on discrete empirical interpolation method. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01060-1.

Akansha:2025:PBN

[3729] S. Akansha. Padé-based nonlinear approximation of bivariate non-smooth functions. BIT Numerical Mathematics, 65(2):??, June 2025. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.com/
article/10.1007/s10543-025-01059-
8.

Fermo:2025:GMS

[3730] Luisa Fermo, Domenico Mezzanotte, and Donatella Occorsio. A global method for solving second-kind Volterra–Fredholm integral equations. BIT Numerical Mathematics, 65(2): ??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01061-0.

Xu:2025:CAT

[3731] Xuefeng Xu. Convergence analysis of two-level methods with general coarse solvers. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01054-z.

Dunster:2025:NAC

[3732] T. M. Dunster, A. Gil, and J. Segura. A numerical algorithm for computing the zeros of parabolic cylinder functions in the complex plane. BIT Numerical Mathematics, 65(2): ??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01065-w.

Huhtanen:2025:GQE

[3733] Marko Huhtanen and Olavi Nevanlinna. Gradients of quotients and eigenvalue problems. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01064-x.

Butcher:2025:BSC

[3734] John C. Butcher, Taketomo Mitsui, and Shun Sato. On the Bseries composition theorem. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01063-y.

Dallerit:2025:SRE

[3735] Valentin Dallerit and Mayya Tokman. Stiffness resilient exponential integrators and φ-order conditions. BIT Numerical Mathematics, 65(2):

??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01062-z.

DellAccio:2025:TGH

[3736] Francesco Dell’Accio, Allal Guessab, and Federico Nudo. Truncated Gegenbauer–Hermite weighted approach for the enrichment of the Crouzeix–Raviart finite element. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01069-6.

Tominec:2025:WPS

[3737] Igor Tominec, Josefin Ahlkrona, and Malte Braack. Well-posedness of the Stokes problem under modified pressure Dirichlet boundary conditions. BIT Numerical Mathematics, 65(2): ??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01068-7.

Cheng:2025:SOE

[3738] Xinyu Cheng. Second order energy stable semi-implicit schemes for the 2D Allen–Cahn equation. BIT Numerical Mathematics, 65(2): ??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01072-x.

Kovacs:2025:NNT

[3739] Mihály Kovács and Mihály Vághy. Neumann–Neumann type domain decomposition of elliptic problems on metric graphs. BIT Numerical Mathematics, 65(2):??, June 2025. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10543-025-01067-

8. See corrigendum [3745].

Li:2025:TRA

[3740] Jiao fen Li, Jing Zhou, and Jia shuo Song. A trust-region approach for iteration solution of the direct fitting metric MDS. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01066-9.

Campagna:2025:ACP

[3741] Rosanna Campagna, Serena Crisci, and Marco Viola. An algorithm for a constrained P-spline. BIT Numerical Mathematics, 65(2):??, June 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01071-y.

Gao:2025:OEB

[3742] Weiguo Gao and Rentao Xu. An optimal error bound for shiftedCholeskyQR3 in oblique inner product. BIT Numerical Mathematics, 65(3):??, September 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01070-z.

Kusch:2025:SOR

[3743] Jonas Kusch. Second-order robust parallel integrators for dynamical low-rank approximation. BIT Numerical Mathematics, 65(3):??, September 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01073-w.

Li:2025:PAV

[3744] Wei Li and Pengzhan Huang. Partitioned algorithm with variable time step for fluid–fluid interaction. BIT Numerical Mathematics, 65(3):??, September 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01074-9.

Kovacs:2025:CNN

[3745] Mihály Kovács and Mihály Vághy.

Correction: Neumann–Neumann type domain decomposition of elliptic problems on metric graphs. BIT Numerical Mathematics, 65(3):??, September 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

//link.springer.com/article/10.
1007/s10543-025-01076-7.  See

[3739].

Kircheis:2025:SRR

[3746] Melanie Kircheis, Daniel Potts, and Manfred Tasche. Some remarks on regularized Shannon sampling formulas. BIT Numerical Mathematics, 65(4):??, December 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01081-w.

Pandur:2025:DLD

[3747] Marija Miloloza Pandur and Ivana Kuzmanović Ivicić. Detecting large definite Hermitian matrix pairs by θ-subspace algorithms. BIT Numerical Mathematics, 65(4):??, December 2025. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10543-025-01082-9.

Gill:1951:PSS

[3748] S. Gill. A process for the step-by-step integration of differential equations in an automatic digital computing machine. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 47:96– 108, 1951. CODEN PCPSA4. ISSN 0008-1981.

Kahan:1965:PFR

[3749] W. Kahan. Pracniques: Further remarks on reducing truncation errors. Communications of the Association for Computing Machinery, 8(1):40, January 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:1983:VSP

[3750] D. T. Lee. Visibility of a simple polygon. Computer Vision, Graphics, and Image Processing, 22(2):207–221, May 1983. CODEN CVGPDB. ISSN 0734-189X. See corrections [1618].

Dobosiewicz:1978:SDP

[3751] Włodzimierz Dobosiewicz. Sorting by distributive partitioning. Information Processing Letters, 7(1): 1–6, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https:

//www.sciencedirect.com/science/
article/abs/pii/0020019078900285.

See debate [?, ?, ?, ?, 1607].

Jonassen:1978:TAW

[3752] Arne T. Jonassen and Donald E.

Knuth. A trivial algorithm whose analysis isn’t. Journal of Computer and System Sciences, 16(3):301–322, 1978. CODEN JCSSBM. ISSN 0022-0000. See continuation [1764].

Fagin:1979:EHF

[3753] Ronald Fagin, Jürg Nievergelt, Nicholas Pippenger, and H. Raymond Strong. Extendible hashing — a fast access method for dynamic files. ACM Transactions on Database Systems, 4(3): 315–344, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1979-4-3/p315-fagin/p315-fagin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-

3/p315-fagin/. Also published in/as: IBM Research Report RJ2305, Jul. 1978. See [1456].

Cox:1997:ASN

[3754] Anthony J. Cox and Nicholas J.

Higham. Accuracy and stability of the null space method for solving the equality constrained least squares problem. Numerical Analysis Report 306, Manchester Centre for Computational Mathematics, Manchester, England, August 1997. 20 pp. Published in [2321].

Cox:1998:BEB

[3755] Anthony J. Cox and Nicholas J. Higham. Backward error bounds for constrained least squares problems. Numerical Analysis Report 321, Manchester Centre for Computational Mathematics, Manchester, England, April 1998. 20 pp. Published in [2331].